A Complete Bibliography of Publications in Fisheries Research (2000–2009)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

30 August 2024 Version 1.09

Title word cross-reference

$105 [Ste05]. $130 [McI03e]. $140 [McI03b]. $144.99 [How03b]. $25 [Gra02a]. $25.00 [McI06b]. $27.50 [McI05c]. $29.50 [Ste03a]. $325 [McI05a]. $34.95 [Mil02a]. $395 [How06]. $40 [Coo00a, Coo01, Gra02a, McI01, Ore03]. $40.00 [McI02c]. $42.95 [McI02e]. $45.00 [McI06a]. $47 [Lon02a]. $59.50 [Ste03a]. $70 [McI09d]. $84.95 [How03a]. $90 [McI09b]. $99 [Ano01c]. ^{210} [KKJ00]. ^{222} [KF02]. ^{226} [KKJ00]. 86 [MK09]. ^{87} [MK09]. _{2} [CS07b]. $A [Kes00]. a [GMA02]. F_{0.1} [CXCD07, GC04]. F_{max} [CXCD07, GC04]. L_{opt} [Sve15]. Z [DFD02].

İskenderun [Ism03, IYI07].

//www.blackwellpublishing.com [McI07b].

0 [Ano00b, Ano00c, Coo00b, Cur00, Gue05, How06, Kes00, McI02f, McI07b, McI08, McI09c, Mil06, Sol05, Ste05, Sym08c]. 0-1-84103-004-X [Bar00].

0-12-088414-3 [Mil06]. 0-12-350438-4 [Ano00b]. 0-226-06618-5 [McI03e]. 0-231-12488-0 [Ste03a]. 0-231-12489-9 [Ste03a]. 0-295-97840-6 [Mil02a]. 0-309-05526-1 [McI02f]. 0-309-06030-3 [Coo00b]. 0-309-07286-7 [McI02e]. 0-412-03541-3 [Ano00c]. 0-444-51027-3 [Ste05]. 0-632-05355-0 [McI03a]. 0-632-05431-X [Sut02]. 0-632-05733-5 [Loc02]. 0-632-05754-8 [Mil02b]. 0-632-06391-2 [Mil03b]. 0-632-06457-9 [Sol05]. 0-8020-7959-8 [Ano01d]. 0-8493-1916-1 [How06]. 0-8493-9061-3 [How03a]. 0-85238-241-3 [Cur00]. 0-85238-246-4 [She02]. 0-85238-250-2 [Win02]. 0-85238-273-1 [McI02d]. 0-85238-280-4 [McI07b]. 0-85238-286-3 [How03b]. 0-85976-410-9 [Ano01f]. 0-958703-5-0 [Kes00]. 0-group [Din05, KA05].

1 [How06, McI02a, McI02f, Sym08a]. 1-56612 [McI06b]. 1-56612-038-1 [Gra02a]. 1-56612-049-7 [McI01]. 1-56612-057-8 [Coo00a]. 1-56612-061-6 [Coo01]. 1-56612-062-4 [McI02c]. 1-56612-077-2 [Ore03]. 1-57808-013-4 [Ano01e]. 1-85233-117-8 [Ell00]. 1-85233-118-6 [Van01]. 1-886106-31-2 [Ano01c]. 1-888569-15-8 [Lon02a]. 1-888569-18-2 [McI02a]. 10-year [SMD04]. 140511194 [McI05b]. 15 [McI03e]. 17th [GKUO07]. 1835 [Vil01]. 1880s [BSM07, EMKG07]. 18th [LDK^{+}07, LAL07]. 1990s [CPW^{+}06, HP09a]. 19th [LOT07a, Ore03]. 1st [Ano03k, Ano07a].

2 [Ano00d, BFB^{+}03, Gue05, McI02a, McI05a, McI06a]. 2-7099-1354-2 [Ano00d]. 2-7099-1517-0 [Gue05]. 2001 [Bla06]. 2004 [VL08]. 20th [LDK^{+}07, LAL07]. 26^{o} [EBV^{+}03]. 2H [LK00]. 2J [LK00]. 2J-3K [LK00]. 2Teleostei [PJR^{+}05].

3 [Ano00c, Ano00e, Coo00b, Cur00, DDP^{+}06c, McI05a, Mil06, Ste05, McI09d]. 3-540-42854-2 [McI05a]. 3-540-61365-X [Ano00e]. 34^{o} [dSH01]. 3K [LK00].

4 [Ano00b, McI07b, McI07a]. 40^{o} [CBC02]. 42 [Bas00]. 45 [CAJ00a]. 461pp [McI09c]. 48 [GMV01].

5 [Sym08b]. 54 [GMR02, NNA02, vSM02c]. 54-58 [Ano02b, Ano02-28].

6 [McI09a]. 60 [MG03a]. 61^{o} [HG01]. 62 [GBC^{+}03b]. 68 [KI05]. 69.50 [McI03a].

7 [McI09b]. 71 [FCL^{+}05a]. 75 [BMB06, McI02d]. 76 [SBD^{+}06]. 79.50 [Mil02b, Win02].

8 [McI05d, Sym07]. 81 [McI06a]. 81-7075-079-2 [McI06a]. 84 [KPL^{+}09]. 89.50 [Loc02].

9 [Pie09, Sol05, Win00]. 90 [Win00]. 90-5410-721-9 [Win00].

90-5972-061-X [McI05c]. 90^{o} [HPK07]. 92 [McI05d, McI07a]. 92-3-103995-4 [McI07a]. 92-5-104825-8 [McI05d]. 95/094 [AKA^{+}01]. 96-03 [Gra02a]. 978 [McI08, McI09c, Pie09, Sym07, Sym08a, Sym08c, Sym08b, McI09a, McI09b, McI09d]. 978-0-8138-2546-5 [Sym08b]. 978-1-4051-3932-8 [Sym07]. 978-1-4051-4161-1 [Sym08a]. 978-1-4292-0446-0 [McI09c]. 978-1-57808-375-6 [McI09a]. 978-1-57808-435-7 [McI09b]. 978-1-57808-529-3 [McI09d]. 978-1-84407-527-0 [McI08]. 978-84-9745-235-9 [Pie09]. 978-92-3-104029-0 [Sym08c].

Âncora [PF00].

abalone [DDHS06, GMP05, HFBM08, LRBH07, MWJ04, PPGD08, RB09, ZLC09]. abbreviata [MC04]. abdominalis [WMS04]. abiotic [RPC^{+}06]. absence [PEAAdLJB07]. Absorption [Dia00]. Abundance [APS07, Kat05, ASR05, dSHG08, AM02, BTC08a, BK09, BEA^{+}08, BPW01, BM09, Cam04, CS04, CPW^{+}06, CdML^{+}07, CF07, Cor01, CBC02, DOQ08, Din05, DKKS07, EF01, FAL^{+}01, GBG^{+}03, HK06a, HJ06, HMH^{+}04, HD00, IHF08, KLU01, KM08b, KLJ06, KHMS09, KCK03, LHMO09, LD01, MvdK06, MS03, MR05, MCB^{+}07, MC06, NGF06, NLJS01, OQE03, PR03, PB03, PCHM07, PCS05a, RPFR06, RVS02, SHKB09, SC07, SHDO04, SCM09, SWC^{+}05, TAS03, Tia09, TSOG01, WN03, YCB^{+}01, YPDS05, YLYL09, ZPR01]. abundances [CC01a, HdS07]. abundant [LRC^{+}08]. abuse [LM09]. Academic [Ano00b, Bar00, McI05c, Mil06]. Academy [Coo00b, McI02e, McI02f]. acanthias [DS07, MF07]. Acanthopagrus [GMY^{+}09, GAFA04b, GCJ07, GCMJ08]. access [LFA08, NGF06, ZR08]. according [MBG^{+}07, VH06]. Accounting [QPR08, SW09, DDP^{+}06a]. accumulation [OKPR04, BM03c]. Accuracy [FUM02, GDFH07, HFS02, HCS^{+}03, MM06a, PWHM06, RHF^{+}08, WIKM05]. ACFM [CL04]. Achievements [GML06]. acid [CCA^{+}03]. acidification [HK02]. Acknowledgement [Ano04x, Ano01s, Ano01t, Ano02z, Ano03-31, Ano05-29, Ano06-31, Ano09-32]. Acknowledgements [Ano01a, RYG01]. Acoustic [BDB^{+}03, DJG^{+}06, KTC00, KLJ06, LRG^{+}00, PR03, SHDO04, TS06, BMW07, CMPM01, DBB00, DK05, FK04, GDDJ09, GMR01, GMR02, HD00, JL00, KS02, KHMS09, MHS^{+}05, MSSJ05, RVS02, TKFJ09, UHN07, vSM02c, vSM02a, vSM02b, vSSK07, vSS09]. acoustical [AD03]. acoustics [KCRT03, MTRM05, MTGH00, TK00]. across [ABF08, APP^{+}07, CSM^{+}09, GWB07, LLV^{+}08, LMR^{+}07, SCM^{+}08, TOBSJH06]. Act [DJGC09]. action [KAO04]. activity [CHQ05, SOH06]. acusticus [TLMN00]. AdaBoost [KMELC05]. adaptive [KP07]. added [LBKG08]. additive [BPW01, HMH^{+}04, KMELC05, WK01, WKM02, WIKM05]. address [KM08a]. adjacent [AAHE04, BPH^{+}06, Mar06a, Mey07, MM06b, MMLR08, PH09, PBM06]. adjoining [CHN08]. Adjusting [ALN03, FKW06]. Adriatic [CUT06, CAM^{+}08, ŠV04, CSDP02, DKGC00, FSB^{+}02, JPVKS^{+}01, LDGA09, MFAM05, MAG^{+}09, TGS^{+}00]. adult [BNSA05, BNS07, BR05, BFM08, BC04, Fra07, GFM02, HMRG08, PAP04, UHN07, UL01, WCM04, WGGN06]. Advances [Ano01e, BDIP08, Sym08a]. Advection [PNJS06, PTPS09]. advice [MMZ09]. Advisory [Ano02h, Ano02j, Ano02k, Ano03k, Ano03i, Ano03j, Ano03f, Ano03g, Ano03h, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano04a, Ano04c, Ano04e, Ano04f, Ano04j, Ano04k, Ano04l, Ano04m, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06r, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano08f, Ano08g, Ano08h, Ano08i, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano03m, Ano04b]. Aegean [KVS04, SE02, KTL06, MTU^{+}04, SME02, TOÖD06]. aeglefinus [AFVSR^{+}09, GM07, Hal02, HMH^{+}04, LR08, ÖW02]. aequidens [HGSP01]. Aequipecten [NS04]. Aerial [CARA^{+}07]. affect [Sal02, SW01]. affected [CLGR^{+}04, KJEF03]. affecting [Daw08, DDP^{+}06c, GWB07, JI07, LAL07, Lan04, SSB03, TPC05]. affects [SC08, SOH06]. affinis [ABFV02, PGMM01]. affinity [ASD09]. Africa [CWJS09, CMGB01, CARA^{+}07, GMS06, HGSP01, NBA^{+}09, RMBG02, RB09, SLA00, YFL08, ZCdG06, ZCTA^{+}08, tHDC06, Bar00]. African [HBBG06, BW08, BBJG04, BDBD09, BGB06, Eze02, GB06, HGF06, RB00, WLS07, dGvZM^{+}06]. after [BLH04, EH04b, FFM08, LBKG08, Mar06a, Mar06b, MTU^{+}04, NGM^{+}09, PBDP00, ROS04, SLB04, SB08, iTMI06, VL08]. against [BA06]. Age [ACPP06, AHAAD01, AHABAA^{+}02, ATB^{+}02, AdBP02, ASAGR06, ARVC01, BLV06, BG04, Bur02, Caz00, CS08c, CMGB01, CMC09, DSPLR06, DKGC00, GBC^{+}03a, GBC^{+}03b, Gon06, GAF06, GCC^{+}08, Hoo06, Hor02, HLH08, IMSS06, Ism03, IYI07, JLC04, JCLL08, LLM05, LS00, LSH04, LDN04, LLJC09, LE01, MRCH04, MQVSN04, MMB07b, MOOH00, MK06, MOGG06, MNM00, MMN03, MC04, NT05, NCW00b, NCW00a, OMF^{+}03, PLDS03, PH09, PH04b, RMBG02, RCGMQV05, RLG^{+}01, RRJ09, STCH09, SGS09, TT01, YST^{+}02, AFSF05, AFG^{+}05, ACG^{+}01, BMC07, Ber07, BF00b, CAJ00a, CAJ00b, CP00, CAHAF07, CCB02, CHSA02, DBK07, Dou04, FL05, Fab06, FDG^{+}07, FUM02, FBJM00, FFL09b, FBS02, GMY^{+}09, GMP05, HH06, JLHO01, JCS01, Jon00, JMGC06, KR00, KK09, KLU01, LK00, LB08, LCFP^{+}08, LMR^{+}05]. age [LMR^{+}07, MPM^{+}06, MM06a, MCAO^{+}05, MN01, MB08, MG03a, MG03b, New02, OSG08, PL02, PMG^{+}06, PC03, PGK03, PCAS04, PKK^{+}06, RMRB06, RYSN08, Smi03, SRTP06, SDS09, SKKP05, SG01b, SWC^{+}05, Sym06, TS09, TEYA02, WSPY05, WL06b, WSPY07, WY06, WKCC06, WB08, XM00, YLYL09]. age- [XM00]. age-0 [KLU01, SWC^{+}05]. age-based [HH06, LMR^{+}05, Smi03]. age-readers [PKK^{+}06]. Age-reading [ATB^{+}02]. Age-specific [MMB07b]. age-structure [CAJ00a, CAJ00b]. age-structured [CAHAF07, CHSA02, DBK07, FFL09b, JMGC06, MG03a, MG03b, WSPY05, WSPY07]. Age-validation [RRJ09]. Ageing [ABC^{+}08, FAN03, MF00, PWHM06, SSS^{+}03]. ages [EPA00, KKJ00, Ker00]. aggregate [KM08b]. aggregating [DJG^{+}06]. aggregation [Dem04, Ste03b]. aggregations [BR01, DJG^{+}06, RB00]. AIC [WL06a]. air [GCH^{+}07, HBBG06, WSC08]. airborne [San00]. AK [Coo00a, Coo01, McI01, McI02c, Ore03, Ore03]. AK-SG-02-01 [Ore03]. AK-SG-97-01 [McI01]. AK-SG-98-01 [Coo00a]. AK-SG-99-01 [Coo01]. AK-SG-99-03 [McI02c]. Akaike [WL06a]. al [BNS07, Bet04, Sve14, Sve15, VL08]. al. [Fra07, Lon03a, MC00b]. Alabama [MM04]. Alacranes [BRL^{+}05, RLSJIA07]. alalunga [RM07, SEPY02, WCC^{+}09]. Alan [Lai00]. Alaska [Coo00a, Coo01, Gra02a, McI01, McI02c, McI06b, Ore03, RB04, APC07, BK09, Bro05b, CMOM00, DA07, ESC08, JDAM08, Loh08, RGSB08, SW09]. Alaskan [ZK03]. albacares [DHM08, LDN04, NC04]. albacore [CHL01, RM07, SEPY02, WCC^{+}09]. Albatross [TVB^{+}09]. albidus [AB09b]. Alboran [RCGC04, VYMGM^{+}09]. Albula [DACS08, KG04, Meh06, SG06, Sch09]. Alexander [Kes00]. alfonsino [RLG^{+}01]. Algarve [CF04, EGB^{+}01, SB01a, SGVM02, SSGM03a, SSGM03b]. Alicante [GRFE06]. alien [RGGB07]. alizarin [Dou08, LZZN09]. alleviate [NGM^{+}09]. allocation [OGSG07, PBMB03, TB04, VD04a]. Allometric [MDGR05, TY02]. Allometry [AC02]. allow [PHGT06]. Allozyme [SMRH04]. along [FB06, GTB04, HG01, KWK^{+}09, KIT^{+}04, LCFP^{+}08, LMRQNM^{+}05, MHS^{+}05, MTC06, MACS08, PTR^{+}05, RRJ08, RRJ09, RCG02, TKFJ09, UL01, VSK05, dSH01]. Alopias [LCNJ06]. alpinus [EF01, HMB^{+}00, WFJB09]. Alta [TNFF03]. alterations [Dah00]. Altering [Sym05a]. Alternative [OA04, Cad03, FFM01, GLH07, HPD08, HGSP01, IWBJ08, MvdK06, SBD^{+}05, SBD^{+}06, VH06, Zho07]. America [SCSC07]. American [Lon02a, McI02a, McI05d, BHH08, CGT08, MBK02, Row02, SCW04, SPS^{+}07, VL09, WT00]. americanus [BHH08, He03, IWW03, Laz08, PH04b, ST03]. Ames [How03b]. Ammodytes [BHJ02, SBD^{+}05, SBD^{+}06]. Among [HWB06, CG04, CMAS08, Dic04, IJD04]. Amsterdam [Mil06]. anadromous [BTC08a]. analis [Bur02]. analyse [Bas99, Bas00]. Analyses [SEPY02, BHH08, CL09, CJ00, KA05, PRdPG08, SM04b, TLH07, WK01]. Analysing [BVW04]. Analysis [BS04, HMH^{+}04, MDS03, Mes03, PVP04, RT07, SSB03, WIKM05, AAFLP^{+}07, APP^{+}07, Bal09, BBRS06, BPRM00, BRL^{+}05, BGH05, BP01b, BBK08, CHL01, CY06, CG04, CdMLGC09, CHSA02, DGP02, FCC^{+}09, GFSM07, GRFE06, GKV06, GOAARJRP09, HPM04, HS03, IS07, JC04, JHSG04, Jun08, KW03a, Kir01, KPL^{+}07, KPL^{+}09, LB08, LCNJ06, MGTB08, MBT08, MLH^{+}05, MT08, MMS09, MK01, MOK^{+}05, MOGG06, MPA08, NJL^{+}00, OF03, OGPC09, Ove02, PPL04, Pál03, Pál05, PH04a, Pou01, PLSI09, PM09, RSP^{+}00, RLSJIA07, SCUADJ07, SDNL09, SÁR^{+}05, STK^{+}05, Sho08, SDH00, SMSZ08, SBF^{+}08, STH03, TS06, TSS^{+}04, TC09, TPC05, TLD06, TSTK08, UMK01, Vas03, WG03, WYR06, WLY06, WL06b, WBMM05, WB08, WH08, XM00, XW07, YEA09]. analytic [CHSA02]. analyze [Tri05, YAHAB01]. Analyzing [LK02, Xia04a, Xia04b]. anchoita [CC03b, PCM07]. anchor [BW08, McP02]. Anchorage [Ore03]. anchored [TF09]. anchoveta [Ara09]. Anchovy [CC03b, ACE^{+}08, AM00, BGP^{+}04, BPRM00, CC01c, CBC02, CRC^{+}07, DUR05, FZHL09, FAW04, GESY^{+}07, JS02, JHK08, Jun08, Kno07, KM08c, LDGA09, PCM07, TOK^{+}05, WLY06]. and/or [Cor01]. angel [BAH09]. Angeles [CTEARD07, NE07, RDEACT07]. angelfish [SDS09]. angle [Gri06]. angled [BR05, GCH^{+}07]. angler [CGLZ04, JDS00, RKEJR06]. anglerfish [CCM^{+}09, DALP01, GRPLE05, LD08, LPDA01, LQD^{+}08, WWG03]. anglers [AM03, ABF08, DACS08]. angles [TMK^{+}06]. Angling [MDJR05, Aló09, AKC^{+}09, GCJ07, MM05, PIdlHGV05, PWB05, RCA08, THN^{+}04, TNL07]. Anguilla [GT04, PRM04]. Animal [McI08, BBRS06, RMA^{+}05]. Animals [Lon02a]. Anisakidae [MFC^{+}08]. Anisakis [MFC^{+}08]. Annecy [GJ09]. ANNEX [Ano01b]. Announcement [Ano02a, Ano05a, Ano05b, Ano06a]. Annual [Ano09a, BPW01, CG04, CHQ05, KA05, KNSSKF09, KKR^{+}05, WWG03, tHDC06, vZRMvD02]. annularis [FSB^{+}02]. annuli [BM00, BBK08, MMP^{+}02]. anomalies [PSS^{+}07]. Anoplopoma [BM00, CS07b, GJSW04]. Anostomidae [AAB05]. Antalya [BK02]. Antarctic [BFB^{+}03, HD00, Hor02]. Antarctica [Hor02]. antennatus [CA03, DCM^{+}05, DMC^{+}09]. Antilles [DJG^{+}06, Gob00]. Antimora [FB06, Mag01]. anything [JIGI06]. Aphanopus [PGS^{+}08]. Apollonio [Ste03a]. apparatus [LBMCE00]. apparent [CPW^{+}06]. appendage [BGB06]. Application [AFVSR^{+}09, Bas99, Bas00, CA03, DPD^{+}03, How03a, STK^{+}05, Sho08, TLD06, TBRL00, WSPY07, WMdSB09, WH08, YPD^{+}04, YPM^{+}06, ZC02, BEBA03, Bet04, BR01, Bro05b, FHI07, GG06, HMM00a, HH06, IHF08, JC04, JCW05, KM08a, KM02, KS02, Kop06, LK02, Mai09, MB08, NC04, RYSN08, TdlMH^{+}03, UVKS05, VVH04, VMSR08, WKM02, WHWM07, ZG08, vHMAO^{+}06]. Applications [Tal09, VD04b]. applied [ED07, Fab06, GKV06, GYMdL^{+}06, Pra02, RSMZ06]. Applying [SKMT02, Tri05, GGVLSR03]. Appraisal [MAG00, CAM^{+}08, NGF07, PP01]. approach [AMC^{+}08a, ASM^{+}09, AIC08, BEBA03, Bet04, But03, Cad03, Dia00, FMGG08, FFM08, FFL09b, GRPLE05, Gav09, GOAARJRP09, GESY^{+}07, GEYPC^{+}09, KMF08, KIZ09, MVAMB07, PBVG07, SW09, SMD03, SM04a, SMS^{+}06, SE06b, Tri06, TSJ04, YJV^{+}03, ZP05, ZKG^{+}09]. Approaches [Ano00b, Coo01, AMC^{+}08b, BVW04, BQ06, Cad02, CGEHC07, ERG07, FFL09a, MP04, TPC05, TMP^{+}06, vMBGF05]. approaching [RO00, Sch09]. appropriate [DDP^{+}06b]. April [Ano00r, Ano01j, Ano02n, Ano03t, Ano04w, Ano05q, Ano07s, Ano08r, Ano09v]. Aptenodytes [PPH01]. Apulian [GTB04]. Aquaculture [Van01, GFMO06, McI03d, Lai00]. aquarium [BSWL06b]. Aquatic [Mil03b, BBRS06, McI02a]. Arabian [GAFA04b, GAFA04a, GAFA05b, GAFA05a, GAF06, Hoo06, JS02, vHMAO^{+}06]. Arauco [LC06]. archipelago [DSPLR06, LD06, RMMP06, SÁR^{+}05, GCS^{+}06, HSSR06, TOBSJH06]. archipelagos [RLG^{+}01]. archival [Loh08, NHT07]. archives [LOT07a]. Arctic [KI05, DSO^{+}05, EF01, GM00, HMB^{+}00, KI04, SGL08, SBF^{+}08, WFJB09]. Arctica [TJ05]. Arcto [HBM^{+}00]. Arcto-Norwegian [HBM^{+}00]. arcuatus [FW06, SDS09]. Area [Güç08, Mey07, ABF08, BWMA02, BMF^{+}10, BG04, BPH^{+}06, CCCC04, CBC02, GESY^{+}07, HHC06, JHGS04, KM02, LHMO09, LZC^{+}08, MWJ04, MRCH04, Moo03, NGF06, NV05, PBW^{+}09, PA09, PLSI09, SLV^{+}08, SSH^{+}06, SHDO04, vS04]. areas [AA03, BPH^{+}06, CXQ^{+}09, CMAS08, CPR07, DCM^{+}05, GCS^{+}06, GBFG06, JCJ^{+}09, KVS04, LDA01, LD01, MH08, McI02e, MMLR08, PH09, PBLP08, PBM06, PKP04, WM04]. argentata [YTK06]. Argentina [AAB05, Gon06, MC04, Nar06, NGF07, Caz00, GNC07, MAL02, MM06b, NS04, TCPC06, VLD^{+}00]. Argentine [CS08c, Gon06, MPE04, MPM^{+}06, PMM05, PCM07, PGF07, VMQ08, VE03]. Argentinian [MHS^{+}05]. argentinus [AM02, BQN05, Lap02, MHS^{+}05, MCP01, WGR08]. Argopecten [OS05]. argus [BRL^{+}05, Ehr08, GYMdL^{+}06, PVMdL05, RLSJIA07, WVW^{+}08]. Argyrops [GAFA04b]. Argyrosomus [Kir01, SGS09]. Argyrozona [BG04, BG05]. Arhynchobatidae [CGP05]. Ariake [NT05, YTK06]. Aristaeomorpha [PK03a, RBD02]. Aristeus [CA03, DCM^{+}05, DMC^{+}09]. arm [YSS^{+}09]. Arnason [Sym08a]. array [DSO^{+}05]. Arrhenius [MC00b]. Arrignon [Win00]. Arripidae [CWL00]. Arripis [ABE^{+}04, CWL00]. art [MN01]. artedi [YSS^{+}09]. artificial [DVAL^{+}09, JLC^{+}03, MLJT09, MSVF05, RSMZ06, How03a]. artificially [YMT^{+}06]. Artisanal [BSMFH07, NGF07, ASIW05, BTC09, BSMF^{+}09, BH01b, CP06, CCCC04, Daw08, GNC07, Güç08, Gus02, Kro04, LHMO09, Mos00, MUO02, MGNB05, OGPC09, PSPCVCD09, SPS^{+}07, SGS02, TMN07, ULDG02, VD06, VMPP07, WC07]. Arve [Cur00]. ascending [TNL07]. Asia [Bar00, GSA^{+}06, SSG06]. Asian [Abl06, Bla06, OLN^{+}02, SST^{+}06]. asinina [MWJ04]. Askale [TEYA02]. aspect [BMW07]. Aspects [CPT06, CE06, Jak01, PGMM01]. Assemblage [SET^{+}09, CFH03, MSB07]. assemblages [ESB09, GSA^{+}06, GFM02, GRC^{+}09, GBFG06, KVS04, KM08b, MCL04, MC05, MBG^{+}07, SAG05]. assess [KA06]. assessed [YSS^{+}09]. Assessing [CH03, Dou04, DMBF08, ESB09, FFL09a, FFL09b, GR07, HE08a, Hor05, KCH05, KJ09, KF04, Mat07, PSPCVCD09, PWP02, RSMZ06, SCP01, SLJ07, SE06b, CdML^{+}07, EF01, JCW05, TSJ04, CCD01]. Assessment [ASIW05, Ano06-30, Bla06, CF07, FML^{+}07, LDA01, MH02, MM04, McP02, MB04, RB00, SBFR00, WFJB09, ZG08, ASM^{+}09, AB09a, APS07, ABD^{+}09, AIC08, AHS^{+}06, BC02, BRL^{+}05, BACP09, BG01, BQ06, Cad02, CRD06, CMGB01, CL09, CCT^{+}08, Coo00a, Cor02, DBK07, ET00, ED07, FZHL09, FRGF07, FW06, FBS02, GG06, GB06, Gob00, GNC07, GAOMC06, GAFA04b, GAFA04a, GAFA05b, GAFA05a, Gro00, GBG^{+}03, HC07, HGSP01, IMOO06, JR07, JCW05, KCH05, Kro04, Lab05, LMASMASZ02, MKGH05, MVAMB07, MT08, Mau03, ML04a, MSSJ05, ML04b, Mes03, MDJR05, MBCMNMHH01, MPG^{+}08, MGJB09, NJE00, PR03, PAB05, PAP06, PW05, PPdA05, PPGD08, PVMdL05, Pun03, RFWH09, SCSC07, SDH00, SA03, SJ05, TCS05, TF09, TdlMH^{+}03, WSPY05, WSPY07, WMdS09, YLD^{+}05, ZDB^{+}08, ZKG^{+}09, dSBAC06, Cur00, McI02a]. assessments [FAN03, PSTM06, Coo00b]. assignment [GM06]. assisted [Fab06]. associated [BSAM09, BAH09, Dem04, HWW^{+}09, LG02, OSC^{+}05, RBH^{+}08, SW00, WRK06a, WRK06b]. association [Ell00, PSS^{+}07]. associations [BQN05, BLR^{+}08, BMB05, BMB06, UVKS05, VBSK00]. assumptions [DMBF08, FHI07, WMdS09]. Asturias [FRGF07]. at-sea [CXCD07]. Atherina [LS00, MMN03, MCS05, TGS^{+}00]. Atherinidae [LS00, MMN03, MCS05]. Atlantic [ABK03, Ano00e, CE07, How06, JIAD03, KN01, McI05d, SDNL09, Sol02, SBD^{+}06, AMC^{+}08b, ADP^{+}08, ATM08, dSHG08, All01, AFVSR^{+}09, AC02, AS04, ASD09, AKK^{+}03, AB09b, BAH09, BBM^{+}06, BLR^{+}08, BC04, BM08, BDBD09, BM03b, BM03c, BSG02, BBK08, BRRL09, CMJ05, CMC^{+}06, CS04, CC03b, CPSBRD^{+}02, CPW^{+}06, CMAS08, CMC09, CK02, DSO^{+}05, DGP02, DR01, DTSR07, DALP01, EMN07, EBV^{+}03, FR03, FH03a, FB06, FR08, GCS^{+}06, GR02, HLL^{+}08, HMO01, Hal02, HP02, HP09a, HG01, HHTE05, HE08a, He05, HP04, HMRG08, IJD04, JLHO01, JSB04, Jok02, JIDM02, JIAD03, JJKN^{+}03, JJJ03, KNJJS06, KNSSKF09, KC06, KFO^{+}04, LDK^{+}07, LPDA01, LPP02, LQD^{+}08, LJP05, LRG^{+}00, LK00, LDN04, Løk03, LLV^{+}08, MKGH05, MM07, MCM^{+}08, MNML00, MGR^{+}02, MFC^{+}08, MR05]. Atlantic [MAC^{+}06, MÁ08, MD02, MSA00, MAL^{+}01, MMN03, MCS05, NW03, Nas03, NP08, ØA08, OGSG07, Ove02, PLDS03, PGS^{+}08, PA02, PH04b, PIdlHGV05, PCB01, PH01, PCMC08, Pra02, QP00, RL03, RM07, RLGLO09, SF05, SSIE06, SCM^{+}08, SHBC02, SKKP05, SMSZ08, SEPY02, SBD^{+}05, SSC^{+}03, Tal09, TNFF03, TNL07, TP00, TOBSJH06, UL01, VP05, VAF04, WIS03, WGR08, WPS^{+}07, WM00, WF07, WD02, WR05, WNRd00, How06]. atlanticus [CAFT00, Cla01, KM02, SR06, SRH^{+}02]. Atol [GCS^{+}06]. Atractoscion [APS07, HGSP01]. atricauda [MSA00]. attempt [But08, HS00, KPBD08]. attendance [APM^{+}09]. attraction [AA03]. attractors [HHTE05]. attrition [GHM04]. audax [ACMMCEGM02, OGPDOM08]. August [Ano00l, Ano02r, Ano03y, Ano04p, Ano05v, Ano07q, Ano08s, Ano09-29, Ano09y]. aurata [FGJD08, HAO00]. auratus [HMM00b, HMM00a, MC06, ODdGV02, Ste08, SJ05, SOKS08, WMTJ07]. auriga [PMG^{+}06]. aurita [Gue05, ZCTA^{+}08]. auritus [BC02]. aurorubens [All07]. austral [GOAARJRP09, PPH01]. australasicus [Tze04]. Australia [CCD01, Fri07, Kes00, ABD^{+}09, AHS^{+}06, AWY02, ABE^{+}04, Bal09, BMBM05, BW02, BGW06, BHM^{+}06, BKD^{+}02, BPW^{+}03, BPH^{+}06, CS09, CCMS^{+}03, CTC^{+}06, Dem04, DPD^{+}03, DDP^{+}06a, DDP^{+}06b, DDP^{+}06c, EL09, GSS09, GKH^{+}01, GKH03, GWB07, GFvdV06, GCMJ08, HFBM08, HBW00, HW00, HC09, LPHH09, MLH^{+}05, Mat07, MSTC04, MCB^{+}07, NCW00a, New02, NS07, OCT^{+}03, OL07, OSG08, PG06, PBW^{+}09, PWP02, PSTM06, PHGT06, SBFR00, SGS09, SB05, SDH00, SST^{+}08, SLJ07, Ste03c, SH09, SJ05, SOKS08, Sva05, SRS08, Tay07, TGH^{+}08, TÁTWD09, TLH07, VD04a, WMTJ07, WSS02, WEK^{+}09, WDC02, WMB08, Wes02, XM00, XK04, Xia04a, YLD^{+}05, YPDS05, vHAN^{+}09]. Australian [MG03a, ABE^{+}04, BMK^{+}04, CWL00, DSK^{+}08, GBM09, GK03, JS05, LWM07, MG03b, OLN^{+}02, RM05, SMM07, SMJS01, WR02, ZG08]. australiensis [RW03]. australis [ASD09, ASD09, BBL07, SLJ07]. Author [Ano00a, Ano02b, Ano04g]. Authors [Ano02i, Ano03l, Ano04h, Ano05k, Ano06o]. autocorrelation [NC04]. autodiametric [AFVSR^{+}09]. Automated [FL05, WSS06, MLJT09]. autonomous [BFB^{+}03, RDM^{+}09]. Autosub [BFB^{+}03]. Autosub-2 [BFB^{+}03]. autumn [IMOO06, YM00]. auxiliary [BZRO05]. availability [NHT07]. available [DV00, McI07b]. Aveiro [CHQ05]. avgotaracho} [KKR^{+}05]. Avoidance [RO00, BFB^{+}03, DK05, JHGS04, SG06, VOH02]. Azores [ACPP06, ATM08, FMB^{+}05, MSA00, MAL^{+}01, PGMM01, RMMP06].

B [McI02a, McI05a, McI07a, McI09a, McI09b, Ore03, Sym08c]. B.V [Ste05]. BACI [FFM08, PBW^{+}09]. back [Hor00, MSB^{+}08, PZH02]. back-calculated [Hor00]. backcalculation [DSPLR06]. backscatter [vSS09]. Bacoma [OGK^{+}08]. Bad [HK02]. Baer [LOT07a]. bag [TS06, Wil09]. bag-net [TS06]. Baglinière [Ell00]. bags [KOO04]. Bahamas [ED07]. Bahia [CTEARD07, LMASMASZ02, RDEACT07, ARCLBPdP07, Caz00, NE07]. Bain [McI02a]. Baird [Kes00, Sym08c]. bairdi [Ste03b]. Bait [SB03, WM07, AAP^{+}09, BH01b, FRF08, GCJ07, LSW06, PCMC08, SOH06, WM09, WBHG01, YKO09]. bait-dependent [SOH06]. baited [MAG^{+}09, RSM05, YCB^{+}01]. Baiting [KJEF03, EJA00]. baits [AAT08]. Baja [ARCLBPdP07, NE07, RCGMQV05, Vel03]. bajad [GAFA04a]. balance [WM09]. Balancing [MLP^{+}08]. Balearic [CAM^{+}03a, CLGR^{+}04, MMN00a]. Balkema [Win00]. ballan [FMB^{+}05]. Baltic [BSM07, CAJ00a, Eer04, GKUO07, HK02, KPL^{+}09, LOT07a, SLV^{+}08, VKD03, BVH^{+}08, BLKR09, CAJ00b, EMKG07, FGOF08, GWR07, HH01, HWM02, Hor05, JJKN^{+}03, KNJJS06, KNSSKF09, KSS05, KK04, KPL^{+}07, MBO^{+}07, MHF02, NLJS01, NJJ06, OAKM07, RPP04, RSST07, SLJ05, TH03, TL03, UVKS05, VN00]. balticus [KK04]. ban [PBDP00]. banded [GBC^{+}03a, GBC^{+}03b]. Banggi [TCS05]. Bangladesh [CHBB04, HIRM06, eJSK09, RC06]. Bank [BD04, GHDM08, SR06, dSBAC06, WKCC06, AFVSR^{+}09, BTC08b, FKW06, FRF08, GHCPPLC^{+}05, KMB01, LG02, MGR^{+}02, Ove02]. banks [MR08a]. bar [BW08, McP02]. Baras [Mil02b]. barbarus [EKU02]. barbata [TCY01]. barbatus [FSB^{+}02, MTU^{+}04]. Barcelona [MDS03]. barcoding [HSW09]. Barents [BG01, Din05, GLH07, HF07, HMM^{+}09, LDK^{+}07, NS06, PTPS09, TH03]. Bari [She02]. Bariloche [VLD^{+}00]. barnacle [MF03]. Barnard [McI08]. barndoor [GHDM08]. barramundi [Bal09, Dou04, SM04b]. barred [GAFA05b]. Barrier [BMBM05, BPW^{+}03, BPH^{+}06, Fri07, Gri03, GWB07, HWW^{+}09, HW00, MLP^{+}08, NCW00b, NCW00a, PBW^{+}09, WMB08, WMD07, GK03, Sei07]. bartramii [BI05, CC03c, CCT^{+}08, IMOO06, YM00]. based [AAFLP^{+}07, Ber07, BCMM09, BG04, But03, CTEARD07, CNH09, CWJS09, CY06, CA07, CRSM04, Dun06, ERG09, FCC^{+}09, FML^{+}07, FSPWG08, HH06, HFHO07, HPK07, IMOO06, JLHO01, Jun08, JCJ^{+}09, KKR^{+}05, KW05, KP07, Kir01, Koe08, LNS08, LMR^{+}05, LMR^{+}07, MAASMB05, MAdAMASCM09, MFB08, MK09, MOK^{+}05, MF03, OJ06, Pál03, PBM06, PK06, RSP^{+}00, RM05, RFWH09, RDEACT07, Smi03, UMK01, WK01, WKM02, WIKM05, WG05, WBMM05, YLYL09, YJV^{+}03, ZKG^{+}09]. bases [HP02]. Basic [KW05, PKT^{+}02]. Basin [VKD03, Eze02]. basins [LDK^{+}07]. basis [AGS^{+}08, Cad00a, DBK07]. basking [PM09]. Basque [MPG^{+}08, PLSI09]. bass [CSSO03, FML^{+}07, HCH09, JWR05, LDE^{+}02, MM06a, MMB07b, MRT03, NGM^{+}09, OSC^{+}05, PPS05, PBLP08, PKP04, SMS02, WSC08, ZSP03]. bassensis [YEA07]. batch [GSMT04, YFT^{+}02]. bathyal [MCL04, PGS^{+}08]. bathymetric [CPCD05, GD03, MBG^{+}07]. Batoids [TCPC06]. Bay [AAA^{+}06, BK02, BP01b, Eer04, GBL^{+}05, GMY^{+}09, GKH^{+}01, LR00a, MR05, NT05, Row02, SKF08, TVB^{+}09, WHML04, AMUE01, AC05, BKD^{+}02, CMPM01, DUR05, FML^{+}07, Ism03, IYI07, KA05, MGTB08, Pou01, Sid03, UL01, VOS01]. Bayesian [APP^{+}07, BQ06, IS07, PLPN09, PCS05a, SE06b, Tri06, WLY06, WL06a, ZP05, ZLC09]. bays [FAW04]. BC [EMN07]. Be [Mil03a, ABRB00, AC00, DO06, IWW03, Kes00, LOT07a, MC00b, ROS04, YSS^{+}09]. Be-all-you-can-be [Mil03a]. Beach [PPH01, BWM07, CDC03, GKH^{+}01, GK03, MB07]. beach-seine [GK03]. beach-seining [GKH^{+}01]. beam [AR05, ASIW05, ABRB00, AC00, BMSJ05, CXMM04, DBAÖ06, FP02, GC03a, GCL03, Kru04, Mat07, MSSJ05, MC00a, MC00b, Pol00, RBH^{+}08, RJ05, RDH05, Sch09, SW00, WRGS09, vM03, vMBGF05, vSM02c, vSM02a, vSM02b]. beam-trawling [RBH^{+}08]. beaming [KS02, RHF^{+}08]. Beamish [Fra07]. beard [CP00]. bearing [TMK^{+}06]. became [SP09]. beds [LLV^{+}08]. before [Cad00b, EH04b, FFM08, SLB04, SS00a]. before-after [FFM08]. beginning [LAL07]. behavior [AKA06, AAKT07, KLS^{+}07, MSSJ05, PL04, Qui07, RB06, Rye08, STH03]. Behavioral [ROS04, WSC08]. Behaviour [Cur00, GFSM07, MNML00, ANS^{+}01, GM00, GSL08, HBBG06, He03, HCH09, JSO08, JHGS04, KJEF03, KW05, KMK09, MvdK06, NGM^{+}09, ORVMP06, ÖW02, PTE09, QPR08, RSM05, SK08a, SKP03, YEA09, McI09b]. Behavioural [MSVF05, AKC^{+}09, CAH^{+}03, DNN^{+}03, Mar07b]. Beibu [CQJX08, CXQ^{+}09]. Belgian [PCV04]. bellied [WMS04]. beneath [IJ06]. benefit [MGTB08, WLDH08]. Benefits [BR00, KS02, GFvdV06, PWP02, SBFR00]. Benguela [GEH^{+}07]. benthic [BPW^{+}03, CC09, CC01a, RJ05, RGSB08, WRGS09]. benthopelagic [All01]. benthos [DDP^{+}08, FP02, RJ05]. bentincki [CBC02, CRC^{+}07]. Benue [NJL^{+}00, NWML00]. Berg [Gon06]. Bergen [Kor04]. berglax [KN01]. bergylta [FMB^{+}05]. Bering [BMB06, McI02c, BQ04, BMB05, DA07, SWC^{+}05, Zhe03, vSS05, vSSK07, vSS09]. Berlin [Ano00e, ABF08]. Berycidae [DG02]. Beryx [DG02, RLG^{+}01, SÁR^{+}05]. best [BZRO05]. Bethesda [Lon02a, McI02a]. better [vOPvDM02]. Between [EPA00, ABC^{+}08, AFG^{+}05, ACG^{+}01, AM02, BM09, BSG02, CCMS^{+}03, CXMM04, DCM^{+}05, Dun09, EGJ00, ESB09, FGOF08, GDDJ09, GJBY05, Gro00, HWB06, HO05, HMM^{+}09, HK02, JDMN00, JPVKS^{+}01, JJKN^{+}03, KHMS09, LPHH09, LSW06, LWM07, MLH^{+}05, MPQ07, MBM04, MJS00, MGJB09, NGF06, OCT^{+}03, OMW^{+}03, OM07, ODR00, OLN^{+}02, PPL04, PB04, PBM06, PKK^{+}06, QDH09, RL03, RMN08, RDM^{+}09, SdPFAL09, STI08, SPS^{+}07, Sho09, SRH^{+}02, SFCK03, TYM^{+}01, TOK^{+}05, TOÖD06, Tze04, VYMGM^{+}09, vSS09]. between-haul [HO05, MBM04]. Between-reader [EPA00]. Beverton [Jen00]. Beware [PCHM08]. Beyond [DDP^{+}08, Dic04, NBF09]. Bhavna [McI05a]. bias [KM08b, Kop06, PKK^{+}06]. biased [GM07, HBJ03, MGJB09, PIdlHGV05]. biases [AB09a, BACP09]. BIC [WL06a]. biddy [IMSS06, IOS07]. bifasciatus [GAFA04b]. bifurcations [VVH04]. Big [TB04, BC02, MK06, WMS04]. big-bellied [WMS04]. big-eye [BC02]. bigeye [AdBP02, BDB^{+}03, CHL^{+}06, CHW^{+}08, SYS^{+}08, WMdSB09, WMdS09]. Bight [APS07]. bigtooth [BSAM09]. bilinearis [HC00]. Billard [Van01]. billfish [BDBD09, OA04]. Billions [Hol09]. binary [OF03]. binoculata [ESC08, MK06]. binomial [MLCGRV07, OF03, PL06]. bio [CAM^{+}08, GC03b, TTP^{+}09]. bio-economic [CAM^{+}08, GC03b, TTP^{+}09]. Bioadhesion [Ano01e]. Biochemical [DNN^{+}03]. Biodiversity [McI06a, HMR^{+}08, KAO04, McI06a]. Bioeconomic [PVMdL05, YLD^{+}05, GGVLSR03, KPL^{+}07, KPL^{+}09, MMS09, ULDG02]. bioelectrical [WH08]. bioenergetics [PF00]. Biofilms [Ano01e]. biofouling [SSS^{+}03, Ano01e]. Biological [Jak01, LZCR08, PGS^{+}08, SND^{+}08, SH09, BAEF06, Boo04, CCH^{+}00, CPSBRD^{+}02, CXCD07, DDP^{+}08, GC04, MCM^{+}08, Mag01, MFC^{+}08, ØT02, Sid03, TJ05, Tri05, UM02, VSS07]. biologists [Cer00]. Biology [Dem04, GAFA04b, McI06b, McI09d, Mil08, Nar06, Ore03, PdAO02, RMF^{+}06, Van01, BSMFH07, CMJ05, CUT06, CPT06, CMC09, ESC08, Eze02, FB06, GRPLE05, GAFA04a, GAFA05b, GAFA05a, HMB^{+}00, HAO00, IOS07, IYI07, MSN01, Mar06b, MBCMNMHH01, MCS05, NNA01, NNA02, PK03a, PGMM01, RRJ08, RWMBHH06, SR07, Ste07, TJ05, TLH07, Vil01, WSS02, WGDP06, XBS^{+}07, Yil02, Ell00, Sut02]. bioluminescence [JGB^{+}06]. Biomass [NMHHMB^{+}00, PPdA05, AFJC08, BM03a, BB01, CRC^{+}07, DV00, Dia00, FH03a, FGJD08, GM07, GHCPPLC^{+}05, GD03, HCM05, Jen02, JAKM05, KM02, MVAMB07, Mat07, MK01, ML04b, Orl03, PBDP00, PK03b, SE06b, SEPY02, SWPY05, YSS^{+}09, ZC02, vSSK07]. biometric [OSG08, SÁR^{+}05]. biota [PBW^{+}09]. Biotechnology [Ano01e, Lai00]. bird [Løk03]. bird-scaring [Løk03]. birdbeak [MF00]. Birkeland [Ano00c]. Biscay [AMUE01, AC05, CMPM01, DUR05, FML^{+}07, KA05, MGTB08, Pou01, UL01, VOS01]. bivalve [Cer00, CC01a]. bivalves [WHWM07]. Bivalvia [CC01b, TSM03]. Bizerta [LBABH09]. Bjørndal [Sym08a]. Black [GCMJ08, BACP09, CCM^{+}09, GMY^{+}09, GCJ07, LD08, LQD^{+}08, MAL02, MD02, PGS^{+}08, SMS02, DS07, Kno07, LBMCE00, TOÖD06]. blackspot [CPT06, GAF06]. blacktail [MSA00]. Blackwell [Cur00, Loc02, McI03a, McI05b, McI07b, Mil02b, Mil03b, Mil08, She02, Sol05, Sut02, Sym07, Sym08a, Sym08b, Win02]. blacodes [GOAARJRP09, PGF07]. bladder [MKGH05]. bladders [NGM^{+}09]. Blanca [Caz00]. Blankenship [McI05b]. bleach [WVW^{+}08]. bleeding [MWB^{+}08]. bleekeri [Tia09]. blennoides [CP00]. Bloch [CMC^{+}06, Dou04]. blocking [BR00]. bloodworm [SABW09]. blotched [DKGC00]. blue [ÅCST01, dSHG08, ASD09, BH01a, BdL05, CS09, CMPM01, CTB01, CAH^{+}03, CATH04, DYWHZ04, FAL^{+}01, FM04, FB06, Hen09, KMB01, LSH04, LMRQNM^{+}05, MNM00, Pál05, PKK^{+}06, WK01, WKM02, WIKM05, XK04, McP02]. blue-mouth [MNM00]. Bluefin [But08, KPBD08, MB08, ADP^{+}08, AD03, FDG^{+}07, GCC^{+}08, HCS^{+}03, Kar04, MM07, NP08, STI08, STCH09, WH08]. Bluefish [TOÖD06, RRJ08, RRJ09, WD02]. bluegill [FMC^{+}09]. bluemouth [ACPP06, MF09]. board [Ano04b, SPS^{+}07, Ano02h, Ano02j, Ano02k, Ano03k, Ano03i, Ano03j, Ano03f, Ano03g, Ano03h, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano04a, Ano04c, Ano04d, Ano04e, Ano04f, Ano04j, Ano04k, Ano04l, Ano04m, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06r, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q]. boat [DK05, LZCR08, LSW06]. boats [BCLP00]. Boca [How06]. bocaccio [PWHM06]. Bodden [GWR07]. Body [BMF^{+}10, ARLB07, Dun06, FKBF05, HCS^{+}03, Hur07, MM05, MPA08, RMN08]. bogaraveo [CPT06]. bogue [PBR05]. bohar [MM06a, MMB07b]. bola [EGB^{+}01]. Boleophthalmus [NT05]. bomb [PWHM06]. Bonaparte [TGH^{+}08]. Bonavista [Row02]. bone [KK09]. bonefish [DACS08]. bonito [CMC^{+}06]. Book [Ano00b, Ano00c, Ano00d, Ano00e, Ano01c, Ano01d, Ano01e, Ano01f, Bar00, Coo00a, Coo01, Cur00, Ell00, Gra02a, How03a, How03b, How06, Kes00, Lai00, Loc02, Lon02a, McI01, McI02a, McI02c, McI02e, McI02d, McI02f, McI03a, McI03e, McI03b, McI05a, McI05b, McI05c, McI06b, McI06a, McI07b, McI07a, McI08, McI09a, McI09b, McI09c, McI09d, Mil02a, Mil02b, Mil03b, Mil08, Ore03, Pie09, She02, Sol05, Ste03a, Ste05, Sut02, Sym07, Sym08a, Sym08c, Sym08b, Van01, Win00, Win02]. Books [Cur00, McI02d, Win02, She02]. Boops [PBR05]. bootstrap [CSR06]. borealis [KCKP02, PAR^{+}03, Wie04, WSP06]. Borgese [McI03e]. born [BVH^{+}08]. Bornholm [BSM07, VKD03]. boscii [LPP02]. Bostock [McI05c]. Botany [GKH^{+}01]. Bothnia [JJJ03]. bottlenose [GGA08]. Bottom [FCG02, MVV^{+}01, WSP06, AE03, AB09a, ABRB00, AC00, BSC^{+}08, CFFP07, CRSM04, DWW^{+}02, FCML05, HdS07, Hal02, IJ06, JR07, LR08, LIW04, MAG^{+}09, MC00b, NHT07, SLB07, SDM04, SE06a, SOS02, Som03, TÖT04, WK08, WPOB09]. bottom-trawl [FCML05]. bottomfish [ZDB^{+}08]. bottoms [DMC^{+}09]. Boulenger [Chi00, Man00]. Bouvier [He05, PGMM01]. Bowdich [AAHE04]. boyeri [LS00, TGS^{+}00]. brachydactyla [CF07]. Brachydeuterus [BC02]. Brachyura [FW06, RMA^{+}05]. brackish [KAN00, Lan04]. brackishwater [Mos00]. Braithwaite [McI09b]. Brama [BSAM09, LE01]. Brandt [McI07b]. bratramii [TCC^{+}09]. Brazil [MAG00, SB01b, BH01b, BNL^{+}08, GCS^{+}06, GMA02, HFB^{+}06, LHMO09, LSH04, MHS^{+}05, MGNB05, PA05, PW05, PPdA05, SB05, VG01a, VH06, VD06, dSH01]. Brazilian [CdCBdR03, FFL09a, FFL09b, Kru04, MGNB05, PdAO02, Vas03]. BRDs [HJ07]. break [MHS^{+}05]. breaking [AOMK^{+}04]. bream [GBC^{+}03a, GBC^{+}03b, GMY^{+}09, GCJ07, GCMJ08, LE01]. Breeding [EKU02, GFM02, LPG^{+}09, PPH01]. Bremec [VL08]. Breton [MMH05]. Brevoortia [RT07, VSS07]. Bridle [Som03]. bridles [Som03]. brief [Bur02]. Bringing [DDP^{+}08]. brink [PZH02]. Bristol [Sid03]. British [BNSA05, Fra07, ZP05, AP05, BNS07, BCM08b, CB00, CK08, MK06, PZH02, YJV^{+}03, ZC02]. broad [MLH^{+}05]. broad-scale [MLH^{+}05]. broadcast [BLR^{+}08]. Broderip [VL08, BLH04]. broodstock [PAP04]. brook [WN03]. Brown [Ell00, ARLB07, ALN03, AKK^{+}03, CHN08, GC03a, Gra03, JR02, JRML00, LF02, LMASHV^{+}03, NCW00a, Pol02, PCV04, PDV05a, PDV05b, RH04b, RMRB06, WN03]. brown-shrimp [GC03a]. brown-stripe [NCW00a]. Brunei [SG04]. Brünnich [CP00]. Buccinanops [Nar06]. Buccinum [MB04, PMAK07, SALnM07]. budegassa [CCM^{+}09, DALP01, GRPLE05, LD08, LPDA01, LQD^{+}08]. budget [KJ08]. Buenaventura [DOQ08]. Buoyancy [ÅCST01, VN00]. burbot [HK02]. Bureau [McI09a]. Burkenroad [RRAS03]. Burrishoole [BPD^{+}04]. Burundi [vZRMvD02]. Bush [CK02, PCS05a]. Bushehr [NAD^{+}07]. butcheri [GCJ07, GCMJ08]. butterfish [WMTJ07]. Butterworth [MR08b]. Buybacks [Sym08b]. By-catch [Gra02a, Gra03, AAFLP^{+}07, ASIW05, CFE03, CF04, CAM^{+}03a, FCL^{+}05a, FCML05, FCL^{+}05b, FP02, GL05, JRML00, Pál05, Sva05, SRS08, VD06]. by-product [SLJ07]. Bycatch [BS06, BSC^{+}08, CGP05, ORVMP06, SMJS01, TF09, WLS07, YAM00, ZCdG06, BM08, BHM^{+}06, BKD^{+}02, BNL^{+}08, CXCD07, CTC^{+}06, CHDGM06, DKH08, FRF08, FHJL08, GFL03, GLK00, GKH03, GMS06, HJ03, HJ07, He05, HBKJ07, HGW^{+}08, KHMG06, LET07, Løk03, MVAMB07, MLCGRV07, PW05, RM07, SBFR00, SK08b, TGH^{+}08, WLDH08, WEK^{+}09, WGDP06, YKO09, ZG08]. bycatches [BTC09, MTHL06].

C [Ano00b, Ano00c, Ano00e, Coo00a, McI09b, McI09c, Mil02b, Ore03, Sym07]. C. [ODdGV02, RH04b]. C.-I [Coo00a]. ca [EMN07, MK09]. CAB [McI03b]. cabrilla [TT01]. Caddy [SS00b]. cadenati [PL02]. Cádiz [BPRM00, SGS02, SGB00, JSR04, VVH04]. caeruleus [MAASMB05, MAdAMASCM09, QVNMGM00]. cages [KFO^{+}04]. calamar [MHS^{+}05]. calamary [SLJ07]. calcarifer [Bal09, Dou04]. calcea [MF00]. calceus [CCB02]. calcified [NBF09]. calcium [KAN00]. calculated [Hor00]. calculating [LLJ^{+}08, TMK^{+}06]. Calculation [Pal08, Hur07]. calculations [HH06]. Caledonia [GLCC09]. Caledonian [Ano00d]. California [ACMMCEGM02, APS07, FOGKT08, LMRQNM^{+}05, MQVSN04, MBCMNMHH01, NE07, NMHHMB^{+}00, RMVJGMMV07, RBHP03, RCGMQV05, RAHHGM^{+}02, Vel03, ASAGR06, ARCLBPdP07, LRBH07, LC07, MSN01, Mar06a, Mar06b, MAdAMASCM09, QVNMGM00]. californiensis [LMASHV^{+}03]. Callianassidae [RW03]. Callinectes [FW06]. Callorhinus [MKB01]. camera [DSO^{+}05, HCS^{+}03, He03, KMK09, LWS04, LIW04, MC06]. cameras [YCB^{+}01]. campechanus [AFSF05, BLB07, EF02]. Campeche [BD04, GHCPPLC^{+}05]. camtschaticus [GFL03, GFS03, NS06]. Can [ABRB00, AC00, DJGC09, HMO01, IWW03, JI07, Mar07a, RGGB07, SOS02, SMSK03, YSS^{+}09, FSPWG08, GM07, LOT07a, Mil03a, MC00b]. Canada [Fra07, Gra02a, Tal09, BNSA05, BNS07, BFM08, BCM08b, CK08, JSB04, OJ06, PZH02, SWM09, ZP05]. Canadian [BM08, CMJ05, MS05b, Mil03a]. Canal [CHQ05]. Canarian [DSPLR06, TOBSJH06]. canariensis [HGF06]. Canary [ABFV02, HGF06, DG02, HGHLCH02, MRCH04, PL02, PLDS03, PMG^{+}06, PGS^{+}08]. Cancer [AA03, BEBA03, EBA^{+}03, UHN07]. canicula [RDH05, RCSFO04]. Cannibalism [ABL02, JMGC06, PCM07]. Cantabrian [RCSFO04, SPB^{+}03]. Cap [LZCR08]. capable [MCB^{+}07]. capacity [MT08, Paw03, RD03, RJ05, TPM03, VG01a, VSK03]. Cape [RB09, GEH^{+}07, HBBG06, MMH05, SE06b, VGBF02]. capelin [BG01, JHGS04, PTPS09]. capensis [Joh03, VGBF02]. Capiz [Bat04]. capoeta [TEYA02, TEYA02]. Captain [Skr08]. captive [CS07b, ZW09]. Capture [Ste08, CAH^{+}03, FCS^{+}05, GG04, HP02, JS05, ORVMP06, RB06, SMS02, SJ05, WSM02, YEA09, vSS05]. captured [CSSO03, GDFH07, HLL^{+}08, MNML00, MHWL01, SB07, SSIE06, VAYD04]. Carangidae [GAFA04a]. Carangoides [GAFA04a]. Caranx [WHML04]. carbo [PGS^{+}08]. Carbon [JSB04]. Carcharhinus [GCS^{+}06, JLC04, JCLL08, SDNL09]. Cardinale [PCHM08]. Caretta [BM08]. Caribbean [CMM^{+}07, GDAM07, GCS^{+}06, GBL^{+}05, GB07, ML04b, PR03, RW07, SCSC07, SR07, WM04, ZW09]. Carlos [VLD^{+}00]. Carolina [BB04, JDS00]. carp [AM03, Bas99, Bas00, RCA08, Van01]. Carpentaria [SBFR00]. Carpenter [McI05d]. carpio [AM03, RCA08]. carponotatus [NCW00a]. carried [JPVKS^{+}01]. carrying [VG01a]. cascading [CP06]. Case [DSK^{+}08, AGB07, Bat04, BTC09, BGH05, CR06, DDP^{+}06a, DDP^{+}06b, DDP^{+}06c, FKW06, FF01, GPG^{+}03, GB07, HKF^{+}09, IWW03, IPH05, KMF08, LFA08, MGTB08, MLH^{+}05, MS05b, MC05, MPG^{+}08, NBA^{+}09, OSG08, PKT^{+}02, PTPS09, PEAAdLJB06, PSTM06, Row07, RE07, RMA^{+}05, RR03b, SSS^{+}03, SSML02, TSTK06, VSK03, WD02, YFL08]. case-study [MLH^{+}05]. Caspian [FGJD08, FZHL09]. Castellammare [PBDP00]. Castellanos [MHS^{+}05]. Castelnau [GCC^{+}08]. Castilla [Sym07]. Catalan [ASA04, OQE03]. Catch [BQ04, CCB02, GK03, Kar04, KTL06, MOK^{+}05, SPB^{+}03, SABW09, TSTK08, WBHG01, AAFLP^{+}07, AP05, Aló09, ASIW05, AKC^{+}09, ARCLBPdP07, BB04, BEA^{+}08, BVW04, BS04, CFE03, CFH03, CF04, CGLZ04, CAM^{+}03a, CNH09, CSR06, CP08b, Chi00, CL04, CLGR^{+}04, DMA07, DPD^{+}03, DFD02, DKH08, EJA00, FH03a, FOGKT08, FCL^{+}05a, FCML05, FCL^{+}05b, FP02, FHJL08, GRFE06, GSS09, Gra02a, Gra03, GL05, GCJ07, GESY^{+}07, HMO01, HWB06, HJ03, HG01, HBW00, HBJ03, HBKJ07, HP04, Hen09, Her05a, Her05b, HH01, HR09, JPYW08, JS05, JRML00, KSS05, KM02, LAL07, LR08, LB08, LSW06, LWM07, MLJT09, MWB^{+}08, MROGVR09, ML04a, MP04, MDS03, MOGG06, MMN00a, MUO02, OF03, OKWM05, OAKM07, OMR09, OA04, OSC^{+}05, OR07, Pál05, PG06, PBLP08, PLPN09, PSvDPM01]. catch [PCMC08, PWTP00, RMN08, SMD04, SSB03, Sca03, Sho08, SMS^{+}06, SYS^{+}08, SLJ05, SSK^{+}06, Sva05, SRS08, THN^{+}04, TNL07, TSJ04, TMN07, iTM04, iTMI06, UVKS05, VD04a, VD06, WK01, WKM02, WL06b, WSP06, WPOB09, WB08, Xia04a, Xia04b, YAHAB01, YLYL09, YKM06, YKO09, YPM^{+}06, You03, ZDB^{+}08, vM03, vZRMvD02, LLJ^{+}08, Mes03]. catch-and-release [Aló09, AKC^{+}09, Hen09, THN^{+}04, TNL07, iTMI06]. catch-at-age [LB08, WL06b, WB08, YLYL09]. catch-per-unit-effort [CP08b, FH03a]. catch-per-unit-of-effort [ML04a]. catch-restricted [PBLP08]. Catch-Survey [Mes03]. Catchability [WMTJ07, ALN03, BYA^{+}08, FHI07, GQR03, GMV00, GMV01, MAdAMASCM09, NJE00, OKPR04, SKP03, SCM09, ZDB^{+}07, ZFJ03]. catchability-at-length [MAdAMASCM09]. Catches [AMAOMM04, AAT08, BTC09, HLHF03, AE03, ASA04, AAP^{+}09, BH01b, CGT08, GKH^{+}01, GJBY05, HHTE05, HBW00, HBJ03, HB01, KE05, KME^{+}02, LDK^{+}07, LLJ^{+}08, Meh06, MSTC04, MD02, OIK^{+}06, OGPDOM08, PAB05, RM05, RKEJR06, RYSN08, VG01a, VKČ^{+}09, WEK^{+}09, WRK06a, WRK06b, tHDC06]. Catching [McI07b, TEG03, LIW04, MvDM02, Paw03]. categorization [BMC07]. catfish [BW08, dGvZM^{+}06]. Catla [Bas99, Bas00, KK09]. Catostylus [PK03b]. caught [AC02, CAM^{+}03b, DPS^{+}01, ECEG09, FFL09a, FFL09b, GCJ07, GCMJ08, HSGF06, JRML00, MKB01, MGNB05, NGM^{+}09, PWB05, SMB^{+}08, SLJ05, TF09, dSH01]. Caulolatilus [BLC08]. cause [MBK02]. caused [CCG^{+}02, SB03]. Causes [SSGM03a, LC07, MBO^{+}07]. cavalla [BSG02]. cavity [ARLB07]. Cellular [Ano00b]. Celtic [ERG07, FAL^{+}01, KA05, Pou01]. CEN [DVAL^{+}09]. census [CARA^{+}07]. Centracanthidae [DKGC00]. Central [CUT06, How06, McI05d, PGS^{+}08, ATM08, AB09b, BCC09b, BCSA08, Bat04, BCM08b, CDC03, CLGR^{+}04, CBC02, CRC^{+}07, CPR^{+}08, DKGC00, KIT^{+}04, LLM05, LC06, MMV^{+}04, MHK^{+}06, MAG^{+}09, MMN03, MCS05, MKB01, NCW00b, NCW00a, PMSD06, PCAS04, SN01, SYS^{+}08, Vel03, VMPP07, BNSA05, CAM^{+}08, PLDS03, ŠV04]. Central-east [PGS^{+}08, MMN03, MCS05, PLDS03]. central-south [PCAS04]. central-southern [CBC02]. centropomid [Dou04]. Centropristis [SMS02]. Centroscyllium [Jak01]. Centrostephanus [BW02]. Centuries [LAL07]. Century [LDK^{+}07, GKUO07, LOT07a]. Cephalopod [Ano06-30, DPS^{+}01, PPH01, VG01b, GPG^{+}03, PAP06, RV03]. Cephalopoda [NNA02, BQN05, CUT06, FRGF07, Hat00, Lap02, MQVSN04, MCP01, NNA01, PdAO02, RW07, SN01, TVY01]. Cephalopods [dSH01, FCG02, FCAD^{+}08, HH06, PPdC01a, SCP01, TASF09, VOS01]. cephalus [KKR^{+}05, LRW09]. CEPHSTOCK [Ano06-30]. cernuus [Ogl09]. certainty [Lon07]. certification [She09]. cervinus [PLDS03, PLDS03]. Cetorhinus [PM09]. Cetti [DSPLR06]. Chaceon [ABFV02, PGMM01]. Chad [NJL^{+}00, NWML00]. chain [JTdH06, JHSG04, PPdC01a, GWR07]. chalcogramma [APC07, BQ04, Dou08, ODR00, RO00, vSSK07, vSS09]. Challenges [SCSC07]. chambered [BSWL06a, BSWL06b]. Chamelea [DNN^{+}03, MFAM05, PPL04]. Champlain [SRTP06]. Champsocephalus [MBDM05]. Chang [LD06]. change [But03, FM04, FHI07, HP09a, Mar07a, OGPC09, SLB04, YFL08]. change-in-ratio [FHI07]. changed [SB08]. Changes [BCC09a, Dah00, FK04, LG02, MCL04, APC07, AFJC08, CMH01, EH04b, EMN07, FÖB08, FHG03, Gri03, GLCC09, HMM00b, JPVKS^{+}01, KF04, LPG^{+}09, MTRM05, MWB^{+}08, MM04, MDGR05, ÖW02, RdFBS01, SOS02, TAS03, WY06, WMB08, Wie04]. Channa [KK09]. Channel [MB04, Kru04, PB04, DPS^{+}01, ERG07, FML^{+}07, RDH05, RPFR06, ULDG02]. Chaoborus [KLJ06]. chaos [KW05]. Chapman [Ano00c, DFD02]. char [HMB^{+}00]. Characterisation [GRFE06, SDM04, AM03, DJG^{+}06, JNDH00, NJL^{+}00]. characterise [BR01]. Characteristics [HLL^{+}08, He05, ABFV02, BAEF06, FAW04, GK03, KIT^{+}04, SH09, SJ05, Tri05, TDK^{+}05, VSS07, VBSK00]. Characterization [GLCC09, AD03, MFC^{+}08]. characterize [KM08b]. Characterizing [VMPP07]. Charles [Mil03b]. Charlotte [ZP05]. charr [EF01, iTM04, iTMI06, WFJB09, YMT^{+}06]. charts [Sca03]. Charybdis [AAH06, AKA06, AAKT07, SKF08]. Chatham [CAFT00]. ‘Cheap [KC06]. chemistry [ASD09, BMBM05, HBBG06, VBSK00]. Chen [AC00]. Chicago [McI03e]. Chichester [Ell00, Van01]. Chile [CSCC07, CBC02, CRC^{+}07, CPR^{+}08, GCN02, GESY^{+}07, LC06, OFOM08, PCAS04, SPS^{+}07]. Chilean [AA03, ACG^{+}01, BBL07, CSM^{+}09, CHSA02, JMGC06, LC06, OB02, OS05, QDH09, RV03]. China [CP08b, CHL^{+}06, CC01c, LFA08, WLL08, YST^{+}02, YFT^{+}02, YLCS^{+}09, WZDY06]. chinensis [COHP02, WZDY06]. Chinese [CLC08]. Chinook [CB00, Kop06, MK09, PB04, QDV05, VAYD04, WCM04, ZB00]. Chionoecetes [BFM08, CWS^{+}05, HMM^{+}01, MD09, Ste03b, Zhe03]. Chircop [McI03e]. chirp [ET00]. Chiswell [AFJC08]. Chitala [SND^{+}08]. chloride [CS07b]. chlorophyll [GMA02]. choice [PL04]. chokka [RB00, SLA00]. Chondrichthyes [VMQ08]. Choosing [BZRO05, Gro00, DDP^{+}06b]. Chorus [GCN02]. Chronic [CCA^{+}03]. chronology [OM07]. Chrysichthys [ODdGV02]. chrysonotus [WBMM05]. chub [IFH^{+}09, SWA08, WY06]. chum [SN09]. cichlid [Boo04]. cichlid-specific [Boo04]. Cichlidae [WBMM05]. cichlids [THN^{+}04]. circa [RdFBS01]. circle [BR05, CSSO03, KG06, OSC^{+}05, WEK^{+}09, YKM06]. circumference [GBM09, HPK07, OGK^{+}08]. cirrhosa [RMF^{+}06, SMD04]. cisco [YSS^{+}09]. CL [Lab05]. clam [CRG^{+}02, CCA^{+}03, GPI04, MFAM05, MJS00, PPL04, WRWK02, ZC02]. clams [CCG^{+}02, GML06]. Clarias [BW08, Eze02]. Clariidae [Eze02]. Clark [BFB^{+}03]. Claroteidae [ODdGV02]. class [BMC07]. classes [ANS^{+}01]. classification [Ber07, NBF09, vSM02c, vSM02a, vSM02b]. classifiers [PBR05]. classifying [TSTK06]. Claude [Ano00d]. Climate [ZCTA^{+}08, Bal09, COLB07, HP09b, TAS03, Tia09]. climatic [GWR07]. close [ASD09]. closed [AGB07, BPH^{+}06, PBW^{+}09]. closures [PA09]. Clupea [BBK08, FUM02, FBJM00, HHC06, NDCK09, NLJS01, Oja06, Ove02]. clupeid [CND02]. Clupeidae [AAHE04, PCM07, RC06]. Clupeonella [FZHL09]. Cluster [MBT08, AKA^{+}02]. Clyde [BWMA02, Moo03]. co [FR08, MB07, Tri05, VKD03, WCM04]. co-management [MB07, Tri05]. co-occurring [VKD03]. co-transported [WCM04]. co-variation [FR08]. coarse [PKŘ^{+}09]. Coast [RRJ08, RRJ09, Ano00e, BCC09b, BAEF06, BTC09, BEBA03, BBJG04, BGB06, Bur02, BCM08b, CDC03, CFE02, CFE03, CFH03, CFG05, COHP02, CARA^{+}07, CE06, CRSM04, CdCBdR03, CCD01, DALP01, EBA^{+}03, EBV^{+}03, FMCS05, FCAD^{+}08, GJSW04, GBC^{+}03a, GBC^{+}03b, GR00, GTB04, HGHLCH02, HP09b, KIT^{+}04, LOT07a, LCFP^{+}08, LMRQNM^{+}05, MHS^{+}05, MTC06, MBG^{+}07, MF09, MACS08, New02, OL07, OQE03, PH09, PNJS06, PTPS09, PTR^{+}05, QP00, RFWH09, RCGMQV05, RCG02, SGVM02, SLA00, SHDO04, Ste03c, SAK05, SWA08, SOKS08, TYM^{+}01, TC09, Vil01, VSK05, YFL08, PDH08, UHN07]. Coastal [Sym08c, ASM^{+}09, ABAA^{+}07, ABCB^{+}07, ASIW05, BF00a, BC02, BNSA05, BNS07, BBLL06, BP01a, BHJ02, BP08, Bla06, BLR^{+}08, CC03a, CC01a, CMAOAA05, CCCC04, CMOM00, ED07, Fra07, GC03b, GRC^{+}09, JHK08, Jun08, JCJ^{+}09, KSS05, LR00a, LK00, LZC^{+}08, MLH^{+}05, MOOH00, MCAO^{+}05, MAL^{+}01, Nar06, NAD^{+}07, NV05, OMF^{+}03, PMG^{+}06, Par04, PMAK07, PCD03, PSPCVCD09, RD03, RGSB08, SN09, SC08, SMM07, SST^{+}06, SSG06, SBF^{+}08, TCPC06, TF09, Wes02, WF07, vOPvDM02]. CoastFish [Ano07a]. coasts [ASA04, OJ06, PSS^{+}07, vHAN^{+}09]. Cod [FRF08, OKWM05, AE03, ABD^{+}09, AFVSR^{+}09, AD00, BM03a, BP01a, BG01, BLR^{+}08, BC04, BS04, CFE02, CFE03, CFH03, CAJ00a, CAJ00b, CS04, CTK^{+}07, CTB01, CAH^{+}03, CATH04, CCBG02, EMKG07, FR08, FHJL08, GM00, GFL03, GOF^{+}04, GF06, GHDS08, GLH07, Hal02, HKG^{+}08, HMH^{+}04, HBM^{+}00, Hen09, Her05a, Her05b, HO05, HPK06, HO06, HFHO07, HPK07, HKF^{+}09, HH01, HWM02, HMRG08, IJD04, JSB04, JSO08, JIDM02, JIAD03, JIGI06, JCJ^{+}09, KI04, KI05, KOO04, LR08, LR00a, LK00, LR00b, MBO^{+}07, MHF02, MTHL06, MGR^{+}02, MR05, NHT07, OMW^{+}03, OKF06, OGK^{+}08, ØA08, PEB03, PP03, PBM06, PCMC08, PCHM07, RBD02, RL03, RR03a, RSST07, SF05, SGL08, SW02, SCM09, SBF^{+}08, SLJ05, SSC^{+}03, Tal09, TH03, TL03, VN00, VBSC08, VKD03, WM00, WF07]. cod [WPOB09, WR05, WGGN06, WNRd00, vM03, vSSK07]. Cod-end [OKWM05, GOF^{+}04, GF06, HO05, HPK06, HFHO07, HPK07, JSO08, KOO04, OMW^{+}03, OKF06, OGK^{+}08, RBD02]. cod-ends [Her05a, Her05b, HO06]. Code [Cad00a, Qui07]. coded [BLB07, CB00, CMOM00, CK02]. coded-wire [BLB07, CB00]. Codend [FCM07, Pol00, BSS06, Bet04, CRD06, FCG02, GBM09, KI04, KI05, MMB^{+}07a, MHMP01, MH02, MSB^{+}08, MTU^{+}04, ÖW02, SLJ05, TÖT04]. codends [BMK^{+}04, DBAÖ06, GSL08, GLS^{+}09, HC00, He07, JIGI06, KFM08, ÖT03, ROS04, SLB07, SLPF08, TASF09]. coefficient [AC02]. coefficients [MBT08]. Coelho [How03b]. Coherent [ZR08]. Coho [QDV05]. cohort [CCT^{+}08, HH06, IMOO06, Jun08, WMD07, XW07, YM00]. cohort-specific [WMD07]. cohorts [CC01c, KWK^{+}09]. coindetii [CUT06]. coioides [GAFA05a]. Cold [Ore03]. colhuapiensis [CS08c]. colias [CTB01, CAH^{+}03, CATH04]. collaboration [IK07]. collapse [EH04b, Lon07, Ove02]. collapsible [AAH06, AAKT07, AAT08, JPYW08]. Collares [How03b]. Collares-Pereira [How03b]. collected [AKA^{+}01, VG01b]. Collection [Gue05, LSW06, PKT^{+}02]. College [Coo00a, Gra02a, McI06b]. Collett [Jak01]. Colloques [Gue05]. Cololabis [TAS03, USKK04]. Colombia [BBV07, GDAM07, Rue07]. Colombian [PR03]. color [YKO09]. Columbia [BNSA05, BNS07, BCM08b, CK08, Fra07, Mil06, PZH02, Ste03a, WCM04, ZP05, AP05, CB00, MK06, ZC02]. comb [Kno07]. comb-jelly [Kno07]. comber [MSA00]. combination [Bra04, EH04a]. combined [AIC08, HPK07]. combining [PRdPG08]. coming [Sym06]. commensals [OS05]. Comment [BCM08a, Fra07, MR08b, MC00b, VL08, BNS07, Hol01, SM04b]. commentary [But08]. Comments [AC00, Sve14]. Commercial [KFM08, MSM^{+}08, SRA^{+}02, AKA^{+}02, ARCLBPdP07, BB04, Bal09, BS06, BVW04, BYA^{+}08, Bra04, BLH04, BNL^{+}07, CNH09, CAFT00, CdMLGC09, DPD^{+}03, FBV06, GKH^{+}01, GK03, Gri03, HFBM08, HCM05, HP04, IJ06, JI07, KM02, KI04, KI05, LHMO09, MMV^{+}04, MSVF05, MJS00, OIK^{+}06, PPL04, Pou01, QPR08, RM05, RFWH09, RDH06, RYSN08, SG01a, SSIE06, TASF09, UVKS05, VL08, VAYD04, WK01, WIKM05, WMB08, WPOB09, dGvZM^{+}06]. commercial-vessel-based [RFWH09]. commercially [ESC08, GWB07, MTC06]. commerson [CMAOAA05, GAOMC06, GAFA05b, MCAO^{+}05, vHMAO^{+}06]. commersonnianus [GFvdV06]. Committee [McI02e, McI02f]. Common [VMQ08, AM03, BACP09, Cad09, CLH^{+}08, CS07a, CBC02, CRC^{+}07, Ism03, IYI07, Lev05, MMP^{+}02, RCA08, SALnM07, She03, Sym09]. communication [eJSK09]. communities [BB01, BHM^{+}06, BGH05, BPH^{+}06, JMGL04, LZC^{+}08, MVAMB07, RJ05, SB01a, YFL08]. Community [CTEARD07, CWJS09, JK07, LG02, MF03, MMLR08, OMR09, PSvDPM01, RE07, VOS01, XJ05]. Community-based [CTEARD07, CWJS09, MF03]. Company [McI09c]. Comparative [AAH06, Gob00, SE02, ABC^{+}08, BGH05, SCW04, VAF04]. compared [AAT08, GCL03, JIGI06, SK03a, dGvZM^{+}06]. Comparing [DYWHZ04, PLPN09, VKD03, FF01, MAG^{+}09, OR07]. Comparison [BSAM09, DCM^{+}05, DFD02, GBFG06, JDAM08, KK09, LP00, MB08, NGF06, OCT^{+}03, OMR09, ÖT03, Pra02, QDH09, RW03, RVS02, SdPFAL09, TOK^{+}05, WKM02, WL06a, WSS02, Wes02, WN03, YKM06, ATB^{+}02, BVW04, BPH^{+}06, BMSJ05, CS07a, CBGP06, CTB01, CA07, DKH08, EF01, EGJ00, ESB09, GHS04, GJSW04, HdS07, Hal02, HFS02, HMM^{+}09, HR09, JPVKS^{+}01, KMELC05, LDK^{+}07, LLJ^{+}08, LSW06, MLJT09, MJS00, PCS05a, RM05, SFCK03, vHAN^{+}09, vT04, vSS09, He07]. Comparisons [LCTB^{+}03, Dun09, HWB06, MC06, Sho09, TY02]. compartment [KMK09]. compensation [KG04, NDCK09]. compensatory [SLE09]. competition [AM02, SME02]. compilation [PKT^{+}02]. complex [AR05, Ove02, PPGD08]. Complexities [Ste03a]. complexity [CC09, SHBÅ07]. complexone [Dou08, LZZN09]. compliance [Qui07, TB04]. component [MBG^{+}07]. components [CFFP07, LP06, OSB09, SGS02]. composite [BKD^{+}02, EH04a, JCW05, PH04a]. Composition [TDK^{+}05, APC07, BSAM09, BEA^{+}08, BWMA02, CT00, CCA^{+}03, FCAD^{+}08, HBJ03, HP04, HMAOAS07, MR05, MMN00a, MB08, PSvDPM01, RYSN08, RAHHGM^{+}02, SW00, SME02, SMS^{+}06, TGH^{+}08, TMN07, WSP06, WH08]. compositions [Dah00, MGNB05]. compute [KM02]. computer [Fab06, SD01, WSS06]. computer-assisted [Fab06]. concentrations [TYM^{+}01]. concept [BGH05, MHMP01, SAA^{+}03]. conceptual [CS02, GG06]. conchilega [RBH^{+}08]. concordance [ODR00]. concurrent [Meh06, SF05]. Condition [LR00b, Chi00, FÖB08, GCH^{+}07, Hur07, KAHR02, MPM^{+}06, MM05, RL03, WH08]. conditions [ACE^{+}08, AMUE01, BGP^{+}04, CC01a, HFS02, HKF^{+}09, KNJJS06, LC06, LWW02, NJJ06, ODR00, Orl03, OGPDOM08, RSST07]. Conduct [Cad00a, Qui07]. conducting [CA07]. Conference [Ano07a, Ano07w, Ano09r]. configuration [CFH03, PGF07, STK^{+}05, TSS^{+}04]. configurations [OSC^{+}05]. confirmation [WKCC06]. conflicts [KJS^{+}06, eJSK09]. confronted [GLCC09]. congeneric [WSC08]. conger [CMC09, OMF^{+}03, CMC09, KIT^{+}04, OMF^{+}03]. congruency [MGJB09]. connectivity [Abl06, BLR^{+}08, SK03a]. consequence [Dah00, NV05, TSJ04]. Consequences [Eer04, FGOF08, HCH09, TKFJ09, AKC^{+}09, GCD^{+}09, HKG^{+}08]. Conservation [Ano08l, Güç08, Lon02a, Mil08, He08b, IK07, MLP^{+}08, PCMC08, Row07, How03b]. Considerations [AMC^{+}08a, Gra02a, Sei07, HP04, ST03]. consistency [CRSM04]. consistent [GPI04]. Constant [HH01, MAdAMASCM09]. constraints [DJGC09, KJ09, MBT08]. Constructing [YPDS05]. construction [AAB05, Cam04]. consumed [DV00]. consumption [DPS^{+}01, PF00, Sva05, SRS08]. contact [SW01, Som03, WK08, Wil09, vS04]. content [HHC06]. Contents [Ano01g, Ano01h, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano05c, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano08a, Ano08b, Ano08c, Ano08e, Ano08d, Ano09e, Ano09f, Ano09g, Ano09d, Ano09b, Ano09c, AB08b, Ano03-33, Ano03-34, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano05-30, Ano05-31, Ano05-32, Ano05-33, Ano05-34]. context [Gav09]. contiguity [MBT08]. continental [AAFLP^{+}07, CAFT00, FFM01, FCG02, HLHF03, VE03]. continue [OL07]. continuous [BEBA03]. contribute [FSPWG08]. contributing [NGF07]. Contribution [GJ09, WM09, OIK^{+}06]. contributions [CS02]. control [AM03, CMAS08, DB08, FFM08, LZS01, MEA^{+}03, RMVJGMMV07, Sca03, She03, SWM09, VVH04]. control-impact [FFM08]. controls [DDP^{+}06b]. controversy [Ano02-29]. conventional [AAT08, BMK^{+}04, GR07, KKJ00, MC06, WH08]. convergence [Ano00e]. Cool [Ore03]. Cooperation [DTSR07, Lon09, Tri05]. Cooperative [HGSP01, AFG^{+}05, Bat04]. Copadichromis [WBMM05]. copper [BM03c]. Coral [Ano00c, BMBM05, CGT08, GC03b, HW00, KVZ05, LMR^{+}05, LMR^{+}07, PSvDPM01, SK03a, SK08a, SMB^{+}08, WM04]. core [VBSK00]. coregonids [KG04, UMV^{+}09]. Coregonus [KG04, LWW02, Meh06, SG06, Sch09, YSS^{+}09]. cores [MK09]. Cormorant [BLKR09]. Corn [Daw08]. Cornish [RCA^{+}07]. Corporation [McI06a]. Correlating [vSSK07]. corresponding [NCW00b]. Corrigendum [BMB06, MG03a, SBD^{+}06]. Corrosion [Ano01e]. Corsica [MBG^{+}07]. Coryphaena [FOGKT08, MROGVR09, PMSD06]. Coryphaenoides [LDA01]. cosmopolitan [SB05]. cosmos [GL05]. cost [MGTB08]. Costa [FW06, DG08]. Costello [Ano00e]. costs [SGP02]. could [She03]. Coull [Ano01f]. Council [Coo00b]. count [CA07, KGB09, MSB07]. counterpoint [ZR04]. counting [LBKG08]. countries [Abl06, Bla06, SST^{+}06]. course [Sym05a]. covariance [Joh03]. Cover [Ano03p, Ano03q, Ano03r, Ano03s, Ano04j, Ano04k, Ano04l, Ano04m, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano06q, Ano06r, Gue05, MHMP01, MH02, McI02e, McI02d, McI07a, Sol05]. covered [MHMP01]. cownose [BSMFH07]. Cowx [How03b]. Cox [Win02]. CPUE [LLJ^{+}08, Sho09, Cam04, CGEHC07, Dun09, HK06a, JI07, KM08b, LD01, MS03, MOK^{+}05, NC04, Par04, QPR08, SG01a, Sho08, SOS02, SM04a, TCC^{+}09]. Crab [PNJS06, ABFV02, AA03, AKA06, BK09, BdL05, BFM08, CWS^{+}05, CF07, DA07, EBA^{+}03, FM04, FHJL08, GFL03, GFS03, HMM^{+}01, JPYW08, MMLT05, MD09, NS04, NS06, PGMM01, Sal02, Sid03, SWM09, SKF08, UHN07, XK04, ZK03, Zhe03, McP02]. Crab2001 [Ore03]. Crabs [Ore03, AAH06, AAKT07, AAT08, BEBA03, BMB05, BMB06, CS09, DYWHZ04, He05, OIK^{+}06, Ste03b, Wah03]. Crangon [ALN03, Gra03, JR02, LF02, Pol02, PCV04, PDV05a, PDV05b, RH04b]. crappie [BACP09, MM04]. Crayfish [Sut02]. CRC [How03a]. Creel [RM05, Wil09]. creels [MAG^{+}09]. Cretan [TT01]. cretense [AMS08, TOBSJH06]. Creus [LZCR08]. Crisis [Mil02a]. criteria [GM06, Gro00]. criterion [WL06a, WB08]. critical [Cor02]. croaker [NV05, VH06, YTK06]. cromis [MAL02]. Cross [AAHE04]. Crustacea [DMC^{+}09, FFM01, RMA^{+}05]. Crustacean [TFA03, Cad03, CFE03, CF04, FCL^{+}05a, FCL^{+}05b, FCM07, FHG03, MDS03, MMN00a, QDH09, SSB03, SA03]. crustacean-trawl [FCL^{+}05b]. crustaceans [GCL03, McI09d]. CSA [Mes03]. Cuban [BA06, CdMLGC09, GYMdL^{+}06, PVMdL05]. Cubitt [MR08b]. cucumber [PK06, SK08a]. Culture [Van01]. cultured [GMY^{+}09, PAP04, Tri06]. Cumulative [Kop06, Sca03]. Cunningham [McI05c]. curema [SR07]. current [BGM^{+}09, DBB00, PCB01, Ste03b, KWK^{+}09, WYR06]. currents [KMB01, WNRd00]. Curtis [Sym08b]. curve [Bas99, Bas00, CHSA02, DFD02, TS09]. curves [Jon00]. CUSUM [Sca03]. Cut [FMC^{+}09, DG08]. cutaway [RDH06]. cutting [iTMI06]. cuttlefish [DR01]. Cuvier [AHAAD01, Caz00, Gon06, MMN03, MCS05, PJR^{+}05]. Cyclades [SE02, SME02]. cycle [HT05, OGSG07, Vil01, WYR06, YTK06, YFT^{+}02]. cycles [AR01, SR00]. Cyclopterus [ATB^{+}02]. cygnus [CCMS^{+}03, MSTC04, PMSC03]. Cymbula [AGB07]. Cynoscion [Caz00, MM06b]. cyprinid [BNS00]. Cyprinidae [BP06]. Cyprinus [AM03, RCA08]. Cyrino [McI09a].

D [Ano01c, McI08, McI09a, McI09c, Ore03, Sol05, Sym08a, Sym08b]. D. [Hor02]. dab [FAN03]. dactylopterus [RMCdS06, ACPP06, MNM00, MF09, RMCdS06]. Dahm [McI07b]. Daily [Dou08, MBDM05, ACE^{+}08, CS07b, CRC^{+}07, Jun08, MK01, MÁ08, MN00, NMF^{+}08, SKMT02]. Dalälven [BLKR09]. dall [PMAK07]. Dam [Yil02, AAB05]. Damage [Güç08, KSS05, SBFR00, SSK^{+}06]. damming [BA06]. Danish [CL04, CR06, EMN07, FHM09, MH01, NM06, VSK03]. dark [ODR00]. Darss [GWR07]. dart [BW08]. Darussalam [SG04]. Dasyatis [Ism03]. Data [WGDP06, AKA^{+}01, AIC08, APP^{+}07, Bas99, Bas00, BEBA03, Bet04, BVW04, BQ06, BS04, CWL00, CWJS09, CP08b, DPD^{+}03, DSK^{+}08, DFD02, ED07, Ehr08, FDG^{+}07, FML^{+}07, GNL^{+}09, GM00, GDFH07, GK01, GKF03, HMM00a, HP00, IMOO06, JL00, Jon00, KM02, KJ08, Koe08, Lab05, LDK^{+}07, LRBH07, MVAMB07, MTA^{+}02, MGR^{+}02, ML04a, MP04, Mes03, MOGG06, MD09, NC04, OF03, OSG08, ØT02, Pal08, PH04a, Pou01, PWTP00, PSTM06, QPR08, RSP^{+}00, RMA^{+}05, RSMZ06, Sho08, SLJ07, SM04a, SE06b, SYS^{+}08, SMB^{+}08, SG01b, Tal09, TCC^{+}09, TPC05, UVKS05, VMPP07, WKM02, WIKM05, WYR06, WL06b, WMdS09, WRWK02, Wil09, XM00, Xia04a, Xia04b, YLYL09, ZLC09, vT04, vSSK07]. data-deficient [ED07]. data-limited [SE06b]. data-loggers [GNL^{+}09]. data-poor [DSK^{+}08, VMPP07]. data-sparse [AIC08]. database [GSS^{+}06, WRK06a]. date [AC05]. dates [QVNMGM00]. Datta [Ano01c]. David [McI02d]. Dawe [Ore03]. day [AB09a, GCL03]. daytime [TKFJ09]. DC [Coo00b, McI02f]. Deania [CCB02, MF00]. Death [Ano00c]. decadactylus [DG02]. decadal [MBO^{+}07]. decades [BCC09a]. decapod [AB09a, MMN00a, GWB07]. Decapoda [ASA04, COHP02, DMC^{+}09, FFM01, KTL06, RW03]. decapods [SK03a]. Decapterus [JCS01, SE06b]. December [Ano00n, Ano02p, Ano05-28, Ano06v, Ano07o]. decision [KW05]. decision-making [KW05]. decisions [CR06, Gav09]. Decline [SST^{+}06, GSS09, HP09a, MBK02]. declines [FHG03, HBKJ07, WY06]. declining [IFH^{+}09]. Decomposition [Pie09]. decussatus [CC01a, CC01b]. Dedication [Skr08]. Deep [FFM01, Gor01, ABFV02, All01, BR05, CFH03, DMC^{+}09, FCM07, HMO01, HG01, HP04, Jak01, JNDH00, JDMN00, LD01, MDS03, MN01, MMN00a, MD01, PPdA05, PCB01, PGMM01, SB01a, SRA^{+}02, SGB00, SKKP05, iTMI06, WKCC06, Wes02]. deep-hooked [iTMI06]. deep-sea [ABFV02, DMC^{+}09, JDMN00, MMN00a, SKKP05]. Deep-water [FFM01, Gor01, All01, FCM07, HMO01, HP04, Jak01, LD01, MDS03, MN01, MD01, PPdA05, PCB01, PGMM01, SB01a, SRA^{+}02, SGB00, WKCC06, Wes02]. deeply [FMC^{+}09]. Deeps [BEA^{+}08, CRD06, RDH06, RWT01]. deepwater [ABK03, Cla01, EGB^{+}01, FBV06, He05, HLHF03, RBD02, VG01b]. deficient [ED07]. define [SMSK03]. Defining [VFCC06, JSR04]. Definition [CFFP07, SGS02, Cor09]. deformable [TBRL00]. deformities [TGS^{+}00]. del [Pie09]. delagoae [Gro00]. Delaroche [ACPP06]. delay [DPD^{+}03]. delay-difference [DPD^{+}03]. deliberately [GFS03]. deliberations [WD02]. delimitation [FML^{+}07]. delineating [MMH05]. deltoides [MJS00]. DeLury [GYMdL^{+}06]. demand [CND02]. Demersal [GDAM07, GLH07, SAG05, BSS06, BB01, BRO05a, BNL^{+}07, BPH^{+}06, BMB05, BMB06, CC09, CAM^{+}08, CR06, ECEG09, FBV06, GSA^{+}06, GOAARJRP09, GK01, GKF03, GBFG06, JS05, JRG08, JPVKS^{+}01, KMK09, KOO04, OKF06, RMHW03, RMMP06, SLPF08, SG04, SF03, SST^{+}06, UM02, VOS01, WLS07, vSS05]. Demographic [CY06, FDG^{+}07, SDNL09, MMB07b]. Demographical [GOAARJRP09]. Denmark [BSM07, PHM07]. dense [SG06]. Densen [Bar00]. densities [FUM02, NJE00, RR03b, WCM04]. Density [HB01, JAKM05, SSC^{+}03, BEBA03, BCM08b, CXMM04, GDAM07, JSO08, LWS04, MPQ07, MF09, NLJS01, OMR09, SB08, SW00, ZDB^{+}07, ZFJ03]. density-dependent [ZDB^{+}07, ZFJ03]. Density-versus [SSC^{+}03]. dentex [MMP^{+}02, MMP^{+}02]. dependence [FKBF05]. dependent [BGH05, CSCC07, FBJM00, JHK08, JCJ^{+}09, Lor06, SOH06, SSC^{+}03, XM00, Xia06, ZDB^{+}07, ZFJ03]. depensation [NDCK09]. depleted [FR08, KIZ09]. depletion [BPW^{+}03, CS04, GYMdL^{+}06, LC07, YPD^{+}04]. deploying [HP00]. deployment [BGM^{+}09]. DEPM [MK01, SKMT02]. deposition [ACE^{+}08]. depth [BMPK06, HCS^{+}03, JS05, LZS01, Ste08]. depurator [RMA^{+}05]. derived [Cad09, Ehr08, MK01, TCC^{+}09]. Descomposición [Pie09]. Description [HFL08, WG05]. descriptors [BF00a, NBF09]. Design [McI02e, PKT^{+}02, USKK04, AKA^{+}02, BG05, BR00, CRSM04, FP00, GSS^{+}06, GL05, HP04, KP07, LZS01, LRC^{+}08, Pri09, PLD05, QDH09, ST03, Smi03, SSK^{+}06, UMV^{+}09]. designing [GRC^{+}09]. designs [AAKT07, CTK^{+}07, JR07, YPDS05]. destructive [BP06]. detect [WVW^{+}08]. Detecting [HBKJ07, MM04]. Detection [AD03, LWW02, HZZG06, TBRL00]. determinant [Hat00]. determinants [FK03]. Determinate [GCA^{+}08]. Determination [Boo04, Has06, JDS00, Sid03, BLV06, CAJ00a, CAJ00b, MK09, MN01, MMN03, New02, PKK^{+}06, STCH09]. determine [BMBM05, Dou04, JHSG04]. determined [DSO^{+}05, WM09, YEA09, ZSP03]. Determining [GMP05, SK03a, SRH^{+}02]. deterrent [KS08]. deterrents [GGA08, SK08b]. develop [But08, CAHAF07, KPBD08]. Developing [DBK07, DSK^{+}08, RFWH09, Abl06, SST^{+}06]. Development [EMKG07, GD03, HZZG06, KM08a, MTHL06, MF03, MD09, PZH02, RGSB08, SHKB09, vHMAO^{+}06, Gue05, BSM07, CLC08, Dem04, FCS^{+}05, FK04, GSMT04, Her05a, LMMD06, MM07, MAC^{+}06, NGF06, Oja06, OM07, RE07, WRK06a]. developmental [Bas99, Bas00]. Developments [Ara09, FH03b]. deviance [WB08]. device [DACS08, HM00, KH05, PCV04, HGW^{+}08]. devices [BHM^{+}06, CTC^{+}06, CHDGM06, Dem04, DJG^{+}06, FHM09, HJ07, SBFR00, VD06]. DEVIS [JL00]. dhufish [JS05]. diadromous [OR07]. diameter [GBM09]. diamond [BM05, BMK^{+}04, CFE02, CFE03, GOF^{+}04, GSL08, HC00, He07, Her05a, Her05b, HO05, HPK06, HO06, JIGI06, MOK^{+}05, MOGG06, SLPF08, TASF09]. diamond- [BMK^{+}04, SLPF08]. Dianne [Ano01d]. Dicentrarchus [FML^{+}07, PPS05, PBLP08, PKP04]. Diel [KHMS09, KE05, VKČ^{+}09, You03, AE03, JI07, JRG08, KG04, YTK06]. diel-reproductive [YTK06]. Diet [MHWL01, RAHHGM^{+}02, APC07, MSA00, MKB01, MCP01, PMC^{+}07, SB08, ZR08, dSH01]. diets [SCP01]. difference [DPD^{+}03]. Differences [FGOF08, JJKN^{+}03, LLV^{+}08, ODR00, AD00, CRG^{+}02, Cro01, GM07, HMM^{+}09, JLHO01, KE05, LPHH09, LCFP^{+}08, RR03a, SKP03]. different [AMC^{+}08b, ACE^{+}08, BdL05, Eer04, FUM02, FRF08, GLS^{+}09, GBFG06, JIDM02, KFM08, LMR^{+}05, LMR^{+}07, OCT^{+}03, SBFR00, SF05, SM04b, TNL07, Tia09, TTP^{+}09, TCY01, VP05, VGBF02, YPDS05]. Differential [GQR03]. Differentiation [MTC06, MDGR05, MGJB09, ØA08]. differently [Ker00]. Difficulties [BAH09]. diffuse [BR01]. Digestive [McI09a]. digging [CHQ05]. Digital [HFBM08, JL00, LRG^{+}00, TMK^{+}06, TBRL00]. digitised [FAN03]. dimensional [DBB00, PBVG07, TBRL00]. Diopatra [CHQ05]. Diplodus [Arc03, DSPLR06, FSB^{+}02, FMB^{+}05, GBC^{+}03a, GBC^{+}03b, PL02, PLDS03, PA02]. Dipturus [GHDM08]. Direct [BACP09, GCC^{+}08, FDG^{+}07, NGF06]. directed [BSC^{+}08, RCD07, YPS^{+}06]. directions [Bla06, GSS^{+}06, LYKH06, Sym05a]. dirty [KC06]. disappearance [CPW^{+}06]. Discard [BWMA02, KFM08, MF07, Mau03, PSTM06, WMB08]. discarded [AKA^{+}01, DO06, GKH^{+}01, GJBY05, LF02, MMV^{+}04, PSPCVCD09, RDH05, Ste08, SRS08, VL08]. Discarding [BRO05a, CFG05, ERG07, ERG09, CXCD07, CCB02, CCBG02, FKW06, Gra02b, WLS07]. Discards [CDC03, GSH^{+}07, ABK03, AKA^{+}01, AKA^{+}02, BZRO05, ERG09, GC03a, HW00, MGTB08, MVV^{+}01, MFAM05, Pál03, RH04b, SDM04, WRGS09, YAM00, tHDC06]. discontinuity [SOKS08]. discourses [GHDS08]. discreetness [VH06]. Discrete [BLR^{+}08]. discretisation [DMBF08]. Discriminating [PPL04, Zho07]. Discrimination [PBM06, Dic04, OB02, OFOM08, WG05]. discussion [Bur02]. disequilibrium [KKJ00]. Dispersal [AM00, PAR^{+}03]. dispersion [HD00, RLGLO09]. display [WF07]. dissection [ATB^{+}02]. Dissostichus [Hor02, OFOM08, PAB05, TdlMH^{+}03, WR02, YCB^{+}01]. distance [Pal08, WW06, WRWK02]. distant [CLC08, SYS^{+}08]. distant-water [CLC08, SYS^{+}08]. distended [NGM^{+}09]. distinct [GFM02]. distinguish [DGP02, VP05, VBSK00]. distorting [Hol01, Sch01]. Distribution [CCM^{+}09, CLH^{+}08, Cur00, FB06, KA05, Mag01, NLJS01, Pou01, RLSJIA07, TYM^{+}01, ÅCST01, AMUE01, BDB^{+}03, BHH08, BSR05, CMH01, Cor01, Daw08, FAL^{+}01, GCB^{+}06, GWB07, HMB^{+}00, HG01, HE08a, HS03, JR02, Joh03, Kat05, KMB01, LR00a, LHMO09, LLV^{+}08, MvdK06, MCB^{+}07, NMHHMB^{+}00, OQE03, OA04, QDV05, RWT01, RMA^{+}05, SR06, Sho08, SHDO04, SWC^{+}05, UMV^{+}09, VKČ^{+}09, VBSC08, VMQ08, WPS^{+}07]. Distributional [ASR05]. distributions [AC05, ABS^{+}03, Dic04, HMH^{+}04, KTC00, MKGH05]. disturbance [GPI04, LBABH09, RB00]. Diurnal [BYA^{+}08, Lap02, SK08a]. diver [CTB01]. divergent [WF07]. divers [HFS02, SG06]. diverse [BHM^{+}06, HBKJ07]. diversity [HMAOAS07, MBG^{+}07, SMJS01]. Division [CP00, PVP04, RRAS03, SGS02]. Divisions [LK00, MAASMB05]. DM [Ano00e]. DNA [CHL^{+}06, CHW^{+}08, CMAS08, HSW09, IJD04, JIAD03, PMSD06, RCGC04, WCC^{+}09]. Do [Hat00, RPC^{+}06, Sal02, RMHW03]. Does [BB04, HBW00, MBGW02, PEAAdLJB07, RMCdS06]. dog [Bra04]. dogfish [DS07, KS08, MF00, MF07, RDH05, RCSFO04, SK08b]. dolomieu [WSC08]. dolphinfish [FOGKT08, MROGVR09, PMSD06]. dolphins [GGA08, Sva05]. Donald [Ano01f]. Donax [MJS00]. done [Kes00, LBKG08]. donkey [MWJ04]. Donsol [Qui07]. Don’t [LBKG08]. Doppler [DBB00]. d’Orbigny [NMHHMB^{+}00, SLA00, Vil01]. dorsal [KK09, LDN04, MF00]. Dosidicus [ARVC01, MSN01, MQVSN04, Mar06a, Mar06b, MBCMNMHH01, NMHHMB^{+}00, NMMTCV^{+}06, NNA01, NNA02, ORVMP06, RWMBHH06, SCUADJ07, SN01, TVY01, TYM^{+}01, WYR06]. double [ÖT03, SET^{+}09]. double-pass [SET^{+}09]. Doubt [Lon07]. douglasii [SH09]. Dover [Cro01]. down [BP08, KGB09, PEAAdLJB06, PEAAdLJB07]. down-regulated [KGB09]. Dr. [BNS07, SS00b]. drag [OKWM05, PAF^{+}08]. dredge [CC09, CRG^{+}02, Gas03, HLL^{+}08, MTRM05, MFAM05, WRWK02]. dredge-tracks [CRG^{+}02]. dredging [CCG^{+}02, CCA^{+}03, DNN^{+}03, MCJB02, MJB02, SGRK09, WRK06b]. dredging-induced [CCA^{+}03]. drift [CC09, RM07, STK^{+}05]. driftnet [IMOO06, YAHAB01]. drive [RPC^{+}06]. driven [CK08, FMGG08, GWR07, SW09]. drobachiensis [CH03, KCH05]. droebachiensis [GC04, PZH02, ZP05]. drought [GSS09]. drum [BR05, HGF06, JDS00, MAL02, PMAK07]. drum-net [PMAK07]. dry [MPM^{+}06]. dual [SW00]. Dubious [SM04b]. due [CS04, SBFR00]. dumeril [BAH09]. Dunn [Sho09]. duorarum [PCD03, RRAS03, RRASLB03]. duration [CXMM04, GCH^{+}07, JRG08, MAASMB05, MM05, RRAS03, SOS02, WSP06]. duration-in-beam [CXMM04]. durations [CATH04, WSC08]. during [AKA^{+}01, BSM07, Bas99, Bas00, BK09, CC01c, CLGR^{+}04, EH04b, EMN07, FAL^{+}01, FK04, GSS09, GDFH07, GLS^{+}09, HCH09, KS02, KG04, LDGA09, LR00b, LWW02, MM07, MSB^{+}08, MNML00, MAND05, MC06, NJE00, OAKM07, ORVMP06, PPH01, Sva05, TNL07, Tia09, TSOG01, iTM04, TKFJ09, WW06, WH08, XJ05, YEA09]. Dutch [tHDC06]. Dutta [Ano01c]. DUV [CC09]. dwelling [NNE^{+}05]. Dynamic [KLS^{+}07, LDK^{+}07, HKG^{+}08, Jen00, NAD^{+}07, PCS05a, QDH09, SE06b]. Dynamics [TVY01, ACMMCEGM02, ASA04, ABAA^{+}07, AAHE04, BCSA08, BLR^{+}08, CCM^{+}09, COHP02, CCD01, DCM^{+}05, DG08, EL09, FCAD^{+}08, Ger09, GFvdV06, GWR07, HKG^{+}08, KLU01, LNG04, ML04a, ML04b, NDCK09, OAKM07, PdAO02, QPR08, RPP04, RSST07, SB01b, SWA08, TMP^{+}06, TDK^{+}05, WGR08, WLL08, WIJB08, Cur00, McI02c].

ear [MWJ04]. Early [GBL^{+}05, GEH^{+}07, KIT^{+}04, LD08, PRM04, YM00, Bas99, Bas00, HP09a, IWW03, Lan04, Oja06, RPC^{+}06, SJ05, How06]. earth [KS08, SK08b]. Earthscan [McI08]. East [NJL^{+}00, NWML00, Bal09, BEBA03, Bur02, CFG05, CCD01, EBA^{+}03, GM00, IJD04, KIT^{+}04, KI04, KI05, LMRQNM^{+}05, MBG^{+}07, MMN03, MCS05, OJ06, OL07, PGS^{+}08, PH09, SLA00, SN01, Ste03c, SOKS08, vHAN^{+}09, CC01c, CLH^{+}08, GRB01, PMAK07, RRJ08, RRJ09, WLL08, WBHG01, YST^{+}02, YFT^{+}02, PLDS03, SGL08]. Eastern [Gue05, BAEF06, BQ04, BK02, BFM08, BGW06, CS04, CSDP02, CdCBdR03, CTC^{+}06, DMA07, DGP02, DA07, DOQ08, DKGC00, EMKG07, EL09, GSS09, GBM09, GCMJ08, HH08, Ism03, IYI07, IS07, JSB04, JCJ^{+}09, LCNJ06, LLJC09, MBO^{+}07, MMV^{+}04, MCL04, MROGVR09, MCB^{+}07, MSA00, MAL^{+}01, Mos00, MUO02, NHSD08, ORVMP06, PMSD06, PSTM06, RB04, SGS09, SRH^{+}02, SST^{+}08, SH09, SFCK03, SWC^{+}05, TÁTWD09, TMP^{+}06, TOBSJH06, VL09, VBSC08, WMdSB09, WMdS09, Wes02, Zhe03, vSS05, vSSK07, vSS09, CCM^{+}09, GB07, KR00, PA02, RB09, RSST07, RWMBHH06, TSOG01, CSM^{+}09]. easy [JJWO09]. ebriensis [Eze02]. Eburon [McI05c]. EC [AKA^{+}01]. echo [JL00, Kor04, MC00a, RSP^{+}00]. echosounder [KLJ06, Kru04]. echosounding [Sch09]. Eco [She09, Ste06]. Eco-certification [She09]. eco-metrics [Ste06]. Ecología [Pie09]. Ecological [Mil03b, AHS^{+}06, But03, CQJX08, HMO01, TCS05, ZG08]. ecologically [FCS^{+}05]. Ecology [Ell00, Lon02a, Pie09, Win02, CJ00, FGJD08, FMB^{+}05, Gro00, Mar07b, MMLT05, NE07, Ste07, Wah03]. Ecomorphological [VE03]. Economic [CLPP05, EH02, FMGG08, Gus02, LFA08, MT08, Mil03b, NS08, AM03, BCLP00, BRRL09, CAM^{+}08, CL04, Daw08, Eer04, GC03b, GLCC09, Kro04, MKGH05, RE07, SW09, TCS05, TTP^{+}09]. economical [IV08]. economically [CMGB01]. Economics [Ore03, Sym08a, DDP^{+}08, Gro00]. economy [MS05b]. Ecosystem [CHDGM06, Koe08, McI02f, SS00a, YLCS^{+}09, BP01b, BRRL09, CXQ^{+}09, Ein09, FSPWG08, GDAM07, Gav09, Gri03, KMF08, KIZ09, MMLT05, MM06c, OJ06, Sei07, SAA^{+}03, SWT09, SAC^{+}02, WM09, ZKG^{+}09, McI03b, Coo01]. Ecosystem-based [Koe08, FSPWG08, OJ06, ZKG^{+}09]. ecosystem-level [MM06c]. ecosystemic [GEYPC^{+}09]. Ecosystems [Cad00b, McI06a, Ste05, GG06, HMR^{+}08, Lon03b, VG01a, WIS03, WM04, McI01, McI02e]. ecotourism [CTEARD07, RDEACT07, RE07]. Ed [Ano00c, Ano00d, Ano01f, How03a, Kes00, Lon02a, McI02d, McI05d, McI06b, McI09d, Mil02a, Mil06, Sol05, Ste03a, Sut02, Van01, Win02, McI07b]. edge [BH01a, RWT01, Rei01, UL01, WPS^{+}07, Sol05]. Edible [EBA^{+}03, ASM^{+}09, BEBA03, GTB04, PK03b, UHN07]. Edinburgh [Ano01f]. Editions [Gue05, Ano00d]. Editor [KW03b, Lon03a, Sho09, AOLLP^{+}07, Pra03]. Editorial [Ano04b, McI03c, McI07c, Ano03m, Ano02h, Ano02j, Ano02k, Ano03k, Ano03i, Ano03j, Ano03f, Ano03g, Ano03h, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano04a, Ano04c, Ano04d, Ano04e, Ano04f, Ano04j, Ano04k, Ano04l, Ano04m, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06r, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q]. Eds [Ano00b, Ano00e, Ano01c, Ano01d, Bar00, Coo00a, Cur00, Ell00, Gue05, How03b, Lai00, Loc02, McI02c, McI03a, McI03e, McI03b, McI05b, McI05c, McI07b, McI08, McI09a, McI09b, McI09c, Mil03b, Mil08, Ore03, She02, Ste05, Sym07, Sym08a, Sym08c, Sym08b]. edulis [WLL08]. Edwards [He05, PGMM01]. edwardsii [BM09, HSGF06, LPHH09, PHGT06, ZFJ03]. eel [CMC09, HB01, NM06, OMF^{+}03, PRM04, PHM07, PLD05, WT00]. eels [GT04]. EEZ [CHN08, Orl03]. Effect [AAKT07, HHTE05, He06, Her05a, Her05b, KAHR02, MWB^{+}08, ÖW02, PGF07, TÖT04, TOBSJH06, WSP06, WPOB09, YKO09, AKA06, CTK^{+}07, CRG^{+}02, CS07a, GOF^{+}04, GCJ07, HJ03, HPK06, HPK07, HWM02, Hor00, JPYW08, KNSSKF09, KSS05, KOO04, MH02, MS05a, MvDM02, OGK^{+}08, Ogl09, PG06, RL03, RMA^{+}05, RT07, SB08, SMS02, SGB00, SBGG^{+}05, SLJ05, TH06, VOH02, VL09, WSM02, vM03, vSM02c, vSM02a, vSM02b, vS04, vSS05]. effective [AA03, AGB07, EH04a, LPG^{+}09, SB03]. Effectiveness [APM^{+}09, BSWL06a, HJ07, RH06, GAOMC06, MWJ04]. Effects [ADP^{+}08, BGM^{+}09, BA06, BCSP03, BCMM09, CC01a, Cur00, DGS^{+}05, EJA00, GBM09, GTB04, HH08, HBBG06, HWW^{+}09, Hur07, JLC^{+}03, JRG08, KG06, MCJB02, MAC^{+}06, NS04, NJJ06, ST00, TNFF03, THN^{+}04, WIJB08, ZG08, vZRMvD02, AOMK^{+}04, BGB06, CP06, CFH03, CSR06, CXQ^{+}09, CCA^{+}03, CAFT00, CdMLGC09, CAH^{+}03, CCD01, CTC^{+}06, DNN^{+}03, DMA07, DDP^{+}08, EPA00, FF01, GCB^{+}06, GCH^{+}07, Gri06, GPI04, HMM00b, Hen09, JJJ03, KM03, KHMS09, Kop06, KF04, LH04, MJB02, MSVF05, Oja06, OKPR04, PMSC03, PBW^{+}09, Qui07, RPP04, RR03b, SJ05, TS09, TNL07, iTM04, VN00, WM07, WEK^{+}09, WMB08, WD02, ZR08, ZCTA^{+}08, McI03a]. Efficacy [JWR05, KA06, WVW^{+}08, CAM^{+}03b, ERG09, RCD07]. Efficiency [HMM^{+}01, APP^{+}07, BMSJ05, CG04, CATH04, CSSO03, FKW06, FK03, GHM04, JK07, LIW04, MTRM05, MT08, MS02, PFHC08, Pri09, RB06, SET^{+}09, Som03, TEG03, TPC05, VMSR08, WSM02, vSS05]. Efficient [KHMG06]. efficiently [KJ08]. Effort [LLJ^{+}08, AFG^{+}05, BQ04, BVW04, CNH09, CXMM04, CP08b, Chi00, CdMLGC09, Daw08, DPD^{+}03, DKH08, FH03a, GGVLSR03, HWB06, Kar04, LRC^{+}08, LP06, MBT08, ML04a, MP04, MDS03, OAKM07, Par04, PLPN09, PA09, PWTP00, SMD04, She03, SYS^{+}08, TB04, TCC^{+}09, Xia04a, Xia04b, YPM^{+}06, vOPvDM02]. Egg [PCM07, CRC^{+}07, KK04, MPE04, MK01, MSB07, SKMT02, VN00]. eggs [ÅCST01, ABL02, AMUE01, CMH01, KGB09, KMB01, MAC^{+}06, OQE03]. elapsed [Cad03]. elasmobranch [BSMF^{+}09, MF07, SK08b, ZG08]. Elasmobranchii [CGP05]. elasmobranchs [LBMCE00]. elastomer [BLB07, WMS04, ZW09]. electralume [HHTE05]. electric [ESB09]. electrical [PDV05a, PDV05b]. electrofishing [KA06, NJE00, SET^{+}09, WN03]. electronic [vT04]. Eledone [RMF^{+}06, SMD04]. eleginoides [Hor02, OFOM08, PAB05, TdlMH^{+}03, WR02, YCB^{+}01]. elemental [CCH^{+}00]. elements [MD01]. elephas [BMF^{+}10, GQR03, GG04, TSP^{+}04]. eliminates [FHJL08]. eliminating [BCMM09]. Elliott [Loc02]. elliptical [TLD06]. Elner [Ore03]. elongation [AOMK^{+}04]. Elsevier [Mil06, Ste05]. embayment [WMTJ07]. emerging [BLKR09]. Emmelichthyidae [EL09]. Emmelichthys [EL09]. emperor [PGK03, WMD07]. emphasis [HD00, MMLT05]. Empirical [PB03, CWS^{+}05, CJ00, GDDJ09, CSCC07]. employing [KA06]. employment [HF07]. encircling [Rue07]. encrasicolus [ACE^{+}08, BGP^{+}04, BPRM00, Kno07, KM08c, LDGA09]. end [CFH03, GOF^{+}04, GF06, HO05, HPK06, HFHO07, HPK07, JSO08, KOO04, OMW^{+}03, OKWM05, OKF06, OGK^{+}08, RBD02]. endangered [SND^{+}08]. endemic [GCN02, LC07, MN03, dGvZM^{+}06]. endogenous [GR02, Lon09]. endpoints [FMC^{+}09]. ends [CFE02, CFE03, Her05a, Her05b, HO06]. energetics [ACMMCEGM02]. energy [OGSG07, Pri09]. Enfield [McI09a, McI09b, McI09d]. enforcement [MWJ04]. Engineering [LMMD06]. England [EF01, The07, AKA^{+}01, BEBA03, BEA^{+}08, BCLM08, EBA^{+}03, PKP04, VG01b, WN03]. English [RDH05, CFG05, CRD06, CCBG02, DPS^{+}01, ERG07, FML^{+}07, RPFR06, ULDG02]. engrauliformis [FZHL09]. Engraulis [ACE^{+}08, AM00, BGP^{+}04, BPRM00, CC03b, CSCC07, CBC02, CRC^{+}07, FAW04, GESY^{+}07, JHK08, Jun08, Kno07, KM08c, LDGA09, PCM07, WLY06]. enhance [SSML02]. Enhancement [MFB08, McI05b, She02, BBRS06, BBLL06, BLB07, GAC^{+}06, GFMO06, HK06b, HCHO^{+}03, KK06, LYKH06, Lor06, NW03, Uki06, WZDY06, War06, YLD^{+}05]. enrichment [Kno07]. ENSO [BBV07, WYR06]. entomelas [RFWH09]. entrainment [JHSG04]. entrance [AKA06]. entry [Sal02]. envelopment [PH04a, TPC05]. environment [BM03b, BM03c, Gue05, HMB^{+}00, HPD09, PPdC01a, SAK05, Sym05b]. Environmental [BQN05, DMA07, GPI04, JMGL04, McI03d, Mil02a, BCSA08, But03, BPD^{+}04, CFG05, CPW^{+}06, CG04, CC01a, DUR05, DKKS07, GMV00, GMV01, HP09a, HT05, LWW02, MOK^{+}05, OGPDOM08, PHM07, RSST07, VYMGM^{+}09, Vel03, VGBF02, VMQ08, ZPR01, ZFJ03]. environments [AR01, Rod01, Ano00e]. eperlanus [JAKM05, PMT06]. epidemic [MG03a, MG03b]. epinepheline [BMBM05]. Epinephelus [BD04, ED07, GHCPPLC^{+}05, GAFA05a, HS03, LCFP^{+}08]. episodic [WHWM07]. EPO [TSOG01]. equations [BVW04]. equatorial [HHTE05, LDN04, PMC^{+}07, SDNL09]. eques [Mag01]. Equilibrium [CHSA02, CA03]. equula [YFT^{+}02]. equulus [IMSS06, IOS07]. Eritrea [TMN07]. Erlend [She02]. Ernst [LOT07a]. erratic [KW03a]. Erratum [Bas00, CAJ00a, FCL^{+}05a, GBC^{+}03b, GMV01, GMR02, KPL^{+}09, KI05, NNA02, vSM02c]. error [Dic04, DFD02, GDDJ09, LB08, OA04, Pal08]. Errors [PC03]. erythropterus [NCW00b]. Escape [MHF02, SLJ05, AAKT07, GLS^{+}09, HBBG06, JPYW08, MS05a, MTU^{+}04, MB00, ÖW02, ROS04, SMS02]. escaped [DO06]. Escapement [IJ06, LRC^{+}08]. escaping [GFSM07]. escolar [MD02]. esculentus [YLD^{+}05]. Esox [AKC^{+}09, LCD^{+}02, LPL^{+}06, LPG^{+}09]. Essential [BHB^{+}04]. established [Loh08]. estimate [BPRM00, BBJG04, CF07, CCBG02, GHM04, HD00, KKR^{+}05, MVAMB07, MS05a, MM06c, NHT07, PCHM07, TÁTWD09, YPD^{+}04, YPM^{+}06, ZP05, ZLC09]. estimated [DSPLR06, EPA00, MF07, MOGG06, OIK^{+}06, PF00, PRdPG08]. Estimates [HCM05, ABK03, BACP09, BG05, FUM02, GM07, HBJ03, JLHO01, Joh03, JI07, KK09, KM02, KTC00, KLJ06, LRBH07, LLJC09, MM06a, MK01, MF09, NCW00b, OKF06, ODdGV02, RPP04, RHF^{+}08, RVS02, Tal09, Zhe03, vS04, vSSK07]. Estimating [AA03, AP05, BP06, Din05, Ehr08, GCMJ08, HMM00a, KK04, LWS04, LIW04, LP06, MC06, MS02, PMSC03, TPM03, VL08, VGBF02, YCB^{+}01, BEBA03, BVW04, BMSJ05, FH03b, HdS07, HP04, HH06, JWR05, Kop06, LRC^{+}08, LB08, MB08, OMR09, RMRB06, RYSN08, SM04a, WN03, WH08]. Estimation [APP^{+}07, BC04, BF00b, CWL00, CHQ05, HFS02, Ker00, Kno07, MBR06, PH04a, WL06b, WRWK02, Wil09, YLYL09, AKA^{+}02, AFSF05, AFVSR^{+}09, CP00, CXCD07, CCB02, CBC02, CRC^{+}07, Dia00, DDP^{+}06c, FL05, Fab06, FBJM00, GCD^{+}09, GC04, GCC^{+}08, HPD09, PL02, PLDS03, PMG^{+}06, PAB05, RM05, SG01b, UMK01, UMV^{+}09]. estimations [BZRO05, MAdAMASCM09]. estimator [IHF08, PCAS04, WHWM07]. estimators [Cad09, CJ00, DFD02, FHI07, HPD08]. Estonian [LOT07a]. estuaries [GK03, Laz08, RM05, SMM07, Loc02]. Estuarine [GSS09, BMK^{+}04, BMSJ05, Gra02b, GKH03, JRG08, MC06, OSG08, Rue07, SR00, SBGG^{+}05, Wes02]. Estuary [SBGG^{+}05, BPCR^{+}05, CWJS09, EAT05, EKU02, GC03a, JMGL04, JDS00, MAL02, MM06b, SBD^{+}05, SBD^{+}06, TGS^{+}00, VFCC06, AAHE04, JWR05]. Ethiopia [dGvZM^{+}06]. Ethmalosa [AAHE04]. Ethnoecology [SB05]. Etienne [Mil02b]. Etmopterus [CE07, Jak01]. Etolikon [LS00]. Euphausia [HD00]. euphausiids [HD00]. EUR [McI05c, McI07a, Ste05, Sym08c, Win00]. Europe [McI02d, BTC08a, OM07, Sym05a]. European [ASR05, ACE^{+}08, AMUE01, AC05, AFG^{+}05, BGP^{+}04, BLV06, BSR05, CMC09, EGB^{+}06, FKBF05, GGVLSR03, GSH^{+}07, GQR03, KA05, KAO04, LDGA09, LCTB^{+}03, MM07, MNBM05, OMF^{+}03, OQE03, PRdPG07, PKŘ^{+}09, ŠV04, SMS^{+}06, Sym09, TSP^{+}04, UL01, VKČ^{+}09, WT00]. eutrophic [OMR09]. eutrophication [AM03]. evacuation [AB08a, TH03]. evaluate [PP01]. evaluated [SCUADJ07]. Evaluating [BRRL09, CP08b, FHI07, HPD08, HPD09, IWBJ08, LB08, Rue07, TCC^{+}09, VMSR08, WMdSB09, BLR^{+}08, Cad00a, DDP^{+}06a, GOAARJRP09, MWJ04, Zho07]. Evaluation [AKA^{+}01, FHM09, GCH^{+}07, GSL08, How03a, IHF08, JJWO09, KFM08, MK02, McI02e, OR07, PWHM06, PCV04, PDH08, VD06, WSPY05, BDIP08, BP06, BG04, BR00, DACS08, DDP^{+}08, FMC^{+}09, FP00, FFM08, GAOMC06, HBM^{+}00, HZZG06, KCRT03, Løk03, MLP^{+}08, Mes03, Nee08, Pal08, TP06, Tre01, TTP^{+}09, YMT^{+}06, Mil03b]. evaluations [DBK07]. event [Mar06a, Mar06b]. events [SMM07]. Evermann [BSMFH07]. Evidence [ACG^{+}01, DHM08, FFM08, FBV06, JTdH06, LBABH09, MBO^{+}07, MTC06, NGF06, SOKS08, vHAN^{+}09, CFH03, CS04, CSCC07]. evidences [DMC^{+}09]. evolution [CLGR^{+}04, CCCC04]. exact [TdlMH^{+}03]. examination [TP00]. examine [CC09]. Examining [CRSM04, HMM00b, SJ05]. example [AMC^{+}08a, Cla01, HH01, Joh03, JCJ^{+}09, LD06, VD04a, WRWK02, dGvZM^{+}06]. Examples [GM07, MPQ07]. exchange [AFG^{+}05]. excluder [BHM^{+}06]. exercise [WSC08]. existing [SdPFAL09]. existing- [SdPFAL09]. exit [GSL08, GLH07]. Exmouth [YLD^{+}05]. expansion [BP08]. Experience [eJSK09, SST^{+}08, AFG^{+}05]. Experiences [OJ06, WFJB09]. experiment [ARLB07, BPW^{+}03, GFSM07, LK02, PBW^{+}09, UL01, WLDH08]. Experimental [RBH^{+}08, RGGB07, GSH^{+}07, GBG^{+}03, PLD05, SdPFAL09, Wah03]. experimental-otterboard [SdPFAL09]. experimentally [CS07b]. experiments [Dou08, Ehr08, LK02, Lor06, MH01, MH02, MS05a, PEB03, Pol02, PDV05a, PSPCVCD09, VU08]. expert [GB07]. explained [ZFJ03]. explicit [But03, CP08b, JL00, MMZ09]. Exploitation [RCA08, Ste05, CWL00, CAFT00, GBL^{+}05, Gue05, MBK02, PH01, PMWC03, SG04, UM02]. Exploited [SWT09, AAHE04, CP08b, DCM^{+}05, DMC^{+}09, ESC08, GWB07, GBG^{+}03, Has06, MN03, MMLR08, PPdA05, PCB01, RBHP03, RCG02, VH06, WM04]. exploratory [VG01b]. Exploring [Daw08]. exposure [GCH^{+}07, HBBG06, WSC08]. extended [YLYL09]. extension [KMK09, AM00]. extent [MLH^{+}05, SK03a]. external [Oja06]. extinct [CMGB01]. Exxon [RR03b]. eye [BC02].

F [Coo00a]. fabricii [Bjø01, Jak01]. Fabricius [CS07a, GQR03, GG04]. face [Pun03]. facilitate [PBLP08]. facing [PTE09]. factor [KAHR02, Lon02b, SKF08]. Factors [DDP^{+}06c, GWB07, JI07, Lan04, QGBS04, TPC05, WK08, CFG05, Chi00, Daw08, GCMJ08, JMGL04, KNSSKF09, LAL07, MOK^{+}05, MUO02, NGF07, RPC^{+}06, SSB03, SAK05, SMB^{+}08, VYMGM^{+}09, VLD^{+}00]. FAD [Dem04]. FADs [Dem04]. failure [RB09]. Fairbanks [Coo00a, Coo01, McI01, McI02c, McI06b, Ore03]. falciformis [JCLL08]. Falkland [PPH01, AM02, Lap02, LB05, MCP01]. Falklands [PAB05]. fall [ABAA^{+}07, BBV07, Lot07b]. falsifiable [Cor02]. familiarity [WR05]. Family [AHAAD01, GAFA04a]. fangi [HLH08]. FAO [McI03b, McI05d]. far [Gri03]. Farfantepenaeus [LMASHV^{+}03, PCD03, RRAS03, RRASLB03, VD06]. farming [NW03]. Farn [CRD06, BEA^{+}08]. Farne [RDH06]. Faroe [MGR^{+}02, SCM09]. fast [PRdPG07, RMHW03]. fat [HHC06]. Fate [SRS08, iTMI06, CPR07, HW00, LF02, NGM^{+}09]. fatigue [CAH^{+}03]. fauna [BPW^{+}03, EMN07, PMC^{+}07, RBH^{+}08]. feasibility [SRA^{+}02, SLJ07]. features [Ber07, NLJS01, WPS^{+}07]. February [Ano00h, Ano01l, Ano02y, Ano03z, Ano04s, Ano05w, Ano06z, Ano07r]. Fecundity [AFVSR^{+}09, GRB01, JHK08, AB09b, GSMT04, GCA^{+}08, JWR05, JCJ^{+}09, KGB09, KJ08, LCTB^{+}03, MFB08, MM06b, YFT^{+}02]. feed [IYI07]. Feeding [CMC^{+}06, DG02, FMB^{+}05, MSA00, ABL02, BSR05, JHGS04, Lap02, Loh08, Mar06a, SOH06, Tay07, VKD03, McI09a]. Female [PIdlHGV05, CSCC07, Cro01, GQR03, MCB^{+}07, YFT^{+}02]. females [XK04]. fence [BMSJ05]. Fernando [GCS^{+}06]. Festschrift [Sym08a]. FF [Ano00d]. Fidelity [WGGN06, BLC08, HKG^{+}08, Loh08]. Field [KS08, AB08b, CA07, FBL^{+}07, HFS02, ODR00, SFCK03, vT04]. Fijian [KVZ05]. filter [ABL02, Tay07]. filter-feeding [Tay07]. filtering [LNS08]. fimbria [BM00, CS07b, GJSW04]. fimbriata [AAHE04]. fin [KK09, Lap02, LDN04]. Fine [BLC08, ASD09, DHM08]. Fine-scale [BLC08]. finfish [KK06, LG02, OSG08, WRGS09]. fingerprints [CCH^{+}00]. finite [SHBÅ07, STH03]. Finland [HH08, HSSR06, Jok02]. finned [PdAO02]. Finnish [JRML00, JAKM05]. fins [HSW09]. Fiorentino [Bet04]. First [Ano07w, PRdPG07, HFS02, LOT07a, LMRQNM^{+}05, RMN08, vSSK07, GSS^{+}06]. Firth [LF02, YPS^{+}06]. Fish [Ano00b, Bas00, BHB^{+}04, BMC07, Ber07, BMB06, BCM08a, CAJ00a, Cor02, Cur00, DK05, FCL^{+}05a, GBC^{+}03b, GMV01, GMR02, KVS04, KPL^{+}09, KI05, McI07b, Mil03b, MMN00a, MG03a, NNA02, PF00, PBDP00, SD01, TKFJ09, vSM02c, APM^{+}09, ABL02, AFJC08, AMAOMM04, AKA^{+}01, AKA^{+}02, AAP^{+}09, ASIW05, Ano02l, AFG^{+}05, AAT08, ABS^{+}03, ABRB00, AC00, BDB^{+}03, BCC09a, BM00, BNS00, BBK08, BF00b, BRRL09, CCH^{+}00, CP06, CFH03, CNH09, Cer00, CSR06, CMGB01, CXMM04, CBGP06, CdML^{+}07, CA07, Dah00, DPS^{+}01, DO06, DVAL^{+}09, DBB00, Dem04, Din05, DJG^{+}06, Dou04, Dun06, EAT05, EMN07, ESB09, FL05, Fab06, FCC^{+}09, FÖB08, FKBF05, FBS02, FHJL08, GFSM07, GSA^{+}06, GDAM07, GG06, GCH^{+}07, Gob00, GCD^{+}09, Gor01, GF06, GBM09, GLK00, GRC^{+}09, GSL08, GBFG06, GCL03, HP02, HG01]. fish [HFS02, HGW^{+}08, HPK07, HMR^{+}08, HM00, IJ06, IPH05, JMGL04, Jen02, JHSG04, JRG08, JSO08, JCJ^{+}09, KAN00, KJEF03, Kat06, KM03, KW03a, KW05, KA06, KP07, KM08b, KFO^{+}04, KLJ06, KHMS09, KMK09, Kru04, KIZ09, KE05, KHM^{+}09, KAHR02, LM09, LZC^{+}08, Lon02b, LMR^{+}05, LMR^{+}07, MvdK06, MVAMB07, MCL04, MWB^{+}08, MPQ07, MSVF05, MK02, MM06a, MMB07b, Mat07, MC05, MSSJ05, MM05, MAC^{+}06, MÁ08, Mes03, MC00a, MN01, MAL^{+}01, MD01, MvDM02, MC00b, NHT07, NBA^{+}09, OSG08, OKPR04, OMR09, Orl03, OLN^{+}02, Ove02, OR07, ØT02, PMG^{+}06, PGS^{+}08, PR03, PH04a, PSvDPM01, PGK03, PWB05, PL00, RL03, RKEJR06, RMRB06, RMMP06, RSM05, RVS02, ROS04, SB01a, SGVM02, SCLM06, SND^{+}08, ST00, SLV^{+}08, SMS02, SB05, SHDO04, SBGG^{+}05, SMM07, SSML02, SFCK03]. fish [SMB^{+}08, SLJ05, STH03, SBD^{+}05, SBD^{+}06, Tre01, USKK04, VOS01, VMSR08, VE03, WG03, WDC02, WKCC06, Wes02, WSS06, WFJB09, XJ05, YEA07, YFL08, YKO09, You03, YSS^{+}09, HM00, Ano01c, Bar00, McI09b]. fish-kill [SMM07]. FISH-LIFT [HM00]. fished [IWW03, Mey07, WW06]. Fisher [GB07, MPQ07, Tri05, Sym08c]. Fisheries [APS07, Ano08e, Bar00, Cad00a, Coo01, Cur00, FW06, Gav09, Kes00, Kir01, LOT07a, Lon02a, McI02a, McI02f, McI03b, McI05c, Mil02a, Mil03b, San00, Sol02, Sym06, Sym08a, Sym08c, Sym09, Van01, VG01a, Abl06, AP05, ASIW05, Ano01f, Ano02c, Ano02d, Ano02e, Ano07y, Ara09, AIC08, AM03, ABF08, AHS^{+}06, BPCR^{+}05, BSM07, BA06, BBLL06, BHB^{+}04, BP08, BSMFH07, Bla06, BM08, BGH05, BNL^{+}08, BP01b, BRRL09, Cad02, CS08b, CMM^{+}07, CJ00, CLC08, CQJX08, CL04, CR06, Cla01, CdCBdR03, CK08, CHBB04, DMA07, Daw08, DSK^{+}08, EMKG07, Ein09, EGB^{+}06, FCS^{+}05, FK03, FSPWG08, FHG03, GHM04, GAC^{+}06, GSS^{+}06, GSA^{+}06, GRFE06, GFMO06, GKUO07, GJ09, GG06, GSH^{+}07, Gor01, Gra03, GC03b, GK03, Gri03, Gro00, GLCC09, GMP05, Gus02, HP02, HCM05]. fisheries [HF07, HSSR06, HMR^{+}08, HT05, HIRM06, HS00, IFH^{+}09, JL00, JCW05, JTdH06, JSR04, KAO04, KC06, KJS^{+}06, KMF08, KJ09, Kro04, KIZ09, KVZ05, Lab05, LDK^{+}07, LAL07, Lev05, LPHH09, LFA08, Lon09, Lor06, LSW06, LD06, MKGH05, Mai09, MLH^{+}05, MT08, MPQ07, MTA^{+}02, MS05b, MN03, MS03, MDS03, MTGH00, MB07, McI02d, MWF06, MHK^{+}06, MR05, Mil03a, MAG00, MF03, MDJR05, Moo03, MMN00a, Mos00, MUO02, MPG^{+}08, MJS00, eJSK09, NGF07, NJL^{+}00, NWML00, NHSD08, NM06, OM07, OAKM07, Pál03, PKT^{+}02, PBLP08, PAP06, PA05, PEAAdLJB06, PSvDPM01, PPdC01a, PKP04, PCB01, PP01, PHM07, PSTM06, QDH09, QPR08, RCA08, Rei01, RCA^{+}07, RV03, Rod01, RGSB08, Rue07, RSMZ06, RT07, SLPF08, SCSC07, SB03, Sca03, Sei07, SW09, She09, She03, SAC^{+}02, SRH^{+}02]. fisheries [SA03, SST^{+}08, SBGG^{+}05, SH03, Ste06, SMM07, SME02, SMSK03, SST^{+}06, SSG06, SHBÅ07, SBR^{+}08, Sym05a, SP09, TCS05, TP06, TPM03, TPC05, Tri05, TMN07, TFA03, TDK^{+}05, TSTK06, ULDG02, VD04b, VSK03, WL06b, WSS02, WM07, WC07, WR02, WD02, XPMQ04, Xia06, YPM^{+}06, YJV^{+}03, ZCdG06, ZCTA^{+}08, ZR04, ZDB^{+}08, ZKG^{+}09, NWML00, McI02f, McI09c, Win00, Win02, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano05c, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano08a, Ano08b, Ano08c, Ano08d, Ano09e, Ano09f, Ano09g, Ano09d, Sym07, Sym08b]. fisheries-dependent [BGH05]. Fishermen [CR06]. fishers [BHB^{+}04, Daw08, HFBM08, Mey07, PL04, WLDH08]. Fishery [Coo00b, GMA02, Gri06, GLH07, McI05d, PLD05, RWMBHH06, ST03, TJ05, VSS07, XBS^{+}07, ADP^{+}08, ABK03, APS07, ABD^{+}09, AAP^{+}09, Ara09, AIC08, BB04, BSS06, Bal09, BC02, BCSP03, BTC09, BWMA02, BS06, BSC^{+}08, BCMM09, BVW04, BSMF^{+}09, BBV07, BRO05a, BM05, BI05, Bra04, BHM^{+}06, BH01b, BKD^{+}02, BMK^{+}04, BD04, BMSJ05, CDC03, CWS^{+}05, CHN08, CMJ05, Cam04, CS09, CFE03, CF04, CFFP07, CA03, CAM^{+}03a, CFG05, CRD06, CGP05, CAM^{+}08, CAHAF07, CXQ^{+}09, CCCC04, CTC^{+}06, CHDGM06, CHSA02, Dem04, DR01, DBAÖ06, DTSR07, ED07, EGB^{+}01, FKW06, FRGF07, FÖB08, FM04, FCL^{+}05a, FCML05, FCL^{+}05b, FCM07, FRF08, FHM09, FR08, GHS04, GC03a, GGVLSR03, GSS09, GFL03, GOAARJRP09, GNC07, GC04, GDFH07, GAFA05b, GB07, Gra02b, GJBY05, Gri03, GFvdV06]. fishery [GL05, GMS06, GWR07, Güç08, GTB04, GD03, HLL^{+}08, HdS07, HC07, HWB06, HPD08, HPD09, HJ07, HMM00b, HMM00a, HPM04, HMAOAS07, HS03, HTF^{+}09, HH01, IMOO06, IHF08, KS08, KCH05, Kar04, KSS05, KG06, KTL06, KFM08, KPL^{+}07, KPL^{+}09, KHMG06, KF04, LF02, LB05, LR08, LHMO09, LBKG08, LZCR08, Løk03, Lon07, LMASMASZ02, LD01, LP06, LWM07, MM07, MBO^{+}07, MLJT09, MGTB08, MVAMB07, MH01, MHF02, MS05a, MLP^{+}08, MMS09, MAdAMASCM09, ML04b, MBCMNMHH01, MFAM05, MAG^{+}09, MGNB05, MCP01, MD09, NS08, Nar06, NPC^{+}08, OCT^{+}03, OL07, ORVMP06, OA04, PEB03, Pál05, PFHC08, PPS05, PEAAdLJB07, PW05, PPdA05, PZH02, PMSC03, Pol02, PCV04, PGF07, PSPCVCD09, PBMB03, PL06, PCS05a, PVMdL05, PWTP00, PWP02, PVP04, PAP04, RB09, RPP04, RdFBS01, RMF^{+}06]. fishery [RDH05, RDH06, RCD07, RH06, RGGB07, RB00, RPFR06, SMD04, SSB03, SCW04, SALnM07, SGS02, SB01b, SHBC02, SGL08, Smi03, SGB00, SLJ07, SF03, SH09, SMJS01, SYS^{+}08, SJ05, SRS08, SABW09, TCPC06, TB04, iTM04, iTMI06, TSTK08, VL08, VAYD04, Vas03, VL09, VD06, VLD^{+}00, VVH04, Vil01, VMPP07, WLS07, WK01, WKM02, WIKM05, WYR06, WGR08, WMdSB09, WMdS09, WEK^{+}09, WMB08, WGDP06, WR02, WVW^{+}08, WBHG01, YAM00, YAHAB01, YPDS05, YKM06, YKO09, YPS^{+}06, You03, ZC02, ZG08, dGvZM^{+}06, dSBAC06, vOPvDM02, vZRMvD02, DPD^{+}03, DDP^{+}06a, DDP^{+}06b, DDP^{+}06c, VD04a, Coo00a]. Fishery-independent [ST03, GDFH07, PWP02, SHBC02]. Fishes [How03b, How06, McI01, All01, AB08b, AB08a, BMBM05, BB01, BMB05, BMB06, CC09, CdMLGC09, FFM08, FBV06, GDFH07, HK06a, HC09, IV08, JLC^{+}03, LP00, MKGH05, MDGR05, MMM^{+}03, PMC^{+}07, PKŘ^{+}08, SK03a, SET^{+}09, SGP02, SR00, SWT09, SG04, SW00, SK03b, SKF08, SG01b, TRL06, VKČ^{+}09, WM04, McI09a, Mil02b, Loc02]. Fisheye [HGW^{+}08]. Fishing [Ano01d, Bat04, BTC08a, Cur00, FF01, Kro04, McI02d, McI03a, OL07, PA09, She02, She03, Sym05b, VSK05, VMQ08, WRK06a, WRK06b, Win02, ZG08, AA03, ABCB^{+}07, AMAOMM04, AOMK^{+}04, AS04, AAH06, AGB07, AAA^{+}06, BDB^{+}03, BCC09b, BM03a, BLH04, Cad00b, CLPP05, CP06, CS09, CS04, COLB07, CSDP02, CRG^{+}02, CdMLGC09, CLGR^{+}04, CCBG02, DKH08, EH02, EH04a, ECEG09, ESB09, FH03b, FBV06, GCB^{+}06, GHCPPLC^{+}05, GPI04, HH08, HMO01, HP09a, HJ03, HMM00a, He03, HMM^{+}01, HCH09, HWW^{+}09, HW00, HWM02, HLHF03, JR02, JJJ03, KW03a, KVZ05, LK02, LG02, LZCR08, Lot07b, MvdK06, MBT08, MLP^{+}08, MHWL01, MSM^{+}08, NS08, NGF06, NNE^{+}05, OCT^{+}03, OAKM07, OGPC09, OGPDOM08, Pal08, PK03a, Par04, Paw03, PEAAdLJB06, PEAAdLJB07, PDV05a, PDV05b, PHM07, PLSI09, RD03, RMCdS06]. fishing [RSM05, SSGM03b, SPS^{+}07, STK^{+}05, SS00a, SOH06, SAK05, TB04, The07, TH06, TCC^{+}09, Tia09, TL03, TMP^{+}06, TOBSJH06, VL08, VG01b, WG03, WM07, WPOB09, Xia04b, YLCS^{+}09, BP08, Ano01d]. fishing-down [PEAAdLJB06, PEAAdLJB07]. FishMeter [ØT02]. fitted [GLS^{+}09]. Fitting [Jon00, MS03]. five [JAKM05]. fixed [ABCB^{+}07, BR05, PWP02, SE02]. fixed-station [PWP02]. fizzing [NGM^{+}09]. fjord [BP01a, PP03, SF05, ØA08]. fjords [BBL07]. FL [How06]. Flake [McI09c]. flatfish [FP02, MTHL06, MTC06, RB06, Rye08, Som03, TC09]. flatfishes [MS02]. flathead [KKR^{+}05, MOOH00, YEA07]. flatheads [YEA09]. flavescens [SPS^{+}07]. flavicans [KLJ06]. flavobrunneum [MD02]. Fleet [BNL^{+}08, FCAD^{+}08, BQ04, BD04, CFFP07, FFL09b, MMB^{+}07a, MT08, OGPC09, Par04, SGS02, ULDG02, VSK03]. fleets [BDBD09, PCB01, SC08]. Fleksibest [FBS02]. flexi [OGK^{+}08]. flexi-grid [OGK^{+}08]. flexible [FCS^{+}05, LMMD06]. fling [CS08a]. Floated [FHJL08]. floating [KH05]. floodplain [CHBB04, JJWO09, PA05]. floor [NHT07, WK08]. Florida [Bur02, Ehr08, LCFP^{+}08]. flounder [He03, Laz08, LZZN09, SKF08]. floundering [But08, KPBD08]. flow [BA06, BSG02, EF02, RDM^{+}09]. flows [MS05b]. fluctuating [IPH05]. fluctuation [SAK05]. fluctuations [LAL07, MOK^{+}05]. fluorescent [WMS04]. fluviatilis [PTE09, PLMDB06]. flying [ARVC01, BI05, CC03c, CCT^{+}08, IMOO06, TYM^{+}01, YM00]. FM [ET00]. Foca [Güç08]. focus [Kor04]. foliacea [PK03a, RBD02]. following [ARLB07, Aló09, GMY^{+}09, JS05, MNML00, MG03a, MG03b, SMM07]. Food [Mar06a, PF00, AAFLP^{+}07, BSAM09, BP08, Cor01, Ism03, LK00, Man00, PEAAdLJB07, PPdC01a, SCM09, VSK05]. foothill [BNS00]. footrope [SW01, WSM02, vS04]. Forage [PMC^{+}07, SC07, VOS01, YSS^{+}09, McI01]. Foraging [NE07, MPQ07]. forbesi [BPW01, CPW^{+}06, PB03, YPD^{+}04]. forced [SE06a]. forecast [MOK^{+}05]. forecasting [CWS^{+}05, CGEHC07, GKV06, GESY^{+}07, PCS05a, Rod01]. forecasts [MSTC04]. foregone [CS02]. Forerunner [EMN07]. Foreword [FKM00, Sol03]. fork [CP00]. fork-beard [CP00]. form [TLD06]. formation [MN00, RC06, WWG03]. formulation [JCJ^{+}09]. Forsgren [McI09b]. Forsskål [BK02, GAFA04b, GAFA04a, GAF06, MM06a, MMB07b]. fortunes [Kro04]. four [ASM^{+}09, BGH05, CFE03, CFH03, CBGP06, GR00, LPP02, LLV^{+}08, PBDP00, RM05]. four-year [PBDP00]. Fourier [NBF09, SHBÅ07, TLD06, WG05]. fourth [McI07b]. Foveaux [CC09]. fragile [ROS04]. frame [KMK09]. framework [BBRS06, CS02, FCS^{+}05, KCH05, PZH02]. France [GJ09, DVAL^{+}09]. Francis [How06]. Fraser [MMS09]. Frazer [BNS07]. free [BSAM09, GNL^{+}09, HCH09, vSS09]. free-running [vSS09]. free-swimming [BSAM09, GNL^{+}09, HCH09]. Freeman [McI09c]. freezer [tHDC06]. freezer-trawlers [tHDC06]. freezing [Ogl09]. French [ABK03, BDB^{+}03, DR01, LD01, Pou01]. Fréon [Cur00, Gue05]. frequencies [ATB^{+}02, DHM08, GCD^{+}09, KLJ06, PC03]. frequency [BPRM00, COLB07, CSCC07, Ehr08, Hor00, JWR05, Joh03, QP00, RSM05, YFT^{+}02]. Freshwater [How03b, Mil02b, Sut02, Win00, ABS^{+}03, BCSP03, Bas99, Bas00, Bro05b, Chi00, DOQ08, Has06, JK07, KAN00, SLV^{+}08]. freshwaters [VBSK00]. Fridman [Kes00]. Froese [Sve14, Sve15]. Front [Ano03p, Ano03q, Ano03r, Ano03s, Ano04j, Ano04k, Ano04l, Ano04m, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano06q, Ano06r, PM09]. frontier [TPC05]. fry [JK07]. FSAMLR [Ano09r]. fuel [PFHC08]. Full [Ano03f, Ano03g, Ano03h, Ano04e, Ano04f, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano06n, SdPFAL09]. full-scale [SdPFAL09]. Fulltitle [Ano02h, Ano03k, Ano03i, Ano03j]. Fulton [KAHR02]. fulviflamma [GAF06]. function [HFL08, HHC06, OKWM05, PL00]. functional [GG06]. Functionally [GFM02]. functioning [YLCS^{+}09]. Functions [McI09a]. Funk [Coo00a]. Furgaleus [SDH00]. furnieri [ABL02, NV05, VH06]. Further [DMC^{+}09, GKF03]. fuscus [RBHP03]. future [Bla06, EMN07, FFM01, GSS^{+}06, Sym05a, How03b]. fuzzy [LZS01]. fyke [BP06, BMSJ05, PLD05]. fyke-net [PLD05]. fyke-nets [BP06].

G [Ell00, How03b, McI03b, McI05a, McI09a, McI09b, McI09c, Mil08, Ore03, Ste05, Sym08c, Win02]. Gabriel [McI07b]. gadid [MKGH05]. gadoid [IJ06]. Gadus [AFVSR^{+}09, BP01a, BG01, BLR^{+}08, BC04, CAJ00a, CAJ00b, CS04, EMKG07, GFL03, Hal02, HMH^{+}04, HBM^{+}00, HKF^{+}09, HH01, HMRG08, JSB04, JIDM02, JCJ^{+}09, LR08, LR00a, LK00, LR00b, MR05, ØA08, PP03, PBM06, PCHM07, RL03, SF05, SCM09, SLJ05, SSC^{+}03, Tal09, VN00, VBSC08, VKD03, WF07, WPOB09, WR05, vSSK07]. gag [SFCK03]. gahi [AM02, Hat00, Vil01]. Gaimard [NCW00a]. gained [JIGI06]. Gaining [Lor06]. galatheid [CS07a]. Galicia [MF03]. Galician [GPG^{+}03, OGSG07]. gall [MKGH05]. gallina [DNN^{+}03, MFAM05, PPL04]. galloprovincialis [RGGB07]. Gambia [EAT05]. game [Tri05]. gamefish [LSW06, LWM07]. gamefish-tournament [LSW06, LWM07]. gamma [Dic04]. gammarus [LCTB^{+}03]. GAMS [XPMQ04, VD04b]. Ganga [BNS00]. gariepinus [BW08]. Gashua [NJL^{+}00, NWML00]. Gastric [TH03]. gastrophysus [PW05, PPdA05]. Gastropod [GCN02]. gauge [FBL^{+}07]. gayi [OB02, LC06, OB02]. GBR [Gri03]. GBR-prawn [Gri03]. Gear [TS09, BGM^{+}09, BdL05, BMPK06, BACP09, BP06, BR00, DWW^{+}02, EH04a, ERG09, GJBY05, GBG^{+}03, Güç08, HFHO07, JGB^{+}06, KSS05, KLS^{+}07, MHWL01, NS08, OSB09, SRA^{+}02, ST00, SHL09, SGB00, SE02, SME02, SSK^{+}06, WRK06a, WRK06b]. gear-based [ERG09]. gear-related [GJBY05]. gears [JR02, KW03a, Paw03, PWB05, RSM05, WG03, WBCP08]. geelbek [HGSP01]. gene [BSG02, DHM08]. genera [KKJ00]. general [Bet04, Cam04, NC04, OSG08]. generalised [BPW01, BVW04, HMH^{+}04, MDS03, VD04a]. Generalized [WK01, BS04, Dic04, HPM04, JC04, KMELC05, OSG08, PEAAdLJB06, Pra02, WKM02, WIKM05, Xia04a, Xia04b, YAHAB01]. generate [PAP04]. generated [SDM04]. Genetic [CSM^{+}09, CBRN08, CMAS08, GCN02, JNDH00, JIDM02, MGJB09, PTR^{+}05, YMT^{+}06, Arc03, FML^{+}07, JDMN00, KN01, KIT^{+}04, MFC^{+}08, OIK^{+}06, OLN^{+}02, PD08, PMSD06, SCUADJ07, Tri06, WF07, vHMAO^{+}06]. Genetics [Abl06, YJV^{+}03, vHAN^{+}09]. genotypes [JS02, JIDM02]. Genus [PPGD08, CCD01, PD08, TLMN00, TRL06]. Genypterus [GOAARJRP09, PGF07]. geodesics [NBF09]. geoduck [MC04]. Geoffr [GBC^{+}03a, GBC^{+}03b]. Geographic [MM06a, SCM^{+}08, HC07, SMRH04, TCY01]. geographical [CC03c, DHM08, KVS04, MKGH05, PPL04]. geolocation [LNS08]. Georges [FKW06, Ove02, AFVSR^{+}09, BTC08b, FRF08, GHDM08, LG02]. Georgia [ZB00]. georgiana [ABE^{+}04]. Geostatistical [JR07, ASM^{+}09, PR03, RMA^{+}05, RSMZ06]. geostatistics [BR01, PLPN09]. Germany [AM03, ABF08]. Gerres [IMSS06, IOS07]. Ghana [NBA^{+}09, ODdGV02]. Ghanaian [BC02]. Ghost [AAA^{+}06, CS09, TL03, AMAOMM04, BCC09b, HMM^{+}01, HLHF03, RW03, SSGM03b]. ghost-fishing [HLHF03]. giant [GML06, JNDH00, KCRT03, PK03a, WHML04]. giganteus [GCN02]. gigas [ARVC01, MSN01, MQVSN04, Mar06a, Mar06b, MBCMNMHH01, NMHHMB^{+}00, NMMTCV^{+}06, NNA01, NNA02, ORVMP06, RWMBHH06, SCUADJ07, SN01, TVY01, TYM^{+}01, WYR06]. gilchristi [GMS06]. Gill [FMCS05, PLMDB06, SME02, BP06, FSB^{+}02, Gra02b, HWM02, KJEF03, LLJ^{+}08, MNML00, MAG00, RCA^{+}07, SSGM03b, STK^{+}05, SE02, TEG03, TL03, VAYD04, VSK03, WFJB09]. Gill-net [FMCS05, BP06, MNML00, VSK03]. gill-nets [RCA^{+}07]. Gillnet [CC03a, OMR09, APP^{+}07, BS06, GSS09, GFL03, GJBY05, KAHR02, OKPR04, PW05, PKŘ^{+}08, Rue07, VKČ^{+}09]. gillnets [AAA^{+}06, DVAL^{+}09, EJA00, GRC^{+}09, He05, He06, HSSR06, HLHF03, LET07, PKŘ^{+}09, RD03, SB07, TF09]. gillnetting [ESB09]. girth [SCLM06, SK03b]. gladius [BH01b, CPSBRD^{+}02, DMA07, HHTE05, MGJB09, Pra02, RYSN08, SWPY05, WSPY05, WSPY07]. glauca [dSHG08, LSH04, WK01]. Glaucosoma [JS05]. Glenelg [GCMJ08]. GLMM [BBJG04]. GLMMs [VD04b, XPMQ04]. GLMs [VD04b, XPMQ04]. global [Ano07y, EH02, Lon07, WRK06a, WRK06b]. globulosum [Nar06]. Gnathanodon [GAFA04a]. gnomonic [MAASMB05, RRAS03]. Goals [BCLM08]. Gobiidae [EKU02]. Gobiidea [LBABH09]. Gobius [LBABH09]. God [GHDS08]. Godø [KW03b]. Going [BA06]. golden [BSMFH07, FGJD08, SKKP05]. Gómez [YPM^{+}06]. gonad [LBABH09]. Gonadosomatic [LBABH09]. Gonatus [Bjø01]. goose [MF03]. gorbuscha [BCM08b]. Gordon [Sym08a, Ano01i]. GOV [Dah00]. GOV-trawl [Dah00]. Governance [CMM^{+}07, Sym06]. GPS [Pal08]. grab [HB01, TB04]. gradient [ABF08]. granatina [AGB07]. Grant [Coo00a, Coo01, Gra02a, McI01, McI02c, McI06b, Ore03]. grate [Gra03, RH06]. Gravimetric [MFB08]. gray [SDS09]. grayling [NNE^{+}05]. Great [BMBM05, BPW^{+}03, BPH^{+}06, Fri07, Gri03, GWB07, HWW^{+}09, HW00, MLP^{+}08, NCW00b, NCW00a, PBW^{+}09, QGBS04, WMB08, WMD07]. greater [CP00]. Greece [HAO00, LS00, CPT06, FK03, KVS04, PKT^{+}02, TDK^{+}05, TSTK06, TSTK08]. Greek [GKV06, PK03a]. green [ABAA^{+}07, CH03, GC04, KCH05, NAD^{+}07, PZH02, ZP05]. Greenland [WSP06, ANS^{+}01, FP00, FB06, GRB01, HLHF03, KGB09, LR08, LR00b, Wie04, WSP06, WBHG01]. greenlip [DDHS06]. Grenada [GB07, MB07]. grenadier [KN01, LDA01, PCAS04]. grey [FGJD08]. Grid [KHMG06, BKD^{+}02, CRD06, EH04a, FCL^{+}05a, FCML05, FCL^{+}05b, Gra03, GOF^{+}04, GF06, GL05, Gri06, KME^{+}02, KI04, KI05, MH01, OGK^{+}08, VU08, Gri06]. grids [GFSM07, GSL08, JIGI06, LMMD06, MGS02, MB00, Pol02, SGL08]. grimaldii [He05]. gripping [DACS08]. Groot [McI03a]. GROTAG [Tal09]. Ground [MTGH00, AS04, JHGS04, NGF06, PK03a, RD03, RR03a]. groundfish [BHB^{+}04, BCLM08, BS04, CFH03, CRSM04, CK08, HP02, HP09a, HP09b, MBG^{+}07, PH01, PDH08, SW02]. groundgear [Dah00]. grounds [AMAOMM04, ASD09, BEA^{+}08, He03, SLA00, TVY01, VBSC08, WGGN06]. Group [How06, Din05, KA05, ØA08]. grouper [BB04, BD04, ED07, GHCPPLC^{+}05, GBL^{+}05, GAFA05a, HS03, LCFP^{+}08, ML04b]. groups [CC03c, NV05, PPL04, SRTP06]. growing [Ker00]. Growth [AC05, BHJ02, BK02, CHN08, CP00, CC01c, FBJM00, Fri07, GR00, GT04, LDGA09, LPDA01, LPP02, LMRQNM^{+}05, LDE^{+}02, NMMTCV^{+}06, New02, PL02, PMG^{+}06, PRdPG08, PJR^{+}05, QVNMGM00, RCG02, VAF04, Vil01, ACPP06, ARLB07, AHABAA^{+}02, AdBP02, ASAGR06, AD00, ARVC01, BH01a, BAH09, BGP^{+}04, Bas99, Bas00, BLV06, BPRM00, BP01a, BMF^{+}10, BBJG04, BG04, BG05, BGB06, Bur02, BF00b, CC03b, CS08c, CMGB01, CY06, CC01a, CCB02, CMC09, DSPLR06, DMBF08, DGS^{+}05, DKGC00, Ehr08, EKU02, Fab06, FCC^{+}09, GRPLE05, GBC^{+}03a, GBC^{+}03b, Gon06, GR07, GAF06, GCC^{+}08, GMP05, Hat00, HMM^{+}09, Hoo06, Hor02, Hur07, HLH08, IMSS06, Ism03, IYI07, JCS01, Jon00, JIDM02, JLC04, JCLL08, JHK08, KA05, KR00, Kat06, KCK03, LLM05, LP00, LJP05, LS00, LSH04]. growth [LDN04, LLJC09, LE01, LCFP^{+}08, LLV^{+}08, LMASHV^{+}03, LCD^{+}02, MRCH04, MQVSN04, MOOH00, MK06, MCAO^{+}05, MR08a, MOGG06, MNM00, MN00, MC04, NT05, NCW00b, NCW00a, OMF^{+}03, PLDS03, PH09, PH04b, PBM06, PRdPG07, PRM04, Pun03, PHGT06, QGBS04, RMBG02, RC06, RPC^{+}06, RCGC04, RCGMQV05, RLG^{+}01, RRJ09, RR03b, STCH09, SGS09, SCM^{+}08, SFCK03, SM04b, SSC^{+}03, Tal09, TS09, TP00, TÁTWD09, TBRL00, TT01, TEYA02, WT00, WKCC06, XM00, YM00, YST^{+}02, ZR08, ZW09, ZLC09, NS06]. Growth-dependent [FBJM00, JHK08]. grunt [BC02]. Guadalquivir [SBGG^{+}05]. guatucupa [Caz00, MM06b]. Guaymas [NMMTCV^{+}06]. Guerra [Pie09]. guggenheim [VMQ08]. Guide [Ano02i, Ano03l, Ano04h, Ano04g, Ano05k, Ano06o, How06, McI05d]. Guinea [AAHE04, FBV06]. Guinean [Mai09]. guitarfish [IYI07]. Gulf [BD04, GRFE06, Gon06, GNC07, McI06b, McI07a, NS04, Nar06, NGF07, Ste03a, Sva05, TSTK06, TSTK08, VSS07, YLD^{+}05, ACMMCEGM02, APC07, All07, AAHE04, ASAGR06, BPRM00, BFM08, BSG02, CWS^{+}05, CHN08, CPCD05, CQJX08, CXQ^{+}09, DGP02, ESC08, FW06, FOGKT08, GKUO07, GHDM08, GAFA04b, GAFA04a, GAFA05b, GAFA05a, GAF06, HMM00b, HMM00a, He06, He07, HMM^{+}01, Hoo06, HMRG08, JS02, JSR04, JDAM08, JJJ03, LC06, LC07, Loh08, LMRQNM^{+}05, MSN01, MQVSN04, Mar06a, Mar06b, MAdAMASCM09, MMH05, MR08a, MBCMNMHH01, NMHHMB^{+}00, NAD^{+}07, Ove02, PBDP00, QVNMGM00, RMVJGMMV07, RRAS03, RRASLB03, RBHP03, RH06, RAHHGM^{+}02, RB04, SBFR00, SWM09, SGS02, SGB00, SFCK03, SRS08, SSC^{+}03, Tal09, TGH^{+}08, VSS07, VVH04, WBCP08, XM00, Xia04a]. Gullmar [ØA08]. gunnari [MBDM05]. gunnel [HLH08]. Günther [Mag01]. guttatus [ASAGR06]. Gymnocephalus [Ogl09].

H [Ano01c, McI05b, McI09c, Ore03]. Haan [NAD^{+}07]. Habitat [AKK^{+}03, BHB^{+}04, BMB05, BMB06, GCS^{+}06, Laz08, LHMO09, McI02a, MR08a, VFCC06, ATM08, BGP^{+}04, BRL^{+}05, BLC08, BLR^{+}08, CC09, GCL03, HCHO^{+}03, HC09, PR03, RLSJIA07, SB08, SJ05, TCC^{+}09, ZR08]. habitats [AR05, ESB09, LZC^{+}08, SGP02, How03a, McI03a]. habits [CMC^{+}06, DG02, MSA00, VKD03]. Haddock [Ano08l, He08b, AFVSR^{+}09, BSC^{+}08, BTC08b, CTK^{+}07, CCBG02, DJGC09, FÖB08, FRF08, GM07, GK01, GKF03, GLH07, Hal02, HMH^{+}04, Her05b, HO05, HO06, HS00, JSO08, LR08, Nee08, NPC^{+}08, OGK^{+}08, ÖW02, Pál03, PCMC08, SGL08, Ano08l]. hadrah [ABCB^{+}07]. Hae [KW03b]. haemolymph [HBBG06]. Haemulidae [AHAAD01]. Haggan [Sym08c]. Hake [Bet04, SSGM03b, ASR05, AMUE01, AC05, BCSA08, BLV06, BSR05, BBL07, EGB^{+}01, FAL^{+}01, FB06, GGVLSR03, GNC07, GMV00, GMV01, GEH^{+}07, HC00, JI07, JMGC06, KA05, LC06, MPE04, MPM^{+}06, MLC03, MMH05, MNBM05, OB02, OQE03, PMM05, PRdPG07, PRdPG08, Pou01, RCA^{+}07, WPS^{+}07]. hakes [VGBF02]. half [LOT07a]. halfbeak [NMF^{+}08]. halibut [ANS^{+}01, CGEHC07, GB03, GRB01, HLHF03, KS08, KF02, KGB09, LR08, Loh08, MS05b, SOH06, SK08b, WSP06, WBHG01]. Haliotis [HFBM08, LRBH07, MWJ04, PPGD08, RB09, ZLC09]. Hall [Ano00c]. Ham [Bas00, Bas99]. Hamilton [GAFA05a, KK09]. hamiltoni [JS02]. hand [PEB03]. hand-line [PEB03]. Handbook [McI05a]. handline [MLJT09, PVP04]. harbour [LET07]. hard [McI02e, McI02d, Sol05]. Hardback [McI06b, McI06a, McI07b, McI08, McI09a, McI09b, McI09c, Mil06, Ste03a, Ste05, Sym07, Sym08a, Sym08b, Ano00b, Ano00d, Ano00e, Ano01c, Ano01d, Ano01e, Ano01f, Coo01, Cur00, Loc02, McI01, McI02f, McI03a, McI03e, McI03b, McI05a, McI09d, Mil02a, Mil02b, Mil03b, Pie09, Win02]. Hardcover [How06, How03a, How03b, McI02c, Ore03, She02]. harengus [BBK08, FUM02, FBJM00, HHC06, NDCK09, NLJS01, Ove02]. Harris [Mil08]. Harvest [Jen00, ABD^{+}09, BDIP08, CP08a, CK08, CGT08, CHQ05, DB08, DSK^{+}08, GOAARJRP09, HCM05, HH01, IWBJ08, Jen02, MAdAMASCM09, PBR05, PPGD08, PCS05b, RM05, SWM09, SST^{+}08, Sve15, Vas03, WIJB08]. harvested [BB01, Kop06]. Harvesting [Ano08l, BP01b, CS02, GPI04, He08b, RGGB07, SSIE06]. harvests [Hen09, PIdlHGV05]. hatch [AC05]. hatch-date [AC05]. hatchery [CPR07, DDHS06, MNBM05, PB04, QDV05, SLE09, SKF08]. hatchery-produced [DDHS06]. hatchery-reared [CPR07, SKF08]. hatching [AD00, HK02, MOGG06, QVNMGM00]. haul [HO05, MSB^{+}08, MBM04]. haul-back [MSB^{+}08]. hauling [LR08]. Hauraki [HMM00b, HMM00a]. having [LK02, RB00, YAHAB01]. Hawaii [BCMM09, CLPP05, CGLZ04, PL04, PL06, WK01, WKM02, WIKM05, WHML04, ZDB^{+}08]. Hawaii-based [BCMM09, WK01, WKM02, WIKM05]. Hawaiian [MHWL01]. hbk [Ell00, Kes00, Sut02]. headed [Lon07]. headline [He06, MBGW02]. health [BBRS06, FR03]. healthy [PEAAdLJB07]. hebraicum [JS05]. Heckel [TEYA02]. Hector [SH09]. Heidelberg [Ell00]. Heifetz [Coo00a]. height [DOQ08, HJ03, He06, JRG08]. held [Bla06]. Helicolenus [ACPP06, MNM00, MF09, PH09, RMCdS06]. helminths [SN01]. Hemingway [Loc02]. Hempel [Ste05]. Hendala [HJ06]. heptadactylus [PJR^{+}05]. herding [RB06]. Herpetologists [McI05d]. Herring [LAL07, ABE^{+}04, BBK08, CL04, Cor01, EPA00, EH04b, FGOF08, FUM02, FBJM00, HHC06, Hor05, KPL^{+}07, KPL^{+}09, NDCK09, NLJS01, Oja06, Ove02, PHM07, RPP04, SC07, SSH^{+}06, VOH02]. Hess [MCB^{+}07, MSB07]. Heterodontus [TÁTWD09]. heterogeneity [ABF08, DHM08, GPI04, JIDM02, Pun03, WRWK02]. heterogeneous [SKMT02]. hexagonal [TASF09]. hierarchical [APP^{+}07, ZLC09, Ste03a]. Hierro [Pie09]. High [WM00, ABAA^{+}07, BW02, CS04, Cor01, IMOO06, JWR05, KJ08, KCRT03, RSM05, RGSB08, RVS02, RR03b]. high- [JWR05]. high-frequency [RSM05]. high-quality [BW02, KJ08]. High-resolution [WM00, KCRT03]. high-speed [RGSB08, RVS02]. Higher [SCM09]. highly [SKMT02]. Hilsa [RC06, SMRH04]. Himalayan [BNS00]. Hindustan [McI06a]. Hippocampus [WMS04]. hippoglossoides [ANS^{+}01, GRB01, HLHF03, KGB09, LR08, WSP06, WBHG01]. Hippoglossus [GB03, KF02, Loh08]. hippurus [FOGKT08, PMSD06]. Hiroshima [GMY^{+}09]. hispidus [MRCH04]. histopathology [LBABH09]. Historic [GDAM07, CGT08, LC07]. Historical [CdMLGC09, GHCPPLC^{+}05, MSM^{+}08, OM07, PSS^{+}07, dSHG08, BK09, CNH09, ERG09]. histories [FCC^{+}09]. History [McI08, Mil02a, AGS^{+}08, ABFV02, AGB07, BTC08a, CCMS^{+}03, CC03c, DS07, FF01, FR08, GEH^{+}07, HGF06, HLH08, KAN00, KIT^{+}04, LD08, MAASMB05, MRT03, PCS05b, RRAS03, SJ05, ZB00]. Hoagland [ZR04]. hobbled [SB07]. Hokkaido [Uki06]. Holdich [Sut02]. holistic [AMC^{+}08a, Cad02]. Hollingworth [Mil03b]. Holm [McI08]. Holothuria [PK06, SK08a]. Holothuroidea [RBHP03]. Holt [Jen00]. Homarus [BHH08, IWW03, LCTB^{+}03, ST03]. Homing [Loh08, CB00, WR05]. homogeneity [KIT^{+}04, PMSD06]. HOMSIR [AMC^{+}08a]. Honduras [GBL^{+}05]. Honour [Sym08a]. hook [Aló09, BB04, BNL^{+}08, FMC^{+}09, GCJ07, JS05, MWB^{+}08, OSC^{+}05, RCA08, SE02, TNFF03, WBHG01]. hook-and-line [BNL^{+}08]. hook/bait [GCJ07]. hooked [FMC^{+}09, iTMI06]. hooking [AAP^{+}09, AWY02, BR05, CSSO03, GCJ07, IV08, MWB^{+}08]. hooks [BR05, BCMM09, CSSO03, KG06, WEK^{+}09, YKM06]. Hopf [VVH04]. Hoplostethus [CAFT00, Cla01, KM02, SR06, SRH^{+}02]. hoppers [HBJ03]. Horizon [Ano01c]. horizontal [BMW07, KS02, RHF^{+}08, RG07, TKFJ09]. horizontal-aspect [BMW07]. Horse [Aba08, AGS^{+}08, AMC^{+}08b, CBRN08, CMAS08, CMH01, FAL^{+}01, GCA^{+}08, KM08a, MCM^{+}08, MFC^{+}08, MACS08, MPA08, RdFBS01, SMSZ08, ZC02]. host [GPG^{+}03, MFC^{+}08]. hot [Hat00]. Howell [She02]. http [McI07b]. hubbsi [GNC07, MPE04, MPM^{+}06, PMM05, WPS^{+}07]. Hudson [ZSP03]. Humacao [MDJR05]. human [VLD^{+}00, FR03]. Humboldt [RWMBHH06]. Hutton [HSGF06]. Hydraulic [MFAM05, DNN^{+}03]. hydro [MSSJ05]. hydro-acoustic [MSSJ05]. Hydroacoustic [GCD^{+}09, KFO^{+}04, KG04, PMT06, PTE09, SRTP06, CXMM04, ET00, KH05, Meh06]. hydroacoustics [ABRB00, AC00, MC00b, SG06, SW00, WFJB09]. hydroclimatic [MUO02]. Hydrodredge [SGRK09]. hydrographical [AMUE01, NLJS01]. Hydrography [RWT01, KMB01]. hydrological [Chi00]. hydrology [ZR08]. hydropower [WCM04]. hydrostasis [GR02]. Hyporhamphus [NMF^{+}08]. hypotheses [VH06]. hypothesis [GR02, PRdPG07].

IA [How03b]. Ianelli [Coo00a]. Ibacus [Ste03c]. Iberian [FH03a, PRdPG07, CP00, CMC09, DALP01, LPDA01, MACS08, PRdPG08]. Ibero [DTSR07]. Ibero-Atlantic [DTSR07]. icefish [MBDM05]. Iceland [JIDM02, PEB03, SKP03, TJ05]. Icelandic [BM03a, BS04, IJD04, Jak01, Mag01, Pál03, Pál05, PBM06]. ICES [CP00, NLJS01, SGS02, ERG07, LDA01, LD01, PVP04]. ichthyofauna [PSPCVCD09]. Ichthyologists [McI05d]. Ichthyology [McI05a]. ichthyoplankton [JDAM08]. ichtyoplankton [BR01]. iconic [SGS09]. Iddya [Lai00]. ideas [Lev05]. Identification [Aba08, HSW09, How06, MHS^{+}05, McI05d, AMC^{+}08a, AMC^{+}08b, GR07, LRG^{+}00, MCM^{+}08, MFC^{+}08, MTGH00, MPA08, OIK^{+}06, OS05, PD08, SÁR^{+}05, TLD06, TRL06]. identified [CMOM00]. identify [BHB^{+}04]. Identifying [BW02, BBM^{+}06, TSTK06, MBT08, War06]. identity [AMC^{+}08b]. if [MBGW02]. IFC [Ano02j, Ano02k, Ano03n, Ano03o, Ano03m, Ano04i, Ano06p]. ignobilis [WHML04]. II [Coo00a, BM03b, Her05b, NWML00, NS06, PDV05b, PM09, WRK06b]. III [Mil02a, BM03c]. ilisha [RC06, SMRH04]. ill [Kes00]. illecebrosus [HH06, PBMB03]. illegal [AP05, RB09, WVW^{+}08]. Illex [AM02, BQN05, CUT06, HH06, Lap02, MHS^{+}05, MCP01, PBMB03, WGR08]. illicia [WWG03]. Illumination [JGB^{+}06, RB06]. illustration [WMdS09]. image [FCC^{+}09, LRG^{+}00, MFB08, TBRL00]. image-based [FCC^{+}09, MFB08]. images [BMC07, FL05, Fab06]. imaging [RSM05, RGSB08]. immersed [CS07b]. immersion [LZZN09]. Imo [EKU02]. Impact [CK02, FM04, PSvDPM01, PPdC01a, RBH^{+}08, Tia09, ZW09, Aló09, AWY02, BLKR09, BHM^{+}06, BRRL09, BCM08b, CL04, DMBF08, FFM08, GC03b, GBG^{+}03, Jen02, MM05, MD02, NS08, OSB09, RB00, RE07, RPFR06, SdPFAL09, SBD^{+}05, SBD^{+}06, WDC02, ZDB^{+}08]. Impacts [HSSR06, Bal09, CLPP05, CC09, CXCD07, CHDGM06, EH02, FMGG08, Gri03, HF07, HT05, Mey07, PHM07, RJ05, SGRK09, TOK^{+}05, WMdSB09, YLCS^{+}09, WIS03]. impairment [DO06, ROS04]. impedance [WH08]. implant [BLB07, WMS04, ZW09]. implementation [BDIP08, DDP^{+}06a, OJ06]. implementing [SST^{+}08]. Implication [KLJ06, CHL01]. Implications [GRC^{+}09, JTdH06, SC08, SLE09, SOH06, TLH07, VN00, WMdS09, YEA07, AAP^{+}09, AM03, ABF08, BG01, Bur02, CRG^{+}02, Dem04, DV00, EBA^{+}03, FZHL09, GSA^{+}06, Gra02b, HBKJ07, KCKP02, LZCR08, MMB07b, MAG00, PKP04, QGBS04, RL03, RSST07, RB06, SALnM07, Ste08, SMJS01, TB04, TSTK08, Vel03, WM04]. Importance [HMR^{+}08, BM08, CRSM04, MKGH05, OGPDOM08, RCA08, SCP01, War06]. important [CdMLGC09, MTC06, SC07, WG03]. improve [GLK00, HS00, JI07, MHF02, Pri09, RHF^{+}08, RCD07, Smi03]. Improved [LR08, MSTC04, KM02, vS04]. Improvement [BSS06]. improvements [DUR05, Kor04]. Improving [Dun09, FAN03, Joh03, LNS08, MMZ09, Sho09, vM03, ET00, MGTB08, PFHC08, PCV04, vSSK07]. inactivity [MC06]. Incidence [TGS^{+}00, BR05, MGJB09, PSS^{+}07]. Incidental [MROGVR09, MD02, SJ05, WKM02]. incidentally [HLL^{+}08, dSH01]. inclined [EF02]. inclinometer [WRWK02]. Including [PSTM06, Joh03, JJWO09, LNG04, NJL^{+}00, NWML00, WMdS09]. incorporates [Hor05]. Incorporating [BRL^{+}05, But03, GC04, HP09b, JMGC06, NC04, CHL01, DUR05]. increase [DJGC09, MBGW02, PBDP00, TSJ04]. increased [CGLZ04]. increases [GDFH07, OL07]. increasing [AGB07]. increment [CS07b, Ehr08, Ker00, MN00, MBDM05, RC06, WWG03]. increments [ACE^{+}08, Dou08, MÁ08, NMF^{+}08, NS06]. independent [APS07, GDFH07, LBKG08, PWP02, SHBC02, ST03]. indeterminate [GCA^{+}08]. Index [Ano00f, Ano00g, Ano02b, Ano02-28, dSHG08, Ano00a, Ano00t, FHI07, GBG^{+}03, HK06a, HTF^{+}09, IHF08, LBABH09, OGPC09, PH04a, SLJ07, VFCC06, YLYL09]. index-removal [FHI07, IHF08]. India [Bas00, McI06a, Bas99, BP08, PJR^{+}05]. Indian [CHL01, CHW^{+}08, DHM08, JCS01, MGJB09, PMC^{+}07, Row07, RG07, SND^{+}08, VSK05, XBS^{+}07]. indicated [LRW09]. indicating [KG04]. indications [GB03]. indicator [KAN00, KCK03]. Indicators [YFL08, CAM^{+}08, Ein09]. indices [Cam04, DUR05, Din05, HBM^{+}00, Hur07, LD01, MS03, MD09, PAB05, PLPN09, TCC^{+}09, VGBF02, YPDS05]. indicus [MOOH00]. indirect [BACP09, CCG^{+}02]. Individual [KPL^{+}07, KPL^{+}09, AB08b, AB08a, Ara09, But03, FCC^{+}09, PHGT06, PK06, WNRd00, Xia04b]. individual-based [But03, PK06]. individuals [BW02, Tri06]. Indo [Hoo06]. Indonesia [WGDP06, vOPvDM02]. Indonesian [PSvDPM01, WC07]. Indrani [Lai00]. induced [CCA^{+}03, KSS05, LG02, OMW^{+}03, SSK^{+}06]. inductive [Cor02]. industrial [CL04, EH04a, KHMG06, ZCdG06, dSBAC06, vZRMvD02]. industry [EH02, NW03]. inexpensively [SDS09]. inference [Kat06, RLSJIA07]. inferred [BHH08, CHL^{+}06, CHW^{+}08, FDG^{+}07, KM08c, MRT03, WCC^{+}09]. infestation [BCM08b]. inflated [MLCGRV07]. Influence [BG05, KP07, KLU01, MPM^{+}06, Orl03, RB06, SG06, SAK05, VMQ08, WYR06, YEA07, AAP^{+}09, AAB05, BB04, DWW^{+}02, HO06, JS05, JSO08, Kno07, Mar07a, MUO02, RO00, SLB07, Tri05, WPS^{+}07]. influenced [DACS08, JR02]. Influences [BH01b, BPD^{+}04, CPW^{+}06, GJBY05, GWR07, MM06a]. influencing [SMB^{+}08, WK08]. influential [GCMJ08]. informal [GLCC09]. Information [LHMO09, AB08b, CUT06, GSS^{+}06, HC07, HP09b, JL00, JI07, Lor06, MMZ09, Vas03, WL06a, WB08]. informs [Moo03]. infracommunities [PBR05]. ingestion [AB08b, AB08a]. inhabiting [DSPLR06, Yil02]. Inherent [KM08b]. initial [AGS^{+}08, BS06]. initiative [SSG06]. injection [ARLB07]. injuries [DGS^{+}05]. Injury [CSSO03, BB04, BCSP03, DACS08, MWB^{+}08, OSC^{+}05]. Inland [AIC08, NJL^{+}00, NWML00, AM03, GAC^{+}06, HIRM06, eJSK09]. inner [WF07]. inodorus [Kir01]. input [Cad02, DOQ08, DDP^{+}06b]. inputs [UM02]. insect [KLJ06]. inserting [GOF^{+}04, OGK^{+}08]. inshore [LPHH09, SLA00, SC08, TOK^{+}05]. Insights [CHDGM06, McI08]. instantaneous [STI08, YLYL09]. Institute [Gue05]. institutes [AFG^{+}05]. instrumental [Cor02]. instrumentation [EGJ00]. insular [GCS^{+}06]. insularis [LHMO09]. intake [BSAM09]. integrated [SMD03]. Integrating [AMC^{+}08b, ML04a]. integration [MC00a, SLV^{+}08]. integrator [Kor04]. intensity [GDDJ09, RO00]. intensive [GMY^{+}09, HH08]. inter [CG04, KA05, VMSR08, vZRMvD02]. inter- [VMSR08]. inter-annual [CG04, KA05, vZRMvD02]. interacting [GGA08]. interaction [MMLR08, PPdC01a, PAF^{+}08, SPS^{+}07]. Interactions [BM03b, BM03c, FR03, NW03, Nas03, PB04, WIS03, KHMS09, McI03d, OM07, PL06, WR02]. interactive [GCH^{+}07]. Interannual [AD00, FOGKT08, LMASHV^{+}03, Tia09, PB03]. interdecadal [Tia09]. interdependency [SW02]. interest [MSVF05]. internal [DYWHZ04, LLV^{+}08]. International [Ano07w, Ano08l, IK07, McI03b, Din05, EMKG07]. interpretation [ATB^{+}02, BM00]. interruptus [ARCLBPdP07, Vel03]. interspecific [AM02, SAK05]. interspecific-relationships [SAK05]. intertemporal [MPG^{+}08]. intertidal [HCM05]. interventions [CdMLGC09]. interview [SPS^{+}07]. Intra [MMB^{+}07a, BPW01, PCS05b, VMSR08]. intra-annual [BPW01]. Intra-fleet [MMB^{+}07a]. intra-specific [PCS05b, VMSR08]. intractable [DA07]. intrinsic [CP08b]. introduced [BK02, NS06]. introducing [BR00]. introduction [CHDGM06, Gor01, KMELC05, Paw03, PLD05, McI09c]. invasive [AAH06, AKA06, AAKT07, Kno07]. Inverse [AM02]. invertebrates [JRG08, WRGS09]. investigate [CSR06, MTRM05]. investigated [PAR^{+}03]. Investigating [BVH^{+}08]. Investigation [SHL09, SMB^{+}08, BFB^{+}03, HFHO07, KS08, KJS^{+}06, OMW^{+}03, RW07, SK08b, tHDC06]. Investigations [GMR01, GMR02]. involvement [RE07]. iodide [WVW^{+}08]. Ionian [PK03a, CPT06, MCL04]. Ionic [Ano00b]. Iowa [How03b]. ir [BCLP00]. Iranian [FGJD08, FZHL09]. IRD [Gue05]. Ireland [AKA^{+}01, BRO05a, BPD^{+}04, CCB02, CK02, FAL^{+}01, HTF^{+}09, SR06]. Irish [ERG07, OMF^{+}03, RCA^{+}07]. Irminger [JNDH00, JDMN00]. iron [LET07]. iron-oxide [LET07]. ISBN [Ano00b, Ano00c, Ano00d, Ano00e, Ano01c, Ano01d, Ano01e, Ano01f, Bar00, Coo00a, Coo00b, Coo01, Cur00, Ell00, Gra02a, Gue05, How03a, How03b, How06, Kes00, Loc02, Lon02a, McI01, McI02a, McI02c, McI02e, McI02d, McI02f, McI03a, McI03e, McI03b, McI05a, McI05b, McI05c, McI05d, McI06b, McI06a, McI07b, McI07a, McI08, McI09a, McI09b, McI09c, McI09d, Mil02a, Mil02b, Mil03b, Mil06, Ore03, Pie09, She02, Sol05, Ste03a, Ste05, Sut02, Sym07, Sym08a, Sym08c, Sym08b, Van01, Win00, Win02]. ISBN-13 [McI08, Sym07, Sym08a, Sym08b]. Island [GCB^{+}06, PH09, SKF08, KVZ05, LHMO09, AFJC08, MB04, TVB^{+}09, TdlMH^{+}03]. islandica [TJ05]. Islands [ABFV02, BMB06, CLGR^{+}04, Daw08, GBL^{+}05, MMN00a, PGS^{+}08, CGT08, BMB05, CAM^{+}03a, DG02, HGHLCH02, LB05, LJP05, MRCH04, MCP01, PL02, PLDS03, PMG^{+}06, PPH01, SWC^{+}05]. Isles [MB04, YJV^{+}03]. isochronal [JCJ^{+}09]. isolated [LPG^{+}09, YMT^{+}06]. Isolating [BWM07]. isolation [MGR^{+}02, vHAN^{+}09]. Isostichopus [RBHP03]. isotope [GB03, WM00]. isotopes [BBM^{+}06, GMP05]. isotopic [GJSW04, JSB04]. issue [Ano03k, Ano07x]. Issues [HIRM06, McI02d, KAO04, SSG06]. Istiophorus [Hoo06, RAHHGM^{+}02]. Isurus [CMJ05, CL09, RCGMQV05]. Itaipava [BNL^{+}08]. Itaipu [MAG00]. Italian [CCCC04, LPL^{+}06, LPG^{+}09]. Italy [ASM^{+}09, GCB^{+}06, GTB04, LCD^{+}02, LDE^{+}02, MAG^{+}09, MFAM05, RCG02]. ITI [EGJ00]. ITQs [Bra04]. IV [WIS03]. IXa [CP00, SGS02]. Izmir [AAA^{+}06].

J [Ano00b, Ano00e, Ano01c, Ano01f, Bar00, BNS07, Coo00a, Ell00, Fra07, How03b, How06, McI02a, McI03a, McI08, McI09a, Mil03b, Ore03, Sym07, Win00, KG06]. J-style [KG06]. jack [ACG^{+}01, CSM^{+}09, CHSA02, CPR^{+}08, KWK^{+}09, YFT^{+}02]. Jackknife [RYSN08]. jacks [MHWL01]. Jackson [TÁTWD09]. jacobaeus [Kat05]. Jamaica [Gus02]. Jamieson [Ore03]. January [Ano00p, Ano01p, Ano02u, Ano03-29, Ano03w, Ano04n, Ano05r, Ano06t, Ano07n, Ano08m, Ano09u, Ano09z, Ore03]. Japan [CLH^{+}08, BM05, BI05, FAW04, GMY^{+}09, HK06b, IMSS06, IOS07, IFH^{+}09, KIT^{+}04, MOOH00, MOK^{+}05, MOGG06, NT05, OIK^{+}06, SN09, STCH09, SAK05, SWA08, SKF08, Tia09, WY06, YTK06]. Japanese [AM00, AKA06, CC01c, CLH^{+}08, FAW04, IMSS06, IOS07, IFH^{+}09, KK06, LZZN09, NMF^{+}08, SAK05, SWA08, SKF08, TOK^{+}05, Uki06]. japonica [AAH06, AKA06, AAKT07, SKF08]. japonicus [AM00, FAW04, HK06b, JHK08, Jun08, KWK^{+}09, SGS09, SWA08, TY02, WLY06, WY06]. Jasus [BM09, BBJG04, BGB06, DGS^{+}05, HSGF06, HBBG06, LPHH09, PHGT06, ZFJ03]. Java [CNH09]. jelly [Kno07]. jellyfish [PK03b]. Jenkins [BSMFH07]. Jenyns [GEYPC^{+}09]. Jeremy [Gue05]. Jersey [MB04]. jig [RB00]. jigging [CLC08, MCP01]. John [Ano01f, Lon02a]. jordani [GHS04, HJ03, HJ07]. Joseph [Mil02a, TGH^{+}08]. journals [Xia06]. Jr [How03a]. July [Ano00o, Ano01n, Ano02m, Ano03u, Ano04q, Ano05-27, Ano06-27, Ano07t, Ano08n, Ano09w]. jumbo [ARVC01, MSN01, MQVSN04, Mar06a, Mar06b, NMHHMB^{+}00, NMMTCV^{+}06, NNA01, NNA02, ORVMP06, SCUADJ07, SN01, TVY01, TYM^{+}01]. June [Ano00i, Ano02w, Ano03-27, Ano05u, Ano06u, Ano07p, Ano08t, Ano09t]. juvenile [ARLB07, APS07, AD00, BSR05, BLB07, BCM08b, CC01b, CMOM00, CS07b, DYWHZ04, DDHS06, Dou08, FK04, GJSW04, GCS^{+}06, GT04, HJLL07, KWK^{+}09, LDGA09, LF02, MGS02, MK09, MMLT05, NS08, NMF^{+}08, NJE00, PCD03, PRdPG08, QVNMGM00, RPC^{+}06, RDM^{+}09, RR03b, RO00, SMD04, SBGG^{+}05, SFCK03, SKF08, SJ05, VFCC06, WCM04, WMTJ07, ZR08, ZB00]. juveniles [ASR05, AM00, BLV06, GMY^{+}09, SG01b, VMSR08].

K. [McI05d]. kaakan [AHAAD01]. Kaiser [McI03a]. Kaiwarinus [YFT^{+}02]. kakaan [AHABAA^{+}02]. Kalman [LNS08]. Kamchatka [Oja06]. kamtschatkana [ZLC09]. Kaneohe [WHML04]. Kapoor [McI05a, McI09a, McI09b]. Karasu [TEYA02]. Kariba [Man00, Chi00]. Karl [Cor09, LOT07a]. Karumba [TVB^{+}09]. Karunasagar [Lai00]. Kathirestan [McI06a]. Katsuwonus [AC02, AS04, TSOG01]. Kattegat [CS04, FHM09, KFM08, VBSC08]. kerathurus [KTL06]. keta [SN09]. Key [Bla06, SSG06, BCSA08, BMSJ05, Lor06, Wah03]. Khanna [McI05a]. kHz [GCD^{+}09]. Kiener [Nar06]. kilka [FZHL09]. kill [SMM07]. Kim [KW03b]. King [PNJS06, VL08, BK09, BSG02, CTC^{+}06, DGP02, DA07, FHJL08, GFL03, GFS03, IS07, MCB^{+}07, NS06, PPH01, Sid03, XM00, Xia04a, BLH04]. Kingclip [PGF07]. kingfish [CMAOAA05, GAOMC06, MCAO^{+}05]. Kino [LMASMASZ02]. Kishinouye [Xia04a]. kisutch [QDV05]. Kitada [McI05b]. kite [MHMP01]. Klisova [HAO00]. know [Cer00]. Knowledge [Sym08c, BHB^{+}04, GJ09, GB07, HMO01, HC07, Moo03, SB05]. known [Dou04]. known-age [Dou04]. Ko [LD06]. kob [Kir01]. Kodiak [BK09]. Korea [COHP02, JCJ^{+}09]. Korean [JHK08, Jun08, PMAK07, ZKG^{+}09]. Kriging [MF09]. krill [BFB^{+}03, HD00]. Krøyer [BCM08b]. Kruse [Ore03]. Kuroshio [AM00]. kuruma [HK06b, TY02]. Kutahya [Yil02]. Kuwait [ABCB^{+}07, AHABAA^{+}02, BYA^{+}08, CAHAF07, YAM00, YAHAB01]. Kuwaiti [AHAAD01]. KwaZulu [CARA^{+}07, RMBG02]. KwaZulu-Natal [CARA^{+}07, RMBG02]. Kyrönjoki [HK02]. Kyushu [NT05, IMSS06, IOS07, MOOH00].

L [Bar00, Ell00, Loc02, McI05b, McI09c]. L. [ADP^{+}08, ACE^{+}08, AMUE01, AC05, AM03, AKC^{+}09, BLV06, BEA^{+}08, BdL05, BPRM00, BP01a, BLR^{+}08, BPD^{+}04, CC01b, CMH01, CK02, DS07, DKGC00, Eer04, FSB^{+}02, FUM02, FAL^{+}01, FBJM00, FML^{+}07, GM06, GM07, HBM^{+}00, HH01, Hor00, Ism03, JLHO01, Jok02, JIDM02, JJKN^{+}03, KN01, KM08c, LPDA01, LQD^{+}08, LR00a, LPL^{+}06, MCM^{+}08, MNML00, MWJ04, MLC03, McP02, MBR06, MR05, MTU^{+}04, MB04, NCW00b, NCW00a, OMF^{+}03, PPL04, PP03, PMT06, PBM06, PKP04, PRdPG07, PBR05, PVP04, RCA08, SSGM03b, SF05, SG06, SLE09, TSM03, UHN07, VBSC08, WF07, Yil02]. Labeo [BP06, KK09]. Labeobarbus [dGvZM^{+}06]. Laboratory [ARLB07, FBL^{+}07, SK08b, GG04, ODR00, PB04, RO00, SFCK03, PDV05a]. Labrador [LK00, MBK02, MD09]. labrax [FML^{+}07, PPS05, PBLP08, PKP04]. Labrus [FMB^{+}05]. Lacépède [CMAOAA05, GAFA05b, LDE^{+}02, ZR08]. Lack [SB07]. laevis [GHDM08]. Lagodon [EF02]. Lagoon [PLD05, PLMDB06, VD06, CC01a, GWR07, HAO00, LBABH09, MDJR05, PCD03, PSPCVCD09, Ano00d]. lagoons [LS00]. Lagrangian [PAR^{+}03, PNJS06]. Lake [CND02, NJL^{+}00, NWML00, RVS02, Yil02, ESB09, HH08, JJWO09, JRML00, Kat05, KHM^{+}09, NHSD08, OMR09, PCS05b, Chi00, GJ09, IWBJ08, LCD^{+}02, LDE^{+}02, Man00, NHSD08, ODdGV02, SRTP06, WG05, WBMM05, WIJB08, YSS^{+}09, YLCS^{+}09, dGvZM^{+}06, vZRMvD02]. Lakes [Bar00, CdML^{+}07, DVAL^{+}09, GDDJ09, GRC^{+}09, JJWO09, JAKM05, KS02, LLJ^{+}08, UMV^{+}09]. lalandii [BBJG04, BGB06, DGS^{+}05, HBBG06, MGNB05]. Lamarck [OS05]. lamprey [BTC08a]. lanceolatus [LLJC09]. lancetfish [PMC^{+}07]. landed [Sca03]. Landing [GDFH07, BCSP03, TSTK08]. landings [BCC09a, BAEF06, FCAD^{+}08, GKV06, GEYPC^{+}09, LP06, LWM07, MAND05, TC09, VYMGM^{+}09, WC07]. landlocked [JRML00]. landscape [ABS^{+}03]. landscape-level [ABS^{+}03]. Lane [Lon02a, MR08a]. Lange [McI07b]. Lanice [RBH^{+}08]. Lanka [DHM08, HJ06]. lantern [CE07]. Lao [GAC^{+}06]. Large [BR05, WLDH08, BNS00, CA07, ESB09, GNL^{+}09, GDFH07, GFvdV06, HH08, He07, IMOO06, KME^{+}02, Lon03b, MM06a, MMB07b, Ove02, PBW^{+}09, PMC^{+}07, RCA08, SGP02, SALnM07, SAA^{+}03, SWT09, SMM07, The07, THN^{+}04, UMV^{+}09, WMTJ07, Wes02, vM03, Ste05]. Large-scale [WLDH08, CA07, IMOO06, The07]. Largemouth [OSC^{+}05, CS08c, CSSO03, LDE^{+}02, WSC08]. larger [MvDM02]. larvae [ÅCST01, ACE^{+}08, AMUE01, AC05, AM00, BH01a, CMH01, EBA^{+}03, FBJM00, KMB01, MFC^{+}08, MÁ08, OQE03, PNJS06, PTPS09, RCGC04, RCG02, VMSR08, VKD03]. Larval [ØA08, PAR^{+}03, ARCLBPdP07, CC03b, Dou08, FAL^{+}01, KLJ06, LDGA09, LC06, MBDM05, NMF^{+}08, OR07, PRM04, QGBS04, RLGLO09, RVS02, STI08, Ste03b, UMV^{+}09, WT00]. lascaris [PTR^{+}05]. laser [OSB09]. last [BCC09a]. late [EMKG07, GKUO07]. latent [Mai09]. later [LC06]. Lates [Bal09, Dou04]. Latin [BR00, SCSC07]. latisulcatus [XM00, Xia04a]. Latitudinal [HMM^{+}09, GD03]. Latreille [GYMdL^{+}06]. Latris [TLH07]. lavaretus [KG04, LWW02]. law [Lon02b]. Lawrence [SWM09, BFM08, CWS^{+}05, HMM^{+}01, MMH05, SSC^{+}03]. layout [TSS^{+}04]. lead [SOS02]. leader [LWW02]. leaders [BR05, WLDH08]. leadership [JTdH06]. learned [But08, GAC^{+}06, GML06, KK06, Koe08, KPBD08, LOT07a]. learning [FL05, Fab06, TMP^{+}06]. Lebanese [BAEF06]. Leber [McI05b]. legal [GFvdV06, HMM00b, WD02]. leidyi [Kno07]. Leiostomus [RPC^{+}06, ZR08]. lemon [AA03]. Length [EAT05, LK00, MAL^{+}01, SCLM06, SK03b, Wie04, ATB^{+}02, AC02, Bas99, Bas00, Ber07, BGW06, CA07, Dah00, DMBF08, Dun06, FBS02, GM07, HFS02, HCS^{+}03, HP04, Hor00, How02, Joh03, LH04, MM04, MAdAMASCM09, MMM^{+}03, MGNB05, Ogl09, ØT02, Pál03, PC03, PCD03, RMMP06, SGVM02, ST00, Som03, TSTK08, WSS06, YEA07]. length-based [CA07, Pál03]. length-structured [DMBF08]. lengths [CSR06, Hor00]. Lepeophtheirus [BCM08b]. Lepidion [Mag01]. Lepidocybium [MD02]. Lepidorhombus [LPP02]. Leporinus [AAB05]. leptocephali [KIT^{+}04, WT00]. Lesser [DJG^{+}06, RDH05, RCSFO04, SBD^{+}05, SBD^{+}06, Gob00]. lesser-spotted [RDH05]. Lessons [GAC^{+}06, KK06, MB07, But08, GML06, KPBD08]. lethal [FMC^{+}09]. Lethrinus [PGK03, WMD07, vHAN^{+}09]. Letter [AOLLP^{+}07, AC00, KW03b, Pra03, Sho09, Lon03a]. leucomaenis [iTM04, YMT^{+}06]. level [ABS^{+}03, BCC09a, DOQ08, IWW03, JDS00, Lon09, MAND05, MM06c, NWML00, SET^{+}09]. levels [BM09, FHM09, GKH03, VGBF02]. Lévi [Ano00d]. Libya [MSM^{+}08]. lice [BNSA05, BNS07, Fra07]. licence [MSTC04]. Lichtenstein [Bjø01]. Life [AGS^{+}08, ABFV02, AR01, DS07, FR08, HT05, Lon02a, MRT03, RRAS03, BTC08a, CCMS^{+}03, CC03c, DKKS07, FF01, GEH^{+}07, HGF06, HLH08, IWW03, KAN00, KIT^{+}04, LD08, MAASMB05, PTPS09, PCS05b, SR00, SJ05, Vil01, ZB00, Ano00c]. life-history [GEH^{+}07, KAN00]. LIFT [HM00]. liftnet [vOPvDM02]. light [CWS^{+}05, HHTE05, LNS08, MSVF05, ODR00, RO00, VMSR08, WYR06]. light-fishery [WYR06]. light-sticks [HHTE05]. lighter [Koe03]. lights [Cad02, vOPvDM02]. lightweight [BMW07]. Ligurian [FSB^{+}02, RMF^{+}06]. like [Hat00]. likelihood [BQ06, Tal09]. Limfjord [PHM07]. Limit [Cad02, MM04, Wil09]. limitations [ASM^{+}03, BACP09]. limited [ROS04, SE06b]. limits [BR01, CdML^{+}07]. Limnothrissa [Chi00, Man00]. limpet [AGB07]. line [BNL^{+}08, FMC^{+}09, HJ03, Løk03, MLP^{+}08, PEB03, SJ05, SMB^{+}08, iTMI06, WMB08]. line-caught [SMB^{+}08]. Linear [CGEHC07, LH04, BVW04, BS04, Cam04, CJ00, Dic04, GESY^{+}07, HPM04, JC04, MBT08, MDS03, NC04, OSG08, RSMZ06, VD04a, Xia04a, Xia04b, YAHAB01]. lineata [TLH07]. linefish [HGSP01, Kir01]. ling [CS08a, GOAARJRP09, PCHM07]. Linkage [MPQ07, CLPP05, GWR07, SW09]. linked [RdFBS01]. Linking [BGP^{+}04]. links [MLH^{+}05]. Linnaeus [Aló09, DALP01, EKU02, HHTE05, IYI07, LSH04, LCD^{+}02, MRCH04, MAG^{+}09, Pou01, PTE09, Sch09, TSOG01, TT01, XK04]. Linné [SSML02]. Liocarcinus [RMA^{+}05]. lion [AFJC08]. lions [SPS^{+}07]. lip [DACS08]. lip-gripping [DACS08]. lipid [CCA^{+}03]. List [Ano09s]. literature [Ste07]. Lithognathus [Arc03, FSB^{+}02, PA02]. Litopenaeus [DOQ08, DKKS07, LMRQNM^{+}05]. littoral [MMM^{+}03, TSM03]. live [HM00]. lived [BM00, DDP^{+}06a, DDP^{+}06b, DDP^{+}06c, Lon02a, MMB07b]. lividus [ASM^{+}09, GCB^{+}06, GTB04]. Living [McI05d, CMM^{+}07, HFS02]. Liza [FGJD08, HAO00]. lizardfish [YST^{+}02]. loading [STK^{+}05, TSS^{+}04]. Lobos [MC04]. Lobster [WVW^{+}08, ARCLBPdP07, BEA^{+}08, BRL^{+}05, BMF^{+}10, BM09, BHH08, Bra04, BBJG04, BSWL06a, BGB06, But03, CFE02, CCMS^{+}03, Daw08, DGS^{+}05, Ehr08, Fri07, FHI07, GFMO06, GQR03, GYMdL^{+}06, Gro00, GBG^{+}03, HBBG06, IHF08, LPHH09, LCTB^{+}03, MB07, MSTC04, Mil03a, PVMdL05, RLSJIA07, Row02, RLGLO09, SCW04, TB04, TSP^{+}04, Vel03, WM09, YPDS05, ZFJ03]. lobsters [BSWL06b, CCD01, FH03b, GMS06, HSGF06, PG06, PHGT06, ST03, Ste03c, Wah03]. Local [HC07, SB05, ADP^{+}08, RG07, WIS03, WPS^{+}07]. location [AAP^{+}09, GCJ07, MWB^{+}08, PBR05, SC08]. loci [WF07]. lock [MK02]. locus [MGR^{+}02, SF05]. logbook [SM04a, WKM02, WIKM05]. logbooks [MTA^{+}02]. Loggerhead [BM08]. loggers [GNL^{+}09, HP00, MTA^{+}02, TMK^{+}06]. logic [LZS01]. logistic [Pra02]. lognormal [Dic04]. loliginid [GKV06]. Loliginidae [Hat00, PdAO02, RW07]. Loligo [AM02, BPW01, CPW^{+}06, GB06, Hat00, PdAO02, PB03, RB00, SLA00, ŠV04, Tia09, Vil01, YPD^{+}04]. London [Ano00c]. Long [BD04, JPVKS^{+}01, KVZ05, Lon02a, MD02, NNE^{+}05, Ste03c, TNL07, WY06, Bal09, BM00, BNL^{+}07, FAN03, HTF^{+}09, MMB07b, NJE00, PdAO02, PHM07, vZRMvD02]. long-finned [PdAO02]. Long-lived [Lon02a, BM00, MMB07b]. Long-term [BD04, JPVKS^{+}01, KVZ05, MD02, NNE^{+}05, Ste03c, TNL07, WY06, Bal09, HTF^{+}09, NJE00, PHM07, vZRMvD02]. longevity [Lon02b]. longirostris [CFE02, DBAÖ06, SGB00]. longitudinal [TKFJ09]. longline [BDB^{+}03, BGM^{+}09, BCMM09, BMPK06, BM08, BH01b, CLPP05, DMA07, DKH08, EGB^{+}01, GDFH07, GB07, HdS07, HHTE05, HS00, KG06, LR08, Løk03, NC04, OA04, PGF07, PL04, PL06, SB03, SYS^{+}08, VL09, WK01, WKM02, WIKM05, WMdS09, WM07, WLDH08, WEK^{+}09, WBHG01, YKM06, YKO09]. longliners [LFA08]. longlines [FRF08, Hal02, MD02, RMHW03, SCM09, SE02, SME02]. Longlining [PCMC08]. longnose [Ger09, MK06]. Lophius [CCM^{+}09, DALP01, GRPLE05, LD08, LPDA01, LQD^{+}08, LJP05, MGS02, PW05, PPdA05, SPB^{+}03, WWG03]. loss [BGB06, KF02, SB03, SSGM03a, WM07]. losses [TSJ04]. lost [AMAOMM04, GFS03, Paw03, RD03, SSS^{+}03, SPB^{+}03, TL03]. Lota [HK02]. lough [KHMS09]. Loughlin [McI02c]. Louisiana [CHN08]. love [Koe08]. Low [COLB07, GKH03, MC05, Cor01, DSK^{+}08, JWR05]. low-frequency [JWR05]. low-value [DSK^{+}08]. Lowe [MF00, PLDS03, PGS^{+}08, RLG^{+}01]. Lowell [Ore03]. lower [MMM^{+}03]. Lowestoft [BGH05]. lowland [GT04]. Ltd [Ell00, McI03a, McI05b, McI07b, Mil02b, Mil03b, Sym07, Sym08a, Van01]. Lucas [Mil02b, SGB00]. lucioperca [Eer04, GWR07, JRML00]. lucius [AKC^{+}09, LCD^{+}02, LPL^{+}06, LPG^{+}09]. lumpsucker [ATB^{+}02]. lumpus [ATB^{+}02]. Lunar [LWM07, RC06, OGPDOM08]. lure [HCH09]. lures [MLJT09]. Lutjanus [AFSF05, ASAGR06, BLB07, Bur02, EF02, GAF06, MM06a, MMB07b, MR08a, NCW00b, NCW00a, New02, vHAN^{+}09].

M [Ano00b, How03b, Loc02, McI02a, McI03a, McI03e, McI03b, McI05b, McI08, Sut02]. M. [WSC08, JRML00]. macarellus [SE06b]. Macaronesian [RLG^{+}01, SÁR^{+}05]. maccoyii [FDG^{+}07, GCC^{+}08, HCS^{+}03, MB08, WH08]. macdonaldi [LC07]. machine [TMP^{+}06]. machine-learning [TMP^{+}06]. machines [BMC07]. mackerel [Aba08, AGS^{+}08, AMC^{+}08b, ACG^{+}01, BR01, BSG02, CSM^{+}09, CBRN08, CMAS08, CMH01, CHSA02, CPR^{+}08, DGP02, FAL^{+}01, GCA^{+}08, GAFA05b, IFH^{+}09, KWK^{+}09, KM08a, KHMS09, KME^{+}02, MCM^{+}08, MFC^{+}08, MAC^{+}06, MÁ08, MB00, MBDM05, MACS08, MPA08, PVP04, RWT01, RdFBS01, SE06b, SMSZ08, SWA08, Tze04, UL01, VAF04, WY06]. macki [SDH00]. macleayi [MMB^{+}07a]. Macquarie [TdlMH^{+}03]. macrocephalus [JCJ^{+}09, PSS^{+}07, vSSK07]. Macrofauna [CRG^{+}02]. macroscopic [GM06]. macrourid [LP00, SG01b]. Macrourus [KN01]. Macruronus [PCAS04]. maculatus [NS04]. Madagascar [JH07, MWF06]. made [AC00, CdML^{+}07, MvdK06]. Madeira [AdBP02, LLM05]. maderensis [LLM05]. maena [DKGC00]. magellanicus [KCRT03, PCAS04, WHWM07]. magnetic [SK08b]. Magnhagen [McI09b]. magnificus [GMS06]. magnitude [RPP04]. Magnuson [DJGC09]. mahseer [BNS00]. mahsena [PGK03]. main [LLV^{+}08, LD01]. Maine [Ove02, Tal09, CH03, GHDM08, GC04, He06, He07, HMRG08, KCH05, Laz08, RH06, SCW04, SABW09, Ste03a]. maintains [FHJL08]. Maireta [SRA^{+}02]. Maisse [Ell00]. Maja [CF07]. major [Bas99, Bas00, EH04b, SKF08, ZK03]. Majorca [RMN08]. Makaira [WIKM05]. makes [GPG^{+}03]. Making [Sym09, KW05, Mil02a]. mako [CMJ05, CL09, RCGMQV05]. malabaricus [NCW00b, New02]. Malacca [ABAA^{+}07]. Málaga [TSM03]. Malawi [WBMM05]. Malaysia [ABAA^{+}07, Bla06]. male [BFM08, BBJG04, GQR03]. Mallotus [BG01]. Malmquist [OGPC09]. Malombe [WBMM05]. mammals [SCP01, dSH01]. man [CdML^{+}07, OM07]. man-made [CdML^{+}07]. manage [PPGD08]. managed [LPG^{+}09, She09]. Management [CHL01, DDP^{+}06a, DDP^{+}06b, DDP^{+}06c, Gra02b, HCHO^{+}03, HIRM06, IFH^{+}09, IPH05, JJJ03, MLP^{+}08, McI02d, McI02f, McI08, Nee08, Ore03, Sym07, Sym08c, Sym09, WBMM05, YJV^{+}03, AGS^{+}08, APM^{+}09, AAP^{+}09, Ara09, AM03, ABF08, AHS^{+}06, BBRS06, Bla06, Bur02, But08, Cad02, Cad09, CTEARD07, CAM^{+}03b, CQJX08, CXQ^{+}09, CdMLGC09, DUR05, DBK07, DDP^{+}08, Eer04, Ein09, FZHL09, FRGF07, FHM09, FSPWG08, GSS^{+}06, GSA^{+}06, GGVLSR03, GFM02, Gav09, GAOMC06, Gri03, GLCC09, GMP05, HPD08, HPD09, HCM05, He08b, HF07, HMR^{+}08, HGSP01, JTdH06, KC06, KPBD08, LHMO09, LZCR08, Lon09, Lor06, MMB07b, MMS09, MB07, Mil03a, MF03, MMZ09, NWML00, NM06, OJ06, PBLP08, PKP04, PWP02, PDH08, RB09, RL03, RDEACT07, SCSC07, Sca03, Sei07, SC08, SALnM07, SMD03, Ste08, TJ05]. management [Tia09, TLH07, Tri05, TTP^{+}09, TSTK08, VSS07, Vel03, VFCC06, WMB08, dSBAC06, Ano08l, Coo01, McI05c, Mil08, Sym08c, Win00, Win02]. management-related [PWP02]. Managing [eJSK09, PMWC03, Rod01, Bat04, SWM09, Lon02a, McI02d]. mandates [WD02]. Mangrove [McI06a, Kru04, MLH^{+}05, MMLT05, McI06a]. manipulation [SB08]. Mano [SKF08]. Manta [JDAM08]. manufactured [PCMC08]. many [Hol09]. Maori [HSGF06]. maorum [HSGF06]. Mapping [SAG05, KCRT03, SHDO04]. marbled [BK02]. March [Ano00m, Ano01o, Ano02q, Ano03v, Ano04r, Ano05z, Ano06-28, Ano08u, Ano09x, Bla06]. Margaritifera [Has06]. marginal [Ker00]. mariculture [MKGH05]. Marie [Ore03]. Marine [Ano01e, Ano09a, BPH^{+}06, How03a, KNSSKF09, Lon02a, MWJ04, McI01, McI02e, McI02f, McI03b, McI05d, McI08, Mey07, PBW^{+}09, QDV05, Ste03a, Ste05, AP05, AAP^{+}09, BBRS06, BMF^{+}10, BP08, BLR^{+}08, BLB07, BHM^{+}06, BG04, CMM^{+}07, CXQ^{+}09, CK02, FFM08, GCN02, GCS^{+}06, GG06, GDFH07, KNJJS06, Kat05, KK06, KHMS09, LHMO09, LZCR08, LZC^{+}08, Lon03b, Lot07b, LD06, MH08, MM06b, MMLR08, Mos00, NGF06, OLN^{+}02, Ove02, RSST07, SCP01, SCLM06, SAA^{+}03, SWT09, SAC^{+}02, SK03b, SBD^{+}05, SBD^{+}06, VG01a, Vel03, VSK05, WM09, WMTJ07, WM04, WRK06a, WRK06b, WKCC06, WHML04, dSH01, McI02e]. marinus [BTC08a, BHJ02, SKKP05, SBD^{+}05, SBD^{+}06]. mark [APP^{+}07, BEBA03, BG04, FML^{+}07, LRBH07, UHN07]. mark-recapture [APP^{+}07, BEBA03, BG04, FML^{+}07, LRBH07, UHN07]. markers [KM08a, LM09, MGJB09, PD08, vHMAO^{+}06]. market [BCC09a, Bra04, JTdH06, MPQ07]. markets [SLV^{+}08]. marking [LZZN09, McP02]. Markov [JHSG04]. marks [CMOM00, YST^{+}02]. marlin [ACMMCEGM02, AB09b, OGPDOM08, WIKM05]. MARMAP [JDAM08]. Marmara [DG08, TOÖD06]. Martin [Mil02b]. Martinique [DJG^{+}06]. marulius [KK09]. Maryland [McI02e, FM04]. masked [FHG03]. mass [BGW06, GC03b, MG03a, MG03b, WM09]. Masturus [LLJC09]. matching [HK02]. material [DBAÖ06, KSS05, TÖT04]. materials [TH06]. Maternal [VN00, MPM^{+}06, MK09, VBSK00]. mathematical [STH03]. Matías [Gon06, GNC07, NS04, Nar06, NGF07]. matrix [GOAARJRP09, SW09]. matter [HBW00]. matters [FSPWG08, FWG15, Sve14]. maturation [MQVSN04, ŠV04, TVY01, WLL08]. mature [HMH^{+}04]. Maturity [ANS^{+}01, EH04b, AB09b, BP01a, CCB02, Cro01, GM06, GBC^{+}03a, GBC^{+}03b, LPHH09, LCTB^{+}03, LMRQNM^{+}05, MBR06, NS07, SGS09, WY06, XK04, YTK06]. Maunder [Pra03]. Mauritania [tHDC06]. mawsoni [Hor02]. maxima [SSML02]. Maximum [BQ06, Jen02, PWB05, HCS^{+}03, MM06c, PCS05b, Tal09, YEA07]. maximus [NJJ06, PM09]. May [Ano00k, Ano02s, Ano04u, Ano05x, Ano06x, Ano07u, Ano08o, Ano09-27, DO06, ROS04]. Mayan [AIC08]. McClanahan [Sym07]. McConnell [McI03e]. McKinley [MR08b]. MD [Lon02a, McI02a]. Mead [Cur00]. mean [FGOF08, MAND05, PEAAdLJB07, SOS02]. means [AGB07, CQJX08, FP02, GHM04, KFM08, TMP^{+}06]. measure [ABRB00, AC00, CAM^{+}03b, HP00, MC00b, OSB09, SB03]. Measurement [BPW^{+}03, NP08, Dun06, FBLO07, WSS06]. Measurements [DBB00, WW06, FAN03, GCD^{+}09, HCS^{+}03, KFO^{+}04, KHMS09, Ogl09, vS04]. measures [BCLP00, DDP^{+}08, ERG09, Løk03, PPS05, TTP^{+}09]. Measuring [CATH04, HJ03, SW09, VSK03, JWR05, LBKG08, Som03]. mechanical [DACS08]. mechanically [JGB^{+}06]. mechanism [WT00]. mechanisms [BWM07]. mechanized [LR08]. media [Sym05b]. Mediterranean [ASR05, ADP^{+}08, ASM^{+}09, BLV06, BK02, CCM^{+}09, CUT06, CLGR^{+}04, GCB^{+}06, GTB04, LZCR08, MDS03, MBG^{+}07, MMN00a, PBDP00, RBD02, RMN08, VYMGM^{+}09, AMC^{+}08b, ASA04, Arc03, BSS06, BF00a, BAEF06, BCSA08, BSR05, CA03, CAM^{+}03a, CAM^{+}08, CMAS08, DCM^{+}05, DMC^{+}09, DMA07, GSMT04, GRPLE05, GR00, Ism03, IYI07, KR00, Kat05, KM08c, LLM05, LD08, LZC^{+}08, MVV^{+}01, MMV^{+}04, MCL04, MT08, MLC03, MNM00, MN01, MMM^{+}03, OQE03, PA02, PPL04, PKT^{+}02, RMCdS06, SLB07, SLPF08, SB03, SDM04, SSB03, SCM^{+}08, SMSK03, SMSZ08, TASF09, TMP^{+}06, TTP^{+}09, TOÖD06, TSTK06, TSTK08]. mediterraneus [KR00]. medium [IV08]. medium-sized [IV08]. Meeting [Ano09a]. megabenthos [WDC02]. Megafauna [RM07, ZCdG06]. megalops [BGW06, HFB^{+}06]. megapixel [RGSB08]. megrim [LPP02]. Melanogrammus [AFVSR^{+}09, GM07, Hal02, HMH^{+}04, LR08, ÖW02]. melanostictus [SAK05, SWA08]. Melbourne [Kes00]. Melicertus [IS07, KTL06, MCB^{+}07, MSB07]. Mendoza [Gue05]. menhaden [RT07, VSS07]. Mente [McI09d]. mentella [JDMN00, SKKP05]. meristic [TOÖD06]. Merlangius [GM06, GM07]. merlangus [GM06, GM07]. Merlucciidae [BBL07, TLMN00]. Merluccius [ASR05, AMUE01, AC05, BCSA08, BLV06, Bet04, BSR05, BBL07, FAL^{+}01, GNC07, HC00, Joh03, LC06, MPE04, MPM^{+}06, MLC03, MNBM05, OB02, OQE03, PMM05, PRdPG07, Pou01, RCA^{+}07, SSGM03b, TLMN00, VGBF02, WPS^{+}07]. Mesh [SF03, BSS06, BCSP03, BKD^{+}02, BMK^{+}04, BNL^{+}07, CFE02, CFE03, CFH03, CF04, CRD06, DVAL^{+}09, EH04a, FBL^{+}07, FBLO07, Gas03, GK01, GKF03, GOF^{+}04, GRC^{+}09, HC00, He07, HSSR06, Her05a, Her05b, HO05, HPK06, HO06, JIGI06, KFM08, MMB^{+}07a, MvDM02, OKF06, RCD07, SLPF08, SGB00, TASF09, TSJ04, VMSR08, vM03]. mesh-size [HSSR06]. meshes [GFSM07, GSL08, HPK07]. Mesoamerican [Sei07]. Mesolongi [LS00]. Messolonghi [HAO00]. meta [HKG^{+}08]. meta-population [HKG^{+}08]. metabolic [WH08]. metal [KS08, SK08b]. metamorphosis [WT00]. Metapenaeopsis [TCY01]. Metapenaeus [MMB^{+}07a].

Metapopulation [SW02]. metapopulations [Sei07, SC08]. Metazoan [OB02, OS05, OFOM08]. method [AFVSR^{+}09, AHS^{+}06, BBK08, CSDP02, CXMM04, CAH^{+}03, HFHO07, HR09, IV08, JCW05, KM02, MK01, Mes03, Pun03, RYSN08, SHL09, SKMT02, UMK01, WSPY05, WSPY07, WH08, YLYL09, ZG08]. methodological [RMA^{+}05]. methodologies [FBLO07]. Methodology [HKF^{+}09, RFWH09]. Methods [McI07b, SA03, AKA^{+}01, BPRM00, BG04, CMPM01, CS07a, CBGP06, CdML^{+}07, CA07, CF07, Dou04, DKH08, ED07, EF01, GMR01, GMR02, HdS07, JSR04, JJWO09, KM08c, LLJ^{+}08, LB08, MFB08, MB08, RHF^{+}08, RW03, San00, SHKB09, USKK04, YPD^{+}04, vT04]. Methusalas [Lon02a]. métier [TPM03]. métiers [SMS^{+}06, TSTK06]. metrics [DVAL^{+}09, LMR^{+}05, LMR^{+}07, Ste06]. metropolitan [ABF08]. Mexican [BSMFH07, MROGVR09, PEAAdLJB06]. Mexico [AAFLP^{+}07, ACMMCEGM02, All07, AIC08, BSMF^{+}09, BSG02, BD04, CHN08, CTEARD07, CPCD05, DGP02, FOGKT08, HS03, LMASMASZ02, LMRQNM^{+}05, MQVSN04, MR08a, MBCMNMHH01, NMHHMB^{+}00, NMMTCV^{+}06, PCD03, RMVJGMMV07, RRAS03, RRASLB03, RCGMQV05, RDEACT07, RAHHGM^{+}02, SFCK03, VSS07, Vel03, WBCP08, BRL^{+}05, NE07, RBHP03]. Michigan [IWBJ08, WIJB08]. micro [SALnM07]. micro-management [SALnM07]. microchemistry [MRT03, VP05, ZSP03]. microlepis [SFCK03]. Micromesistius [ÅCST01, ASD09, BH01a, FAL^{+}01, PKK^{+}06]. Micropogonias [ABL02, NV05, VH06]. Micropterus [CSSO03, LDE^{+}02, WSC08]. Microsatellite [BSG02, LPL^{+}06, KM08a, vHMAO^{+}06]. microstructure [LD08, MÁ08, MNBM05, PRdPG08, TP00, ZB00]. mid [FB06, USKK04, HG01]. mid-Atlantic [FB06, HG01]. mid-water [USKK04]. midae [RB09]. middle [TGS^{+}00]. migrating [BLKR09, CK02]. Migration [Jok02, SLE09, UL01, WR05, BPD^{+}04, CWL00, CXMM04, CLH^{+}08, DSO^{+}05, Hor05, JJKN^{+}03, KKR^{+}05, KG04, Kru04, Lan04, MNML00, NV05, TNL07, WT00, ZSP03, Mil02b]. migrations [MACS08]. Migratory [GM00, ASD09, PMWC03, VBSK00]. Miguel [BP01b]. Miller [RVS02]. Mills [Sol05]. Milne [He05, PGMM01]. Milne-Edwards [He05, PGMM01]. Milner [Mil08]. mince [AAT08]. miniatus [WMD07, vHAN^{+}09]. minimal [Cad02]. minimum [HMM00b, Lon09, MM04, TSTK08]. minisatellite [MGR^{+}02]. minutus [MGR^{+}02, MGR^{+}02]. miodon [Chi00, Man00]. Mira [CHQ05]. misapplication [PCHM08]. Mismanagement [CS08b, Hol01, Sch01]. Misund [Cur00]. mitigation [BM08, Løk03, SB03, TF09]. mitochondrial [CHL^{+}06, CHW^{+}08, CMAS08, PMSD06, RMVJGMMV07, WCC^{+}09]. mixed [AAP^{+}09, BVW04, HPM04, JSR04, LH04, LP06, MC05, QPR08]. mixed-effects [LH04]. mixed-species [AAP^{+}09, JSR04]. mm [BSS06, GOF^{+}04, HC00, SGL08]. Mnemiopsis [Kno07]. mobile [MH08]. mobulid [WGDP06]. Model [Kat06, Bas99, Bas00, Bro05b, CC01b, CHSA02, DPD^{+}03, DMBF08, FBS02, GHM04, GG06, GYMdL^{+}06, HKG^{+}08, HPM04, HH06, Hor05, IS07, Jen00, JC04, JMGC06, KKR^{+}05, Lab05, Lor06, MBT08, Mai09, MN03, Mau03, ML04a, MLCGRV07, MC04, MG03a, MG03b, NC04, PAR^{+}03, PP03, PNJS06, PF00, PCD03, PL06, PAF^{+}08, PCS05a, QDH09, RRAS03, SdPFAL09, SHBÅ07, SWA08, STH03, Tal09, TdlMH^{+}03, ULDG02, UMV^{+}09, VD04a, WM09, WK01, WKM02, WIKM05, WMdS09, WB08, Xia04b, YPM^{+}06, ZC02, Zho07, Her05a]. Modeled [AB09a]. Modeling [CXQ^{+}09, MLCGRV07, NHSD08, PL04, RLGLO09, BAH09, BD04, FCC^{+}09, LZS01, ZP05]. Modelling [BCSA08, BPW01, CS04, Kat06, MBM04, PHGT06, RPFR06, SMD04, ÅCST01, Cad03, CQJX08, FFL09a, Gri03, HH01, KW05, OF03, PB03, PVMdL05, PK06, RMA^{+}05, SE06b, TVB^{+}09, TSJ04, YLD^{+}05]. Models [XPMQ04, BPW01, BVW04, BS04, But03, Cam04, CA03, CAHAF07, DBK07, Dic04, Dun09, FH03b, GHS04, HJ06, HMH^{+}04, JC04, KMELC05, LH04, LYKH06, MS03, ML04a, MDS03, MOK^{+}05, MMZ09, MM06c, OSG08, OA04, Pra02, PHGT06, Sho09, TH03, TÁTWD09, VFCC06, XM00, Xia04a, Xia06, YAHAB01, ZLC09, Zho07, Coo00a]. moderate [SSH^{+}06]. moderately [WM04]. Modern [CGT08, GG04]. modification [RH04b, SSK^{+}06]. Modifications [AR05, FP02]. Modified [GYMdL^{+}06, FCL^{+}05a, FCL^{+}05b, Gri06]. modulates [ZCTA^{+}08]. modulators [PTPS09]. Mokness [She02]. mola [LLJC09]. Molecular [Ano00b, SÁR^{+}05]. mollusc [CC01a]. Mollusca [CC01b, FRGF07]. mollusk [NGF07]. molt [Ehr08]. Molva [PCHM07]. Monacanthidae [MRCH04]. monitored [BDB^{+}03, SG06]. Monitoring [CBGP06, LD06, McI02e, CWS^{+}05, CWJS09, HBKJ07, KFO^{+}04, MD09, NJE00, QPR08, RFWH09, SET^{+}09, Smi03, SMJS01]. monk [Güç08]. monkfish [LJP05, MGS02, PW05, PPdA05, SPB^{+}03]. monofilament [AAA^{+}06, SE02, TH06]. Monthly [CBC02, GMV00, GMV01, GESY^{+}07]. moored [Dem04, DJG^{+}06]. Moray [YPS^{+}06]. mordax [SRTP06]. morhua [AFVSR^{+}09, BP01a, BG01, BLR^{+}08, BC04, CAJ00a, CAJ00b, CS04, EMKG07, GFL03, Hal02, HMH^{+}04, HBM^{+}00, HKF^{+}09, HH01, HMRG08, JSB04, JIDM02, LR08, LR00a, LK00, LR00b, MR05, ØA08, PP03, PBM06, PCHM07, RL03, SF05, SCM09, SLJ05, SSC^{+}03, Tal09, VN00, VBSC08, VKD03, WF07, WPOB09, WR05]. morid [Mag01]. Moridae [FB06]. morio [BD04, GHCPPLC^{+}05, HS03, LCFP^{+}08]. mormyrus [Arc03, FSB^{+}02, PA02]. Mornington [TVB^{+}09]. Moroccan [EBV^{+}03, KMF08]. Morone [JWR05, MRT03]. Morphological [PA02, Tze04, Ber07, HKF^{+}09, PTR^{+}05, SMRH04]. Morphology [Ano01c, KR00, LLV^{+}08, MTC06, MMN00b]. Morphometric [TSP^{+}04, TOÖD06, TCY01]. morphometrics [SDS09]. Morris [Bar00]. Mortality [Aló09, GC03a, JS05, JRML00, AHABAA^{+}02, AC05, ASAGR06, AWY02, BCSP03, BM03a, BS06, BMF^{+}10, BC04, BG05, BD04, Bur02, CCG^{+}02, CSSO03, DO06, DKGC00, EKU02, FH03b, GHDM08, GHCPPLC^{+}05, GFS03, GBC^{+}03a, GBC^{+}03b, GC03b, GAF06, HMM00a, Has06, Hen09, HH06, JCS01, JHK08, JCJ^{+}09, LDGA09, LRBH07, LS00, Lor06, MF07, MAASMB05, MK01, MAC^{+}06, MR08a, MBGW02, MMLT05, MG03a, MG03b, NMMTCV^{+}06, NCW00b, NCW00a, New02, PP03, PCD03, PGF07, PJR^{+}05, RPP04, RRAS03, RCG02, SE06a, STI08, ST00, Ste08, SAK05, SKF08, SLJ05, TT01, WHWM07, WL06b, WMD07, YLYL09, Zhe03, vMBGF05]. mortem [CAH^{+}03]. Morton [BCM08a]. morwong [SH09]. mosaicus [PK03b]. mother [PAR^{+}03]. motion [GNL^{+}09, OMW^{+}03]. motion-sensitive [GNL^{+}09]. motivation [SOH06]. moulting [Cad03, NS06]. Mountain [CND02]. mounted [KI04, KI05]. mouth [MNM00]. Movement [BFM08, LRW09, UHN07, BLC08, BHH08, PK06, SMB^{+}08]. Movements [LQD^{+}08, LJP05, CMPM01, JHSG04, LR00a, RG07]. Mozambique [dSBAC06]. MPA [GCB^{+}06]. mud [JPYW08, MMLT05, OIK^{+}06]. mudskipper [EKU02, NT05]. Mugil [KKR^{+}05, LRW09, SR07]. Mugilidae [HAO00]. mullet [FGJD08, KKR^{+}05, LRW09, MTU^{+}04, SR07]. Mulligan [AC00]. Mullus [FSB^{+}02, MTU^{+}04]. Multi [MBO^{+}07, RR03b, BMC07, BD04, CRSM04, DVAL^{+}09, FKW06, FFL09b, FH03b, Gra02b, GJBY05, GRC^{+}09, HWB06, HBW00, HPM04, IHF08, Kat06, KMK09, MLP^{+}08, MBG^{+}07, MC00a, MM06c, NS08, SLPF08, TH03, TPM03, ULDG02, dSBAC06]. multi-beam [MC00a]. multi-class [BMC07]. multi-compartment [KMK09]. multi-component [MBG^{+}07]. Multi-decadal [MBO^{+}07]. multi-fleet [BD04, FFL09b, ULDG02]. multi-gear [NS08]. multi-mesh [DVAL^{+}09, GRC^{+}09]. multi-métier [TPM03]. multi-model [Kat06]. multi-purpose [TPM03]. multi-sector [MLP^{+}08]. multi-species [FKW06, Gra02b, GJBY05, HWB06, HBW00, MM06c, NS08, SLPF08, TH03, ULDG02, dSBAC06]. multi-vessel [CRSM04, HPM04]. Multi-year [RR03b, FH03b, IHF08]. multibeam [PBVG07]. multidens [OLN^{+}02]. Multidimensional [MPG^{+}08]. Multidisciplinary [TP06]. MULTIFAN [Lab05]. MULTIFAN-CL [Lab05]. multifilament [AAA^{+}06, TH06]. multimesh [OKPR04, PKŘ^{+}09]. Multiple [KWK^{+}09, SRH^{+}02, APP^{+}07, DDP^{+}06a, IS07, KA06, Kop06, KM08c, ML04a, YEA09]. multiprobe [HP00]. multispecies [DA07, FCML05, Gob00, KP07, SM04a]. Multistage [AKA^{+}02]. Multivariate [TY02, MDS03, Smi03, SMS^{+}06]. Mumbai [JCS01, PJR^{+}05]. Mundy [McI06b]. Munida [CS07a]. Muñoz [YPM^{+}06]. Munro [Sym08a]. Munshi [Ano01c]. Muricidae [GCN02]. Murphy [Lon02b]. murphyi [ACG^{+}01, CSM^{+}09, CPR^{+}08]. Murray [ABD^{+}09, Mat07]. Muscat [AMAOMM04]. Musick [Lon02a]. mussel [ABAA^{+}07, Has06, RGGB07]. mutton [Bur02]. Mycteroperca [SFCK03]. Myliobatiformes [WGDP06]. myriaster [KIT^{+}04]. mystex [JS02]. myth [Lon06]. Mytilus [RGGB07]. Myxosporean [MKGH05].

N [Coo00a, McI02a, Mil06, Mil08, Sym08c, EBV^{+}03, HG01]. NAFO [LK00]. nagroor [AHAAD01, AHABAA^{+}02]. NaLA [STK^{+}05, TSS^{+}04]. namaycush [PCS05b]. Namibia [VGBF02]. Namibian [GMV01, DBK07, GMV00, Joh03, JI07, Kir01, KM02]. NAO [HTF^{+}09]. narrow [GAFA05b]. narrow-barred [GAFA05b]. nascent [BTC08b]. Nassau [ED07]. natal [HKG^{+}08, LRC^{+}08, VP05, CARA^{+}07, RMBG02]. National [Coo00b, McI02e, McI02f]. Natural [How03a, MAASMB05, MDJR05, AAP^{+}09, DVAL^{+}09, FH03b, He03, MR08a, MEA^{+}03, New02, PB04, RRAS03, SSML02, TH06, TGS^{+}00, WL06b, YLYL09, Zhe03]. nature [OM07]. navigation [MK02]. neapolitana [CHQ05]. near [AMAOMM04, HW00, RSM05, Tre01]. near-surface [Tre01]. nearshore [APS07, Bat04, HK06a]. necessary [HCM05]. need [KMF08]. needs [KHM^{+}09, Vas03]. negative [MLCGRV07, OF03, PL06, RB00, SGRK09]. Negombo [HJ06]. Negro [VLD^{+}00, CS08c]. neighbouring [GM07, LDA01]. Neil [BNS07]. Neis [Sym08c]. neither [HCM05]. Nekton [TK00]. Nemadactylus [SH09]. Nematoda [MFC^{+}08]. Nemipteridae [WMTJ07]. Neolithodes [He05]. neon [CC03c, CCT^{+}08, IMOO06, YM00]. Nephropidae [ASA04]. Nephrops [ASA04, BEA^{+}08, BWMA02, CFE02, CAM^{+}03b, CFG05, CRD06, CTK^{+}07, FH03a, FHM09, GF06, HTF^{+}09, KFM08, LMMD06, MGTB08, MBR06, MAG^{+}09, RDH06, RCD07, VU08]. Neretva [TGS^{+}00]. nerka [LRC^{+}08]. net [BCC09b, BCSP03, BTC09, BP06, Bro05b, FSB^{+}02, FMCS05, Gra02b, HFL08, JRG08, KSS05, LLJ^{+}08, LWW02, MNML00, MBGW02, MS02, NW03, ODR00, PMAK07, PCV04, PLD05, PLMDB06, RO00, SSGM03a, STK^{+}05, SSIE06, SME02, SMS^{+}06, SSK^{+}06, TS06, TSS^{+}04, TH06, VSK03, WK08, vSS05]. net-pen [NW03]. Netbiblo [Pie09]. Netherlands [vM03]. nets [AOMK^{+}04, BP06, BMSJ05, EGB^{+}06, GGA08, GSH^{+}07, GLK00, GCL03, HWM02, KJEF03, MAG00, MP03, RH04a, RCA^{+}07, RM07, SSS^{+}03, SPB^{+}03, SSGM03b, STK^{+}05, SE02, TEG03, TL03, VAYD04]. netters [FK03]. netting [GLK00, KSS05, PAF^{+}08, SLB04, WFJB09]. network [GEYPC^{+}09, SD01]. networks [MMLR08, SBR^{+}08]. Neural [SBR^{+}08, GEYPC^{+}09, SD01]. neustonic [JDAM08]. neutral [WF07]. Newell [Ano01d]. Newfoundland [AD00, BLR^{+}08, JSB04, LR00a, MR05, MBK02, MD09, Row02, WNRd00]. News [Cur00, McI02d, She02, Win02]. Nguru [NJL^{+}00, NWML00]. NH [Ano01e, McI09a, McI09b, McI09d]. Nicaragua [Daw08]. Nichols [ACG^{+}01, CSM^{+}09]. Nicoya [FW06]. niger [LBABH09]. Nigeria [ASIW05, EKU02, Mos00, MUO02, NJL^{+}00, NWML00]. night [AB09a, GCL03, SB03, SDNL09]. nigricans [WIKM05]. nigrodigitatus [ODdGV02]. Nile [dGvZM^{+}06]. niloticus [PAP04]. nimbaria [TP00]. nimis [FWG15, Sve14]. Ningaloo [NS07, Tay07]. Niño [Mar06a, Mar06b]. nitidus [EL09]. No [Coo00a, Gra02a, McI05d, FFM08, PK06, Row02]. no-take [FFM08, PK06, Row02]. nobilis [APS07]. nocturnal [BYA^{+}08, MC06, Tay07]. noise [ET00]. Non [GESY^{+}07, McI03a, ABL02, BCMM09, BP06, CA03, Cor02, CGEHC07, DTSR07, DFD02, GOAARJRP09, HCM05, Hol01, HGSP01, KAO04, KG06, MHK^{+}06, OA04, OSC^{+}05, Sch01, TB04, VL09, WF07, vMBGF05]. non-commercial [HCM05]. non-compliance [TB04]. non-cooperation [DTSR07]. non-cooperative [HGSP01]. non-destructive [BP06]. non-equilibrium [CA03]. non-falsifiable [Cor02]. non-filter [ABL02]. non-harvest [GOAARJRP09]. Non-linear [GESY^{+}07, CGEHC07]. non-neutral [WF07]. non-offset [OSC^{+}05]. non-precautionary [Hol01, Sch01]. non-sampling [DFD02]. Non-target [McI03a, BCMM09, KAO04, KG06, MHK^{+}06, OA04, VL09, vMBGF05]. noncompliance [Wil09]. nor [HCM05]. Nordic [GL05, Gri06]. Nordmore [RH06, BKD^{+}02, FCL^{+}05a, FCML05, FCL^{+}05b, Gra03, Gri06]. Nordmøre-grid [Gri06, BKD^{+}02]. Noronha [GCS^{+}06]. Norte [NE07]. North [CS04, CL04, How06, Lot07b, QP00, RDH06, SBD^{+}05, SBD^{+}06, ATM08, Bal09, CP00, CFG05, CCB02, DHM08, FRGF07, GM00, GR00, GESY^{+}07, KI04, KI05, Løk03, MM07, MSA00, MAL^{+}01, New02, PF00, Pra02, VAF04, WDC02, AE03, dSHG08, Ano09a, AM00, BB04, BHJ02, BM08, BGH05, CTK^{+}07, CC03c, Cor01, CCBG02, EH02, ERG09, FH03a, GHDS08, HMO01, HP02, HKG^{+}08, HMH^{+}04, HMM^{+}09, IMOO06, JR02, JIAD03, KVS04, KN01, KC06, KHMG06, LNG04, MKGH05, MH01, MPQ07, MKB01, NDCK09, Nee08, NJL^{+}00, NWML00, OKF06, OGK^{+}08, PNJS06, PSS^{+}07, PH01, Pol00, Pol02, PCHM07, PVP04, RdFBS01, RH04b, RCD07, SC07, SHBC02, SGL08, Sol02, TH03, UM02, VBSC08, WSPY05, WSPY07, WM00, WPOB09, YM00]. north-east [Bal09, CFG05, GM00, KI04, KI05, SGL08]. north-eastern [MSA00, MAL^{+}01]. north-west [FRGF07, WDC02]. north-western [GR00, New02]. northeast [All01, All07, CPW^{+}06, Ger09, HP04, LPP02, LQD^{+}08, MBK02, VAF04, WNRd00, AMC^{+}08b, ABK03, BSC^{+}08, MÁ08, PCB01, SKKP05, SMSZ08, SBF^{+}08, Coo00b]. northeastern [BH01b, HFB^{+}06, JLC04, JCLL08, LJP05, LSH04, MVV^{+}01, MLH^{+}05, PCHM07, SCM^{+}08, TOÖD06, WY06, FFL09a, FFL09b]. Northern [Cor01, CK02, RH06, WSP06, APC07, AKC^{+}09, BLV06, BP01a, BHJ02, BSMFH07, BBV07, BHM^{+}06, CP08b, CSCC07, CARA^{+}07, CPR07, GFL03, Gri03, GWB07, GFvdV06, HBW00, HH06, HW00, Jok02, JJJ03, KSS05, LPL^{+}06, LPG^{+}09, MM07, MKB01, Nar06, OM07, OS05, OLN^{+}02, PCM07, RPP04, RMCdS06, SPB^{+}03, SSB03, SW02, TNFF03, VYMGM^{+}09, WBCP08, WF07, Wie04, ZLC09, vZRMvD02, AKA^{+}01, BQN05, DPD^{+}03, DDP^{+}06a, DDP^{+}06b, DDP^{+}06c, GFS03, SLV^{+}08, VD04a]. Northwest [Mil02a, tHDC06, CP00, EF01, He05, JS02, VAF04, XBS^{+}07, BM03b, BM03c, CL09, CCT^{+}08, FR03, FH03a, HP09a, LRG^{+}00, NW03, Nas03, PRdPG07, PRdPG08, WIS03, ZCdG06, ZCTA^{+}08]. northwestern [CA03, MR08a, SB03, SDM04, STI08, TAS03, TCC^{+}09, TSP^{+}04, TGH^{+}08, HJLL07, WCC^{+}09]. norvegicus [ASA04, BEA^{+}08, CFE02, CFG05, GF06, HTF^{+}09, KFM08, MGTB08, MBR06, MAG^{+}09, RCD07]. Norway [SBD^{+}06, BEA^{+}08, BP01a, CFE02, EH02, GFL03, GFS03, KHMG06, NM06, PNJS06, SBD^{+}05, TNFF03, WF07]. Norwegian [BHJ02, Bjø01, EH04b, FGOF08, HF07, HBM^{+}00, HLHF03, JLHO01, KS02, LR08, PP03, PTPS09, RWT01, SHDO04, SBF^{+}08, VOH02]. note [Ano03-32, Ano04y, Ano09-31]. Notes [HFB^{+}06]. notices [Ano02-27, McI02b]. Notopteridae [SND^{+}08]. Notothenidae [OFOM08]. Nova [BFM08]. novel [BMK^{+}04, IV08, PPGD08]. November [Ano00j, Ano01q, Ano02t, Ano02x, Ano03-30, Ano04o, Ano05y, Ano06w, Ano09-28]. NSW [IS07, LSW06]. Nuclear [IJD04, JIAD03, PD08]. nucleic [CCA^{+}03]. number [HPK07, MWB^{+}08]. Numerical [Pri09]. nursery [PKP04]. nutrient [Kno07]. NW [BLV06, LZCR08, MDS03, MF03, PBDP00, RMN08, BSS06, BSR05, MLC03, OQE03, You03]. NY [How03a, McI09c]. nylon [SLB04, WLDH08].

O [McI07b]. obesus [AdBP02, BDB^{+}03, CHL^{+}06, CHW^{+}08, SYS^{+}08, WMdSB09, WMdS09]. Objective [JSR04, FBL^{+}07]. objectives [MLP^{+}08, SP09]. obligatory [GJ09]. oblique [OR07]. Observation [EBV^{+}03, LB08]. Observations [MvdK06, OJ06, ÅCST01, CXMM04, CAH^{+}03, IWW03, KG04, MAG^{+}09, ODR00, PTE09, SPS^{+}07, SK08a, YEA09]. observe [RSM05]. observed [EGJ00, He03, Sch09, SBGG^{+}05]. observer [WKM02]. obtained [DVAL^{+}09]. obtaining [vT04]. obtusidens [AAB05]. occidentalis [DOQ08]. Occurrence [JH07, Row07, Sva05, MGNB05]. occurring [VKD03]. Ocean [TSOG01, BLC08, COLB07, GHS04, HJ03, HJ07, McI02e, NP08, TJ05, WNRd00, All01, ABK03, AM00, BSG02, CSM^{+}09, CHL01, CL09, CCT^{+}08, CHL^{+}06, CHW^{+}08, DHM08, DGP02, GCS^{+}06, Ger09, JIAD03, KN01, MROGVR09, MD02, MKB01, MGJB09, ORVMP06, PMC^{+}07, Pra02, RWMBHH06, Row07, RG07, SDNL09, STI08, SC07, SYS^{+}08, TCC^{+}09, TP00, WSPY05, WSPY07, WMdSB09, WMdS09, WCC^{+}09, XBS^{+}07, YM00, McI03e]. oceanic [ASD09, CPR^{+}08, GCS^{+}06, JDMN00, LHMO09, TAS03, WPS^{+}07]. oceanic-climate [TAS03]. oceanographic [AR01]. oceanography [AR01, Rei01, San00, McI06b]. OCEANS [McI08]. ocellatus [JDS00]. October [Ano00q, Ano01m, Ano02o, Ano03x, Ano04v, Ano05s, Ano06s, Ano08p, Ano09-30]. octopus [BSWL06a, BSWL06b, CSSO03, HSGF06, LHMO09, LP06, NGF06, FRGF07, GMS06, HSGF06, HGHLCH02, LHMO09, NGF06, OGSG07, VYMGM^{+}09]. Odebrecht [Ano00e]. off [AAFLP^{+}07, ABFV02, ASA04, BCC09b, BAEF06, BH01b, CFE02, CFE03, CFH03, CF04, CC03a, CSCC07, Cro01, CBC02, CRC^{+}07, CPR^{+}08, DOQ08, DG02, EBA^{+}03, EBV^{+}03, FFM01, FP00, FMCS05, FCAD^{+}08, FB06, GJSW04, GDAM07, GFS03, GMS06, HJ06, HFB^{+}06, HGHLCH02, HTF^{+}09, HC09, HK02, JIDM02, KE05, LK00, LSH04, LR00b, MDS03, MCB^{+}07, MBK02, MC04, MF09, NMMTCV^{+}06, New02, PL02, PGS^{+}08, PH09, PW05, PRdPG07, PGMM01, PHGT06, PDH08, QP00, RG07, SB01a, SSGM03a, SSGM03b, STCH09, SHDO04, SGB00, SAG05, SAK05, SWA08, TYM^{+}01, TJ05, TÁTWD09, TCY01, TY02, Tze04, VG01b, Vel03, VAF04, Vil01, VGBF02, VE03, WY06, Wie04, XK04, ZCdG06, ZCTA^{+}08]. officinalis [DR01, RPFR06]. offset [OSC^{+}05]. offshore [LPHH09, LFA08, RCA^{+}07, TOK^{+}05]. Ohtani [McI02c]. oil [RR03b]. Ole [Cur00]. Olifants [CWJS09]. oligotrophic [HH08, WM09]. olivacea [JPYW08, MMLT05]. olivaceus [LZZN09, SKF08]. Oman [AMAOMM04, CMAOAA05, GAOMC06, HMAOAS07, MCAO^{+}05]. OMEGA [FBL^{+}07]. Ommastrephes [BI05, CC03c, CCT^{+}08, IMOO06, YM00]. ommastrephid [AR01, GKV06]. Ommastrephidae [NNA02, BQN05, CUT06, Lap02, MQVSN04, MCP01, NNA01, SN01, TVY01]. Ommatrephes [TCC^{+}09]. Ommer [Ano01d]. on-board [SPS^{+}07]. Oncorhynchus [BCM08b, CB00, Kop06, LRC^{+}08, MK09, QDV05, SN09]. one [Cor02]. Oneida [RVS02]. only [JIGI06, SSH^{+}06]. onset [MBR06]. ontogenetic [CMH01, MSA00]. Ontogenic [MACS08]. oocyte [GSMT04, MPM^{+}06]. open [BPH^{+}06, HIRM06, LFA08, NGF06, NP08, RD03, SN01]. open-access [LFA08, NGF06]. operation [MSB^{+}08, USKK04]. Operational [SPS^{+}07, Cad00a, DMA07, HFS02]. operations [The07]. opercular [KK09]. opilio [BFM08, CWS^{+}05, HMM^{+}01, MD09, Zhe03]. opisoplectum [PMAK07]. opportunities [ASM^{+}03, BM08, GDFH07]. opposed [GHDS08]. Optimal [CXMM04, AKA^{+}02, CATH04, VVH04]. optimisation [Pri09]. Optimising [Hen09]. optimization [CQJX08]. Optimizing [LRC^{+}08, FÖB08]. Options [How03b, CQJX08, CL04, HIRM06, RMA^{+}05]. Optomotor [KW03a]. orange [CAFT00, Cla01, GAFA05a, KM02, MC05, SR06, SRH^{+}02]. orange-spotted [GAFA05a]. orcini [BSAM09]. Oregon [Cro01, GHS04, GJSW04, KE05]. Organic [BM03b]. organisation [Lon09]. Organised [RB09]. organization [Ano07y, Ano09a]. organs [PL00]. orientalis [STI08, STCH09]. orientation [BH01b, ODR00, PM09, TKFJ09]. Origin [CMOM00, JLHO01]. ornamental [PA05]. ornate [YPDS05]. ornatus [YPDS05]. ORSTOM [Ano00d]. Oscillation [Sol02]. Oscillatory [HMB^{+}00]. Osmerus [JAKM05, PMT06, SRTP06]. Osney [Cur00]. Osteichthyes [Eze02, MNM00]. Osteoglossiformes [SND^{+}08]. Otaria [SPS^{+}07]. other [BH01b, BRRL09, CMM^{+}07, Dou04, Mag01, Mey07, PGK03, RDH06, SCP01, SWC^{+}05]. Otley [Bar00]. OTMS [SRA^{+}02]. Otolith [ASD09, CCH^{+}00, CT00, KR00, MNBM05, MMN00b, PCAS04, SMSZ08, SFCK03, TP00, VBSK00, ACG^{+}01, BMBM05, BMC07, Ber07, BBK08, CAJ00a, CAJ00b, CMOM00, CS07b, DGP02, Dou08, EPA00, FL05, Fab06, FCC^{+}09, FAN03, HMM^{+}09, KAN00, KK09, LD08, LP00, LMR^{+}05, LMR^{+}07, MÁ08, MK09, MN00, MBDM05, MRT03, NMF^{+}08, PBM06, PGK03, PRdPG08, PRM04, PL00, ST00, SR00, SDS09, SBF^{+}08, TLMN00, TLD06, TRL06, VP05, YST^{+}02, ZB00, ZSP03]. Otolithic [LBMCE00]. Otoliths [MG00, ABC^{+}08, ATB^{+}02, ACE^{+}08, BM00, BG04, DSPLR06, Dou04, FCC^{+}09, GB03, GJSW04, GCC^{+}08, HMB^{+}00, JSB04, KF02, Ker00, MMP^{+}02, MDGR05, MMN03, NCW00b, RC06, TBRL00, WWG03]. otter [BPW^{+}03, FKW06, FRF08, HC00, Hal02, JRG08]. otterboard [SdPFAL09]. otterboards [PAF^{+}08]. ottertrawl [MTHL06]. Otto [Ore03]. oualaniensis [XBS^{+}07]. outcomes [GAC^{+}06, SSG06]. outer [CGT08, MHS^{+}05, SBF^{+}08, WF07]. outline [WG05]. outlook [GDAM07]. output [AGB07]. ovaries [MSB07]. ovary [JWR05]. ovatus [Aló09]. over-exploitation [MBK02]. overall [IJ06]. Overestimation [PKŘ^{+}08]. overexploited [GCN02, WM04]. overfished [HP09b]. Overfishing [Ano02l, CHN08, SGS09, dGvZM^{+}06]. overlooked [MSB^{+}08]. overseas [Gra02a]. Overview [RV03, PPdC01a, Rei01, TK00, VD04b]. Oxford [Cur00, Loc02, McI02d, Mil02b, Mil03b, Mil08, She02, Sol05, Sut02, Win02]. oxide [LET07]. oxygen [GMP05]. oxyrinchus [CMJ05, CL09, RCGMQV05]. Oyashio [AM00].

P [Ano00e, Coo00a, McI06b, McI08, McI09a]. PA [DBAÖ06, SLB04, TÖT04]. Pacific [Ano09a, AM00, CSM^{+}09, CL09, CCT^{+}08, Fra07, GLCC09, MKB01, RWMBHH06, SC07, TSOG01, WSPY05, WSPY07, WCC^{+}09, YM00, ASD09, BNSA05, BNS07, BSMFH07, Bro05b, BM03b, BM03c, CGLZ04, CC03c, CHL^{+}06, CMOM00, DOQ08, FR03, GB03, Ger09, GEYPC^{+}09, Hoo06, HJLL07, IMOO06, IPH05, JHK08, Jun08, JCJ^{+}09, KS08, KF02, Loh08, MS05b, MAASMB05, MAdAMASCM09, MROGVR09, NW03, Nas03, NHT07, ORVMP06, PMSD06, QVNMGM00, STI08, STCH09, SN01, SC07, SOH06, SK08b, SYS^{+}08, SAK05, SWA08, TAS03, TCC^{+}09, USKK04, VL09, VBSK00, WIS03, WMdSB09, WMdS09, vSSK07]. pacificus [CLH^{+}08]. paddlefish [BS06, SB07]. Page [Ano03f, Ano03g, Ano03h, Ano04e, Ano04f, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano06n]. Pagellus [CPT06]. pages [Ano00b, Ano00c, Ano00d, Ano00e, Ano01c, Ano01d, Ano01e, Ano01f, Bar00, Coo00a, Coo00b, Coo01, Cur00, How03a, How03b, Lai00, Loc02, Lon02a, McI01, McI02a, McI02c, McI02e, McI02d, McI02f, McI03a, McI03e, McI03b, Mil02a, Mil02b, Mil03b, Ore03, She02, Sut02, Van01, Win00, Win02, Ano00r, Ano00l, Ano00n, Ano00h, Ano00p, Ano00i, Ano00o, Ano00m, Ano00k, Ano00j, Ano00q, Ano00s, Ano01j, Ano01l, Ano01p, Ano01n, Ano01o, Ano01q, Ano01m, Ano01k, Ano02n, Ano02r, Ano02p, Ano02y, Ano02u, Ano02w, Ano02m, Ano02q, Ano02s, Ano02t, Ano02x, Ano02o, Ano02v, Ano03t, Ano03y, Ano03z, Ano03-29, Ano03w, Ano03-27, Ano03u, Ano03v, Ano03-30, Ano03x, Ano03-28, Ano04w, Ano04p, Ano04s, Ano04n, Ano04q, Ano04r, Ano04u, Ano04o, Ano04v, Ano04t, Ano05q]. Pages [Ano05v, Ano05-28, Ano05w, Ano05r, Ano05u, Ano05-27, Ano05z, Ano05x, Ano05y, Ano05s, Ano05t, Ano06v, Ano06z, Ano06t, Ano06u, Ano06-27, Ano06-28, Ano06x, Ano06w, Ano06s, Ano06y, Ano07s, Ano07q, Ano07o, Ano07r, Ano07n, Ano07p, Ano07t, Ano07u, Ano07v, Ano08r, Ano08s, Ano08m, Ano08t, Ano08n, Ano08u, Ano08o, Ano08p, Ano08q, Ano08v, Ano09v, Ano09-29, Ano09y, Ano09u, Ano09z, Ano09t, Ano09w, Ano09x, Ano09-27, Ano09-28, Ano09-30]. Pagrus [ATM08, HMM00b, HMM00a, MC06, PMG^{+}06, Ste08, SJ05, SOKS08, WMTJ07]. pagurus [BEBA03, EBA^{+}03, UHN07]. painted [Fri07]. pair [Sch09]. pair-trawl [Sch09]. paired [HFHO07, SHL09]. paired-gear [HFHO07, SHL09]. Palinurus [BMF^{+}10, GQR03, GG04, Gro00, GMS06, TSP^{+}04]. Pallas [Dou08]. pallasi [Oja06]. Pan [SF05]. pandalid [KCK03]. Pandalus [GHS04, HJ03, HJ07, KCKP02, PAR^{+}03, Wie04, WSP06]. panel [BNL^{+}07, GKF03, GOF^{+}04, OKF06, OGK^{+}08]. panels [BKD^{+}02, CF04, GK01, GKF03, KFM08, OKF06, RJ05, RCD07, vM03]. panmixia [vHAN^{+}09]. Panopea [MC04]. Panulirus [ARCLBPdP07, BRL^{+}05, CCMS^{+}03, Ehr08, Fri07, GYMdL^{+}06, MSTC04, PMSC03, PVMdL05, RLSJIA07, Vel03, WVW^{+}08, YPDS05]. Paperback [McI05c, Ste03a, Coo00b, McI02a, McI05d]. papers [Bla06]. Papua [FBV06, Mai09]. Paracentrotus [ASM^{+}09, GCB^{+}06, GTB04]. paradox [CP08a]. Paraguay [MAG00]. Paralichthys [LZZN09, SKF08]. Paralithodes [GFL03, GFS03, NS06]. paramamosain [OIK^{+}06]. parameter [CHL01, CY06, ODdGV02]. parameters [AGS^{+}08, Cad09, CC03c, CE07, DKKS07, EPA00, FGJD08, FKBF05, HP00, Kop06, Mag01, MMB07b, PGS^{+}08, PHGT06, RBHP03, Row02, SND^{+}08, SG04, TOK^{+}05, ZLC09]. Paraná [GMA02, PA05]. Parapenaeus [CFE02, SGB00]. Parapenaus [DBAÖ06]. Parapercis [CTB01, CAH^{+}03, CATH04]. Parasite [PBR05, GPG^{+}03, LRW09, MMH05]. Parasites [MCM^{+}08, CPSBRD^{+}02, LK00, LM09, MKGH05, OB02, OS05, OFOM08]. Parasitic [SN01]. parasitology [MTC06]. parental [Oja06]. Paris [Gue05, McI07a]. Park [BPH^{+}06, PBW^{+}09]. Pärnu [Eer04]. parr [Lan04, PB04, VP05]. parrotfish [AMS08, TOBSJH06]. part [CLH^{+}08, HBW00]. partial [BGW06]. participation [Bat04, Lon09]. particle [PAR^{+}03]. particle-tracking [PAR^{+}03]. partition [CXMM04]. partitioning [BM03a, HC09, MHWL01]. Pascual [Pie09]. pass [KA06, SET^{+}09]. passive [AD03, CBGP06]. passive-acoustical [AD03]. PAST [McI08, IMOO06]. pastinaca [Ism03]. Patagonia [Gon06, GNC07, MC04, Nar06, NGF07, CS08c]. Patagonian [BQN05, VL08, BBM^{+}06, BLH04, CGP05, Hat00, Hor02, LB05, LLV^{+}08, MPE04, PMM05, PAB05, PCAS04, TdlMH^{+}03, VL08, WPS^{+}07, WR02, YCB^{+}01]. patagonica [BBM^{+}06, BLH04, GD03, LLV^{+}08, VL08]. patagonicus [PPH01]. patches [STI08]. Patinopecten [Uki06]. Patos [VD06]. Patraikos [TSTK06, TSTK08]. patronus [VSS07]. Pattern [GSMT04, LLV^{+}08, MPE04, MWB^{+}08, Ove02, PMM05, PMG^{+}06, PH01]. Patterns [AE03, CB00, CMAOAA05, Kru04, MCB^{+}07, MCP01, ZSP03, AMS08, AM02, BLC08, BHH08, BSR05, CPCD05, FDG^{+}07, GCB^{+}06, GD03, KSS05, KVZ05, LAL07, LPP02, MBT08, PKP04, PRM04, PSPCVCD09, QDV05, RRASLB03, RLSJIA07, SWT09, SGL08, SAG05, Ste03b, TP00, UM02, VAF04, VE03, WSS02, WMD07, ZR08, ZPR01]. paucispinis [PWHM06]. Paul [Ore03]. paulensis [VD06]. pay [CGLZ04]. payao [BSAM09]. payao-associated [BSAM09]. Pb [KKJ00]. Pb [?]Kastelle:2000:UPR. PDR [GAC^{+}06]. PE [DBAÖ06, TÖT04]. pearl [Has06]. pearls [But08, KPBD08]. Pecten [Kat05]. pectinirostris [NT05]. pedra [EGB^{+}01]. pedra-bola [EGB^{+}01]. Pelagic [BMPK06, Cur00, AFJC08, AD00, BDB^{+}03, BGM^{+}09, BF00a, BCMM09, BM08, CNH09, DV00, DJG^{+}06, DKH08, EGB^{+}01, GFvdV06, HC09, HS00, JRML00, JK07, KG06, KME^{+}02, LCNJ06, OA04, PR03, PGF07, PMC^{+}07, Sch09, SHBC02, SHDO04, USKK04, VKČ^{+}09, WM07, WLDH08, WEK^{+}09, YKM06, YKO09, YSS^{+}09, ZCdG06, tHDC06, vZRMvD02]. pelagics [vOPvDM02]. pelagicus [AAH06, AAKT07, BdL05, CS09, LCNJ06, McP02, XK04]. pelamis [AC02, AS04, TSOG01]. Pellegrin [Eze02]. pen [NW03]. Penaeid [BVW04, BMK^{+}04, FFM01, GWB07, HJ06, RM05, SMJS01, TEG03, You03]. penaeid-trawl [BMK^{+}04]. Penaeidae [COHP02, GWB07, KTL06]. penaeids [AAFLP^{+}07]. penaeoid [CPCD05]. Penaeus [COHP02, CTC^{+}06, HK06b, NAD^{+}07, TY02, WZDY06, XM00, Xia04a, YLD^{+}05]. Penang [Bla06]. penguins [PPH01]. penicillata [LMASMASZ02]. Peninsula [CP00, PRdPG07, vHMAO^{+}06]. Peninsular [ABAA^{+}07]. Pennahia [YTK06]. Pentapodus [WMTJ07]. People [Ano01d]. per-recruit [AGB07, GAOMC06, HH06, TLH07, WBMM05]. Perca [PTE09, PLMDB06]. percentages [GLS^{+}09]. perception [Mar07a, vZRMvD02]. Perceptions [MvdK06]. perch [CS08c, FK04, HH08, IWBJ08, JRML00, KLU01, PH09, PTE09, PLMDB06, SHKB09, SB08, WIJB08]. Percichthys [CS08c]. percid [PKŘ^{+}08]. Percidae [PKŘ^{+}08]. Perciformes [LC07, vHMAO^{+}06]. Percoidei [CWL00]. percoides [PH09]. Pereira [How03b]. perezi [GCS^{+}06]. Performance [GG04, SdPFAL09, WB08, BCLP00, BDBD09, DB08, DDP^{+}06c, DDP^{+}08, ET00, GR00, LFA08, MD09, Pun03, Rue07, RSMZ06, SLE09, WIJB08, vSM02c, vSM02a, vSM02b]. peri [CAH^{+}03]. peri-mortem [CAH^{+}03]. period [Bas99, Bas00, CS04, CLGR^{+}04, EMN07, HCH09, JRG08, MAND05, SMD04, WW06, XJ05]. periodicity [CS07b, RC06, YTK06]. periods [MC06]. Periophthalmus [EKU02]. permanent [Mey07, RSST07]. Perna [ABAA^{+}07]. peronii [Ste03c]. Persian [NAD^{+}07]. Persistence [MH08, vZRMvD02]. perspective [Dun06, GAC^{+}06, HMO01, PHM07]. perspective-based [Dun06]. Perspectives [PA05, Ste03a, MSM^{+}08, MF03]. perturbations [ADP^{+}08]. Peru [Ara09, WYR06]. Peruvian [ARVC01, TVY01, TYM^{+}01, Vil01]. Peter [Sho09]. Peterhead [BGH05]. petroleum [SW00]. Petromyzon [BTC08a]. phase [LWM07, OGPDOM08, RYSN08]. phases [GLS^{+}09]. phenomenon [MSB^{+}08, PEAAdLJB06]. Phil [Skr08]. Philippine [CHL^{+}06]. Philippines [Bat04, BP01b, GML06, MWJ04, Qui07, VMPP07]. philopatry [PBLP08]. Phocoena [LET07]. Pholis [HLH08]. Phosphoglucomutase [JS02]. Photichthyidae [TP00]. photo [GR07]. photographed [TMK^{+}06]. Photololigo [WLL08]. Phycis [CP00].

Physeter [PSS^{+}07]. physical [OSB09, QDH09, SLB04]. Physiological [AKC^{+}09, MM05, WSC08, ZFJ03]. physiology [HCH09, MMN00b]. picarel [DKGC00]. picture [KHM^{+}09, TMK^{+}06]. pictures [FAN03]. Pierce [Ano01r]. Pierre [Cur00, Gue05]. pike [AKC^{+}09, CPR07, JRML00, LCD^{+}02, LPL^{+}06, LPG^{+}09, SB08]. pike-perch [JRML00]. piked [BGW06]. pikeperch [Eer04, GWR07, HSSR06, KLU01]. Pilbara [New02]. pilchard [MG03a, MG03b]. pilchardus [QP00, RCGC04, RCG02]. pile [JLC^{+}03]. pillage [RB09]. Pilot [Güç08, MTA^{+}02, PBR05]. pinfish [EF02]. Pingers [GGA08]. Pinguipedidae [CTB01]. pink [BCM08b, GOAARJRP09, PCD03, RRAS03, RRASLB03, VD06]. pinnotherid [NS04]. Piracicaba [SB01b]. piscatorius [DALP01, LPDA01, LQD^{+}08, LJP05, WWG03]. Pisces [AAB05, BP06, BG04, BG05, BBL07, DKGC00, FB06, HAO00, LS00, MRCH04, MSA00, MMN03, MCS05, NV05, ODdGV02, OFOM08, RMBG02, TOBSJH06, WBMM05]. piscivorous [SKF08]. piscivory [LG02]. PIT [ARLB07, CPR07, HZZG06]. Pitcher [Mil03b]. pitfall [HK02]. PITpacks [HZZG06]. Placentia [LR00a, MR05]. Places [Ano01d]. Placopecten [KCRT03, WHWM07]. plaice [CAJ00a, CAJ00b, GM06, MBK02, MP03, SKP03, TBRL00, vM03]. Plains [QGBS04]. plan [KAO04]. plankton [OR07, TK00]. planktonic [ABL02, CC01a]. planning [Gav09, eJSK09, NJL^{+}00]. plans [HP09b]. plastic [GG04]. Plata [JMGL04, MAL02, MM06b, NV05]. Plateau [SCM09]. platessa [CAJ00a, CAJ00b, GM06, SKP03, TBRL00]. platform [BMW07]. platforms [SW00]. Platycephalidae [YEA09]. Platycephalus [MOOH00, YEA07]. platypterus [Hoo06, RAHHGM^{+}02]. plebejus [CTC^{+}06, IS07, MCB^{+}07, MSB07]. plei [PdAO02]. Pleuronectes [CAJ00a, CAJ00b, GM06, He03, SKP03, TBRL00]. plumb [AR05]. plumbeus [JLC04]. Poey [SDNL09]. Pogonias [MAL02]. point [CA07, MSB07]. point-count [MSB07]. points [Boo04, Cad02, CXCD07, GC04, HPD08, HPD09, Jen00, Koe03, Sid03, TJ05, UM02, VSS07, ZP05]. Poisson [PL06]. Poland [PLD05, PLMDB06]. pole [IV08]. pole-hooking [IV08]. policies [ABD^{+}09, BDIP08, CP08a, CK08, DB08, HP02, IWBJ08, PPGD08]. Policy [CS08b, Lev05, Mar07a, She03, Sym05a, SP09, WIJB08, Sym09]. Polish [GDDJ09, Orl03]. Pollicipes [MF03]. pollock [APC07, BQ04, Dou08, IFH^{+}09, ODR00, RO00, SC07, SWC^{+}05, TMK^{+}06, vSSK07, vSS09]. polyamide [SLB07, TH06]. Polydactylus [CGLZ04]. polymorphism [LPL^{+}06]. Polynemidae [PJR^{+}05]. Polynemus [PJR^{+}05]. Polynesia [BDB^{+}03]. polynomial [Bas99, Bas00]. Polyprion [PH04b]. Polysteganus [CMGB01]. Pomacanthus [SDS09]. Pomadasys [AHAAD01, AHABAA^{+}02]. Pomatomidae [SB05]. Pomatomus [AWY02, RRJ08, RRJ09, SB05, TOÖD06]. pomfret [BSAM09]. pontoon [HFL08]. pool [Jen00]. poor [DSK^{+}08, MGR^{+}02, VMPP07]. Pope [XW07].

Popper [Cor09]. Population [ATM08, ABAA^{+}07, BNS00, CHL^{+}06, CHW^{+}08, CE07, CCD01, DG08, FGJD08, GAFA04a, GAFA05a, GFvdV06, GWR07, KN01, KM08c, Lor06, Mar06b, MMLT05, NAD^{+}07, ODdGV02, RBHP03, Row02, SCUADJ07, ŠV04, SG04, UMV^{+}09, WLL08, WCC^{+}09, ABFV02, All07, AAHE04, AAB05, ARVC01, ASD09, BEBA03, BW08, BD04, But03, CSM^{+}09, CCM^{+}09, Caz00, COHP02, CL09, CF07, DCM^{+}05, DMC^{+}09, DMBF08, ED07, EPA00, EH04b, FÖB08, GCN02, GJ09, Ger09, GR07, GD03, HH08, HKG^{+}08, HTF^{+}09, IWW03, Jen02, JC04, KM08a, KK04, LPG^{+}09, MGR^{+}02, ML04a, MB04, MC06, MG03a, MG03b, MGJB09, NDCK09, NNE^{+}05, PCM07, PAB05, PP03, PdAO02, RMCdS06, RMRB06, RSST07, RCSFO04, SK03a, SG06, SW02, SWA08, UMK01, War06, WY06, Xia06, YJV^{+}03, vHMAO^{+}06, vHAN^{+}09]. populations [BBK08, BMSJ05, CB00, CBGP06, CA07, EAT05, HP09a, Has06, HWW^{+}09, LCTB^{+}03, Lon02b, LMR^{+}05, LMR^{+}07, LPL^{+}06, MTC06, MG00, MEA^{+}03, MRT03, OLN^{+}02, PAR^{+}03, PBM06, PMSD06, QDV05, RW03, SF05, SHKB09, SSML02, SFCK03, TGS^{+}00, Tze04, WFJB09, YMT^{+}06, McI08]. porcupine [He05, KMB01, SR06]. porgy [ATM08]. porpoise [LET07]. Porsuk [Yil02]. Port [SC08, SCW04, TÁTWD09]. portable [HZZG06]. porteri [AA03]. Porto [CdCBdR03]. Portugal [CF04, EGB^{+}01, GBC^{+}03b, SGVM02, SSGM03a, SSGM03b, CDC03, CE06, CE07, GC03a, GBC^{+}03a, LE01, OGPC09, PF00, SB01a, SAG05]. Portuguese [FCL^{+}05a, BCC09a, BCC09b, BTC09, CFE02, CFE03, CFH03, CFFP07, CC01a, FFM01, FCG02, FMCS05, FCML05, FCL^{+}05b, FCM07, FCAD^{+}08, JR07, MTC06, MF09, Par04, PFHC08, PTR^{+}05, TC09]. Portunidae [FW06]. Portunus [AAH06, AAKT07, BdL05, CS09, McP02, XK04]. portusjacksoni [TÁTWD09]. Position [MTA^{+}02, GK01, GKF03]. positional [Pal08]. positive [PSS^{+}07]. positively [CTK^{+}07]. Possible [JHGS04, LC07, AM02, BHB^{+}04, CPW^{+}06, GB03, Hat00]. Post [ARCLBPdP07, GCMJ08, Kor04, MR08a, NGM^{+}09, PRM04, SMB^{+}08]. Post-larval [ARCLBPdP07, PRM04]. post-processing [Kor04]. post-release [GCMJ08, NGM^{+}09, SMB^{+}08]. post-settlement [MR08a]. pot [AKA06, AAKT07, CS09, CTB01, FHJL08]. Potential [Ano06-29, BBLL06, BACP09, BNL^{+}08, CAJ00a, CAJ00b, DUR05, FCML05, RB06, SALnM07, SSIE06, ZDB^{+}08, CTK^{+}07, CSSO03, FÖB08, LPHH09, LMASMASZ02, MC00a, MCB^{+}07, New02, WM07]. pots [AAH06, AAT08, CATH04, GFS03, JPYW08]. potting [CAH^{+}03]. pour [Gue05]. pout [KHMG06, NM06]. poutassou [ÅCST01, BH01a, FAL^{+}01, PKK^{+}06]. power [HWM02, OCT^{+}03, OL07, WM07]. Powers [Coo00a]. pp [Ell00, Gue05, How06, Kes00, McI05a, McI05c, McI05d, McI06b, McI06a, McI07b, McI07a, McI09a, McI09b, McI09d, Mil06, Pie09, Sol05, Ste05, Sym07, Sym08a, Sym08c, Sym08b]. PPS [CCBG02]. Practical [CK08, FCS^{+}05, YJV^{+}03]. practice [PLD05]. Practices [McI02d]. Prawn [DPD^{+}03, DDP^{+}06a, DDP^{+}06b, DDP^{+}06c, VD04a, BHM^{+}06, BKD^{+}02, BPW^{+}03, CTC^{+}06, GKH03, Gri03, HK06b, HGW^{+}08, HW00, IS07, LYKH06, NAD^{+}07, PBW^{+}09, Sva05, SRS08, TGH^{+}08, TCY01, WMTJ07, XM00, Xia04a, YLD^{+}05, Gri03]. prawn-trawl [BKD^{+}02]. prawns [DPD^{+}03, GWB07, MMB^{+}07a, MCB^{+}07, RM05, SBFR00, TEG03, TVB^{+}09, You03, ZDB^{+}07]. Praxis [Ell00, Van01]. Pre [SSH^{+}06, PCS05a, SLJ07]. pre-fishery [PCS05a]. pre-recruit [SLJ07]. Pre-spawning [SSH^{+}06]. Precautionary [CP08a, Hol01, PZH02, Sch01]. precipitation [DOQ08]. precise [Ker00]. Precision [AFSF05, AB08a, PKK^{+}06, DVAL^{+}09, FP00, FAN03, HFS02, HCS^{+}03, Joh03, MM06a]. Predation [BG01, HSGF06, SKF08, BSWL06a, BSWL06b, GMS06, SB08, SBD^{+}05, SBD^{+}06, VMSR08]. Predator [CND02]. Predators [Bjø01, BBM^{+}06, DV00, SCP01]. predatory [AB08b, AB08a, ORVMP06]. predict [LMR^{+}05, LMR^{+}07, SDS09]. Predicting [AB08b, ABS^{+}03, DKKS07, OSG08, CATH04, MSTC04]. Prediction [HKF^{+}09, Meh06, CC01b, GEYPC^{+}09, UMV^{+}09]. predictions [HE08a]. Predictive [HJ06, KMELC05]. predictor [PGK03]. Predictors [Par04, AB08a, DO06, DUR05, PBR05]. Preface [CSCS07, PPdC01b, RYG01]. preference [SJ05]. preferences [ATM08, HC09, NM06]. preferential [RLSJIA07]. prehistoric [CGT08]. Preliminary [ABK03, GAFA05b, MAG^{+}09, PAB05, PLD05, GRPLE05, RW07, TCS05, YCB^{+}01]. presbyter [MMN03, MCS05]. Presence [NV05, SF05, SBD^{+}05, SBD^{+}06]. Present [DR01, DFD02, MSM^{+}08]. preservation [MFB08]. Press [Ano00b, Ano01d, Coo00b, How03a, How03b, McI02e, McI02f, McI03e, Mil02a, Mil06, Ste03a]. pressure [CP06, CS04, GBFG06, LZC^{+}08, RMCdS06, SB08, TOBSJH06]. prevalence [BCM08b]. prevent [VMSR08]. previous [DKKS07]. Prey [TH03, APC07, AB08b, AB08a, BSAM09, CND02, DV00, FUM02, PPH01, SBD^{+}05, SBD^{+}06]. Prey-specific [TH03]. Pribilof [BMB06, BMB05, SWC^{+}05]. Price [How06, McI05c, McI06a, McI07b, McI07a, McI09a, McI09b, Mil06, Sym07, Sym08a, Sym08c, Sym08b, BCC09a, JTdH06, RMN08, SMD04]. primary [FFL09a, FFL09b, KJ09, Ste07]. Prince [KTC00]. princeps [BLC08, Jak01]. principles [SSK^{+}06]. Prionace [dSHG08, LSH04, WK01]. Pristipomoides [OLN^{+}02]. probability [NS06, Sal02, VGBF02, XK04]. probable [HW00]. probe [JWR05]. problem [VVH04]. Problems [Ano06-29, BBLL06, PD08, SCP01]. procedure [But08, Cad00a, CWS^{+}05, DBK07, KPBD08]. Procedures [KJ08]. Proceedings [Ore03, Ano06-30, Ano07w]. process [BLH04, But03, GLS^{+}09, LB08, VL08, YEA09]. processes [MN00, ZFJ03]. processing [JL00, Kor04, LRG^{+}00, TBRL00]. produced [DDHS06]. producing [KJ08]. product [SLJ07]. production [CA03, CHSA02, CRC^{+}07, Gus02, JC04, KKR^{+}05, KJ09, KK04, MPE04, Mai09, MLH^{+}05, MK01, MM06c, New02, Pra02, SKMT02, TPC05, WM09].

Productivity [OGPC09, MUO02, RL03]. products [MS05b]. Professor [Sym08a]. Profile [RB09, TCS05, TKFJ09]. profiler [DBB00, OSB09]. Program [Coo00a, Gra02a, MN03]. Programme [McI06b, KJS^{+}06, NJE00, GHDS08]. programmes [BBRS06, KK06, War06]. programs [SCW04, SLE09]. Progress [Ano06-29, Sym07, BBLL06, KHM^{+}09]. project [AMC^{+}08a, AKA^{+}01]. prolonged [CS04]. properties [DWW^{+}02, JIDM02, MAG00, RH04a, RDH06, SLB04]. Property [RDEACT07]. Proposed [SWM09]. prospect [GNL^{+}09]. Protected [Mey07, BMF^{+}10, BG04, CXQ^{+}09, GCS^{+}06, LHMO09, LZC^{+}08, MWJ04, MH08, McI02e, NGF06, SSH^{+}06, WM04]. Protecting [Hol01, Sch01]. Protection [Ste05]. protective [PLD05]. protocol [Hur07]. protocols [AFG^{+}05, RSP^{+}00]. provided [BRRL09]. providing [LYKH06]. Province [RB09]. provision [GFMO06]. Proximity [NHT07]. Psetta [SSML02]. Pseudopercis [Gon06]. Pseudopleuronectes [Laz08]. pseudoreplication [MA04]. psychology [Koe08]. Publication [McI05d]. publications [Kes00]. published [Xia06]. Publisher [Ano01s, Ano01t, Ano02z, Ano03-31, Ano04x, Ano05-29, Ano06-31, Ano09-32, Ano09-31, Ano03-32, Ano04y]. Publishers [Ano01c, Ano01e, Ano01f, Lai00, McI05c, McI09a, McI09b, McI09d, Van01]. Publishing [Bar00, Ell00, McI05b, McI06a, McI07b, Mil08, Sym07, Sym08a, Sym08c, Sym08b]. Puerto [MDJR05, MC04]. pueruli [PMSC03]. puerulus [BM09, GFMO06]. Puget [QDV05]. Pulau [TCS05]. pulse [Mey07]. puntazzo [DSPLR06, PA02]. purpleback [XBS^{+}07]. purpose [TPM03]. purposes [SM04a, McI05d]. purpuratus [OS05]. purse [BAEF06, Kar04, KLS^{+}07, MROGVR09, MB00, MBGW02, ORVMP06, RT07, Vas03, vZRMvD02]. purse-seine [Kar04, ORVMP06, RT07, vZRMvD02]. pushnets [OR07]. pusillus [CE07].

Qasim [McI06a]. quahogs [TJ05]. qualitative [AHS^{+}06, PLSI09]. quality [BW02, KJ08]. quantify [AKA^{+}01, MSB07]. Quantifying [CTC^{+}06, MHK^{+}06, DV00]. Quantitative [NJE00, WG03, YEA09, KW03a, PLSI09, USKK04, ZG08]. Queen [ZP05]. queenfish [GFvdV06]. Queensland [Bal09, CS09, CCD01, CTC^{+}06, OCT^{+}03, OL07, WSS02]. questionable [BQ06]. quickly [SDS09]. quid [FWG15, Sve14]. Quinn [Coo00a]. Quintana [AIC08]. quite [How02]. quota [Ara09]. quotas [KPL^{+}07, KPL^{+}09]. Quoy [NCW00a].

R [Ano01f, BNS07, Fra07, McI02c, McI06b, Ore03, She02, Sym08a, Sym08b, Van01]. R. [CK02, ESC08, Mil06]. Ra [KKJ00]. radiation [AOMK^{+}04]. radio [MNML00]. radiocarbon [PWHM06]. Radiometric [SKKP05, WKCC06]. radius [ST00]. rainbow [MM05, SRTP06]. rainforest [Eze02]. Raja [CE06, ESC08, Ger09, MK06]. Rajidae [CGP05, ECEG09]. Ram [Tay07]. Ranching [McI05b, She02]. range [CPW^{+}06, LLV^{+}08, MGR^{+}02, SMRH04]. ranges [Hor00]. Ranong [MMLT05]. Rao [Bas00, Bas99]. RAPD [SCUADJ07]. RAPFISH [PP01, TP06]. Rapid [TMN07, PP01, PPGD08]. rare [KS08, SK08b]. rare-earth [KS08]. rate [AKA06, AAKT07, BB04, BH01a, BBJG04, BPW^{+}03, FH03b, GR07, HH01, KM02, MWB^{+}08, MAdAMASCM09, New02, STI08, SG04, WSP06, YLYL09, You03, ZW09]. rates [AD00, BS06, BCLP00, CGLZ04, CWL00, CC01a, CSSO03, DMA07, EJA00, FOGKT08, GRFE06, GHDM08, GSS09, HG01, HBKJ07, JCS01, KTL06, LDGA09, LWM07, MLJT09, MK01, NCW00b, OA04, OSC^{+}05, PG06, PCS05b, RCGC04, SN09, SPB^{+}03, SSGM03a, SSB03, SMS^{+}06, SFCK03, TH03, WCM04, WK01, WKM02, WT00, WMB08, WPOB09, WBHG01, YAHAB01, vM03, vZRMvD02]. ratio [FHI07, KTL06, XK04]. rational [Cor02]. ratios [AB09b, KAN00, RCGC04]. Raton [How06]. ray [BSMFH07, CE06, HSW09, KK09, WC07, LE01]. rays [WGDP06]. Re [BLR^{+}08]. Re-evaluating [BLR^{+}08]. reaction [KW03a, SG06, SSH^{+}06, WG03]. Reactions [Sch09]. reactivity [RB06]. reader [EPA00]. readers [PKK^{+}06]. reading [ATB^{+}02, AFG^{+}05, DSPLR06, RMRB06]. real [YEA09]. real-time [YEA09]. really [Lon07]. reared [ACE^{+}08, CPR07, FUM02, Jok02, JJKN^{+}03, KNJJS06, KNSSKF09, MNBM05, SSML02, SKF08]. Rebuilding [KIZ09, BCLM08, HP09b]. Reburial [CCG^{+}02]. recapture [APP^{+}07, BEBA03, BG04, Ehr08, FML^{+}07, LRBH07, PRdPG07, RMRB06, SLA00, SMB^{+}08, TdlMH^{+}03, UHN07, XM00, ZLC09]. recaptures [JLHO01, Ste03c]. Recherche [Gue05]. reciprocal [PCD03]. recognition [AGS^{+}08, SD01]. recommendations [MTA^{+}02, Tia09, WBMM05]. Reconstructing [CNH09, FCC^{+}09, OM07]. Reconstruction [BK09]. recorded [HB01]. Recording [MC00a, ØT02]. recordings [MC00a]. records [JSB04, MHS^{+}05, MSM^{+}08, WM00]. recovered [CMOM00]. Recovery [GHDS08, BTC08b, CWL00, CAHAF07, MN03, MG03a, MG03b, Ove02, PPS05, SW02]. Recreational [CGLZ04, ABD^{+}09, AAP^{+}09, Aló09, ABF08, BCSP03, BRRL09, DACS08, GCB^{+}06, Hen09, LZCR08, MSTC04, Mey07, MJS00, OF03, PBLP08, PWB05, RCA08, RM05, SMM07, SMB^{+}08, VLD^{+}00, Wil09, ZDB^{+}08, Mil03b]. recreationally [GCJ07, GCMJ08]. recruit [AHABAA^{+}02, AGB07, AWY02, Cad09, CL09, GHS04, GBC^{+}03a, GBC^{+}03b, GAOMC06, HH06, Kir01, Kop06, LCNJ06, PJR^{+}05, SLJ07, SEPY02, SWPY05, TLH07, WBMM05, ZC02]. recruited [BM09]. Recruitment [RRASLB03, AS04, BCSA08, BK09, BP01a, CS04, CC01a, CC01b, CC01c, DMC^{+}09, DUR05, FH03a, FHG03, GHCPPLC^{+}05, HBM^{+}00, Kno07, Lon02b, MLC03, MMLT05, NDCK09, NMMTCV^{+}06, PAB05, PKP04, QGBS04, RB00, Sol02, VN00, VGBF02, Wah03, WL06a, WM04, ZK03, Zho07, dGvZM^{+}06]. recruitment-overfishing [dGvZM^{+}06]. Red [PNJS06, You03, ATM08, AFSF05, BK09, BR05, BM09, BI05, BLB07, BD04, CA03, CGP05, DA07, EF02, FHJL08, GHCPPLC^{+}05, GFL03, GFS03, HS03, JDS00, JTdH06, LRBH07, LZZN09, LCFP^{+}08, MM06a, MMB07b, MTU^{+}04, NCW00b, NS06, PK03a, RBD02, Sid03, TCY01, WBCP08, WMD07, TP06, TMN07]. red-spot [TCY01]. redbait [EL09]. redfish [GR02, SKKP05]. redistribution [PA09]. reduce [BR05, BKD^{+}02, GLK00, KFM08, Løk03, MTHL06, PCMC08, RJ05, WLDH08]. Reduced [GFL03, HGW^{+}08, Hur07]. reduces [BSWL06b, KHMG06, SCM09]. Reducing [FP02, RH04b, SGRK09, SK08b, SHBÅ07, WRGS09, APM^{+}09, BSWL06a, ERG09, GL05, HPK07]. Reduction [GMY^{+}09, HGW^{+}08, LET07, SSK^{+}06, BSC^{+}08, BHM^{+}06, CF04, CTC^{+}06, CHDGM06, FCL^{+}05a, FCML05, FCL^{+}05b, Gra03, HJ07, SBFR00, VD06, vMBGF05]. Redundancy [She09]. reed [ESB09]. Reef [BPW^{+}03, BPH^{+}06, CP06, CdCBdR03, Fri07, GWB07, HW00, How03a, NCW00a, PBW^{+}09, Tay07, BRL^{+}05, BMBM05, CdMLGC09, CA07, CGT08, FFM08, FFL09a, FFL09b, GCS^{+}06, Gob00, GC03b, HK06a, HFS02, HWW^{+}09, KVZ05, LMR^{+}05, LMR^{+}07, MWB^{+}08, MM06a, MMB07b, OLN^{+}02, PSvDPM01, RW07, RLSJIA07, SK03a, Sei07, SWT09, SK08a, SMB^{+}08, TCS05, TMN07, WM04, WMB08, ZW09, Gri03, HWW^{+}09, MLP^{+}08, NCW00b, NS07, WMD07]. reefs [HW00, JLC^{+}03, RBH^{+}08, Ano00c]. Reevaluation [BM00]. Referees [Ano07x, Ano09s]. reference [Boo04, Cad02, CXCD07, GC04, HPD08, HPD09, Jen00, Koe03, Sid03, TJ05, UM02, VSS07, ZP05]. refinements [GB06]. reflect [FKW06, ZR08]. reflects [VBSK00]. reflex [DO06]. Reform [Sym09]. regime [GB03, RdFBS01]. regimes [Eer04, Tia09]. Region [TEYA02, BMB05, BMB06, CMAS08, GEH^{+}07, MWF06, RMVJGMMV07, Tal09, ZP05]. Regional [HGF06, LCFP^{+}08, Lon09, RG07, SN09, Tal09, GSS^{+}06, KSS05, Row07, SSG06, TDK^{+}05]. Regions [Ore03, AM00, LNG04, SWC^{+}05]. regression [CJ00, DFD02, MLCGRV07, PL06, WK01]. regulated [BNS00, KGB09]. regulation [FMGG08, HSSR06, JJJ03, MN00, PH01, Ano00b]. regulations [CLPP05, GAOMC06, Kir01, NNE^{+}05, TB04]. rehabilitation [HCHO^{+}03]. Reilly [Lai00]. Reinhardt [Jak01]. Reinhardtius [ANS^{+}01, GRB01, HLHF03, KGB09, LR08, WSP06, WBHG01]. related [DGS^{+}05, GJBY05, KMF08, LC07, PWP02, QGBS04, Ste08, Vel03]. relating [GKF03]. relation [AMUE01, BCSA08, BEA^{+}08, Gas03, GK01, JR02, KNJJS06, KMB01, LMASHV^{+}03, MR05, NLJS01, RKEJR06, SN09, ST00, SHDO04, Ste03b, TAS03, TSTK08, WCM04, WY06, Wie04]. relations [PR03]. Relationship [DOQ08, LWM07, RMN08, SWC^{+}05, AC02, ACG^{+}01, ARCLBPdP07, CG04, Chi00, GDDJ09, GMV00, GMV01, HWB06, JDMN00, Kno07, MGNB05, OMW^{+}03, SFCK03]. Relationships [LSW06, VYMGM^{+}09, BM09, BGW06, Cad09, CCMS^{+}03, EAT05, GM07, GG06, GJBY05, LH04, LK00, LMR^{+}05, LMR^{+}07, MAL^{+}01, MMM^{+}03, PL00, RDM^{+}09, RMMP06, SB01a, SGVM02, SCLM06, SRH^{+}02, SK03b, SAK05, TSP^{+}04, Wah03, WL06a, ZPR01, ZK03]. Relative [BM08, BM03a, BMSJ05, DB08, GMA02, HK06a, HdS07, OCT^{+}03, OGPDOM08, Rye08, SET^{+}09, TSOG01]. relatively [DA07, EAT05]. release [Aló09, AKC^{+}09, GCMJ08, Hen09, JS05, Lor06, MGS02, NGM^{+}09, RJ05, SN09, Ste03b, SMB^{+}08, TNFF03, THN^{+}04, TNL07, iTM04, iTMI06, TdlMH^{+}03, ZCdG06]. released [CPR07, GCH^{+}07, MBGW02, SSIE06, SKF08, Tri06, VAYD04]. Releases [PAP04, GMY^{+}09, SSML02]. releasing [CAM^{+}03b]. Relevance [IJ06, BCLP00]. relevant [FHM09]. Reliability [CWJS09]. Remedies [MA04]. remote [CC09, Dun06, KVZ05, PM09, San00]. Remotely [WGR08, WPS^{+}07, WYR06]. removal [FHI07, IHF08]. remove [FMC^{+}09]. removing [PMSC03]. renewal [DMC^{+}09]. reparameterization [SHBÅ07]. repeated [MJB02]. replacing [Cor02]. replenishment [PAP04]. Reply [KW03b, Lon03a, PEAAdLJB07, Pra03, SH04, Sve15, Sho09, SS00b]. Report [Coo00a, Gra02a, McI02e, McI02f]. reported [MROGVR09]. reporting [FCS^{+}05, FM04, FH03b, JDS00]. reports [WKM02]. represent [DKH08]. representatives [GAFA04a]. Reproduction [COHP02, DALP01, HAO00, MAL02, TSM03, TEYA02, CMAOAA05, Fri07, GAF06, HFB^{+}06, HGHLCH02, Ism03, JCLL08, LMRQNM^{+}05, LDE^{+}02, LBABH09, ŠV04, TVY01]. Reproductive [All01, CS02, CE06, ESC08, EL09, FAW04, IOS07, MSN01, MCS05, OGSG07, PMM05, RRJ08, SR07, TLH07, Vel03, YTK06, Yil02, YFT^{+}02, AMS08, AGB07, BSMFH07, CMC09, HCH09, IYI07, LPHH09, Mar06b, MGR^{+}02, MCB^{+}07, NJJ06, NV05, OMF^{+}03, TOK^{+}05, WGDP06, McI09d]. requirements [AKK^{+}03]. Res [Bas00, BMB06, BCM08a, CAJ00a, FCL^{+}05a, GBC^{+}03b, GMV01, GMR02, KPL^{+}09, KI05, MG03a, NNA02, SBD^{+}06, vSM02c]. Research [Ano08e, Coo00b, dSBAC06, ALN03, Ano02c, Ano02d, Ano02e, BM05, BI05, Cad00a, GRC^{+}09, LYKH06, MTA^{+}02, NHT07, RGSB08, SBR^{+}08, VD04b, XPMQ04, Ano01c, Ste05, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano05c, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano08a, Ano08b, Ano08c, Ano08d, Ano09e, Ano09f, Ano09g, Ano09d]. Reserve [MWJ04, MDJR05, FFM08, LZCR08, RSST07, WHML04]. Reserves [McI02e, MMLR08, Row02]. Reservoir [AAB05, MAG00, KHM^{+}09, QGBS04, TKFJ09, VKČ^{+}09]. Reservoirs [Bar00, BP06, DV00, GMA02, MM04]. Residency [HMRG08, RR03a]. resilience [Mar07a]. resist [RMCdS06]. resolution [KCRT03, WM00]. Resolving [KJS^{+}06, PD08]. resource [FFM01, GCN02, GSS^{+}06, GB06, GAOMC06, GBG^{+}03, HPM04, KCKP02, Lot07b, MN03, MHWL01, MUO02]. Resources [McI05d, Mil03b, Bat04, COLB07, CMM^{+}07, CGT08, Gob00, HIRM06, JPVKS^{+}01, KMF08, LZCR08, PPGD08, RCA08, SC08, SST^{+}06]. Response [BNS07, BCM08a, PCHM08, RDM^{+}09, WNRd00, ASR05, CP06, Cor01, EF02, FWG15, KW03a, MM05, PA09]. responses [GPI04, NNE^{+}05, WG03, WSC08]. Responsible [Cad00a, SAC^{+}02, McI03b]. Restocking [ASM^{+}03, BBLL06, PK06, GML06, War06]. restore [DDHS06]. Restoring [Mil06]. restricted [PBLP08]. restrictive [NNE^{+}05]. result [GM07, HMB^{+}00]. resultant [PSvDPM01]. resulting [CHQ05, Qui07]. Results [SPS^{+}07, SHBC02, AMC^{+}08b, GDDJ09, MTA^{+}02, PRdPG07, PCHM08, PLD05]. Retained [GKH^{+}01, GJBY05]. Retention [BW08, ARLB07, CCBG02, HCH09, PTPS09]. retrieval [WW06]. Retrospective [MMS09, CMGB01, Jon00]. return [BCLP00, SN09, WCM04, Mil06]. returning [APS07]. revealed [Kru04, MvdK06]. Revealing [Wah03]. reveals [ASD09, PM09, SF05]. revenue [GHS04]. Review [Ano00b, Ano00c, Ano00d, Ano00e, Ano01c, Ano01d, Ano01e, Ano01f, Bar00, CCMS^{+}03, Coo00a, Coo00b, Coo01, Cur00, Ell00, FBLO07, Gra02a, How03a, How03b, How06, Kes00, Lai00, Loc02, Løk03, Lon02a, McI01, McI02a, McI02c, McI02e, McI02d, McI02f, McI03a, McI03e, McI03b, McI05a, McI05b, McI05c, McI06b, McI06a, McI07b, McI07a, McI08, McI09a, McI09b, McI09c, McI09d, Mil02a, Mil02b, Mil03b, Mil08, MBCMNMHH01, MN00, Ore03, Pie09, She02, Sol05, Ste03a, Ste05, Sut02, Sym07, Sym08a, Sym08c, Sym08b, Van01, Win00, Win02, BM05, BI05, CLC08, DB08, Ein09, HP02, HK06b, KHM^{+}09, Mar07b, MP04, Mos00, NNA01, NNA02, PCB01, Rye08, San00, Ste07, SG01b, VAF04]. revisit [XW07]. revisited [Lon02b]. reynaudii [GB06, RB00, SLA00]. RFLP [IJD04, JIAD03]. Rhabdosargus [RMBG02]. rhina [ESC08, Ger09, MK06]. Rhincodon [CARA^{+}07, GR07, HJLL07, JH07, Mar07b, NE07, NS07, Qui07, RMVJGMMV07, Row07, Ste07, Tay07]. Rhinobatos [IYI07]. Rhinoptera [BSMFH07]. Rhizoprionodon [MGNB05]. Rhizostomeae [PK03b]. rhomboides [EF02]. Rhomboplites [All07]. rhombus [BM05, MOK^{+}05, MOGG06]. rhythm [Lap02]. Ria [CHQ05]. Rica [FW06]. Richards [How06]. Richardson [JS05, NCW00a]. richness [KA06]. Rico [MDJR05]. Ridge [FB06, HG01]. Riga [GKUO07]. rigged [DBAÖ06]. rights [RDEACT07]. rights-based [RDEACT07]. rigid [Gra03, GOF^{+}04, JIGI06, MGS02]. Rímov [TKFJ09]. ring [BMSJ05, TBRL00, YST^{+}02]. ringens [CSCC07, CBC02, CRC^{+}07, GESY^{+}07]. Rio [VLD^{+}00, JMGL04, MAL02, MM06b, NV05]. Rise [CAFT00, Lot07b, BBV07]. risk [AHS^{+}06, BBRS06, DDP^{+}06a, JCW05, Nas03, PVMdL05, SE06a, SDH00, YLD^{+}05, ZG08]. Risso [CA03, DCM^{+}05, FAL^{+}01, HAO00, LS00, PTR^{+}05, PKK^{+}06, RBD02, TGS^{+}00]. River [AAHE04, Bro05b, CWJS09, DSO^{+}05, EKU02, GMA02, JWR05, Mat07, Mil06, NJL^{+}00, NWML00, PA05, PF00, SB01b, TEYA02, Win02, BA06, BNS00, BLKR09, CS08c, Eze02, HK02, Jok02, JJJ03, LRC^{+}08, SGP02, TGS^{+}00, Wes02, BRRL09, CG04, GCMJ08, MMS09, Mil06, PCS05a, TNFF03, THN^{+}04, WCM04, ZSP03]. riverine [CXMM04, SGP02]. rivers [AKK^{+}03, SND^{+}08, SM04b]. rivulatus [BK02]. Rn [KF02]. RNA [RCGC04]. RNA/DNA [RCGC04]. roach [Hor00, SB08]. Robson [DFD02]. Rocas [GCS^{+}06]. rock [BM09, Bra04, BBJG04, BSWL06a, BSWL06b, BGB06, CCMS^{+}03, DGS^{+}05, FHI07, GFMO06, GBG^{+}03, HSGF06, IHF08, LPHH09, LMASMASZ02, MSTC04, PG06, PHGT06, YPDS05, ZFJ03].

Rockall [LDA01, MD01, NPC^{+}08]. rockfish [MF09, PWHM06, RFWH09, WKCC06]. rocky [GTB04, LZC^{+}08]. rod [BRRL09, MNML00]. rod-captured [MNML00]. rodgersii [BW02]. roe [BW02]. roes [KKR^{+}05]. rohita [KK09]. role [AAFLP^{+}07, CFG05, DMC^{+}09, PPS05, VOS01]. roles [HP09a]. rolled [MBGW02]. Rome [McI05d]. Roo [AIC08]. rookery [AFJC08]. rose [ASAGR06, CFE02, DBAÖ06, SGB00]. Rosemary [Ano01d]. Ross [BFB^{+}03, Hor02]. rostrata [FB06, Mag01]. rotenone [MK02]. Rotterdam [Win00]. rough [FAN03]. roughhead [KN01]. roughy [CAFT00, Cla01, KM02, MC05, SR06, SRH^{+}02]. round [HPK06, HPK07]. roundfish [MS05a]. roundnose [LDA01]. route [WR05]. routine [CQJX08]. Routledge [BCM08a]. RRS [BFB^{+}03]. Rs [McI06a]. rubberlip [SH09]. Ruditapes [CC01a, CC01b]. rufescens [LRBH07]. ruffe [Ogl09]. rufus [WKCC06]. rugosa [CS07a]. Ruhle [Skr08]. rule [SWM09]. rules [DB08]. run [MK09]. running [vSS09]. rupestris [LDA01]. Ruppell [JCS01]. rural [ABF08]. rush [GG04]. Russ [SH04]. russelli [JCS01]. Rutilus [Hor00].

S [Ano01c, KK04, McI03a, McI05b, McI05c, McI06a, Ore03, Pie09, CBC02, LZZN09, dSH01]. S. [SKKP05]. S.I. [Ano01i, Ano01r]. Sabah [TCS05]. sablefish [CS07b, GJSW04]. Sado [SKF08]. SAFE [ZG08, HFL08, SSK^{+}06]. safer [MFB08]. safety [FR03]. sagax [CSCC07, GFM02, GEYPC^{+}09, Ker00, MBGW02, MG03a, MG03b, WSS02]. Sagay [MWJ04]. sagitta [MDGR05, VE03, WWG03]. Sagittal [TRL06, PRdPG08, TLMN00]. sailfish [Hoo06, RAHHGM^{+}02]. Sainte [Ore03]. Sainte-Marie [Ore03]. saira [TAS03, USKK04]. saithe [MB00, OGK^{+}08]. sajori [NMF^{+}08]. salar [JRML00, BLKR09, CK02, JLHO01, Jok02, JJKN^{+}03, KFO^{+}04, MNML00, SBD^{+}05, SBD^{+}06, VP05]. sale [RMN08]. sales [WIKM05]. salinity [HHC06, NJJ06]. Salmo [ARLB07, BLKR09, BPD^{+}04, CK02, JLHO01, Jok02, JRML00, JJKN^{+}03, KFO^{+}04, Lan04, MNML00, RMRB06, SLE09, SBD^{+}05, SBD^{+}06, VP05]. salmoides [CSSO03, LDE^{+}02, WSC08]. Salmon [Mil02a, Mil06, AKK^{+}03, BNSA05, BNS07, BLKR09, Bro05b, BM03b, BM03c, BRRL09, BCM08b, CB00, CWL00, CMOM00, CK02, DSO^{+}05, FR03, Fra07, JLHO01, Jok02, JRML00, JJKN^{+}03, JJJ03, KNJJS06, KNSSKF09, KFO^{+}04, Kop06, LDK^{+}07, LRC^{+}08, MNML00, MMS09, MK09, MEA^{+}03, NW03, Nas03, NJE00, PB04, PIdlHGV05, PCS05a, QDV05, RDM^{+}09, RR03b, SN09, SSIE06, SBD^{+}05, SBD^{+}06, TNFF03, TNL07, VAYD04, VP05, WCM04, WIS03, ZB00, NW03, Sol05]. salmonid [HCHO^{+}03, SMD03, YJV^{+}03]. salmonids [ASM^{+}03, CG04, PMWC03, VBSK00]. salmonis [BCM08b]. saltatrix [AWY02, RRJ08, RRJ09, SB05, TOÖD06]. Salvaged [KPBD08, But08]. Salvelinus [EF01, HMB^{+}00, PCS05b, iTM04, WFJB09, YMT^{+}06]. SAM [FMGG08]. Sameoto [JDAM08]. Samoa [CGT08]. sample [AKA^{+}02, BG05, CCBG02, MFB08, TS09]. sampled [GRC^{+}09, HBW00]. sampler [RVS02]. samplers [JDAM08]. samples [OMR09, PC03, SLB04]. Sampling [JRG08, PP03, AMC^{+}08a, AR05, AKA^{+}01, AKA^{+}02, APP^{+}07, BdL05, BACP09, BP06, BZRO05, CSR06, CXMM04, CTB01, DFD02, GDDJ09, GDFH07, GCL03, HD00, HM00, JR07, JJWO09, JK07, KP07, KH05, LRC^{+}08, LBKG08, MK02, MFB08, MSSJ05, MC06, PKŘ^{+}08, PKŘ^{+}09, RYSN08, RW03, SCW04, SGP02, Smi03, USKK04, UMV^{+}09, WBCP08, Wes02]. San [BP01b, Gon06, GNC07, NS04, Nar06, NGF07, VLD^{+}00]. sand [HB01, MMN03, MCS05, NM06, YEA07]. sand-smelt [MMN03]. sandbar [JLC04]. sandeel [BHJ02, LNG04, MTRM05, SBD^{+}05, SBD^{+}06]. Sander [GWR07, PCS05b]. Sandperch [Gon06]. sandsmelt [TGS^{+}00]. sarba [RMBG02]. Sarda [CMC^{+}06]. Sardina [EBV^{+}03, QP00, RCGC04, RCG02]. sardine [Chi00, CBC02, CRC^{+}07, DBK07, DTSR07, GSMT04, GFM02, Gue05, GEYPC^{+}09, IFH^{+}09, Ker00, Man00, MAASMB05, MAdAMASCM09, QVNMGM00, RCGC04, SCM^{+}08, SAK05, SWA08, Vas03, WSS02]. Sardinella [EBV^{+}03, Gue05, ZCTA^{+}08]. Sardinia [ASM^{+}09, ADP^{+}08, TSP^{+}04]. Sardinops [CSCC07, GFM02, GEYPC^{+}09, Ker00, MAASMB05, MAdAMASCM09, MBGW02, MG03a, MG03b, QVNMGM00, SAK05, SWA08, WSS02]. sargus [PL02, FMB^{+}05, PA02]. Satellite [GCLA07, HJLL07, Loh08, PM09, San00]. satellite-transmitting [Loh08]. saturation [GBG^{+}03]. Saurida [YST^{+}02]. saury [TAS03, USKK04]. Savannah [JWR05]. saxatilis [JWR05, MRT03]. scabbardfish [PGS^{+}08]. scabra [PK06]. scad [JCS01, SE06b]. scalar [SR00]. Scale [RMRB06, ADP^{+}08, All07, BLC08, BP01b, CA07, DHM08, EGB^{+}06, GRFE06, GSH^{+}07, HC07, IMOO06, KK09, LD06, MBO^{+}07, MLH^{+}05, PBW^{+}09, RGGB07, RG07, SdPFAL09, SCSC07, SK03a, Sei07, SPS^{+}07, The07, TDK^{+}05, TSTK06, TSTK08, VMPP07, WLDH08, WG05, dGvZM^{+}06]. scaled [IWW03]. scales [ABC^{+}08, ASR05, BLR^{+}08, GRC^{+}09, GCC^{+}08, LMR^{+}05, LMR^{+}07, MMP^{+}02, SALnM07]. Scalet [McI09c]. Scaling [Cad03]. scallop [BBM^{+}06, BLH04, GD03, HLL^{+}08, Kat05, KCRT03, LK02, LLV^{+}08, NS04, OS05, RB04, SGRK09, Uki06, VL08]. scallops [MCJB02, MJB02, VL08, WHWM07]. scanning [PBVG07]. scantrol [ØT02]. Scaridae [TOBSJH06]. scaring [Løk03]. scarlet [New02]. scattering [KLJ06]. Schaefer [Mau03]. schedules [Cad03, Cro01]. schemes [CSR06]. schlegelii [GMY^{+}09]. Schneider [GOAARJRP09, New02]. school [DBB00, MMB^{+}07a, PBVG07, RSP^{+}00]. schooling [MC00a, STH03, ZDB^{+}07]. schools [BF00a, MC00a, NP08]. Schweder [Hol01]. Schweigert [Coo00a]. sciaenid [MDGR05, SGS09]. Sciaenidae [ABL02, HGF06, LC07, NV05, YTK06]. Sciaenops [JDS00]. Science [Ano01c, Ano01e, Ano09a, Lai00, Loc02, McI03a, McI09a, McI09b, McI09d, Mil02b, Mil03b, She02, Sol05, Sut02, Sym08c, CS08b, Cor02, Cor09, GAC^{+}06, IK07, KC06, KJS^{+}06, Lon07, Moo03, Sym06, WD02]. Scientific [Bar00, Win02, HP02, IWW03, MvdK06, YPDS05]. scientist [Mau03]. scientists [BHB^{+}04]. scnDNA [SF05]. Scomber [CMH01, FAL^{+}01, PVP04, SWA08, Tze04, VAF04, WY06]. Scomberoides [GFvdV06]. Scomberomorus [BSG02, CMAOAA05, GAOMC06, GAFA05b, MCAO^{+}05, vHMAO^{+}06]. scombrus [CMH01, FAL^{+}01, PVP04, VAF04]. Scophthalmus [NJJ06]. Scorpaena [LLM05]. Scorpaenidae [MNM00, MF09]. Scorpaenids [KKJ00]. scorpionfish [LLM05]. Scotia [BFM08]. Scotian [SSC^{+}03]. Scotland [Moo03, BWMA02, BRRL09, HKG^{+}08, YPM^{+}06, Ano01f]. Scottish [BPW01, DPS^{+}01, KOO04, LCTB^{+}03, NPC^{+}08, PB03, WGGN06, YPD^{+}04, ZPR01]. scuba [SG06]. Scyliorhinus [RDH05, RCSFO04]. Scylla [JPYW08, MMLT05, OIK^{+}06]. Scyllarid [CCD01, Ste03c]. Scyphozoa [PK03b]. SDSAM [SW09]. SE [GRFE06, GTB04]. Sea [BSM07, BNSA05, BNS07, BLV06, BK02, BMB06, CS04, CUT06, CPT06, Coo00a, Coo01, Eer04, Ell00, Fra07, GKUO07, Gra02a, GHDS08, HSSR06, Hor02, KVS04, KPL^{+}09, Lot07b, McI01, McI02c, McI05b, McI06b, MBG^{+}07, Moo03, MFAM05, MAG^{+}09, Ore03, PBDP00, RBD02, SPB^{+}03, She02, SE02, VYMGM^{+}09, ABFV02, AFJC08, ASM^{+}09, Ano00e, Ano01f, BTC08a, BW02, BRRL09, BPD^{+}04, CH03, CXCD07, DMC^{+}09, DOQ08, FML^{+}07, GCB^{+}06, GMY^{+}09, GC04, GTB04, HLL^{+}08, JDMN00, JJKN^{+}03, KNJJS06, KCH05, KLS^{+}07, KFO^{+}04, LK02, Lan04, LRW09, MFB08, MMN00a, NHT07, OMW^{+}03, PH09, PPS05, PBLP08, PZH02, PKP04, Pou01, PL06, PK06, RKEJR06, SdPFAL09, SE06a, SCW04, SPS^{+}07, SLE09, SMS02, SK08a, SKKP05, VBSK00, WHWM07, WK08, YKO09, ZP05, AMC^{+}08b, ASR05, AE03, Arc03, BF00a]. Sea [BQ04, BVH^{+}08, BWMA02, BHJ02, Bjø01, BG01, BLKR09, BMB05, CMC^{+}06, CNH09, CCM^{+}09, CTK^{+}07, CAM^{+}08, COHP02, CP08b, CHL^{+}06, CC01c, CLH^{+}08, CL04, Cor01, CCBG02, DCM^{+}05, DMA07, DS07, DG08, DA07, Din05, ERG09, FGJD08, FZHL09, FAL^{+}01, GDAM07, GLH07, GTB04, HKG^{+}08, HMH^{+}04, HF07, HH01, HMM^{+}09, HK02, HLH08, IYI07, JLHO01, JR02, JNDH00, JDMN00, JPVKS^{+}01, JJKN^{+}03, KA05, KNJJS06, KNSSKF09, KSS05, KTL06, Kno07, KM08c, KPL^{+}07, KHMG06, LD08, LDGA09, LDK^{+}07, LOT07a, LAL07, LNG04, LFA08, Lot07b, LBMCE00, MVV^{+}01, MMV^{+}04, MH01, MCL04, MPQ07, MTU^{+}04, MOK^{+}05, MOGG06, NDCK09, Nee08, NS06, NJJ06, OKF06, OGK^{+}08, OAKM07, PA02, PK03a, PMAK07, PTPS09, PSS^{+}07, Pol00, Pol02, PDV05b, PCHM07, RPP04, RCGC04, RdFBS01, RMF^{+}06, RH04b, RDH06, RCD07]. Sea [RCSFO04, SSB03, SLV^{+}08, ŠV04, SCM^{+}08, SRH^{+}02, SME02, SWC^{+}05, TH03, TP06, Tia09, TL03, TMN07, UM02, UVKS05, VBSC08, VKD03, WLY06, WLL08, WPOB09, XJ05, YST^{+}02, YFT^{+}02, You03, Zhe03, vSS05, vSSK07, vSS09, McI02c, Mil08]. sea-trial [KLS^{+}07]. seabass [APS07]. seabed [KCRT03, OSB09, PBW^{+}09, SdPFAL09, vSM02c, vSM02a, vSM02b]. Seabird [WR02, APM^{+}09, Løk03, PGF07]. seabirds [BNL^{+}08, Ein09, SB03, Sva05]. seabream [CPT06, DSPLR06, FMB^{+}05, JTdH06, PLDS03]. seabreams [ABC^{+}08]. seagrass [GCL03]. seahorse [MHK^{+}06, WMS04]. Seahorses [VMPP07]. Seal [Güç08, HFL08, KSS05, SSK^{+}06, WR02]. seal-induced [KSS05, SSK^{+}06]. seal-safe [HFL08, SSK^{+}06]. Seals [Moo03, Güç08, MKB01]. Seaman [How03a]. seamount [FBV06]. seaperch [New02]. search [vOPvDM02]. Searching [Ste03a]. seas [ERG07, FSB^{+}02, HJ06, IMOO06, SKMT02, TOÖD06]. seasnakes [HGW^{+}08]. season [MOGG06]. Seasonal [APC07, AFJC08, ASA04, AS04, CMH01, GJBY05, JLHO01, LR00a, MPE04, MLC03, MR05, MP03, RCGC04, SE06a, SR06, SC07, SB01b, ACMMCEGM02, BHH08, FÖB08, FOGKT08, GR07, MSA00, ÖW02, RSST07, SAG05, TSOG01, XM00, tHDC06]. seasonality [AB09b, GMV00, GMV01, HMAOAS07, TVY01, vZRMvD02]. seasons [AGB07]. Seattle [Mil02a]. seawater [HBJ03]. sebae [NCW00b, vHAN^{+}09]. sebago [JRML00]. Sebastes [GR02, JNDH00, JDMN00, KKJ00, PD08, PWHM06, RFWH09, SKKP05, WKCC06]. Sebastolobus [KKJ00, LWS04, LIW04]. Section [BF00b, CT00, MG00, MMN00b, BNS00]. sectioned [BG04, Dou04]. sector [MLP^{+}08]. sectors [OCT^{+}03]. seeding [DDHS06]. Seeliger [Ano00e]. SEFOS [Rei01]. segregative [WT00]. Seguro [CdCBdR03]. Seine [GCL03, BAEF06, CDC03, GLK00, GK03, Kar04, KLS^{+}07, MB07, MBGW02, ORVMP06, Par04, RT07, Vas03, vZRMvD02]. seiners [MROGVR09, RT07]. seines [BWM07, MB00]. seining [GKH^{+}01, GKH03]. seismic [PG06, SHDO04]. selected [SGVM02]. Selecting [CPSBRD^{+}02, KJS^{+}06]. selection [APC07, BLC08, BLH04, BWM07, FCM07, GF06, GBM09, GSL08, GLS^{+}09, HC00, Hal02, Her05b, HPK07, HS00, JIGI06, Kat06, KME^{+}02, KHMG06, MGS02, OMW^{+}03, PLSI09, SLPF08, TS09, VL08, WL06a, WB08]. Selective [BP01b, CAH^{+}03, RDH06, BVH^{+}08, DWW^{+}02, FHM09, GLH07, MTHL06, MAG00, RH04a, Rue07, SSIE06, VU08, VAYD04]. selectivities [ÖT03]. Selectivity [BdL05, BMK^{+}04, DBAÖ06, GOF^{+}04, GLS^{+}09, He07, MS05a, MSB^{+}08, Pol02, TASF09, iTM04, ABCB^{+}07, APP^{+}07, BSS06, Bet04, BACP09, BP06, Bro05b, CFE02, CFE03, CFH03, CC03a, EH04a, EGB^{+}06, FSB^{+}02, FÖB08, FCG02, FMCS05, FHG03, Gas03, GK01, GKF03, GLK00, HMM00a, He06, Hen09, Her05a, Her05b, HO05, HPK06, HO06, HFHO07, HKF^{+}09, HWM02, IJ06, KAHR02, KF04, KI04, KI05, KOO04, LIW04, LMMD06, MMB^{+}07a, MGTB08, MHMP01, MH02, MHF02, MBM04, MP03, OKF06, OGK^{+}08, PMAK07, Pol00, PCV04, PSPCVCD09, PKŘ^{+}09, PLMDB06, RBD02, RCA^{+}07, SLB07, ST00, SB07, SGL08, SHL09, SGB00, SE02, SF03, Sve15, TS09, TÖT04, WMTJ07, WBCP08, WBHG01, YEA07, vM03]. selectivity-improving [PCV04]. semelparous [HH06]. semi [EGB^{+}01, PGF07]. semi-pelagic [EGB^{+}01, PGF07]. semifasciata [Gon06]. Séminaires [Gue05]. semisulcatus [NAD^{+}07]. senegalensis [VFCC06]. sensed [WYR06, WGR08, WPS^{+}07]. sensing [PM09, San00]. sensitive [GNL^{+}09]. Sensitivity [Cad09, PCS05b, UM02, CHSA02, MK01]. sensor [RDM^{+}09]. sensors [WRWK02]. Separate [OKF06]. separating [GFSM07].

Separation [GF06, SBF^{+}08, Gra03, SRTP06, SMSZ08]. separator [CF04, KMK09]. Sepia [DR01, RPFR06]. sepioidea [RW07, ZW09]. Sepioteuthis [RW07, SLJ07, ZW09]. September [Ano00s, Ano01k, Ano02v, Ano03-28, Ano04t, Ano05t, Ano06y, Ano07v, Ano08q, Ano08v]. sequence [RMVJGMMV07]. sequential [JC04]. Series [Van01, GRFE06, GKV06, SHBÅ07, TC09]. serranid [BMBM05]. Serranidae [MSA00]. Serranus [MSA00, TT01, TRL06]. service [BRRL09]. set [CATH04, How06, TS06]. set-net [TS06]. setiferus [DKKS07]. sets [BCMM09]. setting [Løk03, SB03]. Settlement [WM04, ARCLBPdP07, BM09, MR08a]. Several [Xia06, RDH06, SK03b]. severe [CS04]. Sex [AB09b, MGJB09, RR03a, SWPY05, XK04, Ber07, KTL06, KCK03, MGNB05, WSPY05, WSPY07, Wie04]. Sex-biased [MGJB09]. Sex-specific [SWPY05, WSPY05, WSPY07]. sexfilis [CGLZ04]. Sexual [SKP03, AB09b, BP01a, MBR06, XK04, YTK06]. Seychelles [GC03b, RG07, RE07]. SG [Coo00a, Coo01, McI01, McI02c, Ore03]. shad [RC06, SMRH04]. shallow [APS07, BCMM09, BMW07, CTC^{+}06, CPR07, DK05, ESB09, GCL03, HP00, KH05, dSBAC06]. shallow-water [dSBAC06]. Shape [BBK08, NBF09, DGP02, FCC^{+}09, Her05a, Her05b, KAHR02, MDGR05, MPA08, NP08, PPL04, PBM06, SMSZ08, SBF^{+}08, TRL06]. shared [DTSR07, HGSP01]. sharing [MMZ09]. Shark [Ano07w, PWP02, YKM06, dSHG08, BAH09, CL09, CCB02, CE07, HMAOAS07, HWW^{+}09, HSW09, JLC04, LSH04, LCNJ06, MWF06, MLCGRV07, MGNB05, NS07, PM09, PWTP00, Qui07, RCGMQV05, RDEACT07, Row07, RE07, SDNL09, SDH00, Ste07, TS09, TÁTWD09, WK01, WKM02, WLDH08, WC07, BKD^{+}02]. sharks [CMJ05, CAM^{+}03a, CC03a, CY06, CARA^{+}07, GCS^{+}06, GCLA07, GNL^{+}09, GR07, HJLL07, IK07, Jak01, JH07, JCLL08, Mar07b, NE07, RMVJGMMV07, RG07, SHBC02, Tay07, TF09]. sharpnose [MGNB05]. sharpsnout [DSPLR06]. sharptail [LLJC09]. sharptooth [BW08]. shear [BGM^{+}09]. shelf [AAFLP^{+}07, BH01a, MHS^{+}05, PLDS03, PMG^{+}06, RWT01, Rei01, SWC^{+}05, TT01, UL01, VE03, WPS^{+}07, WDC02, BQN05, GYMdL^{+}06, MMH05, SSC^{+}03, WNRd00]. shell [DG08, LLV^{+}08]. shellfish [HCM05]. shells [Cer00]. sheltering [SK08a]. Sherman [Ste05, Lon03a]. Shetland [BGH05, LJP05, MLJT09]. Shields [BGH05]. Shift [SB08, GB03, MSA00, RdFBS01]. shifts [TMN07, UM02]. Shirley [Ore03]. shoaling [BGM^{+}09, BMPK06, PTE09, WNRd00]. shooter [Løk03]. shooting [SHDO04]. shoreline [ZR08]. Short [AWY02, CGEHC07, BR05, BC04, CRG^{+}02, DACS08, DDP^{+}06a, DDP^{+}06b, DDP^{+}06c, Lap02, MF07, TSJ04]. short-fin [Lap02]. short-lived [DDP^{+}06a]. Short-term [AWY02, CGEHC07, BC04, CRG^{+}02, DACS08, MF07, TSJ04]. Shorter [Ano02-27, McI02b]. shortfin [BQN05, CMJ05, CL09, HH06, RCGMQV05]. should [Cer00]. showed [SSH^{+}06]. shows [WLDH08]. Shrimp [Gri06, ASIW05, BPW^{+}03, CHN08, CFE02, CA03, CGP05, CAHAF07, CHDGM06, DBAÖ06, DOQ08, DKKS07, FP00, GHS04, GC03a, Gra03, Gri03, GWB07, GL05, HJ03, HJ07, HJ06, JR02, KCK03, LYKH06, LMASMASZ02, LMASHV^{+}03, LMRQNM^{+}05, MVAMB07, MH01, PK03a, Pol00, Pol02, PCV04, PDV05a, PDV05b, PSPCVCD09, RBD02, RRAS03, RRASLB03, RH04b, RH06, RW03, SGB00, TY02, VD06, WZDY06, Wes02, Wie04, WSP06, YAM00, dSBAC06]. shrimp-trawl [MVAMB07]. shrimps [ALN03, BYA^{+}08, CPCD05, FFM01, LF02, PCD03]. Shuttleworth [Ano00b]. Sicily [PBDP00, RBD02, GBFG06]. Sicyonia [LMASMASZ02]. side [Koe03]. sidescan [Tre01]. Sierra [Pie09]. sieve [PCV04, RH04a]. sieves [PLD05]. Siganidae [BK02]. Siganus [BK02]. signal [ET00]. signal-to-noise [ET00]. signals [ET00, WG05]. signature [WF07]. signatus [SDNL09]. Significant [GM07]. significantly [ET00]. silky [JCLL08]. silver [HC00, IMSS06, IOS07, Kir01]. silver-biddy [IMSS06, IOS07]. Simojoki [Jok02, JJJ03]. simple [Bet04, CHL01, HR09, KFM08, Smi03, UMV^{+}09]. Simplified [LZS01, PAF^{+}08]. simulate [MG03a, MG03b]. Simulated [KM03, GFSM07, Lab05, Mes03]. Simulating [CQJX08]. Simulation [HFHO07, HPK07, Jun08, AGB07, CXCD07, CXQ^{+}09, CHDGM06, KLS^{+}07, QDH09, ULDG02]. Simulation-based [HFHO07, HPK07, Jun08]. Simultaneous [GFMO06, PM09]. Sinclair [McI03b]. single [FSPWG08, GFM02, MGR^{+}02, ÖT03, SET^{+}09, SRA^{+}02, SF05, YSS^{+}09, vSM02c, vSM02a, vSM02b]. single-pass [SET^{+}09]. single-species [FSPWG08]. single-warp [SRA^{+}02]. sink [RMHW03, SC08, WIJB08]. SIT [KMK09]. site [BLC08, Loh08, MCL04]. sites [JIDM02]. situ [ABRB00, AC00, Dun06, IV08, MC00b, PMT06, VMSR08]. situation [SE06b]. six [KLJ06]. Size [CFE02, CFE03, CFH03, EGB^{+}06, FSPWG08, FWG15, Gas03, GSL08, HC00, JIGI06, JCJ^{+}09, KME^{+}02, MGS02, MGNB05, NS07, PMAK07, PKŘ^{+}09, SLPF08, SALnM07, SGL08, Sve14, WBCP08, BB04, BACP09, BLH04, Chi00, CSCC07, CdML^{+}07, DWW^{+}02, Ehr08, GMY^{+}09, GK01, GR07, GLK00, GFvdV06, GCJ07, Hal02, HSSR06, Hen09, Her05a, Her05b, HPK07, HWM02, HS00, JLC^{+}03, JPYW08, Jon00, KTL06, KLU01, KCK03, KHMG06, LIW04, LRBH07, LPHH09, LCTB^{+}03, LCFP^{+}08, LMRQNM^{+}05, Lor06, LPG^{+}09, MMB^{+}07a, MPM^{+}06, MHF02, MWB^{+}08, MN03, MBR06, MM05, MBM04, MC06, NP08, OKWM05, PWB05, Pun03, RCA08, RMN08, RKEJR06, SN09, SLB07, SHKB09, SB07, SGB00, SOS02, SE02, SFCK03, SLJ05, Sve15, TS09, TSJ04, UMV^{+}09, VL08, VN00, Wie04, WSP06, WN03, XK04, Xia06, YFL08]. size [YPD^{+}04, You03, FWG15]. size-at-age [GMY^{+}09, Jon00]. Size-dependent [JCJ^{+}09, Lor06, Xia06]. size-specific [LRBH07]. size-structured [MN03, Pun03]. sized [IV08]. Sizes [DV00, MMV^{+}04, AKA^{+}02, HMM00b, KK04, MvDM02, TSTK08]. Sizing [PK06]. Skagerak [MS05a]. Skagerrak [CS04, FHM09, KFM08, PCHM07]. skate [GHDM08, Ger09, MK06]. skates [CGP05, ESC08, ECEG09]. skewed [TB04]. skipjack [AC02, AS04, TSOG01]. skipper [RT07]. Skippers [RT07]. sled [LWS04, LIW04]. slide [ET00]. sling [GDFH07]. slope [CAFT00, FFM01, HLHF03, MMM^{+}03, vSM02c, vSM02a, vSM02b]. slopes [JNDH00]. Slow [Lon02a]. Small [All07, JJWO09, The07, BMF^{+}10, BCLP00, BP06, BP01b, CNH09, CC03a, CBGP06, EGB^{+}06, GRFE06, GSH^{+}07, HC07, HP02, IV08, KJ08, LD06, Mat07, Mey07, PR03, RGGB07, SCSC07, Sei07, SPS^{+}07, SALnM07, SKMT02, SE02, TDK^{+}05, TSTK06, TSTK08, VMPP07, dGvZM^{+}06, vOPvDM02]. Small- [The07]. small-hook [SE02]. Small-scale [All07, BP01b, EGB^{+}06, GRFE06, GSH^{+}07, HC07, LD06, RGGB07, SCSC07, Sei07, SPS^{+}07, TDK^{+}05, TSTK06, TSTK08, VMPP07, dGvZM^{+}06]. smaller [MvDM02]. smallest [Jen02]. smallmouth [HCH09, NGM^{+}09, WSC08]. smelt [JAKM05, MMN03, MCS05, PMT06, SRTP06]. Smith [CND02, GR07]. Smitt [OFOM08]. Smolt [WCM04, BPD^{+}04, CG04, KNJJS06, KNSSKF09]. Smolt-to-adult [WCM04]. smolts [DSO^{+}05, SLE09, SBD^{+}05, SBD^{+}06]. smooth [CE07]. smoothing [MLCGRV07]. snail [GCN02]. Snake [WCM04]. snapper [AFSF05, All07, ASAGR06, BLB07, Bur02, EF02, GBL^{+}05, GAF06, HMM00b, HMM00a, ML04b, MR08a, NCW00a, Ste08, SJ05, SOKS08, WMTJ07, WBCP08]. snappers [FFL09a, FFL09b, NCW00b]. snorkeling [WN03]. snow [BFM08, CWS^{+}05, HMM^{+}01, MD09, SWM09, Zhe03]. Social [Mil03b, CFG05, GAC^{+}06, Mar07a, SW09, Sym06, SP09]. Society [Lon02a, McI02a, McI05d]. Socio [AM03, GLCC09, Kro04, KPL^{+}07, KPL^{+}09, RE07, TCS05]. socio-bioeconomic [KPL^{+}07, KPL^{+}09]. Socio-economic [AM03, GLCC09, Kro04, RE07, TCS05]. Socioeconomic [BDBD09, LZCR08]. Sociometric [AAFLP^{+}07]. sockeye [LRC^{+}08, MMS09, RR03b]. sod [GHDS08]. Sofala [dSBAC06]. Soft [McI07a, CF07, Gue05]. Softback [Sym08c]. solar [AOMK^{+}04]. sole [Cro01, vM03]. Solea [PTR^{+}05, VFCC06]. soles [VFCC06]. solida [CCG^{+}02, CCA^{+}03, Gas03]. solutions [SCP01]. Solving [Gra02a]. Solway [LF02]. somatic [BBJG04, LP00]. Some [CUT06, GB06, PGMM01, Dah00, FKBF05, GAOMC06, GK01, Hat00, HP04, KCKP02, Mag01, MPQ07, MTA^{+}02, RH04a, SLB04, SND^{+}08, UM02]. sonar [Mat07, MC00a, PBVG07, RSM05, Tre01]. Sonora [LMASMASZ02, BSMF^{+}09]. sorting [EH04a, GSL08, KME^{+}02, MGS02, MB00, Pol02, SGL08]. Sound [YTK06, KTC00, NLJS01, QDV05]. sounder [RSP^{+}00]. source [SC08, WIJB08]. Sourcebooks [Sym08c]. sources [Kop06, Pal08]. South [AHS^{+}06, BW02, CF04, CWJS09, CARA^{+}07, Dem04, GKH03, GLCC09, OSG08, RMBG02, RB09, SF03, Sva05, BGW06, CFE02, CFE03, CPW^{+}06, CE06, EBV^{+}03, EL09, GBC^{+}03a, GBC^{+}03b, GBM09, GCMJ08, IJD04, JIDM02, LE01, MHS^{+}05, Mos00, MUO02, PCAS04, PSTM06, SLA00, SGS09, SDH00, SST^{+}08, UL01, YFL08, AC02, BBJG04, BGB06, CSM^{+}09, CUT06, CMGB01, CP08b, CHL^{+}06, GSA^{+}06, GB06, GMS06, HBBG06, HE08a, HGF06, HGSP01, JDS00, LPHH09, LFA08, PH09, RB00, SLA00, SPS^{+}07, SLJ07, SSG06, SEPY02, SRS08, VL09, WLS07, WSS02, YFL08]. south-east [IJD04, SLA00]. south-eastern [BGW06, EL09, GBM09, Mos00, MUO02, PSTM06, SGS09, SST^{+}08, CSM^{+}09]. south-western [SDH00, CUT06]. Southbank [Kes00]. Southeast [GSA^{+}06, SSG06, ABD^{+}09, Bar00, MHS^{+}05, OLN^{+}02, SWT09]. southeastern [ASAGR06, CC03a, CMOM00, ML04b, MGNB05, RM05, dSH01]. Southern [But08, CC09, KPBD08, McI02d, MB08, PWP02, ACMMCEGM02, ASD09, BFM08, BSWL06a, BBL07, CMC^{+}06, CB00, CE07, CCCC04, CBC02, CRC^{+}07, DS07, EGB^{+}01, EGB^{+}06, FDG^{+}07, FOGKT08, FCM07, FHI07, GAC^{+}06, GGVLSR03, GSH^{+}07, GAFA05b, GAFA05a, GAF06, GWR07, GCC^{+}08, HSGF06, HCS^{+}03, HC09, HMM^{+}09, IWBJ08, JR02, Jun08, LPHH09, Lot07b, MMH05, MC04, MF09, OFOM08, PdAO02, PW05, PPdA05, QGBS04, RRAS03, RRASLB03, RBHP03, RAHHGM^{+}02, SGVM02, SSGM03a, SSGM03b, SLJ07, SMS^{+}06, TSM03, VG01a, VH06, VOS01, WLL08, WSS02, WIJB08, WN03, WH08, XK04, XJ05, ZFJ03, dSH01, ASM^{+}09, APS07, CAM^{+}08, GAFA04b, GAFA04a, SR07]. southwest [AC02, AS04, Ano00e, CFH03, GYMdL^{+}06, HHTE05, MWF06, MGJB09, WGR08]. Southwestern [VYMGM^{+}09, ABE^{+}04, CPCD05, CLH^{+}08, GCS^{+}06, MD02, PH04b, SWM09, Tia09, SDNL09]. Sp [Ano01c, HFBM08, YST^{+}02]. Space [ZFJ03, STH03, ZP05]. spacing [Gas03]. Spain [BPRM00, FRGF07, GRFE06, MF03, Pie09, SPB^{+}03, SGS02, SGB00, SBGG^{+}05, TSM03, AKA^{+}01, PIdlHGV05, PVP04, VAF04]. Spanish [GRPLE05, GAFA05b, HT05, JTdH06, JSR04, PCB01, QP00, SGS02, VL09]. sparid [CMGB01, MC06, PMG^{+}06]. Sparidae [Arc03, BG04, BG05, CSDP02, PL02, PA02, RMBG02, WMTJ07]. Sparids [GAFA04b]. Sparisoma [AMS08, TOBSJH06]. sparse [AIC08]. spat [ABAA^{+}07]. spat-fall [ABAA^{+}07]. Spatial [AMS08, AMUE01, BF00a, BH01a, BM03a, CPCD05, Cro01, DSO^{+}05, Daw08, GSA^{+}06, GRC^{+}09, HSGF06, HF07, HS03, KCKP02, LAL07, LRBH07, LPHH09, Man00, MCAO^{+}05, OQE03, OLN^{+}02, PMSD06, RLSJIA07, SLV^{+}08, SWT09, TSOG01, WMB08, WMD07, ZPR01, ASR05, All07, BRL^{+}05, BR01, BP08, BLR^{+}08, CXQ^{+}09, CRG^{+}02, CPR^{+}08, DMA07, DDP^{+}06a, GJBY05, HE08a, HMH^{+}04, Joh03, KA05, Kat05, LK02, LMR^{+}05, MBT08, MLC03, NC04, Ove02, Pun03, SK03a, SC08, SALnM07, VBSC08, WGR08, War06, WRWK02]. spatially [But03, Cam04, CP08b, DKH08, GPI04, JL00, MMZ09, PPGD08, PWTP00]. spatially-explicit [But03, JL00]. spatially-structured [PWTP00]. Spatio [OAKM07, TVB^{+}09, TOBSJH06]. Spatio-temporal [OAKM07, TVB^{+}09, TOBSJH06]. spawner [BLR^{+}08, Kop06]. spawners [RR03b]. Spawning [BBL07, CRC^{+}07, LC06, LCNJ06, MM06b, QP00, ANS^{+}01, ASD09, AB09b, BM03a, CMAOAA05, CSCC07, CPR^{+}08, EH04b, EBV^{+}03, FGOF08, FUM02, HH06, JS02, JIDM02, JHGS04, KKR^{+}05, MCB^{+}07, NV05, PBM06, RR03a, SLA00, SSH^{+}06, SKP03, SEPY02, SWPY05, TVY01, TNL07, VOH02, VBSC08, WSS02, WGGN06, YFT^{+}02, YSS^{+}09, ZC02]. Spear [CLGR^{+}04, Mey07, Tia09]. Spearfishing [LZC^{+}08]. Special [McI05d, Ano07x]. specialised [AM03, RCA08]. Species [HMAOAS07, McI03a, McI05d, TGH^{+}08, VU08, AAP^{+}09, ASIW05, Arc03, BTC09, BCMM09, BH01b, Bro05b, BRRL09, CFE03, CFH03, CSDP02, CLGR^{+}04, DNN^{+}03, Dah00, DDP^{+}06a, DDP^{+}06b, DDP^{+}06c, FKW06, FSPWG08, FKBF05, GR00, Gra02b, GJBY05, GPI04, HWB06, He06, HBW00, HBKJ07, HP04, HP09b, JS02, JSR04, JCJ^{+}09, KAN00, KAO04, KVS04, KKR^{+}05, KG06, KA06, KM08b, LRG^{+}00, LD01, MMV^{+}04, MTHL06, Mag01, Mai09, MH08, MWB^{+}08, MSVF05, MC05, MBG^{+}07, MAL^{+}01, MD01, MM06c, NS08, NGF06, NS06, OA04, OLN^{+}02, PD08, PGK03, PKŘ^{+}09, RMN08, RMMP06, ROS04, SLPF08, SGVM02, SCLM06, SC07, SW00, SME02, SMSK03, SMS^{+}06, SD01, SWA08, TH03, ULDG02, UVKS05, VG01b, VL09, VFCC06, WKCC06, WSS06, YKO09, You03, dSBAC06]. Species-selective [VU08, MTHL06]. Specific [CSDP02, Boo04, LRBH07, MMB07b, PCS05b, SWPY05, TH03, VMSR08, WSPY05, WSPY07, WMD07]. speciosus [GAFA04a]. spectra [CdML^{+}07]. spectroscopy [Dia00]. speed [ABRB00, AC00, CXMM04, DWW^{+}02, JSO08, MJB02, MC00b, OKWM05, RGSB08, RVS02, SW01, WSM02, YEA07, vSM02c, vSM02a, vSM02b]. Spencer [Ste03a, Sva05, SRS08]. sperm [PSS^{+}07]. Spey [BRRL09]. Spicara [DKGC00]. spider [CF07]. spill [RR03b]. spillover [FFM08]. spinal [TGS^{+}00]. spinefoot [BK02]. spines [LDN04, MF00]. spinifer [GAFA04b]. Spinola [CCM^{+}09, DALP01, GRPLE05, LD08]. Spiny [WVW^{+}08, ARCLBPdP07, BRL^{+}05, BMF^{+}10, BSWL06a, But03, DS07, Ehr08, Fri07, GQR03, GYMdL^{+}06, Gro00, GMS06, KS08, MF07, PVMdL05, RLSJIA07, RLGLO09, SK08b, TB04, TSP^{+}04, Vel03]. Spisula [CCG^{+}02, CCA^{+}03, DG08, Gas03]. splendens [DG02, RLG^{+}01, SÁR^{+}05]. split [ABRB00, AC00, Kru04, Mat07, MSSJ05, MC00b, Sch09]. split-beam [ABRB00, AC00, Kru04, Mat07, MSSJ05, MC00b, Sch09]. sponge [MSM^{+}08]. sponges [WDC02, Ano00d]. sport [OGPDOM08, VLD^{+}00]. sportfish [GFvdV06]. spot [LPP02, RPC^{+}06, TCY01, ZR08]. spotted [ASAGR06, GAFA05a, RDH05, RCSFO04, iTM04, iTMI06, Tze04, YMT^{+}06]. spotters [RT07]. spp [CPW^{+}06, CCD01, DACS08, GR02, GT04, LWS04, LIW04, MROGVR09, MFC^{+}08, RT07, SPB^{+}03]. sprat [BVH^{+}08, HHC06, KHMS09, KK04, VKD03]. sprattus [KK04, BVH^{+}08, HHC06, KK04, VKD03]. spread [vSS05]. spring [BVH^{+}08, CCT^{+}08, EH04b, FGOF08, FUM02, KTC00, VOH02, VAYD04]. spring-born [BVH^{+}08]. spring-spawning [EH04b, FGOF08]. Springer [Ano00e, Ell00, McI05a, Van01]. Springer-Verlag [McI05a]. spurdog [BGW06]. squalid [CCB02, Jak01]. Squalus [BGW06, DS07, HFB^{+}06, MF07]. Square [GK01, GKF03, OKF06, BSS06, BKD^{+}02, BMK^{+}04, BNL^{+}07, CFE02, CFE03, CF04, CRD06, EH04a, GFSM07, GOF^{+}04, HC00, He07, KFM08, MMB^{+}07a, RCD07, SLPF08, TASF09, BR00]. square-mesh [BKD^{+}02, BMK^{+}04, CRD06, KFM08, MMB^{+}07a, RCD07, SLPF08]. Squatina [BAH09, VMQ08]. Squatinidae [VMQ08]. Squid [MKB01, AR01, ARVC01, AM02, BQN05, BPW01, Bjø01, BM05, BI05, CC03c, CCT^{+}08, CLC08, CLH^{+}08, GB06, GMR01, GMR02, Hat00, HH06, IMOO06, Lap02, MSN01, MQVSN04, Mar06a, Mar06b, MOK^{+}05, MOGG06, NMHHMB^{+}00, NMMTCV^{+}06, NNA01, NNA02, ORVMP06, PdAO02, PB03, RW07, Rod01, RWMBHH06, RB00, SCUADJ07, SLA00, SN01, ŠV04, TVY01, TYM^{+}01, Tia09, WLL08, XBS^{+}07, YM00, YPD^{+}04, YPM^{+}06, YPS^{+}06, ZW09]. Squires [Sym08b]. Sr [MK09, MK09]. Sr. [Skr08]. Sr/ [MK09]. Sr/Ca [MK09]. SrCl [CS07b]. Sri [HJ06, DHM08]. St [BFM08, CWS^{+}05, HMM^{+}01, MMH05, SWM09, SSC^{+}03, XM00, Xia04a]. Stability [DVAL^{+}09, CBRN08]. Stable [GB03, GJSW04, BBM^{+}06, GMP05, WM00]. staff [AR05]. stage [MAASMB05, RRAS03]. Stages [How06, CCMS^{+}03, DKKS07, GM06, GBL^{+}05, IWW03, LDGA09, TNL07]. stake [ABCB^{+}07]. Stakeholder [NM06, CK08]. stakeholder-driven [CK08]. Standard [How02, RSP^{+}00, JI07, PWB05]. Standardisation [SG01a, Cam04]. standardised [LP06, MS03]. Standardization [PWTP00, BQ04, ML04a, OA04, Par04]. standardized [CNH09, PKŘ^{+}09]. Standardizing [MP04, QPR08, SYS^{+}08]. standing [JLC^{+}03]. starch [WVW^{+}08]. starch-iodide [WVW^{+}08]. Starkey [McI08]. State [How03b, Bra04, MN01, OMW^{+}03, ZP05]. States [McI02e, GOAARJRP09, CC03a, GHDM08, RRJ08, RRJ09, SWT09, WD02]. Static [PCS05a, Paw03]. station [PWP02]. stationary [ABRB00, AC00, BMW07, CA07, KH05, MC00b]. stations [KP07]. Statistical [CG04, Fab06, MSSJ05, BR00, FL05, FP00, HR09, LB08, MUO02, PBR05, WL06b, WB08]. statistics [GJ09, PKT^{+}02, SABW09]. statolith [MOGG06]. Status [SMM07, VH06, CP08b, DR01, JCW05, MSM^{+}08, NS07, OMF^{+}03, PCB01, PP01]. Steele [ZR04]. steelhead [WCM04]. steerable [Tre01]. Steindachner [KR00]. steindachneri [BSMFH07]. Steller [AFJC08]. stenolepis [GB03, KF02, Loh08]. step [vSSK07]. Stephanolepis [MRCH04]. stereo [Dun06, HFS02, HCS^{+}03]. stereo-video [Dun06, HFS02]. Stevens [DJGC09]. Stevenson [McI02a]. Sthenoteuthis [XBS^{+}07]. sticks [HHTE05]. still [FWG15, TMK^{+}06, WFJB09]. still-picture [TMK^{+}06]. Stimpson [LMRQNM^{+}05]. stimulated [JGB^{+}06]. stingray [Ism03]. Stizostedion [Eer04, JRML00]. stochastic [DFD02, TPC05, TÁTWD09, ZC02]. Stock [AMC^{+}08b, ABE^{+}04, BC02, BHH08, CL09, Coo00b, Cur00, FZHL09, Gro00, IMOO06, KCRT03, LNG04, LMASMASZ02, MPA08, NDCK09, SDH00, Uki06, WZDY06, WG05, ZK03, AMC^{+}08a, AGS^{+}08, ASM^{+}09, Arc03, BBRS06, BM03a, BBLL06, BRL^{+}05, BMBM05, BG01, BQ06, BM09, BLB07, BTC08b, Cad02, Cad09, Cam04, CMGB01, CH03, CAHAF07, Coo00a, Cor02, CRC^{+}07, DBK07, DTSR07, EBA^{+}03, Eer04, FH03a, FML^{+}07, FBS02, FR08, GAC^{+}06, GGVLSR03, GFM02, GG06, GCD^{+}09, GYMdL^{+}06, GAFA04b, HK06b, JCW05, KCH05, KK06, Kno07, KHM^{+}09, LC07, LM09, LYKH06, Lon09, LDA01, MCM^{+}08, MPE04, MWJ04, MN03, MFC^{+}08, Mau03, ML04a, MK01, MR05, Mes03, MBCMNMHH01, MBK02, MGJB09, OIK^{+}06, OB02, OS05, OFOM08, PMM05, PH04a, PPdA05, PMSC03, Pun03, PSTM06, QPR08, RPP04, RFWH09, RKEJR06]. stock [RCSFO04, RPFR06, SWM09, SRH^{+}02, SA03, SMSZ08, SAK05, SOKS08, SEPY02, SWPY05, TLD06, TdlMH^{+}03, VN00, VH06, Wah03, WLY06, WL06a, WMdS09, War06, Wie04, YLD^{+}05, YPD^{+}04, ZDB^{+}08, ZC02, Zho07, McI05b, She02]. stock-recovery [MN03]. stocked [OIK^{+}06, SSIE06]. stocking [JJJ03]. stocks [Aba08, BCLM08, CCH^{+}00, CC03b, CTK^{+}07, DGP02, DDHS06, FH03a, FFL09a, FFL09b, GM07, GMR01, GMR02, HFBM08, HCHO^{+}03, Hor05, HGSP01, IPH05, IS07, JLC^{+}03, KM03, Kop06, KIZ09, KF04, Lon07, MPQ07, MMH05, PF00, PH01, RL03, SÁR^{+}05, SKMT02, TOÖD06, UM02, ZP05, ZK03, vHMAO^{+}06]. stomach [AB08b, OFOM08]. stomachs [AB08a]. stop [Koe08, LBKG08]. storage [GM00]. straddling [Lon09]. Strait [CC09, ZP05, GBFG06, RBD02, YPDS05, ZB00]. stranded [dSH01]. strandings [PSS^{+}07]. Strangomera [CBC02, CRC^{+}07]. straps [HPK06]. strata [MBG^{+}07]. strategic [CR06, Gav09]. Strategies [PFHC08, AMC^{+}08a, APM^{+}09, All01, BCLM08, DDP^{+}06a, DDP^{+}06b, DDP^{+}06c, DSK^{+}08, HGSP01, KAN00, MvdK06, MRT03, PDH08, SST^{+}08, Vas03, Vel03]. strategy [BGM^{+}09, CAHAF07, DDP^{+}08, GEH^{+}07, MLP^{+}08, Nee08, Smi03, Sve15]. Stratified [UMV^{+}09, VKČ^{+}09]. straying [CB00]. stream [KA06, NNE^{+}05, PB04, SET^{+}09, McI07a]. stream-dwelling [NNE^{+}05]. streams [AKK^{+}03, CBGP06, GT04, MEA^{+}03, VP05, WN03]. Street [DOQ08]. strength [AOMK^{+}04, FGOF08, FK04, FKBF05, KFO^{+}04, MC05, Meh06, PMT06, RHF^{+}08, TH06]. strengthening [KOO04]. stress [CCA^{+}03]. striata [SMS02]. striatus [ED07]. strip [CA07]. stripe [NCW00a, OSB09]. striped [ACMMCEGM02, JWR05, MM06b, MRT03, OGPDOM08, TLH07, VVH04, ZSP03]. stripey [NCW00a]. Strong [BM09]. Strongylocentrotus [CH03, GC04, KCH05, PZH02, ZP05]. strontium [CS07b, KAN00, SR00]. structuration [MBG^{+}07]. Structure [BAEF06, CCCC04, PL00, ATM08, All07, Arc03, ARVC01, ASD09, ABE^{+}04, BF00a, BMBM05, BNS00, BHH08, BD04, Cad00b, CP06, CSM^{+}09, CAJ00a, CAJ00b, CCM^{+}09, CPCD05, Caz00, CHL^{+}06, CHW^{+}08, CBRN08, CF07, CPR^{+}08, DCM^{+}05, DG08, DDP^{+}06a, EBA^{+}03, FÖB08, GCN02, GSA^{+}06, GR07, GD03, KM08a, KTL06, KCKP02, KM08c, MCL04, Mar06b, MC06, MGJB09, NP08, OMR09, PSvDPM01, SCUADJ07, SHKB09, SWT09, ŠV04, SW02, SS00a, VN00, War06, WN03, WCC^{+}09, XJ05, YFL08, YLCS^{+}09, Zho07]. structured [CAHAF07, CHSA02, DBK07, DMBF08, FFL09b, FBS02, JMGC06, MN03, MG03a, MG03b, PWTP00, Pun03, WSPY05, WSPY07]. structures [IS07, JLHO01, NBF09]. structuring [JMGL04]. studied [GM00]. Studies [JCS01, JDMN00, AdBP02, BGH05, CRG^{+}02, DSK^{+}08, DKH08, HFHO07, HR09, KN01, MHMP01, MG00, PSTM06, Rei01, SLA00, SHL09, SE02, TS09]. Study [CMPM01, TMP^{+}06, ABC^{+}08, Abl06, BTC09, BPH^{+}06, CJ00, CXCD07, CR06, CHSA02, CPR07, EH02, FK03, FF01, FBV06, HF07, Her05b, HO05, HPK06, HO06, HPK07, HKF^{+}09, HW00, IWW03, KMK09, KAHR02, MLH^{+}05, MTA^{+}02, MWF06, MMLT05, MPG^{+}08, OSG08, PA02, PKT^{+}02, PNJS06, PTPS09, PBR05, RE07, RMA^{+}05, RR03b, SSS^{+}03, SCW04, SSML02, TSTK06, YCB^{+}01, YFL08, ZB00, vHMAO^{+}06]. Studying [UVKS05, BDB^{+}03]. style [KG06]. stylirostris [LMRQNM^{+}05]. Sub [LDA01, LD01, BH01b, DSO^{+}05, Dou08, KVS04, LNG04, PAP04, SJ05, WSS02]. sub-adult [PAP04]. sub-Arctic [DSO^{+}05]. Sub-areas [LDA01, LD01, KVS04]. sub-daily [Dou08]. sub-regions [LNG04]. sub-surface [BH01b]. sub-tropical [SJ05, WSS02]. subantarctic [Hor02]. subarea [ERG07]. subcutaneous [RW07]. Subdivision [NLJS01, OLN^{+}02]. Subject [Ano00t, Ano02-28]. sublethal [FMC^{+}09]. submergence [SE06a]. subpopulations [SW02]. Subsampling [HBW00, HBJ03, Hur07]. subsetting [SM04a]. Subsistence [CGT08, KVZ05]. substantial [BSG02]. substrates [CF07]. subtropical [THN^{+}04, Ano00e]. subtruncata [DG08]. success [NJJ06]. Successful [DDHS06, McI05c]. sufficient [HCM05]. suggest [MGR^{+}02]. suggested [Cad00a]. suggests [BSG02]. Suitability [SAA^{+}03, TCC^{+}09, VFCC06]. suitable [GPG^{+}03]. suite [CAM^{+}08]. sulcus [TLMN00]. Sullivan [Coo00a]. Sultanate [AMAOMM04, CMAOAA05, HMAOAS07, MCAO^{+}05]. sum [Sca03]. Sumaila [Sym08a]. summer [BVH^{+}08, LC06, Loh08, YSS^{+}09]. summer- [BVH^{+}08]. Sunken [Hol09]. sunlight [TH06]. superba [HD00]. Superior [NHSD08, YSS^{+}09]. supplementing [MC00a]. supply [FMGG08, SW09]. supply-driven [FMGG08, SW09]. support [BMC07, Gav09, PRdPG07, RGGB07]. supportive [LPG^{+}09]. supportive-breeding [LPG^{+}09]. suppositions [NJL^{+}00]. surf [MJS00]. surface [BH01b, HHTE05, KE05, OR07, PTE09, Tre01, USKK04]. surface-longline [HHTE05]. surplus [Mai09, MM06c, Pra02]. surplus-production [Pra02]. Survey [HP04, AB09a, CSR06, CRSM04, CCBG02, DA07, Din05, DK05, FP00, GMR01, GMR02, HPM04, JI07, KP07, KCKP02, KS02, LK02, LIW04, MvdK06, MWJ04, MUO02, MS02, RM05, RFWH09, RB04, RMA^{+}05, SHBC02, SW01, Som03, WRWK02, WSM02, WK08, Wil09, WFJB09, YSS^{+}09, vSS05, Mes03, WSP06]. surveying [SSH^{+}06]. surveys [ALN03, AE03, BMW07, BS04, CA07, EBA^{+}03, GDDJ09, GDFH07, GRC^{+}09, HdS07, HFBM08, JR07, Joh03, JPVKS^{+}01, LBKG08, PG06, PWP02, RSP^{+}00, SPS^{+}07, ST03, SOS02, Tre01, TKFJ09, WPOB09, YPDS05]. Survival [BLH04, KNJJS06, MTU^{+}04, MB00, PEB03, SSIE06, VAYD04, ARLB07, BVH^{+}08, CWL00, CS07a, CK02, DACS08, ECEG09, GCH^{+}07, GCMJ08, KNSSKF09, LC06, QDV05, QGBS04, RDH05, RR03b, SMB^{+}08, VL08]. survivors [YLYL09]. sustainability [Ano07y, GNC07, KM03, Lon06, MPG^{+}08, NHSD08, PP01, SC08, TCS05, TP06, ZG08, Sym07]. Sustainable [McI02f, BPCR^{+}05, Cla01, FCS^{+}05, HCM05, IK07, MM06c, PCS05b, SH03, ZR04]. sustainably [She09]. sustaining [McI02e, McI02f]. Svåsand [She02]. Svasand [McI05b]. Svedäng [FWG15, PCHM08]. SW [ASM^{+}09, BPRM00, SGS02, SGB00, SBGG^{+}05, BBM^{+}06, CC03b, LLV^{+}08, WPS^{+}07]. swab [WVW^{+}08]. Swedish [CRD06, OAKM07, UHN07, VU08]. swept [KM02, vS04]. swim [NGM^{+}09]. swimbladder [FGOF08, KG04]. Swimmer [McP02, BdL05, CS09, XK04]. Swimming [He03, AAH06, AKA06, AAKT07, AAT08, ABRB00, AC00, BSAM09, GNL^{+}09, HCH09, MC00b, ODR00, SKF08, YEA07]. switching [KM03]. Swordfish [MGJB09, BH01b, CPSBRD^{+}02, DMA07, HHTE05, HE08a, PMC^{+}07, Pra02, RYSN08, SWPY05, TMP^{+}06, TTP^{+}09, VL09, WSPY05, WSPY07]. swordtip [WLL08]. symbolism [Lon03b]. Symes [McI02d]. symmetricus [ACG^{+}01]. Symposium [Ano08l, Lon02a, Ore03]. synagris [MR08a]. synchrony [SN09]. synergy [She09]. synopsis [Nas03]. system [Ara09, BP01a, BPD^{+}04, GSS^{+}06, GB07, GLH07, HCS^{+}03, JL00, KAN00, Kor04, KMK09, LZS01, MH01, MDJR05, MD09, PKT^{+}02, RGSB08, STK^{+}05, TSS^{+}04, WCM04, vSM02c, vSM02a, vSM02b]. systematic [LRC^{+}08]. Systems [Ste03a, ET00, FM04, FFL09a, FFL09b, GLS^{+}09, HC07, VL09]. Szczecin [PLMDB06].

T [Ano00b, Bar00, Coo00a, McI02c, McI05b, McI05c, McI07b, Ore03, Sym07, Sym08a, BW08, LZZN09, McP02]. T-bar [BW08]. T. [JS02]. TAC [FMGG08]. tactical [CR06, Gav09]. tactics [TEYA02]. Tag [PRdPG07, SLA00, ARLB07, CWL00, DGS^{+}05, Ehr08, FH03b, GHM04, HZZG06, JLHO01, SMB^{+}08, XM00, ZLC09]. tag-attrition [GHM04]. tag-recapture [SMB^{+}08, ZLC09]. tag-related [DGS^{+}05]. tag/recapture [XM00]. tagged [CB00, LJP05, RMRB06, Ste03c]. tagging [BC04, CPSBRD^{+}02, CS07a, CCD01, CK02, DGS^{+}05, FH03b, GNL^{+}09, GR07, IV08, MNML00, RW07, Tal09, UL01, WH08, WMS04]. tags [BW08, BLB07, CMOM00, CK02, DYWHZ04, GHM04, GCLA07, GM00, Loh08, MCM^{+}08, MFC^{+}08, McP02, MMH05, NHT07, ZW09]. Tagus [BPCR^{+}05, GC03a, VFCC06]. Taihu [YLCS^{+}09]. tail [Bra04]. tailor [AWY02]. Taiwan [JLC04, JCLL08, LCNJ06, LLJC09, STCH09, SWPY05, TCY01, TY02, Tze04]. Taiwanese [SYS^{+}08, WMdS09]. take [FFM08, PK06, Row02]. Talang [GFvdV06]. Tana [SBD^{+}06, DSO^{+}05, SBD^{+}05, dGvZM^{+}06]. Tanganyika [vZRMvD02]. tangle [SPB^{+}03, VAYD04]. tanks [GG04]. Tanner [Ste03b]. Target [FKBF05, BTC09, BCMM09, BDBD09, DNN^{+}03, DA07, FGOF08, FK04, FHJL08, KAO04, KM03, KG06, KFO^{+}04, LD01, MTGH00, MC05, McI03a, MHK^{+}06, Meh06, OA04, PMT06, RHF^{+}08, RMN08, RCA^{+}07, SMSK03, VL09, WEK^{+}09, ZP05, vMBGF05]. target-strength [RHF^{+}08]. targeting [Mai09, PVP04, QPR08, WBCP08]. targets [Mil03a]. Tasman [SRH^{+}02]. Tasmania [TLH07, ZFJ03]. Tasmanian [Bra04, HSGF06, IHF08]. tax [GGVLSR03]. taxa [CP08b]. Taylor [How06, Mil02a]. teabags [AAT08]. Technical [FK03, ERG09, FKW06, FP02, HP02, PPS05, TPC05]. technique [ET00, GCJ07, HP00, IV08, KMELC05, LR08, MF00, MSB07, OIK^{+}06, PP01, Smi03, TMK^{+}06, WVW^{+}08]. techniques [AD03, Aló09, BR00, GKV06, MC06, SRH^{+}02]. technology [CPR07]. Tehuelche [NS04, NGF06]. tehuelchus [NS04, NGF06]. telemetry [BDB^{+}03]. Teleost [LBABH09]. Teleostei [BK02, EKU02, vHMAO^{+}06]. temperate [AMS08, FF01, WSS02]. Temperature [SOH06, TOK^{+}05, ACE^{+}08, Chi00, GCH^{+}07, JR02, KLU01, LNS08, LMASHV^{+}03, MM05, MAC^{+}06, NJJ06, ÖW02, PSS^{+}07, RB06, SN09, SLJ05, SSC^{+}03, Wie04, YEA07, Hat00]. temperature-dependent [SSC^{+}03]. temperatures [Cor01]. template [TBRL00]. Temporal [BSMF^{+}09, BB01, BPD^{+}04, KSS05, LPG^{+}09, PK03b, ASR05, AMUE01, CC01c, CBRN08, DMA07, DSO^{+}05, HSGF06, LAL07, LRBH07, LMR^{+}07, MBT08, Man00, OAKM07, OQE03, PMSD06, SWT09, SAG05, TGH^{+}08, TVB^{+}09, TOBSJH06, TDK^{+}05, WMB08]. tench [Yil02]. Tenerife [ABFV02]. Tenualosa [RC06, SMRH04]. tenuis [MMH05]. Terje [She02]. term [AWY02, Bal09, BC04, BD04, CRG^{+}02, CGEHC07, DACS08, HTF^{+}09, JPVKS^{+}01, KVZ05, MF07, MD02, NNE^{+}05, NJE00, PHM07, Ste03c, TNL07, TSJ04, WY06, vZRMvD02]. test [PBW^{+}09]. tested [MH01]. Testing [KF02, Lab05, RSST07, KFM08, MTHL06, ML04a, QDH09, SdPFAL09]. tethered [GCLA07]. Tetrapturus [ACMMCEGM02, AB09b, OGPDOM08]. Texas [CHN08]. Thailand [LD06, MMLT05]. Thalassinidea [RW03]. Their [MKGH05, MFC^{+}08, CRG^{+}02, CC01c, JCJ^{+}09, KHMS09, LRC^{+}08, LLV^{+}08, MMB07b, PPdC01a, Pri09, RL03, Sva05, WPS^{+}07]. Thenus [CCD01]. Theoretical [HO05, HPK06, HO06, Pal08, Her05b]. theory [SW02, Tri05, VD04b]. Theragra [APC07, BQ04, Dou08, ODR00, RO00, vSSK07, vSS09]. there [Ste06]. thermal [CMOM00, Orl03, PM09, WPS^{+}07]. thickness [HO06, HWM02, SLB07]. thornyheads [LWS04, LIW04]. those [BPH^{+}06]. threadfin [CGLZ04]. three [All01, BMBM05, CS07a, CdML^{+}07, DBB00, DSK^{+}08, EF01, FHM09, GLS^{+}09, GBFG06, JK07, KS02, Løk03, PBVG07, PLPN09, PMC^{+}07, SHKB09, SW00, SST^{+}06, TASF09, UM02, You03, ZK03]. three-dimensional [DBB00, PBVG07]. thresher [LCNJ06]. threshold [PDH08]. throat [WMD07]. throughout [MGR^{+}02, SMRH04]. Thryssa [JS02]. Thunnus [ADP^{+}08, AdBP02, BDB^{+}03, CHL^{+}06, CHW^{+}08, DHM08, FDG^{+}07, GCC^{+}08, HCS^{+}03, LDN04, MM07, MB08, NC04, RM07, STI08, STCH09, SYS^{+}08, SEPY02, WMdSB09, WMdS09, WH08, WCC^{+}09]. thynnus [ADP^{+}08, MM07]. Thysanoteuthis [BM05, MOK^{+}05, MOGG06]. tidal [Kru04, Ste03b]. tier [GF06]. tiger [DPD^{+}03, NAD^{+}07, YLD^{+}05]. tilapia [BBV07, dGvZM^{+}06]. tilt [TMK^{+}06]. Time [GKV06, TC09, AB08a, Cad03, CCG^{+}02, GRFE06, JS02, MAASMB05, Mau03, MK09, PBMB03, RRAS03, TdlMH^{+}03, XM00, YEA09, ZFJ03]. time-dependent [XM00]. times [AB08b, AD00]. Timing [Ste03b, DGS^{+}05, HMM^{+}09, WT00]. Tinca [Yil02]. Title [Ano03f, Ano03g, Ano03h, Ano04e, Ano04f, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano06n]. Todarodes [CLH^{+}08]. Today [Gra02a]. together [PM09]. Toliara [MWF06]. Tomorrow [Gra02a]. Tonga [Kro04]. Tony [Mil03b]. tool [FCC^{+}09, GMP05, OB02, OFOM08, SMSZ08, TLD06]. toolbox [Mau03]. tools [GSS^{+}06, MTA^{+}02, McI02e, OS05, RSMZ06]. tooth [Gas03]. toothfish [Hor02, LB05, PAB05, TdlMH^{+}03, WR02, YCB^{+}01]. top [SCP01, vM03]. Topographic [HC09]. Tore [Hol01]. Toronto [Ano01d]. Torres [YPDS05]. Tortugas [ARCLBPdP07]. Total [BGW06, GHDM08, KI04, KI05, BB01, RMN08, ST00]. totoaba [LC07, LC07]. tourism [IK07]. Tourist [Qui07]. tournament [LSW06, LWM07, NGM^{+}09]. tournament-caught [NGM^{+}09]. tournaments [MHWL01]. tow [CSR06, JRG08, MJB02, SOS02, WRWK02, WSP06]. towed [JR02, KW03a, KW05, LWS04, LIW04, MC06, OSB09, PTE09, WG03]. towing [DWW^{+}02, GLS^{+}09, JSO08, OKWM05]. tows [OR07]. Trace [MD01]. tracers [CCH^{+}00]. Trachurus [AGS^{+}08, AMC^{+}08b, ACG^{+}01, CSM^{+}09, CBRN08, CMAS08, CMH01, CPR^{+}08, FAL^{+}01, KWK^{+}09, KR00, KM08a, MCM^{+}08, MPA08, SMSZ08]. Trachynotus [Aló09]. tracked [PM09]. Tracking [MS05b, BBK08, GCLA07, HJLL07, PAR^{+}03, Tri06, UHN07]. tracks [CRG^{+}02, EGJ00]. trade [Ano07y]. Traditional [FFL09a, GG04, MWF06, Moo03]. traffic [Cad02, CWS^{+}05]. trailer [LSW06]. trailer-boat [LSW06]. trait [PTPS09]. traits [AMS08, BTC08a, FF01, KNSSKF09]. Trammel [BCC09b, FSB^{+}02, SMS^{+}06, BTC09, EGB^{+}06, FK03, GGA08, GSH^{+}07, MP03, TEG03]. trammelnet [RMN08]. trammelnets [GQR03]. transducer [BMW07, PTE09]. transect [CA07, KP07]. transect-based [KP07]. transferable [KPL^{+}07, KPL^{+}09]. transforming [PC03]. transition [AM00, KCK03, LDGA09, Wie04]. Transitive [BR01]. translocation [WZDY06, YMT^{+}06]. transmitting [Loh08]. transparent [GLK00]. transplanted [WR05]. transported [WCM04]. transversal [PRdPG08]. trap [ADP^{+}08, ABCB^{+}07, BSWL06a, BSWL06b, CG04, FHG03, Gro00, GMS06, HSGF06, HFL08, KSS05, SSIE06, ST03, SF03, SSK^{+}06, VMSR08]. trap-caught [HSGF06]. trap-net [HFL08, KSS05, SSIE06, SSK^{+}06]. trapping [BEBA03]. traps [AA03, AMAOMM04, GQR03, GG04, HMM^{+}01, PMAK07, Sal02, SMS02, Xia06]. trash [NBA^{+}09]. Trasimeno [LCD^{+}02, LDE^{+}02]. traveled [Pal08]. Trawl [BEA^{+}08, RBD02, SGB00, WSP06, AR05, APM^{+}09, AE03, ASIW05, BSS06, BSC^{+}08, BVW04, BHM^{+}06, BKD^{+}02, BMK^{+}04, BNL^{+}07, CF04, CFFP07, CAM^{+}03a, CTK^{+}07, CAM^{+}08, CRSM04, CTC^{+}06, CHDGM06, DPS^{+}01, DBAÖ06, DA07, EGJ00, FKW06, FÖB08, FP00, FCG02, FCL^{+}05a, FCML05, FCL^{+}05b, FCM07, GC03a, Gri03, Gro00, HJ03, HJ07, He07, HBW00, HBJ03, HBKJ07, HGW^{+}08, IJ06, JGB^{+}06, JR07, JSR04, Joh03, JSO08, JPVKS^{+}01, KTL06, KW05, KCKP02, KMK09, KE05, KME^{+}02, KHMG06, KI04, KI05, KOO04, LIW04, LZS01, MMB^{+}07a, MGTB08, MVV^{+}01, MVAMB07, MH02, MHF02, MSB^{+}08, Meh06, MTU^{+}04, MAG^{+}09, MS02, MPG^{+}08, OCT^{+}03, OL07, OMR09, ODR00, PFHC08, PBDP00, Pol00, PAF^{+}08, QDH09, RH04b, RDH05, RDH06, RMA^{+}05, ROS04, RB06, SLB07, SLPF08, SRA^{+}02, SSB03, Sch09, SW01, SOS02]. trawl [Som03, SMSK03, SJ05, SLJ05, TCPC06, Tia09, TÖT04, TGH^{+}08, USKK04, WW06, WLS07, WSM02, WK08, WR02, YEA07, YEA09, You03, ZG08, vS04, vSS05, vSSK07, Dah00]. trawled [MF07]. trawler [ZCdG06]. trawlers [AKA^{+}02, ECEG09, HW00, PLSI09, WRGS09, tHDC06]. Trawling [FCAD^{+}08, AB09a, BPH^{+}06, BMSJ05, DDP^{+}08, FP02, GFSM07, GBFG06, JRML00, PBW^{+}09, RBH^{+}08, SDM04, Sva05, VU08, WMTJ07, WRK06b, vMBGF05, vSS09]. trawls [Bet04, BPW^{+}03, DWW^{+}02, FRF08, GK01, GKF03, GBM09, GCL03, HC00, Hal02, HM00, JRG08, JK07, MBM04, MvDM02, NHT07, OKF06, Pri09, RJ05, RDH06, Rye08, SE06a, TMK^{+}06, WDC02, Wes02, vM03]. treasure [NBA^{+}09]. tree [WK01]. trend [BBJG04]. Trends [CPW^{+}06, FH03a, PAP06, RKEJR06, Ste05, WRK06b, AS04, CNH09, GHCPPLC^{+}05, HSGF06, KVZ05, PB03, PSS^{+}07, vZRMvD02]. trevally [WHML04]. trial [BNL^{+}07, KLS^{+}07, SRA^{+}02]. trials [AAH06, BSWL06b, BR00, FBL^{+}07, PDV05b, SdPFAL09]. triggers [Sal02]. trip [CNH09, PL04]. Trisopterus [MGR^{+}02]. Trochus [PAP04]. Trophic [ACMMCEGM02, SB01a, BCC09a, MAND05, PEAAdLJB06]. trophy [RKEJR06]. Tropical [HGW^{+}08, RLGLO09, BHM^{+}06, BD04, CdML^{+}07, CHDGM06, DOQ08, EAT05, FF01, FBV06, GHM04, HBW00, HWW^{+}09, LHMO09, MVAMB07, MWB^{+}08, MM06a, NS08, NCW00b, OLN^{+}02, PGK03, Rue07, SMRH04, SMJS01, SJ05, TP00, WSS02]. tropicalization [Ano02l]. trough [DG08, LDA01, MD01]. Trout [Mil08, ARLB07, AKK^{+}03, BLKR09, BRRL09, BPD^{+}04, JRML00, KNJJS06, Lan04, MM05, MEA^{+}03, NHSD08, PCS05b, RMRB06, SLE09, WN03, Ell00]. true [KHM^{+}09]. trumpeter [TLH07]. truncated [OF03]. truncation [Mai09]. truth [MTGH00]. trutta [ARLB07, BLKR09, BPD^{+}04, JRML00, Lan04, RMRB06, SLE09]. truttaceus [CWL00]. Trypaea [RW03]. TS [KFO^{+}04]. tshawytscha [CB00, Kop06, MK09, QDV05]. Tsushima [KWK^{+}09]. Tucannon [CG04]. Tumidotheres [NS04]. Tuna [But08, KPBD08, MB08, ADP^{+}08, AD03, AdBP02, AC02, AS04, BDB^{+}03, BCMM09, CHL^{+}06, CHW^{+}08, DHM08, FDG^{+}07, GHM04, GCC^{+}08, HCS^{+}03, HT05, Kar04, Lab05, LDN04, MM07, MROGVR09, MD02, NP08, NC04, ORVMP06, OA04, PMC^{+}07, RM07, STI08, STCH09, SYS^{+}08, TSOG01, WMdSB09, WMdS09, WH08, YKM06]. tunas [FF01]. Tunisia [LBABH09]. turbot [He05, NJJ06, SSML02]. turbulent [RDM^{+}09]. Turkey [BK02, Yil02, AAA^{+}06, DG08, Güç08, TEYA02]. Turkish [DBAÖ06, Kar04, TÖT04]. turning [HPK07]. turtle [BM08, BHM^{+}06, PL06, SE06a, YKO09]. turtles [BNL^{+}08, HLL^{+}08]. Tweedie [Sho08]. twin [DBAÖ06]. twine [GBM09, HO06, HWM02, SLB07]. Two [GHDS08, How06, Lev05, PSTM06, RHF^{+}08, SWA08, AKA06, AAKT07, Arc03, BSWL06a, BSWL06b, CC03c, CC01c, CdML^{+}07, DYWHZ04, DV00, ESC08, Eer04, ED07, FAW04, GHM04, GCS^{+}06, GCD^{+}09, GBC^{+}03a, GBC^{+}03b, GF06, GK03, Jak01, JS02, JIDM02, KA06, LP00, LCTB^{+}03, Mag01, MM04, NV05, OSG08, PH09, QDV05, RYSN08, SC07, SMM07, TBRL00, WN03]. two-chambered [BSWL06a, BSWL06b]. two-dimensional [TBRL00]. two-pass [KA06]. two-phase [RYSN08]. Two-species [SWA08]. two-tier [GF06]. Two-volume [How06]. type [AAP^{+}09, BB04, BCSP03, BH01b, JS05, LSW06, OGK^{+}08, WPOB09]. types [Aló09, AKA06, BdL05, DYWHZ04, FRF08, GHM04, ML04a, Xia04b]. typology [PBVG07]. typus [CARA^{+}07, GR07, HJLL07, JH07, Mar07b, NE07, NS07, Qui07, RMVJGMMV07, Row07, Ste07, Tay07]. Tyrrhenian [BLV06, CMC^{+}06, SSB03].

U [Ano00e, Sym08a]. U.S. [CRSM04, FKW06, HLL^{+}08, HP09b, PDH08, RFWH09, VSS07]. UK [Bar00, Van01, BH01a, BGH05, ECEG09, LF02, MLJT09, PPS05, RD03, WRGS09, WFJB09, YPD^{+}04]. ultrasonic [JWR05]. umbla [TEYA02]. umbratus [BP06]. Umbria [LCD^{+}02, LDE^{+}02]. Umbrina [HGF06]. Unaccounted [GFS03, RPP04, HMM00a]. Uncertainties [Zhe03, LNG04]. uncertainty [CP08a, CHL01, DDP^{+}06a, GC04, Hol01, Kat06, MK01, Nas03, Sch01, UM02, Zho07]. undatum [MB04, SALnM07]. undersized [GFSM07, MCJB02, MJB02, PEB03]. Understanding [ABF08, DB08]. underwater [BFB^{+}03, CC09, CA07, HCS^{+}03, He03, Løk03, OSB09]. undisturbed [EAT05]. undulata [CE06]. undulate [CE06]. undulosus [CMGB01]. UNESCO [McI07a, Sym08c]. uneven [AR05]. unexploited [DCM^{+}05, PK03a]. Union [Sym09]. unique [RCSFO04]. Unit [LLJ^{+}08, BQ04, BZRO05, CNH09, CP08b, Chi00, FH03a, GFM02, LP06, ML04a, MDS03, PLPN09, vOPvDM02]. United [GHDM08, McI02e, RRJ08, RRJ09, CC03a, SWT09, WD02]. units [AGS^{+}08, APP^{+}07, GFM02, HZZG06, KA06]. univariate [CGEHC07, GESY^{+}07]. University [Ano01d, Coo00a, Coo01, Gra02a, How03b, McI01, McI02c, McI03e, McI06b, Mil02a, Ore03, Ste03a]. unregulated [AP05]. unreported [AP05, PAB05]. unscented [LNS08]. unsustainability [NGF07]. until [EMKG07]. upon [AOMK^{+}04, KHMS09, RB06]. upper [PA05, GMA02, NJL^{+}00, NWML00]. upstream [MNML00]. uptake [HHC06]. upward [PTE09]. upward-facing [PTE09]. upwelling [GDAM07, GEH^{+}07, KMF08]. upwelling-related [KMF08]. urban [ABF08, HCM05]. urchin [ASM^{+}09, BW02, CH03, GC04, GTB04, KCH05, PZH02, ZP05]. urchins [GCB^{+}06]. Urophycis [MMH05]. ursinus [MKB01]. Uruguay [GD03, MM06b]. Uruguayan [MAND05, VMQ08]. US$88 [McI09a]. USA [Ano01c, Ano01e, Gra02a, Lai00, McI02a, McI09a, McI09b, McI09d, Tal09, BSC^{+}08, CH03, KCH05, RB04, The07, WCM04]. USA/Canada [Tal09]. usage [MSTC04]. Use [BPRM00, BLB07, CdML^{+}07, GLK00, MMH05, OF03, PAB05, RSM05, Sca03, TÁTWD09, TBRL00, WHML04, XM00, Xia04a, Xia04b, YAHAB01, ZB00, ZP05, ZLC09, AHS^{+}06, BMBM05, BMW07, CF04, CAJ00a, CAJ00b, CPSBRD^{+}02, Dou04, Ein09, FCML05, GM00, GDFH07, HD00, LM09, LZZN09, LYKH06, Lor06, Lot07b, MFC^{+}08, MR08a, MSB07, PBLP08, PPGD08, RW07, SLB04, SSS^{+}03, SBFR00, SB08, SSML02, TH03, VD06, WVW^{+}08]. used [ABRB00, AC00, DACS08, GHM04, HMM00b, MC00b, Pol00, RDH06, RCD07, RCA^{+}07, Rue07, TRL06, vSS05]. useful [SR00]. Using [AIC08, BHB^{+}04, BBJG04, BKD^{+}02, CWS^{+}05, CC09, CSR06, CAHAF07, CPR07, DGP02, FCL^{+}05a, FCL^{+}05b, KKJ00, LMR^{+}05, LMR^{+}07, MTRM05, MM06c, MG03a, MG03b, NHT07, SDS09, SE06b, VP05, VBSK00, AA03, AKA^{+}01, AD03, AdBP02, AAH06, AAT08, AGB07, ABS^{+}03, APP^{+}07, BSS06, BPW01, BMC07, BP06, BBM^{+}06, BFB^{+}03, BS04, But03, Cam04, CWL00, CNH09, CMPM01, CL09, CXQ^{+}09, CP08b, DVAL^{+}09, DDP^{+}06b, Dou04, Dun06, FL05, FÖB08, FAN03, FFM08, FH03b, GM06, GCLA07, GMR01, GMR02, Gra03, GF06, GR07, GMP05, HC07, HMM00a, HBJ03, HMH^{+}04, HSW09, IWW03, IS07, Joh03, JRG08, KM08b, KCRT03, KMK09, KME^{+}02, LD08, Lab05, LWS04, LIW04, LRG^{+}00, LZS01, LDN04, Loh08, MGS02, MK02, ML04a, Mes03, MC06, MD09, NCW00b, OIK^{+}06, OSG08, Pal08, PD08]. using [PBVG07, PH04a, PLPN09, PZH02, PWHM06, PBR05, PHGT06, RRAS03, RYSN08, San00, SHKB09, SW00, SLJ07, SD01, SHBÅ07, SMB^{+}08, STH03, TP06, VAYD04, VGBF02, WM09, WL06b, WRWK02, Wil09, WFJB09, WWG03, YCB^{+}01, YST^{+}02, vT04]. Ustica [GCB^{+}06]. utilisation [TPM03]. utility [GSS^{+}06, PGK03]. utilization [VSK03].

V [LDA01, LD01, McI09b, Sym08a, FR03]. Val [BC02]. Valdez [RR03b]. Valdimarsson [McI03b]. Valenciennes [AAB05, LLM05, SR07]. validate [KKJ00]. validated [Dou08]. validating [MC00a]. Validation [ACE^{+}08, CS07b, MMP^{+}02, MÁ08, NMF^{+}08, AHAAD01, MN01, MMN03, MBDM05, RMRB06, RRJ09, SKKP05, SG01b]. Validity [TSS^{+}04]. Value [LBKG08, DSK^{+}08, JTdH06, MAG00]. Value-added [LBKG08]. values [YAHAB01]. Vanuatu [PAP04]. Variability [BP01a, GM06, MMLR08, TLMN00, AR01, BH01a, Bal09, BB01, COLB07, CG04, GMV00, GMV01, Hat00, HTF^{+}09, LK02, Laz08, MBO^{+}07, MMB^{+}07a, MGR^{+}02, MLC03, MBM04, PHM07, PCS05b, RPFR06, SCM^{+}08, TOBSJH06, Vel03, VAF04, You03, vZRMvD02]. variable [ALN03, AR01, BZRO05, HPD09, Rod01]. variables [ABS^{+}03, BCSA08, CG04, VFCC06, VMQ08, ZPR01]. variance [RYSN08]. variances [LB08]. Variation [KAN00, MBG^{+}07, RL03, SW00, AE03, All07, BCC09a, BPW01, BSMF^{+}09, BSG02, BPD^{+}04, CC01c, CMAS08, EPA00, FOGKT08, FR08, GCN02, GRC^{+}09, HO05, HGF06, IJD04, JI07, JIAD03, LRBH07, LMASHV^{+}03, MCAO^{+}05, MR05, PTR^{+}05, PK03b, PHGT06, SMRH04, SALnM07, SMJS01, TS06, TGH^{+}08, TSOG01, TOÖD06, TCY01, Tze04, VKČ^{+}09, WMB08, ZFJ03]. Variations [CC03c, RMVJGMMV07, STI08, TAS03, XJ05, ASR05, AC02, GB03, Man00, MP03, RCGC04, Tia09]. various [WSC08]. varying [Cam04, CSR06, LWW02, OGK^{+}08, RSST07]. vector [BMC07]. vegetated [CPR07]. vehicle [BFB^{+}03]. velocities [DBB00]. vendace [GDDJ09, Meh06, SG06, Sch09]. Veneridae [PPL04, TSM03]. Venezuela [Gue05]. Venezuelan [OA04]. venting [JS05]. vents [JPYW08, SMS02]. venus [VVH04, TSM03]. Veracruz [AAFLP^{+}07]. Verany [CUT06]. Verde [SE06b]. Verification [WWG03, KLS^{+}07]. Verlag [McI05a]. vermilion [All07]. verrucosa [TSM03]. versatile [Bro05b]. versicolor [Fri07]. versus [BDB^{+}03, Dic04, FF01, GCA^{+}08, GHDS08, HGSP01, KG06, PCS05a, PLD05, SSC^{+}03, The07, TPC05]. vertebrae [AdBP02, GCC^{+}08, KK09]. Vertical [BSR05, ÅCST01, BDB^{+}03, CMH01, JR02, KG04, PBVG07, RG07, UMV^{+}09, VKČ^{+}09]. Vessel [PBMB03, APM^{+}09, Ara09, CRSM04, EGJ00, HPM04, JHGS04, MD09, OMW^{+}03, RFWH09, SSH^{+}06, VOH02, vSM02c, vSM02a, vSM02b, vSS09]. vessels [BLH04, CCBG02, HBJ03, MvdK06, Pal08, VL08, WKM02, WPOB09]. VI [Nas03]. Vic [Kes00]. Victoria [PG06, PHGT06, TÁTWD09]. Video [RB04, STH03, CC09, DSO^{+}05, Dun06, HFBM08, HFS02, HCS^{+}03, He03, LWS04, LIW04]. Vietnam [MHK^{+}06]. Vietnamese [LFA08]. view [Cor02]. vii [Ell00, ERG07, LDA01, LD01]. VIIIbc [PVP04]. VIIIc [CP00]. village [NWML00]. villosus [BG01]. Vincent [XM00, Xia04a]. Vinciguerria [TP00]. virgin [PK03b]. Virginia [CND02]. viridis [ABAA^{+}07]. virtual [CL09]. Visible [WMS04, BLB07, ZW09]. vision [SD01, WSS06]. Vistula [PLD05]. visual [CA07, PPGD08]. visualization [JL00]. vitreus [PCS05b]. vitta [NCW00a, WMTJ07]. VNTR [MGR^{+}02]. Voituriez [McI07a]. Vol [Ano01e, Lon02a, Mil03b, Sym08c, McI05d]. volcanic [FBV06]. Volta [ODdGV02]. Volume [Ano01g, Ano01h, Ano02e, Ano02f, Ano02g, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03-33, Ano03-34, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano05c, Ano05-30, Ano05-31, Ano05-32, Ano05-33, Ano05-34, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano06g, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano08a, Ano08b, Ano08c, Ano08d, Ano09e, Ano09f, Ano09g, Ano09d, Ano02c, Ano02d, Ano03k, FGOF08, How06, JWR05, KG04, TS06]. Volumes [Ano02b, Ano02-28, Ano00a, Ano00t]. voluntary [GJ09]. Volunteer [PPH01]. vomerinus [MGS02]. vs [SET^{+}09]. vulgaris [GB06, RB00, SLA00, Arc03, FRGF07, GBC^{+}03a, GBC^{+}03b, HGHLCH02, OGSG07, ŠV04, VYMGM^{+}09]. Vulnerability [FBV06, dGvZM^{+}06, CP08b, MHWL01].

W [ADP^{+}08, Bar00, HAO00, How06, KGB09, LS00, McI09c]. Wadden [Lot07b]. wagging [Bra04]. Wakefield [Ore03]. Walb [QP00]. Walbaum [GRB01, RCG02, WBHG01]. Wales [AHS^{+}06, BW02, Dem04, GKH03, OSG08, PKP04, SF03]. walleye [APC07, BQ04, Dou08, IFH^{+}09, ODR00, PCS05b, RO00, SC07, TMK^{+}06, WG05, vSSK07, vSS09]. walleyes [QGBS04]. Wallingford [McI03b]. Walsh [KW03b]. war [GHDS08]. Warm [KWK^{+}09, EMN07]. warming [EH02]. warp [SRA^{+}02]. warps [PAF^{+}08]. Was [MBK02]. Washington [Coo00b, McI02f, Mil02a, GJSW04, KE05]. waste [APM^{+}09]. wastes [BM03b]. Water [HHC06, Ore03, ABAA^{+}07, All01, BMW07, CLC08, Cor01, CTC^{+}06, EF02, FFM01, FCM07, GCH^{+}07, Gor01, HMO01, HG01, HP04, HIRM06, HP00, Jak01, JR02, JNDH00, KAN00, Lan04, LD01, MDS03, MN01, MD01, PPdA05, PCB01, PGMM01, SB01a, SRA^{+}02, SGB00, SW01, SYS^{+}08, SLJ05, USKK04, WKCC06, Wes02, YEA07, dSBAC06]. waters [ABCB^{+}07, AHAAD01, AHABAA^{+}02, APS07, AD00, ARVC01, AM02, BC02, BNSA05, BNS07, BPW01, BHJ02, BRO05a, CMJ05, Caz00, CMAOAA05, CMC09, DPS^{+}01, DK05, ECEG09, FGJD08, FZHL09, FCG02, FCM07, Fra07, GJSW04, GKV06, GPG^{+}03, GRB01, HSGF06, Hor02, HGF06, IJD04, JCS01, Jak01, JSR04, JLC04, JCLL08, JHK08, Jun08, KWK^{+}09, Kar04, Kir01, KH05, LPDA01, Lap02, LB05, LCNJ06, LLJC09, Mag01, Mar06a, MOOH00, MK06, MCAO^{+}05, MM06b, MCB^{+}07, Nar06, NAD^{+}07, NHSD08, OMF^{+}03, OGSG07, PMAK07, PdAO02, PB03, PJR^{+}05, RC06, RD03, RV03, SN01, SME02, SMS^{+}06, SWPY05, TVY01, TOK^{+}05, TCY01, WY06, WFJB09, WGGN06, YPD^{+}04, ZPR01, McI02d]. wavelet [WG05]. way [MGJB09]. weakfish [Caz00, MM06b]. web [AAFLP^{+}07, BP08, PEAAdLJB06, PEAAdLJB07, VSK05]. weekly [DPD^{+}03]. Weight [MMM^{+}03, RMMP06, SGVM02, UMK01, AC02, ACG^{+}01, Ber07, CAJ00a, CAJ00b, EAT05, GM07, LH04, LK00, LMR^{+}05, LMR^{+}07, MPM^{+}06, MAL^{+}01, MGNB05, Ogl09, ØT02, PGK03, PCAS04]. Weight-based [UMK01]. weights [BR05]. Wendt [McI07b]. West [ASM^{+}09, PDH08, UHN07, BH01a, BBJG04, BGB06, BPD^{+}04, CCT^{+}08, CCB02, CLGR^{+}04, CRSM04, FRGF07, FAL^{+}01, HP09b, LCFP^{+}08, MOOH00, OJ06, RWT01, RFWH09, SR06, SHDO04, SKP03, WDC02, vHAN^{+}09, BDBD09, FP00, HKG^{+}08, HTF^{+}09, JS05, NBA^{+}09, Wie04, WSP06]. Westburry [Bar00]. Western [GCB^{+}06, How06, McI05d, Tay07, dSHG08, AM00, AB09b, BM08, BSG02, CWL00, CCMS^{+}03, CAM^{+}03a, COHP02, CHL^{+}06, ESC08, FMCS05, GR00, HMRG08, IMSS06, IOS07, LDN04, MSTC04, MOGG06, MNM00, MMN00a, MFAM05, MMM^{+}03, MKB01, New02, NLJS01, PG06, PMC^{+}07, RCGMQV05, RCG02, SSB03, SDH00, SHBC02, SYS^{+}08, WMTJ07, XM00, Xia04a, YSS^{+}09, ASR05, ASA04, AWY02, BKD^{+}02, CCMS^{+}03, CUT06, ERG07, HFBM08, HC09, MSTC04, NS07, RDH05, WMTJ07]. wetlands [NJL^{+}00, NWML00]. Whale [Ano07w, RDEACT07, Ste07, WC07, CARA^{+}07, GCLA07, GR07, HJLL07, IK07, JH07, Mar07b, NE07, NS07, Qui07, RMVJGMMV07, Row07, RG07, RE07, Tay07]. whales [Hol01, PSS^{+}07, Sch01]. whaling [Ano02-29]. Whatever [SP09]. whelk [MB04, Nar06, PMAK07, SALnM07]. which [HBW00]. whiskery [SDH00]. White [LDK^{+}07, APS07, AB09b, DOQ08, DKKS07, FMB^{+}05, LQD^{+}08, MMH05, SR07, iTM04, iTMI06, VH06, WWG03, YTK06, YMT^{+}06, LAL07]. white-spotted [iTM04, iTMI06, YMT^{+}06]. whitefin [YFT^{+}02]. whitefish [BLC08, GJ09, LWW02, PHM07]. whitemouth [NV05]. whiting [ÅCST01, ASD09, BH01a, CMPM01, CTK^{+}07, CCBG02, FAL^{+}01, GM06, GM07, GKF03, IPH05, KMB01, Pál05, PKK^{+}06, ZPR01, vM03]. Whitley [SDH00]. whitmaei [SK08a]. whole [BG04, JJWO09, Pol00]. whole-lake [JJWO09]. Wider [CMM^{+}07]. widow [RFWH09]. width [WK08]. widths [Ker00]. wild [BW08, CK02, DSO^{+}05, DDHS06, FCS^{+}05, Jok02, JJKN^{+}03, MM05, SLE09, iTM04]. Wildlife [McI09c]. Will [PBLP08]. William [How03a, KTC00]. Williams [Mil06]. willingness [CGLZ04]. Willis [McI09c]. Windermere [EF01]. window [EH04a]. windows [CF04, GSL08, MHF02, MS05a, GLH07]. Winnipeg [WG05]. winter [CCT^{+}08, He03, Laz08, PPH01, XJ05]. wintering [HB01, VOH02]. wire [BLB07, CB00, CMOM00, CK02]. Wisconsin [NHSD08]. within [BMF^{+}10, BSWL06a, BSWL06b, CRG^{+}02, DJGC09, GFM02, KAN00, LK00, LRC^{+}08, MK09, NE07, PKK^{+}06, SF05, SALnM07, SBGG^{+}05, TS06]. within-trap [BSWL06a, BSWL06b]. without [KI04, KI05, PC03]. Wood [Ano00b]. Woodby [Ore03]. words [GHDS08]. Work [Sym09, RCD07, She03]. workshop [Bla06, Ano06-30]. World [Ste05, Ano07y, McI07b, Kes00]. worrying [Koe08]. Wounding [DO06]. wrasse [FMB^{+}05, GR00]. wreckfish [PH04b]. wrecks [RD03]. written [vT04]. www.blackwellpublishing.com [McI07b, Mil08].

X [Bar00, McI05c, Sut02, Ano00e, McI06b]. xanthurus [RPC^{+}06, ZR08]. xi [Win02]. Xiphias [BH01b, CPSBRD^{+}02, DMA07, HHTE05, MGJB09, Pra02, RYSN08, SWPY05, WSPY05, WSPY07]. xvii [Sut02].

Yacyretá [AAB05]. Yarrow [HGW^{+}08]. year [BNL^{+}07, FH03b, IHF08, KNJJS06, Laz08, PBDP00, PTE09, RR03b, SMD04, SHKB09]. year-long [BNL^{+}07]. Yearbook [McI03e]. years [FAL^{+}01, GRFE06]. yellow [IWBJ08, WIJB08, COHP02, HLH08, WLY06, XJ05]. yellowfin [AD03, DHM08, LDN04, NC04, PMC^{+}07]. yessoensis [Uki06]. yield [AHABAA^{+}02, AWY02, CHL01, CHSA02, DJGC09, EKU02, FÖB08, GHS04, GMA02, GBC^{+}03a, GBC^{+}03b, HCM05, KM03, Kir01, KVZ05, LMASMASZ02, MAG00, NMMTCV^{+}06, NCW00b, PJR^{+}05, SJ05, SEPY02, SWPY05]. yield-curve [CHSA02]. yield-per-recruit [AHABAA^{+}02, GBC^{+}03a, GBC^{+}03b, Kir01]. yields [MM06c]. Yong [KW03b]. Yong-Hae [KW03b]. York [Ano00b, How03a, Ste03a, RVS02]. young [Laz08, PTE09, SHKB09]. young-of-the-year [Laz08, PTE09, SHKB09]. you’re [LBKG08]. YOY [CPR07]. Yucatan [BRL^{+}05, HS03, RLSJIA07]. Yukon [Bro05b].

Z [McI06a]. Zambezi [THN^{+}04]. Zealand [BM09, CC09, CAFT00, Cla01, GT04, HMM00b, HMM00a, Hor02, KJ09, PH09]. zebra [PLDS03]. Zeller [SH04]. zero [MLCGRV07, Sho08, YAHAB01]. zero-catch [Sho08]. zeroinflated [MLCGRV07]. Zhang [Coo00a]. Zimbabwe [Man00]. zinc [BM03c, HMB^{+}00]. Zingst [GWR07]. Zone [VMQ08, PCMC08, AIC08]. zones [ASM^{+}09, JHSG04, PK06, WF07]. zooplankton [KTC00]. Zygochlamys [BBM^{+}06, BLH04, GD03, LLV^{+}08, VL08].

References

Aedo:2003:EAA

[AA03] Gustavo Aedo and Hugo Arancibia. Estimating attraction areas and effective fishing areas for Chilean lemon crab (Cancer porteri) using traps. Fisheries Research, 60(2–3):267–272, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001777.

Ayaz:2006:GFM

[AAA^{+}06] Adnan Ayaz, Deniz Acarli, Ugur Altinagac, Ugur Ozekinci, Ali Kara, and Ozcan Ozen. Ghost fishing by monofilament and multifilament gillnets in Izmir Bay, Turkey. Fisheries Research, 79(3):267–271, July 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001391.

Araya:2005:IDC

[AAB05] Patricia R. Araya, Angelo A. Agostinho, and José A. Bechara. The influence of dam construction on a population of Leporinus obtusidens (Valenciennes, 1847) (Pisces, Anostomidae) in the Yacyretá Reservoir (Argentina). Fisheries Research, 74(1–3):198–209, August 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000640.

Abarca-Arenas:2007:SAR

[AAFLP^{+}07] L. G. Abarca-Arenas, J. Franco-Lopez, M. S. Peterson, N. J. Brown-Peterson, and E. Valero-Pacheco. Sociometric analysis of the role of penaeids in the continental shelf food web off Veracruz, Mexico based on bycatch. Fisheries Research, 87(1):46–57, October 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001464.

Archdale:2006:CFT

[AAH06] Miguel Vazquez Archdale, Cherry Pilapil Añasco, and Shinji Hiromori. Comparative fishing trials for invasive swimming crabs Charybdis japonica and Portunus pelagicus using collapsible pots. Fisheries Research, 82(1–3):50–55, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606003018.

Ama-Abasi:2004:DEP

[AAHE04] Daniel Ama-Abasi, Sieghard Holzloehner, and Udeme Enin. The dynamics of the exploited population of Ethmalosa fimbriata (Bowdich, 1825, Clupeidae) in the Cross River Estuary and adjacent Gulf of Guinea. Fisheries Research, 68(1–3):225–235, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000219.

Archdale:2007:ETC

[AAKT07] Miguel Vazquez Archdale, Cherry Pilapil Añasco, Yousuke Kawamura, and Sanae Tomiki. Effect of two collapsible pot designs on escape rate and behavior of the invasive swimming crabs Charybdis japonica and Portunus pelagicus. Fisheries Research, 85(1–2):202–209, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000513.

Alos:2009:ITN

[AAP^{+}09] Josep Alós, Robert Arlinghaus, Miquel Palmer, David March, and Itziar Álvarez. The influence of type of natural bait on fish catches and hooking location in a mixed-species marine recreational fishery, with implications for management. Fisheries Research, 97(3):270–277, May 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000708.

Archdale:2008:CSC

[AAT08] Miguel Vazquez Archdale, Cherry Pilapil Añasco, and Yoshihito Tahara. Catches of swimming crabs using fish mince in “teabags” compared to conventional fish baits in collapsible pots. Fisheries Research, 91(2–3):291–298, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607003529.

Andersen:2008:PITb

[AB08a] Niels G. Andersen and Jan E. Beyer. Precision of ingestion time and evacuation predictors for individual prey in stomachs of predatory fishes. Fisheries Research, 92(1):11– 22, July 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607003542.

Andersen:2008:PITa

[AB08b] Niels G. Andersen and Jan E. Beyer. Predicting ingestion times of individual prey from information about stomach contents of predatory fishes in the field. Fisheries Research, 92(1):1–10, July 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003505.

Aguzzi:2009:MDN

[AB09a] J. Aguzzi and N. Bahamon. Modeled day–night biases in decapod assessment by bottom trawling survey. Fisheries Research, 100(3):274–280, November 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002318.

Arocha:2009:SRS

[AB09b] Freddy Arocha and Alexander Bárrios. Sex ratios, spawning seasonality, sexual maturity, and fecundity of white marlin (Tetrapturus albidus) from the western central Atlantic. Fisheries Research, 95(1):98–111, January 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002300.

Abaunza:2008:HMI

[Aba08] Pablo Abaunza. Horse mackerel: Identification of stocks. Fisheries Research, 89(2):101–103, February 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002354.

Al-Barwani:2007:PDG

[ABAA^{+}07] S. M. Al-Barwani, A. Arshad, S. M. Nurul Amin, S. B. Japar, S. S. Siraj, and C. K. Yap. Population dynamics of the green mussel Perna viridis from the high spat-fall coastal water of Malacca, Peninsular Malaysia. Fisheries Research, 84(2):147–152, April 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003468.

Abecasis:2008:ASC

[ABC^{+}08] D. Abecasis, L. Bentes, R. Coelho, C. Correia, P. G. Lino, P. Monteiro, J. M. S. Gonçalves, J. Ribeiro, and K. Erzini. Ageing seabreams: a comparative study between scales and otoliths. Fisheries Research, 89(1):37–48, January 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002123.

Al-Baz:2007:FSH

[ABCB^{+}07] Ali F. Al-Baz, W. Chen, J. M. Bishop, M. Al-Husaini, and S. A. Al-Ayoub. On fishing selectivity of hadrah (fixed stake trap) in the coastal waters of Kuwait. Fisheries Research, 84(2):202–209, April 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600381X.

Allen:2009:ARF

[ABD^{+}09] M. S. Allen, P. Brown, J. Douglas, W. Fulton, and M. Catalano. An assessment of recreational fishery harvest policies for Murray cod in southeast Australia. Fisheries Research, 95(2–3):260–267, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003135.

Ayvazian:2004:SSA

[ABE^{+}04] Suzanne G. Ayvazian, Trevor P. Bastow, John S. Edmonds, Jason How, and Gabrielle B. Nowara. Stock structure of Australian herring (Arripis georgiana) in southwestern Australia. Fisheries Research, 67(1):39–53, March 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002686.

Arlinghaus:2008:UHR

[ABF08] Robert Arlinghaus, Martin Bork, and Erik Fladung. Understanding the heterogeneity of recreational anglers across an urban–rural gradient in a metropolitan area (Berlin, Germany), with implications for fisheries management. Fisheries Research, 92(1):53–62, July 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003645.

Abellan:2002:LHC

[ABFV02] L. J. López Abellán, E. Balguerıás, and V. Fernández-Vergaz. Life history characteristics of the deep-sea crab Chaceon affinis population off Tenerife (Canary Islands). Fisheries Research, 58(2):231–239, November 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003848.

Allain:2003:PEF

[ABK03] V. Allain, A. Biseau, and B. Kergoat. Preliminary estimates of French deepwater fishery discards in the Northeast Atlantic Ocean. Fisheries Research, 60(1):185–192, January 30, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602000310.

Acha:2002:CPE

[ABL02] Eduardo Marcelo Acha, Claudia Bremec, and Carlos Lasta. Cannibalism on planktonic eggs by a nonfilter feeding fish, Micropogonias furnieri (Sciaenidae). Fisheries Research, 56(3):321–326, June 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003265.

Ablan:2006:GSF

[Abl06] M. A. Carmen A. Ablan. Genetics and the study of fisheries connectivity in Asian developing countries. Fisheries Research, 78(2–3):158–168, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000609.

Arrhenius:2000:CSB

[ABRB00] Fredrik Arrhenius, Bastiaan J. A. M. Benneheij, Lars G. Rudstam, and Daniel Boisclair. Can stationary bottom split-beam hydroacoustics be used to measure fish swimming speed in situ? Fisheries Research, 45(1):31–41, February 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699001022.

Argent:2003:PFF

[ABS^{+}03] David G. Argent, Joseph A. Bishop, Jay R. Stauffer, Robert F. Carline, and Wayne L. Myers. Predicting freshwater fish distributions using landscape-level variables. Fisheries Research, 60(1):17–32, January 30, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000760.

Arrhenius:2000:LCM

[AC00] F. Arrhenius and M. Cardinale. Letter on “Comments made Mulligan and Chen ‘Can stationary bottom split-beam hydroacoustics be used to measure fish swimming speed in situ?”’. Fisheries Research, 49(1):97, November 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002617.

Andrade:2002:ACV

[AC02] Humber Agrelli Andrade and Rodrigo Oliveira Campos. Allometry coefficient variations of the length–weight relationship of skipjack tuna (Katsuwonus pelamis) caught in the southwest South Atlantic. Fisheries Research, 55 (1–3):307–312, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783601003058.

Alvarez:2005:GMH

[AC05] Paula Alvarez and Unai Cotano. Growth, mortality and hatch-date distributions of European hake larvae, Merluccius merluccius (L.), in the Bay of Biscay. Fisheries Research, 76(3):379–391, December 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002122.

Aldanondo:2008:VDI

[ACE^{+}08] N. Aldanondo, U. Cotano, E. Etxebeste, X. Irigoien, P. Álvarez, A. Martínez de Murguía, and D. L. Herrero. Validation of daily increments deposition in the otoliths of European anchovy larvae (Engraulis encrasicolus L.) reared under different temperature conditions. Fisheries Research, 93(3):257–264, September 28, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001379.

Araya:2001:ERB

[ACG^{+}01] Miguel Araya, Luis A. Cubillos, Margarita Guzmán, Jessica Peñailillo, and Aquiles Sepúlveda. Evidence of a relationship between age and otolith weight in the Chilean jack mackerel, Trachurus symmetricus murphyi (Nichols). Fisheries Research, 51(1):17–26, April 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002332.

Abitia-Cardenas:2002:TDS

[ACMMCEGM02] L. A. Abitıá-Cárdenas, A. Muhlia-Melo, V. Cruz-Escalona, and F. Galván-Magaña. Trophic dynamics and seasonal energetics of striped marlin Tetrapturus audax in the southern Gulf of California, Mexico. Fisheries Research, 57(3):287–295, September 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003502.

Abecasis:2006:AGB

[ACPP06] D. Abecasis, A. R. Costa, J. G. Pereira, and M. R. Pinho. Age and growth of bluemouth, Helicolenus dactylopterus (Delaroche, 1809) from the Azores. Fisheries Research, 79(1–2):148–154, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000725.

Aadlandsvik:2001:BVD

[ÅCST01] B. Ådlandsvik, S. Coombs, S. Sundby, and G. Temple. Buoyancy and vertical distribution of eggs and larvae of blue whiting (Micromesistius poutassou): observations and modelling. Fisheries Research, 50(1–2):59–72, February 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600002423.

Anderson:2000:IDH

[AD00] John T. Anderson and Edgar L. Dalley. Interannual differences in hatching times and growth rates of pelagic juvenile cod in Newfoundland waters. Fisheries Research, 46(1–3):227–238, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360000148X.

Allen:2003:DCY

[AD03] Scott Allen and David A. Demer. Detection and characterization of yellowfin and bluefin tuna using passive-acoustical techniques. Fisheries Research, 63(3):393–403, September 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603000961.

Alves:2002:AGS

[AdBP02] Adriana Alves, Pedro de Barros, and Mário R. Pinho. Age and growth studies of bigeye tuna Thunnus obesus from Madeira using vertebrae. Fisheries Research, 54(3): 389–393, February 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783601002685.

Addis:2008:ELS

[ADP^{+}08] Piero Addis, John Mark Dean, Paola Pesci, Ivan Locci, Rita Cannas, Stefano Corrias, and Angelo Cau. Effects of local scale perturbations in the Atlantic bluefin tuna (Thunnus thynnus L.) trap fishery of Sardinia (W. Mediterranean). Fisheries Research, 92(2–3):242–254, August 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000441.

Adlerstein:2003:PDV

[AE03] Sara Adlerstein and Siegfried Ehrich. Patterns in diel variation of cod catches in North Sea bottom trawl surveys. Fisheries Research, 63(2):169–178, August 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000742.

Appelberg:2005:CEE

[AFG^{+}05] M. Appelberg, N. Formigo, A. J. Geffen, C. Hammer, W. McCurdy, J. Modin, E. Moksness, H. Mosegaard, B. Morales-Nin, H. Troadec, and P. Wright. A cooperative effort to exchange age reading experience and protocols between European fish institutes. Fisheries Research, 76(2):167–173, November 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001700.

Adams:2008:SCP

[AFJC08] Charles F. Adams, Robert J. Foy, Devin S. Johnson, and Kenneth O. Coyle. Seasonal changes in pelagic fish biomass around the Chiswell Island Steller sea lion rookery in 2003. Fisheries Research, 93(1–2):179–185, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001161.

Allman:2005:PAE

[AFSF05] Robert J. Allman, Gary R. Fitzhugh, Karl J. Starzinger, and Robert A. Farsky. Precision of age estimation in red snapper (Lutjanus campechanus). Fisheries Research, 73 (1–2):123–133, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604003285.

Alonso-Fernandez:2009:FEA

[AFVSR^{+}09] Alexandre Alonso-Fernández, Ann Carole Vallejo, Fran Saborido-Rey, Hilario Murua, and Edward A. Trippel. Fecundity estimation of Atlantic cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) of Georges Bank: Application of the autodiametric method. Fisheries Research, 99(1):47–54, July 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001039.

Arendse:2007:CFS

[AGB07] Clement J. Arendse, Anesh Govender, and George M. Branch. Are closed fishing seasons an effective means of increasing reproductive output?: a per-recruit simulation using the limpet Cymbula granatina as a case history. Fisheries Research, 85(1–2):93–100, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000082.

Abaunza:2008:LHP

[AGS^{+}08] P. Abaunza, L. S. Gordo, M. T. García Santamaría, S. A. Iversen, A. G. Murta, and E. Gallo. Life history parameters as basis for the initial recognition of stock management units in horse mackerel (Trachurus trachurus).

Fisheries Research, 89(2):167–180, February 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002445.

Al-Husaini:2001:AVN

[AHAAD01] M. Al-Husaini, S. Al-Ayoub, and J. Dashti. Age validation of nagroor, Pomadasys kaakan (Cuvier, 1830) (family: Haemulidae) in Kuwaiti waters. Fisheries Research, 53(1): 71–81, September 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360000268X.

Al-Husaini:2002:AGM

[AHABAA^{+}02] M. Al-Husaini, A. Al-Baz, S. Al-Ayoub, S. Safar, Z. Al-Wazan, and S. Al-Jazzaf. Age, growth, mortality, and yieldper-recruit for nagroor, Pomadasys kakaan, in Kuwait’s waters. Fisheries Research, 59(1–2):101–115, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601004179.

Astles:2006:EMQ

[AHS^{+}06] K. L. Astles, M. G. Holloway, A. Steffe, M. Green, C. Ganassin, and P. J. Gibbs. An ecological method for qualitative risk assessment and its use in the management of fisheries in New South Wales, Australia. Fisheries Research, 82(1–3):290–303, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600213X.

Arce-Ibarra:2008:IFM

[AIC08] A. Minerva Arce-Ibarra and Anthony T. Charles. Inland fisheries of the Mayan Zone in Quintana Roo, Mexico: Using a combined approach to fishery assessment for datasparse fisheries. Fisheries Research, 91(2–3):151–159, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607003360.

Allen:2001:ESM

[AKA^{+}01] Michelle Allen, David Kilpatrick, Mike Armstrong, Richard Briggs, Nélida Pérez, and Grant Course. Evaluation of sampling methods to quantify discarded fish using data collected during discards project EC 95/094 by Northern Ireland, England and Spain. Fisheries Research, 49(3): 241–254, January 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783600002071.

Allen:2002:MCS

[AKA^{+}02] Michelle Allen, David Kilpatrick, Mike Armstrong, Richard Briggs, Grant Course, and Nélida Pérez. Multistage cluster sampling design and optimal sample sizes for estimation of fish discards from commercial trawlers. Fisheries Research, 55(1–3):11–24, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003083.

Archdale:2006:ETP

[AKA06] Miguel Vazquez Archdale, Lika Kariyazono, and Cherry Pilapil Añasco. The effect of two pot types on entrance rate and entrance behavior of the invasive Japanese swimming crab Charybdis japonica. Fisheries Research, 77(3): 271–274, March 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606000130.

Arlinghaus:2009:PBC

[AKC^{+}09] Robert Arlinghaus, Thomas Klefoth, Steven J. Cooke, Andrew Gingerich, and Cory Suski. Physiological and behavioural consequences of catch-and-release angling on northern pike (Esox lucius L.). Fisheries Research, 97 (3):223–233, May 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783609000472.

Armstrong:2003:HRA

[AKK^{+}03] J. D. Armstrong, P. S. Kemp, G. J. A. Kennedy, M. Ladle, and N. J. Milner. Habitat requirements of Atlantic salmon and brown trout in rivers and streams.

Fisheries Research, 62(2):143–170, May 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001601.

Allain:2001:RST

[All01] Valérie Allain. Reproductive strategies of three deepwater benthopelagic fishes from the northeast Atlantic Ocean. Fisheries Research, 51(2–3):165–176, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002430.

Allman:2007:SSS

[All07] Robert J. Allman. Small-scale spatial variation in the population structure of vermilion snapper (Rhomboplites aurorubens) from the northeast Gulf of Mexico. Fisheries Research, 88(1–3):88–99, December 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001865.

Addison:2003:AVC

[ALN03] J. T. Addison, A. R. Lawler, and M. D. Nicholson. Adjusting for variable catchability of brown shrimps (Crangon crangon) in research surveys. Fisheries Research, 65(1–3):285–294, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002492.

Alos:2009:MIR

[Aló09] Josep Alós. Mortality impact of recreational angling techniques and hook types on Trachynotus ovatus (Linnaeus, 1758) following catch-and-release. Fisheries Research, 95(2–3):365–369, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002464.

Aoki:2000:DLJ

[AM00] Ichiro Aoki and Kazushi Miyashita. Dispersal of larvae and juveniles of Japanese anchovy Engraulis japonicus in the Kuroshio Extension and Kuroshio–Oyashio transition regions, western North Pacific Ocean. Fisheries Research, 49(2):155–164, December 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001971.

Arkhipkin:2002:IPA

[AM02] Alexander I. Arkhipkin and David A. J. Middleton. Inverse patterns in abundance of Illex argentinus and Loligo gahi in Falkland waters: possible interspecific competition between squid? Fisheries Research, 59(1–2):181–196, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601004143.

Arlinghaus:2003:SEC

[AM03] Robert Arlinghaus and Thomas Mehner. Socio-economic characterisation of specialised common carp (Cyprinus carpio L.) anglers in Germany, and implications for inland fisheries management and eutrophication control. Fisheries Research, 61(1–3):19–33, March 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002436.

Al-Masroori:2004:CLF

[AMAOMM04] H. Al-Masroori, H. Al-Oufi, J. L. McIlwain, and E. McLean. Catches of lost fish traps (ghost fishing) from fishing grounds near Muscat, Sultanate of Oman. Fisheries Research, 69(3):407–414, October 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001444.

Abaunza:2008:CSS

[AMC^{+}08a] P. Abaunza, A. G. Murta, N. Campbell, R. Cimmaruta, A. S. Comesaña, G. Dahle, E. Gallo, M. T. García Santamaría, L. S. Gordo, S. A. Iversen, K. MacKenzie, A. Magoulas, S. Mattiucci, J. Molloy, G. Nascetti, A. L. Pinto, R. Quinta, P. Ramos, A. Ruggi, A. Sanjuan, A. T. Santos, C. Stransky, and C. Zimmermann. Considerations on sampling strategies for an holistic approach to stock identification: the example of the HOMSIR project. Fisheries Research, 89(2):104–113, February 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002366.

Abaunza:2008:SIH

[AMC^{+}08b] P. Abaunza, A. G. Murta, N. Campbell, R. Cimmaruta, A. S. Comesaña, G. Dahle, M. T. García Santamaría, L. S. Gordo, S. A. Iversen, K. MacKenzie, A. Magoulas, S. Mattiucci, J. Molloy, G. Nascetti, A. L. Pinto, R. Quinta, P. Ramos, A. Sanjuan, A. T. Santos, C. Stransky, and C. Zimmermann. Stock identity of horse mackerel (Trachurus trachurus) in the Northeast Atlantic and Mediterranean Sea: Integrating the results from different stock identification approaches. Fisheries Research, 89(2):196–209, February 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607002470.

Afonso:2008:SPR

[AMS08] Pedro Afonso, Telmo Morato, and Ricardo Serrão Santos. Spatial patterns in reproductive traits of the temperate parrotfish Sparisoma cretense. Fisheries Research, 90(1–3):92– 99, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607002500.

Alvarez:2001:STD

[AMUE01] Paula Alvarez, Lorenzo Motos, Adolfo Uriarte, and Joseba Egaña. Spatial and temporal distribution of European hake, Merluccius merluccius (L.), eggs and larvae in relation to hydrographical conditions in the Bay of Biscay. Fisheries Research, 50(1–2):111–128, February 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002459.

Anonymous:2000:AIV

[Ano00a] Anonymous. Author index, volumes 39–43. Fisheries Research, 44(3):281–289, January 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600900213.

Anonymous:2000:BRC

[Ano00b] Anonymous. Book review: Cellular and Molecular Approaches to Fish Ionic Regulation: C. M. Wood, T. J. Shuttleworth (Eds.); Academic Press, New York, 1995, 352 pages, hardback, ISBN 0-12-350438-4. Fisheries Research, 48(3):292, October 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002265.

Anonymous:2000:BRL

[Ano00c] Anonymous. Book review: Life and Death of Coral Reefs: C. Birkeland (Ed.); Chapman & Hall, London, 1997, 536 pages, ISBN 0-412-03541-3. Fisheries Research, 48(3):291, October 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600002253.

Anonymous:2000:BRSa

[Ano00d] Anonymous. Book review: Sponges of the New Caledonian Lagoon: Claude Lévi (Ed.); ORSTOM éditions, 1998, 214 pages, 270 FF hardback, ISBN 2-7099-1354-2. Fisheries Research, 48(3):291, October 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002241.

Anonymous:2000:BRSb

[Ano00e] Anonymous. Book review: Subtropical convergence environments: the coast and sea in the southwest Atlantic: U. Seeliger, C. Odebrecht, J. P. Costello (Eds.); Springer, Berlin, 1997, 308 pages, 178 DM hardback, ISBN 3-540-61365-X. Fisheries Research, 48(3):292, October 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002277.

Anonymous:2000:Ia

[Ano00f] Anonymous. Index. Fisheries Research, 48(3):295–305, October 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600002526.

Anonymous:2000:Ib

[Ano00g] Anonymous. Index. Fisheries Research, 48(3):307–330, October 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600002538.

Anonymous:2000:PF

[Ano00h] Anonymous. Pages 1–104 (February 2000). Fisheries Research, 45(1):??, February 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2000:PJb

[Ano00i] Anonymous. Pages 1–104 (June 2000). Fisheries Research, 47(1):??, June 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2000:PN

[Ano00j] Anonymous. Pages 1–104 (November 2000). Fisheries Research, 49(1):??, November 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2000:PMb

[Ano00k] Anonymous. Pages 1–373 (May 2000). Fisheries Research, 46(1–3):??, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2000:PAb

[Ano00l] Anonymous. Pages 1–95 (August 2000). Fisheries Research, 48(1):??, August 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2000:PMa

[Ano00m] Anonymous. Pages 105–203 (March 2000). Fisheries Research, 45(2):??, March 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2000:PD

[Ano00n] Anonymous. Pages 105–206 (December 2000). Fisheries Research, 49(2):??, December 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2000:PJc

[Ano00o] Anonymous. Pages 105–262 (July 2000). Fisheries Research, 47(2–3):??, July 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2000:PJa

[Ano00p] Anonymous. Pages 201–316 (January 2000). Fisheries Research, 44(3):??, January 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2000:PO

[Ano00q] Anonymous. Pages 205–334 (October 2000). Fisheries Research, 48(3):??, October 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2000:PAa

[Ano00r] Anonymous. Pages 207–314 (April 2000). Fisheries Research, 45(3):??, April 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2000:PS

[Ano00s] Anonymous. Pages 99–203 (1 September 2000). Fisheries Research, 48(2):??, September 1, 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2000:SIV

[Ano00t] Anonymous. Subject index, volumes 39–43. Fisheries Research, 44(3):291–313, January 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600900237.

Anonymous:2001:Aa

[Ano01a] Anonymous. Acknowledgements. Fisheries Research, 52(1– 2):3, June 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601002259.

Anonymous:2001:Ab

[Ano01b] Anonymous. ANNEX 1. Fisheries Research, 52(1–2):141– 142, June 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601002375.

Anonymous:2001:BRFa

[Ano01c] Anonymous. Book review: Fish Morphology, Horizon of New Research: J. S. Datta Munshi, H. D. Dutta (Eds.); Sp Science Publishers, USA, 1996, 300 pages, hardback, ISBN 1-886106-31-2 ($99). Fisheries Research, 51(1):99, April 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600003131.

Anonymous:2001:BRFb

[Ano01d] Anonymous. Book review: Fishing Places, Fishing People: Dianne Newell, Rosemary E. Ommer (Eds.); University of Toronto Press, Toronto, 1999, 374 pages, hardback, ISBN 0-8020-7959-8 (£16.50). Fisheries Research, 51(1):100, April 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600003167.

Anonymous:2001:BRR

[Ano01e] Anonymous. Book review: Recent Advances in Marine Biotechnology, Vol. 3. Biofilms, Bioadhesion, Corrosion, and Biofouling: Science Publishers, Inc., NH, USA, 1999, 312 pages, hardback, ISBN 1-57808-013-4 (£58). Fisheries Research, 51(1):99–100, April 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600003155.

Anonymous:2001:BRS

[Ano01f] Anonymous. Book review: The sea fisheries of Scotland: J. R. Coull (Ed.); John Donald Publishers, Edinburgh, 1996, 308 pages, ISBN 0-85976-410-9, £30 (hardback). Fisheries Research, 51(1):99, April 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600003143.

Anonymous:2001:CVa

[Ano01g] Anonymous. Contents of volume. Fisheries Research, 51 (2–3):415–417, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783601002909.

Anonymous:2001:CVb

[Ano01h] Anonymous. Contents of volume 53. Fisheries Research, 53 (3):309–310, November 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360100368X.

Anonymous:2001:G

[Ano01i] Anonymous. (Gordon S.I.). Fisheries Research, 51(2–3): ??, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2001:PAa

[Ano01j] Anonymous. Pages 1–103 (April 2001). Fisheries Research, 51(1):??, April 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2001:PS

[Ano01k] Anonymous. Pages 1–104 (September 2001). Fisheries Research, 53(1):??, September 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2001:PF

[Ano01l] Anonymous. Pages 1–204 (February 2001). Fisheries Research, 50(1–2):??, February 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2001:PO

[Ano01m] Anonymous. Pages 105–210 (October 2001). Fisheries Research, 53(2):??, October 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2001:PJb

[Ano01n] Anonymous. Pages 143–232 (July 2001). Fisheries Research, 52(3):??, July 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2001:PM

[Ano01o] Anonymous. Pages 205–304 (March 2001). Fisheries Research, 50(3):??, March 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2001:PJa

[Ano01p] Anonymous. Pages 207–310 (January 2001). Fisheries Research, 49(3):??, January 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2001:PN

[Ano01q] Anonymous. Pages 211–310 (November 2001). Fisheries Research, 53(3):??, November 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2001:P

[Ano01r] Anonymous. (Pierce S.I.). Fisheries Research, 52(1–2): ??, June 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2001:PAb

[Ano01s] Anonymous. Publisher’s acknowledgement. Fisheries Research, 51(2–3):413, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0165783601002910.

Anonymous:2001:PAc

[Ano01t] Anonymous. Publisher’s acknowledgement. Fisheries Research, 52(3):231–232, July 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0165783601003150.

Anonymous:2002:A

[Ano02a] Anonymous. Announcement. Fisheries Research, 56(1): 115, April 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360200022X.

Anonymous:2002:AIV

[Ano02b] Anonymous. Author index, volumes 54-58. Fisheries Research, 58(3):413–424, November 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001960.

Anonymous:2002:CFRa

[Ano02c] Anonymous. Contents of fisheries research volume 56. Fisheries Research, 56(3):327–328, June 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000504.

Anonymous:2002:CFRb

[Ano02d] Anonymous. Contents of fisheries research, volume 57. Fisheries Research, 57(3):313–314, September 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001054.

Anonymous:2002:CFRc

[Ano02e] Anonymous. Contents of fisheries research, volume 58. Fisheries Research, 58(3):453–455, November 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002047.

Anonymous:2002:CVa

[Ano02f] Anonymous. Contents of volume 54. Fisheries Research, 54 (3):419–421, February 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783601004064.

Anonymous:2002:CVb

[Ano02g] Anonymous. Contents of volume 55. Fisheries Research, 55 (1–3):341–342, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783601004167.

Anonymous:2002:FEA

[Ano02h] Anonymous. Fulltitle (Editorial Advisory Board). Fisheries Research, 59(1–2):iii, December 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002242.

Anonymous:2002:GA

[Ano02i] Anonymous. Guide for authors. Fisheries Research, 54 (2):299–302, January 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783601004052.

Anonymous:2002:IEAa

[Ano02j] Anonymous. IFC — (Editorial Advisory Board). Fisheries Research, 58(2):ifc, November 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001509.

Anonymous:2002:IEAb

[Ano02k] Anonymous. IFC — (Editorial Advisory Board). Fisheries Research, 58(3):ifc, November 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360200200X.

Anonymous:2002:OTF

[Ano02l] Anonymous. Overfishing, tropicalization of fish. Fisheries Research, 55(1–3):1–9, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100279X.

Anonymous:2002:PJc

[Ano02m] Anonymous. Pages 1–102 (July 2002). Fisheries Research, 57(1):??, July 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2002:PAa

[Ano02n] Anonymous. Pages 1–116 (April 2002). Fisheries Research, 56(1):??, April 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2002:PO

[Ano02o] Anonymous. Pages 1–118 (October 2002). Fisheries Research, 58(1):??, October 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2002:PD

[Ano02p] Anonymous. Pages 1–304 (30 December 2002). Fisheries Research, 59(1–2):??, December 30, 2002. CODEN FIS-RDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2002:PMa

[Ano02q] Anonymous. Pages 1–344 (March 2002). Fisheries Research, 55(1–3):??, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2002:PAc

[Ano02r] Anonymous. Pages 103–210 (August 2002). Fisheries Research, 57(2):??, August 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2002:PMb

[Ano02s] Anonymous. Pages 117–226 (May 2002). Fisheries Research, 56(2):??, May 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2002:PNa

[Ano02t] Anonymous. Pages 119–254 (November 2002). Fisheries Research, 58(2):??, November 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2002:PJa

[Ano02u] Anonymous. Pages 145–304 (January 2002). Fisheries Research, 54(2):??, January 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2002:PS

[Ano02v] Anonymous. Pages 211–314 (September 2002). Fisheries Research, 57(3):??, September 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2002:PJb

[Ano02w] Anonymous. Pages 227–328 (June 2002). Fisheries Research, 56(3):??, June 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2002:PNb

[Ano02x] Anonymous. Pages 255–458 (30 November 2002). Fisheries Research, 58(3):??, November 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2002:PF

[Ano02y] Anonymous. Pages 305–422 (February 2002). Fisheries Research, 54(3):??, February 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2002:PAb

[Ano02z] Anonymous. Publisher’s acknowledgement. Fisheries Research, 56(1):113, April 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783602000218.

Anonymous:2002:SN

[Ano02-27] Anonymous. Shorter notices. Fisheries Research, 56 (2):219–220, May 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360200005X.

Anonymous:2002:SIV

[Ano02-28] Anonymous. Subject index, volumes 54-58. Fisheries Research, 58(3):425–451, November 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001972.

Anonymous:2002:WC

[Ano02-29] Anonymous. The whaling controversy. Fisheries Research, 54(2):145–151, January 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0165783600002939.

Anonymous:2003:CFRa

[Ano03a] Anonymous. Contents of Fisheries Research, volume. Fisheries Research, 59(3):447–449, January 3, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002291.

Anonymous:2003:CFRb

[Ano03b] Anonymous. Contents of Fisheries Research, volume 60. Fisheries Research, 60(2–3):593–595, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000079.

Anonymous:2003:CFRc

[Ano03c] Anonymous. Contents of Fisheries Research, volume 61. Fisheries Research, 61(1–3):155, March 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000389.

Anonymous:2003:CFRd

[Ano03d] Anonymous. Contents of Fisheries Research, volume 62. Fisheries Research, 62(3):389–390, June 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001188.

Anonymous:2003:CFRe

[Ano03e] Anonymous. Contents of Fisheries Research, volume 63. Fisheries Research, 63(3):455–457, September 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001711.

Anonymous:2003:FTPa

[Ano03f] Anonymous. Full title page (Editorial Advisory Board). Fisheries Research, 63(1):iii, July 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http:

//www.sciencedirect.com/science/article/pii/S0165783603001383.

Anonymous:2003:FTPb

[Ano03g] Anonymous. Full title page (Editorial Advisory Board). Fisheries Research, 64(1):iii, October 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002819.

Anonymous:2003:FTPc

[Ano03h] Anonymous. Full title page (Editorial Advisory Board). Fisheries Research, 65(1–3):iii, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603003345.

Anonymous:2003:FEAb

[Ano03i] Anonymous. Fulltitle — (Editorial Advisory Board). Fisheries Research, 61(1–3):iii, March 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000298.

Anonymous:2003:FEAc

[Ano03j] Anonymous. Fulltitle — (Editorial Advisory Board). Fisheries Research, 62(1):iii, April 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000523.

Anonymous:2003:FEAa

[Ano03k] Anonymous. Fulltitle — (Editorial Advisory Board) (1st issue of volume). Fisheries Research, 60(1):iii, January 30, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602002667.

Anonymous:2003:GA

[Ano03l] Anonymous. Guide for authors. Fisheries Research, 60(1): 193–196, January 30, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360200259X.

Anonymous:2003:IEAc

[Ano03m] Anonymous. IFC — (editorial advisory). Fisheries Research, 62(2):ifc, May 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783603000845.

Anonymous:2003:IEAa

[Ano03n] Anonymous. IFC — (Editorial Advisory Board). Fisheries Research, 59(3):ifc, January 3, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002333.

Anonymous:2003:IEAb

[Ano03o] Anonymous. IFC — (Editorial Advisory Board). Fisheries Research, 60(2–3):ifc, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000110.

Anonymous:2003:IFCa

[Ano03p] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 62(3):ifc, June 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001140.

Anonymous:2003:IFCb

[Ano03q] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 63(2):ifc, August 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001589.

Anonymous:2003:IFCc

[Ano03r] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 63(3):ifc, September 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360300167X.

Anonymous:2003:IFCd

[Ano03s] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 64(2–3):ifc, November 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603003163.

Anonymous:2003:PAa

[Ano03t] Anonymous. Pages 1–108 (April 2003). Fisheries Research, 62(1):??, April 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2003:PJd

[Ano03u] Anonymous. Pages 1–148 (July 2003). Fisheries Research, 63(1):??, July 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2003:PM

[Ano03v] Anonymous. Pages 1–156 (March 2003). Fisheries Research, 61(1–3):??, March 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2003:PJb

[Ano03w] Anonymous. Pages 1–198 (30 January 2003). Fisheries Research, 60(1):??, January 30, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2003:PO

[Ano03x] Anonymous. Pages 1–98 (October 2003). Fisheries Research, 64(1):??, October 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2003:PAc

[Ano03y] Anonymous. Pages 149–302 (August 2003). Fisheries Research, 63(2):??, August 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2003:PF

[Ano03z] Anonymous. Pages 199–598 (February 2003). Fisheries Research, 60(2–3):??, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2003:PJc

[Ano03-27] Anonymous. Pages 235–392 (June 2003). Fisheries Research, 62(3):??, June 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2003:PS

[Ano03-28] Anonymous. Pages 303–460 (September 2003). Fisheries Research, 63(3):??, September 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2003:PJa

[Ano03-29] Anonymous. Pages 305–450 (3 January 2003). Fisheries Research, 59(3):??, January 3, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2003:PNa

[Ano03-30] Anonymous. Pages 99–214 (November 2003). Fisheries Research, 64(2–3):??, November 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2003:PAb

[Ano03-31] Anonymous. Publisher’s acknowledgement. Fisheries Research, 62(1):105, April 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783603000456.

Anonymous:2003:PNb

[Ano03-32] Anonymous. Publisher’s note. Fisheries Research, 64(2– 3):99, November 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783603003229.

Anonymous:2003:VCa

[Ano03-33] Anonymous. Volume contents. Fisheries Research, 64(2–3): 211–212, November 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783603003126.

Anonymous:2003:VCb

[Ano03-34] Anonymous. Volume contents. Fisheries Research, 65(1–3): 507–508, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783603003400.

Anonymous:2004:EABa

[Ano04a] Anonymous. Editorial Advisory Board. Fisheries Research, 67(1):iii, March 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604000098.

Anonymous:2004:EABb

[Ano04b] Anonymous. Editorial advisory board. Fisheries Research, 68(1–3):iii, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604001006.

Anonymous:2004:EABc

[Ano04c] Anonymous. (Editorial Advisory Board). Fisheries Research, 69(1):iii, August 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S016578360400116X.

Anonymous:2004:EB

[Ano04d] Anonymous. Editorial Board. Fisheries Research, 67(2): ifc, April 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604000396.

Anonymous:2004:FTPa

[Ano04e] Anonymous. Full title page (Editorial Advisory Board). Fisheries Research, 66(1):iii, January 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603003552.

Anonymous:2004:FTPb

[Ano04f] Anonymous. Full title page (Editorial Advisory Board). Fisheries Research, 70(1):iii, November 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002437.

Anonymous:2004:GAb

[Ano04g] Anonymous. Guide for author. Fisheries Research, 68 (1–3):375–379, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604000918.

Anonymous:2004:GAa

[Ano04h] Anonymous. Guide for authors. Fisheries Research, 66 (1):121–125, January 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783603003473.

Anonymous:2004:I

[Ano04i] Anonymous. IFC. Fisheries Research, 67(3):ifc, May 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000542.

Anonymous:2004:IFCa

[Ano04j] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 66(2–3):ifc, February 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603003643.

Anonymous:2004:IFCb

[Ano04k] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 69(2):ifc, September 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002024.

Anonymous:2004:IFCc

[Ano04l] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 69(3):ifc, October 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002218.

Anonymous:2004:IFCd

[Ano04m] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 70(2–3):ifc, December 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002723.

Anonymous:2004:PJa

[Ano04n] Anonymous. Pages 1–128 (January 2004). Fisheries Research, 66(1):??, January 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2004:PNb

[Ano04o] Anonymous. Pages 1–136 (November 2004). Fisheries Research, 70(1):??, November 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2004:PAc

[Ano04p] Anonymous. Pages 1–146 (August 2004). Fisheries Research, 69(1):??, August 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2004:PJb

[Ano04q] Anonymous. Pages 1–384 (July 2004). Fisheries Research, 68(1–3):??, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2004:PMa

[Ano04r] Anonymous. Pages 1–98 (March 2004). Fisheries Research, 67(1):??, March 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2004:PF

[Ano04s] Anonymous. Pages 129–390 (February 2004). Fisheries Research, 66(2–3):??, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2004:PS

[Ano04t] Anonymous. Pages 147–302 (September 2004). Fisheries Research, 69(2):??, September 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2004:PMb

[Ano04u] Anonymous. Pages 253–336 (May 2004). Fisheries Research, 67(3):??, May 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2004:PO

[Ano04v] Anonymous. Pages 303–430 (October 2004). Fisheries Research, 69(3):??, October 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2004:PAa

[Ano04w] Anonymous. Pages 99–252 (April 2004). Fisheries Research, 67(2):??, April 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2004:PAb

[Ano04x] Anonymous. Publisher’s acknowledgement. Fisheries Research, 67(2):249, April 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604000359.

Anonymous:2004:PNa

[Ano04y] Anonymous. Publisher’s note. Fisheries Research, 66(1):1, January 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603003485.

Anonymous:2004:VCa

[Ano04z] Anonymous. Volume contents. Fisheries Research, 66(2– 3):387–388, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783603003606.

Anonymous:2004:VCb

[Ano04-27] Anonymous. Volume contents. Fisheries Research, 67 (3):335–336, May 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604000517.

Anonymous:2004:VCc

[Ano04-28] Anonymous. Volume contents. Fisheries Research, 68(1– 3):381–382, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360400092X.

Anonymous:2004:VCd

[Ano04-29] Anonymous. Volume contents. Fisheries Research, 69(3): 427–429, October 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604002176.

Anonymous:2004:VCe

[Ano04-30] Anonymous. Volume contents. Fisheries Research, 70(2–3): 427–428, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604002693.

Anonymous:2005:Aa

[Ano05a] Anonymous. Announcement. Fisheries Research, 73(1–2): 269, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605001050.

Anonymous:2005:Ab

[Ano05b] Anonymous. Announcement. Fisheries Research, 75(1– 3):1, September 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783605001803.

Anonymous:2005:CFR

[Ano05c] Anonymous. Contents of Fisheries Research, volume 76. Fisheries Research, 76(3):483–485, December 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002535.

Anonymous:2005:FTPa

[Ano05d] Anonymous. Full title page (Editorial Advisory Board). Fisheries Research, 71(1):iii, January 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002942.

Anonymous:2005:FTPb

[Ano05e] Anonymous. Full title page (Editorial Advisory Board). Fisheries Research, 72(1):iii, April 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360500055X.

Anonymous:2005:FTPc

[Ano05f] Anonymous. Full title page (Editorial Advisory Board). Fisheries Research, 73(1–2):iii, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001013.

Anonymous:2005:FTPd

[Ano05g] Anonymous. Full title page (Editorial Advisory Board). Fisheries Research, 74(1–3):iii, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001591.

Anonymous:2005:FTPe

[Ano05h] Anonymous. Full title page (Editorial Advisory Board). Fisheries Research, 75(1–3):iii, September 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001931.

Anonymous:2005:FTPf

[Ano05i] Anonymous. Full title page (Editorial Advisory Board). Fisheries Research, 76(1):iii, October 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002274.

Anonymous:2005:FTPg

[Ano05j] Anonymous. Full title page (Editorial Advisory Board). Fisheries Research, 76(3):iii, December 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002626.

Anonymous:2005:GA

[Ano05k] Anonymous. Guide for authors. Fisheries Research, 71 (1):133–137, January 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604002863.

Anonymous:2005:IFCa

[Ano05l] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 71(2):CO2, February 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604003200.

Anonymous:2005:IFCb

[Ano05m] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 71(3):CO2, March 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000123.

Anonymous:2005:IFCc

[Ano05n] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 72(2–3):CO2, May 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000767.

Anonymous:2005:IFCd

[Ano05o] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 73(3):CO2, July 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001177.

Anonymous:2005:IFCe

[Ano05p] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 76(2):CO2, November 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002420.

Anonymous:2005:PAa

[Ano05q] Anonymous. Pages 1–138 (April 2005). Fisheries Research, 72(1):??, April 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2005:PJa

[Ano05r] Anonymous. Pages 1–140 (January 2005). Fisheries Research, 71(1):??, January 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2005:PO

[Ano05s] Anonymous. Pages 1–166 (October 2005). Fisheries Research, 76(1):??, October 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2005:PS

[Ano05t] Anonymous. Pages 1–208 (September 2005). Fisheries Research, 75(1–3):??, September 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2005:PJb

[Ano05u] Anonymous. Pages 1–270 (June 2005). Fisheries Research, 73(1–2):??, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2005:PAc

[Ano05v] Anonymous. Pages 1–316 (August 2005). Fisheries Research, 74(1–3):??, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2005:PF

[Ano05w] Anonymous. Pages 141–256 (February 2005). Fisheries Research, 71(2):??, February 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2005:PMb

[Ano05x] Anonymous. Pages 141–368 (May 2005). Fisheries Research, 72(2–3):??, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2005:PN

[Ano05y] Anonymous. Pages 167–306 (November 2005). Fisheries Research, 76(2):??, November 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2005:PMa

[Ano05z] Anonymous. Pages 259–382 (March 2005). Fisheries Research, 71(3):??, March 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2005:PJc

[Ano05-27] Anonymous. Pages 273–384 (July 2005). Fisheries Research, 73(3):??, July 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2005:PD

[Ano05-28] Anonymous. Pages 307–486 (December 2005). Fisheries Research, 76(3):??, December 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2005:PAb

[Ano05-29] Anonymous. Publisher’s acknowledgement. Fisheries Research, 72(1):137–138, April 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0165783605000470.

Anonymous:2005:VCa

[Ano05-30] Anonymous. Volume contents. Fisheries Research, 71 (3):379–381, March 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783605000081.

Anonymous:2005:VCb

[Ano05-31] Anonymous. Volume contents. Fisheries Research, 72(2– 3):365–367, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783605000718.

Anonymous:2005:VCc

[Ano05-32] Anonymous. Volume contents. Fisheries Research, 73 (3):383–384, July 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360500113X.

Anonymous:2005:VCd

[Ano05-33] Anonymous. Volume contents. Fisheries Research, 74(1– 3):314–315, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783605001517.

Anonymous:2005:VCe

[Ano05-34] Anonymous. Volume contents. Fisheries Research, 75(1–3): 206–207, September 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783605001852.

Anonymous:2006:A

[Ano06a] Anonymous. Announcement. Fisheries Research, 81(2–3):I, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606003316.

Anonymous:2006:CFRa

[Ano06b] Anonymous. Contents of Fisheries Research, volume 77. Fisheries Research, 77(3):356–357, March 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600049X.

Anonymous:2006:CFRb

[Ano06c] Anonymous. Contents of Fisheries Research, volume 78. Fisheries Research, 78(2–3):392–394, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000968.

Anonymous:2006:CFRc

[Ano06d] Anonymous. Contents of Fisheries Research, volume 79. Fisheries Research, 79(3):353–354, July 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001871.

Anonymous:2006:CFRd

[Ano06e] Anonymous. Contents of Fisheries Research, volume 80. Fisheries Research, 80(2–3):355–356, September 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002530.

Anonymous:2006:CFRe

[Ano06f] Anonymous. Contents of Fisheries Research, volume 81. Fisheries Research, 81(2–3):345–346, November 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003237.

Anonymous:2006:CFRf

[Ano06g] Anonymous. Contents of Fisheries Research, volume 82. Fisheries Research, 82(1–3):332–333, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003602.

Anonymous:2006:EABa

[Ano06h] Anonymous. Editorial Advisory Board. Fisheries Research, 79(1–2):iii, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606001561.

Anonymous:2006:EABb

[Ano06i] Anonymous. Editorial Advisory Board. Fisheries Research, 79(3):CO2, July 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606001846.

Anonymous:2006:EABc

[Ano06j] Anonymous. Editorial Advisory Board. Fisheries Research, 80(1):CO2, August 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606001962.

Anonymous:2006:EABd

[Ano06k] Anonymous. Editorial Advisory Board. Fisheries Research, 81(1):iii, October 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606002700.

Anonymous:2006:EABe

[Ano06l] Anonymous. Editorial Advisory Board. Fisheries Research, 81(2–3):CO2, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0165783606003201.

Anonymous:2006:EABf

[Ano06m] Anonymous. Editorial Advisory Board. Fisheries Research, 82(1–3):iii, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606003535.

Anonymous:2006:FTP

[Ano06n] Anonymous. Full title page (Editorial Advisory Board). Fisheries Research, 77(1):iii, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003152.

Anonymous:2006:GA

[Ano06o] Anonymous. Guide for authors. Fisheries Research, 77 (1):125–128, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783605003206.

Anonymous:2006:I

[Ano06p] Anonymous. IFC. Fisheries Research, 77(3):CO2, March 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606000464.

Anonymous:2006:IFCa

[Ano06q] Anonymous. Inside front cover. Fisheries Research, 77 (2):CO2, February 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606000038.

Anonymous:2006:IFCb

[Ano06r] Anonymous. Inside front cover (Editorial Advisory Board). Fisheries Research, 78(1):CO2, April 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000798.

Anonymous:2006:PO

[Ano06s] Anonymous. Pages 1–106 (October 2006). Fisheries Research, 81(1):??, October 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2006:PJa

[Ano06t] Anonymous. Pages 1–128 (January 2006). Fisheries Research, 77(1):??, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2006:PJb

[Ano06u] Anonymous. Pages 1–232 (June 2006). Fisheries Research, 79(1–2):??, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2006:PD

[Ano06v] Anonymous. Pages 1–334 (December 2006). Fisheries Research, 82(1–3):??, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2006:PN

[Ano06w] Anonymous. Pages 107–348 (November 2006). Fisheries Research, 81(2–3):??, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2006:PMb

[Ano06x] Anonymous. Pages 107–394 (May 2006). Fisheries Research, 78(2–3):??, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2006:PS

[Ano06y] Anonymous. Pages 129–358 (September 2006). Fisheries Research, 80(2–3):??, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2006:PF

[Ano06z] Anonymous. Pages 131–264 (February 2006). Fisheries Research, 77(2):??, February 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2006:PJc

[Ano06-27] Anonymous. Pages 233–354 (July 2006). Fisheries Research, 79(3):??, July 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2006:PMa

[Ano06-28] Anonymous. Pages 265–358 (March 2006). Fisheries Research, 77(3):??, March 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2006:PPP

[Ano06-29] Anonymous. Potential, problems and progress. Fisheries Research, 80(1):??, August 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2006:PCC

[Ano06-30] Anonymous. Proceedings of the CEPHSTOCK Cephalopod Assessment Workshop. Fisheries Research, 78(1):??, April 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2006:PA

[Ano06-31] Anonymous. Publisher’s acknowledgement. Fisheries Research, 78(2–3):390–391, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http:

//www.sciencedirect.com/science/article/pii/S0165783606000610.

Anonymous:2007:CC

[Ano07a] Anonymous. 1st CoastFish Conference. Fisheries Research, 87(1):??, October 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2007:CFRa

[Ano07b] Anonymous. Contents of Fisheries Research, volume 83. Fisheries Research, 83(2–3):364–365, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004450.

Anonymous:2007:CFRb

[Ano07c] Anonymous. Contents of Fisheries Research, volume 84. Fisheries Research, 84(3):395–396, May 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000641.

Anonymous:2007:CFRc

[Ano07d] Anonymous. Contents of Fisheries Research, volume 86. Fisheries Research, 85(3):334–335, July 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000999.

Anonymous:2007:CFRd

[Ano07e] Anonymous. Contents of Fisheries Research, volume 86. Fisheries Research, 86(2–3):289–290, September 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001804.

Anonymous:2007:CFRe

[Ano07f] Anonymous. Contents of Fisheries Research, volume 87. Fisheries Research, 87(2–3):260–261, November 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002664.

Anonymous:2007:CFRf

[Ano07g] Anonymous. Contents of Fisheries Research, volume 88. Fisheries Research, 88(1–3):157–158, December 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002858.

Anonymous:2007:EABa

[Ano07h] Anonymous. Editorial Advisory Board. Fisheries Research, 83(1):iii, January 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606004176.

Anonymous:2007:EABb

[Ano07i] Anonymous. Editorial Advisory Board. Fisheries Research, 84(2):i, April 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783607000458.

Anonymous:2007:EABc

[Ano07j] Anonymous. Editorial Advisory Board. Fisheries Research, 84(3):i, May 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783607000628.

Anonymous:2007:EABd

[Ano07k] Anonymous. Editorial Advisory Board. Fisheries Research, 85(1–2):iii, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783607000781.

Anonymous:2007:EABe

[Ano07l] Anonymous. Editorial Advisory Board. Fisheries Research, 86(1):iii, August 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360700135X.

Anonymous:2007:EABf

[Ano07m] Anonymous. Editorial Advisory Board. Fisheries Research, 88(1–3):iii, December 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783607002834.

Anonymous:2007:PJa

[Ano07n] Anonymous. Pages 1–122 (January 2007). Fisheries Research, 83(1):??, January 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2007:PD

[Ano07o] Anonymous. Pages 1–158 (December 2007). Fisheries Research, 88(1–3):??, December 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2007:PJb

[Ano07p] Anonymous. Pages 1–238 (June 2007). Fisheries Research, 85(1–2):??, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2007:PAb

[Ano07q] Anonymous. Pages 1–84 (August 2007). Fisheries Research, 86(1):??, August 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2007:PF

[Ano07r] Anonymous. Pages 123–366 (February 2007). Fisheries Research, 83(2–3):??, February 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2007:PAa

[Ano07s] Anonymous. Pages 137–274 (April 2007). Fisheries Research, 84(2):??, April 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2007:PJc

[Ano07t] Anonymous. Pages 239–336 (July 2007). Fisheries Research, 85(3):??, July 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2007:PM

[Ano07u] Anonymous. Pages 275–396 (May 2007). Fisheries Research, 84(3):??, May 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2007:PS

[Ano07v] Anonymous. Pages 85–292 (September 2007). Fisheries Research, 86(2–3):??, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2007:PFI

[Ano07w] Anonymous. Proceedings of the First International Whale Shark Conference. Fisheries Research, 84(1):??, March 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2007:RSI

[Ano07x] Anonymous. Referees for this special issue. Fisheries Research, 87(1):iv, October 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0165783607002093.

Anonymous:2007:WTO

[Ano07y] Anonymous. The world trade organization and global fisheries sustainability. Fisheries Research, 88(1–3):1–4, December 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607002202.

Anonymous:2008:CFRa

[Ano08a] Anonymous. Contents of Fisheries Research, volume 61. Fisheries Research, 89(3):311–313, March 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000258.

Anonymous:2008:CFRb

[Ano08b] Anonymous. Contents of Fisheries Research, volume 90. Fisheries Research, 90(1–3):318–319, April 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000775.

Anonymous:2008:CFRc

[Ano08c] Anonymous. Contents of Fisheries Research, volume 91. Fisheries Research, 91(2–3):354–355, June 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360800129X.

Anonymous:2008:CFRe

[Ano08d] Anonymous. Contents of Fisheries Research, volume 94. Fisheries Research, 94(3):391–393, December 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003457.

Anonymous:2008:CFRd

[Ano08e] Anonymous. Contents of fisheries research. Fisheries Research, 93(3):360–362, September 28, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002397.

Anonymous:2008:EABa

[Ano08f] Anonymous. Editorial Advisory Board. Fisheries Research, 89(1):iii, January 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360700327X.

Anonymous:2008:EABb

[Ano08g] Anonymous. Editorial Advisory Board. Fisheries Research, 90(1–3):iii, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608000726.

Anonymous:2008:EABc

[Ano08h] Anonymous. Editorial Advisory Board. Fisheries Research, 91(1):iii, May 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608001045.

Anonymous:2008:EABd

[Ano08i] Anonymous. Editorial Advisory Board. Fisheries Research, 94(1):iii, October 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608002518.

Anonymous:2008:EBa

[Ano08j] Anonymous. Editorial Board. Fisheries Research, 92(1): iii, July 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608001501.

Anonymous:2008:EBb

[Ano08k] Anonymous. Editorial Board. Fisheries Research, 93(1–2): iii, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608002063.

Anonymous:2008:HIS

[Ano08l] Anonymous. Haddock 2007: International Symposium on Haddock Conservation, Harvesting and Management. Fisheries Research, 94(2):??, November 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2008:PJa

[Ano08m] Anonymous. Pages 1–100 (January 2008). Fisheries Research, 89(1):??, January 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2008:PJc

[Ano08n] Anonymous. Pages 1–114 (July 2008). Fisheries Research, 92(1):??, July 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2008:PMb

[Ano08o] Anonymous. Pages 1–118 (May 2008). Fisheries Research, 91(1):??, May 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2008:PO

[Ano08p] Anonymous. Pages 1–118 (October 2008). Fisheries Research, 94(1):??, October 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2008:PSa

[Ano08q] Anonymous. Pages 1–256 (1 September 2008). Fisheries Research, 93(1–2):??, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2008:PAa

[Ano08r] Anonymous. Pages 1–320 (April 2008). Fisheries Research, 90(1–3):??, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2008:PAb

[Ano08s] Anonymous. Pages 115–352 (August 2008). Fisheries Research, 92(2–3):??, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2008:PJb

[Ano08t] Anonymous. Pages 119–356 (June 2008). Fisheries Research, 91(2–3):??, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2008:PMa

[Ano08u] Anonymous. Pages 211–314 (March 2008). Fisheries Research, 89(3):??, March 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2008:PSb

[Ano08v] Anonymous. Pages 257–362 (28 September 2008). Fisheries Research, 93(3):??, September 28, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:AMN

[Ano09a] Anonymous. Annual meeting of the North Pacific Marine Science Organization. Fisheries Research, 100(1): ??, September 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:Ca

[Ano09b] Anonymous. Contents. Fisheries Research, 95(2–3):386– 387, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608003743.

Anonymous:2009:Cb

[Ano09c] Anonymous. Contents. Fisheries Research, 97(3):278–279, May 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609000848.

Anonymous:2009:CFRd

[Ano09d] Anonymous. Contents of Fisheries Research, volume 100. Fisheries Research, 100(3):281–282, November 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002380.

Anonymous:2009:CFRa

[Ano09e] Anonymous. Contents of Fisheries Research, volume 96. Fisheries Research, 96(2–3):324–326, March 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000265.

Anonymous:2009:CFRb

[Ano09f] Anonymous. Contents of Fisheries Research, volume 98. Fisheries Research, 98(1–3):113, June 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001222.

Anonymous:2009:CFRc

[Ano09g] Anonymous. Contents of Fisheries Research, volume 99. Fisheries Research, 99(3):248–249, August 15, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001702.

Anonymous:2009:EABa

[Ano09h] Anonymous. Editorial Advisory Board. Fisheries Research, 95(1):iii, January 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608003597.

Anonymous:2009:EABb

[Ano09i] Anonymous. Editorial Advisory Board. Fisheries Research, 95(2–3):i, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360800372X.

Anonymous:2009:EABc

[Ano09j] Anonymous. Editorial Advisory Board. Fisheries Research, 97(1–2):iii, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783609000538.

Anonymous:2009:EABd

[Ano09k] Anonymous. Editorial Advisory Board. Fisheries Research, 97(3):i, May 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783609000824.

Anonymous:2009:EABe

[Ano09l] Anonymous. Editorial Advisory Board. Fisheries Research, 98(1–3):iii, June 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783609001209.

Anonymous:2009:EABf

[Ano09m] Anonymous. Editorial Advisory Board. Fisheries Research, 99(1):iii, July 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783609001325.

Anonymous:2009:EABg

[Ano09n] Anonymous. Editorial Advisory Board. Fisheries Research, 99(2):i, August 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783609001532.

Anonymous:2009:EABh

[Ano09o] Anonymous. Editorial Advisory Board. Fisheries Research, 99(3):i, August 15, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360900188X.

Anonymous:2009:EABi

[Ano09p] Anonymous. Editorial Advisory Board. Fisheries Research, 100(2):i, October 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360900215X.

Anonymous:2009:EABj

[Ano09q] Anonymous. Editorial Advisory Board. Fisheries Research, 100(3):i, November 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783609002367.

Anonymous:2009:FC

[Ano09r] Anonymous. FSAMLR Conference. Fisheries Research, 96 (1):??, February 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:LR

[Ano09s] Anonymous. List of referees. Fisheries Research, 96(1):iv, February 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608002725.

Anonymous:2009:PJc

[Ano09t] Anonymous. Pages 1–114 (June 2009). Fisheries Research, 98(1–3):??, June 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:PJa

[Ano09u] Anonymous. Pages 1–138 (1 January 2009). Fisheries Research, 95(1):??, January 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:PAa

[Ano09v] Anonymous. Pages 1–154 (April 2009). Fisheries Research, 97(1–2):??, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:PJd

[Ano09w] Anonymous. Pages 1–62 (July 2009). Fisheries Research, 99(1):??, July 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:PMa

[Ano09x] Anonymous. Pages 119–328 (March 2009). Fisheries Research, 96(2–3):??, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:PAd

[Ano09y] Anonymous. Pages 137–250 (15 August 2009). Fisheries Research, 99(3):??, August 15, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:PJb

[Ano09z] Anonymous. Pages 139–388 (14 January 2009). Fisheries Research, 95(2–3):??, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:PMb

[Ano09-27] Anonymous. Pages 155–280 (May 2009). Fisheries Research, 97(3):??, May 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:PNb

[Ano09-28] Anonymous. Pages 185–282 (November 2009). Fisheries Research, 100(3):??, November 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:PAc

[Ano09-29] Anonymous. Pages 63–136 (1 August 2009). Fisheries Research, 99(2):??, August 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:PO

[Ano09-30] Anonymous. Pages 99–184 (October 2009). Fisheries Research, 100(2):??, October 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic).

Anonymous:2009:PNa

[Ano09-31] Anonymous. Publisher note. Fisheries Research, 95(1):v, January 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608003524.

Anonymous:2009:PAb

[Ano09-32] Anonymous. Publisher’s acknowledgement. Fisheries Research, 97(1–2):151–152, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http:

//www.sciencedirect.com/science/article/pii/S0165783609000368.

Albert:2001:MCS

[ANS^{+}01] O. T. Albert, E. M. Nilssen, A. Stene, A. C. Gundersen, and K. H. Nedreaas. Maturity classes and spawning behaviour of Greenland halibut (Reinhardtius hippoglossoides). Fisheries Research, 51(2–3):217–228, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002478.

Aburto-Oropeza:2007:LE

[AOLLP^{+}07] Octavio Aburto-Oropeza, Luis G. López-Lemus, Gustavo Paredes, Miriam Reza, Andrea Sáenz-Arroyo, and Enric Sala. Letter to the Editor. Fisheries Research, 85(1–2):233– 234, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606003730.

Al-Oufi:2004:ESR

[AOMK^{+}04] H. Al-Oufi, E. McLean, A. S. Kumar, M. Claereboudt, and M. Al-Habsi. The effects of solar radiation upon breaking strength and elongation of fishing nets. Fisheries Research, 66(1):115–119, January 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001036.

Ainsworth:2005:EIU

[AP05] C. H. Ainsworth and T. J. Pitcher. Estimating illegal, unreported and unregulated catch in British Columbia’s marine fisheries. Fisheries Research, 75(1–3):40–55, September 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605001402.

Adams:2007:SCD

[APC07] Charles F. Adams, Alexei I. Pinchuk, and Kenneth O. Coyle. Seasonal changes in the diet composition and prey selection of walleye pollock (Theragra chalcogramma) in the northern Gulf of Alaska. Fisheries Research, 84(3):378– 389, May 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606004279.

Abraham:2009:EFW

[APM^{+}09] Edward R. Abraham, Johanna P. Pierre, David A. J. Middleton, John Cleal, Nathan A. Walker, and Susan M. Waugh. Effectiveness of fish waste management strategies in reducing seabird attendance at a trawl vessel. Fisheries Research, 95(2–3):210–219, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002658.

Askey:2007:EGE

[APP^{+}07] Paul J. Askey, John R. Post, Eric A. Parkinson, Etienne Rivot, Andrew J. Paul, and Peter A. Biro. Estimation of gillnet efficiency and selectivity across multiple sampling units: a hierarchical Bayesian analysis using mark-recapture data. Fisheries Research, 83(2–3):162–174, February 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606003377.

Allen:2007:FIA

[APS07] Larry G. Allen, Daniel J. Pondella, and Michael A. Shane. Fisheries independent assessment of a returning fishery: Abundance of juvenile white seabass (Atractoscion nobilis) in the shallow nearshore waters of the Southern California Bight, 1995–2005. Fisheries Research, 88(1–3):24–32, December 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607001646.

Anderson:2001:LCO

[AR01] Cairistiona I. H. Anderson and Paul G. Rodhouse. Life cycles, oceanography and variability: ommastrephid squid in variable oceanographic environments. Fisheries Research, 54(1):133–143, December 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003782.

Abookire:2005:MPS

[AR05] Alisa A. Abookire and Craig S. Rose. Modifications to a plumb staff beam trawl for sampling uneven, complex habitats. Fisheries Research, 71(2):247–254, February 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001341.

Aranda:2009:DFM

[Ara09] Martin Aranda. Developments on fisheries management in Peru: the new individual vessel quota system for the anchoveta fishery. Fisheries Research, 96(2–3):308–312, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608003809.

Arculeo:2003:SGS

[Arc03] M. Arculeo. The stock genetic structure of two Sparidae species, Diplodus vulgaris and Lithognathus mormyrus, in the Mediterranean Sea. Fisheries Research, 63(3):339– 347, September 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783603001024.

Arteaga-Rios:2007:PLS

[ARCLBPdP07] Lilia Danae Arteaga-Ríos, Jorge Carrillo-Laguna, Jorge Belmar-Pérez, and Sergio A. Guzman del Proo. Post-larval settlement of California spiny lobster Panulirus interruptus in Bahía Tortugas, Baja California and its relationship to the commercial catch. Fisheries Research, 88(1–3):51–55, December 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607001683.

Acolas:2007:LES

[ARLB07] M. L. Acolas, J. M. Roussel, J. M. Lebel, and J. L. Baglinière. Laboratory experiment on survival, growth and tag retention following PIT injection into the body cavity of juvenile brown trout (Salmo trutta). Fisheries Research, 86(2–3):280–284, September 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001221.

Arguelles:2001:AGP

[ARVC01] J. Argüelles, P. G. Rodhouse, P. Villegas, and G. Castillo. Age, growth and population structure of the jumbo flying squid Dosidicus gigas in Peruvian waters. Fisheries Research, 54(1):51–61, December 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003800.

Andrade:2004:STR

[AS04] Humber Agrelli Andrade and João Antônio Teixeira Santos. Seasonal trends in the recruitment of skipjack tuna (Katsuwonus pelamis) to the fishing ground in the southwest Atlantic. Fisheries Research, 66(2–3):185–194, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001991.

Aguzzi:2004:SDN

[ASA04] Jacopo Aguzzi, Francesc Sardà, and Rosario Allué. Seasonal dynamics in Nephrops norvegicus (Decapoda: Nephropidae) catches off the Catalan coasts (Western Mediterranean). Fisheries Research, 69(2):293– 300, September 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604001298.

Amezcua:2006:AGM

[ASAGR06] Felipe Amezcua, Clarissa Soto-Avila, and Yanira Green-Ruiz. Age, growth, and mortality of the spotted rose snapper Lutjanus guttatus from the southeastern Gulf of California. Fisheries Research, 77(3):293–300, March 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003048.

Arkhipkin:2009:OCR

[ASD09] Alexander I. Arkhipkin, Pia C. Schuchert, and Leonid Danyushevsky. Otolith chemistry reveals fine population structure and close affinity to the Pacific and Atlantic oceanic spawning grounds in the migratory southern blue whiting (Micromesistius australis australis). Fisheries Research, 96(2–3):188–194, March 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003512.

Ambrose:2005:AFC

[ASIW05] Eyo E. Ambrose, B. B. Solarin, C. E. Isebor, and A. B. Williams. Assessment of fish by-catch species from coastal artisanal shrimp beam trawl fisheries in Nigeria. Fisheries Research, 71(1):125–132, January 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001559.

Aprahamian:2003:RSO

[ASM^{+}03] M. W. Aprahamian, K. Martin Smith, P. McGinnity, S. McKelvey, and J. Taylor. Restocking of salmonids — opportunities and limitations. Fisheries Research, 62(2):211– 227, May 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602001637.

Addis:2009:GAS

[ASM^{+}09] Piero Addis, Marco Secci, Andrea Manunza, Stefano Corrias, Alessio Niffoi, and Angelo Cau. A geostatistical approach for the stock assessment of the edible sea urchin, Paracentrotus lividus, in four coastal zones of Southern and West Sardinia (SW Italy, Mediterranean Sea). Fisheries Research, 100(3):215–221, November 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002069.

Abella:2005:DRV

[ASR05] Alvaro Abella, Fabrizio Serena, and Michela Ria. Distributional response to variations in abundance over spatial and temporal scales for juveniles of European hake (Merluccius merluccius) in the Western Mediterranean Sea. Fisheries Research, 71(3):295–310, March 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002127.

Albert:2002:ARL

[ATB^{+}02] O. T. Albert, E. Torstensen, B. Bertelsen, S. T. Jonsson, I. H. Pettersen, and J. C. Holst. Age-reading of lumpsucker (Cyclopterus lumpus) otoliths: dissection, interpretation and comparison with length frequencies. Fisheries Research, 55(1–3):239–252, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002818.

Afonso:2008:PSH

[ATM08] Pedro Afonso, Fernando Tempera, and Gui Menezes. Population structure and habitat preferences of red porgy (Pagrus pagrus) in the Azores, central north Atlantic. Fisheries Research, 93(3):338–345, September 28, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001872.

Ayvazian:2002:STH

[AWY02] Suzanne G. Ayvazian, Brent S. Wise, and Glen C. Young. Short-term hooking mortality of tailor (Pomatomus saltatrix) in Western Australia and the impact on yield per recruit. Fisheries Research, 58(2):241–248, November 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003873.

Baisre:2006:GAF

[BA06] Julio A. Baisre and Zenaida Arboleya. Going against the flow: Effects of river damming in Cuban fisheries. Fisheries Research, 81(2–3):283–292, November 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002426.

Binion:2009:DIE

[BACP09] Greg R. Binion, Micheal S. Allen, Matt J. Catalano, and William E. Pine. Direct and indirect estimates of black crappie size selectivity to a common sampling gear: Potential biases and limitations for assessment. Fisheries Research, 95(1):47–54, January 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002208.

Bariche:2006:SBC

[BAEF06] M. Bariche, N. Alwan, and M. El-Fadel. Structure and biological characteristics of purse seine landings off the Lebanese coast (eastern Mediterranean). Fisheries Research, 82(1–3):246–252, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002220.

Baremore:2009:DAM

[BAH09] Ivy Elizabeth Baremore, Kate Irene Andrews, and Loraine Frances Hale. Difficulties associated with modeling growth in the Atlantic angel shark (Squatina dumeril). Fisheries Research, 99(3):203–209, August 15, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001623.

Balston:2009:AIL

[Bal09] Jacqueline Balston. An analysis of the impacts of longterm climate variability on the commercial barramundi (Lates calcarifer) fishery of north-east Queensland, Australia. Fisheries Research, 99(2):83–89, August 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001076.

Baras:2000:BRF

[Bar00] Etienne Baras. Book review: Fish and Fisheries of Lakes and Reservoirs in southeast Asia and Africa: W. L. T. van Densen, J. Morris (Eds.); Westburry Academic & Scientific Publishing, Otley, UK, 1999, 432 pages, ISBN 0-1-84103-004-X (£35). Fisheries Research, 48(3):287–288, October 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360000223X.

Basu:1999:ARP

[Bas99] T. K. Basu. Application of Rao’s polynomial growth curve model to analyse the length growth data of a freshwater major carp of India, Catla catla (Ham) during the early developmental period. Fisheries Research, 42(3): 303–307, September 1999. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578369900051X.

See erratum [Bas00].

Basu:2000:EAR

[Bas00] T. K. Basu. Erratum to “Application of Rao’s polynomial growth curve model to analyse the length growth data of a freshwater major carp of India, Catla catla (Ham) during the early developmental period”: [Fish. Res. 42 (1999) 303–307]. Fisheries Research, 45(2):201, March 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001752. See [Bas99].

Baticados:2004:FCP

[Bat04] Didi B. Baticados. Fishing cooperatives’ participation in managing nearshore resources: the case in Capiz, central Philippines. Fisheries Research, 67(1):81–91, March 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002273.

Blanchard:2001:TVT

[BB01] Fabian Blanchard and Jean Boucher. Temporal variability of total biomass in harvested communities of demersal fishes. Fisheries Research, 49(3):283–293, January 2001.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002034.

Bacheler:2004:DHT

[BB04] Nathan M. Bacheler and Jeffrey A. Buckel. Does hook type influence the catch rate, size, and injury of grouper in a North Carolina commercial fishery? Fisheries Research, 69(3):303–311, October 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001407.

Brandao:2004:UGE

[BBJG04] A. Brandão, D. S. Butterworth, S. J. Johnston, and J. P. Glazer. Using a GLMM to estimate the somatic growth rate trend for male South African west coast rock lobster, Jasus lalandii. Fisheries Research, 70(2–3):339–349, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001766.

Burke:2008:SAO

[BBK08] Nóirín Burke, D. Brophy, and P. A. King. Shape analysis of otolith annuli in Atlantic herring (Clupea harengus); a new method for tracking fish populations. Fisheries Research, 91(2–3):133–143, June 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003220.

Bustos:2007:SSH

[BBL07] Claudia A. Bustos, Fernando Balbontín, and Mauricio F. Landaeta. Spawning of the southern hake Merluccius australis (Pisces: Merlucciidae) in Chilean fjords. Fisheries Research, 83(1):23–32, January 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003031.

Bell:2006:RSE

[BBLL06] Johann D. Bell, Devin M. Bartley, Kai Lorenzen, and Neil R. Loneragan. Restocking and stock enhancement of coastal fisheries: Potential, problems and progress. Fisheries Research, 80(1):1–8, August 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001147.

Botto:2006:IPS

[BBM^{+}06] Florencia Botto, Claudia Bremec, Angel Marecos, Laura Schejter, Mario Lasta, and Oscar Iribarne. Identifying predators of the SW Atlantic Patagonian scallop Zygochlamys patagonica using stable isotopes. Fisheries Research, 81(1):45–50, October 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002219.

Bartley:2006:RAF

[BBRS06] Devin M. Bartley, Melba G. Bondad-Reantaso, and Rohana P. Subasinghe. A risk analysis framework for aquatic animal health management in marine stock enhancement programmes. Fisheries Research, 80(1):28–36, August 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606001172.

Blanco:2007:ERF

[BBV07] Jacobo A. Blanco, Juan Carlos Narváez Barandica, and Efraín A. Viloria. ENSO and the rise and fall of a tilapia fishery in northern Colombia. Fisheries Research, 88(1–3):100–108, December 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001877.

Bannerman:2002:SAB

[BC02] P. O. Bannerman and I. G. Cowx. Stock assessment of the big-eye grunt (Brachydeuterus auritus, Val.) fishery in Ghanaian coastal waters. Fisheries Research, 59(1–2):197–207, December 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601004040.

Brattey:2004:EST

[BC04] John Brattey and Noel Cadigan. Estimation of short-term tagging mortality of adult Atlantic cod (Gadus morhua). Fisheries Research, 66(2–3):223–233, February 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002030.

Baeta:2009:CTL

[BCC09a] Filipa Baeta, Maria José Costa, and Henrique Cabral. Changes in the trophic level of Portuguese landings and fish market price variation in the last decades. Fisheries Research, 97(3):216–222, May 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000460.

Baeta:2009:TNG

[BCC09b] Filipa Baeta, Maria José Costa, and Henrique Cabral. Trammel nets’ ghost fishing off the Portuguese central coast. Fisheries Research, 98(1–3):33–39, June 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000769.

Brodziak:2008:GSR

[BCLM08] Jon Brodziak, Steve X. Cadrin, Christopher M. Legault, and Steve A. Murawski. Goals and strategies for rebuilding New England groundfish stocks. Fisheries Research, 94(3):355–366, December 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000970.

Boncoeur:2000:IRR

[BCLP00] Jean Boncoeur, Louisa Coglan, Bertrand Le Gallic, and Sean Pascoe. On the (ir)relevance of rates of return measures of economic performance to small boats. Fisheries Research, 49(2):105–115, December 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002095.

Butterworth:2008:RCB

[BCM08a] Kevin Butterworth, K. Fiona Cubitt, and R. Scott McKinley. Response to “Comment on Butterworth, Cubitt and McKinley (2008)” by Morton and Routledge [Fish. Res. (2008)]. Fisheries Research, 92(2–3):347–349, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608000933.  See

[BCM08b, MR08b].

Butterworth:2008:PDI

[BCM08b] Kevin G. Butterworth, K. Fiona Cubitt, and R. Scott McKinley. The prevalence, density and impact of Lepeophtheirus salmonis (Krøyer) infestation on juvenile pink salmon (Oncorhynchus gorbuscha) from the central coast of British Columbia, Canada. Fisheries Research, 91(1):35– 41, May 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607003104.  See

comment [MR08b] and response [BCM08a].

Beverly:2009:EES

[BCMM09] Steve Beverly, Daniel Curran, Michael Musyl, and Brett Molony. Effects of eliminating shallow hooks from tuna longline sets on target and non-target species in the Hawaiibased pelagic tuna fishery. Fisheries Research, 96(2–3): 281–288, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608003950.

Bartolino:2008:MRD

[BCSA08] V. Bartolino, F. Colloca, P. Sartor, and G. Ardizzone. Modelling recruitment dynamics of hake, Merluccius merluccius, in the central Mediterranean in relation to key environmental variables. Fisheries Research, 92(2–3):277–288, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608000477.

Barthel:2003:ELN

[BCSP03] B. L. Barthel, S. J. Cooke, C. D. Suski, and D. P. Philipp. Effects of landing net mesh type on injury and mortality in a freshwater recreational fishery.

Fisheries Research, 63(2):275–282, August 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000596.

Burgos:2004:LTP

[BD04] Rogerio Burgos and Omar Defeo. Long-term population structure, mortality and modeling of a tropical multi-fleet fishery: the red grouper Epinephelus morio of the Campeche Bank, Gulf of Mexico. Fisheries Research, 66(2–3):325–335, February 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001929.

Bach:2003:ATV

[BDB^{+}03] Pascal Bach, Laurent Dagorn, Arnaud Bertrand, Erwan Josse, and Christophe Misselis. Acoustic telemetry versus monitored longline fishing for studying the vertical distribution of pelagic fish: bigeye tuna (Thunnus obesus) in French Polynesia. Fisheries Research, 60(2–3):281–292, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602001807.

Brinson:2009:SPW

[BDBD09] Ayeisha A. Brinson, David J. Die, Paul O. Bannerman, and Youssouph Diatta. Socioeconomic performance of West African fleets that target Atlantic billfish. Fisheries Research, 99(1):55–62, July 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001040.

Bence:2008:RAE

[BDIP08] James R. Bence, Martin W. Dorn, Brian J. Irwin, and André E. Punt. Recent advances in the evaluation and implementation of harvest policies. Fisheries Research, 94(3):207–209, December 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003147.

Bellchambers:2005:SDG

[BdL05] L. M. Bellchambers and S. de Lestang. Selectivity of different gear types for sampling the blue swimmer crab, Portunus pelagicus L. Fisheries Research, 73(1–2):21–27, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360500024X.

Bell:2008:TCC

[BEA^{+}08] M. C. Bell, J. M. Elson, J. T. Addison, A. S. Revill, and D. Bevan. Trawl catch composition in relation to Norway lobster (Nephrops norvegicus L.) abundance on the Farn Deeps grounds, NE England. Fisheries Research, 90(1–3):128–137, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002561.

Bell:2003:MRA

[BEBA03] M. C. Bell, D. R. Eaton, R. C. A. Bannister, and J. T. Addison. A mark-recapture approach to estimating population density from continuous trapping data: application to edible crabs, Cancer pagurus, on the east coast of England. Fisheries Research, 65(1–3):361–378, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002558.

Bermejo:2007:FACb

[Ber07] Sergio Bermejo. Fish age classification based on length, weight, sex and otolith morphological features. Fisheries Research, 84(2):270–274, April 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000070.

Bethke:2004:SGA

[Bet04] Eckhard Bethke. A simple general approach to codend selectivity of trawls and its application to the data of Fiorentino et al. (1998) for hake (Merluccius merluccius). Fisheries Research, 70(1):113–119, November 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001584.

Bahri:2000:SSC

[BF00a] Tarûb Bahri and Pierre Fréon. Spatial structure of coastal pelagic schools descriptors in the Mediterranean Sea. Fisheries Research, 48(2):157–166, September 1, 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001764.

Butler:2000:SEF

[BF00b] John L. Butler and Arild Folkvord. Section 2: Estimation of fish age and growth. Fisheries Research, 46(1–3):51– 52, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600001326.

Brierley:2003:IAA

[BFB^{+}03] Andrew S. Brierley, Paul G. Fernandes, Mark A. Brandon, Frederick Armstrong, Nicholas W. Millard, Steven D. McPhail, Peter Stevenson, Miles Pebody, James Perrett, Mark Squires, Douglas G. Bone, and Gwyn Griffiths. An investigation of avoidance by Antarctic krill of RRS James Clark Ross using the Autosub-2 autonomous underwater vehicle. Fisheries Research, 60(2–3):569–576, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602001443.

Biron:2008:MAM

[BFM08] Michel Biron, Christine Ferron, and Mikio Moriyasu. Movement of adult male snow crab, Chionoecetes opilio, in the southern Gulf of St. Lawrence and eastern Nova Scotia, Canada. Fisheries Research, 91(2–3):260–270, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003487.

Bogstad:2001:PCG

[BG01] B. Bogstad and H. Gjøsæter. Predation by cod (Gadus morhua) on capelin (Mallotus villosus) in the Barents Sea: implications for capelin stock assessment. Fisheries Research, 53(2):197–209, October 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002885.

Brouwer:2004:AGA

[BG04] Stephen L. Brouwer and Marc H. Griffiths. Age and growth of Argyrozona argyrozona (Pisces: Sparidae) in a marine protected area: an evaluation of methods based on whole otoliths, sectioned otoliths and markrecapture. Fisheries Research, 67(1):1–12, March 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002662.

Brouwer:2005:ISD

[BG05] Stephen L. Brouwer and Marc H. Griffiths. Influence of sample design on estimates of growth and mortality in Argyrozona argyrozona (Pisces: Sparidae). Fisheries Research, 74(1–3):44–54, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001244.

Brouwer:2006:EAL

[BGB06] Stephen L. Brouwer, Johan C. Groeneveld, and Bradley Blows. The effects of appendage loss on growth of South African west coast rock lobster Jasus lalandii. Fisheries Research, 78(2–3):236–242, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000282.

Brookfield:2005:CFD

[BGH05] Katherine Brookfield, Tim Gray, and Jenny Hatchard. The concept of fisheries-dependent communities: a comparative analysis of four UK case studies: Shetland, Peterhead, North Shields and Lowestoft. Fisheries Research, 72(1):55– 69, April 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604002619.

Bach:2009:EGD

[BGM^{+}09] Pascal Bach, Daniel Gaertner, Christophe Menkes, Evgeny Romanov, and Paulo Travassos. Effects of the gear deployment strategy and current shear on pelagic longline shoaling. Fisheries Research, 95(1):55–64, January 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360800221X.

Basilone:2004:LHC

[BGP^{+}04] G. Basilone, C. Guisande, B. Patti, S. Mazzola, A. Cuttitta, A. Bonanno, and A. Kallianiotis. Linking habitat conditions and growth in the European anchovy (Engraulis encrasicolus). Fisheries Research, 68(1–3):9–19, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000669.

Braccini:2006:TPL

[BGW06] J. Matías Braccini, Bronwyn M. Gillanders, and Terence I. Walker. Total and partial length–length, mass– mass and mass–length relationships for the piked spurdog (Squalus megalops) in south-eastern Australia. Fisheries Research, 78(2–3):385–389, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600021X.

Bailey:2001:SVG

[BH01a] M. C. Bailey and M. R. Heath. Spatial variability in the growth rate of blue whiting (Micromesistius poutassou) larvae at the shelf edge west of the UK. Fisheries Research, 50(1–2):73–87, February 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002435.

Broadhurst:2001:ITO

[BH01b] Matt K. Broadhurst and Fabio H. V. Hazin. Influences of type and orientation of bait on catches of swordfish (Xiphias gladius) and other species in an artisanal sub-surface longline fishery off northeastern Brazil.

Fisheries Research, 53(2):169–179, October 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002976.

Bergmann:2004:UKF

[BHB^{+}04] M. Bergmann, H. Hinz, R. E. Blyth, M. J. Kaiser, S. I. Rogers, and M. Armstrong. Using knowledge from fishers and fisheries scientists to identify possible groundfish ‘essential fish habitats’. Fisheries Research, 66(2–3):373–379, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603002674.

Bowlby:2008:SSS

[BHH08] Heather D. Bowlby, J. Mark Hanson, and Jeffrey A. Hutchings. Stock structure and seasonal distribution patterns of American lobster, Homarus americanus, inferred through movement analyses. Fisheries Research, 90(1–3):279–288, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607003050.

Bergstad:2002:GSA

[BHJ02] O. A. Bergstad, Å. S. Høines, and T. Jørgensen. Growth of sandeel, Ammodytes marinus, in the northern North Sea and Norwegian coastal waters. Fisheries Research, 56(1):9– 23, April 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601003174.

Brewer:2006:ITE

[BHM^{+}06] David Brewer, Don Heales, David Milton, Quinton Dell, Gary Fry, Bill Venables, and Peter Jones. The impact of turtle excluder devices and bycatch reduction devices on diverse tropical marine communities in Australia’s northern prawn trawl fishery. Fisheries Research, 81(2–3):176–188, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606002797.

Bower:2005:RFS

[BI05] John R. Bower and Taro Ichii. The red flying squid (Ommastrephes bartramii): a review of recent research and the fishery in Japan. Fisheries Research, 76(1):39–55, October 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605001475.

Bjorke:2001:PSG

[Bjø01] Herman Bjørke. Predators of the squid Gonatus fabricii (Lichtenstein) in the Norwegian Sea. Fisheries Research, 52(1–2):113–120, June 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783601002351.

Bilecenoglu:2002:GMS

[BK02] Murat Bilecenoglu and Murat Kaya. Growth of marbled spinefoot Siganus rivulatus Forsskål, 1775 (Teleostei: Siganidae) introduced to Antalya Bay, eastern Mediterranean Sea (Turkey). Fisheries Research, 54(2):279–285, January 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600002964.

Bechtol:2009:RHA

[BK09] William R. Bechtol and Gordon H. Kruse. Reconstruction of historical abundance and recruitment of red king crab during 1960–2004 around Kodiak, Alaska. Fisheries Research, 100(1):86–98, September 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002701.

Broadhurst:2002:UCS

[BKD^{+}02] Matt K. Broadhurst, Mervi I. Kangas, Cristiana Damiano, Scott A. Bickford, and Steven J. Kennelly. Using composite square-mesh panels and the Nordmøre-grid to reduce bycatch in the Shark Bay prawn-trawl fishery, Western Australia. Fisheries Research, 58(3):349–365, November 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601003903.

Blaber:2006:AMF

[Bla06] Stephen J. M. Blaber. Assessment, management and future directions for coastal fisheries in Asian countries: Key papers from a workshop held in Penang, Malaysia in March 2001. Fisheries Research, 78(2–3):107–108, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000555.

Brennan:2007:UCW

[BLB07] Nathan P. Brennan, Kenneth M. Leber, and Brett R. Blackburn. Use of coded-wire and visible implant elastomer tags for marine stock enhancement with juvenile red snapper Lutjanus campechanus. Fisheries Research, 83(1):90–97, January 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606003122.

Bellquist:2008:FSM

[BLC08] Lyall F. Bellquist, Christopher G. Lowe, and Jennifer E. Caselle. Fine-scale movement patterns, site fidelity, and habitat selection of ocean whitefish (Caulolatilus princeps). Fisheries Research, 91(2–3):325–335, June 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003621.

Bremec:2004:SPS

[BLH04] C. S. Bremec, M. L. Lasta, and D. Hernández. Survival of Patagonian scallop (Zygochlamys patagonica, King and Broderip, 1832) after the size selection process on commercial fishing vessels. Fisheries Research, 66(1):49–52, January 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603001796.

Bostrom:2009:CIT

[BLKR09] Maria Kristina Boström, Sven-Gunnar Lunneryd, Lars Karlsson, and Bjarne Ragnarsson. Cormorant impact on trout (Salmo trutta) and salmon (Salmo salar) migrating from the river Dalälven emerging in the Baltic Sea. Fisheries Research, 98(1–3):16–21, June 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000745.

Bradbury:2008:DSD

[BLR^{+}08] I. R. Bradbury, B. J. Laurel, D. Robichaud, G. A. Rose, P. V. R. Snelgrove, R. S. Gregory, D. Cote, and M. J. S. Windle. Discrete spatial dynamics in a marine broadcast spawner: Re-evaluating scales of connectivity and habitat associations in Atlantic cod (Gadus morhua L.) in coastal Newfoundland. Fisheries Research, 91(2–3):299–309, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607003530.

Belcari:2006:ADG

[BLV06] P. Belcari, A. Ligas, and C. Viva. Age determination and growth of juveniles of the European hake, Merluccius merluccius (L., 1758), in the northern Tyrrhenian Sea (NW Mediterranean). Fisheries Research, 78(2–3):211–217, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606000336.

Beamish:2000:RIA

[BM00] R. J. Beamish and G. A. McFarlane. Reevaluation of the interpretation of annuli from otoliths of a long-lived fish, Anoplopoma fimbria. Fisheries Research, 46(1–3):105–111, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001375.

Begg:2003:SPR

[BM03a] Gavin A. Begg and Gudrun Marteinsdottir. Spatial partitioning of relative fishing mortality and spawning stock biomass of Icelandic cod. Fisheries Research, 59(3):343–362, January 3, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000255.

Brooks:2003:IASa

[BM03b] Kenneth M. Brooks and Conrad V. W. Mahnken. Interactions of Atlantic salmon in the Pacific Northwest environment: II. Organic wastes. Fisheries Research, 62(3):255– 293, June 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360300064X.

Brooks:2003:IASb

[BM03c] Kenneth M. Brooks and Conrad V. W. Mahnken. Interactions of Atlantic salmon in the Pacific Northwest environment: III. Accumulation of zinc and copper. Fisheries Research, 62(3):295–305, June 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000651.

Bower:2005:DST

[BM05] John R. Bower and Kazutaka Miyahara. The diamond squid (Thysanoteuthis rhombus): a review of the fishery and recent research in Japan. Fisheries Research, 73(1–2): 1–11, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605000408.

Brazner:2008:LTC

[BM08] John C. Brazner and James McMillan. Loggerhead turtle (Caretta caretta) bycatch in Canadian pelagic longline fisheries: Relative importance in the western North Atlantic and opportunities for mitigation. Fisheries Research, 91(2–3):310–324, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360700361X.

Booth:2009:SRB

[BM09] John D. Booth and Andy McKenzie. Strong relationships between levels of puerulus settlement and recruited stock abundance in the red rock lobster (Jasus edwardsii) in New Zealand. Fisheries Research, 95(2–3):161–168, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608002567.

Busby:2005:HAD

[BMB05] Morgan S. Busby, Kathryn L. Mier, and Richard D. Brodeur. Habitat associations of demersal fishes and crabs in the Pribilof Islands region of the Bering Sea. Fisheries Research, 75(1–3):15–28, September 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001396.  See corrigendum

[BMB06].

Busby:2006:CHA

[BMB06] Morgan S. Busby, Kathryn L. Mier, and Richard D. Brodeur. Corrigendum to “Habitat associations of demersal fishes and crabs in the Pribilof Islands region of the Bering Sea” [Fish. Res. 75 (2005) 15–28]. Fisheries Research, 77(1):122–123, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002766. See [BMB05].

Bergenius:2005:UOC

[BMBM05] Mikaela A. J. Bergenius, Bruce D. Mapstone, Gavin A. Begg, and Cameron D. Murchie. The use of otolith chemistry to determine stock structure of three epinepheline serranid coral reef fishes on the Great Barrier Reef, Australia. Fisheries Research, 72(2–3):253–270, May 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002504.

Bermejo:2007:FACa

[BMC07] Sergio Bermejo, Brais Monegal, and Joan Cabestany. Fish age categorization from otolith images using multi-class support vector machines. Fisheries Research, 84(2):247– 253, April 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606003857.

Bevacqua:2010:BGM

[BMF^{+}10] Daniele Bevacqua, Paco Melià, Maria C. Follesa, Giulio A. De Leo, Marino Gatto, and Angelo Cau. Body growth and mortality of the spiny lobster Palinurus elephas within and outside a small marine protected area. Fisheries Research, 106(3):543–549, December 2010. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783610002560. See erratum

[KPL^{+}07].

Broadhurst:2004:SCD

[BMK^{+}04] Matt K. Broadhurst, Russell B. Millar, Steven J. Kennelly, William G. Macbeth, Damian J. Young, and Charles A. Gray. Selectivity of conventional diamond- and novel square-mesh codends in an Australian estuarine penaeidtrawl fishery. Fisheries Research, 67(2):183–194, April 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603002984.

Bigelow:2006:PLG

[BMPK06] Keith Bigelow, Michael K. Musyl, Francois Poisson, and Pierre Kleiber. Pelagic longline gear depth and shoaling. Fisheries Research, 77(2):173–183, February 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002973.

Butcher:2005:CRE

[BMSJ05] Adam Butcher, David Mayer, Darren Smallwood, and Mathew Johnston. A comparison of the relative efficiency of ring, fyke, fence nets and beam trawling for estimating key estuarine fishery populations. Fisheries Research, 73(3):311–321, July 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000342.

Boswell:2007:LTP

[BMW07] Kevin M. Boswell, Mark W. Miller, and Charles A. Wilson. A lightweight transducer platform for use in stationary shallow water horizontal-aspect acoustic surveys. Fisheries Research, 85(3):291–294, July 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000719.

Bullough:2007:YLT

[BNL^{+}07] L. W. Bullough, I. R. Napier, C. H. Laurenson, D. Riley, R. J. Fryer, R. S. T. Ferro, and R. J. Kynoch. A yearlong trial of a square mesh panel in a commercial demersal trawl. Fisheries Research, 83(1):105–112, January 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003353.

Bugoni:2008:PBS

[BNL^{+}08] Leandro Bugoni, Tatiana S. Neves, Nilamon O. Leite, Demétrio Carvalho, Gilberto Sales, Robert W. Furness, Carlos E. Stein, Fabiano V. Peppes, Bruno B. Giffoni, and Danielle S. Monteiro. Potential bycatch of seabirds and turtles in hook-and-line fisheries of the Itaipava Fleet, Brazil. Fisheries Research, 90(1–3):217–224, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002779.

Bhatt:2000:PSH

[BNS00] J. P. Bhatt, P. Nautiyal, and H. R. Singh. Population structure of Himalayan mahseer, a large cyprinid fish in the regulated foothill section of the river Ganga. Fisheries Research, 44(3):267–271, January 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000831.

Beamish:2007:RDN

[BNS07] R. J. Beamish, C. M. Neville, and R. M. Sweeting. Response to Dr. Neil Frazer’s comment on “Sea lice on adult Pacific salmon in the coastal waters of British Columbia, Canada” by R. J. Beamish et al. (2005). Fisheries Research, 85(3):332–333, July 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000161.

Beamish:2005:SLA

[BNSA05] R. J. Beamish, C. M. Neville, R. M. Sweeting, and N. Ambers. Sea lice on adult Pacific salmon in the coastal waters of Central British Columbia, Canada. Fisheries Research, 76(2):198–208, November 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001785.

Booth:2004:DCS

[Boo04] Anthony J. Booth. Determination of cichlid-specific biological reference points. Fisheries Research, 67(3):307–316, May 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603003035.

Berg:2001:VRG

[BP01a] Erik Berg and Torstein Pedersen. Variability in recruitment, growth and sexual maturity of coastal cod (Gadus morhua L.) in a fjord system in northern Norway. Fisheries Research, 52(3):179–189, July 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002563.

Bundy:2001:SHS

[BP01b] Alida Bundy and Daniel Pauly. Selective harvesting by small-scale fisheries: ecosystem analysis of San Miguel Bay, Philippines. Fisheries Research, 53(3):263–281, November 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600002952.

Booth:2006:EGN

[BP06] Anthony J. Booth and Warren M. Potts. Estimating gill-net selectivity for Labeo umbratus (Pisces: Cyprinidae), and an evaluation of using fyke-nets as a non-destructive sampling gear in small reservoirs. Fisheries Research, 79(1–2):202– 209, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360600110X.

Bhathal:2008:FMF

[BP08] Brajgeet Bhathal and Daniel Pauly. ‘Fishing down marine food webs’ and spatial expansion of coastal fisheries in India, 1950–2000. Fisheries Research, 91(1):26–34, May 2008.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003098.

Baeta:2005:FTE

[BPCR^{+}05] Filipa Baeta, Ana Pinheiro, Madalena Corte-Real, José Lino Costa, Pedro Raposo de Almeida, Henrique Cabral, and Maria José Costa. Are the fisheries in the Tagus estuary sustainable? Fisheries Research, 76(2):243–251, November 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605001992.

Byrne:2004:TEI

[BPD^{+}04] C. J. Byrne, R. Poole, M. Dillane, G. Rogan, and K. F. Whelan. Temporal and environmental influences on the variation in sea trout (Salmo trutta L.) smolt migration in the Burrishoole system in the west of Ireland from 1971 to 2000. Fisheries Research, 66(1):85–94, January 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001462.

Burridge:2006:CDC

[BPH^{+}06] C. Y. Burridge, C. R. Pitcher, B. J. Hill, T. J. Wassenberg, and I. R. Poiner. A comparison of demersal communities in an area closed to trawling with those in adjacent areas open to trawling: a study in the Great Barrier Reef Marine Park, Australia. Fisheries Research, 79(1–2):64–74, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606000208.

Bellido:2000:UFA

[BPRM00] J. M. Bellido, G. J. Pierce, J. L. Romero, and M. Millán. Use of frequency analysis methods to estimate growth of anchovy (Engraulis encrasicolus L. 1758) in the Gulf of Cádiz (SW Spain). Fisheries Research, 48(2):107–115, September 1, 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001831.

Bellido:2001:MIA

[BPW01] J. M. Bellido, G. J. Pierce, and J. Wang. Modelling intra-annual variation in abundance of squid Loligo forbesi in Scottish waters using generalised additive models. Fisheries Research, 52(1–2):23–39, June 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002284.

Burridge:2003:MRD

[BPW^{+}03] C. Y. Burridge, C. R. Pitcher, T. J. Wassenberg, I. R. Poiner, and B. J. Hill. Measurement of the rate of depletion of benthic fauna by prawn (shrimp) otter trawls: an experiment in the Great Barrier Reef, Australia. Fisheries Research, 60(2–3):237–253, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001790.

Battaile:2004:CPU

[BQ04] Brian C. Battaile and Terrance J. Quinn. Catch per unit effort standardization of the eastern Bering Sea walleye pollock (Theragra chalcogramma) fleet. Fisheries Research, 70(2–3):161–177, December 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360400164X.

Booth:2006:MLB

[BQ06] Anthony J. Booth and Terrance J. Quinn. Maximum likelihood and Bayesian approaches to stock assessment when data are questionable. Fisheries Research, 80(2–3):169–181, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001743.

Bazzino:2005:EAS

[BQN05] Gastón Bazzino, Renato A. Quiñones, and Walter Norbis. Environmental associations of shortfin squid Illex argentinus (Cephalopoda: Ommastrephidae) in the Northern Patagonian Shelf. Fisheries Research, 76(3):401–416, December 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605002109.

Burridge:2000:BSB

[BR00] C. Y. Burridge and J. B. Robins. Benefits of statistical blocking techniques in the design of gear evaluation trials: introducing the Latin Square design. Fisheries Research, 47(1):69–79, June 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001253.

Bez:2001:TGC

[BR01]  N. Bez and J. Rivoirard.  Transitive geostatistics

to characterise spatial aggregations with diffuse limits: an application on mackerel ichtyoplankton. Fisheries Research, 50(1–2):41–58, February 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002411.

Beckwith:2005:LCH

[BR05] George H. Beckwith and Peter S. Rand. Large circle hooks and short leaders with fixed weights reduce incidence of deep hooking in angled adult red drum. Fisheries Research, 71(1):115–120, January 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001973.

Bradshaw:2004:CSM

[Bra04] Matt Bradshaw. A combination of state and market through ITQs in the Tasmanian commercial rock lobster fishery: the tail wagging the dog? Fisheries Research, 67(2):99–109, April 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000189.

Bello:2005:ISA

[BRL^{+}05] P. Javier Bello, L. Veronica Rios, C. M. A. Liceaga, M. Carlos Zetina, C. Kenneth Cervera, B. Patricia Arceo, and N. Hector Hernandez. Incorporating spatial analysis of habitat into spiny lobster (Panulirus argus) stock assessment at Alacranes reef, Yucatan, México. Fisheries Research, 73(1–2):37–47, June 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000226.

Borges:2005:DDF

[BRO05a] L. Borges, E. Rogan, and R. Officer. Discarding by the demersal fishery in the waters around Ireland. Fisheries Research, 76(1):1–13, October 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001384.

Bromaghin:2005:VNS

[Bro05b] Jeffrey F. Bromaghin. A versatile net selectivity model, with application to Pacific salmon and freshwater species of the Yukon River, Alaska. Fisheries Research, 74(1–3): 157–168, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360500086X.

Butler:2009:EES

[BRRL09] J. R. A. Butler, A. Radford, G. Riddington, and R. Laughton. Evaluating an ecosystem service provided by Atlantic salmon, sea trout and other fish species in the River Spey, Scotland: the economic impact of recreational rod fisheries. Fisheries Research, 96(2–3):259–266, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608003913.

Brynjarsdottir:2004:ACC

[BS04] Jenný Brynjarsdóttir and Gunnar Stefánsson. Analysis of cod catch data from Icelandic groundfish surveys using generalized linear models. Fisheries Research, 70(2–3):195–208, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001663.

Bettoli:2006:BRI

[BS06] Phillip W. Bettoli and George D. Scholten. Bycatch rates and initial mortality of paddlefish in a commercial gillnet fishery. Fisheries Research, 77(3):343–347, March 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002924.

Babaran:2009:CFI

[BSAM09] Ricardo P. Babaran, Cornelio M. Selorio, Kazuhiko Anraku, and Tatsuro Matsuoka. Comparison of the food intake and prey composition of payao-associated and free-swimming bigtooth pomfret Brama orcini. Fisheries Research, 95(1):132–137, January 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002233.

Beutel:2008:BRN

[BSC^{+}08] David Beutel, Laura Skrobe, Kathleen Castro, Philip Ruhle, Philip Ruhle, James O’Grady, and Jonathan Knight. Bycatch reduction in the Northeast USA directed haddock bottom trawl fishery. Fisheries Research, 94(2): 190–198, November 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608002579.

Broughton:2002:MVS

[BSG02] Richard E. Broughton, Leah B. Stewart, and John R. Gold. Microsatellite variation suggests substantial gene flow between king mackerel (Scomberomorus cavalla) in the western Atlantic Ocean and Gulf of Mexico. Fisheries Research, 54(3):305–316, February 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002752.

Bager:2007:DFB

[BSM07] M. Bager, M. Karnøe Søndergaard, and B. R. MacKenzie. The development of fisheries at Bornholm, Denmark (Baltic Sea) during 1880s–1914. Fisheries Research, 87(2–3):146–154, November 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001919.

Bizzarro:2009:TVA

[BSMF^{+}09] Joseph J. Bizzarro, Wade D. Smith, J. Fernando Márquez-Farías, John Tyminski, and Robert E. Hueter. Temporal variation in the artisanal elasmobranch fishery of Sonora, Mexico. Fisheries Research, 97(1–2):103–117, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000332.

Bizzarro:2007:AFR

[BSMFH07] Joseph J. Bizzarro, Wade D. Smith, J. Fernando Márquez-Farías, and Robert E. Hueter. Artisanal fisheries and reproductive biology of the golden cownose ray, Rhinoptera steindachneri Evermann and Jenkins, 1891, in the northern Mexican Pacific. Fisheries Research, 84(2):137–146, April 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606003444.

Bozzano:2005:VDF

[BSR05] Anna Bozzano, Francesc Sardà, and José Ríos. Vertical distribution and feeding patterns of the juvenile European hake, Merluccius merluccius in the NW Mediterranean. Fisheries Research, 73(1–2):29–36, June 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000238.

Bahamon:2006:ITS

[BSS06] Nixon Bahamon, Francesc Sardà, and Petri Suuronen. Improvement of trawl selectivity in the NW Mediterranean demersal fishery by using a 40 mm square mesh codend. Fisheries Research, 81(1):15–25, October 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002256.

Brock:2006:ETC

[BSWL06a] Daniel John Brock, Thor Mayo Saunders, Timothy Mark Ward, and Adrian Joseph Linnane. Effectiveness of a two-chambered trap in reducing within-trap predation by octopus on southern spiny rock lobster. Fisheries Research, 77(3):348–355, March 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002857.

Brock:2006:TCT

[BSWL06b] Daniel John Brock, Thor Mayo Saunders, Timothy Mark Ward, and Adrian Joseph Linnane. A two-chambered trap reduces within-trap predation by octopus on rock lobsters in aquarium trials. Fisheries Research, 80(2–3): 129–135, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606001792.

Beaulaton:2008:FAL

[BTC08a] L. Beaulaton, C. Taverny, and G. Castelnaud. Fishing, abundance and life history traits of the anadromous sea lamprey (Petromyzon marinus) in Europe. Fisheries Research, 92(1):90–101, July 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000179.

Brodziak:2008:NRG

[BTC08b] Jon Brodziak, Michele L. Traver, and Laurel A. Col. The nascent recovery of the Georges Bank haddock stock. Fisheries Research, 94(2):123–132, November 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000969.

Batista:2009:CTS

[BTC09] Marisa I. Batista, Célia M. Teixeira, and Henrique N. Cabral. Catches of target species and bycatches of an artisanal fishery: the case study of a trammel net fishery in the Portuguese coast. Fisheries Research, 100(2):167–177, October 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609002033.

Burton:2002:AGM

[Bur02] Michael L. Burton. Age, growth and mortality of mutton snapper, Lutjanus analis, from the east coast of Florida, with a brief discussion of management implications. Fisheries Research, 59(1–2):31–41, December 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000073.

Butler:2003:IEP

[But03] Mark J. Butler. Incorporating ecological process and environmental change into spiny lobster population models using a spatially-explicit, individual-based approach. Fisheries Research, 65(1–3):63–79, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002352.

Butterworth:2008:CSP

[But08] D. S. Butterworth. A commentary on: salvaged pearls: lessons learned from a floundering attempt to develop a management procedure for southern bluefin tuna. Fisheries Research, 94(3):351–354, December 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003196.

Baumann:2008:ISS

[BVH^{+}08] Hannes Baumann, Rudi Voss, Hans-Harald Hinrichsen, Volker Mohrholz, Jörn O. Schmidt, and Axel Temming. Investigating the selective survival of summerover spring-born sprat, Sprattus sprattus, in the Baltic Sea. Fisheries Research, 91(1):1–14, May 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003074.

Bishop:2004:ACC

[BVW04] Janet Bishop, W. N. Venables, and You-Gan Wang. Analysing commercial catch and effort data from a penaeid trawl fishery: a comparison of linear models, mixed models, and generalised estimating equations approaches. Fisheries Research, 70(2–3):179–193, December 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001651.

Blount:2002:IIS

[BW02] Craig Blount and Duncan Worthington. Identifying individuals of the sea urchin Centrostephanus rodgersii with high-quality roe in New South Wales, Australia. Fisheries Research, 58(3):341–348, November 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100399X.

Booth:2008:RBA

[BW08] Anthony J. Booth and Olaf L. F. Weyl. Retention of T-bar anchor and dart tags by a wild population of African sharptooth catfish, Clarias gariepinus. Fisheries Research, 92(2–3):333–339, August 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000556.

Broadhurst:2007:ISM

[BWM07] Matt K. Broadhurst, Michael E. L. Wooden, and Russell B. Millar. Isolating selection mechanisms in beach seines. Fisheries Research, 88(1–3):56–69, December 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001701.

Bergmann:2002:DCN

[BWMA02] M. Bergmann, S. K. Wieczorek, P. G. Moore, and R. J. A. Atkinson. Discard composition of the Nephrops fishery in the Clyde Sea area, Scotland. Fisheries Research, 57 (2):169–183, August 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783601003459.

Bishop:2008:DNC

[BYA^{+}08] J. M. Bishop, Y. Ye, A. H. Alsaffar, H. M. Al-Foudari, and S. Al-Jazzaf. Diurnal and nocturnal catchability of Kuwait’s commercial shrimps. Fisheries Research, 94(1):58–72, October 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001963.

Borges:2005:CBS

[BZRO05] L. Borges, A. F. Zuur, E. Rogan, and R. Officer. Choosing the best sampling unit and auxiliary variable for discards estimations. Fisheries Research, 75(1–3):29–39, September 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605001372.

Carbonell:2003:ANE

[CA03] A. Carbonell and M. Azevedo. Application of nonequilibrium production models to the red shrimp (Aristeus antennatus, Risso, 1816) fishery in the northwestern Mediterranean. Fisheries Research, 65(1–3):323–334, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603002522.

Colvocoresses:2007:LSF

[CA07] James Colvocoresses and Alejandro Acosta. A largescale field comparison of strip transect and stationary point count methods for conducting length-based underwater visual surveys of reef fish populations. Fisheries Research, 85(1–2):130–141, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000318.

Caddy:2000:CCR

[Cad00a] J. F. Caddy. The code of conduct for responsible fisheries as a basis for evaluating fisheries research: a suggested operational procedure. Fisheries Research, 48(3):205–211, October 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600002307.

Caddy:2000:ESB

[Cad00b]  John Caddy.  Ecosystems structure before fishing?

Fisheries Research, 47(1):101, June 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001909.

Caddy:2002:LRP

[Cad02] J. F. Caddy. Limit reference points, traffic lights, and holistic approaches to fisheries management with minimal stock assessment input. Fisheries Research, 56(2):133–137, May 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601003435.

Caddy:2003:SET

[Cad03] J. F. Caddy. Scaling elapsed time: an alternative approach to modelling crustacean moulting schedules? Fisheries Research, 63(1):73–84, July 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002771.

Cadigan:2009:SCE

[Cad09] N. G. Cadigan. Sensitivity of common estimators of management parameters derived from stock–recruit relationships. Fisheries Research, 96(2–3):195–205, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003536.

Clark:2000:ECE

[CAFT00] Malcolm R. Clark, Owen F. Anderson, R. I. C. Chris Francis, and Dianne M. Tracey. The effects of commercial exploitation on orange roughy (Hoplostethus atlanticus) from the continental slope of the Chatham Rise, New Zealand, from 1979 to 1997. Fisheries Research, 45(3):217–238, April 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699001216.

Cole:2003:SCB

[CAH^{+}03] R. G. Cole, N. K. Alcock, S. J. Handley, K. R. Grange, S. Black, D. Cairney, J. Day, S. Ford, and A. R. Jerrett. Selective capture of blue cod Parapercis colias by potting: behavioural observations and effects of capture method on peri-mortem fatigue. Fisheries Research, 60(2–3):381–392, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602001339.

Chen:2007:UAS

[CAHAF07] Weizhong Chen, Mohsen Al-Husaini, and Hussain M. Al-Foudari. Using age-structured models to develop a stock recovery strategy for Kuwait’s shrimp fishery. Fisheries Research, 83(2–3):276–284, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600364X.

Cardinale:2000:EPU

[CAJ00a] M. Cardinale, F. Arrhenius, and B. Johnsson. Erratum to “Potential use of otolith weight for the determination of age-structure of Baltic cod (Gadus morhua) and plaice (Pleuronectes platessa)” [Fish. Res. 45 (2000) 239– 252]. Fisheries Research, 48(3):293, October 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002186. See [CAJ00b].

Cardinale:2000:PUO

[CAJ00b] M. Cardinale, F. Arrhenius, and B. Johnsson. Potential use of otolith weight for the determination of age-structure of Baltic cod (Gadus morhua) and plaice (Pleuronectes platessa). Fisheries Research, 45(3):239–252, April 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001228. See erratum [Bas99].

Carbonell:2003:CSW

[CAM^{+}03a] A. Carbonell, F. Alemany, P. Merella, A. Quetglas, and E. Román. The by-catch of sharks in the western Mediterranean (Balearic Islands) trawl fishery. Fisheries Research, 61(1–3):7–18, March 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002424.

Castro:2003:ERC

[CAM^{+}03b] Margarida Castro, Artur Araújo, Pedro Monteiro, Ana Margarida Madeira, and William Silvert. The efficacy of releasing caught Nephrops as a management measure. Fisheries Research, 65(1–3):475–484, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002625.

Campbell:2004:CSC

[Cam04] Robert A. Campbell. CPUE standardisation and the construction of indices of stock abundance in a spatially varying fishery using general linear models. Fisheries Research, 70(2–3):209–227, December 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001675.

Ceriola:2008:BEI

[CAM^{+}08] Luca Ceriola, Paolo Accadia, Piero Mannini, Fabio Massa, Nicoletta Milone, and Nicola Ungaro. A bio-economic indicators suite for the appraisal of the demersal trawl fishery in the Southern Adriatic Sea (Central Mediterranean). Fisheries Research, 92(2–3):255–267, August 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000453.

Cliff:2007:ACW

[CARA^{+}07] Geremy Cliff, Michael D. Anderson-Reade, Andrew P. Aitken, Graeme E. Charter, and Victor M. Peddemors. Aerial census of whale sharks (Rhincodon typus) on the northern KwaZulu-Natal coast, South Africa. Fisheries Research, 84(1):41–46, March 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003997.

Cole:2004:MEP

[CATH04] Russell G. Cole, Niki K. Alcock, Anna Tovey, and Sean J. Handley. Measuring efficiency and predicting optimal set durations of pots for blue cod Parapercis colias. Fisheries Research, 67(2):163–170, April 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002911.

Cazorla:2000:ASP

[Caz00] A. Lopez Cazorla. Age structure of the population of weakfish Cynoscion guatucupa (Cuvier) in the Bahía Blanca waters, Argentina. Fisheries Research, 46(1–3):279–286, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001521.

Candy:2000:PHS

[CB00] John R. Candy and Terry D. Beacham. Patterns of homing and straying in southern British Columbia coded-wire tagged Chinook salmon (Oncorhynchus tshawytscha) populations. Fisheries Research, 47(1):41–56, June 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001241.

Cubillos:2002:MAE

[CBC02] Luis A. Cubillos, Doris A. Bucarey, and Mariella Canales. Monthly abundance estimation for common sardine Strangomera bentincki and anchovy Engraulis ringens in the central-southern area off Chile (34–40^{o}S). Fisheries Research, 57(2):117–130, August 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100340X.

Clavero:2006:MSF

[CBGP06] M. Clavero, F. Blanco-Garrido, and J. Prenda. Monitoring small fish populations in streams: a comparison of four passive methods. Fisheries Research, 78(2–3):243–251, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606000270.

Cimmaruta:2008:GST

[CBRN08] R. Cimmaruta, P. Bondanelli, A. Ruggi, and G. Nascetti. Genetic structure and temporal stability in the horse mackerel (Trachurus trachurus). Fisheries Research, 89(2): 114–121, February 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783607002378.

Chicharo:2001:EEC

[CC01a] Luis Chıcharo and M. Alexandra Chıcharo. Effects of environmental conditions on planktonic abundances, benthic recruitment and growth rates of the bivalve mollusc Ruditapes decussatus in a Portuguese coastal lagoon. Fisheries Research, 53(3):235–250, November 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002903.

Chicharo:2001:JRP

[CC01b] Luis Chıćharo and M. Alexandra Chıćharo. A juvenile recruitment prediction model for Ruditapes decussatus (L.) (Bivalvia: Mollusca). Fisheries Research, 53(3):219–233, November 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600003003.

Chiu:2001:GTV

[CC01c] T. S. Chiu and C. S. Chen. Growth and temporal variation of two Japanese anchovy cohorts during their recruitment to the East China Sea. Fisheries Research, 53(1):1–15, September 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002940.

Carlson:2003:GSS

[CC03a] John K. Carlson and Enric Cortés. Gillnet selectivity of small coastal sharks off the southeastern United States. Fisheries Research, 60(2–3):405–414, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001352.

Castello:2003:ASE

[CC03b] Leandro Castello and Jorge P. Castello. Anchovy stocks (Engraulis anchoita) and larval growth in the SW Atlantic. Fisheries Research, 59(3):409–421, January 3, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000140.

Chen:2003:VLH

[CC03c] Chih-Shin Chen and Tai-Sheng Chiu. Variations of life history parameters in two geographical groups of the neon flying squid, Ommastrephes bartramii, from the North Pacific. Fisheries Research, 63(3):349–366, September 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001012.

Carbines:2009:URD

[CC09] Glen Carbines and Russell G. Cole. Using a remote drift underwater video (DUV) to examine dredge impacts on demersal fishes and benthic habitat complexity in Foveaux Strait, Southern New Zealand. Fisheries Research, 96(2–3):230–237, March 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003834.

Chicharo:2003:CED

[CCA^{+}03] Maria A. Chıćharo, Luis Chıćharo, Ana Amaral, Sílvia Condinho, and Miguel Gaspar. Chronic effects of dredginginduced stress on the clam (Spisula solida): nucleic acid and lipid composition. Fisheries Research, 63(3):447– 452, September 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783603000778.

Clarke:2002:CDA

[CCB02] M. W. Clarke, P. L. Connolly, and J. J. Bracken. Catch, discarding, age estimation, growth and maturity of the squalid shark Deania calceus west and north of Ireland. Fisheries Research, 56(2):139–153, May 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601004192.

Cotter:2002:PSS

[CCBG02] A. J. R. Cotter, G. Course, S. T. Buckland, and C. Garrod. A PPS sample survey of English fishing vessels to estimate discarding and retention of North Sea cod, haddock, and whiting. Fisheries Research, 55(1–3):25–35, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100306X.

Colloca:2004:SEA

[CCCC04] Francesco Colloca, Valerio Crespi, Sandro Cerasi, and Salvatore R. Coppola. Structure and evolution of the artisanal fishery in a southern Italian coastal area. Fisheries Research, 69(3):359–369, October 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001511.

Courtney:2001:PDS

[CCD01] A. J. Courtney, M. G. Cosgrove, and D. J. Die. Population dynamics of scyllarid lobsters of the genus Thenus spp. on the Queensland (Australia) east coast: I. Assessing the effects of tagging. Fisheries Research, 53(3):251–261, November 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600002745.

Chicharo:2002:RTI

[CCG^{+}02] L. Chıćharo, M. Chıćharo, M. Gaspar, J. Regala, and F. Alves. Reburial time and indirect mortality of Spisula solida clams caused by dredging. Fisheries Research, 59(1–2):247–257, December 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000127.

Campana:2000:OEF

[CCH^{+}00] S. E. Campana, G. A. Chouinard, J. M. Hanson, A. Fréchet, and J. Brattey. Otolith elemental fingerprints as biological tracers of fish stocks. Fisheries Research, 46(1–3):343–357, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001582.

Carlucci:2009:DPS

[CCM^{+}09] R. Carlucci, F. Capezzuto, P. Maiorano, L. Sion, and G. D’Onghia. Distribution, population structure and dynamics of the black anglerfish (Lophius budegassa) (Spinola, 1987) in the Eastern Mediterranean Sea. Fisheries Research, 95(1):76–87, January 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002257.

Caputi:2003:RRB

[CCMS^{+}03] N. Caputi, C. Chubb, R. Melville-Smith, A. Pearce, and D. Griffin. Review of relationships between life history stages of the western rock lobster, Panulirus cygnus, in Western Australia. Fisheries Research, 65(1–3):47–61, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603002340.

Chen:2008:AWW

[CCT^{+}08] Xinjun Chen, Yong Chen, Siquan Tian, Bilin Liu, and Weiguo Qian. An assessment of the west winter– spring cohort of neon flying squid (Ommastrephes bartramii) in the Northwest Pacific Ocean. Fisheries Research, 92(2–3):221–230, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000428.

Cabral:2003:DBS

[CDC03] Henrique Cabral, Joel Duque, and Maria José Costa. Discards of the beach seine fishery in the central coast of Portugal. Fisheries Research, 63(1):63–71, July 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000043.

Costa:2003:RFP

[CdCBdR03] Paulo A. S. Costa, Adriana da Costa Braga, and Luis Otávio Frota da Rocha. Reef fisheries in Porto Seguro, eastern Brazilian coast. Fisheries Research, 60(2–3):577–583, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602001455.

Coll:2007:ULT

[CdML^{+}07] Céline Coll, Luis Tito de Morais, Raymond Laë, Anne Lebourges-Dhaussy, Monique Simier, Jean Guillard, Erwan Josse, Jean-Marc Ecoutin, Jean-Jacques Albaret, Jean Raffray, and Justin Kantoussan. Use and limits of three methods for assessing fish size spectra and fish abundance in two tropical man-made lakes. Fisheries Research, 83(2–3):306–318, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003675.

Claro:2009:HAC

[CdMLGC09] Rodolfo Claro, Yvonne Sadovy de Mitcheson, Kenyon C. Lindeman, and Alida R. García-Cagide. Historical analysis of Cuban commercial fishing effort and the effects of management interventions on important reef fishes from 1960–2005. Fisheries Research, 99(1):7–16, July 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000952.

Coelho:2006:RAU

[CE06] Rui Coelho and Karim Erzini. Reproductive aspects of the undulate ray, Raja undulata, from the south coast of Portugal. Fisheries Research, 81(1):80–85, October 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002207.

Coelho:2007:PPS

[CE07] Rui Coelho and Karim Erzini. Population parameters of the smooth lantern shark, Etmopterus pusillus, in southern Portugal (NE Atlantic). Fisheries Research, 86(1):42–57, August 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607000938.

Cerrato:2000:WFB

[Cer00] Robert M. Cerrato. What fish biologists should know about bivalve shells. Fisheries Research, 46(1–3):39–49, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001314.

Campos:2004:USP

[CF04] Aida Campos and Paulo Fonseca. The use of separator panels and square mesh windows for by-catch reduction in the crustacean trawl fishery off the Algarve (South Portugal). Fisheries Research, 69(2):147–156, September 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001389.

Corgos:2007:AME

[CF07] Antonio Corgos and Juan Freire. Assessment of methods to estimate abundance and population structure of the spider crab Maja brachydactyla in soft substrates. Fisheries Research, 83(2–3):297–305, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003663.

Campos:2002:SSD

[CFE02] Aida Campos, Paulo Fonseca, and Karim Erzini. Size selectivity of diamond and square mesh cod ends for rose shrimp (Parapenaeus longirostris) and Norway lobster (Nephrops norvegicus) off the Portuguese south coast. Fisheries Research, 58(3):281–301, November 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003964.

Campos:2003:SSD

[CFE03] Aida Campos, Paulo Fonseca, and Karim Erzini. Size selectivity of diamond and square mesh cod ends for four bycatch species in the crustacean fishery off the Portuguese south coast. Fisheries Research, 60(1):79–97, January 30, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602000619.

Campos:2007:DFC

[CFFP07] Aida Campos, Paulo Fonseca, Tereza Fonseca, and Joaquim Parente. Definition of fleet components in the Portuguese bottom trawl fishery. Fisheries Research, 83(2–3):185–191, February 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606003390.

Catchpole:2005:DEN

[CFG05] T. L. Catchpole, C. L. J. Frid, and T. S. Gray. Discarding in the English north-east coast Nephrops norvegicus fishery: the role of social and environmental factors. Fisheries Research, 72(1):45–54, April 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002632.

Campos:2003:SSF

[CFH03] Aida Campos, Paulo Fonseca, and Victor Henriques. Size selectivity for four fish species of the deep groundfish assemblage off the Portuguese southwest coast: evidence of mesh size, mesh configuration and cod end catch effects. Fisheries Research, 63(2):213–233, August 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000602.

Cheng:2004:SAR

[CG04] Yuk W. Cheng and Michael P. Gallinat. Statistical analysis of the relationship among environmental variables, interannual variability and smolt trap efficiency of salmonids in the Tucannon River. Fisheries Research, 70(2–3):229–238, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604001687.

Czerwinski:2007:STF

[CGEHC07] Ivone Alejandra Czerwinski, Juan Carlos Gutiérrez-Estrada, and José Antonio Hernando-Casal. Short-term forecasting of halibut CPUE: Linear and non-linear univariate approaches. Fisheries Research, 86(2–3):120– 128, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783607001166.

Cantrell:2004:RAW

[CGLZ04] Robert N. Cantrell, Marissa Garcia, PingSun Leung, and David Ziemann. Recreational anglers’ willingness to pay for increased catch rates of Pacific threadfin (Polydactylus sexfilis) in Hawaii. Fisheries Research, 68(1–3):149–158, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604000608.

Cedrola:2005:BSE

[CGP05] P. V. Cedrola, A. M. González, and A. D. Pettovello. Bycatch of skates (Elasmobranchii: Arhynchobatidae, Rajidae) in the Patagonian red shrimp fishery. Fisheries Research, 71(2):141–150, February 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002097.

Craig:2008:SHC

[CGT08] P. Craig, A. Green, and F. Tuilagi. Subsistence harvest of coral reef resources in the outer islands of American Samoa: Modern, historic and prehistoric catches. Fisheries Research, 89(3):230–240, March 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002196.

Chen:2003:AGS

[CH03] Yong Chen and Margaret Hunter. Assessing the green sea urchin (Strongylocentrotus drobachiensis) stock in Maine, USA. Fisheries Research, 60(2–3):527–537, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000826.

Craig:2004:BFF

[CHBB04] J. F. Craig, A. S. Halls, J. J. F. Barr, and C. W. Bean. The Bangladesh floodplain fisheries. Fisheries Research, 66(2–3):271–286, February 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001966.

Criales-Hernandez:2006:EII

[CHDGM06] Maria Isabel Criales-Hernandez, Luis O. Duarte, Camilo B. García, and Luis Manjarrés. Ecosystem impacts of the introduction of bycatch reduction devices in a tropical shrimp trawl fishery: Insights through simulation. Fisheries Research, 77(3):333–342, March 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002961.

Chifamba:2000:RTH

[Chi00] Portia Chiyedza Chifamba. The relationship of temperature and hydrological factors to catch per unit effort, condition and size of the freshwater sardine, Limnothrissa miodon (Boulenger), in Lake Kariba. Fisheries Research, 45(3):271–281, April 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001150.

Chang:2001:MII

[CHL01] Shui-Kai Chang, Chien-Chung Hsu, and Hsi-Chiang Liu. Management implication on Indian Ocean albacore from simple yield analysis incorporating parameter uncertainty. Fisheries Research, 51(1):1–10, April 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002678.

Chiang:2006:PSB

[CHL^{+}06] Hsin-Chieh Chiang, Chien-Chung Hsu, Hung-Du Lin, Gwo Chin Ma, Tzen-Yuh Chiang, and Hsi-Yuan Yang. Population structure of bigeye tuna (Thunnus obesus) in the South China Sea, Philippine Sea and western Pacific Ocean inferred from mitochondrial DNA. Fisheries Research, 79(1–2):219–225, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000749.

Caillouet:2008:GOB

[CHN08] C. W. Caillouet, R. A. Hart, and J. M. Nance. Growth overfishing in the brown shrimp fishery of Texas, Louisiana, and adjoining Gulf of Mexico EEZ. Fisheries Research, 92(2– 3):289–302, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608000490.

Cunha:2005:EDN

[CHQ05] Teresa Cunha, Andreia Hall, and Henrique Queiroga. Estimation of the Diopatra neapolitana annual harvest resulting from digging activity in Canal de Mira, Ria de Aveiro. Fisheries Research, 76(1):56–66, October 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001487.

Cubillos:2002:EYC

[CHSA02] Luis A. Cubillos, Aldo Hernández, Aquiles Sepúlveda, and Dagoberto F. Arcos. Equilibrium yield-curve analysis through an analytic age-structured production model: a sensitivity study for the Chilean jack mackerel fishery. Fisheries Research, 54(3):395–407, February 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002739.

Chiang:2008:PSB

[CHW^{+}08] Hsin-Chieh Chiang, Chien-Chung Hsu, Georgiana Cho-Chen Wu, Shui-Kai Chang, and Hsi-Yuan Yang. Population structure of bigeye tuna (Thunnus obesus) in the Indian Ocean inferred from mitochondrial DNA. Fisheries Research, 90(1–3):305–312, April 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003128.

Chen:2000:ESE

[CJ00] Yong Chen and Donald A. Jackson. An empirical study on estimators for linear regression analyses in fisheries and ecology. Fisheries Research, 49(2):193–206, December 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001958.

Crozier:2002:ITC

[CK02] W. W. Crozier and G. J. A. Kennedy. Impact of tagging with coded wire tags on marine survival of wild Atlantic salmon (Salmo salar L.) migrating from the R. Bush, Northern Ireland. Fisheries Research, 59(1–2):209– 215, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783601004027.

Cox:2008:PSD

[CK08] Sean P. Cox and Allen Robert Kronlund. Practical stakeholder-driven harvest policies for groundfish fisheries in British Columbia, Canada. Fisheries Research, 94(3): 224–237, December 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608001586.

Christensen:2004:EIA

[CL04] Steen Christensen and Hans Lassen. The economic impact of the ACFM catch options for the Danish North Sea herring and industrial fisheries in 1999. Fisheries Research, 68(1–3):21–35, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000670.

Chang:2009:SAS

[CL09] Jui-Han Chang and Kwang-Ming Liu. Stock assessment of the shortfin mako shark (Isurus oxyrinchus) in the Northwest Pacific Ocean using per recruit and virtual population analyses. Fisheries Research, 98(1–3):92–101, June 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000940.

Clark:2001:DFS

[Cla01] Malcolm Clark. Are deepwater fisheries sustainable? — The example of orange roughy (Hoplostethus atlanticus) in New Zealand. Fisheries Research, 51(2–3):123–135, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601002405.

Chen:2008:RDC

[CLC08] Xinjun Chen, Bilin Liu, and Yong Chen. A review of the development of Chinese distant-water squid jigging fisheries. Fisheries Research, 89(3):211–221, March 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002755.

Coll:2004:SFB

[CLGR^{+}04] J. Coll, M. Linde, A. García-Rubies, F. Riera, and A. M. Grau. Spear fishing in the Balearic Islands (west central Mediterranean): species affected and catch evolution during the period 1975–2001. Fisheries Research, 70(1):97–111, November 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604001286.

Choi:2008:DMJ

[CLH^{+}08] Kwangho Choi, Chung Il Lee, Kwangseok Hwang, Sang-Woo Kim, Jong-Hwa Park, and Yeong Gong. Distribution and migration of Japanese common squid, Todarodes pacificus, in the southwestern part of the East (Japan) Sea. Fisheries Research, 91(2–3):281–290, June 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003517.

Cai:2005:ELI

[CLPP05] Junning Cai, PingSun Leung, Minling Pan, and Sam Pooley. Economic linkage impacts of Hawaii’s longline fishing regulations. Fisheries Research, 74(1–3):232–242, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605000627.

Claereboudt:2005:PRS

[CMAOAA05] M. R. Claereboudt, J. L. McIlwain, H. S. Al-Oufi, and A. A. Ambu-Ali. Patterns of reproduction and spawning of the kingfish (Scomberomorus commerson, Lacépède) in the coastal waters of the Sultanate of Oman. Fisheries Research, 73(3):273–282, July 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000664.

Comesana:2008:GVM

[CMAS08] Angel S. Comesaña, Ma. Teresa Martínez-Areal, and Andrés Sanjuan. Genetic variation in the mitochondrial DNA control region among horse mackerel (Trachurus trachurus) from the Atlantic and Mediterranean areas. Fisheries Research, 89(2):122–131, February 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360700238X.

Campo:2006:FHA

[CMC^{+}06] D. Campo, E. Mostarda, L. Castriota, M. P. Scarabello, and F. Andaloro. Feeding habits of the Atlantic bonito, Sarda sarda (Bloch, 1793) in the southern Tyrrhenian Sea. Fisheries Research, 81(2–3):169–175, November 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002803.

Correia:2009:AGR

[CMC09] A. T. Correia, S. Manso, and J. Coimbra. Age, growth and reproductive biology of the European conger eel (Conger conger) from the Atlantic Iberian waters. Fisheries Research, 99(3):196–202, August 15, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001477.

Chale-Matsau:2001:AGR

[CMGB01] J. R. Chale-Matsau, A. Govender, and L. E. Beckley. Age, growth and retrospective stock assessment of an economically extinct sparid fish, Polysteganus undulosus, from South Africa. Fisheries Research, 51(1):87–92, April 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002356.

Coombs:2001:SOC

[CMH01] S. H. Coombs, D. Morgans, and N. C. Halliday. Seasonal and ontogenetic changes in the vertical distribution of eggs and larvae of mackerel (Scomber scombrus L.) and horse mackerel (Trachurus trachurus L.). Fisheries Research, 50(1–2):27–40, February 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360000240X.

Campana:2005:BFS

[CMJ05] Steven E. Campana, Linda Marks, and Warren Joyce. The biology and fishery of shortfin mako sharks (Isurus oxyrinchus) in Atlantic Canadian waters. Fisheries Research, 73(3):341–352, July 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000305.

Chakalall:2007:GFO

[CMM^{+}07] Bisessar Chakalall, Robin Mahon, Patrick McConney, Leonard Nurse, and Derrick Oderson. Governance of fisheries and other living marine resources in the wider Caribbean. Fisheries Research, 87(1):92–99, October 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001476.

Courtney:2000:OJP

[CMOM00] Dean L. Courtney, Donald G. Mortensen, Joseph A. Orsi, and Kristen M. Munk. Origin of juvenile Pacific salmon recovered from coastal southeastern Alaska identified by otolith thermal marks and coded wire tags. Fisheries Research, 46(1–3):267–278, May 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360000151X.

Carrera:2001:SBW

[CMPM01] P. Carrera, M. Meixide, C. Porteiro, and J. Miquel. Study of the blue whiting movements around the Bay of Biscay using acoustic methods. Fisheries Research, 50(1–2):151–161, February 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600002484.

Cyterski:2002:PDC

[CND02] Michael Cyterski, John Ney, and Michael Duval. Predator demand for clupeid prey in Smith Mountain Lake, Virginia. Fisheries Research, 59(1–2):1–16, December 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000103.

Cardinale:2009:RHT

[CNH09] Massimiliano Cardinale, Duto Nugroho, and Lars Hernroth. Reconstructing historical trends of small pelagic fish in the Java Sea using standardized commercial trip based catch per unit of effort. Fisheries Research, 99(3):151–158, August 15, 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609001428.

Cha:2002:RPD

[COHP02] Hyung Kee Cha, Chul-Woong Oh, Sung Yun Hong, and Kyung Yang Park. Reproduction and population dynamics of Penaeus chinensis (Decapoda: Penaeidae) on the western coast of Korea, Yellow Sea. Fisheries Research, 56(1):25–36, April 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003101.

Castro-Ortiz:2007:LFV

[COLB07] Jose Luis Castro-Ortiz and Daniel Lluch-Belda. Low frequency variability of fishing resources, climate, and ocean. Fisheries Research, 85(1–2):186–196, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000392.

Cook:2000:BRF

[Coo00a] Robin Cook. Book review: Fishery stock assessment models: F. Funk, T. J. Quinn II, J. Heifetz, J. N. Ianelli, J. E. Powers, J. F. Schweigert, P. J. Sullivan, C.-I. Zhang (Eds.); Alaska Sea Grant College Program Report No. AK-SG-98-01, University of Fairbanks, 1998, 1037 pages, ISBN 1-56612-057-8 (US $40). Fisheries Research, 48(3):288–289, October 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600002216.

Cook:2000:RNF

[Coo00b] Robin Cook. Review of Northeast Fishery Stock Assessments: National Research Council; National Academy Press, Washington DC, 1998, 128 pages; £30.95, paperback, ISBN 0-309-06030-3. Fisheries Research, 47(1):98– 99, June 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600001211.

Cook:2001:BRE

[Coo01] Robin Cook. Book review: Ecosystem Approaches for Fisheries Management: University of Alaska Sea Grant, AK-SG-99-01, Fairbanks, AK, 1999, 756 pages, hardback, ISBN 1-56612-061-6 ($40). Fisheries Research, 51(1):93–94, April 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600003106.

Corten:2001:NDN

[Cor01] Ad Corten. Northern distribution of North Sea herring as a response to high water temperatures and/or low food abundance. Fisheries Research, 50(1–2):189–204, February 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600002514.

Corkett:2002:FSA

[Cor02] C. J. Corkett. Fish stock assessment as a non-falsifiable science: replacing an inductive and instrumental view with a critical rational one. Fisheries Research, 56(2):117–123, May 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601003526.

Corkett:2009:KPD

[Cor09] Christopher Corkett. Karl Popper’s definition of science. Fisheries Research, 96(2–3):323, March 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003937.

Casas:2000:GAE

[CP00] José Miguel Casas and Carmen Piñeiro. Growth and age estimation of greater fork-beard (Phycis blennoides Brünnich, 1768) in the north and northwest of the Iberian Peninsula (ICES Division VIIIc and IXa). Fisheries Research, 47(1):19–25, June 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783600001089.

Campbell:2006:RFS

[CP06] S. J. Campbell and S. T. Pardede. Reef fish structure and cascading effects in response to artisanal fishing pressure. Fisheries Research, 79(1–2):75–83, June 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000221.

Cadrin:2008:PHP

[CP08a] Steven X. Cadrin and Martin A. Pastoors. Precautionary harvest policies and the uncertainty paradox. Fisheries Research, 94(3):367–372, December 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360800180X.

Cheung:2008:ESE

[CP08b] William W. L. Cheung and Tony J. Pitcher. Evaluating the status of exploited taxa in the northern South China Sea using intrinsic vulnerability and spatially explicit catchper-unit-effort data. Fisheries Research, 92(1):28–40, July 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360700358X.

Castrejon:2005:SSB

[CPCD05] Hugo Castrejón, Roberto Pérez-Castañeda, and Omar Defeo. Spatial structure and bathymetric patterns of penaeoid shrimps in the southwestern Gulf of Mexico. Fisheries Research, 72(2–3):291–300, May 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360400253X.

Cucherousset:2007:UPT

[CPR07] Julien Cucherousset, Jean-Marc Paillisson, and Jean-Marc Roussel. Using PIT technology to study the fate of hatchery-reared YOY northern pike released into shallow vegetated areas. Fisheries Research, 85(1–2):159–164, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607000343.

Cubillos:2008:SSO

[CPR^{+}08] Luis A. Cubillos, Jorge Paramo, Patricia Ruiz, Sergio Núñez, and Aquiles Sepúlveda. The spatial structure of the oceanic spawning of jack mackerel (Trachurus murphyi) off central Chile (1998–2001). Fisheries Research, 90(1–3):261– 270, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607002998.

Castro-Pampillon:2002:SPU

[CPSBRD^{+}02] J. A. Castro-Pampillón, M. Soto-Búa, H. Rodríguez-Domínguez, J. Mejuto-García, C. Arias-Fernández, and J. M. García-Estévez. Selecting parasites for use in biological tagging of the Atlantic swordfish (Xiphias gladius). Fisheries Research, 59(1–2):259–262, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000085.

Chilari:2006:ABB

[CPT06] Anna Chilari, George Petrakis, and Evaggelos Tsamis. Aspects of the biology of blackspot seabream (Pagellus bogaraveo) in the Ionian Sea, Greece. Fisheries Research, 77(1):84–91, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002687.

Chen:2006:ADL

[CPW^{+}06] C. S. Chen, G. J. Pierce, J. Wang, J.-P. Robin, J. C. Poulard, J. Pereira, A. F. Zuur, P. R. Boyle, N. Bailey, D. J. Beare, P. Jereb, S. Ragonese, A. Mannini, and L. Orsi-Relini. The apparent disappearance of Loligo forbesi from the south of its range in the 1990s: Trends in Loligo spp. abundance in the northeast Atlantic and possible environmental influences. Fisheries Research, 78(1):44–54, April 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605003267.

Chen:2008:SFM

[CQJX08] Zuozhi Chen, Yongsong Qiu, Xiaoping Jia, and Shannan Xu. Simulating fisheries management options for the Beibu Gulf by means of an ecological modelling optimization routine. Fisheries Research, 89(3):257–265, March 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002238.

Christensen:2006:FTS

[CR06] Anne-Sofie Christensen and Jesper Raakjær. Fishermen’s tactical and strategic decisions: a case study of Danish demersal fisheries. Fisheries Research, 81(2–3):258–267, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606002463.

Cubillos:2007:SDE

[CRC^{+}07] Luis A. Cubillos, Patricia Ruiz, Gabriel Claramunt, Santiago Gacitúa, Sergio Núñez, Leonardo R. Castro, Katty Riquelme, Carolina Alarcón, Ciro Oyarzún, and Aquiles Sepúlveda. Spawning, daily egg production, and spawning stock biomass estimation for common sardine (Strangomera bentincki) and anchovy (Engraulis ringens) off central southern Chile in 2002. Fisheries Research, 86(2–3):228–240, September 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001300.

Catchpole:2006:ASG

[CRD06] T. L. Catchpole, A. S. Revill, and G. Dunlin. An assessment of the Swedish grid and square-mesh codend in the English (Farn Deeps) Nephrops fishery. Fisheries Research, 81(2–3):118–125, November 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002906.

Chicharo:2002:MSD

[CRG^{+}02] L. Chıćharo, J. Regala, M. Gaspar, F. Alves, and A. Chıćharo. Macrofauna spatial differences within clam dredge-tracks and their implications for short-term fishing effect studies. Fisheries Research, 54(3):349–354, February 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601002727.

Crone:2001:SDM

[Cro01] P. R. Crone. Spatial differences in maturity schedules of female Dover sole off Oregon (1989–1991). Fisheries Research, 51(2–3):393–402, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002636.

Cooper:2004:EIC

[CRSM04] Andrew B. Cooper, Andrew A. Rosenberg, Gunnar Stefánsson, and Marc Mangel. Examining the importance of consistency in multi-vessel trawl survey design based on the U.S. west coast groundfish bottom trawl survey. Fisheries Research, 70(2–3):239–250, December 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001699.

Caddy:2002:RCF

[CS02] J. F. Caddy and J. C. Seijo. Reproductive contributions foregone with harvesting: a conceptual framework. Fisheries Research, 59(1–2):17–30, December 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000115.

Cardinale:2004:MRA

[CS04] Massimiliano Cardinale and Henrik Svedäng. Modelling recruitment and abundance of Atlantic cod, Gadus morhua, in the eastern Skagerrak–Kattegat (North Sea): evidence of severe depletion due to a prolonged period of high fishing pressure. Fisheries Research, 69(2):263–282, September 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604000803.

Claverie:2007:CET

[CS07a] Thomas Claverie and I. Philip Smith. A comparison of the effect of three common tagging methods on the survival of the galatheid Munida rugosa (Fabricius, 1775). Fisheries Research, 86(2–3):285–288, September 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001580.

Courtney:2007:VOI

[CS07b] Dean L. Courtney and Kenneth P. Severin. Validation of otolith increment daily periodicity in captive juvenile sablefish (Anoplopoma fimbria) experimentally immersed in strontium chloride (SrCl_{2}). Fisheries Research, 83(2–3):246–252, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600347X.

Cardinale:2008:FL

[CS08a] M. Cardinale and H. Svedäng. To have a fling on the ling. Fisheries Research, 89(3):305–306, March 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/

science/article/pii/S0165783607002548. See reponse

[PCHM08].

Cardinale:2008:MFP

[CS08b] Massimiliano Cardinale and Henrik Svedäng. Mismanagement of fisheries: Policy or science? Fisheries Research, 93(1–2):244–247, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001628.

Cazorla:2008:AGL

[CS08c] Andrea Lopez Cazorla and Nora Sidorkewicj. Age and growth of the largemouth perch Percichthys colhuapiensis in the Negro river, Argentine Patagonia. Fisheries Research, 92(2–3):169–179, August 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000374.

Campbell:2009:GFP

[CS09] Matthew J. Campbell and Wayne D. Sumpton. Ghost fishing in the pot fishery for blue swimmer crabs Portunus pelagicus in Queensland, Australia. Fisheries Research, 95(2–3):246–253, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003111.

Claramunt:2007:SFD

[CSCC07] G. Claramunt, R. Serra, L. R. Castro, and L. Cubillos. Is the spawning frequency dependent on female size? Empirical evidence in Sardinops sagax and Engraulis ringens off northern Chile. Fisheries Research, 85(3):248–257, July 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607000094.

Charles:2007:P

[CSCS07] Anthony Charles, Silvia Salas, Ratana Chuenpagdee, and Juan Carlos Seijo. Preface. Fisheries Research, 87(1):1–4, October 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360700152X.

Cetinic:2002:SMF

[CSDP02] Perica Cetinić, Alen Soldo, Jakov Dulčić, and Armin Pallaoro. Specific method of fishing for Sparidae species in the eastern Adriatic. Fisheries Research, 55(1–3):131–139, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601002995.

Cardenas:2009:GPS

[CSM^{+}09] Leyla Cárdenas, Andrea X. Silva, Antonios Magoulas, Josefina Cabezas, Elie Poulin, and F. Patricio Ojeda. Genetic population structure in the Chilean jack mackerel, Trachurus murphyi (Nichols) across the South-eastern Pacific Ocean. Fisheries Research, 100(2):109–115, October 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001799.

Cervino:2006:UBI

[CSR06] Santiago Cerviño and Fran Saborido-Rey. Using the bootstrap to investigate the effects of varying tow lengths and catch sampling schemes in fish survey. Fisheries Research, 79(3):294–302, July 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001305.

Cooke:2003:IRH

[CSSO03] Steven J. Cooke, Cory D. Suski, Michael J. Siepker, and Kenneth G. Ostrand. Injury rates, hooking efficiency and mortality potential of largemouth bass (Micropterus salmoides) captured on circle hooks and octopus hooks. Fisheries Research, 61(1–3):135–144, March 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002448.

Campana:2000:SOC

[CT00] S. E. Campana and W.-N. Tzeng. Section 4: Otolith composition. Fisheries Research, 46(1–3):287–288, May 2000.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001533.

Cole:2001:CDP

[CTB01] R. G. Cole, D. S. Tindale, and R. G. Blackwell. A comparison of diver and pot sampling for blue cod (Parapercis colias: Pinguipedidae). Fisheries Research, 52(3):191–201, July 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600002575.

Courtney:2006:QEB

[CTC^{+}06] A. J. Courtney, M. L. Tonks, M. J. Campbell, D. P. Roy, S. W. Gaddes, P. M. Kyne, and M. F. O’Neill. Quantifying the effects of bycatch reduction devices in Queensland’s (Australia) shallow water eastern king prawn (Penaeus plebejus) trawl fishery. Fisheries Research, 80(2–3):136–147, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001780.

Cardenas-Torres:2007:CBM

[CTEARD07] Nirari Cárdenas-Torres, Roberto Enríquez-Andrade, and Natalie Rodríguez-Dowdell. Community-based management through ecotourism in Bahia de los Angeles, Mexico. Fisheries Research, 84(1):114–118, March 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004085.

Catchpole:2007:PNN

[CTK^{+}07] T. L. Catchpole, A. N. Tidd, L. T. Kell, A. S. Revill, and G. Dunlin. The potential for new Nephrops trawl designs to positively effect North Sea stocks of cod, haddock and whiting. Fisheries Research, 86(2–3):262–267, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607001452.

Cury:2000:BRD

[Cur00] P. Cury. Book review: Dynamics of Pelagic Fish Distribution and Behaviour: Effects on Fisheries and Stock Assessment; Pierre Fréon, Ole Arve Misund (Eds.); Fishing News Books. Blackwell, Osney Mead, Oxford, 1998, 360 pages, £59.50, hardback, ISBN 0-85238-241-3. Fisheries Research, 48(1):93–94, August 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001235.

Ceriola:2006:SIB

[CUT06] Luca Ceriola, Nicola Ungaro, and Francesco Toteda. Some information on the biology of Illex coindetii Verany, 1839 (Cephalopoda, Ommastrephidae) in the South-Western Adriatic Sea (Central Mediterranean). Fisheries Research, 82(1–3):41–49, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600302X.

Carvalho:2009:RCB

[CWJS09] Adriana Rosa Carvalho, Samantha Williams, Mariam January, and Merle Sowman. Reliability of community-based data monitoring in the Olifants River estuary (South Africa). Fisheries Research, 96(2–3):119–128, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002324.

Cappo:2000:ERM

[CWL00] Mike Cappo, Carl J. Walters, and Rodney C. Lenanton. Estimation of rates of migration, exploitation and survival using tag recovery data for western Australian “salmon” (Arripis truttaceus: Arripidae: Percoidei). Fisheries Research, 44(3):207–217, January 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000910.

Caddy:2005:UET

[CWS^{+}05] J. F. Caddy, E. Wade, T. Surette, M. Hebert, and M. Moriyasu. Using an empirical traffic light procedure for monitoring and forecasting in the Gulf of St. Lawrence fishery for the snow crab, Chionoecetes opilio. Fisheries Research, 76(1):123–145, October 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001712.

Chen:2007:SSI

[CXCD07] Yong Chen, Liuxiong Xu, Xinjun Chen, and Xiaojie Dai. A simulation study of impacts of at-sea discarding and bycatch on the estimation of biological reference points F_{0.1} and F_{max}. Fisheries Research, 85(1–2):14–22, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004310.

Chen:2004:OPS

[CXMM04] D. G. Chen, Y. Xie, T. J. Mulligan, and D. N. MacLennan. Optimal partition of sampling effort between observations of fish density and migration speed for a riverine hydroacoustic duration-in-beam method. Fisheries Research, 67(3):275–282, May 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603003424.

Chen:2009:MEF

[CXQ^{+}09] Zuozhi Chen, Shannan Xu, Yongsong Qiu, Zhaojin Lin, and Xiaoping Jia. Modeling the effects of fishery management and marine protected areas on the Beibu Gulf using spatial ecosystem simulation. Fisheries Research, 100(3):222–229, November 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609002082.

Chen:2006:DAB

[CY06] Pimao Chen and Weiwen Yuan. Demographic analysis based on the growth parameter of sharks. Fisheries Research, 78(2–3):374–379, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600035X.

Dew:2007:ARK

[DA07] C. Braxton Dew and Roberta G. Austring. Alaska red king crab: a relatively intractable target in a multispecies trawl survey of the eastern Bering Sea. Fisheries Research, 85(1–2):165–173, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000367.

Danylchuk:2008:EIS

[DACS08] Andy J. Danylchuk, Aaron Adams, Steven J. Cooke, and Cory D. Suski. An evaluation of the injury and short-term survival of bonefish (Albula spp.) as influenced by a mechanical lip-gripping device used by recreational anglers. Fisheries Research, 93(1–2):248–252, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001781.

Dahm:2000:CLC

[Dah00] E. Dahm. Changes in the length compositions of some fish species as a consequence of alterations in the groundgear of the GOV-trawl. Fisheries Research, 49(1):39–50, November 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001922.

Duarte:2001:RAL

[DALP01] Rafael Duarte, Manuela Azevedo, Jorge Landa, and Pilar Pereda. Reproduction of anglerfish (Lophius budegassa Spinola and Lophius piscatorius Linnaeus) from the Atlantic Iberian coast. Fisheries Research, 51(2–3):349–361, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601002594.

Daw:2008:SDE

[Daw08] Tim M. Daw. Spatial distribution of effort by artisanal fishers: Exploring economic factors affecting the lobster fisheries of the Corn Islands, Nicaragua. Fisheries Research, 90(1–3):17–25, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002299.

Deroba:2008:RHP

[DB08] Jonathan J. Deroba and James R. Bence. A review of harvest policies: Understanding relative performance of control rules. Fisheries Research, 94(3):210–223, December 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000192.

Deval:2006:SPP

[DBAÖ06] Mehmet Cengiz Deval, Tomris Bök, Celal Ates, and Hüseyin Özbilgin. Selectivity of PE and PA material codends for rose shrimp (Parapenaus longirostris) in Turkish twin rigged beam trawl fishery. Fisheries Research, 81(1):72–79, October 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001809.

Demer:2000:MTD

[DBB00] David A. Demer, Manuel Barange, and Alan J. Boyd. Measurements of three-dimensional fish school velocities with an acoustic Doppler current profiler. Fisheries Research, 47(2–3):201–214, July 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001703.

DeOliveira:2007:DAS

[DBK07] J. A. A. De Oliveira, H. J. Boyer, and C. H. Kirchner. Developing age-structured stock assessment models as a basis for management procedure evaluations for Namibian sardine. Fisheries Research, 85(1–2):148–158, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000331.

DOnghia:2005:CPS

[DCM^{+}05] G. D’Onghia, F. Capezzuto, Ch. Mytilineou, P. Maiorano, K. Kapiris, R. Carlucci, L. Sion, and A. Tursi.

Comparison of the population structure and dynamics of Aristeus antennatus (Risso, 1816) between exploited and unexploited areas in the Mediterranean Sea. Fisheries Research, 76(1):22–38, October 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001463.

Dixon:2006:SSH

[DDHS06] Cameron D. Dixon, Rob W. Day, Sylvain M. H. Huchette, and Scoresby A. Shepherd. Successful seeding of hatcheryproduced juvenile greenlip abalone to restore wild stocks. Fisheries Research, 78(2–3):179–185, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000415.

Dichmont:2006:MSSa

[DDP^{+}06a] Catherine M. Dichmont, Aijun (Roy) Deng, André E. Punt, William Venables, and Malcolm Haddon. Management strategies for short-lived species: the case of Australia’s Northern Prawn Fishery: 1. Accounting for multiple species, spatial structure and implementation uncertainty when evaluating risk. Fisheries Research, 82(1–3):204–220, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606002293.

Dichmont:2006:MSSb

[DDP^{+}06b] Catherine M. Dichmont, Aijun (Roy) Deng, André E. Punt, William Venables, and Malcolm Haddon. Management strategies for short lived species: the case of Australia’s Northern Prawn Fishery: 2. Choosing appropriate management strategies using input controls. Fisheries Research, 82(1–3):221–234, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002268.

Dichmont:2006:MSSc

[DDP^{+}06c] Catherine M. Dichmont, Aijun (Roy) Deng, André E. Punt, William Venables, and Malcolm Haddon. Management strategies for short lived species: the case of Australia’s Northern Prawn Fishery: 3. factors affecting management and estimation performance. Fisheries Research, 82(1–3):235–245, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002281.

Dichmont:2008:BBP

[DDP^{+}08] C. M. Dichmont, A. Deng, A. E. Punt, N. Ellis, W. N. Venables, T. Kompas, Y. Ye, S. Zhou, and J. Bishop. Beyond biological performance measures in management strategy evaluation: Bringing in economics and the effects of trawling on the benthos. Fisheries Research, 94(3):238–250, December 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608001434.

Dempster:2004:BFA

[Dem04] Tim Dempster. Biology of fish associated with moored fish aggregation devices (FADs): implications for the development of a FAD fishery in New South Wales, Australia. Fisheries Research, 68(1–3):189–201, July 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000475.

Dunn:2002:CCR

[DFD02] A. Dunn, R. I. C. C. Francis, and I. J. Doonan. Comparison of the Chapman–Robson and regression estimators of Z from catch–curve data when non-sampling stochastic error is present. Fisheries Research, 59(1–2):149–159, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601004076.

Durr:2002:FHB

[DG02] J. Dürr and J. A. González. Feeding habits of Beryx splendens and Beryx decadactylus (Berycidae) off the Canary Islands. Fisheries Research, 54(3):363–374, February 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002697.

Deval:2008:PSD

[DG08] Mehmet Cengiz Deval and Didem Göktürk. Population structure and dynamics of the cut trough shell Spisula subtruncata (da Costa) in the Sea of Marmara, Turkey. Fisheries Research, 89(3):241–247, March 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002214.

DeVries:2002:UOS

[DGP02] Douglas A. DeVries, Churchill B. Grimes, and Michael H. Prager. Using otolith shape analysis to distinguish eastern Gulf of Mexico and Atlantic Ocean stocks of king mackerel. Fisheries Research, 57(1):51–62, July 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003320.

Dubula:2005:ETR

[DGS^{+}05] O. Dubula, J. C. Groeneveld, J. Santos, D. L. van Zyl, S. L. Brouwer, N. van den Heever, and S. A. McCue. Effects of tag-related injuries and timing of tagging on growth of rock lobster, Jasus lalandii. Fisheries Research, 74(1–3):1–10, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605001256.

deGraaf:2006:VSS

[dGvZM^{+}06] Martin de Graaf, Paul A. M. van Zwieten, Marcel A. M. Machiels, Endale Lemma, Tesfaye Wudneh, Eshete Dejen, and Ferdinand A. Sibbing. Vulnerability to a smallscale commercial fishery of Lake Tana’s (Ethiopia) endemic Labeobarbus compared with African catfish and Nile tilapia: an example of recruitment-overfishing? Fisheries Research, 82(1–3):304–318, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002128.

Dammannagoda:2008:EFG

[DHM08] Sudath T. Dammannagoda, David A. Hurwood, and Peter B. Mather. Evidence for fine geographical scale heterogeneity in gene frequencies in yellowfin tuna (Thunnus albacares) from the north Indian Ocean around Sri Lanka. Fisheries Research, 90(1–3):147–157, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002597.

Diachok:2000:ASN

[Dia00] Orest Diachok. Absorption spectroscopy: a new approach to estimation of biomass. Fisheries Research, 47(2–3):231– 244, July 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600001727.

Dick:2004:BLV

[Dic04] E. J. Dick. Beyond ‘lognormal versus gamma’: discrimination among error distributions for generalized linear models. Fisheries Research, 70(2–3):351–366, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001778.

Dingsor:2005:EAI

[Din05] Gjert Endre Dingsør. Estimating abundance indices from the international 0-group fish survey in the Barents Sea. Fisheries Research, 72(2–3):205–218, May 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002589.

Doray:2006:ACP

[DJG^{+}06] Mathieu Doray, Erwan Josse, Paul Gervain, Lionel Reynal, and Josselin Chantrel. Acoustic characterisation of pelagic fish aggregations around moored fish aggregating devices in Martinique (Lesser Antilles). Fisheries Research, 82(1–3):162–175, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002414.

Dankel:2009:CWI

[DJGC09] Dorothy J. Dankel, Nikki Jacobson, Dan Georgianna, and Steven X. Cadrin. Can we increase haddock yield within the constraints of the Magnuson–Stevens Act? Fisheries Research, 100(3):240–247, November 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002240.

Drastik:2005:FAA

[DK05] Vladislav Draštík and Jan Kubečka. Fish avoidance of acoustic survey boat in shallow waters. Fisheries Research, 72(2–3):219–228, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002577.

Dulcic:2000:AGM

[DKGC00] Jakov Dulčić, Miro Kraljević, Branka Grbec, and Perica Cetinić. Age, growth and mortality of blotched picarel Spicara maena L. (Pisces: Centracanthidae) in the eastern central Adriatic. Fisheries Research, 48(1):69–78, August 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001120.

Dunn:2008:CMS

[DKH08] Daniel C. Dunn, Connie Y. Kot, and Patrick N. Halpin. A comparison of methods to spatially represent pelagic longline fishing effort in catch and bycatch studies. Fisheries Research, 92(2–3):268–276, August 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000465.

Diop:2007:PAW

[DKKS07] Hamady Diop, Walter R. Keithly, Richard F. Kazmierczak, and Richard F. Shaw. Predicting the abundance of white shrimp (Litopenaeus setiferus) from environmental parameters and previous life stages. Fisheries Research, 86(1):31–41, August 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000926.

Damalas:2007:EST

[DMA07] Dimitrios Damalas, Persefoni Megalofonou, and Maria Apostolopoulou. Environmental, spatial, temporal and operational effects on swordfish (Xiphias gladius) catch rates of eastern Mediterranean Sea longline fisheries. Fisheries Research, 84(2):233–246, April 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003845.

Drouineau:2008:AID

[DMBF08] H. Drouineau, S. Mahévas, M. Bertignac, and A. Fertin. Assessing the impact of discretisation assumptions in a length-structured population growth model. Fisheries Research, 91(2–3):160–167, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003372.

DOnghia:2009:FED

[DMC^{+}09] G. D’Onghia, P. Maiorano, F. Capezzuto, R. Carlucci, D. Battista, A. Giove, L. Sion, and A. Tursi. Further evidences of deep-sea recruitment of Aristeus antennatus (Crustacea: Decapoda) and its role in the population renewal on the exploited bottoms of the Mediterranean. Fisheries Research, 95(2–3):236–245, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360800307X.

DaRos:2003:BBE

[DNN^{+}03] L. Da Ros, N. Nesto, C. Nasci, V. Moschino, D. Pampanin, and M. G. Marin. Biochemical and behavioural effects of hydraulic dredging on the target species Chamelea gallina. Fisheries Research, 64(1):71–78, October 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002017.

Davis:2006:WRI

[DO06] M. W. Davis and M. L. Ottmar. Wounding and reflex impairment may be predictors for mortality in discarded or escaped fish. Fisheries Research, 82(1–3):1–6, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360600316X.

Diaz-Ochoa:2008:RPF

[DOQ08] Javier A. Díaz-Ochoa and Renato A. Quiñones. Relationship of precipitation, freshwater input, and sea level height with the abundance of the white shrimp (Litopenaeus occidentalis; Street, 1871) off Buenaventura, eastern tropical Pacific. Fisheries Research, 92(2–3):148–161, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000180.

Dougall:2004:AUS

[Dou04] Andrew Mc Dougall. Assessing the use of sectioned otoliths and other methods to determine the age of the centropomid fish, barramundi (Lates calcarifer) (Bloch), using knownage fish. Fisheries Research, 67(2):129–141, April 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002996.

Dougherty:2008:DSD

[Dou08] A. B. Dougherty. Daily and sub-daily otolith increments of larval and juvenile walleye pollock, Theragra chalcogramma (Pallas), as validated by alizarin complexone experiments. Fisheries Research, 90(1–3):271–278, April 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003013.

Dichmont:2003:AWD

[DPD^{+}03] C. M. Dichmont, A. E. Punt, A. Deng, Q. Dell, and W. Venables. Application of a weekly delay-difference model to commercial catch and effort data for tiger prawns in Australia’s Northern Prawn Fishery. Fisheries Research, 65(1–3):335–350, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002534.

Daly:2001:CCT

[DPS^{+}01] H. I. Daly, G. J. Pierce, M. B. Santos, J. Royer, S. K. Cho, G. Stowasser, J.-P. Robin, and S. M. Henderson. Cephalopod consumption by trawl caught fish in Scottish and English Channel waters. Fisheries Research, 52(1–2):51–64, June 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601002302.

Denis:2001:PSF

[DR01] Vincent Denis and Jean-Paul Robin. Present status of the French Atlantic fishery for cuttlefish (Sepia officinalis). Fisheries Research, 52(1–2):11–22, June 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002272.

Demirhan:2007:LHS

[DS07] Sefa Ayhan Demirhan and Kadir Seyhan. Life history of spiny dogfish, Squalus acanthias (L. 1758), in the southern Black Sea. Fisheries Research, 85(1–2):210–216, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000525.

deSousa:2006:RAM

[dSBAC06] L. Palha de Sousa, A. Brito, S. Abdula, and N. Caputi. Research assessment for the management of the industrial shallow-water multi-species shrimp fishery in Sofala Bank in Mozambique. Fisheries Research, 77(2):207–219, February 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605002882.

dosSantos:2001:CDM

[dSH01] Roberta Aguiar dos Santos and Manuel Haimovici. Cephalopods in the diet of marine mammals stranded or incidentally caught along southeastern and southern Brazil (21–34^{o}S). Fisheries Research, 52(1–2):99–112, June 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100234X.

Aires-da-Silva:2008:HIA

[dSHG08] A. M. Aires da Silva, J. J. Hoey, and V. F. Gallucci. A historical index of abundance for the blue shark (Prionace glauca) in the western North Atlantic. Fisheries Research, 92(1):41–52, July 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003633.

Dowling:2008:DHS

[DSK^{+}08] Natalie A. Dowling, David C. Smith, Ian Knuckey, Anthony D. M. Smith, Phil Domaschenz, Heather M. Patterson, and Wade Whitelaw. Developing harvest strategies for low-value and data-poor fisheries: Case studies from three Australian fisheries. Fisheries Research, 94(3):380–390, December 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608003160.

Davidsen:2005:STM

[DSO^{+}05] Jan Davidsen, Martin-A. Svenning, Panu Orell, Nigel Yoccoz, J. Brian Dempson, Eero Niemelä, Anders Klemetsen, Anders Lamberg, and Jaakko Erkinaro. Spatial and temporal migration of wild Atlantic salmon smolts determined from a video camera array in the sub-Arctic River Tana. Fisheries Research, 74(1–3):210–222, August 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000639.

Dominguez-Seoane:2006:AGS

[DSPLR06] Rosa Domínguez-Seoane, José G. Pajuelo, José M. Lorenzo, and Antonio G. Ramos. Age and growth of the sharpsnout seabream Diplodus puntazzo (Cetti, 1777) inhabiting the Canarian archipelago, estimated by reading otoliths and by backcalculation. Fisheries Research, 81(2–3):142–148, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606002852.

Domiguez-Torreiro:2007:CNC

[DTSR07] Marcos Domíguez-Torreiro and Juan C. Surís-Regueiro. Cooperation and non-cooperation in the Ibero-Atlantic sardine shared stock fishery. Fisheries Research, 83(1):1–10, January 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606002864.

Dunbrack:2006:SMF

[Dun06] R. L. Dunbrack. In situ measurement of fish body length using perspective-based remote stereo-video. Fisheries Research, 82(1–3):327–331, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002840.

Dunn:2009:ICB

[Dun09] Peter K. Dunn. Improving comparisons between models for CPUE. Fisheries Research, 97(1–2):148–149, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000058. See reply [Sho09].

DeOliveira:2005:PIM

[DUR05] José A. A. De Oliveira, Andres Uriarte, and Beatriz A. Roel. Potential improvements in the management of Bay of Biscay anchovy by incorporating environmental indices as recruitment predictors. Fisheries Research, 75(1–3):2–14, September 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001438.

Dennerline:2000:SPC

[DV00] Donald E. Dennerline and Michael J. Van Den Avyle. Sizes of prey consumed by two pelagic predators in US reservoirs: implications for quantifying biomass of available prey. Fisheries Research, 45(2):147–154, March 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001083.

Deceliere-Verges:2009:SPF

[DVAL^{+}09] C. Deceliere-Vergès, C. Argillier, C. Lanoiselée, J. De Bortoli, and J. Guillard. Stability and precision of the fish metrics obtained using CEN multi-mesh gillnets in natural and artificial lakes in France. Fisheries Research, 99(1):17– 25, July 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609000976.

Dahm:2002:ITS

[DWW^{+}02] E. Dahm, H. Wienbeck, C. W. West, J. W. Valdemarsen, and F. G. O’Neill. On the influence of towing speed and gear size on the selective properties of bottom trawls. Fisheries Research, 55(1–3):103–119, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003010.

Davis:2004:CTT

[DYWHZ04] Jana L. D. Davis, Alicia C. Young-Williams, Anson H. Hines, and Oded Zmora. Comparing two types of internal tags in juvenile blue crabs. Fisheries Research, 67(3):265– 274, May 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603003436.

Ecoutin:2005:LWR

[EAT05] J. M. Ecoutin, J. J. Albaret, and S. Trape. Length– weight relationships for fish populations of a relatively undisturbed tropical estuary: The Gambia. Fisheries Research, 72(2–3):347–351, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002565.

Eaton:2003:ECC

[EBA^{+}03] D. R. Eaton, J. Brown, J. T. Addison, S. P. Milligan, and L. J. Fernand. Edible crab (Cancer pagurus) larvae surveys off the east coast of England: implications for stock structure. Fisheries Research, 65(1–3):191–199, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002431.

Ettahiri:2003:OSS

[EBV^{+}03] O. Ettahiri, Am Berraho, G. Vidy, M. Ramdani, and T. Do chi. Observation on the spawning of Sardina and Sardinella off the south Moroccan Atlantic coast (21–26^{o}N). Fisheries Research, 60(2–3):207–222, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001728.

Enever:2009:SSR

[ECEG09] R. Enever, T. L. Catchpole, J. R. Ellis, and A. Grant. The survival of skates (Rajidae) caught by demersal trawlers fishing in UK waters. Fisheries Research, 97(1–2):72–76, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609000046.

Ehrhardt:2007:BNG

[ED07] Nelson M. Ehrhardt and Vallierre K. W. Deleveaux. The Bahamas’ Nassau grouper (Epinephelus striatus) fishery — two assessment methods applied to a data-deficient coastal population. Fisheries Research, 87(1):17–27, October 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001488.

Eero:2004:CMP

[Eer04] Margit Eero. Consequences of management of pikeperch (Stizostedion lucioperca L.) stock in Pärnu Bay (Baltic Sea) under two different economic regimes, 1960–1999. Fisheries Research, 68(1–3):1–7, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000736.

Elliott:2001:CTM

[EF01] J. M. Elliott and J. M. Fletcher. A comparison of three methods for assessing the abundance of Arctic charr, Salvelinus alpinus, in Windermere (northwest England). Fisheries Research, 53(1):39–46, September 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002708.

Engaas:2002:RRS

[EF02] Arill Engås and Daniel Foster. The response of red snapper (Lutjanus campechanus) and pinfish (Lagodon rhomboides) to inclined water flow. Fisheries Research, 58(3):315–321, November 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003988.

Erzini:2001:HDS

[EGB^{+}01] K. Erzini, J. M. S. Gonçalves, L. Bentes, P. G. Lino, and J. Ribeiro. The hake deepwater semi-pelagic (“pedrabola”) longline fishery in the Algarve (southern Portugal). Fisheries Research, 51(2–3):327–336, May 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002569.

Erzini:2006:SST

[EGB^{+}06] Karim Erzini, Jorge M. S. Gonçalves, Luís Bentes, Dimitrios K. Moutopoulos, Jose A. Hernando Casal, Milagrosa C. Soriguer, Esteban Puente, Luis A. Errazkin, and Konstantinos I. Stergiou. Size selectivity of trammel nets in southern European small-scale fisheries. Fisheries Research, 79(1–2):183–201, June 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600107X.

Engaas:2000:CBV

[EGJ00] Arill Engås, Olav Rune Godø, and Terje Jørgensen. A comparison between vessel and trawl tracks as observed by the ITI trawl instrumentation. Fisheries Research, 45(3):297– 301, April 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578369900123X.

Eide:2002:EIG

[EH02] Arne Eide and Knut Heen. Economic impacts of global warming: a study of the fishing industry in North Norway. Fisheries Research, 56(3):261–274, June 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003241.

Eigaard:2004:ESC

[EH04a] Ole Ritzau Eigaard and René Holst. The effective selectivity of a composite gear for industrial fishing: a sorting grid in combination with a square mesh window. Fisheries Research, 68(1–3):99–112, July 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000451.

Engelhard:2004:MCN

[EH04b]  Georg H. Engelhard and Mikko Heino.  Maturity

changes in Norwegian spring-spawning herring before, during, and after a major population collapse. Fisheries Research, 66(2–3):299–310, February 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001954.

Ehrhardt:2008:EGF

[Ehr08] Nelson M. Ehrhardt. Estimating growth of the Florida spiny lobster, Panulirus argus, from molt frequency and size increment data derived from tag and recapture experiments. Fisheries Research, 93(3):332–337, September 28, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608001860.

Einoder:2009:RUS

[Ein09] L. D. Einoder. A review of the use of seabirds as indicators in fisheries and ecosystem management. Fisheries Research, 95(1):6–13, January 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003093.

Engaas:2000:ECR

[EJA00] Arill Engås, Terje Jørgensen, and Karl K. Angelsen. Effects on catch rates of baiting gillnets. Fisheries Research, 45(3):265–270, April 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783699001174.

Murshed-e-Jahan:2009:MFC

[eJSK09] Khondker Murshed e Jahan, Nerissa D. Salayo, and Usha Kanagaratnam. Managing fisheries conflicts through communication planning: Experience from inland fisheries of Bangladesh. Fisheries Research, 99(2):112–122, August 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609001027.

Etim:2002:BGM

[EKU02] Lawrence Etim, Richard P. King, and Mfon T. Udo. Breeding, growth, mortality and yield of the mudskipper Periophthalmus barbarus (Linnaeus 1766) (Teleostei: Gobiidae) in the Imo River estuary, Nigeria. Fisheries Research, 56(3):227–238, June 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003277.

Ewing:2009:RDR

[EL09] G. P. Ewing and J. M. Lyle. Reproductive dynamics of redbait, Emmelichthys nitidus (Emmelichthyidae), from southeastern Australia. Fisheries Research, 97(3):206–215, May 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609000459.

Elliott:2000:BRB

[Ell00] J. Malcolm Elliott. Book review: Biology and Ecology of the Brown and Sea Trout: J. L. Baglinière, G. Maisse (Eds.); Springer, Heidelberg, in association with Praxis Publishing Ltd., Chichester, 1999, vii + 286 pp., ISBN 1-85233-117-8, £45 (hbk). Fisheries Research, 47(1):97–98, June 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360000120X.

Eero:2007:DIF

[EMKG07] Margit Eero, Brian R. MacKenzie, Hrefna M. Karlsdóttir, and Ritma Gaumiga. Development of international fisheries for the eastern Baltic cod (Gadus morhua) from the late 1880s until 1938. Fisheries Research, 87(2–3):155–166, November 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607000586.

Enghoff:2007:DFF

[EMN07] Inge B. Enghoff, Brian R. MacKenzie, and Einar Eg Nielsen. The Danish fish fauna during the warm Atlantic period (ca. 7000–3900 BC): Forerunner of future changes? Fisheries Research, 87(2–3):167–180, November 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000598.

Eklund:2000:BRV

[EPA00] Jan Eklund, Raimo Parmanne, and Gunnar Aneer. Between-reader variation in herring otolith ages and effects on estimated population parameters. Fisheries Research, 46(1–3):147–154, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001417.

Enever:2007:DEC

[ERG07] R. Enever, A. Revill, and A. Grant. Discarding in the English Channel, Western approaches, Celtic and Irish seas (ICES subarea VII). Fisheries Research, 86(2–3):143–152, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001191.

Enever:2009:DNS

[ERG09] R. Enever, A. S. Revill, and A. Grant. Discarding in the North Sea and on the historical efficacy of gearbased technical measures in reducing discards. Fisheries Research, 95(1):40–46, January 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002191.

Eros:2009:AFA

[ESB09] T. Erős, A. Specziár, and P. Bíró. Assessing fish assemblages in reed habitats of a large shallow lake — a comparison between gillnetting and electric fishing. Fisheries Research, 96(1):70–76, February 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360800283X.

Ebert:2008:RBT

[ESC08] David A. Ebert, Wade D. Smith, and Gregor M. Cailliet. Reproductive biology of two commercially exploited skates, Raja binoculata and R. rhina, in the western Gulf of Alaska. Fisheries Research, 94(1):48–57, October 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001951.

Ehrenberg:2000:FSC

[ET00] John E. Ehrenberg and Thomas C. Torkelson. FM slide (chirp) signals: a technique for significantly improving the signal-to-noise performance in hydroacoustic assessment systems. Fisheries Research, 47(2–3):193–199, July 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001697.

Ezenwaji:2002:BCE

[Eze02] H. M. G. Ezenwaji. The biology of Clarias ebriensis Pellegrin, 1920 (Osteichthyes: Clariidae) in an African rainforest river basin. Fisheries Research, 54(2):235–252, January 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600002915.

Fablet:2006:SLA

[Fab06] Ronan Fablet. Statistical learning applied to computerassisted fish age and growth estimation from otolith images. Fisheries Research, 81(2–3):219–228, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002657.

Fives:2001:DAL

[FAL^{+}01] J. M. Fives, S. Acevedo, M. Lloves, A. Whitaker, M. Robinson, and P. A. King. The distribution and abundance of larval mackerel, Scomber scombrus L., horse mackerel, Trachurus trachurus (L.), hake, Merluccius merluccius (L.), and blue whiting, Micromesistius poutassou (Risso, 1826) in the Celtic Sea and west of Ireland during the years 1986, 1989 and 1992. Fisheries Research, 50(1–2):17–26, February 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600002393.

Fossen:2003:IPA

[FAN03] Inge Fossen, Ole Thomas Albert, and Einar M. Nilssen. Improving the precision of ageing assessments for long rough dab by using digitised pictures and otolith measurements. Fisheries Research, 60(1):53–64, January 30, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000632.

Funamoto:2004:RCJ

[FAW04] Tetsuichiro Funamoto, Ichiro Aoki, and Yozo Wada. Reproductive characteristics of Japanese anchovy, Engraulis japonicus, in two bays of Japan. Fisheries Research, 70 (1):71–81, November 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604001596.

Fossen:2006:DBB

[FB06] Inge Fossen and Odd Aksel Bergstad. Distribution and biology of blue hake, Antimora rostrata (Pisces: Moridae), along the mid-Atlantic Ridge and off Greenland. Fisheries Research, 82(1–3):19–29, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003092.

Folkvord:2000:GDA

[FBJM00] Arild Folkvord, Geir Blom, Arne Johannessen, and Erlend Moksness. Growth-dependent age estimation in herring (Clupea harengus L.) larvae. Fisheries Research, 46 (1–3):91–103, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783600001363.

Fonteyne:2007:LFT

[FBL^{+}07] Ronald Fonteyne, Gabriele Buglioni, Iole Leonori, Finbarr G. O’Neill, and Robert J. Fryer. Laboratory and field trials of OMEGA, a new objective mesh gauge. Fisheries Research, 85(1–2):197–201, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000410.

Fonteyne:2007:RMM

[FBLO07] Ronald Fonteyne, Gabriele Buglioni, Iole Leonori, and Finbarr G. O’Neill. Review of mesh measurement methodologies. Fisheries Research, 85(3):279–284, July 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000562.

Froysa:2002:FAL

[FBS02] Kristin Guldbrandsen Frøysa, Bjarte Bogstad, and Dankert W. Skagen. Fleksibest — an age–length structured fish stock assessment model. Fisheries Research, 55(1–3):87–101, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601003071.

Fry:2006:VDD

[FBV06] G. C. Fry, D. T. Brewer, and W. N. Venables. Vulnerability of deepwater demersal fishes to commercial fishing: Evidence from a study around a tropical volcanic seamount in Papua New Guinea. Fisheries Research, 81(2–3):126–141, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606002876.

Fonseca:2008:TCP

[FCAD^{+}08] Tereza Fonseca, Aida Campos, Manuel Afonso-Dias, Paulo Fonseca, and João Pereira. Trawling for cephalopods off the Portuguese coast — fleet dynamics and landings composition. Fisheries Research, 92(2–3):180–188, August 2008.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000386.

Fablet:2009:RIS

[FCC^{+}09] Ronan Fablet, Anatole Chessel, Sebastien Carbini, Abdesslam Benzinou, and Hélène de Pontual. Reconstructing individual shape histories of fish otoliths: a new image-based tool for otolith growth analysis and modeling. Fisheries Research, 96(2–3):148–159, March 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003366.

Fonseca:2002:BTC

[FCG02] Paulo Fonseca, Aida Campos, and Alexandra Garcia. Bottom trawl codend selectivity for cephalopods in Portuguese continental waters. Fisheries Research, 59(1–2): 263–271, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783601004180.

Fonseca:2005:EUM

[FCL^{+}05a] Paulo Fonseca, Aida Campos, Roger B. Larsen, Teresa C. Borges, and Karim Erzini. Erratum to “Using a modified Nordmøre grid for by-catch reduction in the Portuguese crustacean trawl fishery” [Fish. Res. 71 (2005) 223–239]. Fisheries Research, 75(1–3):204–205, September 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001323. See [FCL^{+}05b].

Fonseca:2005:UMN

[FCL^{+}05b] Paulo Fonseca, Aida Campos, Roger B. Larsen, Teresa C. Borges, and Karim Erzini. Using a modified Nordmøre grid for by-catch reduction in the Portuguese crustaceantrawl fishery. Fisheries Research, 71(2):223–239, February 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001833. See erratum

[FCL^{+}05a].

Fonseca:2007:CSD

[FCM07] Paulo Fonseca, Aida Campos, and Russell B. Millar. Codend selection in the deep-water crustacean trawl fishery in Portuguese southern waters. Fisheries Research, 85(1–2):49–60, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004346.

Fonseca:2005:PUN

[FCML05] Paulo Fonseca, Aida Campos, Beatriz Mendes, and Roger B. Larsen. Potential use of a Nordmøre grid for bycatch reduction in a Portuguese bottom-trawl multispecies fishery. Fisheries Research, 73(1–2):49–66, June 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000214.

Fletcher:2005:FPF

[FCS^{+}05] W. J. Fletcher, J. Chesson, K. J. Sainsbury, T. J. Hundloe, and M. Fisher. A flexible and practical framework for reporting on ecologically sustainable development for wild capture fisheries. Fisheries Research, 71(2):175–183, February 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604001869.

Farley:2007:DPS

[FDG^{+}07] Jessica H. Farley, Tim L. O. Davis, John S. Gunn, Naomi P. Clear, and Ann L. Preece. Demographic patterns of southern bluefin tuna, Thunnus maccoyii, as inferred from direct age data. Fisheries Research, 83(2–3):151–161, February 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606003341.

Fromentin:2001:FEL

[FF01] Jean-Marc Fromentin and Alain Fonteneau. Fishing effects and life history traits: a case study comparing tropical versus temperate tunas. Fisheries Research, 53(2):133–150, October 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360000299X.

Fredou:2009:ASPa

[FFL09a] Thierry Frédou, Beatrice P. Ferreira, and Yves Letourneur. Assessing the stocks of the primary snappers caught in Northeastern Brazilian reef systems. 1: Traditional modelling approaches. Fisheries Research, 99(2):90–96, August 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609001088.

Fredou:2009:ASPb

[FFL09b] Thierry Frédou, Beatrice P. Ferreira, and Yves Letourneur. Assessing the stocks of the primary snappers caught in Northeastern Brazilian reef systems. 2. A multi-fleet agestructured approach. Fisheries Research, 99(2):97–105, August 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360900109X.

Figueiredo:2001:DWP

[FFM01] Maria José Figueiredo, Ivone Figueiredo, and Pedro Bordalo Machado. Deep-water penaeid shrimps (Crustacea: Decapoda) from off the Portuguese continental slope: an alternative future resource? Fisheries Research, 51(2–3):321– 326, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601002557.

Francini-Filho:2008:ESR

[FFM08] R. B. Francini-Filho and R. L. Moura. Evidence for spillover of reef fishes from a no-take marine reserve: an evaluation using the before-after control-impact (BACI) approach. Fisheries Research, 93(3):346–356, September 28, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608001896.

Fazli:2008:PEP

[FGJD08] Hasan Fazli, Davod Ghaninejad, Ali Asghar Janbaz, and Reza Daryanabard. Population ecology parameters and biomass of golden grey mullet (Liza aurata) in Iranian waters of the Caspian Sea. Fisheries Research, 93(1–2):222–228, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001355.

Fassler:2008:DSV

[FGOF08] Sascha M. M. Fässler, Natalia Gorska, Egil Ona, and Paul G. Fernandes. Differences in swimbladder volume between Baltic and Norwegian spring-spawning herring: Consequences for mean target strength. Fisheries Research, 92(2–3):314–321, August 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000532.

Farina:2003:TCP

[FH03a] A. C. Fariña and I. González Herraiz. Trends in catchper-unit-effort, stock biomass and recruitment in the North and Northwest Iberian Atlantic Nephrops stocks. Fisheries Research, 65(1–3):351–360, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002546.

Frusher:2003:RDE

[FH03b] S. D. Frusher and J. M. Hoenig. Recent developments in estimating fishing and natural mortality and tag reporting rate of lobsters using multi-year tagging models. Fisheries Research, 65(1–3):379–390, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360300256X.

Frusher:2003:CSM

[FHG03] S. D. Frusher, J. M. Hoenig, and C. Gardner. Have changes in selectivity masked recruitment declines in crustacean trap fisheries? Fisheries Research, 65(1–3):467–474, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603002613.

Frusher:2007:ECA

[FHI07] S. D. Frusher, J. M. Hoenig, and T. F. Ihde. Evaluating catchability assumptions for change-in-ratio and indexremoval estimators, with application to southern rock lobster. Fisheries Research, 84(2):254–262, April 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003869.

Furevik:2008:FFP

[FHJL08] Dag M. Furevik, Odd-Børre Humborstad, Terje Jørgensen, and Svein Løkkeborg. Floated fish pot eliminates bycatch of red king crab and maintains target catch of cod. Fisheries Research, 92(1):23–27, July 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003554.

Frandsen:2009:ETL

[FHM09] Rikke P. Frandsen, René Holst, and Niels Madsen. Evaluation of three levels of selective devices relevant to management of the Danish Kattegat–Skagerrak Nephrops fishery. Fisheries Research, 97(3):243–252, May 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000666.

Fousekis:2003:TED

[FK03] Panos Fousekis and Stathis Klonaris. Technical efficiency determinants for fisheries: a study of trammel netters in Greece. Fisheries Research, 63(1):85–95, July 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000195.

Frouzova:2004:CAT

[FK04] Jaroslava Frouzova and Jan Kubecka. Changes of acoustic target strength during juvenile perch development. Fisheries Research, 66(2–3):355–361, February 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001826.

Frouzova:2005:TSS

[FKBF05] Jaroslava Frouzova, Jan Kubecka, Helge Balk, and Jan Frouz. Target strength of some European fish species and its dependence on fish body parameters. Fisheries Research, 75(1–3):86–96, September 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001426.

Fossum:2000:F

[FKM00] P. Fossum, J. Kalish, and E. Moksness. Foreword. Fisheries Research, 46(1–3):1–2, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001260.

Fare:2006:ATE

[FKW06] Rolf Färe, James E. Kirkley, and John B. Walden. Adjusting technical efficiency to reflect discarding: the case of the U.S. Georges Bank multi-species otter trawl fishery. Fisheries Research, 78(2–3):257–265, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000257.

Fablet:2005:AFA

[FL05] Ronan Fablet and Nicolas Le Josse. Automated fish age estimation from otolith images using statistical learning. Fisheries Research, 72(2–3):279–290, May 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002528.

Fogarty:2004:ICR

[FM04] Michael J. Fogarty and Thomas J. Miller. Impact of a change in reporting systems in the Maryland blue crab fishery. Fisheries Research, 68(1–3):37–43, July 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000694.

Figueiredo:2005:FEW

[FMB^{+}05] Miguel Figueiredo, Telmo Morato, João P. Barreiros, Pedro Afonso, and Ricardo S. Santos. Feeding ecology of the white seabream, Diplodus sargus, and the ballan wrasse, Labrus bergylta, in the Azores. Fisheries Research, 75(1–3):107–119, September 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001645.

Fobert:2009:CLR

[FMC^{+}09] Emily Fobert, Patrick Meining, Alison Colotelo, Constance O’Connor, and Steven J. Cooke. Cut the line or remove the hook? An evaluation of sublethal and lethal endpoints for deeply hooked bluegill. Fisheries Research, 99(1):38– 46, July 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609001015.

Fonseca:2005:GNS

[FMCS05] Paulo Fonseca, Rogélia Martins, Aida Campos, and Preciosa Sobral. Gill-net selectivity off the Portuguese western coast. Fisheries Research, 73(3):323–339, July 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000366.

Fernandez-Macho:2008:EIT

[FMGG08] Javier Fernández-Macho, Carmen Gallastegui, and Pilar González. Economic impacts of TAC regulation: a supplydriven SAM approach. Fisheries Research, 90(1–3):225– 234, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607002950.

Fritsch:2007:ASB

[FML^{+}07] M. Fritsch, Y. Morizur, E. Lambert, F. Bonhomme, and B. Guinand. Assessment of sea bass (Dicentrarchus labrax, L.) stock delimitation in the Bay of Biscay and the English Channel based on mark-recapture and genetic data. Fisheries Research, 83(2–3):123–132, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003134.

Ferro:2008:POY

[FÖB08] R. S. T. Ferro, H. Özbilgin, and M. Breen. The potential for optimizing yield from a haddock trawl fishery using seasonal changes in selectivity, population structure and fish condition. Fisheries Research, 94(2):151–159, November 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608002683.

Flores:2008:ISV

[FOGKT08] Marcela S. Zúñiga Flores, Sofía Ortega-García, and Alexander Klett-Traulsen. Interannual and seasonal variation of dolphinfish (Coryphaena hippurus) catch rates in the southern Gulf of California, Mexico. Fisheries Research, 94(1):13–17, October 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001793.

Folmer:2000:SED

[FP00] Ole Folmer and Michael Pennington. A statistical evaluation of the design and precision of the shrimp trawl survey off West Greenland. Fisheries Research, 49(2):165–178, December 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360000196X.

Fonteyne:2002:RBC

[FP02] Ronald Fonteyne and Hans Polet. Reducing the benthos by-catch in flatfish beam trawling by means of technical modifications. Fisheries Research, 55(1–3):219–230, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601002879.

Fairgrieve:2003:IAS

[FR03] William T. Fairgrieve and Michael B. Rust. Interactions of Atlantic salmon in the Pacific Northwest: V. Human health and safety. Fisheries Research, 62(3):329–338, June 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603000675.

Fudge:2008:LHC

[FR08] Susan B. Fudge and George A. Rose. Life history co-variation in a fishery depleted Atlantic cod stock. Fisheries Research, 92(1):107–113, July 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000623.

Frazer:2007:CSL

[Fra07] L. Neil Frazer. Comment on “Sea lice on adult Pacific salmon in the coastal waters of British Columbia, Canada” by R. J. Beamish et al. Fisheries Research, 85(3):328–331, July 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606003742.

Ford:2008:CBO

[FRF08] Jennifer S. Ford, Tom Rudolph, and Susanna D. Fuller. Cod bycatch in otter trawls and in longlines with different bait types in the Georges Bank haddock fishery. Fisheries Research, 94(2):184–189, November 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001124.

Fernandez-Rueda:2007:OVM

[FRGF07] P. Fernández-Rueda and L. García-Flórez. Octopus vulgaris (Mollusca: Cephalopoda) fishery management assessment in Asturias (north-west Spain). Fisheries Research, 83(2–3):351–354, February 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003687.

Frisch:2007:GRP

[Fri07] Ashley J. Frisch. Growth and reproduction of the painted spiny lobster (Panulirus versicolor) on the Great Barrier Reef (Australia). Fisheries Research, 85(1–2):61–67, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606004425.

Fabi:2002:TNG

[FSB^{+}02] G. Fabi, M. Sbrana, F. Biagi, F. Grati, I. Leonori, and P. Sartor. Trammel net and gill net selectivity for Lithognathus mormyrus (L., 1758), Diplodus annularis (L., 1758) and Mullus barbatus (L., 1758) in the Adriatic and Ligurian seas. Fisheries Research, 54(3):375–388, February 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601002703.

Froese:2008:SMH

[FSPWG08] Rainer Froese, Amanda Stern-Pirlot, Henning Winker, and Didier Gascuel. Size matters: How single-species management can contribute to ecosystem-based fisheries management. Fisheries Research, 92(2–3):231–241, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360800043X. See comments [Sve14,

Sve15] and response [FWG15].

Feet:2002:AAE

[FUM02] Per Øystein Feet, Karl Inne Ugland, and Erlend Moksness. Accuracy of age estimates in spring spawning herring (Clupea harengus L.) reared under different prey densities. Fisheries Research, 56(1):59–67, April 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003137.

Fischer:2006:FAC

[FW06] Sönke Fischer and Matthias Wolff. Fisheries assessment of Callinectes arcuatus (Brachyura, Portunidae) in the Gulf of Nicoya, Costa Rica. Fisheries Research, 77(3):301–311, March 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605003036.

Froese:2015:SSM

[FWG15] Rainer Froese, Henning Winker, and Didier Gascuel. Size still matters. a response to Svedäng (2013): Size matters: Ne quid nimis. Fisheries Research, 164(??):329–330, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783614003014.  See

[FSPWG08, Sve14].

Fazli:2009:SAM

[FZHL09] Hasan Fazli, Chang Ik Zhang, Douglas Edward Hay, and Chun-Woo Lee. Stock assessment and management implications of anchovy kilka (Clupeonella engrauliformis) in Iranian waters of the Caspian Sea. Fisheries Research, 100(2):103–108, October 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360900160X.

Garaway:2006:SSP

[GAC^{+}06] C. J. Garaway, R. I. Arthur, B. Chamsingh, P. Homekingkeo, K. Lorenzen, B. Saengvilaikham, and K. Sidavong. A social science perspective on stock enhancement outcomes: Lessons learned from inland fisheries in southern Lao PDR. Fisheries Research, 80(1):37–45, August 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001184.

Grandcourt:2006:AGM

[GAF06] Edwin Mark Grandcourt, Thabit Zahran Al Abdessalaam, and Franklin Francis. Age, growth, mortality and reproduction of the blackspot snapper, Lutjanus fulviflamma (Forsskål, 1775), in the southern Arabian Gulf. Fisheries Research, 78(2–3):203–210, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000348.

Grandcourt:2004:PBA

[GAFA04a] E. M. Grandcourt, T. Z. Al Abdessalaam, F. Francis, and A. Al Shamsi. Population biology and assessment of representatives of the family Carangidae: Carangoides bajad and Gnathanodon speciosus (Forsskål, 1775), in the Southern Arabian Gulf. Fisheries Research, 69(3):331–341, October 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604001419.

Grandcourt:2004:BSA

[GAFA04b] E. M. Grandcourt, T. Z. Al Abdessalaam, F. Francis, and A. T. Al Shamsi. Biology and stock assessment of the sparids, Acanthopagrus bifasciatus and Argyrops spinifer (Forsskål, 1775), in the Southern Arabian Gulf. Fisheries Research, 69(1):7–20, August 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000888.

Grandcourt:2005:PBA

[GAFA05a] E. M. Grandcourt, T. Z. Al Abdessalaam, F. Francis, and A. T. Al Shamsi. Population biology and assessment of the orange-spotted grouper, Epinephelus coioides (Hamilton, 1822), in the southern Arabian Gulf. Fisheries Research, 74(1–3):55–68, August 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001128.

Grandcourt:2005:PAB

[GAFA05b] E. M. Grandcourt, T. Z. Al Abdessalaam, F. Francis, and A. T. Al Shamsi. Preliminary assessment of the biology and fishery for the narrow-barred Spanish mackerel, Scomberomorus commerson (Lacépède, 1800), in the southern Arabian Gulf. Fisheries Research, 76(2):277–290, November 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360500202X.

Govender:2006:PRA

[GAOMC06] A. Govender, H. Al-Oufi, J. L. McIlwain, and M. C. Claereboudt. A per-recruit assessment of the kingfish (Scomberomorus commerson) resource of Oman with an evaluation of the effectiveness of some management regulations. Fisheries Research, 77(2):239–247, February 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002742.

Gaspar:2003:SSS

[Gas03] M. Gaspar. Size selectivity of the Spisula solida dredge in relation to tooth spacing and mesh size. Fisheries Research, 60(2–3):561–568, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001406.

Gavaris:2009:FMP

[Gav09] Stratis Gavaris. Fisheries management planning and support for strategic and tactical decisions in an ecosystem approach context. Fisheries Research, 100(1):6–14, September 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608003883.

Gao:2003:SIV

[GB03] Yongwen Gao and Richard J. Beamish. Stable isotope variations in otoliths of Pacific halibut (Hippoglossus stenolepis) and indications of the possible 1990 regime shift. Fisheries Research, 60(2–3):393–404, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001340.

Glazer:2006:SRA

[GB06] J. P. Glazer and D. S. Butterworth. Some refinements of the assessment of the South African squid resource, Loligo vulgaris reynaudii. Fisheries Research, 78(1):14–25, April 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605003322.

Grant:2007:FKE

[GB07] Sandra Grant and Fikret Berkes. Fisher knowledge as expert system: a case from the longline fishery of Grenada, the Eastern Caribbean. Fisheries Research, 84(2):162–170, April 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606003766.

Goncalves:2003:AGM

[GBC^{+}03a] J. M. S. Gonçalves, L. Bentes, R. Coelho, C. Correia, P. G. Lino, C. C. Monteiro, J. Ribeiro, and K. Erzini. Age and growth, maturity, mortality and yield-per-recruit for two banded bream (Diplodus vulgaris geoffr.) from the south coast of Portugal. Fisheries Research, 62(3):349–359, June 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602002801. See erratum

[GBC^{+}03b].

Goncalves:2003:EAG

[GBC^{+}03b] J. M. S. Gonçalves, L. Bentes, R. Coelho, C. Correia, P. G. Lino, P. Monteiro, J. Ribeiro, and K. Erzini. Erratum to “Age and growth, maturity, mortality and yield-per-recruit for two banded bream (Diplodus vulgaris Geoffr.) from the south coast of Portugal”: [Fish. Res. 62 (2003) 349–359]. Fisheries Research, 63(2):299, August 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001486. See [GBC^{+}03a].

Gristina:2006:CDF

[GBFG06] M. Gristina, T. Bahri, F. Fiorentino, and G. Garofalo. Comparison of demersal fish assemblages in three areas of the Strait of Sicily under different trawling pressure. Fisheries Research, 81(1):60–71, October 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001822.

Groeneveld:2003:EAI

[GBG^{+}03] Johan C. Groeneveld, Douglas S. Butterworth, Jean P. Glazer, George M. Branch, and Andrew C. Cockcroft. An experimental assessment of the impact of gear saturation on an abundance index for an exploited rock lobster resource. Fisheries Research, 65(1–3):453–465, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002601.

Gobert:2005:ESS

[GBL^{+}05] Bertrand Gobert, Patrick Berthou, Ester Lopez, Patrick Lespagnol, Maria Dolores Oqueli Turcios, Christophe Macabiau, and Pedro Portillo. Early stages of snapper– grouper exploitation in the Caribbean (Bay Islands, Honduras). Fisheries Research, 73(1–2):159–169, June 2005.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604003121.

Graham:2009:ECC

[GBM09] Ken J. Graham, Matt K. Broadhurst, and Russell B. Millar. Effects of codend circumference and twine diameter on selection in south-eastern Australian fish trawls. Fisheries Research, 95(2–3):341–349, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003251.

Gamito:2003:MBS

[GC03a] Rita Gamito and Henrique Cabral. Mortality of brownshrimp discards from the beam trawl fishery in the Tagus estuary, Portugal. Fisheries Research, 63(3):423– 427, September 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783603001085.

Grandcourt:2003:BEI

[GC03b] Edwin M. Grandcourt and Herman S. J. Cesar. The bioeconomic impact of mass coral mortality on the coastal reef fisheries of the Seychelles. Fisheries Research, 60(2–3): 539–550, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360200173X.

Grabowski:2004:IUE

[GC04] Robert Grabowski and Yong Chen. Incorporating uncertainty into the estimation of the biological reference points F_{0.1} and F_{max} for the Maine green sea urchin (Strongylocentrotus droebachiensis) fishery. Fisheries Research, 68(1–3):367–371, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000293.

Gordo:2008:DVI

[GCA^{+}08] L. S. Gordo, A. Costa, P. Abaunza, P. Lucio, A. T. G. W. Eltink, and I. Figueiredo. Determinate versus indeterminate fecundity in horse mackerel. Fisheries Research, 89(2):181–185, February 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002457.

Gianguzza:2006:ERP

[GCB^{+}06] Paola Gianguzza, Mariachiara Chiantore, Chiara Bonaviri, Riccardo Cattaneo-Vietti, Ilaria Vielmini, and Silvano Riggio. The effects of recreational Paracentrotus lividus fishing on distribution patterns of sea urchins at Ustica Island MPA (Western Mediterranean, Italy). Fisheries Research, 81(1):37–44, October 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002232.

Gunn:2008:AGS

[GCC^{+}08] John S. Gunn, Naomi P. Clear, Thor I. Carter, Anthony J. Rees, Clive A. Stanley, Jessica H. Farley, and John M. Kalish. Age and growth in southern bluefin tuna, Thunnus maccoyii (Castelnau): Direct estimation from otoliths, scales and vertebrae. Fisheries Research, 92(2–3):207–220, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608000416.

Godlewska:2009:HMT

[GCD^{+}09] M. Godlewska, M. Colon, L. Doroszczyk, B. Długoszewski, C. Verges, and J. Guillard. Hydroacoustic measurements at two frequencies: 70 and 120 kHz — consequences for fish stock estimation. Fisheries Research, 96(1):11–16, February 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608002750.

Gingerich:2007:EIE

[GCH^{+}07] Andrew J. Gingerich, Steven J. Cooke, Kyle C. Hanson, Michael R. Donaldson, Caleb T. Hasler, Cory D. Suski, and Robert Arlinghaus. Evaluation of the interactive effects of air exposure duration and water temperature on the condition and survival of angled and released fish. Fisheries Research, 86(2–3):169–178, September 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001233.

Grixti:2007:EHB

[GCJ07] Daniel Grixti, Simon D. Conron, and Paul L. Jones. The effect of hook/bait size and angling technique on the hooking location and the catch of recreationally caught black bream Acanthopagrus butcheri. Fisheries Research, 84(3):338–344, May 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606004231.

Guest:2003:SNB

[GCL03] Michaela A. Guest, Rod M. Connolly, and Neil R. Loneragan. Seine nets and beam trawls compared by day and night for sampling fish and crustaceans in shallow seagrass habitat. Fisheries Research, 64(2–3):185–196, November 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001097.

Gifford:2007:STW

[GCLA07] Andrew Gifford, Leonard J. V. Compagno, Marie Levine, and Alex Antoniou. Satellite tracking of whale sharks using tethered tags. Fisheries Research, 84(1):17–24, March 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606003961.

Grixti:2008:EPR

[GCMJ08] Daniel Grixti, Simon D. Conron, Alexander Morison, and Paul L. Jones. Estimating post-release survival and the influential factors for recreationally caught black bream (Acanthopagrus butcheri) in the Glenelg River, south eastern Australia. Fisheries Research, 92(2–3):303–313, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608000507.

Gajardo:2002:GVP

[GCN02] Gonzalo Gajardo, Juan M. Cancino, and Jorge M. Navarro. Genetic variation and population structure in the marine snail Chorus giganteus (Gastropod: Muricidae), an overexploited endemic resource from Chile. Fisheries Research, 55(1–3):329–333, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100282X.

Garla:2006:HJC

[GCS^{+}06] R. C. Garla, D. D. Chapman, M. S. Shivji, B. M. Wetherbee, and A. F. Amorim. Habitat of juvenile Caribbean reef sharks, Carcharhinus perezi, at two oceanic insular marine protected areas in the southwestern Atlantic Ocean: Fernando de Noronha Archipelago and Atol das Rocas, Brazil. Fisheries Research, 81(2–3):236–241, November 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002633.

Gutierrez:2003:DNS

[GD03] Nicolás Gutiérrez and Omar Defeo. Development of a new scallop Zygochlamys patagonica fishery in Uruguay: latitudinal and bathymetric patterns in biomass and population structure. Fisheries Research, 62(1):21–36, April 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002473.

Garcia:2007:DFD

[GDAM07] Camilo B. García, Luis O. Duarte, Jairo Altamar, and Luis M. Manjarrés. Demersal fish density in the upwelling ecosystem off Colombia, Caribbean Sea: Historic outlook. Fisheries Research, 85(1–2):68–73, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000021.

Godlewska:2009:RBS

[GDDJ09] Małgorzata Godlewska, Bronisław Długoszewski, Lech Doroszczyk, and Adam Jóźwik. The relationship between sampling intensity and sampling error — empirical results from acoustic surveys in Polish vendace lakes. Fisheries Research, 96(1):17–22, February 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002762.

Grace:2007:LSU

[GDFH07] Mark A. Grace, William B. Driggers III, Jack Forrester, and Nicholas Hopkins. Landing sling use increases data accuracy and sampling opportunities for large marine fishes captured during fishery-independent longline surveys. Fisheries Research, 88(1–3):146–149, December 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001907.

Grote:2007:ELH

[GEH^{+}07] B. Grote, W. Ekau, W. Hagen, J. A. Huggett, and H. M. Verheye. Early life-history strategy of Cape hake in the Benguela upwelling region. Fisheries Research, 86(2–3):179–187, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001245.

Gertseva:2009:PDL

[Ger09] Vladlena V. Gertseva. The population dynamics of the longnose skate, Raja rhina, in the northeast Pacific Ocean. Fisheries Research, 95(2–3):146–153, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002348.

Gutierrez-Estrada:2007:MCF

[GESY^{+}07] Juan Carlos Gutiérrez-Estrada, Claudio Silva, Eleuterio Yáñez, Nibaldo Rodríguez, and Inmaculada Pulido-Calvo. Monthly catch forecasting of anchovy Engraulis ringens in the north area of Chile: Non-linear univariate approach. Fisheries Research, 86(2–3):188–200, September 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001269.

Gutierrez-Estrada:2009:PSS

[GEYPC^{+}09] Juan Carlos Gutiérrez-Estrada, Eleuterio Yáñez, Inmaculada Pulido-Calvo, Claudio Silva, Francisco Plaza, and Cinthya Bórquez. Pacific sardine (Sardinops sagax, Jenyns 1842) landings prediction. A neural network ecosystemic approach. Fisheries Research, 100(2):116–125, October 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609001805.

Graham:2006:SFN

[GF06] N. Graham and R. J. Fryer. Separation of fish from Nephrops norvegicus into a two-tier cod-end using a selection grid. Fisheries Research, 82(1–3):111–118, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606002943.

Godoy:2003:RBR

[GFL03] Hallvard Godøy, Dag Furevik, and Svein Løkkeborg. Reduced bycatch of red king crab (Paralithodes camtschaticus) in the gillnet fishery for cod (Gadus morhua) in northern Norway. Fisheries Research, 62(3):377–384, June 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002813.

Gaughan:2002:FDA

[GFM02] Daniel J. Gaughan, Warrick J. Fletcher, and John P. McKinlay. Functionally distinct adult assemblages within a single breeding stock of the sardine, Sardinops sagax: management units within a management unit. Fisheries Research, 59(1–2):217–231, December 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601004118.

Gardner:2006:SER

[GFMO06] Caleb Gardner, Stewart Frusher, David Mills, and Megan Oliver. Simultaneous enhancement of rock lobster fisheries and provision of puerulus for aquaculture. Fisheries Research, 80(1):122–128, August 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001275.

Godoy:2003:UMR

[GFS03] Hallvard Godøy, Dag M. Furevik, and Stian Stiansen. Unaccounted mortality of red king crab (Paralithodes camtschaticus) in deliberately lost pots off Northern Norway. Fisheries Research, 64(2–3):171–177, November 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002169.

Gabr:2007:BAU

[GFSM07] Mohamed Gabr, Y. Fujimori, S. Shimizu, and T. Miura. Behaviour analysis of undersized fish escaping through square meshes and separating grids in simulated trawling experiment. Fisheries Research, 85(1–2):112–121, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000148.

Griffiths:2006:PDF

[GFvdV06] S. P. Griffiths, G. C. Fry, and T. D. van der Velde. Population dynamics and fishery benefits of a large legal size of a pelagic sportfish, the talang queenfish, Scomberoides commersonnianus, in northern Australia. Fisheries Research, 82(1–3):74–86, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002980.

Gristina:2004:PTR

[GG04] Michele Gristina and Monica Gagliano. Performance of traditional rush and modern plastic traps on the capture of Palinurus elephas (Fabricius, 1787) in laboratory tanks. Fisheries Research, 67(2):235–239, April 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002893.

Gertseva:2006:CMF

[GG06] Vladlena V. Gertseva and Vladimir I. Gertsev. A conceptual model of fish functional relationships in marine ecosystems and its application for fisheries stock assessment. Fisheries Research, 81(1):9–14, October 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600227X.

Gazo:2008:PDB

[GGA08] Manel Gazo, Joan Gonzalvo, and Alex Aguilar. Pingers as deterrents of bottlenose dolphins interacting with trammel nets. Fisheries Research, 92(1):70–75, July 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003669.

Garza-Gil:2003:EHF

[GGVLSR03] M. Dolores Garza-Gil, Manuel M. Varela-Lafuente, and Juan C. Surís-Regueiro. European hake fishery bioeconomic management (southern stock) applying an effort tax. Fisheries Research, 60(2–3):199–206, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001753.

Gimenez-Hurtado:2005:HBF

[GHCPPLC^{+}05] Enrique Giménez-Hurtado, Raúl Coyula-Pérez-Puelles, Salvador E. Lluch-Cota, Abel A. González-Yañez, Víctor Moreno-García, and Rogerio Burgos de-la Rosa. Historical biomass, fishing mortality, and recruitment trends of the Campeche Bank red grouper (Epinephelus morio). Fisheries Research, 71(3):267–277, March 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002280.

Gedamke:2008:TMR

[GHDM08] Todd Gedamke, John M. Hoenig, William D. DuPaul, and John A. Musick. Total mortality rates of the barndoor skate, Dipturus laevis, from the Gulf of Maine and Georges Bank, United States, 1963–2005. Fisheries Research, 89(1):17–25, January 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002068.

Gray:2008:NCW

[GHDS08] Tim Gray, Jenny Hatchard, Tim Daw, and Selina Stead. New cod war of words: ‘cod is god’ versus ‘sod the cod’ — two opposed discourses on the North Sea Cod Recovery Programme. Fisheries Research, 93(1–2):1–7, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608001331.

Gaertner:2004:TAM

[GHM04] Daniel Gaertner, Jean-Pierre Hallier, and Mark N. Maunder. A tag-attrition model as a means to estimate the efficiency of two types of tags used in tropical tuna fisheries. Fisheries Research, 69(2):171–180, September 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001353.

Gallagher:2004:CYP

[GHS04] Charmaine Marie Gallagher, Robert W. Hannah, and Gilbert Sylvia. A comparison of yield per recruit and revenue per recruit models for the Oregon ocean shrimp, Pandalus jordani, fishery. Fisheries Research, 66(1):71–84, January 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603001474.

Gerdeaux:2009:COV

[GJ09] Daniel Gerdeaux and Muhammad Yamin Janjua. Contribution of obligatory and voluntary fisheries statistics to the knowledge of whitefish population in Lake Annecy (France). Fisheries Research, 96(1):6–10, February 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002749.

Gray:2005:SSG

[GJBY05] Charles A. Gray, Daniel D. Johnson, Matt K. Broadhurst, and Damian J. Young. Seasonal, spatial and gear-related influences on relationships between retained and discarded catches in a multi-species gillnet fishery.

Fisheries Research, 75(1–3):56–72, September 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001657.

Gao:2004:SIC

[GJSW04] Yongwen Gao, Steve H. Joner, Russell A. Svec, and Ken L. Weinberg. Stable isotopic comparison in otoliths of juvenile sablefish (Anoplopoma fimbria) from waters off the Washington and Oregon coast. Fisheries Research, 68(1–3):351– 360, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603003448.

Graham:2001:SMP

[GK01] N. Graham and R. J. Kynoch. Square mesh panels in demersal trawls: some data on haddock selectivity in relation to mesh size and position. Fisheries Research, 49(3):207–218, January 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600002113.

Gray:2003:CCC

[GK03] Charles A. Gray and Steven J. Kennelly. Catch characteristics of the commercial beach-seine fisheries in two Australian barrier estuaries. Fisheries Research, 63(3): 405–422, September 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783603000766.

Graham:2003:SMP

[GKF03] N. Graham, R. J. Kynoch, and R. J. Fryer. Square mesh panels in demersal trawls: further data relating haddock and whiting selectivity to panel position. Fisheries Research, 62(3):361–375, June 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002795.

Gray:2001:RDC

[GKH^{+}01] C. A. Gray, S. J. Kennelly, K. E. Hodgson, C. J. T. Ashby, and M. L. Beatson. Retained and discarded catches from commercial beach-seining in Botany Bay, Australia. Fisheries Research, 50(3):205–219, March 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002289.

Gray:2003:LLB

[GKH03] Charles A. Gray, Steven J. Kennelly, and Kate E. Hodgson. Low levels of bycatch from estuarine prawn seining in New South Wales, Australia. Fisheries Research, 64(1):37–54, October 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603001851.

Gaumiga:2007:GRB

[GKUO07] R. Gaumiga, G. Karlsons, D. Uzars, and H. Ojaveer. Gulf of Riga (Baltic Sea) fisheries in the late 17th century. Fisheries Research, 87(2–3):120–125, November 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000914.

Georgakarakos:2006:TSA

[GKV06] S. Georgakarakos, D. Koutsoubas, and V. Valavanis. Time series analysis and forecasting techniques applied on loliginid and ommastrephid landings in Greek waters. Fisheries Research, 78(1):55–71, April 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003292.

Grimaldo:2005:CGN

[GL05] E. Grimaldo and R. B. Larsen. The cosmos grid: a new design for reducing by-catch in the Nordic shrimp fishery. Fisheries Research, 76(2):187–197, November 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001773.

Guillemot:2009:CMI

[GLCC09] Nicolas Guillemot, Marc Léopold, Marion Cuif, and Pascale Chabanet. Characterization and management of informal fisheries confronted with socio-economic changes in New Caledonia (South Pacific). Fisheries Research, 98(1–3):51– 61, June 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609000782.

Grimaldo:2007:EWA

[GLH07] E. Grimaldo, R. B. Larsen, and R. Holst. Exit Windows as an alternative selective system for the Barents Sea demersal fishery for cod and haddock. Fisheries Research, 85(3):295– 305, July 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607000732.

Gray:2000:UTN

[GLK00] C. A. Gray, R. B. Larsen, and S. J. Kennelly. Use of transparent netting to improve size selectivity and reduce bycatch in fish seine nets. Fisheries Research, 45(2):155–166, March 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699001113.

Grimaldo:2009:SEP

[GLS^{+}09] E. Grimaldo, R. B. Larsen, M. Sistiaga, N. Madsen, and M. Breen. Selectivity and escape percentages during three phases of the towing process for codends fitted with different selection systems. Fisheries Research, 95(2–3):198–205, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002634.

Godo:2000:MBN

[GM00] Olav Rune Godø and Kathrine Michalsen. Migratory behaviour of north-east Arctic cod, studied by use of data storage tags. Fisheries Research, 48(2):127–140, September 1, 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600001776.

Gerritsen:2006:VAM

[GM06] H. D. Gerritsen and D. McGrath. Variability in the assignment of maturity stages of plaice (Pleuronectes platessa L.) and whiting (Merlangius merlangus L.) using macroscopic maturity criteria. Fisheries Research, 77(1):72–77, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605002754.

Gerritsen:2007:SDL

[GM07] H. D. Gerritsen and D. McGrath. Significant differences in the length–weight relationships of neighbouring stocks can result in biased biomass estimates: Examples of haddock (Melanogrammus aeglefinus, L.) and whiting (Merlangius merlangus, L.). Fisheries Research, 85(1–2):106–111, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607000136.

Gomes:2002:FYR

[GMA02] L. C. Gomes, L. E. Miranda, and A. A. Agostinho. Fishery yield relative to chlorophyll a in reservoirs of the Upper Paraná River, Brazil. Fisheries Research, 55(1–3):335–340, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601002788.

Gomez:2006:ALL

[GML06] Edgardo D. Gomez and S. Suzanne Mingoa-Licuanan. Achievements and lessons learned in restocking giant clams in the Philippines. Fisheries Research, 80(1):46–52, August 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606001196.

Gurney:2005:DAG

[GMP05] L. J. Gurney, C. Mundy, and M. C. Porteus. Determining age and growth of abalone using stable oxygen isotopes: a tool for fisheries management. Fisheries Research, 72(2–3):353–360, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604003054.

Goss:2001:ISS

[GMR01] Catherine Goss, David Middleton, and Paul Rodhouse. Investigations of squid stocks using acoustic survey methods. Fisheries Research, 54(1):111–121, December 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003757. See erratum [GMR02].

Goss:2002:EIS

[GMR02] Catherine Goss, David Middleton, and Paul Rodhouse. Erratum to “Investigations of squid stocks using acoustic survey methods”: [Fish. Res. 54 (2001) 111–121]. Fisheries Research, 56(2):223, May 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000309. See [GMR01].

Groeneveld:2006:OMP

[GMS06] J. C. Groeneveld, G. Maharaj, and C. D. Smith. Octopus magnificus predation and bycatch in the trap fishery for spiny lobsters Palinurus gilchristi off South Africa. Fisheries Research, 79(1–2):90–96, June 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000385.

Gordoa:2000:MVC

[GMV00] Ana Gordoa, Mercedes Masó, and Lizette Voges. Monthly variability in the catchability of Namibian hake and its relationship with environmental seasonality. Fisheries Research, 48(2):185–195, September 1, 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001600. See erratum [CAJ00b].

Gordoa:2001:EMV

[GMV01] Ana Gordoa, Mercedes Masó, and Lizette Voges. Erratum to “Monthly variability in the catchability of Namibian hake and its relationship with environmental seasonality” [Fish. Res. 48 (2000) 185–195]. Fisheries Research, 51(1): 101, April 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600002691.  See

[GMV00].

Gonzalez:2009:RSA

[GMY^{+}09] Enrique Blanco Gonzalez, Tomoya Murakami, Takashi Yoneji, Kazuya Nagasawa, and Tetsuya Umino. Reduction in size-at-age of black sea bream (Acanthopagrus schlegelii) following intensive releases of cultured juveniles in Hiroshima Bay, Japan. Fisheries Research, 99(2):130–133, August 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609001052.

Gonzalez:2007:ASH

[GNC07] Raúl A. González, Maite A. Narvarte, and Guillermo M. Caille. An assessment of the sustainability of the hake Merluccius hubbsi artisanal fishery in San Matías Gulf, Patagonia, Argentina. Fisheries Research, 87(1):58–67, October 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360700149X.

Gleiss:2009:NPT

[GNL^{+}09] Adrian C. Gleiss, Brad Norman, Nikolai Liebsch, Clive Francis, and Rory P. Wilson. A new prospect for tagging large free-swimming sharks with motion-sensitive dataloggers. Fisheries Research, 97(1–2):11–16, April 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003986.

Gonzalez-Olivares:2009:DAP

[GOAARJRP09] Eduardo González-Olivares, Adriana Aránguiz-Acuña, Rodrigo Ramos-Jiliberto, and Alejandro Rojas-Palma. Demographical analysis of the pink ling Genypterus blacodes (Schneider 1801) in the austral demersal fishery: a matrix approach evaluating harvest and non-harvest states. Fisheries Research, 96(2–3):216–222, March 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003792.

Gobert:2000:CAM

[Gob00] Bertrand Gobert. Comparative assessment of multispecies reef fish resources in the Lesser Antilles. Fisheries Research, 44(3):247–260, January 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000855.

Graham:2004:SMD

[GOF^{+}04] N. Graham, F. G. O’Neill, R. J. Fryer, R. D. Galbraith, and A. Myklebust. Selectivity of a 120 mm diamond codend and the effect of inserting a rigid grid or a square mesh panel. Fisheries Research, 67(2):151–161, April 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360300287X.

Gonzalez:2006:AGA

[Gon06] Raúl A. C. González. Age and growth of the Argentine sandperch Pseudopercis semifasciata (Cuvier, 1829) Berg, 1899 in the San Matías Gulf (Patagonia, Argentina). Fisheries Research, 79(1–2):120–128, June 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000695.

Gordon:2001:DWF

[Gor01] John D. M. Gordon. Deep-water fish and fisheries: introduction. Fisheries Research, 51(2–3):105–111, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002387.

Gonzalez:2003:WMC

[GPG^{+}03] Angel F. González, Santiago Pascual, Camino Gestal, Elvira Abollo, and Angel Guerra. What makes a cephalopod a suitable host for parasite? The case of Galician waters. Fisheries Research, 60(1):177–183, January 30, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000590.

Gutierrez:2004:EHS

[GPI04] Jorge L. Gutiérrez, M. Gabriela Palomo, and Oscar O. Iribarne. Environmental heterogeneity and species responses to fishing disturbance: are the effects of clam harvesting spatially consistent? Fisheries Research, 67(1):55–70, March 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603002704.

Goni:2003:DCM

[GQR03] Raquel Goñi, Antoni Quetglas, and Olga Reñones. Differential catchability of male and female European spiny lobster Palinurus elephas (Fabricius, 1787) in traps and trammelnets. Fisheries Research, 65(1–3):295–307, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002509.

Gordoa:2000:GPF

[GR00] Ana Gordoa and Balbina Molıánd Nuria Raventós. Growth performance of four wrasse species on the north-western Mediterranean coast. Fisheries Research, 45(1):43–50, February 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000946.

Gauthier:2002:HEH

[GR02] Stéphane Gauthier and George A. Rose. An hypothesis on endogenous hydrostasis in Atlantic redfish (Sebastes spp.). Fisheries Research, 58(2):227–230, November 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100385X.

Graham:2007:ASG

[GR07] Rachel T. Graham and Callum M. Roberts. Assessing the size, growth rate and structure of a seasonal population of whale sharks (Rhincodon typus Smith 1828) using conventional tagging and photo identification. Fisheries Research, 84(1):71–80, March 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004024.

Graham:2002:BRS

[Gra02a] Norman Graham. Book review: Solving By-catch — Considerations for Today and Tomorrow: University of Alaska, Sea Grant College Program, Report No. 96-03, ISBN 1-56612-038-1, $25 (USA and Canada), $40 (overseas). Fisheries Research, 57(1):100–102, July 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000024.

Gray:2002:MID

[Gra02b] Charles A. Gray. Management implications of discarding in an estuarine multi-species gill net fishery. Fisheries Research, 56(2):177–192, May 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003198.

Graham:2003:CRB

[Gra03] Norman Graham. By-catch reduction in the brown shrimp, Crangon crangon, fisheries using a rigid separation Nordmøre grid (grate). Fisheries Research, 59(3):393–407, January 3, 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602000152.

Gundersen:2001:FGH

[GRB01] Agnes Christine Gundersen, Jan Erich Rønneberg, and Jesper Boje. Fecundity of Greenland halibut (Reinhardtius hippoglossoides Walbaum) in East Greenland waters. Fisheries Research, 51(2–3):229–236, May 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100248X.

Gray:2009:SSV

[GRC^{+}09] C. A. Gray, D. Rotherham, M. G. Chapman, A. J. Underwood, and D. D. Johnson. Spatial scales of variation of assemblages of fish in coastal lakes sampled with multi-mesh gillnets: Implications for designing research surveys. Fisheries Research, 96(1):58–63, February 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002816.

Garcia-Rodriguez:2006:CAC

[GRFE06] Mariano García-Rodríguez, Ángel M. Fernández, and Antonio Esteban. Characterisation, analysis and catch rates of the small-scale fisheries of the Alicante Gulf (SE Spain) over a 10 years time series. Fisheries Research, 77(2):226–238, February 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605002808.

Gribble:2003:GPM

[Gri03] Neil A. Gribble. GBR-prawn: modelling ecosystem impacts of changes in fisheries management of the commercial prawn (shrimp) trawl fishery in the far northern Great Barrier Reef. Fisheries Research, 65(1–3):493–506, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603002649.

Grimaldo:2006:EGA

[Gri06] Eduardo Grimaldo. The effects of grid angle on a modified Nordmøre-grid in the Nordic shrimp fishery. Fisheries Research, 77(1):53–59, January 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360500281X.

Groeneveld:2000:SAE

[Gro00] Johan C. Groeneveld. Stock assessment, ecology and economics as criteria for choosing between trap and trawl fisheries for spiny lobster Palinurus delagoae. Fisheries Research, 48(2):141–155, September 1, 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001788.

Garcia-Rodriguez:2005:BGA

[GRPLE05] Mariano García-Rodríguez, Pilar Pereda, Jorge Landa, and Antonio Esteban. On the biology and growth of the anglerfish Lophius budegassa Spinola, 1807 in the Spanish Mediterranean: a preliminary approach. Fisheries Research, 71(2):197–208, February 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002115.

Garces:2006:SSD

[GSA^{+}06] L. R. Garces, I. Stobutzki, M. Alias, W. Campos, N. Koongchai, L. Lachica-Alino, G. Mustafa, S. Nurhakim, M. Srinath, and G. Silvestre. Spatial structure of demersal fish assemblages in South and Southeast Asia and implications for fisheries management. Fisheries Research, 78 (2–3):143–157, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606000592.

Goncalves:2007:DET

[GSH^{+}07] J. M. S. Gonçalves, K. I. Stergiou, J. A. Hernando, E. Puente, D. K. Moutopoulos, L. Arregi, M. C. Soriguer, C. Vilas, R. Coelho, and K. Erzini. Discards from experimental trammel nets in southern European small-scale fisheries. Fisheries Research, 88(1–3):5–14, December 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001427.

Grimaldo:2008:ECS

[GSL08] E. Grimaldo, M. Sistiaga, and R. B. Larsen. Evaluation of codends with sorting grids, exit windows, and diamond meshes: Size selection and fish behaviour. Fisheries Research, 91(2–3):271–280, June 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003499.

Ganias:2004:POD

[GSMT04] Konstantinos Ganias, Stylianos Somarakis, Athanassios Machias, and Athanasios Theodorou. Pattern of oocyte development and batch fecundity in the Mediterranean sardine. Fisheries Research, 67(1):13–23, March 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002881.

Garces:2006:RDM

[GSS^{+}06] L. R. Garces, G. T. Silvestre, I. Stobutzki, F. C. Gayanilo, F. Valdez, M. Saupi, T. Boonvanich, M. Roongratri, P. Thouc, Purwanto, I. Haroon, K. N. Kurup, M. Srinath, H. A. B. Rodrigo, M. D. Santos, F. S. B. Torres, M. K. Tan, and D. Pauly. A regional database management system — the fisheries resource information system and tools (FiRST): Its design, utility and future directions. Fisheries Research, 78(2–3):119–129, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000579.

Gillson:2009:EGF

[GSS09] Jonathan Gillson, James Scandol, and Iain Suthers. Estuarine gillnet fishery catch rates decline during drought in eastern Australia. Fisheries Research, 99(1):26–37, July 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609001003.

Graynoth:2004:GJE

[GT04] Eric Graynoth and Mark Taylor. Growth of juvenile eels (Anguilla spp.) in lowland streams in New Zealand. Fisheries Research, 66(1):95–106, January 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001450.

Guidetti:2004:EES

[GTB04] Paolo Guidetti, Antonio Terlizzi, and Ferdinando Boero. Effects of the edible sea urchin, Paracentrotus lividus, fishery along the Apulian rocky coast (SE Italy, Mediterranean Sea). Fisheries Research, 66(2–3):287–297, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002066.

Guclusoy:2008:DMS

[Güç08] Harun Güçlüsoy. Damage by monk seals to gear of the artisanal fishery in the Foça Monk Seal Pilot Conservation Area, Turkey. Fisheries Research, 90(1–3):70–77, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607002342.

Guerra:2005:PFJ

[Gue05] Angel Guerra. Pierre Fréon and Jeremy Mendoza (Eds.), The sardine (Sardinella aurita). Its environment and exploitation in Eastern Venezuela, IRD Editions, Institute de Recherche pour le Dévelopment, Collection Colloques et Séminaires, Paris, 2003, 549 pp., soft cover, ISBN: 2-7099-1517-0. Fisheries Research, 73(1–2):265–267, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000421.

Gustavson:2002:EPA

[Gus02] Kent Gustavson. Economic production from the artisanal fisheries of Jamaica. Fisheries Research, 57(2):103–115, August 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601003460.

Gribble:2007:FAD

[GWB07] N. A. Gribble, T. J. Wassenberg, and C. Burridge. Factors affecting the distribution of commercially exploited penaeid prawns (shrimp) (Decapod: Penaeidae) across the northern Great Barrier Reef, Australia. Fisheries Research, 85(1–2):174–185, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000379.

Groger:2007:PDP

[GWR07] Joachim Paul Gröger, Helmut Winkler, and Rodney Alan Rountree. Population dynamics of pikeperch (Sander lucioperca) and its linkage to fishery driven and climatic influences in a southern Baltic lagoon of the Darss–Zingst Bodden Chain. Fisheries Research, 84(2):189–201, April 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606003808.

Gonzalez-Yanez:2006:MDD

[GYMdL^{+}06] Akim A. González-Yáñez, Rafael Puga Millán, María Estela de León, Liset Cruz-Font, and Matthias Wolff. Modified DeLury depletion model applied to spiny lobster, Panulirus argus (Latreille, 1804) stock, in the southwest of the Cuban Shelf. Fisheries Research, 79(1–2):155–161, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606000737.

Halliday:2002:CSS

[Hal02] R. G. Halliday. A comparison of size selection of Atlantic cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) by bottom longlines and otter trawls. Fisheries Research, 57(1):63–73, July 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003368.

Hotos:2000:RBL

[HAO00] G. N. Hotos, D. Avramidou, and I. Ondrias. Reproduction biology of Liza aurata (Risso, 1810), (Pisces Mugilidae) in the lagoon of Klisova (Messolonghi, W. Greece). Fisheries Research, 47(1):57–67, June 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001289.

Hastie:2006:DME

[Has06] Lee C. Hastie. Determination of mortality in exploited freshwater pearl mussel (Margaritifera margaritifera) populations. Fisheries Research, 80(2–3):305–311, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606001354.

Hatfield:2000:DSL

[Hat00] Emma M. C. Hatfield. Do some like it hot? Temperature as a possible determinant of variability in the growth of the Patagonian squid, Loligo gahi (Cephalopoda: Loliginidae). Fisheries Research, 47(1):27–40, June 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001277.

Hoines:2001:DWS

[HB01] Å. S. Høines and O. A. Bergstad. Density of wintering sand eel in the sand recorded by grab catches. Fisheries Research, 49(3):295–301, January 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002137.

Haupt:2006:EEA

[HBBG06] Philip Haupt, Stephen L. Brouwer, George M. Branch, and Gerd Gäde. Effects of exposure to air on the escape behaviour and haemolymph chemistry of the South African Cape lobster, Jasus lalandii. Fisheries Research, 81(2–3):210–218, November 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002761.

Heales:2003:STC

[HBJ03] D. S. Heales, D. T. Brewer, and P. N. Jones. Subsampling trawl catches from vessels using seawater hoppers: are catch composition estimates biased? Fisheries Research, 63(1):113–120, July 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002783.

Heales:2007:DDC

[HBKJ07] D. S. Heales, D. T. Brewer, P. M. Kuhnert, and P. N. Jones. Detecting declines in catch rates of diverse trawl bycatch species, and implications for monitoring. Fisheries Research, 84(2):153–161, April 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003705.

Helle:2000:ERI

[HBM^{+}00] Kristin Helle, Bjarte Bogstad, C. Tara Marshall, Kathrine Michalsen, Geir Ottersen, and Michael Pennington. An evaluation of recruitment indices for arcto-Norwegian cod (Gadus morhua L.). Fisheries Research, 48(1):55–67, August 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001193.

Heales:2000:SMS

[HBW00] D. S. Heales, D. T. Brewer, and Y.-G. Wang. Subsampling multi-species trawl catches from tropical northern Australia: Does it matter which part of the catch is sampled? Fisheries Research, 48(2):117–126, September 1, 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360000182X.

Halliday:2000:SSS

[HC00] R. G. Halliday and C. G. Cooper. Size selection of silver hake (Merluccius bilinearis) by otter trawls with square and diamond mesh codends of 55–60 mm mesh size. Fisheries Research, 49(1):77–84, November 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001880.

Hall:2007:LKA

[HC07] G. Brent Hall and C. H. Close. Local knowledge assessment for a small-scale fishery using geographic information systems. Fisheries Research, 83(1):11–22, January 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600289X.

Hobday:2009:TPH

[HC09] Alistair J. Hobday and Geoff Campbell. Topographic preferences and habitat partitioning by pelagic fishes off southern Western Australia. Fisheries Research, 95(2–3):332– 340, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360800324X.

Henry:2009:CFL

[HCH09] Neil A. Henry, Steven J. Cooke, and Kyle C. Hanson. Consequences of fishing lure retention on the behaviour and physiology of free-swimming smallmouth bass during the reproductive period. Fisheries Research, 100(2):178–182, October 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609001854.

Hendry:2003:MHR

[HCHO^{+}03] K. Hendry, D. Cragg-Hine, M. O’Grady, H. Sambrook, and A. Stephen. Management of habitat for rehabilitation and enhancement of salmonid stocks. Fisheries Research, 62(2):171–192, May 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001613.

Hartill:2005:EBS

[HCM05] B. W. Hartill, M. Cryer, and M. A. Morrison. Estimates of biomass, sustainable yield, and harvest: neither necessary nor sufficient for the management of noncommercial urban intertidal shellfish fisheries. Fisheries Research, 71(2):209–222, February 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002103.

Harvey:2003:APU

[HCS^{+}03] Euan Harvey, Mike Cappo, Mark Shortis, Stuart Robson, Jeff Buchanan, and Peter Speare. The accuracy and precision of underwater measurements of length and maximum body depth of southern bluefin tuna (Thunnus maccoyii) with a stereo–video camera system. Fisheries Research, 63(3):315–326, September 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000808.

Hewitt:2000:UAS

[HD00] Roger P. Hewitt and David A. Demer. The use of acoustic sampling to estimate the dispersion and abundance of euphausiids, with an emphasis on Antarctic krill, Euphausia superba. Fisheries Research, 47(2–3):215–229, July 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001715.

Haimovici:2007:CME

[HdS07] Manuel Haimovici and Antônio Olinto Ávila da Silva. A comparison of methods for estimating relative abundances in bottom longline fishery surveys. Fisheries Research, 85 (1–2):226–231, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783607000124.

He:2003:SBW

[He03] Pingguo He. Swimming behaviour of winter flounder (Pleuronectes americanus) on natural fishing grounds as observed by an underwater video camera. Fisheries Research, 60(2–3):507–514, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000863.

He:2005:CBP

[He05] Pingguo He. Characteristics of bycatch of porcupine crabs, Neolithodes grimaldii (Milne-Edwards and Bouvier, 1894) from deepwater turbot gillnets in the northwest Atlantic. Fisheries Research, 74(1–3):35–43, August 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001232.

He:2006:EHH

[He06] Pingguo He. Effect of the headline height of gillnets on species selectivity in the Gulf of Maine. Fisheries Research, 78(2–3):252–256, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000269.

He:2007:SLM

[He07] Pingguo He. Selectivity of large mesh trawl codends in the Gulf of Maine: I. Comparison of square and diamond mesh. Fisheries Research, 83(1):44–59, January 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003067.

Hazin:2008:ASD

[HE08a] Humberto Hazin and Karim Erzini. Assessing swordfish distribution in the South Atlantic from spatial predictions. Fisheries Research, 90(1–3):45–55, April 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002329.

He:2008:HCH

[He08b] Pingguo He. Haddock conservation, harvesting and management. Fisheries Research, 94(2):121–122, November 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608002695.

Henderson:2009:ORH

[Hen09] I. M. Henderson. Optimising recreational harvests of blue cod: the effects of catch-and-release mortality and size selectivity. Fisheries Research, 99(3):184–195, August 15, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609001465.

Herrmann:2005:ECSa

[Her05a] Bent Herrmann. Effect of catch size and shape on the selectivity of diamond mesh cod-ends: I. Model development. Fisheries Research, 71(1):1–13, January 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001924.

Herrmann:2005:ECSb

[Her05b] Bent Herrmann. Effect of catch size and shape on the selectivity of diamond mesh cod-ends: II. Theoretical study of haddock selection. Fisheries Research, 71(1):15–26, January 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001912.

Heen:2007:SEI

[HF07] Knut Heen and Ola Flaaten. Spatial employment impacts of fisheries management: a study of the Barents Sea and the Norwegian Sea fisheries. Fisheries Research, 85(1–2):74– 83, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607000033.

Hazin:2006:NRS

[HFB^{+}06] Fabio H. V. Hazin, Alessandra F. Fischer, Matt K. Broadhurst, Dráusio Veras, Paulo G. Oliveira, and George H. Burgess. Notes on the reproduction of Squalus megalops off northeastern Brazil. Fisheries Research, 79(3):251–257, July 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606001470.

Hart:2008:DVS

[HFBM08] Anthony M. Hart, Frank P. Fabris, Jamin Brown, and Dave Murphy. Digital video surveys of abalone (Haliotis sp.) stocks by commercial fishers in Western Australia. Fisheries Research, 93(3):305–314, September 28, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001768.

Herrmann:2007:SBI

[HFHO07] Bent Herrmann, Rikke P. Frandsen, René Holst, and Finbarr G. O’Neill. Simulation-based investigation of the paired-gear method in cod-end selectivity studies. Fisheries Research, 83(2–3):175–184, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003389.

Hemmingsson:2008:PTD

[HFL08] Malin Hemmingsson, Arne Fjälling, and Sven-Gunnar Lunneryd. The pontoon trap: Description and function of a seal-safe trap-net. Fisheries Research, 93(3):357–359, September 28, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608001914.

Harvey:2002:ERF

[HFS02] Euan Harvey, David Fletcher, and Mark Shortis. Estimation of reef fish length by divers and by stereovideo: a first comparison of the accuracy and precision in the field on living fish under operational conditions. Fisheries Research, 57(3):255–265, September 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003563.

Hareide:2001:DCR

[HG01] Nils-Roar Hareide and Greta Garnes. The distribution and catch rates of deep water fish along the Mid-Atlantic Ridge from 43 to 61^{o}N. Fisheries Research, 51(2–3):297–310, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601002533.

Hutchings:2006:RVL

[HGF06] Kenneth Hutchings, Marc H. Griffiths, and John G. Field. Regional variation in the life history of the canary drum Umbrina canariensis (Sciaenidae), in South African waters. Fisheries Research, 77(3):312–325, March 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003000.

Hernandez-Garcia:2002:ROV

[HGHLCH02] V. Hernández-Garcıá, J. L. Hernández-López, and J. J. Castro-Hdez. On the reproduction of Octopus vulgaris off the coast of the Canary Islands. Fisheries Research, 57(2):197–203, August 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003411.

Hutton:2001:CVN

[HGSP01] T. Hutton, M. H. Griffiths, U. R. Sumaila, and T. J. Pitcher. Cooperative versus non-cooperative management of shared linefish stocks in South Africa: an assessment of alternative management strategies for geelbek (Atractoscion aequidens). Fisheries Research, 51(1):53–68, April 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600002319.

Heales:2008:TPT

[HGW^{+}08] D. S. Heales, R. Gregor, J. Wakeford, Y.-G. Wang, J. Yarrow, and D. A. Milton. Tropical prawn trawl bycatch of fish and seasnakes reduced by Yarrow Fisheye Bycatch Reduction Device. Fisheries Research, 89(1):76–83, January 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607002172.

Hjerne:2001:CCC

[HH01] Olle Hjerne and Sture Hansson. Constant catch or constant harvest rate?: the Baltic Sea cod (Gadus morhua L.) fishery as a modelling example. Fisheries Research, 53(1):57–70, September 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002666.

Hendrickson:2006:ABC

[HH06] Lisa C. Hendrickson and Deborah R. Hart. An age-based cohort model for estimating the spawning mortality of semelparous cephalopods with an application to per-recruit calculations for the northern shortfin squid, Illex illecebrosus. Fisheries Research, 78(1):4–13, April 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003279.

Haakana:2008:EIF

[HH08] Helena Haakana and Hannu Huuskonen. Effects of intensive fishing on the perch population in a large oligotrophic lake in eastern Finland. Fisheries Research, 91(2–3):144– 150, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607003359.

Hjelm:2006:WUH

[HHC06] Joakim Hjelm, Malin Hultgren, and Massimiliano Cardinale. Water uptake in herring (Clupea harengus) and sprat (Sprattus sprattus) as a function of area, salinity and fat content. Fisheries Research, 81(1):94–99, October 2006.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002141.

Hazin:2005:ELS

[HHTE05] H. G. Hazin, F. H. V. Hazin, P. Travassos, and K. Erzini. Effect of light-sticks and electralume attractors on surface-longline catches of swordfish (Xiphias gladius, Linnaeus, 1959) in the southwest equatorial Atlantic. Fisheries Research, 72(2–3):271–277, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002516.

Hossain:2006:MIO

[HIRM06] Md. Monir Hossain, Md. Aminul Islam, Stephen Ridgway, and Takashi Matsuishi. Management of inland open water fisheries resources of Bangladesh: Issues and options. Fisheries Research, 77(3):275–284, March 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000105.

Hannah:2003:MHF

[HJ03] Robert W. Hannah and Stephen A. Jones. Measuring the height of the fishing line and its effect on shrimp catch and bycatch in an ocean shrimp (Pandalus jordani) trawl. Fisheries Research, 60(2–3):427–438, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001388.

Haputhantri:2006:PMP

[HJ06] S. S. K. Haputhantri and P. A. A. T. Jayawardane. Predictive models for penaeid shrimp abundance in the seas off Negombo and Hendala, Sri Lanka. Fisheries Research, 77(1):34–44, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002833.

Hannah:2007:EBR

[HJ07] Robert W. Hannah and Stephen A. Jones. Effectiveness of bycatch reduction devices (BRDs) in the ocean shrimp (Pandalus jordani) trawl fishery. Fisheries Research, 85(1–2):217–225, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000537.

Hsu:2007:STJ

[HJLL07] Hua-Hsun Hsu, Shoou-Jeng Joung, Yih-Yia Liao, and Kwang-Ming Liu. Satellite tracking of juvenile whale sharks, Rhincodon typus, in the Northwestern Pacific. Fisheries Research, 84(1):25–31, March 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003973.

Hudd:2002:BMB

[HK02] Richard Hudd and Jakob Kjellman. Bad matching between hatching and acidification: a pitfall for the burbot, Lota lota, off the river Kyrönjoki, Baltic Sea. Fisheries Research, 55(1–3):153–160, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003034.

Haggarty:2006:CIR

[HK06a] Dana R. Haggarty and Jacquelynne R. King. CPUE as an index of relative abundance for nearshore reef fishes. Fisheries Research, 81(1):89–93, October 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002189.

Hamasaki:2006:RKP

[HK06b] Katsuyuki Hamasaki and Shuichi Kitada. A review of kuruma prawn Penaeus japonicus stock enhancement in Japan. Fisheries Research, 80(1):80–90, August 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001238.

Herrmann:2009:PSM

[HKF^{+}09] Bent Herrmann, Ludvig A. Krag, Rikke P. Frandsen, Niels Madsen, Bo Lundgren, and Karl-Johan Stæhr. Prediction of selectivity from morphological conditions: Methodology and a case study on cod (Gadus morhua). Fisheries Research, 97(1–2):59–71, April 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000034.

Heath:2008:MMP

[HKG^{+}08] Michael R. Heath, Philip A. Kunzlik, Alejandro Gallego, Steven J. Holmes, and Peter J. Wright. A model of metapopulation dynamics for North Sea and West of Scotland cod — the dynamic consequences of natal fidelity. Fisheries Research, 93(1–2):92–116, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000854.

Hwang:2008:AGL

[HLH08] Sun-Do Hwang, Tae-Won Lee, and Sun-Wan Hwang. Age, growth and life history of gunnel, Pholis fangi, in the Yellow Sea. Fisheries Research, 93(1–2):72–76, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608000660.

Humborstad:2003:CGH

[HLHF03] Odd-Børre Humborstad, Svein Løkkeborg, Nils-Roar Hareide, and Dag Magne Furevik. Catches of Greenland halibut (Reinhardtius hippoglossoides) in deepwater ghostfishing gillnets on the Norwegian continental slope. Fisheries Research, 64(2–3):163–170, November 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002157.

Haas:2008:CST

[HLL^{+}08] Heather L. Haas, Erin LaCasella, Robin LeRoux, Henry Milliken, and Brett Hayward. Characteristics of sea turtles incidentally captured in the U.S. Atlantic sea scallop dredge fishery. Fisheries Research, 93(3):289–295, September 28, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608001598.

Holst:2000:FLD

[HM00] Jens Christian Holst and Angus McDonald. FISH-LIFT: a device for sampling live fish with trawls. Fisheries Research, 48(1):87–91, August 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001168.

Henderson:2007:SOS

[HMAOAS07] Aaron C. Henderson, Jennifer L. McIlwain, Hamed S. Al-Oufi, and Sulaiman Al-Sheili. The Sultanate of Oman shark fishery: Species composition, seasonality and diversity. Fisheries Research, 86(2–3):159–168, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360700121X.

Halden:2000:OZD

[HMB^{+}00] Norman M. Halden, Sergio R. Mejia, John A. Babaluk, James D. Reist, Allan H. Kristofferson, John L. Campbell, and William J. Teesdale. Oscillatory zinc distribution in Arctic char (Salvelinus alpinus) otoliths: the result of biology or environment? Fisheries Research, 46(1–3):289– 298, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600001545.

Hedger:2004:ASD

[HMH^{+}04] Richard Hedger, Eddie McKenzie, Mike Heath, Peter Wright, Beth Scott, Alejandro Gallego, and Jessica Andrews. Analysis of the spatial distributions of mature cod (Gadus morhua) and haddock (Melanogrammus aeglefinus) abundance in the North Sea (1980–1999) using generalised additive models. Fisheries Research, 70(1):17–25, November 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001523.

Harley:2000:EUF

[HMM00a] Shelton J. Harley, Russell B. Millar, and Brian H. McArdle. Estimating unaccounted fishing mortality using selectivity data: an application in the Hauraki Gulf snapper (Pagrus auratus) fishery in New Zealand. Fisheries Research, 45(2):167–178, March 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001095.

Harley:2000:EEC

[HMM00b] Shelton J. Harley, Russell B. Millar, and Brian H. McArdle. Examining the effects of changes in the minimum legal sizes used in the Hauraki Gulf snapper (Pagrus auratus) fishery in New Zealand. Fisheries Research, 45(2):179–187, March 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699001101.

Hebert:2001:EGF

[HMM^{+}01] M. Hébert, G. Miron, M. Moriyasu, R. Vienneau, and P. DeGrâce. Efficiency and ghost fishing of snow crab (Chionoecetes opilio) traps in the Gulf of St. Lawrence. Fisheries Research, 52(3):143–153, July 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002599.

Hoie:2009:LDT

[HMM^{+}09] Hans Høie, Richard S. Millner, Sophy McCully, Kjell H. Nedreaas, Graham M. Pilling, and Janicke Skadal. Latitudinal differences in the timing of otolith growth: a comparison between the Barents Sea and southern North Sea. Fisheries Research, 96(2–3):319–322, March 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003901.

Haedrich:2001:CEK

[HMO01] R. L. Haedrich, N. R. Merrett, and N. R. O’Dea. Can ecological knowledge catch up with deep-water fishing? a North Atlantic perspective. Fisheries Research, 51(2– 3):113–122, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783601002399.

Hiddink:2008:IFB

[HMR^{+}08] J. G. Hiddink, B. R. MacKenzie, A. Rijnsdorp, N. K. Dulvy, E. E. Nielsen, D. Bekkevold, M. Heino, P. Lorance, and H. Ojaveer. Importance of fish biodiversity for the management of fisheries and ecosystems. Fisheries Research, 90(1–3):6–8, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003414.

Howell:2008:RAA

[HMRG08] W. Huntting Howell, Michael Morin, Nathan Rennels, and David Goethel. Residency of adult Atlantic cod (Gadus morhua) in the western Gulf of Maine. Fisheries Research, 91(2–3):123–132, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003219.

Herrmann:2005:TSB

[HO05] Bent Herrmann and Finbarr G. O’Neill. Theoretical study of the between-haul variation of haddock selectivity in a diamond mesh cod-end. Fisheries Research, 74(1–3):243–252, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605000457.

Herrmann:2006:TSI

[HO06] Bent Herrmann and Finbarr G. O’Neill. Theoretical study of the influence of twine thickness on haddock selectivity in diamond mesh cod-ends. Fisheries Research, 80(2–3):221–229, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001494.

Holt:2001:CTS

[Hol01] Sidney Holt. A comment on Tore Schweder’s “Protecting whales by distorting uncertainty: non-precautionary mismanagement”. Fisheries Research, 52(3):227–230, July 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600002629.

Holt:2009:SBH

[Hol09] Sidney Holt. Sunken billions — but how many? Fisheries Research, 97(1–2):3–10, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000174.

Hoolihan:2006:AGI

[Hoo06] John P. Hoolihan. Age and growth of Indo–Pacific sailfish, Istiophorus platypterus, from the Arabian Gulf. Fisheries Research, 78(2–3):218–226, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000324.

Horppila:2000:ELF

[Hor00] Jukka Horppila. The effect of length frequency ranges on the back-calculated lengths of roach, Rutilus rutilus (L.). Fisheries Research, 45(1):21–29, February 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000995.

Horn:2002:AGP

[Hor02] P. L. Horn. Age and growth of Patagonian toothfish (Dissostichus eleginoides) and Antarctic toothfish (D. mawsoni) in waters from the New Zealand subantarctic to the Ross Sea, Antarctica. Fisheries Research, 56(3):275–287, June 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601003253.

Horbowy:2005:ABH

[Hor05] Jan Horbowy. Assessing Baltic herring stocks with a model that incorporates migration. Fisheries Research, 76(2):

266–276, November 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783605002018.

Howe:2002:SLQ

[How02] Jeffrey C. Howe. Standard length: not quite so standard. Fisheries Research, 56(1):1–7, April 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003125.

Howe:2003:BRA

[How03a] Jeffrey C. Howe. Book review: Artificial Reef Evaluation with Application to Natural Marine Habitats: William Seaman Jr. (Ed.); CRC Press, New York, NY, 2000, 246 pages, hardcover, ISBN 0-8493-9061-3 (US $84.95). Fisheries Research, 63(2):297–298, August 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001267.

Howe:2003:BRC

[How03b] Jeffrey C. Howe. Book review: Conservation of Freshwater Fishes: Options for the Future: M. J. Collares-Pereira, M. M. Coelho, I. G. Cowx (Eds.); Iowa State University Press, Ames, IA, 2002, 472 pages, hardcover, ISBN 0-85238-286-3, US $144.99. Fisheries Research, 64(1):87–88, October 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603002170.

Howe:2006:BRE

[How06] Jeffrey C. Howe. Book review: Early Stages of Atlantic Fishes: an Identification Guide for the Western Central North Atlantic, W. J. Richards. Two-volume set. Taylor & Francis Group, Boca Raton, FL (2006). 2640 pp., Hardcover, Price: £226, US $395, ISBN: 0-8493-1916-1. Fisheries Research, 80(2–3):353–354, September 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001640.

Howe:2000:TDM

[HP00] Jeffrey C. Howe and Blan A. Page. A technique for deploying multiprobe data loggers to measure shallow water parameters. Fisheries Research, 45(3):303–307, April 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001204.

Halliday:2002:RST

[HP02] R. G. Halliday and A. T. Pinhorn. A review of the scientific and technical bases for policies on the capture of small fish in North Atlantic groundfish fisheries. Fisheries Research, 57(3):211–222, September 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000796.

Helle:2004:SDC

[HP04] Kristin Helle and Michael Pennington. Survey design considerations for estimating the length composition of the commercial catch of some deep-water species in the northeast Atlantic. Fisheries Research, 70(1):55–60, November 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360400147X.

Halliday:2009:RFE

[HP09a] R. G. Halliday and A. T. Pinhorn. The roles of fishing and environmental change in the decline of Northwest Atlantic groundfish populations in the early 1990s. Fisheries Research, 97(3):163–182, May 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000423.

Holt:2009:ICI

[HP09b]  Carrie A. Holt and André E. Punt.  Incorporat-

ing climate information into rebuilding plans for overfished groundfish species of the U.S. west coast. Fisheries Research, 100(1):57–67, September 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000654.

Haltuch:2008:EAE

[HPD08] Melissa A. Haltuch, André E. Punt, and Martin W. Dorn. Evaluating alternative estimators of fishery management reference points. Fisheries Research, 94(3):290–303, December 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608000489.

Haltuch:2009:EEF

[HPD09] Melissa A. Haltuch, André E. Punt, and Martin W. Dorn. Evaluating the estimation of fishery management reference points in a variable environment. Fisheries Research, 100(1):42–56, September 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000642.

Herrmann:2006:TSE

[HPK06] Bent Herrmann, Daniel Priour, and Ludvig A. Krag. Theoretical study of the effect of round straps on the selectivity in a diamond mesh cod-end. Fisheries Research, 80(2–3):148–157, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001767.

Herrmann:2007:SBS

[HPK07] Bent Herrmann, Daniel Priour, and Ludvig A. Krag. Simulation-based study of the combined effect on codend size selection of turning meshes by 90^{o}and reducing the number of meshes in the circumference for round fish. Fisheries Research, 84(2):222–232, April 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003833.

Helser:2004:GLM

[HPM04] Thomas E. Helser, André E. Punt, and Richard D. Methot, Jr. A generalized linear mixed model analysis of a multi-vessel fishery resource survey. Fisheries Research, 70(2–3):251–264, December 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001705.

Holst:2009:SSM

[HR09] René Holst and Andrew Revill. A simple statistical method for catch comparison studies. Fisheries Research, 95(2–3):254–259, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003123.

Huse:2000:AIS

[HS00] Irene Huse and Aud Vold Soldal. An attempt to improve size selection in pelagic longline fisheries for haddock. Fisheries Research, 48(1):43–54, August 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360000117X.

Hernandez:2003:SDA

[HS03] Alvaro Hernandez and Juan C. Seijo. Spatial distribution analysis of red grouper (Epinephelus morio) fishery in Yucatan, Mexico. Fisheries Research, 63(1):135–141, July 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603000390.

Harrington:2006:PTC

[HSGF06] Julian J. Harrington, Jayson M. Semmens, Caleb Gardner, and Stewart D. Frusher. Predation of trap-caught southern rock lobsters, Jasus edwardsii (Hutton, 1875), in Tasmanian waters by the Maori octopus, Octopus maorum (Hutton, 1880): Spatial and temporal trends. Fisheries Research, 77(1):10–16, January 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002948.

Heikinheimo:2006:IMS

[HSSR06] Outi Heikinheimo, Jari Setälä, Kaija Saarni, and Jari Raitaniemi. Impacts of mesh-size regulation of gillnets on the pikeperch fisheries in the Archipelago Sea, Finland.

Fisheries Research, 77(2):192–199, February 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002936.

Holmes:2009:ISR

[HSW09] Bronwyn H. Holmes, Dirk Steinke, and Robert D. Ward. Identification of shark and ray fins using DNA barcoding. Fisheries Research, 95(2–3):280–288, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003172.

Hospido:2005:LCE

[HT05] Almudena Hospido and Peter Tyedmers. Life cycle environmental impacts of Spanish tuna fisheries. Fisheries Research, 76(2):174–186, November 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360500175X.

Herraiz:2009:NIL

[HTF^{+}09] I. González Herraiz, M. A. Torres, A. C. Fariña, J. Freire, and J. R. Cancelo. The NAO index and the long-term variability of Nephrops norvegicus population and fishery off West of Ireland. Fisheries Research, 98(1–3):1–7, June 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360900071X.

Hurley:2007:ERS

[Hur07] Keith L. Hurley. Effects of a reduced subsampling protocol on calculation of growth and body condition indices. Fisheries Research, 83(2–3):360–363, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600378X.

Hill:2000:PFD

[HW00] B. J. Hill and T. J. Wassenberg. The probable fate of discards from prawn trawlers fishing near coral reefs: a study in the northern Great Barrier Reef, Australia.

Fisheries Research, 48(3):277–286, October 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001855.

Halls:2006:RBM

[HWB06] A. S. Halls, R. L. Welcomme, and R. W. Burn. The relationship between multi-species catch and effort: Among fishery comparisons. Fisheries Research, 77(1):78–83, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605002699.

Holst:2002:ETT

[HWM02] R. Holst, D. Wileman, and N. Madsen. The effect of twine thickness on the size selectivity and fishing power of Baltic cod gill nets. Fisheries Research, 56(3):303–312, June 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003289.

Heupel:2009:EFT

[HWW^{+}09] M. R. Heupel, A. J. Williams, D. J. Welch, A. Ballagh, B. D. Mapstone, G. Carlos, C. Davies, and C. A. Simpfendorfer. Effects of fishing on tropical reef associated shark populations on the Great Barrier Reef. Fisheries Research, 95(2–3):350–361, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003317.

Hill:2006:DEP

[HZZG06] Megan S. Hill, Gayle B. Zydlewski, Joseph D. Zydlewski, and James M. Gasvoda. Development and evaluation of portable PIT tag detection units: PITpacks. Fisheries Research, 77(1):102–109, January 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002195.

Ishida:2009:MDJ

[IFH^{+}09] Yukimasa Ishida, Tetsuichiro Funamoto, Satoshi Honda, Keizou Yabuki, Hiroshi Nishida, and Chikako Watanabe. Management of declining Japanese sardine, chub mackerel and walleye pollock fisheries in Japan. Fisheries Research, 100(1):68–77, September 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360900068X.

Ihde:2008:EMY

[IHF08] T. F. Ihde, J. M. Hoenig, and S. D. Frusher. Evaluation of a multi-year index-removal abundance estimator, with application to a Tasmanian rock lobster fishery. Fisheries Research, 89(1):26–36, January 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002081.

Ingolfsson:2006:EGF

[IJ06] Ólafur Arnar Ingólfsson and Terje Jørgensen. Escapement of gadoid fish beneath a commercial bottom trawl: Relevance to the overall trawl selectivity. Fisheries Research, 79(3):303–312, July 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001111.

Imsland:2004:NDR

[IJD04] A. K. Imsland, Ó. D. B. Jónsdóttir, and A. K. Daníelsdóttir. Nuclear DNA RFLP variation among Atlantic cod in south and south-east Icelandic waters. Fisheries Research, 67 (2):227–233, April 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783603003011.

Irvine:2007:ICS

[IK07] Tennille R. Irvine and John K. Keesing. International collaboration in science, conservation and sustainable tourism of whale sharks. Fisheries Research, 84(1):1–3, March 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003936.

Ichii:2006:SAA

[IMOO06] Taro Ichii, Kedarnath Mahapatra, Hiroshi Okamura, and Yoshihiro Okada. Stock assessment of the autumn cohort of neon flying squid (Ommastrephes bartramii) in the North Pacific based on past large-scale high seas driftnet fishery data. Fisheries Research, 78(2–3):286–297, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600018X.

Iqbal:2006:AGJ

[IMSS06] Khan M. Iqbal, Yasuji Masuda, Hiroshi Suzuki, and Akihiko Shinomiya. Age and growth of the Japanese silver-biddy, Gerres equulus, in western Kyushu, Japan. Fisheries Research, 77(1):45–52, January 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002821.

Iqbal:2007:RBJ

[IOS07] Khan M. Iqbal, Jun Ohtomi, and Hiroshi Suzuki. Reproductive biology of the Japanese silver-biddy, Gerres equulus, in western Kyushu, Japan. Fisheries Research, 83(2–3):145–150, February 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003183.

Ishimura:2005:MFF

[IPH05] Gakushi Ishimura, André E. Punt, and Daniel D. Huppert. Management of fluctuating fish stocks: the case of Pacific whiting. Fisheries Research, 73(1–2):201–216, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604003066.

Ives:2007:BAN

[IS07] Matthew C. Ives and James P. Scandol. A Bayesian analysis of NSW eastern king prawn stocks (Melicertus plebejus) using multiple model structures. Fisheries Research, 84(3):314–327, May 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003924.

Ismen:2003:AGR

[Ism03] Ali Ismen. Age, growth, reproduction and food of common stingray (Dasyatis pastinaca L., 1758) in İskenderun Bay, the eastern Mediterranean. Fisheries Research, 60(1): 169–176, January 30, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783602000589.

Tsuboi:2004:SEW

[iTM04] Jun ichi Tsuboi and Kentaro Morita. Selectivity effects on wild white-spotted charr (Salvelinus leucomaenis) during a catch and release fishery. Fisheries Research, 69(2):229–238, September 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001262.

Tsuboi:2006:FDH

[iTMI06] Jun ichi Tsuboi, Kentaro Morita, and Hisatoshi Ikeda. Fate of deep-hooked white-spotted charr after cutting the line in a catch-and-release fishery. Fisheries Research, 79(1–2):226–230, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001056.

Irigoyen:2008:PHM

[IV08] Alejo J. Irigoyen and Leonardo A. Venerus. The ‘polehooking’ method: a novel and economical technique for in situ tagging small to medium-sized fishes. Fisheries Research, 91(2–3):349–353, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003670.

Irwin:2008:EAH

[IWBJ08] Brian J. Irwin, Michael J. Wilberg, James R. Bence, and Michael L. Jones. Evaluating alternative harvest policies for yellow perch in southern Lake Michigan. Fisheries Research, 94(3):267–281, December 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001604.

Incze:2003:CSO

[IWW03] Lewis S. Incze, Nicholas Wolff, and Richard A. Wahle. Can scientific observations of early life stages be scaled up to the level of a fished population? A case study using Homarus americanus. Fisheries Research, 65(1–3):33–46, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603002339.

Ismen:2007:AGR

[IYI07] Ali Ismen, Cigdem Yıgın, and Pınar Ismen. Age, growth, reproductive biology and feed of the common guitarfish (Rhinobatos rhinobatos Linnaeus, 1758) in İskenderun Bay, the eastern Mediterranean Sea. Fisheries Research, 84(2):263– 269, April 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606004462.

Jakobsdottir:2001:BAT

[Jak01] Klara B. Jakobsdóttir. Biological aspects of two deep-water squalid sharks: Centroscyllium fabricii (Reinhardt, 1825) and Etmopterus princeps (Collett, 1904) in Icelandic waters. Fisheries Research, 51(2–3):247–265, May 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002508.

Jurvelius:2005:DBS

[JAKM05] Juha Jurvelius, Heikki Auvinen, Irma Kolari, and Timo J. Marjomäki. Density and biomass of smelt (Osmerus eperlanus) in five Finnish lakes. Fisheries Research, 73(3):353– 361, July 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605000378.

Jiao:2004:AGL

[JC04] Yan Jiao and Yong Chen. An application of generalized linear models in production model and sequential population analysis. Fisheries Research, 70(2–3):367–376, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360400178X.

Jung:2009:SDM

[JCJ^{+}09] Sukgeun Jung, Ilsu Choi, Hongsung Jin, Dong woo Lee, Hyung kee Cha, Yeonghye Kim, and Jeong yong Lee. Size-dependent mortality formulation for isochronal fish species based on their fecundity: an example of Pacific cod (Gadus macrocephalus) in the eastern coastal areas of Korea. Fisheries Research, 97(1–2):77–85, April 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000186.

Joung:2008:AGR

[JCLL08] Shoou-Jeng Joung, Che-Tsung Chen, Hsian-Hau Lee, and Kwang-Ming Liu. Age, growth, and reproduction of silky sharks, Carcharhinus falciformis, in northeastern Taiwan waters. Fisheries Research, 90(1–3):78–85, April 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002482.

Jaiswar:2001:SAG

[JCS01] A. K. Jaiswar, S. K. Chakraborty, and R. P. Swamy. Studies on the age, growth and mortality rates of Indian scad Decapterus russelli (Ruppell) from Mumbai waters. Fisheries Research, 53(3):303–308, November 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003046.

Jiao:2005:ACR

[JCW05] Yan Jiao, Yong Chen, and Joe Wroblewski. An application of the composite risk assessment method in assessing fisheries stock status. Fisheries Research, 72(2–3):173–183, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604002991.

Jump:2008:CSM

[JDAM08] Christina M. Jump, Janet T. Duffy-Anderson, and Kathryn L. Mier. Comparison of the Sameoto, Manta, and MARMAP neustonic ichthyoplankton samplers in the Gulf of Alaska. Fisheries Research, 89(3):222–229, March 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002184.

Johansen:2000:SGR

[JDMN00] T. Johansen, A. K. Daníelsdóttir, K. Meland, and G. Nævdal. Studies of the genetic relationship between deepsea and oceanic Sebastes mentella in the Irminger Sea. Fisheries Research, 49(2):179–192, December 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001934.

Jenkins:2000:DAR

[JDS00] Wallace E. Jenkins, Michael R. Denson, and Theodore I. J. Smith. Determination of angler reporting level for red drum (Sciaenops ocellatus) in a South Carolina estuary. Fisheries Research, 44(3):273–277, January 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000867.

Jensen:2000:HRP

[Jen00] A. L. Jensen. Harvest reference points for the Beverton and Holt dynamic pool model. Fisheries Research, 47(1):93– 96, June 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699001265.

Jensen:2002:MHF

[Jen02] A. L. Jensen. Maximum harvest of a fish population that has the smallest impact on population biomass. Fisheries Research, 57(1):89–91, July 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100337X.

Jamieson:2006:ITG

[JGB^{+}06] Alan J. Jamieson, Olav Rune Godø, Phillip M. Bagley, Julian C. Partridge, and Imants G. Priede. Illumination of trawl gear by mechanically stimulated bioluminescence. Fisheries Research, 81(2–3):276–282, November 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002438.

Jonahson:2007:OWS

[JH07] Mananjo Jonahson and Simon Harding. Occurrence of whale sharks (Rhincodon typus) in Madagascar. Fisheries Research, 84(1):132–135, March 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004115.

Jorgensen:2004:PVA

[JHGS04] Roar Jørgensen, Nils Olav Handegard, Harald Gjøsæter, and Aril Slotte. Possible vessel avoidance behaviour of capelin in a feeding area and on a spawning ground. Fisheries Research, 69(2):251–261, September 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001328.

Jung:2008:FGD

[JHK08] Sukgeun Jung, Sun-Do Hwang, and Jinyeong Kim. Fecundity and growth-dependent mortality of Pacific anchovy (Engraulis japonicus) in Korean coastal waters. Fisheries Research, 93(1–2):39–46, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000593.

Johnson:2004:MCA

[JHSG04] G. E. Johnson, J. B. Hedgepeth, J. R. Skalski, and A. E. Giorgi. A Markov chain analysis of fish movements to determine entrainment zones. Fisheries Research, 69(3):349–358, October 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604001237.

Johnsen:2007:FAD

[JI07] Espen Johnsen and Titus Iilende. Factors affecting the diel variation in commercial CPUE of Namibian hake — can new information improve standard survey estimates? Fisheries Research, 88(1–3):70–79, December 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001713.

Jonsdottir:2003:NDR

[JIAD03] Ó. D. B. Jónsdóttir, A. K. Imsland, Ó. Ý. Atladóttir, and A. K. Daníelsdóttir. Nuclear DNA RFLP variation of Atlantic cod in the North Atlantic Ocean. Fisheries Research, 63(3):429–436, September 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000985.

Jonsdottir:2002:GHG

[JIDM02] Ó. D. B. Jónsdóttir, A. K. Imsland, A. K. Daníelsdóttir, and G. Marteinsdóttir. Genetic heterogeneity and growth properties of different genotypes of Atlantic cod (Gadus morhua L.) at two spawning sites off south Iceland. Fisheries Research, 55(1–3):37–47, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100296X.

Jorgensen:2006:SSC

[JIGI06] Terje Jørgensen, Ólafur Arnar Ingólfsson, Norman Graham, and Bjørnar Isaksen. Size selection of cod by rigid grids — is anything gained compared to diamond mesh codends only? Fisheries Research, 79(3):337–348, July 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000294.

Jutila:2003:MAS

[JJJ03] Eero Jutila, Erkki Jokikokko, and Markku Julkunen. Management of Atlantic salmon in the Simojoki river, northern Gulf of Bothnia: effects of stocking and fishing regulation. Fisheries Research, 64(1):5–17, October 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001073.

Jutila:2003:DSM

[JJKN^{+}03] Eero Jutila, Erkki Jokikokko, Irma Kallio-Nyberg, Irma Saloniemi, and Pentti Pasanen. Differences in sea migration between wild and reared Atlantic salmon (Salmo salar L.) in the Baltic Sea. Fisheries Research, 60(2–3):333–343, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602001698.

Jurajda:2009:SEE

[JJWO09] P. Jurajda, M. Janáč, S. M. White, and M. Ondračková. Small — but not easy: Evaluation of sampling methods in floodplain lakes including whole-lake sampling. Fisheries Research, 96(1):102–108, February 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002877.

Juza:2007:ETF

[JK07] T. Jůza and J. Kubečka. The efficiency of three fry trawls for sampling the freshwater pelagic fry community. Fisheries Research, 85(3):285–290, July 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000707.

Jech:2000:DEV

[JL00] J. Michael Jech and Jiangang Luo. Digital echo visualization and information system (DEVIS) for processing spatially-explicit fisheries acoustic data. Fisheries Research, 47(2–3):115–124, July 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001636.

Jan:2003:EPS

[JLC^{+}03] Rong-Quen Jan, Yu-Hsing Liu, Ching-Yi Chen, Min-Chang Wang, Gwo-Shyh Song, Hong-Cheng Lin, and Kwang-Tsao Shao. Effects of pile size of artificial reefs on the standing stocks of fishes. Fisheries Research, 63(3):327–337, September 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360300081X.

Joung:2004:AGS

[JLC04] Shoou-Jeng Joung, Yih-Yia Liao, and Che-Tsung Chen. Age and growth of sandbar shark, Carcharhinus plumbeus, in northeastern Taiwan waters. Fisheries Research, 70(1): 83–96, November 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604001602.

Jacobsen:2001:SDO

[JLHO01] Jan A. Jacobsen, Roar A. Lund, Lars P. Hansen, and Niall O’Maoileidigh. Seasonal differences in the origin of Atlantic salmon (Salmo salar L.) in the Norwegian Sea based on estimates from age structures and tag recaptures. Fisheries Research, 52(3):169–177, July 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002551.

Jurado-Molina:2006:ICA

[JMGC06] Jesús Jurado-Molina, Claudio Gatica, and Luis A. Cubillos. Incorporating cannibalism into an age-structured model for the Chilean hake. Fisheries Research, 82(1–3):30–40, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606003043.

Jaureguizar:2004:EFS

[JMGL04] Andrés J. Jaureguizar, Roberto Menni, Raul Guerrero, and Carlos Lasta. Environmental factors structuring fish communities of the Río de la Plata estuary. Fisheries Research, 66(2–3):195–211, February 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002005.

Johansen:2000:GCG

[JNDH00] T. Johansen, G. Nævdal, A. K. Daniélsdóttir, and N. R. Hareide. Genetic characterisation of giant Sebastes in the deep water slopes in the Irminger Sea. Fisheries Research, 45(3):207–216, April 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001320.

Johnsen:2003:IPL

[Joh03] Espen Johnsen. Improving the precision of length frequency distribution estimates from trawl surveys by including spatial covariance — using Namibian Merluccius capensis as an example. Fisheries Research, 62(1):7–20, April 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002758.

Jokikokko:2002:MWR

[Jok02] Erkki Jokikokko. Migration of wild and reared Atlantic salmon (Salmo salar L.) in the river Simojoki, northern Finland. Fisheries Research, 58(1):15–23, October 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003642.

Jones:2000:FGC

[Jon00] Cynthia M. Jones. Fitting growth curves to retrospective size-at-age data. Fisheries Research, 46(1–3):123–129, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001399.

Jukic-Peladic:2001:LTC

[JPVKS^{+}01] S. Jukic-Peladic, N. Vrgoc, S. Krstulovic-Sifner, C. Piccinetti, G. Piccinetti-Manfrin, G. Marano, and N. Ungaro. Long-term changes in demersal resources of the Adriatic Sea: comparison between trawl surveys carried out in 1948 and 1998. Fisheries Research, 53(1):95–104, September 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600002320.

Jirapunpipat:2008:EEV

[JPYW08] Kanchana Jirapunpipat, Pisit Phomikong, Masashi Yokota, and Seiichi Watanabe. The effect of escape vents in collapsible pots on catch and size of the mud crab Scylla olivacea. Fisheries Research, 94(1):73–78, October 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001975.

Jeffery:2002:VDS

[JR02] Shaun Jeffery and Andrew Revill. The vertical distribution of southern North Sea Crangon crangon (brown shrimp) in relation to towed fishing gears as influenced by water temperature. Fisheries Research, 55(1–3):319–323, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601002971.

Jardim:2007:GAS

[JR07] Ernesto Jardim and Paulo J. Ribeiro. Geostatistical assessment of sampling designs for Portuguese bottom trawl surveys. Fisheries Research, 85(3):239–247, July 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000574.

Johnson:2008:SEF

[JRG08] Daniel D. Johnson, Douglas Rotherham, and Charles A. Gray. Sampling estuarine fish and invertebrates using demersal otter trawls: Effects of net height, tow duration and diel period. Fisheries Research, 93(3):315–323, September 28, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360800177X.

Jurvelius:2000:MPP

[JRML00] Juha Jurvelius, Raimo Riikonen, Timo J. Marjomäki, and Juha Lilja. Mortality of pike-perch (Stizostedion lucioperca), brown trout (Salmo trutta) and landlocked salmon (Salmo salar m. sebago) caught as by-catch in pelagic trawling in a Finnish lake. Fisheries Research, 45(3):291–296, April 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699001162.

Jawad:2002:PGS

[JS02] Laith. A. Jawad and Peter. J. Smith. Phosphoglucomutase genotypes and spawning time in two species of anchovy, Thryssa hamiltoni and T. mystex, from the northwest Arabian Gulf. Fisheries Research, 59(1–2):273–277, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360100409X.

John:2005:MDW

[JS05] Jill St John and Clinton John Syers. Mortality of the demersal West Australian dhufish, Glaucosoma hebraicum (Richardson 1845) following catch and release: the influence of capture depth, venting and hook type. Fisheries Research, 76(1):106–116, October 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001682.

Jamieson:2004:CIR

[JSB04] R. E. Jamieson, H. P. Schwarcz, and J. Brattey. Carbon isotopic records from the otoliths of Atlantic cod (Gadus morhua) from eastern Newfoundland, Canada. Fisheries Research, 68(1–3):83–97, July 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360400075X.

Jones:2008:ITS

[JSO08] Emma G. Jones, Keith Summerbell, and Finbarr O’Neill. The influence of towing speed and fish density on the behaviour of haddock in a trawl cod-end. Fisheries Research, 94(2):166–174, November 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001884.

Jimenez:2004:OMD

[JSR04] M. P. Jiménez, I. Sobrino, and F. Ramos. Objective methods for defining mixed-species trawl fisheries in Spanish waters of the Gulf of Cádiz. Fisheries Research, 67(2):195– 206, April 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603003072.

Jimenez-Toribio:2006:EMP

[JTdH06] Ramón Jiménez-Toribio and Juan José García del Hoyo. Evidence of market price leadership in the Spanish red seabream value chain — implications for fisheries management. Fisheries Research, 81(1):51–59, October 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001913.

Jung:2008:SBD

[Jun08] Sukgeun Jung. Simulation-based daily cohort analysis of Pacific anchovy (Engraulis japonicus) in southern Korean coastal waters. Fisheries Research, 93(3):280–288, September 28, 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608001446.

Jennings:2005:EHL

[JWR05] Cecil A. Jennings, Ted A. Will, and Thomas R. Reinert. Efficacy of a high- and low-frequency ultrasonic probe for measuring ovary volume and estimating fecundity of striped bass Morone saxatilis in the Savannah River Estuary. Fisheries Research, 76(3):445–453, December 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002067.

Kacher:2005:DGG

[KA05] M. Kacher and R. Amara. Distribution and growth of 0-group European hake in the Bay of Biscay and Celtic Sea: a spatial and inter-annual analyses. Fisheries Research, 71(3):373–378, March 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002139.

Kimmel:2006:ETP

[KA06] William G. Kimmel and David G. Argent. Efficacy of two-pass electrofishing employing multiple units to assess stream fish species richness. Fisheries Research, 82(1–3):

14–18, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606003110.

Kurkilahti:2002:EFS

[KAHR02] M. Kurkilahti, M. Appelberg, T. Hesthagen, and M. Rask. Effect of fish shape on gillnet selectivity: a study with Fulton’s condition factor. Fisheries Research, 54(2):153–170, January 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600003015.

Kafemann:2000:VOS

[KAN00] R. Kafemann, S. Adlerstein, and R. Neukamm. Variation in otolith strontium and calcium ratios as an indicator of life-history strategies of freshwater fish species within a brackish water system. Fisheries Research, 46(1–3):313– 325, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600001569.

Kaiser:2004:EBA

[KAO04] M. J. Kaiser, M. C. V. Austen, and H. Ojaveer. European biodiversity action plan for fisheries: issues for non-target species. Fisheries Research, 69(1):1–6, August 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001043.

Karakulak:2004:CEB

[Kar04] F. Saadet Karakulak. Catch and effort of the bluefin tuna purse-seine fishery in Turkish waters. Fisheries Research, 68(1–3):361–366, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000712.

Katsanevakis:2005:ASD

[Kat05] Stelios Katsanevakis. Abundance and spatial distribution of the Mediterranean scallop, Pecten jacobaeus, in a marine lake. Fisheries Research, 76(3):417–429, December 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002092.

Katsanevakis:2006:MFG

[Kat06] Stelios Katsanevakis. Modelling fish growth: Model selection, multi-model inference and model selection uncertainty. Fisheries Research, 81(2–3):229–235, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606002645.

Kelly:2006:CDF

[KC06] Ciaran J. Kelly and Edward A. Codling. ‘Cheap and dirty’ fisheries science and management in the North Atlantic. Fisheries Research, 79(3):233–238, July 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001135.

Kanaiwa:2005:ASA

[KCH05] Minoru Kanaiwa, Yong Chen, and Margaret Hunter. Assessing a stock assessment framework for the green sea urchin Strongylocentrotus drobachiensis fishery in Maine, USA. Fisheries Research, 74(1–3):96–115, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001074.

Koeller:2003:SST

[KCK03] P. A. Koeller, M. Covey, and M. King. Is size at sex transition an indicator of growth or abundance in pandalid shrimp? Fisheries Research, 65(1–3):217–230, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603002455.

Kingsley:2002:SSR

[KCKP02] M. C. S. Kingsley, D. M. Carlsson, P. Kanneworff, and M. Pennington. Spatial structure of the resource of Pandalus borealis and some implications for trawl survey. Fisheries Research, 58(2):171–183, November 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003885.

Kostylev:2003:SEG

[KCRT03] Vladimir E. Kostylev, Robert C. Courtney, Ginette Robert, and Brian J. Todd. Stock evaluation of giant scallop (Placopecten magellanicus) using high-resolution acoustics for seabed mapping. Fisheries Research, 60(2–3):479–492, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602001005.

Krutzikowsky:2005:DDS

[KE05] Gregory K. Krutzikowsky and Robert L. Emmett. Diel differences in surface trawl fish catches off Oregon and Washington. Fisheries Research, 71(3):365–371, March 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002309.

Kerstan:2000:EPA

[Ker00] M. Kerstan. Estimation of precise ages from the marginal increment widths of differently growing sardine (Sardinops sagax) otoliths. Fisheries Research, 46(1–3):207–225, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001478.

Kesteven:2000:BRW

[Kes00] G. L. Kesteven. Book review: World Fisheries — What is to be done?: Alexander Fridman (Ed.); Baird publications, Southbank, Melbourne, Vic., Australia. 1999, 146 ill., 249 pp., $A (hbk), ISBN: 0-958703-5-0. Fisheries Research, 48(1):94–95, August 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001247.

Kastelle:2002:TLR

[KF02] Craig R. Kastelle and Joan E. Forsberg. Testing for loss of ^{222}Rn from Pacific halibut (Hippoglossus stenolepis) otoliths. Fisheries Research, 57(1):93–98, July 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003393.

Kvamme:2004:AES

[KF04] Cecilie Kvamme and Kristin Guldbrandsen Frøysa. Assessing the effects on stocks of selectivity changes in a fishery. Fisheries Research, 69(2):283–292, September 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000761.

Krag:2008:ESM

[KFM08] Ludvig Ahm Krag, Rikke Petri Frandsen, and Niels Madsen. Evaluation of a simple means to reduce discard in the Kattegat–Skagerrak Nephrops (Nephrops norvegicus) fishery: Commercial testing of different codends and squaremesh panels. Fisheries Research, 91(2–3):175–186, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607003396.

Knudsen:2004:HMF

[KFO^{+}04] F. R. Knudsen, J. E. Fosseidengen, F. Oppedal, Ø. Karlsen, and E. Ona. Hydroacoustic monitoring of fish in sea cages: target strength (TS) measurements on Atlantic salmon (Salmo salar). Fisheries Research, 69(2):205–209, September 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001377.

Knudsen:2004:HOI

[KG04] Frank Reier Knudsen and Karl Øystein Gjelland. Hydroacoustic observations indicating swimbladder volume compensation during the diel vertical migration in coregonids (Coregonus lavaretus and Coregonus albula). Fisheries Research, 66(2–3):337–341, February 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001917.

Kerstetter:2006:ECV

[KG06] D. W. Kerstetter and J. E. Graves. Effects of circle versus J-style hooks on target and non-target species in a pelagic longline fishery. Fisheries Research, 80(2–3):239– 250, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606001433.

Kennedy:2009:WCY

[KGB09] J. Kennedy, A. C. Gundersen, and J. Boje. When to count your eggs: Is fecundity in Greenland halibut (Reinhardtius hippoglossoides w.) down-regulated? Fisheries Research, 100(3):260–265, November 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360900229X.

Krumme:2005:FDS

[KH05] Uwe Krumme and Andreas Hanning. A floating device for stationary hydroacoustic sampling in shallow waters. Fisheries Research, 73(3):377–381, July 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000317.

Kubecka:2009:TPL

[KHM^{+}09] Jan Kubečka, Eva Hohausová, Josef Matěna, Jiří Peterka, Upali S. Amarasinghe, Scott A. Bonar, Jon Hateley, Phil Hickley, Petri Suuronen, Vladimir Tereschenko, Robin Welcomme, and Ian J. Winfield. The true picture of a lake or reservoir fish stock: a review of needs and progress. Fisheries Research, 96(1):1–5, February 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002737.

Kvalsvik:2006:GSN

[KHMG06] Kurt Kvalsvik, Ingvar Huse, Ole Arve Misund, and Kjell Gamst. Grid selection in the North Sea industrial trawl fishery for Norway pout: Efficient size selection reduces bycatch. Fisheries Research, 77(2):248–263, February 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002730.

Knudsen:2009:DIB

[KHMS09] Frank R. Knudsen, Anthony Hawkins, Rob McAllen, and Olav Sand. Diel interactions between sprat and mackerel in a marine lough and their effects upon acoustic measurements of fish abundance. Fisheries Research, 100(2):140–147, October 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001830.

Kvamme:2004:TSC

[KI04] Cecilie Kvamme and Bjørnar Isaksen. Total selectivity of a commercial cod trawl with and without a grid mounted: grid and codend selectivity of north-east Arctic cod. Fisheries Research, 68(1–3):305–318, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000244. See erratum [KI05].

Kvamme:2005:ETS

[KI05] Cecilie Kvamme and Bjørnar Isaksen. Erratum to “Total selectivity of a commercial cod trawl with and without a grid mounted: grid and codend selectivity of north-east Arctic cod” [Fish. Res. 68 (2004) 305–318]. Fisheries Research, 72(2–3):363–364, May 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604003273. See [KI04].

Kirchner:2001:FRB

[Kir01] Carola H. Kirchner. Fisheries regulations based on yieldper-recruit analysis for the linefish silver kob Argyrosomus inodorus in Namibian waters. Fisheries Research, 52 (3):155–167, July 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783600002587.

Kimura:2004:ELH

[KIT^{+}04] Y. Kimura, S. Ishikawa, T. Tokai, M. Nishida, and K. Tsukamoto. Early life history characteristics and genetic homogeneity of Conger myriaster leptocephali along the east coast of central Japan. Fisheries Research, 70 (1):61–69, November 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604001560.

Kruse:2009:RDF

[KIZ09] Gordon H. Kruse, Yukimasa Ishida, and Chang Ik Zhang. Rebuilding of depleted fish stocks through an ecosystem approach to fisheries. Fisheries Research, 100(1):1–5, September 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360900201X.

Klibansky:2008:PEP

[KJ08] Nikolai Klibansky and Francis Juanes. Procedures for efficiently producing high-quality fecundity data on a small budget. Fisheries Research, 89(1):84–89, January 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360700224X.

Knight:2009:APP

[KJ09] Benjamin R. Knight and Weimin Jiang. Assessing primary production constraints in New Zealand fisheries. Fisheries Research, 100(1):15–25, September 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001489.

Kallayil:2003:BGN

[KJEF03] Jawhar K. Kallayil, Terje Jørgensen, Arill Engås, and Anders Fernö. Baiting gill nets — how is fish behaviour affected? Fisheries Research, 61(1–3):125–133, March 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001819.

Kerr:2006:RCS

[KJS^{+}06] S. Kerr, K. Johnson, J. Side, M. Baine, C. Davos, and J. Henley. Resolving conflicts in selecting a programme of fisheries science investigation. Fisheries Research, 79(3):313–324, July 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001093.

Kraus:2004:EBS

[KK04] Gerd Kraus and Friedrich W. Köster. Estimating Baltic sprat (Sprattus sprattus balticus s.) population sizes from egg production. Fisheries Research, 69(3):313–329, October 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001304.

Kitada:2006:LLJ

[KK06] Shuichi Kitada and Hirohisa Kishino. Lessons learned from Japanese marine finfish stock enhancement programmes. Fisheries Research, 80(1):101–112, August 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001251.

Khan:2009:CAE

[KK09] M. Afzal Khan and Shahista Khan. Comparison of age estimates from scale, opercular bone, otolith, vertebrae and dorsal fin ray in Labeo rohita (Hamilton), Catla catla (Hamilton) and Channa marulius (Hamilton). Fisheries Research, 100(3):255–259, November 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002264.

Kastelle:2000:UPR

[KKJ00] Craig R. Kastelle, Daniel K. Kimura, and Shelly R. Jay. Using ^{210}Pb/^{226}Ra disequilibrium to validate conventional ages in Scorpaenids (genera Sebastes and Sebastolobus). Fisheries Research, 46(1–3):299–312, May 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001557.

Katselis:2005:MEA

[KKR^{+}05] George Katselis, Constantin Koutsikopoulos, Yiannis Rogdakis, Thanasis Lachanas, Evagelos Dimitriou, and Kosmas Vidalis. A model to estimate the annual production of roes (avgotaracho) of flathead mullet (Mugil cephalus) based on the spawning migration of species. Fisheries Research, 75(1–3):138–148, September 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001347.

Knudsen:2006:ASL

[KLJ06] Frank Reier Knudsen, Petter Larsson, and Per Johan Jakobsen. Acoustic scattering from a larval insect (Chaoborus flavicans) at six echosounder frequencies: Implication for acoustic estimates of fish abundance. Fisheries Research, 79(1–2):84–89, June 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000233.

Kim:2007:DSB

[KLS^{+}07] Hyun-Young Kim, Chun-Woo Lee, Jong-Keun Shin, Hyung-Seok Kim, Bong-Jin Cha, and Gun-Ho Lee. Dynamic simulation of the behavior of purse seine gear and sea-trial verification. Fisheries Research, 88(1–3):109–119, December 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607001920.

Kjellman:2001:ITS

[KLU01] Jakob Kjellman, Jyrki Lappalainen, and Lauri Urho. Influence of temperature on size and abundance dynamics of age-0 perch and pikeperch. Fisheries Research, 53(1): 47–56, September 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783600002654.

Kirchner:2002:NIM

[KM02] Carola Kirchner and Murdoch McAllister. A new improved method to compute swept area estimates of biomass from commercial catch rate data: application to Namibian orange roughy (Hoplostethus atlanticus). Fisheries Research, 56(1):69–88, April 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003113.

Katsukawa:2003:SET

[KM03] Toshio Katsukawa and Hiroyuki Matsuda. Simulated effects of target switching on yield and sustainability of fish stocks. Fisheries Research, 60(2–3):515–525, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000838.

Kasapidis:2008:DAM

[KM08a] P. Kasapidis and A. Magoulas. Development and application of microsatellite markers to address the population structure of the horse mackerel Trachurus trachurus. Fisheries Research, 89(2):132–135, February 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002391.

Kleiber:2008:IBU

[KM08b] Pierre Kleiber and Mark N. Maunder. Inherent bias in using aggregate CPUE to characterize abundance of fish species assemblages. Fisheries Research, 93(1–2):140–145, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360800091X.

Kristoffersen:2008:PSA

[KM08c] Jon Bent Kristoffersen and Antonios Magoulas. Population structure of anchovy Engraulis encrasicolus L. in the Mediterranean Sea inferred from multiple methods. Fisheries Research, 91(2–3):187–195, June 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003402.

Kloppmann:2001:DBW

[KMB01] Matthias Kloppmann, Christian Mohn, and Joachim Bartsch. The distribution of blue whiting eggs and larvae on Porcupine Bank in relation to hydrography and currents. Fisheries Research, 50(1–2):89–109, February 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002447.

Kvalsvik:2002:SSL

[KME^{+}02] Kurt Kvalsvik, Ole Arve Misund, Arill Engås, Kjell Gamst, Renè Holst, Derek Galbraith, and Håkon Vederhus. Size selection of large catches: using sorting grid in pelagic mackerel trawl. Fisheries Research, 59(1–2):129–148, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601004088.

Kawakita:2005:IPT

[KMELC05] M. Kawakita, M. Minami, S. Eguchi, and C. E. Lennert-Cody. An introduction to the predictive technique AdaBoost with a comparison to generalized additive models. Fisheries Research, 76(3):328–343, December 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360500216X.

Kifani:2008:NEA

[KMF08] Souad Kifani, Hicham Masski, and Abdelmalek Faraj. The need of an ecosystem approach to fisheries: the Moroccan upwelling-related resources case. Fisheries Research, 94(1):36–42, October 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608001938.

Krag:2009:SFB

[KMK09] Ludvig A. Krag, Niels Madsen, and Junita D. Karlsen. A study of fish behaviour in the extension of a demersal trawl using a multi-compartment separator frame and SIT camera system. Fisheries Research, 98(1–3):62–66, June 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000794.

Katsarou:2001:PGS

[KN01] Euthemia Katsarou and Gunnar Nævdal. Population genetic studies of the roughhead grenadier, Macrourus berglax L., in the North Atlantic Ocean. Fisheries Research, 51(2–3):207–215, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002466.

Kallio-Nyberg:2006:SRA

[KNJJS06] Irma Kallio-Nyberg, Eero Jutila, Erkki Jokikokko, and Irma Saloniemi. Survival of reared Atlantic salmon and sea trout in relation to marine conditions of smolt year in the Baltic Sea. Fisheries Research, 80(2–3):295–304, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606001366.

Knowler:2007:ESR

[Kno07] Duncan Knowler. Estimation of a stock–recruitment relationship for Black Sea anchovy (Engraulis encrasicolus) under the influence of nutrient enrichment and the invasive comb-jelly, Mnemiopsis leidyi. Fisheries Research, 84(3):275–281, May 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003870.

Kallio-Nyberg:2009:MSR

[KNSSKF09] Irma Kallio-Nyberg, Matti Salminen, Irma Saloniemi, and Laura Kannala-Fisk. Marine survival of reared Atlantic salmon in the Baltic Sea: the effect of smolt traits and annual factors. Fisheries Research, 96(2–3):289–295, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608003962.

Koeller:2003:LSR

[Koe03] P. Koeller. The lighter side of reference points. Fisheries Research, 62(1):1–6, April 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0165783602002412.

Koeller:2008:EBP

[Koe08] Peter Koeller. Ecosystem-based psychology, or, how I learned to stop worrying and love the data. Fisheries Research, 90(1–3):1–5, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003049.

Kynoch:2004:ESB

[KOO04] R. J. Kynoch, M. C. O’Dea, and F. G. O’Neill. The effect of strengthening bags on cod-end selectivity of a Scottish demersal trawl. Fisheries Research, 68(1–3):249–257, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604000220.

Kope:2006:CEM

[Kop06] Robert Kope. Cumulative effects of multiple sources of bias in estimating spawner–recruit parameters with application to harvested stocks of Chinook salmon (Oncorhynchus tshawytscha). Fisheries Research, 82(1–3):101–110, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606002955.

Korneliussen:2004:BEI

[Kor04] Rolf J. Korneliussen. The Bergen echo integrator postprocessing system, with focus on recent improvements. Fisheries Research, 68(1–3):159–169, July 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000499.

King:2007:IAS

[KP07] Sarah E. King and Eric N. Powell. Influence of adaptive stations in a transect-based sampling design for a multispecies fish survey. Fisheries Research, 86(2–3):241–261, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607001439.

Kolody:2008:SPL

[KPBD08] D. Kolody, T. Polacheck, M. Basson, and C. Davies. Salvaged pearls: lessons learned from a floundering attempt to develop a management procedure for southern bluefin tuna. Fisheries Research, 94(3):339–350, December 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002609.

Kulmala:2007:ITQ

[KPL^{+}07] S. Kulmala, H. Peltomäki, M. Lindroos, P. Söderkultalahti, and S. Kuikka. Individual transferable quotas in the Baltic Sea herring fishery: a socio-bioeconomic analysis. Fisheries Research, 84(3):368–377, May 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763

(electronic).  URL http://www.sciencedirect.com/
science/article/pii/S0165783606004267. See erratum

[KPL^{+}09].

Kulmala:2009:EIT

[KPL^{+}09] S. Kulmala, H. Peltomäki, M. Lindroos, P. Söderkultalahtir, and S. Kuikka. Erratum to “Individual transferable quotas in the Baltic Sea herring fishery: a sociobioeconomic analysis” [Fish. Res. 84 (2007) 368–377]. Fisheries Research, 100(2):183, October 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001866. See [KPL^{+}07].

Karlou-Riga:2000:OMA

[KR00] C. Karlou-Riga. Otolith morphology and age and growth of Trachurus mediterraneus (Steindachner) in the Eastern Mediterranean. Fisheries Research, 46(1–3):69–82, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360000134X.

Kronen:2004:FFS

[Kro04] Mecki Kronen. Fishing for fortunes?: a socio-economic assessment of Tonga’s artisanal fisheries. Fisheries Research, 70(1):121–134, November 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001614.

Krumme:2004:PTM

[Kru04] Uwe Krumme. Patterns in tidal migration of fish in a Brazilian mangrove channel as revealed by a split-beam echosounder. Fisheries Research, 70(1):1–15, November 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001547.

Knudsen:2002:BHB

[KS02] Frank Reier Knudsen and Harald Sægrov. Benefits from horizontal beaming during acoustic survey: application to three Norwegian lakes. Fisheries Research, 56(2):205–211, May 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601003186.

Kaimmer:2008:FIR

[KS08] Steve Kaimmer and Allan W. Stoner. Field investigation of rare-earth metal as a deterrent to spiny dogfish in the Pacific halibut fishery. Fisheries Research, 94(1):43–47, October 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360800194X.

Kauppinen:2005:TRP

[KSS05] Tiina Kauppinen, Antti Siira, and Petri Suuronen. Temporal and regional patterns in seal-induced catch and gear damage in the coastal trap-net fishery in the northern Baltic Sea: effect of netting material on damage. Fisheries Research, 73(1–2):99–109, June 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000056.

Kirsch:2000:AEZ

[KTC00] Jay Kirsch, G. L. Thomas, and R. Ted Cooney. Acoustic estimates of zooplankton distributions in Prince William Sound, spring 1996. Fisheries Research, 47(2–3):245–260, July 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001739.

Kevrekidis:2006:CRS

[KTL06] K. Kevrekidis and M. Thessalou-Legaki. Catch rates, size structure and sex ratio of Melicertus kerathurus (Decapoda: Penaeidae) from an Aegean Sea trawl fishery. Fisheries Research, 80(2–3):270–279, September 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001408.

Kallianiotis:2004:FSA

[KVS04] A. Kallianiotis, P. Vidoris, and G. Sylaios. Fish species assemblages and geographical sub-areas in the North Aegean Sea, Greece. Fisheries Research, 68(1–3):171–187, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604000463.

Kuster:2005:LTT

[KVZ05] C. Kuster, V. C. Vuki, and L. P. Zann. Longterm trends in subsistence fishing patterns and coral reef fisheries yield from a remote Fijian island. Fisheries Research, 76(2):221–228, November 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001979.

Kim:2003:ORE

[KW03a] Yong-Hae Kim and Clem S. Wardle. Optomotor response and erratic response: quantitative analysis of fish reaction to towed fishing gears. Fisheries Research, 60(2–3):455–470, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602001145.

Kim:2003:RYH

[KW03b] Yong-Hae Kim and Clem S. Wardle. Reply from Yong-Hae Kim to the letter to the editor of Walsh and Godø. Fisheries Research, 63(2):293, August 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001292.

Kim:2005:BMF

[KW05] Yong-Hae Kim and Clement S. Wardle. Basic modelling of fish behaviour in a towed trawl based on chaos in decisionmaking. Fisheries Research, 73(1–2):217–229, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604003078.

Kanaji:2009:MCJ

[KWK^{+}09] Yu Kanaji, Yoshiro Watanabe, Tomohiko Kawamura, Songguang Xie, Yoh Yamashita, Chiyuki Sassa, and Youichi Tsukamoto. Multiple cohorts of juvenile jack mackerel Trachurus japonicus in waters along the Tsushima Warm Current. Fisheries Research, 95(2–3):139–145, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608002312.

Labelle:2005:TMC

[Lab05]  Marc Labelle.  Testing the MULTIFAN-CL assess-

ment model using simulated tuna fisheries data. Fisheries Research, 71(3):311–334, March 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002140.

Laird:2000:BRA

[Lai00] Lindsay Laird. Book review: Aquaculture and Biotechnology: Iddya Karunasagar, Indrani Karunasagar, Alan Reilly (Eds.); Science Publishers Inc., USA 1999, 186 pages (£45). Fisheries Research, 48(3):289–290, October 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002228.

Lajus:2007:HFW

[LAL07] Dmitry L. Lajus, Yaroslava I. Alekseeva, and Julia A. Lajus. Herring fisheries in the White Sea in the 18th– beginning of the 20th centuries: Spatial and temporal patterns and factors affecting the catch fluctuations. Fisheries Research, 87(2–3):255–259, November 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001440.

Landergren:2004:FAE

[Lan04] Peter Landergren. Factors affecting early migration of sea trout Salmo trutta parr to brackish water. Fisheries Research, 67(3):283–294, May 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603003060.

Laptikhovsky:2002:DFR

[Lap02] Vladimir Laptikhovsky. Diurnal feeding rhythm of the short-fin squid Illex argentinus (Cephalopoda: Ommastrephidae) in the Falkland waters. Fisheries Research, 59(1–2):233–237, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000061.

Lazzari:2008:HVY

[Laz08] Mark A. Lazzari. Habitat variability in young-of-the-year winter flounder, Pseudopleuronectes americanus, in Maine estuaries. Fisheries Research, 90(1–3):296–304, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003116.

Laptikhovsky:2005:PTF

[LB05] Vladimir Laptikhovsky and Paul Brickle. The Patagonian toothfish fishery in Falkland Islands’ waters. Fisheries Research, 74(1–3):11–23, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360500127X.

Linton:2008:EME

[LB08] Brian C. Linton and James R. Bence. Evaluating methods for estimating process and observation error variances in statistical catch-at-age analysis. Fisheries Research, 94(1):26–35, October 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001926.

Louiz:2009:GIG

[LBABH09] Ibtissem Louiz, Mossadok Ben-Attia, and Oum Kalthoum Ben-Hassine. Gonadosomatic index and gonad histopathology of Gobius niger (Gobiidea, Teleost) from Bizerta lagoon (Tunisia): Evidence of reproduction disturbance.

Fisheries Research, 100(3):266–273, November 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002306.

Link:2008:VAS

[LBKG08] Jason Link, Jay Burnett, Paul Kostovick, and John Galbraith. Value-added sampling for fishery independent surveys: Don’t stop after you’re done counting and measuring. Fisheries Research, 93(1–2):229–233, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001367.

Lychakov:2000:OAB

[LBMCE00] D. V. Lychakov, A. Boyadzhieva-Mikhailova, I. Christov, and I. I. Evdokimov. Otolithic apparatus in Black Sea elasmobranchs. Fisheries Research, 46(1–3):27–38, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001302.

Landaeta:2006:SLS

[LC06] Mauricio F. Landaeta and Leonardo R. Castro. Spawning and larval survival of the Chilean hake Merluccius gayi under later summer conditions in the Gulf of Arauco, central Chile. Fisheries Research, 77(1):115–121, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002705.

Lercari:2007:PCR

[LC07] Diego Lercari and Ernesto A. Chávez. Possible causes related to historic stock depletion of the totoaba, Totoaba macdonaldi (Perciformes: Sciaenidae), endemic to the Gulf of California. Fisheries Research, 86(2–3):136–142, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360700118X.

Lorenzoni:2002:GPE

[LCD^{+}02] Massimo Lorenzoni, Massimiliano Corboli, A. J. Martin Dörr, Mario Mearelli, and Giancarlo Giovinazzo. The growth of pike (Esox lucius Linnaeus, 1798) in Lake Trasimeno (Umbria, Italy). Fisheries Research, 59(1–2):239– 246, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783602000139.

Lombardi-Carlson:2008:RSA

[LCFP^{+}08] Linda Lombardi-Carlson, Gary Fitzhugh, Chris Palmer, Chris Gardner, Robert Farsky, and Mauricio Ortiz. Regional size, age and growth differences of red grouper (Epinephelus morio) along the west coast of Florida. Fisheries Research, 91(2–3):239–251, June 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003463.

Liu:2006:SPR

[LCNJ06] Kwang-Ming Liu, Yu-Ting Chang, I.-Hsun Ni, and Chien-Bang Jin. Spawning per recruit analysis of the pelagic thresher shark, Alopias pelagicus, in the eastern Taiwan waters. Fisheries Research, 82(1–3):56–64, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003006.

Lizarraga-Cubedo:2003:CSM

[LCTB^{+}03] H. A. Lizárraga-Cubedo, I. Tuck, N. Bailey, G. J. Pierce, and J. A. M. Kinnear. Comparisons of size at maturity and fecundity of two Scottish populations of the European lobster, Homarus gammarus. Fisheries Research, 65(1–3):137–152, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002406.

Lorance:2001:CAI

[LD01] Pascal Lorance and Hervé Dupouy. CPUE abundance indices of the main target species of the French deep-water fishery in ICES Sub-areas V–VII. Fisheries Research, 51(2–3):137–149, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002417.

Lunn:2006:MSS

[LD06] Kristin E. Lunn and Philip Dearden. Monitoring smallscale marine fisheries: an example from Thailand’s Ko Chang archipelago. Fisheries Research, 77(1):60–71, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605002791.

LaMesa:2008:ELH

[LD08] Mario La Mesa and Francesca De Rossi. Early life history of the black anglerfish Lophius budegassa Spinola, 1807 in the Mediterranean Sea using otolith microstructure. Fisheries Research, 93(1–2):234–239, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001458.

Lorance:2001:ARG

[LDA01] Pascal Lorance, Hervé Dupouy, and Valérie Allain. Assessment of the roundnose grenadier (Coryphaenoides rupestris) stock in the Rockall Trough and neighbouring areas (ICES Sub-areas V–VII). Fisheries Research, 51(2–3):151– 163, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601002429.

Lorenzoni:2002:GRL

[LDE^{+}02] Massimo Lorenzoni, A. J. Martin Dörr, Rocco Erra, Giancarlo Giovinazzo, Mario Mearelli, and Sabrina Selvi. Growth and reproduction of largemouth bass (Micropterus salmoides Lacépède, 1802) in Lake Trasimeno (Umbria, Italy). Fisheries Research, 56(1):89–95, April 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003095.

LaMesa:2009:GMR

[LDGA09] M. La Mesa, F. Donato, G. Giannetti, and E. Arneri. Growth and mortality rates of European anchovy (Engraulis encrasicolus) in the Adriatic Sea during the transition from larval to juvenile stages. Fisheries Research, 96 (2–3):275–280, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608003949.

Lajus:2007:ASF

[LDK^{+}07] Dmitry L. Lajus, Zoya V. Dmitrieva, Alexei V. Kraikovski, Julia A. Lajus, and Daniel A. Alexandrov. Atlantic salmon fisheries in the white and Barents Sea basins: Dynamic of catches in the 17–18th century and comparison with 19– 20th century data. Fisheries Research, 87(2–3):240–254, November 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607001610.

Lessa:2004:AGY

[LDN04] Rosangela Lessa and Paulo Duarte-Neto. Age and growth of yellowfin tuna (Thunnus albacares) in the western equatorial Atlantic, using dorsal fin spines. Fisheries Research, 69(2):157–170, September 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001365.

Lobo:2001:AGR

[LE01] Carina Lobo and Karim Erzini. Age and growth of Ray’s bream (Brama brama) from the south of Portugal. Fisheries Research, 51(2–3):343–347, May 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002582.

Larsen:2007:RHP

[LET07] Finn Larsen, Ole Ritzau Eigaard, and Jakob Tougaard. Reduction of harbour porpoise (Phocoena phocoena) bycatch by iron-oxide gillnets. Fisheries Research, 85(3):270–278, July 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607000550.

Levi:2005:TNI

[Lev05] D. Levi. Two new ideas for the common fisheries policy. Fisheries Research, 71(2):255–256, February 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002358.

Lancaster:2002:FDJ

[LF02] Jane Lancaster and Christopher L. J. Frid. The fate of discarded juvenile brown shrimps (Crangon crangon) in the Solway Firth UK fishery. Fisheries Research, 58(1):95–107, October 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601003575.

Long:2008:EPO

[LFA08] Le Kim Long, Ola Flaaten, and Nguyen Thi Kim Anh. Economic performance of open-access offshore fisheries — the case of Vietnamese longliners in the South China Sea. Fisheries Research, 93(3):296–304, September 28, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001616.

Link:2002:CPA

[LG02] Jason S. Link and Lance P. Garrison. Changes in piscivory associated with fishing induced changes to the finfish community on Georges Bank. Fisheries Research, 55(1–3):71– 86, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601003009.

Lai:2004:LME

[LH04] Han-Lin Lai and Thomas Helser. Linear mixed-effects models for weight–length relationships. Fisheries Research, 70(2–3):377–387, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001791.

Leite:2009:HDA

[LHMO09] T. S. Leite, M. Haimovici, J. Mather, and J. E. Lins Oliveira. Habitat, distribution, and abundance of the commercial octopus (Octopus insularis) in a tropical oceanic island, Brazil: Information for management of an artisanal fishery inside a marine protected area.

Fisheries Research, 98(1–3):85–91, June 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000939.

Lauth:2004:ESS

[LIW04] R. R. Lauth, J. Ianelli, and W. W. Wakefield. Estimating the size selectivity and catching efficiency of a survey bottom trawl for thornyheads, Sebastolobus spp. using a towed video camera sled. Fisheries Research, 70(1):27–37, November 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604001468.

Laurenson:2005:MGM

[LJP05] C. H. Laurenson, A. Johnson, and I. G. Priede. Movements and growth of monkfish Lophius piscatorius tagged at the Shetland Islands, northeastern Atlantic. Fisheries Research, 71(2):185–195, February 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001857.

Lee:2000:LWA

[LK00] E. M. Lee and R. A. Khan. Length–weight–age relationships, food, and parasites of Atlantic cod (Gadus morhua) off coastal Labrador within NAFO divisions 2H and 2J-3K. Fisheries Research, 45(1):65–72, February 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001010.

Lai:2002:ASE

[LK02] Han-Lin Lai and Daniel K. Kimura. Analyzing survey experiments having spatial variability with an application to a sea scallop fishing experiment. Fisheries Research, 56(3):239–259, June 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100323X.

Lauridsen:2008:CMC

[LLJ^{+}08] Torben L. Lauridsen, Frank Landkildehus, Erik Jeppesen, Torben B. Jørgensen, and Martin Søndergaard. A comparison of methods for calculating Catch Per Unit Effort (CPUE) of gill net catches in lakes. Fisheries Research, 93(1–2):204–211, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360800132X.

Liu:2009:AGE

[LLJC09] Kwang-Ming Liu, Mei-Ling Lee, Shoou-Jeng Joung, and Yung-Chou Chang. Age and growth estimates of the sharptail mola, Masturus lanceolatus, in waters of eastern Taiwan. Fisheries Research, 95(2–3):154–160, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608002452.

LaMesa:2005:AGM

[LLM05] M. La Mesa, G. La Mesa, and M. Micalizzi. Age and growth of Madeira scorpionfish, Scorpaena maderensis Valenciennes, 1833, in the central Mediterranean. Fisheries Research, 74(1–3):265–272, August 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360500041X.

Lomovasky:2008:DSM

[LLV^{+}08] Betina J. Lomovasky, Mario Lasta, Macarena Valiñas, Martín Bruschetti, Pablo Ribeiro, Silvana Campodónico, and Oscar Iribarne. Differences in shell morphology and internal growth pattern of the Patagonian scallop Zygochlamys patagonica in the four main beds across their SW Atlantic distribution range. Fisheries Research, 89(3):266–275, March 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002251.

Lester:2009:UAP

[LM09] R. J. G. Lester and K. MacKenzie. The use and abuse of parasites as stock markers for fish. Fisheries Research, 97(1–2):1–2, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360900006X.

Lopez-Martinez:2003:IVG

[LMASHV^{+}03] Juana López-Martínez, Francisco Arreguín-Sánchez, Sergio Hernández-Vázquez, Alma Rosa Garcıá-Juárez, and Wenceslao Valenzuela-Quiñonez. Interannual variation of growth of the brown shrimp Farfantepenaeus californiensis and its relation to temperature. Fisheries Research, 61(1–3):95–105, March 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002394.

Lopez-Martinez:2002:SAP

[LMASMASZ02] Juana López-Martínez, Francisco Arreguín-Sánchez, Rufino Morales-Azpeitia, and César Salinas-Zavala. Stock assessment and potential yield of the rock shrimp, Sicyonia penicillata, fishery of Bahia Kino, Sonora, Mexico. Fisheries Research, 59(1–2):71–81, December 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100412X.

Loaec:2006:EDF

[LMMD06] H. Loaec, F. Morandeau, M. Meillat, and P. Davies. Engineering development of flexible selectivity grids for Nephrops. Fisheries Research, 79(1–2):210–218, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000361.

Lou:2005:UOW

[LMR^{+}05] Dong C. Lou, Bruce D. Mapstone, Garry R. Russ, Campbell R. Davies, and Gavin A. Begg. Using otolith weight–age relationships to predict age-based metrics of coral reef fish populations at different spatial scales. Fisheries Research, 71(3):279–294, March 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002279.

Lou:2007:UOW

[LMR^{+}07] Dong C. Lou, Bruce D. Mapstone, Garry R. Russ, Gavin A. Begg, and Campbell R. Davies. Using otolith weight– age relationships to predict age based metrics of coral reef fish populations across different temporal scales. Fisheries Research, 83(2–3):216–227, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003420.

Lopez-Martinez:2005:GRS

[LMRQNM^{+}05] Juana López-Martínez, Carlos Rábago-Quiroz, Manuel O. Nevárez-Martínez, Alma Rosa García-Juárez, Gabriel Rivera-Parra, and Jorge Chávez-Villalba. Growth, reproduction, and size at first maturity of blue shrimp, Litopenaeus stylirostris (Stimpson, 1874) along the east coast of the Gulf of California, Mexico. Fisheries Research, 71 (1):93–102, January 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604001274.

Lewy:2004:SDS

[LNG04] Peter Lewy, Anders Nielsen, and Henrik Gislason. Stock dynamics of sandeel in the North Sea and sub-regions including uncertainties. Fisheries Research, 68(1–3):237–248, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604000281.

Lam:2008:ILT

[LNS08] Chi H. Lam, Anders Nielsen, and John R. Sibert. Improving light and temperature based geolocation by unscented Kalman filtering. Fisheries Research, 91(1):15–25, May 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607003086.

Lockwood:2002:BRF

[Loc02] Stephen J. Lockwood. Book review: Fishes in Estuaries: M. Elliott, K. L. Hemingway (Eds.); Blackwell Science, Oxford, 636 pages, hardback, ISBN 0-632-05733-5, £89.50. Fisheries Research, 58(3):410–412, November 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602001674.

Loher:2008:HSF

[Loh08] Timothy Loher. Homing and summer feeding site fidelity of Pacific halibut (Hippoglossus stenolepis) in the Gulf of Alaska, established using satellite-transmitting archival tags. Fisheries Research, 92(1):63–69, July 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003657.

Lokkeborg:2003:RET

[Løk03] Svein Løkkeborg. Review and evaluation of three mitigation measures — bird-scaring line, underwater setting and line shooter — to reduce seabird bycatch in the north Atlantic longline fishery. Fisheries Research, 60(1):11–16, January 30, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602000784.

Longhurst:2002:BRM

[Lon02a] Alan Longhurst. Book review: Managing Methusalas: Life in the Slow Lane: Ecology and Conservation of Long-lived Marine Animals; John A. Musick (Ed.), American Fisheries Society Symposium, Vol. 23; American Fisheries Society, Bethesda, MD, 1999, 262 pages, ISBN 1-888569-15-8, $47. Fisheries Research, 57(3):311–312, September 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000334.

Longhurst:2002:MLR

[Lon02b] Alan Longhurst. Murphy’s law revisited: longevity as a factor in recruitment to fish populations. Fisheries Research, 56(2):125–131, May 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003514.

Longhurst:2003:RLE

[Lon03a] Alan Longhurst. Reply to the letter to the editor by Sherman et al. Fisheries Research, 64(2–3):205, November 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002303.

Longhurst:2003:SLM

[Lon03b] Alan Longhurst. The symbolism of large marine ecosystems. Fisheries Research, 61(1–3):1–6, March 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002539.

Longhurst:2006:SM

[Lon06] Alan Longhurst. The sustainability myth. Fisheries Research, 81(2–3):107–112, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002487.

Longhurst:2007:DCF

[Lon07] Alan Longhurst. Doubt and certainty in fishery science: Are we really headed for a global collapse of stocks? Fisheries Research, 86(1):1–5, August 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000409.

Long:2009:RFM

[Lon09] Le Kim Long. Regional fisheries management organisation with an endogenous minimum participation level for cooperation in straddling stock fisheries. Fisheries Research, 97(1–2):42–52, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608004025.

Lorenzen:2006:PMF

[Lor06] Kai Lorenzen. Population management in fisheries enhancement: Gaining key information from release experiments through use of a size-dependent mortality model.

Fisheries Research, 80(1):19–27, August 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001160.

Lajus:2007:FEB

[LOT07a] Julia Lajus, Henn Ojaveer, and Erki Tammiksaar. Fisheries at the Estonian Baltic Sea coast in the first half of the 19th century: What can be learned from the archives of Karl Ernst Baer? Fisheries Research, 87(2–3):126–136, November 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607001658.

Lotze:2007:RFF

[Lot07b] Heike K. Lotze. Rise and fall of fishing and marine resource use in the Wadden Sea, southern North Sea. Fisheries Research, 87(2–3):208–218, November 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000100.

Labropoulou:2000:COG

[LP00] M. Labropoulou and C. Papaconstantinou. Comparison of otolith growth and somatic growth in two macrourid fishes. Fisheries Research, 46(1–3):177–188, May 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001442.

Lourenco:2006:ESL

[LP06] Sílvia Lourenço and João Pereira. Estimating standardised landings per unit effort for an octopus mixed components fishery. Fisheries Research, 78(1):89–95, April 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003309.

Landa:2001:GAL

[LPDA01] J. Landa, P. Pereda, R. Duarte, and M. Azevedo. Growth of anglerfish (Lophius piscatorius and L. budegassa) in Atlantic Iberian waters. Fisheries Research, 51(2–3):363–376, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601002600.

Lucentini:2009:TCE

[LPG^{+}09] L. Lucentini, A. Palomba, L. Gigliarelli, G. Sgaravizzi, H. Lancioni, L. Lanfaloni, M. Natali, and F. Panara. Temporal changes and effective population size of an Italian isolated and supportive-breeding managed northern pike (Esox lucius) population. Fisheries Research, 96(2–3):139–147, March 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003354.

Linnane:2009:SDS

[LPHH09] Adrian Linnane, Shane Penny, Peter Hawthorne, and Matthew Hoare. Spatial differences in size of maturity and reproductive potential between inshore and offshore fisheries for southern rock lobster (Jasus edwardsii) in South Australia. Fisheries Research, 96(2–3):238–243, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608003846.

Lucentini:2006:MPI

[LPL^{+}06] Livia Lucentini, Antonella Palomba, Hovirag Lancioni, Lilia Gigliarelli, Mauro Natali, and Fausto Panara. Microsatellite polymorphism in Italian populations of northern pike (Esox lucius L.). Fisheries Research, 80(2–3): 251–262, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606001421.

Landa:2002:GPF

[LPP02] J. Landa, N. Pérez, and C. Piñeiro. Growth patterns of the four spot megrim (Lepidorhombus boscii) in the northeast Atlantic. Fisheries Research, 55(1–3):141–152, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003022.

Landa:2008:MBW

[LQD^{+}08] J. Landa, I. Quincoces, R. Duarte, A. C. Fariña, and H. Dupouy. Movements of black and white anglerfish (Lophius budegassa and L. piscatorius) in the northeast Atlantic. Fisheries Research, 94(1):1–12, October 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001197.

Lawson:2000:SDM

[LR00a] Gareth L. Lawson and George A. Rose. Seasonal distribution and movements of coastal cod (Gadus morhua L.) in Placentia Bay, Newfoundland. Fisheries Research, 49 (1):61–75, November 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783600001879.

Lloret:2000:CCG

[LR00b] Josep Lloret and Hans-Joachim Rätz. Condition of cod (Gadus morhua) off Greenland during 1982–1998. Fisheries Research, 48(1):79–86, August 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001119.

Larsen:2008:ICC

[LR08] Roger B. Larsen and Lasse Rindahl. Improved catch on cod (Gadus morhua), haddock (Melanogrammus aeglefinus) and Greenland halibut (Reinhardtius hippoglossoides) with a new hauling technique in the Norwegian mechanized bottom longline fishery. Fisheries Research, 94(2):160–165, November 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360800115X.

Leaf:2007:STS

[LRBH07] Robert T. Leaf, Laura Rogers-Bennett, and Peter L. Haaker. Spatial, temporal, and size-specific variation in mortality estimates of red abalone, Haliotis rufescens, from mark-recapture data in California. Fisheries Research, 83(2–3):341–350, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003729.

Lilja:2008:OSE

[LRC^{+}08] Juha Lilja, Teri Ridley, George M. W. Cronkite, Hermann J. Enzenhofer, and John A. Holmes. Optimizing sampling effort within a systematic design for estimating abundant escapement of sockeye salmon (Oncorhynchus nerka) in their natal river. Fisheries Research, 90(1–3):118–127, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360700255X.

LeFeuvre:2000:ASI

[LRG^{+}00] P. LeFeuvre, G. A. Rose, R. Gosine, R. Hale, W. Pearson, and R. Khan. Acoustic species identification in the Northwest Atlantic using digital image processing. Fisheries Research, 47(2–3):137–147, July 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360000165X.

Lester:2009:MSM

[LRW09] R. J. G. Lester, S. E. Rawlinson, and L. C. Weaver. Movement of sea mullet Mugil cephalus as indicated by a parasite. Fisheries Research, 96(2–3):129–132, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003330.

Leonardos:2000:AGM

[LS00] Ioannis Leonardos and Apostolos Sinis. Age, growth and mortality of Atherina boyeri Risso, 1810 (Pisces: Atherinidae) in the Mesolongi and Etolikon lagoons (W. Greece). Fisheries Research, 45(1):81–91, February 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000971.

Lessa:2004:AGB

[LSH04] Rosângela Lessa, Francisco M. Santana, and Fábio H. Hazin. Age and growth of the blue shark Prionace glauca (Linnaeus, 1758) off northeastern Brazil. Fisheries Research, 66(1):19–30, January 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001930.

Lowry:2006:RBB

[LSW06] Michael Lowry, Aldo Steffe, and Danielle Williams. Relationships between bait collection, bait type and catch: a comparison of the NSW trailer-boat and gamefishtournament fisheries. Fisheries Research, 78(2–3):266–275, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606000245.

Lowry:2007:LLR

[LWM07] Michael Lowry, Danielle Williams, and Yola Metti. Lunar landings — relationship between lunar phase and catch rates for an Australian gamefish-tournament fishery. Fisheries Research, 88(1–3):15–23, December 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001634.

Lauth:2004:EDT

[LWS04] R. R. Lauth, W. W. Wakefield, and K. Smith. Estimating the density of thornyheads, Sebastolobus spp., using a towed video camera sled. Fisheries Research, 70(1):39–48, November 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604001456.

Lunneryd:2002:DLN

[LWW02] S.-G. Lunneryd, H. Westerberg, and M. Wahlberg. Detection of leader net by whitefish Coregonus lavaretus during varying environmental conditions. Fisheries Research, 54(3):355–362, February 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002715.

Loneragan:2006:NDR

[LYKH06] Neil R. Loneragan, Yimin Ye, Rob A. Kenyon, and Mick D. E. Haywood. New directions for research in prawn (shrimp) stock enhancement and the use of models in providing directions for research. Fisheries Research, 80(1):91–100, August 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600124X.

Lloret:2008:SPF

[LZC^{+}08] Josep Lloret, Núria Zaragoza, David Caballero, Toni Font, Margarida Casadevall, and Victòria Riera. Spearfishing pressure on fish communities in rocky coastal habitats in a Mediterranean marine protected area. Fisheries Research, 94(1):84–91, October 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002002.

Lloret:2008:BSI

[LZCR08] Josep Lloret, Núria Zaragoza, David Caballero, and Victòria Riera. Biological and socioeconomic implications of recreational boat fishing for the management of fishery resources in the marine reserve of Cap de Creus (NW Mediterranean). Fisheries Research, 91(2–3):252–259, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607003475.

Lee:2001:STS

[LZS01] Chun-Woo Lee, Chang-Ik Zhang, and Hyeon-Ok Shin. Simplified trawl system modeling and design of a depth control system using fuzzy logic. Fisheries Research, 53(1): 83–94, September 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783600002642.

Liu:2009:UAR

[LZZN09] Q. Liu, X. M. Zhang, P. D. Zhang, and S. A. Nwafili. The use of alizarin red S and alizarin complexone for immersion marking Japanese flounder Paralichthys olivaceus (T.). Fisheries Research, 98(1–3):67–74, June 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000903.

Millar:2004:RP

[MA04] Russell B. Millar and Marti J. Anderson. Remedies for pseudoreplication. Fisheries Research, 70(2–3):397–407, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360400181X.

Mendiola:2008:VDI

[MÁ08] Diego Mendiola and Paula Álvarez. Validation of daily increments in the otolith microstructure of Northeast Atlantic mackerel fish larvae. Fisheries Research, 89(3):300–304, March 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607002986.

Martinez-Aguilar:2005:NML

[MAASMB05] Susana Martínez-Aguilar, Francisco Arreguín-Sánchez, and Enrique Morales-Bojórquez. Natural mortality and life history stage duration of Pacific sardine (Sardinops caeruleus) based on gnomonic time divisions. Fisheries Research, 71(1):103–114, January 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000839.

Mendiola:2006:ETD

[MAC^{+}06] D. Mendiola, P. Alvarez, U. Cotano, E. Etxebeste, and A. Martínez de Murguia. Effects of temperature on development and mortality of Atlantic mackerel fish eggs. Fisheries Research, 80(2–3):158–168, September 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001755.

Murta:2008:OMH

[MACS08] Alberto G. Murta, Pablo Abaunza, Fátima Cardador, and Francisco Sánchez. Ontogenic migrations of horse mackerel along the Iberian coast. Fisheries Research, 89(2):186–195, February 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607002469.

Martinez-Aguilar:2009:CHR

[MAdAMASCM09] Susana Martínez-Aguilar, Juan Antonio de Anda-Montañez, Francisco Arreguín-Sánchez, and Miguel Ángel Cisneros-Mata. Constant harvest rate for the Pacific sardine (Sardinops caeruleus) fishery in the Gulf of California based on catchability-at-length estimations. Fisheries Research, 99(2):74–82, August 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001064.

Miranda:2000:ASP

[MAG00] L. E. Miranda, A. A. Agostinho, and L. C. Gomes. Appraisal of the selective properties of gill nets and implications for yield and value of the fisheries at the Itaipu Reservoir, Brazil–Paraguay. Fisheries Research, 45(2):105–116, March 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699001149.

Magnusson:2001:DSO

[Mag01] Jutta V. Magnússon. Distribution and some other biological parameters of two morid species Lepidion eques (Günther, 1887) and Antimora rostrata (Günther, 1878) in Icelandic waters. Fisheries Research, 51(2–3):267–281, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360100251X.

Morello:2009:FNN

[MAG^{+}09] Elisabetta B. Morello, Bruno Antolini, Maria Emilia Gramitto, R. James A. Atkinson, and Carlo Froglia. The fishery for Nephrops norvegicus (Linnaeus, 1758) in the central Adriatic Sea (Italy): Preliminary observations comparing bottom trawl and baited creels. Fisheries Research, 95(2–3):325–331, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003238.

Mainardi:2009:SPM

[Mai09] Stefano Mainardi. A surplus production model with latent truncation and species targeting, with an application to Papua New Guinean fisheries. Fisheries Research, 95(2–3):296–308, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003214.

Morato:2001:LWR

[MAL^{+}01] T. Morato, P. Afonso, P. Lourinho, J. P. Barreiros, R. S. Santos, and R. D. M. Nash. Length–weight relationships for 21 coastal fish species of the Azores, north-eastern Atlantic. Fisheries Research, 50(3):297–302, March 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002150.

Macchi:2002:RBD

[MAL02] Gustavo J. Macchi, Eduardo M. Acha, and Carlos A. Lasta. Reproduction of black drum (Pogonias cromis) in the Río de la Plata estuary, Argentina. Fisheries Research, 59(1–2):83–92, December 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601004106.

Mandima:2000:STV

[Man00] Jimmiel J. Mandima. Spatial and temporal variations in the food of the sardine Limnothrissa miodon (Boulenger, 1906) in Lake Kariba, Zimbabwe. Fisheries Research, 48(2):197–203, September 1, 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001612.

Milessi:2005:MTL

[MAND05] Andrés C. Milessi, Hugo Arancibia, Sergio Neira, and Omar Defeo. The mean trophic level of Uruguayan landings during the period 1990–2001. Fisheries Research, 74(1–3): 223–231, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783605000445.

Markaida:2006:FFJ

[Mar06a] Unai Markaida. Food and feeding of jumbo squid Dosidicus gigas in the Gulf of California and adjacent waters after the 1997–98 El Niño event. Fisheries Research, 79(1–2):16– 27, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606000634.

Markaida:2006:PSR

[Mar06b] Unai Markaida. Population structure and reproductive biology of jumbo squid Dosidicus gigas from the Gulf of California after the 1997–1998 El Niño event. Fisheries Research, 79(1–2):28–37, June 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000646.

Marshall:2007:CPP

[Mar07a] N. A. Marshall. Can policy perception influence social resilience to policy change? Fisheries Research, 86(2–3): 216–227, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783607001294.

Martin:2007:RBE

[Mar07b] R. Aidan Martin. A review of behavioural ecology of whale sharks (Rhincodon typus). Fisheries Research, 84(1):10–16, March 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360600395X.

Matveev:2007:ABS

[Mat07] Vladimir F. Matveev. Assessing the biomass of small fish with a split-beam sonar in the Murray River, Australia. Fisheries Research, 88(1–3):139–145, December 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001695.

Maunder:2003:ITD

[Mau03] Mark N. Maunder. Is it time to discard the Schaefer model from the stock assessment scientist’s toolbox?

Fisheries Research, 61(1–3):145–149, March 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002734.

Misund:2000:SMS

[MB00] Ole Arve Misund and Arvid K. Beltestad. Survival of mackerel and saithe that escape through sorting grids in purse seines. Fisheries Research, 48(1):31–41, August 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001181.

Morel:2004:AWB

[MB04] G. M. Morel and S. F. Bossy. Assessment of the whelk (Buccinum undatum L.) population around the Island of Jersey, Channel Isles. Fisheries Research, 68(1–3):283–291, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604000268.

McConney:2007:LCM

[MB07] Patrick McConney and Roland Baldeo. Lessons in comanagement from beach seine and lobster fisheries in Grenada. Fisheries Research, 87(1):77–85, October 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001506.

Morton:2008:CME

[MB08] Richard Morton and Mark Bravington. Comparison of methods for estimating age composition with application to southern bluefin tuna (Thunnus maccoyii). Fisheries Research, 93(1–2):22–28, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360800057X.

Morales-Bojorquez:2001:RSA

[MBCMNMHH01] Enrique Morales-Bojórquez, Miguel Angel Cisneros-Mata, Manuel O. Nevárez-Martínez, and Agustín Hernández-Herrera. Review of stock assessment and fishery biology of Dosidicus gigas in the Gulf of California, Mexico. Fisheries Research, 54(1):83–94, December 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003721.

Morley:2005:DOI

[MBDM05] Simon A. Morley, Mark Belchier, Judith Dickson, and Thérèse Mulvey. Daily otolith increment validation in larval mackerel icefish, Champsocephalus gunnari. Fisheries Research, 75(1–3):200–203, September 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001311.

Merigot:2007:MCS

[MBG^{+}07] Bastien Mérigot, Jacques A. Bertrand, Jean-Claude Gaertner, Jean-Pierre Durbec, Nabila Mazouni, and Claude Manté. The multi-component structuration of the species diversity of groundfish assemblages of the east coast of Corsica (Mediterranean Sea): Variation according to the bathymetric strata. Fisheries Research, 88(1–3):120–132, December 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607001932.

Mitchell:2002:DMR

[MBGW02] R. W. Mitchell, S. J. Blight, D. J. Gaughan, and I. W. Wright. Does the mortality of released Sardinops sagax increase if rolled over the headline of a purse seine net? Fisheries Research, 57(3):279–285, September 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100354X.

Morgan:2002:WEC

[MBK02] M. J. Morgan, W. B. Brodie, and D. W. Kulka. Was over-exploitation the cause of the decline of the American plaice stock off Labrador and northeast Newfoundland? Fisheries Research, 57(1):39–49, July 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003319.

Millar:2004:MBH

[MBM04] Russell B. Millar, Matt K. Broadhurst, and William G. Macbeth. Modelling between-haul variability in the size selectivity of trawls. Fisheries Research, 67(2):171–181, April 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603002923.

MacKenzie:2007:MDS

[MBO^{+}07] Brian R. MacKenzie, Maibritt Bager, Henn Ojaveer, Kenneth Awebro, Ulla Heino, Poul Holm, and Aadu Must. Multi-decadal scale variability in the eastern Baltic cod fishery 1550–1860 — evidence and causes. Fisheries Research, 87(2–3):106–119, November 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001622.

McQuaid:2006:ESO

[MBR06] N. McQuaid, R. P. Briggs, and D. Roberts. Estimation of the size of onset of sexual maturity in Nephrops norvegicus (L.). Fisheries Research, 81(1):26–36, October 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002244.

Mahevas:2008:CAL

[MBT08] Stéphanie Mahévas, Lise Bellanger, and Verena M. Trenkel. Cluster analysis of linear model coefficients under contiguity constraints for identifying spatial and temporal fishing effort patterns. Fisheries Research, 93(1–2):29–38, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608000581.

Misund:2000:RFS

[MC00a] Ole Arve Misund and Janet Coetzee. Recording fish schools by multi-beam sonar: potential for validating and supplementing echo integration recordings of schooling fish. Fisheries Research, 47(2–3):149–159, July 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001661.

Mulligan:2000:CCS

[MC00b] T. J. Mulligan and D. G. Chen. Comment on ‘Can stationary bottom split-beam hydroacoustics be used to measure fish swimming speed in situ?’ by Arrhenius et al. Fisheries Research, 49(1):93–96, November 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002605.

Morsan:2004:AGM

[MC04] Enrique Morsán and Néstor F. Ciocco. Age and growth model for the southern geoduck, Panopea abbreviata, off Puerto Lobos (Patagonia, Argentina). Fisheries Research, 69(3):343–348, October 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001481.

McClatchie:2005:LTS

[MC05] S. McClatchie and R. F. Coombs. Low target strength fish in mixed species assemblages: the case of orange roughy. Fisheries Research, 72(2–3):185–192, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604003261.

Morrison:2006:EAS

[MC06] M. Morrison and G. Carbines. Estimating the abundance and size structure of an estuarine population of the sparid Pagrus auratus, using a towed camera during nocturnal periods of inactivity, and comparisons with conventional sampling techniques. Fisheries Research, 82(1–3):150–161, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360600244X.

McIlwain:2005:SVA

[MCAO^{+}05] J. L. McIlwain, M. R. Claereboudt, H. S. Al-Oufi, S. Zaki, and J. S. Goddard. Spatial variation in age and growth of the kingfish (Scomberomorus commerson) in the coastal waters of the Sultanate of Oman.

Fisheries Research, 73(3):283–298, July 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000676.

Montgomery:2007:PDA

[MCB^{+}07] S. S. Montgomery, A. J. Courtney, C. Blount, J. Stewart, D. J. Die, M. Cosgrove, and M. F. O’Neill. Patterns in the distribution and abundance of female eastern king prawns, Melicertus plebejus (Hess, 1865), capable of spawning and reproductive potential in waters off eastern Australia. Fisheries Research, 88(1–3):80–87, December 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001841.

McIntyre:2001:BRF

[McI01] A. D. McIntyre. Book review: Forage Fishes in Marine Ecosystems: University of Alaska Sea Grant, AK-SG-97-01, Fairbanks, 1997, 816 pages, hardback, ISBN 1-56612-049-7 ($40). Fisheries Research, 51(1):94–95, April 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600003118.

McIntyre:2002:BRA

[McI02a] A. D. McIntyre. Book review: Aquatic Habitat Assessment: Bain, M. B., Stevenson, N. J.; American Fisheries Society, Bethesda, MD, USA, 1999, 224 pages, paperback, ISBN 1-888569-18-2. Fisheries Research, 57(2):207–208, August 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602000322.

McIntyre:2002:SN

[McI02b] A. D. McIntyre. Shorter notices. Fisheries Research, 59(1–2):301–302, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001315.

McIntyre:2002:BRD

[McI02c] Alasdair D. McIntyre. Book review: Dynamics of the Bering Sea: T. R. Loughlin, K. Ohtani (Eds.); University of Alaska Sea Grant, AK-SG-99-03, Fairbanks, 1999, 838 pages, hardcover, ISBN 1-56612-062-4, $40.00. Fisheries Research, 56(2):215–217, May 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000048.

McIntyre:2002:BRMb

[McI02d] Alasdair D. McIntyre. Book review: Managing fisheries: Europe’s Southern Waters: Management Issues and Practices; David Symes (Ed.); Fishing News Books, Oxford, 1999, 198 pages, hard cover, ISBN 0-85238-273-1, £75. Fisheries Research, 58(3):407–409, November 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001650.

McIntyre:2002:BRMa

[McI02e] Alasdair D. McIntyre. Book review: Marine protected areas: tools for sustaining ocean ecosystems, Report of the Committee on Evaluation, Design, and Monitoring of Marine Reserves and Protected Areas in the United States; National Academy Press, Maryland, 2001, 272 pages, hard cover, ISBN 0-309-07286-7, $42.95. Fisheries Research, 58(2):249–250, November 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001285.

McIntyre:2002:BRS

[McI02f] Alasdair D. McIntyre. Book review: Sustaining Marine Fisheries: Report of the Committee on Ecosystem Management for Sustainable Marine Fisheries, National Academy Press, Washington, DC, 1999, 164 pages, hardback, £17.95, ISBN 0-309-05526-1. Fisheries Research, 57(1):99–100, July 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602000358.

McIntyre:2003:BRE

[McI03a] A. D. McIntyre. Book review: Effects of Fishing on Non-target Species and Habitats: M. J. Kaiser, S. J. de Groot (Eds.); Blackwell Science Ltd., 2002, 399 pages, hardback, ISBN 0-632-05355-0, £69.50. Fisheries Research, 64(1):95–96, October 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001279.

McIntyre:2003:BRR

[McI03b] A. D. McIntyre. Book review: Responsible Fisheries in the Marine Ecosystem: M. Sinclair, G. Valdimarsson (Eds.); FAO and CAB International, Wallingford, 2003, 448 pages, hardback, ISBN 0 85199 633 7 (CAB International), ISBN 925 104767 7 (FAO), US $140. Fisheries Research, 64(2–3):207–209, November 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002285.

McIntyre:2003:E

[McI03c] A. D. McIntyre. Editorial. Fisheries Research, 60(1):1, January 30, 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602002552.

McIntyre:2003:EIA

[McI03d] A. D. McIntyre. Environmental interactions of aquaculture. Fisheries Research, 62(3):235, June 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000626.

McIntyre:2003:BRO

[McI03e] Alasdair D. McIntyre. Book review: Ocean Yearbook 15: E. M. Borgese, A. Chircop and M. McConnell (Eds.); University of Chicago Press, 2001, 1018 pages, hardback, ISBN 0-226-06618-5, US $130. Fisheries Research, 64(1):92–93, October 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603002194.

McIntyre:2005:BRI

[McI05a] Alasdair D. McIntyre. Book review: Ichthyology Handbook, B. G. Kapoor, Bhavna Khanna. Springer-Verlag (2004), 1059 pp., US $325 (hardback), ISBN: 3-540-42854-2. Fisheries Research, 72(1):134–135, April 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000330.

McIntyre:2005:BRSa

[McI05b] Alasdair D. McIntyre. Book review: Stock Enhancement and Sea Ranching, K. M. Leber, S. Kitada, H. L. Blankenship, T. Svasand (Eds.). Blackwell Publishing Ltd. (2004), 576, £99.50, ISBN 140511194. Fisheries Research, 72(1):133–134, April 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000329.

McIntyre:2005:BRSb

[McI05c] Alasdair D. McIntyre. Book review: Successful Fisheries Management, S. Cunningham, T. Bostock (Eds.). Eburon Academic Publishers (2005), (Paperback, 240 pp., Price EUR 25, US $27.50, ISBN 90-5972-061-X). Fisheries Research, 76(1):163–165, October 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001748.

McIntyre:2005:KCE

[McI05d] Alasdair D. McIntyre. K. E. Carpenter (Ed.), The Living Marine Resources of the Western Central Atlantic, FAO Species Identification Guide for Fishery Purposes, American Society of Ichthyologists and Herpetologists, vol. 1–3, Special Publication No. 5, Rome, FAO, 2002, 2127 pp. (paperback), ISBN 92-5-104825-8. Fisheries Research, 72(2–3):361–362, May 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000597.

McIntyre:2006:BRM

[McI06a] Alasdair D. McIntyre. Book review: Mangrove Biodiversity, K. Kathirestan, S. Z. Qasim. Biodiversity of Mangrove Ecosystems. Hindustan Publishing Corporation, India (2005). 268 pp., Hardback, Price: Rs 875.00, US $45.00, ISBN: 81-7075-079-2. Fisheries Research, 80(2–3):352–353, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001652.

McIntyre:2006:BRG

[McI06b] Alasdair D. McIntyre. Book review: The Gulf of Alaska: Biology and Oceanography, P. R. Mundy (Ed.). Alaska Sea Grant College Programme, University of Alaska Fairbanks (2005), 214 pp., Hardback, $25.00, ISBN: 1-56612 - 090 - X. Fisheries Research, 80(2–3):350, September 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001676.

McIntyre:2007:BRG

[McI07a] Alasdair D. McIntyre. Book review: The Gulf Stream, B. Voituriez. UNESCO, Paris (2006), 223 pp., Soft cover, Price: EUR 18.00, ISBN 10: 92-3-103995-4. Fisheries Research, 88(1–3):155, December 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001944.

McIntyre:2007:BRB

[McI07b] Alasdair D. McIntyre. Book review: Von Brandt’s Fish Catching Methods of the World, fourth ed., O. Gabriel, K. Lange, E. Dahm, T. Wendt (Eds.). Blackwell Publishing

Ltd. (2005), http://www.blackwellpublishing.com, 536

pp., Hardback, Price: £99.50, available at: ISBN: 0-85238-280-4. Fisheries Research, 85(1–2):232, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000501.

McIntyre:2007:E

[McI07c] Alasdair D. McIntyre. Editorial. Fisheries Research, 87 (2–3):101, November 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783607001889.

McIntyre:2008:BRO

[McI08] Alasdair D. McIntyre. Book review: OCEANS PAST: Management Insights from the History of Marine Animal Populations, D. J. Starkey, P. Holm, M. Barnard (Eds.). Earthscan (2008), p. 223, Hardback, ISBN-13:978-1-84407-527-

0. Fisheries Research, 93(1–2):253–254, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001392.

McIntyre:2009:BRFa

[McI09a] Alasdair D. McIntyre. Book review: Feeding and Digestive Functions of Fishes, J. E. Cyrino, D. P. Bureau, B. G. Kapoor (Eds.). Science Publishers, Enfield, NH, USA (2008), 575 pp., Hardback, Price US$88, ISBN: 978-1-57808-375-6. Fisheries Research, 99(2):134–135, August 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360900112X.

McIntyre:2009:BRFb

[McI09b] Alasdair D. McIntyre. Book review: Fish Behaviour, C. Magnhagen, V. A. Braithwaite, E. Forsgren, B. G. Kapoor (Eds.). Science Publishers, Enfield, NH, USA (2008), 648 pp., Hardback, Price: $90, ISBN: 978-1-57808-435-7. Fisheries Research, 99(2):135, August 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001131.

McIntyre:2009:BRI

[McI09c] Alasdair D. McIntyre. Book review: Introduction to Wildlife and Fisheries, D. W. Willis, C. G. Scalet, L. D. Flake (Eds.). W. H. Freeman and Company, NY (2008), 461pp., Hardback, £65.99, ISBN: 13:978-1-4292-0446-0. Fisheries Research, 99(2):135, August 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001143.

McIntyre:2009:BRR

[McI09d] Alasdair D. McIntyre. Book review: Reproductive Biology of Crustaceans, E. Mente (Ed.). Science Publishers, Enfield, NH, USA (2008), 549 pp., $70 (hardback), ISBN: 978-1-57808-529-3. Fisheries Research, 99(2):134, August 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001118.

Maguire:2002:EDU

[MCJB02] Julie A. Maguire, Alison Coleman, Stuart Jenkins, and Gavin M. Burnell. Effects of dredging on undersized scallops. Fisheries Research, 56(2):155–165, May 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003228.

Madurell:2004:CSF

[MCL04] T. Madurell, J. E. Cartes, and M. Labropoulou. Changes in the structure of fish assemblages in a bathyal site of the Ionian Sea (eastern Mediterranean). Fisheries Research, 66(2–3):245–260, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002054.

MacKenzie:2008:PBT

[MCM^{+}08] K. MacKenzie, N. Campbell, S. Mattiucci, P. Ramos, A. L. Pinto, and P. Abaunza. Parasites as biological tags for stock identification of Atlantic horse mackerel Trachurus trachurus L. Fisheries Research, 89(2):136–145, February 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607002408.

Mouat:2001:PDI

[MCP01] Beth Mouat, Martin A. Collins, and Joost Pompert. Patterns in the diet of Illex argentinus (Cephalopoda:

Ommastrephidae) from the Falkland Islands jigging fishery. Fisheries Research, 52(1–2):41–49, June 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002296.

McPherson:2002:ABA

[McP02] Ross McPherson. Assessment of T bar anchor tags for marking the Blue Swimmer Crab Portunus pelagicus (L.). Fisheries Research, 54(2):209–216, January 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600003039.

Moreno:2005:RBS

[MCS05] Teresa Moreno, José Juan Castro, and Juan Socorro. Reproductive biology of the sand smelt Atherina presbyter Cuvier, 1829 (Pisces: Atherinidae) in the central-east Atlantic. Fisheries Research, 72(1):121–131, April 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001572.

Mormede:2001:TED

[MD01] Sophie Mormede and Ian M. Davies. Trace elements in deep-water fish species from the Rockall Trough. Fisheries Research, 51(2–3):197–206, May 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002454.

Milessi:2002:LTI

[MD02] Andrés C. Milessi and Omar Defeo. Long-term impact of incidental catches by tuna longlines: the black escolar (Lepidocybium flavobrunneum) of the southwestern Atlantic Ocean. Fisheries Research, 58(2):203–213, November 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601003654.

Mullowney:2009:DPI

[MD09] D. R. Mullowney and E. G. Dawe. Development of performance indices for the Newfoundland and Labrador snow crab (Chionoecetes opilio) fishery using data from a vessel monitoring system. Fisheries Research, 100(3):248–254, November 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609002252.

Monteiro:2005:ACS

[MDGR05] Leandro R. Monteiro, Ana Paula M. Di Beneditto, Luis H. Guillermo, and Luis A. Rivera. Allometric changes and shape differentiation of sagitta otoliths in sciaenid fishes. Fisheries Research, 74(1–3):288–299, August 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000834.

Montano:2005:AAF

[MDJR05] Orlando J. Ferrer Montaño, Eric D. Dibble, Donald C. Jackson, and Kirk R. Rundle. Angling assessment of the fisheries of Humacao Natural Reserve lagoon system, Puerto Rico. Fisheries Research, 76(1):81–90, October 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605001669.

Maynou:2003:ACP

[MDS03] F. Maynou, M. Demestre, and P. Sánchez. Analysis of catch per unit effort by multivariate analysis and generalised linear models for deep-water crustacean fisheries off Barcelona (NW Mediterranean). Fisheries Research, 65(1–3):257–269, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603002479.

Milner:2003:NCS

[MEA^{+}03] N. J. Milner, J. M. Elliott, J. D. Armstrong, R. Gardiner, J. S. Welton, and M. Ladle. The natural control of salmon and trout populations in streams. Fisheries Research, 62(2):111–125, May 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001571.

Mehner:2006:PHT

[Meh06] Thomas Mehner. Prediction of hydroacoustic target strength of vendace (Coregonus albula) from concurrent trawl catches. Fisheries Research, 79(1–2):162–169, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606000750.

Mesnil:2003:CSA

[Mes03] Benoit Mesnil. The Catch-Survey Analysis (CSA) method of fish stock assessment: an evaluation using simulated data. Fisheries Research, 63(2):193–212, August 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000729.

Meyer:2007:ISO

[Mey07] Carl G. Meyer. The impacts of spear and other recreational fishers on a small permanent marine protected area and adjacent pulse fished area. Fisheries Research, 84(3):301– 307, May 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606003894.

Machado:2000:TAB

[MF00] Pedro Bordalo Machado and Ivone Figueiredo. A technique for ageing the birdbeak dogfish (Deania calcea Lowe, 1839) from dorsal spines. Fisheries Research, 45(1):93–98, February 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783699001009.

Molares:2003:DPC

[MF03] José Molares and Juan Freire. Development and perspectives for community-based management of the goose barnacle (Pollicipes pollicipes) fisheries in Galicia (NW Spain). Fisheries Research, 65(1–3):485–492, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002637.

Mandelman:2007:EST

[MF07] John W. Mandelman and Marianne A. Farrington. The estimated short-term discard mortality of a trawled elasmobranch, the spiny dogfish (Squalus acanthias). Fisheries Research, 83(2–3):238–245, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003456.

Moura:2009:KDE

[MF09] Olga Moura and Ana Cláudia Fernandes. Kriging density estimates for the bluemouth rockfish, Helicolenus dactylopterus (Scorpaenidae) off the southern Portuguese coast. Fisheries Research, 95(1):112–124, January 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002336.

Morello:2005:HDD

[MFAM05] Elisabetta B. Morello, Carlo Froglia, R. James A. Atkinson, and P. Geoffrey Moore. Hydraulic dredge discards of the clam (Chamelea gallina) fishery in the western Adriatic Sea, Italy. Fisheries Research, 76(3):430–444, December 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002079.

McCarthy:2008:EIB

[MFB08] J. L. McCarthy, K. D. Friedland, and J. Brodziak. Enhancement of image-based fecundity methods: Gravimetric sampling at sea and safer sample preservation. Fisheries Research, 93(1–2):47–53, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000635.

Mattiucci:2008:ASL

[MFC^{+}08] S. Mattiucci, V. Farina, N. Campbell, K. MacKenzie, P. Ramos, A. L. Pinto, P. Abaunza, and G. Nascetti. Anisakis spp. larvae (Nematoda: Anisakidae) from Atlantic horse mackerel: Their genetic identification and use as biological tags for host stock characterization.

Fisheries Research, 89(2):146–151, February 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360700241X.

McFarlane:2000:SOS

[MG00] G. A. McFarlane and A. Geffen. Section 3: Otoliths in studies of populations. Fisheries Research, 46(1–3):189– 190, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600001454.

Murray:2003:CUA

[MG03a] Alexander G. Murray and Daniel J. Gaughan. Corrigendum to “Using an age-structured model to simulate the recovery of the Australian pilchard (Sardinops sagax) population following epidemic mass mortality”: [Fish. Res. 60 (2003) 415–426]. Fisheries Research, 63(3):453, September 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001498. See [MG03b].

Murray:2003:UAS

[MG03b] Alexander G. Murray and Daniel J. Gaughan. Using an age-structured model to simulate the recovery of the Australian pilchard (Sardinops sagax) population following epidemic mass mortality. Fisheries Research, 60(2–3):415–426, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602001364. See cor-

rigendum [MG03a].

Muths:2009:GPS

[MGJB09] D. Muths, P. Grewe, C. Jean, and J. Bourjea. Genetic population structure of the swordfish (Xiphias gladius) in the southwest Indian Ocean: Sex-biased differentiation, congruency between markers and its incidence in a way of stock assessment. Fisheries Research, 97(3):263–269, May 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000691.

Motta:2005:SSC

[MGNB05] Fabio S. Motta, Otto B. F. Gadig, Rafael C. Namora, and Francisco M. S. Braga. Size and sex compositions, length–weight relationship, and occurrence of the Brazilian sharpnose shark, Rhizoprionodon lalandii, caught by artisanal fishery from southeastern Brazil. Fisheries Research, 74(1–3):116–126, August 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001116.

Mattiangeli:2002:VVA

[MGR^{+}02] V. Mattiangeli, P. Galvin, A. W. Ryan, J. Mork, and T. F. Cross. VNTR variability in Atlantic poor cod (Trisopterus minutus minutus) throughout its range: single locus minisatellite data suggest reproductive isolation for the Faroe Bank population. Fisheries Research, 58(2):185–191, November 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601003861.

Maartens:2002:SSR

[MGS02] Lima Maartens, Kjell A. Gamst, and Peter M. Schneider. Size selection and release of juvenile monkfish Lophius vomerinus using rigid sorting grids. Fisheries Research, 57(1):75–88, July 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783601003381.

Macher:2008:CBA

[MGTB08] Claire Macher, Olivier Guyader, Catherine Talidec, and Michel Bertignac. A cost–benefit analysis of improving trawl selectivity in the case of discards: the Nephrops norvegicus fishery in the Bay of Biscay. Fisheries Research, 92(1):76–89, July 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000027.

Madsen:2001:DEG

[MH01] N. Madsen and K. E. Hansen. Danish experiments with a grid system tested in the North Sea shrimp fishery. Fisheries Research, 52(3):203–216, July 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002344.

Madsen:2002:ACE

[MH02] Niels Madsen and Renè Holst. Assessment of the cover effect in trawl codend selectivity experiments. Fisheries Research, 56(3):289–301, June 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003307.

Malvadkar:2008:PMS

[MH08] Urmila Malvadkar and Alan Hastings. Persistence of mobile species in marine protected areas. Fisheries Research, 91(1):69–78, May 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783607003153.

Madsen:2002:EWI

[MHF02] Niels Madsen, René Holst, and Leslie Foldager. Escape windows to improve the size selectivity in the Baltic cod trawl fishery. Fisheries Research, 57(3):223–235, September 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601003551.

Meeuwig:2006:QNT

[MHK^{+}06] Jessica J. Meeuwig, Do Huu Hoang, Truong Si Ky, Suresh D. Job, and Amanda C. J. Vincent. Quantifying non-target seahorse fisheries in central Vietnam. Fisheries Research, 81(2–3):149–157, November 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002827.

Madsen:2001:KCN

[MHMP01] N. Madsen, K. E. Hansen, and T. Moth-Poulsen. The kite cover: a new concept for covered codend selectivity studies. Fisheries Research, 49(3):219–226, January 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002101.

Madureira:2005:IAR

[MHS^{+}05] L. Madureira, R. Habiaga, C. Soares, S. Weigert, C. Ferreira, D. Eliseire, and A. C. Duvoisin. Identification of acoustic records of the Argentinian calamar Illex argentinus (Castellanos, 1960) along the outer shelf and shelf break of the south and southeast coast of Brazil. Fisheries Research, 73(1–2):251–257, June 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000354.

Meyer:2001:DRP

[MHWL01] Carl G. Meyer, Kim N. Holland, Bradley M. Wetherbee, and Christopher G. Lowe. Diet, resource partitioning and gear vulnerability of Hawaiian jacks captured in fishing tournaments. Fisheries Research, 53(2):105–113, October 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360000285X.

Mills:2002:BRMa

[Mil02a] Derek Mills. Book review: Making Salmon: an Environmental History of the Northwest Fisheries Crisis; Joseph E. Taylor III (Ed.); University of Washington Press, Seattle, 421 pages, hardback, ISBN 0-295-97840-6, US $34.95. Fisheries Research, 56(2):213–215, May 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000036.

Mills:2002:BRMb

[Mil02b] Derek Mills. Book review: Migration of Freshwater Fishes, Martin C. Lucas and Etienne Baras; Blackwell Science Ltd., Oxford, 440 pages, hardback, ISBN 0-632-05754-8, £79.50. Fisheries Research, 58(2):250–251, November 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001297.

Miller:2003:AYC

[Mil03a] Robert J. Miller. Be-all-you-can-be management targets for Canadian lobster fisheries. Fisheries Research, 64(2–3):179–184, November 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002236.

Mills:2003:BRR

[Mil03b] Derek Mills. Book review: Recreational Fisheries, Ecological, Economic and Social Evaluation: Fish and Aquatic Resources, Vol. 8; Tony J. Pitcher, Charles Hollingworth (Eds.); Blackwell Science Ltd., Oxford, 271 pages, hardback, ISBN 0-632-06391-2 (£79.50). Fisheries Research, 63(2):295–296, August 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001280.

Mills:2006:RWE

[Mil06] Derek Mills. R. N. Williams (Ed.). Return to the River: Restoring Salmon to the Columbia River. Elsevier Academic Press, Amsterdam, 2005, 699 pp., Hardback, Price: £45, ISBN: 0-12-088414-3. Fisheries Research, 80(2–3):350–352, September 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001664.

Mills:2008:BRS

[Mil08] Derek Mills. Book review: Sea Trout. Biology, Conservation & Management, G. Harris, N. Milner (Eds.). Blackwell

Publishing, Oxford (2007), www.blackwellpublishing.com.

Fisheries Research, 89(1):97–99, January 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001968.

Maguire:2002:ERD

[MJB02] Julie A. Maguire, Stuart Jenkins, and Gavin M. Burnell. The effects of repeated dredging and speed of tow on undersized scallops. Fisheries Research, 58(3):367–377, November 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601003927.

Murray-Jones:2000:CBC

[MJS00] Sue Murray-Jones and Aldo S. Steffe. A comparison between the commercial and recreational fisheries of the surf clam, Donax deltoides. Fisheries Research, 44(3):219–233, January 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699000958.

McGarvey:2001:ASS

[MK01] R. McGarvey and M. A. Kinloch. An analysis of the sensitivity of stock biomass estimates derived from the daily egg production method (DEPM) to uncertainty in egg mortality rates. Fisheries Research, 49(3):303–307, January 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002083.

Margraf:2002:EFS

[MK02] F. J. Margraf and Carey T. Knight. Evaluation of fish sampling using rotenone in a navigation lock. Fisheries Research, 55(1–3):297–305, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002983.

McFarlane:2006:AGB

[MK06] G. A. McFarlane and J. R. King. Age and growth of big skate (Raja binoculata) and longnose skate (Raja rhina) in British Columbia waters. Fisheries Research, 78(2–3):169– 178, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606000440.

Miller:2009:DMR

[MK09] J. A. Miller and A. J. R. Kent. The determination of maternal run time in juvenile Chinook salmon (Oncorhynchus tshawytscha) based on Sr/Ca and ^{87}Sr/^{86}Sr within otolith cores. Fisheries Research, 95(2–3):373–378, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608003081.

Mori:2001:SDN

[MKB01] Junta Mori, Tsunemi Kubodera, and Norihisa Baba. Squid in the diet of northern fur seals, Callorhinus ursinus, caught in the western and central North Pacific Ocean. Fisheries Research, 52(1–2):91–97, June 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002338.

MacKenzie:2005:MGB

[MKGH05] K. MacKenzie, C. Kalavati, M. Gaard, and W. Hemmingsen. Myxosporean gall bladder parasites of gadid fishes in the North Atlantic: Their geographical distributions and an assessment of their economic importance in fisheries and mariculture. Fisheries Research, 76(3):454–465, December 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605002055.

Maunder:2004:ISC

[ML04a] Mark N. Maunder and Adam D. Langley. Integrating the standardization of catch-per-unit-of-effort into stock assessment models: testing a population dynamics model and using multiple data types. Fisheries Research, 70(2–3):389–395, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001808.

Mendoza:2004:BDA

[ML04b] Jeremy J. Mendoza and Asdrubal Larez. A biomass dynamics assessment of the southeastern Caribbean snapper– grouper fishery. Fisheries Research, 66(2–3):129–144, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603002261.

Maynou:2003:SSV

[MLC03] F. Maynou, J. Lleonart, and J. E. Cartes. Seasonal and spatial variability of hake (Merluccius merluccius L.) recruitment in the NW Mediterranean. Fisheries Research, 60(1):65–78, January 30, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000620.

Minami:2007:MSB

[MLCGRV07] M. Minami, C. E. Lennert-Cody, W. Gao, and M. Román-Verdesoto. Modeling shark bycatch: the zero-inflated negative binomial regression model with smoothing. Fisheries Research, 84(2):210–221, April 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003821.

Manson:2005:BSA

[MLH^{+}05] F. J. Manson, N. R. Loneragan, B. D. Harch, G. A. Skilleter, and L. Williams. A broad-scale analysis of links between coastal fisheries production and mangrove extent: a case-study for northeastern Australia. Fisheries Research, 74(1–3):69–85, August 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001098.

Macdonald:2009:CCR

[MLJT09] Paul Macdonald, Chevonne H. Laurenson, Arthur Johnson, and Leslie Tait. A comparison of catch rates of artificial lures from an automated handline fishery at Shetland, UK. Fisheries Research, 95(2–3):379–385, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608003263.

Mapstone:2008:MSE

[MLP^{+}08] B. D. Mapstone, L. R. Little, A. E. Punt, C. R. Davies, A. D. M. Smith, F. Pantus, A. D. McDonald, A. J. Williams, and A. Jones. Management strategy evaluation for line fishing in the Great Barrier Reef: Balancing conservation and multi-sector fishery objectives. Fisheries Research, 94(3):315–329, December 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002269.

Martin:2004:ADM

[MM04] Allen D. Martin and Michael J. Maceina. Assessment of detecting minimum length limit changes for crappie in two Alabama reservoirs. Fisheries Research, 68(1–3):293–303, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604000256.

Meka:2005:PRW

[MM05] Julie M. Meka and Stephen D. McCormick. Physiological response of wild rainbow trout to angling: impact of angling duration, fish size, body condition, and temperature. Fisheries Research, 72(2–3):311–322, May 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002553.

Marriott:2006:GIA

[MM06a] Ross J. Marriott and Bruce D. Mapstone. Geographic influences on and the accuracy and precision of age estimates for the red bass, Lutjanus bohar (Forsskål 1775): a large tropical reef fish. Fisheries Research, 80(2–3):322–328, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606001779.

Militelli:2006:SFS

[MM06b] M. I. Militelli and G. J. Macchi. Spawning and fecundity of striped weakfish, Cynoscion guatucupa, in the Río de la Plata estuary and adjacent marine waters, Argentina– Uruguay. Fisheries Research, 77(1):110–114, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002717.

Mueter:2006:UMS

[MM06c] Franz J. Mueter and Bernard A. Megrey. Using multi-species surplus production models to estimate ecosystem-level maximum sustainable yields. Fisheries Research, 81(2–3):189–201, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002785.

MacKenzie:2007:DNE

[MM07] Brian R. MacKenzie and Ransom A. Myers. The development of the northern European fishery for north Atlantic bluefin tuna Thunnus thynnus during 1900–1950. Fisheries Research, 87(2–3):229–239, November 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000355.

Macbeth:2007:IFV

[MMB^{+}07a] William G. Macbeth, Russell B. Millar, Matt K. Broadhurst, Charles W. Hewitt, and Michael E. L. Wooden. Intra-fleet variability in the size selectivity of a square-mesh trawl codend for school prawns (Metapenaeus macleayi). Fisheries Research, 86(2–3):92–98, September 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001099.

Marriott:2007:ASD

[MMB07b] R. J. Marriott, B. D. Mapstone, and G. A. Begg. Age-specific demographic parameters, and their implications for management of the red bass, Lutjanus bohar (Forsskål 1775): a large, long-lived reef fish. Fisheries Research, 83(2–3):204–215, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003419.

Melendy:2005:UPT

[MMH05] J. Melendy, G. McClelland, and T. Hurlbut. Use of parasite tags in delineating stocks of white hake (Urophycis tenuis) from the southern Gulf of St. Lawrence and Cape Breton Shelf. Fisheries Research, 76(3):392–400, December 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002110.

Montano-Moctezuma:2008:VCI

[MMLR08] Gabriela Montaño-Moctezuma, Hiram W. Li, and Philippe A. Rossignol. Variability of community interaction networks in marine reserves and adjacent exploited areas. Fisheries Research, 94(1):99–108, October 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002166.

Moser:2005:PEM

[MMLT05] Sean Moser, Donald Macintosh, Somboon Laoprasert, and Nualanong Tongdee. Population ecology of the mud crab Scylla olivacea: a study in the Ranong mangrove ecosystem, Thailand, with emphasis on juvenile recruitment and mortality. Fisheries Research, 71(1):27–41, January 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001900.

Morey:2003:WLR

[MMM^{+}03] G. Morey, J. Moranta, E. Massutí, A. Grau, M. Linde, F. Riera, and B. Morales-Nin. Weight–length relationships of littoral to lower slope fishes from the western Mediterranean. Fisheries Research, 62(1):89–96, April 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002503.

Moranta:2000:FCC

[MMN00a] Joan Moranta and Enric Massutıánd Beatriz Morales-Nin. Fish catch composition of the deep-sea decapod crustacean fisheries in the Balearic Islands (western Mediterranean). Fisheries Research, 45(3):253–264, April 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001198.

Mosegaard:2000:SOP

[MMN00b] Henrik Mosegaard and Beatriz Morales-Nin. Section 1: Otolith physiology and morphology. Fisheries Research, 46(1–3):3–4, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783600001272.

Moreno:2003:ADV

[MMN03] Teresa Moreno and Beatriz Morales-Nin. Age determination and validation on otoliths of the sand-smelt Atherina presbyter (Cuvier, 1829) (Pisces: Atherinidae) from the central-east Atlantic. Fisheries Research, 62(1):77–87, April 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602002527.

Machias:2002:VAS

[MMP^{+}02] Athanassios Machias, Eleni Maraveyia, Michalis Pavlidis, Stylianos Somarakis, and Pascal Divanach. Validation of annuli on scales and otoliths of common dentex (Dentex dentex). Fisheries Research, 54(2):287–294, January 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600003027.

Marsden:2009:RBA

[MMS09] A. Dale Marsden, Steven J. D. Martell, and U. Rashid Sumaila. Retrospective bioeconomic analysis of Fraser River sockeye salmon fishery management. Fisheries Research, 97(1–2):32–41, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0165783608004013.

Machias:2004:SDC

[MMV^{+}04] A. Machias, P. Maiorano, V. Vassilopoulou, C. Papaconstantinou, A. Tursi, and N. Tsimenides. Sizes of discarded commercial species in the eastern–central Mediterranean Sea. Fisheries Research, 66(2–3):213–222, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002029.

Montenegro:2009:IMA

[MMZ09] Carlos Montenegro, Mark N. Maunder, and Maximiliano Zilleruelo. Improving management advice through spatially explicit models and sharing information. Fisheries Research, 100(3):191–199, November 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002021.

Morales-Nin:2000:RGR

[MN00] Beatriz Morales-Nin. Review of the growth regulation processes of otolith daily increment formation. Fisheries Research, 46(1–3):53–67, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001338.

Morales-Nin:2001:MDW

[MN01] Beatriz Morales-Nin. Mediterranean deep-water fish age determination and age validation: the state of the art. Fisheries Research, 51(2–3):377–383, May 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002612.

Matsuda:2003:SSM

[MN03] Hiroyuki Matsuda and Katsuhiro Nishimori. A sizestructured model for a stock-recovery program for an exploited endemic fisheries resource. Fisheries Research, 60 (2–3):223–236, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0165783602001856.

Morales-Nin:2005:OMH

[MNBM05] Beatriz Morales-Nin, R. M. Bjelland, and Erlend Moksness. Otolith microstructure of a hatchery reared European hake (Merluccius merluccius). Fisheries Research, 74 (1–3):300–305, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783605000822.

Morales-Nin:2000:AGB

[MNM00] Enric Massutıánd Beatriz Morales-Nin and Joan Moranta. Age and growth of blue-mouth, Helicolenus dactylopterus (Osteichthyes: Scorpaenidae), in the western Mediterranean. Fisheries Research, 46(1–3):165–176, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001430.

Makinen:2000:BGN

[MNML00] Teemu S. Mäkinen, Eero Niemelä, Kjell Moen, and Reijo Lindström. Behaviour of gill-net and rod-captured Atlantic salmon (Salmo salar L.) during upstream migration and following radio tagging. Fisheries Research, 45(2):117–127, March 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699001071.

Miyahara:2006:AGH

[MOGG06] Kazutaka Miyahara, Taro Ota, Tsuneo Goto, and Shigeaki Gorie. Age, growth and hatching season of the diamond squid Thysanoteuthis rhombus estimated from statolith analysis and catch data in the western Sea of Japan. Fisheries Research, 80(2–3):211–220, September 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001500.

Miyahara:2005:CFD

[MOK^{+}05] Kazutaka Miyahara, Taro Ota, Nobuhisa Kohno, Yukio Ueta, and John R. Bower. Catch fluctuations of the diamond squid Thysanoteuthis rhombus in the Sea of Japan and models to forecast CPUE based on analysis of environmental factors. Fisheries Research, 72(1):71–79, April 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002656.

Moore:2003:SFC

[Moo03] P. G. Moore. Seals and fisheries in the Clyde Sea area (Scotland): traditional knowledge informs science. Fisheries Research, 63(1):51–61, July 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000031.

Masuda:2000:AGF

[MOOH00] Y. Masuda, T. Ozawa, O. Onoue, and T. Hamada. Age and growth of the flathead, Platycephalus indicus, from the coastal waters of west Kyushu, Japan.

Fisheries Research, 46(1–3):113–121, May 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001387.

Moses:2000:RAM

[Mos00] Boniface S. Moses. A review of artisanal marine and brackishwater fisheries of south-eastern Nigeria. Fisheries Research, 47(1):81–92, June 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001332.

Moth-Poulsen:2003:SVS

[MP03] Thomas Moth-Poulsen. Seasonal variations in selectivity of plaice trammel nets. Fisheries Research, 61(1–3):87–94, March 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602001868.

Maunder:2004:SCE

[MP04] Mark N. Maunder and André E. Punt. Standardizing catch and effort data: a review of recent approaches. Fisheries Research, 70(2–3):141–159, December 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001638.

Murta:2008:SIH

[MPA08] Alberto G. Murta, Ana Luísa Pinto, and Pablo Abaunza. Stock identification of horse mackerel (Trachurus trachurus) through the analysis of body shape. Fisheries Research, 89(2):152–158, February 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002421.

Macchi:2004:SEP

[MPE04] Gustavo J. Macchi, Marcelo Pájaro, and Martín Ehrlich. Seasonal egg production pattern of the Patagonian stock of Argentine hake (Merluccius hubbsi). Fisheries Research, 67(1):25–38, March 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002716.

Murillas:2008:MIS

[MPG^{+}08] A. Murillas, R. Prellezo, E. Garmendia, M. Escapa, C. Gallastegui, and A. Ansuategi. Multidimensional and intertemporal sustainability assessment: a case study of the Basque trawl fisheries. Fisheries Research, 91(2–3):222–238, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607003451.

Macchi:2006:ISA

[MPM^{+}06] Gustavo J. Macchi, Marcelo Pájaro, María I. Militelli, Natalia Radovani, and Lorena Rivas. Influence of size, age and maternal condition on the oocyte dry weight of Argentine hake (Merluccius hubbsi). Fisheries Research, 80(2–3):345–349, September 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001044.

Marchal:2007:LBF

[MPQ07] Paul Marchal, Jan-Jaap Poos, and Floor Quirijns. Linkage between fishers’ foraging, market and fish stocks density: Examples from some North Sea fisheries. Fisheries Research, 83(1):33–43, January 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003055.

Markaida:2004:AGM

[MQVSN04] Unai Markaida, Casimiro Quiñónez-Velázquez, and Oscar Sosa-Nishizaki. Age, growth and maturation of jumbo squid Dosidicus gigas (Cephalopoda: Ommastrephidae) from the Gulf of California, Mexico. Fisheries Research, 66(1):31–47, January 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360300184X.

Mello:2005:SVA

[MR05] L. G. S. Mello and G. A. Rose. Seasonal variation in abundance and stock composition of Atlantic cod (Gadus morhua L.) in Placentia Bay, Newfoundland, in relation to fisheries. Fisheries Research, 74(1–3):142–156, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605000895.

Mikulas:2008:HUG

[MR08a] Joseph J. Mikulas and Jay R. Rooker. Habitat use, growth, and mortality of post-settlement lane snapper (Lutjanus synagris) on natural banks in the northwestern Gulf of Mexico. Fisheries Research, 93(1–2):77–84, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608000672.

Morton:2008:CBC

[MR08b] Alexandra Morton and Richard Routledge. Comment on Butterworth, Cubitt, and McKinley (2008). Fisheries Research, 92(2–3):345–346, August 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001148. See [BCM08b] and re-

sponse [BCM08a].

Mancera-Rodriguez:2004:AGS

[MRCH04] Néstor Javier Mancera-Rodríguez and José Juan Castro-Hernández. Age and growth of Stephanolepis hispidus (Linnaeus, 1766) (Pisces: Monacanthidae), in the Canary Islands area. Fisheries Research, 66(2–3):381–386, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603002248.

Martinez-Rincon:2009:ICD

[MROGVR09] R. O. Martínez-Rincón, S. Ortega-García, and J. G. Vaca-Rodriguez. Incidental catch of dolphinfish (Coryphaena spp.) reported by the Mexican tuna purse seiners in the eastern Pacific Ocean. Fisheries Research, 96(2–3):296–302, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608003974.

Morris:2003:LHS

[MRT03] James A. Morris, Roger A. Rulifson, and Larry H. Toburen. Life history strategies of striped bass, Morone saxatilis, populations inferred from otolith microchemistry. Fisheries Research, 62(1):53–63, April 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002461.

Munro:2002:ENE

[MS02] Peter T. Munro and David A. Somerton. Estimating net efficiency of a survey trawl for flatfishes. Fisheries Research, 55(1–3):267–279, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002806.

Maunder:2003:FFM

[MS03] M. N. Maunder and P. J. Starr. Fitting fisheries models to standardised CPUE abundance indices. Fisheries Research, 63(1):43–50, July 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360300002X.

Madsen:2005:SEE

[MS05a] Niels Madsen and Karl-Johan Stæhr. Selectivity experiments to estimate the effect of escape windows in the Skagerak roundfish fishery. Fisheries Research, 71(2):241–245, February 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604001845.

Marsden:2005:TFF

[MS05b] A. Dale Marsden and Ussif Rashid Sumaila. Tracking flows of fisheries products: the case of Pacific halibut in the Canadian economy. Fisheries Research, 73(1–2):259–264, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360400311X.

Morato:2000:FHS

[MSA00] Telmo Morato, Ricardo Serrão Santos, and José Pedro Andrade. Feeding habits, seasonal and ontogenetic diet shift of blacktail comber, Serranus atricauda (Pisces: Serranidae), from the Azores, north-eastern Atlantic. Fisheries Research, 49(1):51–59, November 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001892.

Montgomery:2007:UPC

[MSB07] S. S. Montgomery, J. Stewart, and C. Blount. The use of a point-count technique to quantify the egg assemblage in Melicertus plebejus (Hess, 1865) ovaries. Fisheries Research, 83(2–3):228–237, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003432.

Madsen:2008:STC

[MSB^{+}08] Niels Madsen, Roar Skeide, Mike Breen, Ludvig A. Krag, Irene Huse, and Aud V. Soldal. Selectivity in a trawl codend during haul-back operation — an overlooked phenomenon. Fisheries Research, 91(2–3):168–174, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003384.

Milanese:2008:CSF

[MSM^{+}08] M. Milanese, A. Sarà, R. Manconi, A. Ben Abdalla, and R. Pronzato. Commercial sponge fishing in Libya: Historical records, present status and perspectives. Fisheries Research, 89(1):90–96, January 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002287.

Markaida:2001:RBJ

[MSN01] Unai Markaida and Oscar Sosa-Nishizaki. Reproductive biology of jumbo squid Dosidicus gigas in the Gulf of California, 1995–1997. Fisheries Research, 54(1):63–82, December 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601003733.

McKinstry:2005:SAF

[MSSJ05] Craig A. McKinstry, Mary Ann Simmons, C. Steve Simmons, and Robert L. Johnson. Statistical assessment of fish behavior from split-beam hydro-acoustic sampling. Fisheries Research, 72(1):29–44, April 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002644.

Melville-Smith:2004:IFR

[MSTC04] Roy Melville-Smith, Adrian W. Thomson, and Nick Caputi. Improved forecasts of recreational western rock lobster (Panulirus cygnus) catches in Western Australia, by predicting licence usage. Fisheries Research, 68(1–3):203– 208, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604000207.

Marchesan:2005:BEA

[MSVF05] Mara Marchesan, Maurizio Spoto, Laura Verginella, and Enrico A. Ferrero. Behavioural effects of artificial light on fish species of commercial interest. Fisheries Research, 73(1–2):171–185, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604003133.

Maravelias:2008:EEA

[MT08] Christos D. Maravelias and Efthymia V. Tsitsika. Economic efficiency analysis and fleet capacity assessment in Mediterranean fisheries. Fisheries Research, 93(1–2):85– 91, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608000842.

Marrs:2002:PDL

[MTA^{+}02] S. J. Marrs, I. D. Tuck, R. J. A. Atkinson, T. D. I. Stevenson, and C. Hall. Position data loggers and logbooks as tools in fisheries research: results of a pilot study and some recommendations. Fisheries Research, 58(1):109–117, October 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601003629.

Marques:2006:DCI

[MTC06] Joana F. Marques, Célia M. Teixeira, and Henrique N. Cabral. Differentiation of commercially important flatfish populations along the Portuguese coast: Evidence from morphology and parasitology. Fisheries Research, 81(2–3): 293–305, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606002402.

McClatchie:2000:GTT

[MTGH00] Sam McClatchie, Richard E. Thorne, Paul Grimes, and Stuart Hanchet. Ground truth and target identification for fisheries acoustics. Fisheries Research, 47(2–3):173–191, July 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001685.

Madsen:2006:DTS

[MTHL06] N. Madsen, V. Tschernij, K. Hansen, and P.-O. Larsson. Development and testing of a species-selective flatfish ottertrawl to reduce cod bycatches. Fisheries Research, 78(2–3):298–308, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000178.

Mackinson:2005:UAI

[MTRM05] S. Mackinson, K. Turner, D. Righton, and J. D. Metcalfe. Using acoustics to investigate changes in efficiency of a sandeel dredge. Fisheries Research, 71(3):357–363, March 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604002292.

Metin:2004:SRM

[MTU^{+}04] Cengiz Metin, Adnan Tokaç, Ali Ulaş, F. Ozan Düzbastılar, Altan Lök, Hüseyin Özbilgin, Gülnur Metin, Zafer Tosunoğlu, Hakan Kaykaç, and Celalettin Aydın. Survival of red mullet (Mullus barbatus L., 1758) after escape from a trawl codend in the Aegean Sea. Fisheries Research, 70(1):49–53, November 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360400150X.

Moses:2002:SSA

[MUO02] Boniface S. Moses, Okon M. Udoidiong, and Aniefioknkpong O. Okon. A statistical survey of the artisanal fisheries of south-eastern Nigeria and the influence of hydroclimatic factors on catch and resource productivity. Fisheries Research, 57(3):267–278, September 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003538.

Madrid-Vera:2007:AAE

[MVAMB07] Juan Madrid-Vera, Felipe Amezcua, and Enrique Morales-Bojórquez. An assessment approach to estimate biomass of fish communities from bycatch data in a tropical shrimptrawl fishery. Fisheries Research, 83(1):81–89, January 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606003109.

Mackinson:2006:PFD

[MvdK06] Steven Mackinson and Jeroen van der Kooij. Perceptions of fish distribution, abundance and behaviour: Observations revealed by alternative survey strategies made by scientific and fishing vessels. Fisheries Research, 81(2–3):306–315, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606002384.

Mous:2002:ESM

[MvDM02] P. J. Mous, W. L. T. van Densen, and M. A. M. Machiels. The effect of smaller mesh sizes on catching larger fish with trawls. Fisheries Research, 54(2):171–179, January 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600003040.

Machias:2001:BTD

[MVV^{+}01] A. Machias, V. Vassilopoulou, D. Vatsos, P. Bekas, A. Kallianiotis, C. Papaconstantinou, and N. Tsimenides. Bottom trawl discards in the northeastern Mediterranean Sea. Fisheries Research, 53(2):181–195, October 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002988.

Mapleston:2008:ECH

[MWB^{+}08] Amos Mapleston, David Welch, Gavin A. Begg, Mark McLennan, David Mayer, and Ian Brown. Effect of changes in hook pattern and size on catch rate, hooking location, injury and bleeding for a number of tropical reef fish species. Fisheries Research, 91(2–3):203–211, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003438.

McVean:2006:TSF

[MWF06] Angus R. McVean, Ryan C. J. Walker, and Eibleis Fanning. The traditional shark fisheries of southwest Madagascar: a study in the Toliara region. Fisheries Research, 82(1–3):280–289, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002153.

Maliao:2004:SSD

[MWJ04] Ronald J. Maliao, Edward L. Webb, and Kathe R. Jensen. A survey of stock of the donkey’s ear abalone, Haliotis asinina L. in the Sagay Marine Reserve, Philippines: evaluating the effectiveness of marine protected area enforcement. Fisheries Research, 66(2–3):343–353, February 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001814.

Niamaimandi:2007:PDG

[NAD^{+}07] Nassir Niamaimandi, Aziz Bin Arshad, Siti Khalijah Daud, Ross Cheroos Saed, and Bahram Kiabi. Population dynamic of green tiger prawn, Penaeus semisulcatus (De Haan) in Bushehr coastal waters, Persian Gulf. Fisheries Research, 86(2–3):105–112, September 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001142.

Narvarte:2006:BFW

[Nar06] Maite Andrea Narvarte. Biology and fishery of the whelk Buccinanops globulosum (Kiener, 1834) in northern coastal waters of the San Matías Gulf (Patagonia, Argentina). Fisheries Research, 77(2):131–137, February 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003061.

Nash:2003:IASb

[Nas03] Colin E. Nash. Interactions of Atlantic salmon in the Pacific Northwest: VI. a synopsis of the risk and uncertainty. Fisheries Research, 62(3):339–347, June 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000687.

Nunoo:2009:WTF

[NBA^{+}09] F. K. E. Nunoo, Josephine O. Boateng, Angela M. Ahulu, Kwame A. Agyekum, and Ussif Rashid Sumaila. When trash fish is treasure: the case of Ghana in West Africa. Fisheries Research, 96(2–3):167–172, March 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360800338X.

Nasreddine:2009:SGC

[NBF09] Kamal Nasreddine, Abdesslam Benzinou, and Ronan Fablet. Shape geodesics for the classification of calcified structures: Beyond Fourier shape descriptors. Fisheries Research, 98(1–3):8–15, June 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000721.

Nishida:2004:ISA

[NC04] Tom Nishida and Ding-Geng Chen. Incorporating spatial autocorrelation into the general linear model with an application to the yellowfin tuna (Thunnus albacares) longline CPUE data. Fisheries Research, 70(2–3):265–274, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001717.

Newman:2000:AGMb

[NCW00a] Stephen J. Newman, Michael Cappo, and David McB. Williams. Age, growth and mortality of the stripey, Lutjanus carponotatus (Richardson) and the brown-stripe snapper, L. vitta (Quoy and Gaimard) from the central Great Barrier Reef, Australia. Fisheries Research, 48(3): 263–275, October 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783600001843.

Newman:2000:AGMa

[NCW00b] Stephen J. Newman, Michael Cappo, and David McB. Williams. Age, growth, mortality rates and corresponding yield estimates using otoliths of the tropical red snappers, Lutjanus erythropterus, L. malabaricus and L. sebae, from the central Great Barrier Reef. Fisheries Research, 48(1):1–14, August 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001156.

Nash:2009:SRN

[NDCK09] Richard D. M. Nash, Mark Dickey-Collas, and Laurence T. Kell. Stock and recruitment in North Sea herring (Clupea harengus); compensation and depensation in the population dynamics. Fisheries Research, 95(1):88–97, January 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608002282.

Nelson:2007:FEW

[NE07] Jonathan D. Nelson and Scott A. Eckert. Foraging ecology of whale sharks (Rhincodon typus) within Bahía de Los Angeles, Baja California Norte, México. Fisheries Research, 84(1):47–64, March 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004000.

Needle:2008:MSE

[Nee08] Coby L. Needle. Management strategy evaluation for North Sea haddock. Fisheries Research, 94(2):141–150, November 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608000878.

Newman:2002:GRA

[New02] Stephen J. Newman. Growth rate, age determination, natural mortality and production potential of the scarlet seaperch, Lutjanus malabaricus Schneider 1801, off the Pilbara coast of north-western Australia. Fisheries Research, 58(2):215–225, November 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003678.

Narvarte:2006:CTO

[NGF06] Maite Narvarte, Raúl González, and Miriam Fernández. Comparison of Tehuelche octopus (Octopus tehuelchus) abundance between an open-access fishing ground and a marine protected area: Evidence from a direct development species. Fisheries Research, 79(1–2):112–119, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606000683.

Narvarte:2007:AMF

[NGF07] Maite Narvarte, Raúl González, and Pablo Filippo. Artisanal mollusk fisheries in San Matías Gulf (Patagonia, Argentina): an appraisal of the factors contributing to unsustainability. Fisheries Research, 87(1):68–76, October 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001518.

Nguyen:2009:PRB

[NGM^{+}09] Vivian Nguyen, Marie-Ange Gravel, Amos Mapleston, Kyle C. Hanson, and Steven J. Cooke. The postrelease behaviour and fate of tournament-caught smallmouth bass after ‘fizzing’ to alleviate distended swim bladders. Fisheries Research, 96(2–3):313–318, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003858.

Nieland:2008:MSL

[NHSD08] Julie L. Nieland, Michael J. Hansen, Michael J. Seider, and Jonathan J. Deroba. Modeling the sustainability of lake trout fisheries in eastern Wisconsin waters of Lake Superior. Fisheries Research, 94(3):304–314, December 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002178.

Nichol:2007:PPC

[NHT07] Daniel G. Nichol, Taina Honkalehto, and Grant G. Thompson. Proximity of Pacific cod to the sea floor: Using archival tags to estimate fish availability to research bottom trawls. Fisheries Research, 86(2–3):129–135, September 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001178.

Niemela:2000:QEJ

[NJE00] Eero Niemelä, Markku Julkunen, and Jaakko Erkinaro. Quantitative electrofishing for juvenile salmon densities: assessment of the catchability during a long-term monitoring programme. Fisheries Research, 48(1):15–22, August 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001132.

Nissling:2006:EST

[NJJ06] Anders Nissling, Ulrika Johansson, and Marie Jacobsson. Effects of salinity and temperature conditions on the reproductive success of turbot (Scophthalmus maximus) in the Baltic Sea. Fisheries Research, 80(2–3):230–238, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606001469.

Neiland:2000:IFNa

[NJL^{+}00] A. E. Neiland, S. Jaffry, B. M. B. Ladu, M. T. Sarch, and S. P. Madakan. Inland fisheries of North East Nigeria including the Upper River Benue, Lake Chad and the Nguru– Gashua wetlands: I. Characterisation and analysis of planning suppositions. Fisheries Research, 48(3):229–243, October 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600001806.

Nielsen:2001:DDA

[NLJS01] J. Rasmus Nielsen, Bo Lundgren, Torben F. Jensen, and Karl-Johan Stæhr. Distribution, density and abundance of the western Baltic herring (Clupea harengus) in the Sound (ICES Subdivision 23) in relation to hydrographical features. Fisheries Research, 50(3):235–258, March 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002204.

Nielsen:2006:SPD

[NM06] Jesper Raakjær Nielsen and Christoph Mathiesen. Stakeholder preferences for Danish fisheries management of sand eel and Norway pout. Fisheries Research, 77(1):92–101, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605002675.

Nakaya:2008:VOD

[NMF^{+}08] Mitsuhiro Nakaya, Taizo Morioka, Kyohei Fukunaga, Sachio Sekiya, Tadao Jinbo, Hiroshi Nagashima, and Yasuhiro Ueno. Validation of otolith daily increments for larval and juvenile Japanese halfbeak Hyporhamphus sajori. Fisheries Research, 93(1–2):186–189, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001173.

Nevarez-Martinez:2000:BDJ

[NMHHMB^{+}00] Manuel O. Nevárez-Martínez, Agustín Hernández-Herrera, Enrique Morales-Bojórquez, Alejandro Balmori-Ramírez, Miguel A. Cisneros-Mata, and Rufino Morales-Azpeitia. Biomass and distribution of the jumbo squid (Dosidicus gigas; d’Orbigny, 1835) in the Gulf of California, Mexico. Fisheries Research, 49(2):129–140, December 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001983.

Nevarez-Martinez:2006:GMR

[NMMTCV^{+}06] Manuel O. Nevárez-Martínez, Francisco J. Méndez-Tenorio, Celio Cervantes-Valle, Juana López-Martínez, and Myrna L. Anguiano-Carrasco. Growth, mortality, recruitment, and yield of the jumbo squid (Dosidicus gigas) off Guaymas, Mexico. Fisheries Research, 79(1–2):38–47, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606000658.

Nigmatullin:2001:RBJ

[NNA01] Ch. M. Nigmatullin, K. N. Nesis, and A. I. Arkhipkin. A review of the biology of the jumbo squid Dosidicus gigas (Cephalopoda: Ommastrephidae). Fisheries Research, 54(1):9–19, December 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360100371X. See erratum [vSM02a].

Nigmatullin:2002:ERB

[NNA02] Ch. M. Nigmatullin, K. N. Nesis, and A. I. Arkhipkin. Erratum to “A review of the biology of the jumbo squid Dosidicus gigas (Cephalopoda: Ommastrephidae)”: [Fish. Res. 54 (2001) 9–19]. Fisheries Research, 56(2):221, May 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000292. See [NNA01].

Naslund:2005:LTR

[NNE^{+}05] Ingemar Näslund, Fredrik Nordwall, Torleif Eriksson, David Hannersjö, and Lars-Ove Eriksson. Long-term responses of a stream-dwelling grayling population to restrictive fishing regulations. Fisheries Research, 72(2–3):323– 332, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360400298X.

Newlands:2008:MSS

[NP08] Nathaniel K. Newlands and Tracy A. Porcelli. Measurement of the size, shape and structure of Atlantic bluefin tuna schools in the open ocean. Fisheries Research, 91 (1):42–55, May 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360700313X.

Newton:2008:RSH

[NPC^{+}08] A. W. Newton, K. J. Peach, K. A. Coull, M. Gault, and C. L. Needle. Rockall and the Scottish haddock fishery. Fisheries Research, 94(2):133–140, November 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000994.

Narvarte:2004:EPC

[NS04] Maite Andrea Narvarte and Marıá Natalia Saiz. Effects of the pinnotherid crab Tumidotheres maculatus on the Tehuelche scallop Aequipecten tehuelchus in the San Matías Gulf, Argentina. Fisheries Research, 67(2):207–214, April 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603002935.

Nilssen:2006:ISR

[NS06] Einar M. Nilssen and Jan H. Sundet. The introduced species red king crab (Paralithodes camtschaticus) in the Barents Sea: II. Growth increments and moulting probability. Fisheries Research, 82(1–3):319–326, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001706.

Norman:2007:SMS

[NS07] Bradley M. Norman and John D. Stevens. Size and maturity status of the whale shark (Rhincodon typus) at Ningaloo Reef in Western Australia. Fisheries Research, 84(1):81–86, March 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004036.

Najmudeen:2008:EIJ

[NS08] T. M. Najmudeen and R. Sathiadhas. Economic impact of juvenile fishing in a tropical multi-gear multi-species fishery. Fisheries Research, 92(2–3):322–332, August 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000544.

Nanami:2005:AGM

[NT05] Atsushi Nanami and Takeshi Takegaki. Age and growth of the mudskipper Boleophthalmus pectinirostris in Ariake Bay, Kyushu, Japan. Fisheries Research, 74(1–3):24–34, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605001268.

Norbis:2005:PTW

[NV05] Walter Norbis and José Verocai. Presence of two whitemouth croaker (Micropogonias furnieri, Pisces: Sciaenidae) groups in the Río de la Plata spawning coastal area as consequence of reproductive migration. Fisheries Research, 74(1–3):134–141, August 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000883.

Nash:2003:IASa

[NW03] Colin E. Nash and F. William Waknitz. Interactions of Atlantic salmon in the Pacific Northwest: I. Salmon enhancement and the net-pen farming industry. Fisheries Research, 62(3):237–254, June 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000638.

Neiland:2000:IFNb

[NWML00] A. E. Neiland, J. Weeks, S. P. Madakan, and B. M. B. Ladu. Inland fisheries of North East Nigeria including the Upper River Benue, Lake Chad and the Nguru– Gashua wetlands: II. Fisheries management at village level. Fisheries Research, 48(3):245–261, October 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001818.

Ortiz:2004:AED

[OA04] Mauricio Ortiz and Freddy Arocha. Alternative error distribution models for standardization of catch rates of non-target species from a pelagic longline fishery: billfish species in the Venezuelan tuna longline fishery. Fisheries Research, 70(2–3):275–297, December 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001729.

Oresland:2008:LGD

[ØA08] Vidar Øresland and Carl André. Larval group differentiation in Atlantic cod (Gadus morhua) inside and outside the Gullmar Fjord. Fisheries Research, 90(1–3):9–16, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607002275.

Ojaveer:2007:SBS

[OAKM07] Henn Ojaveer, Kenneth Awebro, Hrefna M. Karlsdóttir, and Brian R. MacKenzie. Swedish Baltic Sea fisheries during 1868–1913: Spatio-temporal dynamics of catch and fishing effort. Fisheries Research, 87(2–3):137–145, November 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607001725.

Oliva:2002:MPC

[OB02] Marcelo E. Oliva and Isabel Ballón. Metazoan parasites of the Chilean hake Merluccius gayi gayi as a tool for stock discrimination. Fisheries Research, 56(3):313–320, June 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601003290.

ONeill:2003:CRF

[OCT^{+}03] Michael F. O’Neill, Anthony J. Courtney, Clive T. Turnbull, Norm M. Good, Kate M. Yeomans, Jonathan Staunton Smith, and Celeste Shootingstar. Comparison of relative fishing power between different sectors of the Queensland trawl fishery, Australia. Fisheries Research, 65(1–3):309–321, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002510.

Ofori-Danson:2002:PPE

[ODdGV02] P. K. Ofori-Danson, G. J. de Graaf, and C. J. Vanderpuye. Population parameter estimates for Chrysichthys auratus and C. nigrodigitatus (Pisces: Claroteidae) in Lake Volta, Ghana. Fisheries Research, 54(2):267–277, January 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002927.

Olla:2000:DOS

[ODR00] B. L. Olla, M. W. Davis, and C. Rose. Differences in orientation and swimming of walleye pollock Theragra chalcogramma in a trawl net under light and dark conditions: concordance between field and laboratory observations. Fisheries Research, 44(3):261–266, January 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000934.

ONeill:2003:UBT

[OF03] Michael F. O’Neill and Malcolm J. Faddy. Use of binary and truncated negative binomial modelling in the analysis of recreational catch data. Fisheries Research, 60(2–3):471–477, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001017.

Oliva:2008:MPS

[OFOM08] M. E. Oliva, I. Fernández, C. Oyarzún, and C. Murillo. Metazoan parasites of the stomach of Dissostichus eleginoides Smitt 1898 (Pisces: Notothenidae) from southern Chile: a tool for stock discrimination? Fisheries Research, 91(2–3):119–122, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003190.

ONeill:2008:EVC

[OGK^{+}08] F. G. O’Neill, N. Graham, R. J. Kynoch, R. S. T. Ferro, P. A. Kunzlik, and R. J. Fryer. The effect of varying cod-end circumference, inserting a ‘flexi-grid’ or inserting a bacoma type panel on the selectivity of North Sea haddock and saithe. Fisheries Research, 94(2):175–183, November 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001859.

Ogle:2009:EFL

[Ogl09] D. H. Ogle. The effect of freezing on the length and weight measurements of ruffe (Gymnocephalus cernuus). Fisheries Research, 99(3):244–247, August 15, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001763.

Oliveira:2009:PCA

[OGPC09] M. M. Oliveira, M. B. Gaspar, J. P. Paixão, and A. S. Camanho. Productivity change of the artisanal fishing fleet in Portugal: a Malmquist index analysis. Fisheries Research, 95(2–3):189–197, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002610.

Ortega-Garcia:2008:RIL

[OGPDOM08] S. Ortega-Garcia, G. Ponce-Diaz, R. O’Hara, and J. Merilä. The relative importance of lunar phase and environmental conditions on striped marlin (Tetrapturus audax) catches in sport fishing. Fisheries Research, 93(1–2):190–194, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608001185.

Otero:2007:RCE

[OGSG07] Jaime Otero, Ángel F. González, M. Pilar Sieiro, and Ángel Guerra. Reproductive cycle and energy allocation of Octopus vulgaris in Galician waters, NE Atlantic. Fisheries Research, 85(1–2):122–129, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360700015X.

Obata:2006:CSM

[OIK^{+}06] Yasuhiro Obata, Hideyuki Imai, Toshihide Kitakado, Katsuyuki Hamasaki, and Shuichi Kitada. The contribution of stocked mud crabs Scylla paramamosain to commercial catches in Japan, estimated using a genetic stock identification technique. Fisheries Research, 80(1):113–121, August 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606001263.

OBoyle:2006:OIE

[OJ06] Robert O’Boyle and Glen Jamieson. Observations on the implementation of ecosystem-based management: Experiences on Canada’s east and west coasts. Fisheries Research, 79(1–2):1–12, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600138X.

Ojaveer:2006:EPE

[Oja06] E. Ojaveer. On the external and parental effects in early development of herring (Clupea pallasi) at the NE Kamchatka. Fisheries Research, 81(1):1–8, October 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002311.

ONeill:2006:SMP

[OKF06] F. G. O’Neill, R. J. Kynoch, and R. J. Fryer. Square mesh panels in North Sea demersal trawls: Separate estimates of panel and cod-end selectivity. Fisheries Research, 78(2–3):333–341, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000142.

Olin:2004:EFA

[OKPR04] M. Olin, M. Kurkilahti, P. Peitola, and J. Ruuhijärvi. The effects of fish accumulation on the catchability of multimesh gillnet. Fisheries Research, 68(1–3):135–147, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604000657.

ONeill:2005:CED

[OKWM05] F. G. O’Neill, L. H. Knudsen, D. A. Wileman, and S. J. McKay. Cod-end drag as a function of catch size and towing speed. Fisheries Research, 72(2–3):163–171, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604003030.

ONeill:2007:FPI

[OL07] Michael F. O’Neill and George M. Leigh. Fishing power increases continue in Queensland’s east coast trawl fishery, Australia. Fisheries Research, 85(1–2):84–92, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000057.

Ovenden:2002:SGS

[OLN^{+}02] Jennifer R. Ovenden, Julie Lloyd, Stephen J. Newman, Clive P. Keenan, and Leigh S. Slater. Spatial genetic subdivision between northern Australian and southeast Asian populations of Pristipomoides multidens: a tropical marine reef fish species. Fisheries Research, 59(1–2):57–69, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601004155.

Ojaveer:2007:HDF

[OM07] Henn Ojaveer and Brian R. MacKenzie. Historical development of fisheries in northern Europe — reconstructing chronology of interactions between nature and man.

Fisheries Research, 87(2–3):102–105, November 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001890.

OSullivan:2003:AGR

[OMF^{+}03] S. O’Sullivan, C. Moriarty, R. D. FitzGerald, J. Davenport, and M. F. Mulcahy. Age, growth and reproductive status of the European conger eel, Conger conger (L.) in Irish coastal waters. Fisheries Research, 64(1):55–69, October 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603001000.

Olin:2009:GCE

[OMR09] M. Olin, T. Malinen, and J. Ruuhijärvi. Gillnet catch in estimating the density and structure of fish community — comparison of gillnet and trawl samples in a eutrophic lake. Fisheries Research, 96(1):88–94, February 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002853.

ONeill:2003:IRB

[OMW^{+}03] F. G. O’Neill, S. J. McKay, J. N. Ward, A. Strickland, R. J. Kynoch, and A. F. Zuur. An investigation of the relationship between sea state induced vessel motion and cod-end selection. Fisheries Research, 60(1):107–130, January 30, 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602000565.

Olivar:2003:STD

[OQE03] M. P. Olivar, G. Quílez, and M. Emelianov. Spatial and temporal distribution and abundance of European hake, Merluccius merluccius, eggs and larvae in the Catalan coast (NW Mediterranean). Fisheries Research, 60(2–3):321–331, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360200139X.

Overton:2007:EPS

[OR07] Anthony S. Overton and Roger A. Rulifson. Evaluation of plankton surface pushnets and oblique tows for comparing the catch of diadromous larval fish. Fisheries Research, 86(2–3):99–104, September 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001105.

Orensanz:2003:BRC

[Ore03] J. M. Orensanz. Book review: Cool Crabs: Crabs in Cold Water Regions: Biology, Management, and Economics, Proceedings of the Symposium Crab2001 (19th Lowell Wakefield Symposium), January 17–20, 2001, Anchorage, AK: A. J. Paul, E. G. Dawe, R. Elner, G. S. Jamieson, G. H. Kruse, R. S. Otto, B. Sainte-Marie, T. C. Shirley, D. Woodby (Eds.); University of Alaska Sea Grant AK-SG-02-01, Fairbanks, 2002, 876 pages, hardcover, ISBN 1-56612-077-2, US $40. Fisheries Research, 64(1):89–92, October 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603002182.

Orlowski:2003:ITC

[Orl03]  Andrzej Orlowski.  Influence of thermal conditions

on biomass of fish in the Polish EEZ. Fisheries Research, 63(3):367–377, September 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000973.

Olson:2006:BJS

[ORVMP06] Robert J. Olson, Marlon H. Román-Verdesoto, and Glenthon L. Macías-Pita. Bycatch of jumbo squid Dosidicus gigas in the tuna purse-seine fishery of the eastern Pacific Ocean and predatory behaviour during capture. Fisheries Research, 79(1–2):48–55, June 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600066X.

Oliva:2005:MPC

[OS05] Marcelo E. Oliva and Monique F. Sánchez. Metazoan parasites and commensals of the northern Chilean scallop Argopecten purpuratus (Lamarck, 1819) as tools for stock identification. Fisheries Research, 71(1):71–77, January 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001948.

ONeill:2009:ULS

[OSB09] F. G. O’Neill, K. Summerbell, and M. Breen. An underwater laser stripe seabed profiler to measure the physical impact of towed gear components on the seabed. Fisheries Research, 99(3):234–238, August 15, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001441.

Ostrand:2005:LBC

[OSC^{+}05] Kenneth G. Ostrand, Michael J. Siepker, Steven J. Cooke, William F. Bauer, and David H. Wahl. Largemouth bass catch rates and injury associated with non-offset and offset circle hook configurations. Fisheries Research, 74(1–3):306–311, August 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000858.

Ochwada:2008:PAF

[OSG08] Faith A. Ochwada, James P. Scandol, and Charles A. Gray. Predicting the age of fish using general and generalized linear models of biometric data: a case study of two estuarine finfish from New South Wales, Australia. Fisheries Research, 90(1–3):187–197, April 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002731.

Ovredal:2002:SFR

[ØT02] Jan Tore Øvredal and Bjørn Totland. The scantrol Fish-Meter for recording fish length, weight and biological data.

Fisheries Research, 55(1–3):325–328, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002740.

Ozbilgin:2003:CSD

[ÖT03] Hüseyin Özbilgin and Zafer Tosunoğlu. Comparison of the selectivities of double and single codends. Fisheries Research, 63(1):143–147, July 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000055.

Overholtz:2002:GMG

[Ove02] W. J. Overholtz. The Gulf of Maine–Georges Bank Atlantic herring (Clupea harengus): spatial pattern analysis of the collapse and recovery of a large marine fish complex. Fisheries Research, 57(3):237–254, September 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003599.

Ozbilgin:2002:EST

[ÖW02] H. Özbilgin and C. S. Wardle. Effect of seasonal temperature changes on the escape behaviour of haddock, Melanogrammus aeglefinus, from the codend. Fisheries Research, 58(3):323–331, November 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003940.

Palma:2002:MSD

[PA02] J. Palma and J. P. Andrade. Morphological study of Diplodus sargus, Diplodus puntazzo, and Lithognathus mormyrus (Sparidae) in the Eastern Atlantic and Mediterranean Sea. Fisheries Research, 57(1):1–8, July 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003356.

Pelicice:2005:POF

[PA05] Fernando Mayer Pelicice and Angelo Antonio Agostinho. Perspectives on ornamental fisheries in the upper Paraná River floodplain, Brazil. Fisheries Research, 72(1):109–119, April 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604002322.

Powers:2009:FER

[PA09] Joseph E. Powers and Shane M. Abeare. Fishing effort redistribution in response to area closures. Fisheries Research, 99(3):216–225, August 15, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001751.

Payne:2005:PAF

[PAB05] Adam G. Payne, David J. Agnew, and Anabela Brandão. Preliminary assessment of the Falklands Patagonian toothfish (Dissostichus eleginoides) population: Use of recruitment indices and the estimation of unreported catches. Fisheries Research, 76(3):344–358, December 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002158.

Prat:2008:SMI

[PAF^{+}08] J. Prat, J. Antonijuan, A. Folch, A. Sala, A. Lucchetti, F. Sardà, and A. Manuel. A simplified model of the interaction of the trawl warps, the otterboards and netting drag. Fisheries Research, 94(1):109–117, October 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360800218X.

Palsson:2003:LBA

[Pál03] Ólafur K. Pálsson. A length-based analysis of haddock discards in Icelandic fisheries. Fisheries Research, 59(3):437–446, January 3, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000176.

Palsson:2005:ACI

[Pál05] Ólafur K. Pálsson. An analysis of by-catch in the Icelandic blue whiting fishery. Fisheries Research, 73(1–2):135–146, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605000020.

Palmer:2008:CDT

[Pal08] Michael C. Palmer. Calculation of distance traveled by fishing vessels using GPS positional data: a theoretical evaluation of the sources of error. Fisheries Research, 89(1):57–64, January 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607002159.

Purcell:2004:RCS

[PAP04] S. W. Purcell, M. J. Amos, and K. Pakoa. Releases of cultured sub-adult Trochus niloticus generate broodstock for fishery replenishment in Vanuatu. Fisheries Research, 67(3):329–333, May 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603003059.

Payne:2006:TAC

[PAP06] Adam G. Payne, David J. Agnew, and Graham J. Pierce. Trends and assessment of cephalopod fisheries. Fisheries Research, 78(1):1–3, April 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003358.

Pedersen:2003:LDM

[PAR^{+}03] O. P. Pedersen, M. Aschan, T. Rasmussen, K. S. Tande, and D. Slagstad. Larval dispersal and mother populations of Pandalus borealis investigated by a Lagrangian particletracking model. Fisheries Research, 65(1–3):173–190, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360300242X.

Parente:2004:PCS

[Par04] Joaquim Parente. Predictors of CPUE and standardization of fishing effort for the Portuguese coastal seine fleet. Fisheries Research, 69(3):381–387, October 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001250.

Pawson:2003:CCL

[Paw03] M. G. Pawson. The catching capacity of lost static fishing gears: introduction. Fisheries Research, 64(2–3):101–105, November 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360300208X.

Pierce:2003:EMI

[PB03] G. J. Pierce and P. R. Boyle. Empirical modelling of interannual trends in abundance of squid (Loligo forbesi) in Scottish waters. Fisheries Research, 59(3):305–326, January 3, 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602000280.

Peery:2004:IBN

[PB04] Christopher A. Peery and Theodore C. Bjornn. Interactions between natural and hatchery Chinook salmon parr in a laboratory stream channel. Fisheries Research, 66(2– 3):311–324, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783603001942.

Pipitone:2000:FBI

[PBDP00] Carlo Pipitone, Fabio Badalamenti, Giovanni D’Anna, and Bernardo Patti. Fish biomass increase after a fouryear trawl ban in the Gulf of Castellammare (NW Sicily, Mediterranean Sea). Fisheries Research, 48(1):23–30, August 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001144.

Pawson:2008:WPS

[PBLP08] M. G. Pawson, M. Brown, J. Leballeur, and G. D. Pickett. Will philopatry in sea bass, Dicentrarchus labrax, facilitate the use of catch-restricted areas for management of recreational fisheries? Fisheries Research, 93(1–2):240–243, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360800088X.

Petursdottir:2006:DBI

[PBM06] Groa Petursdottir, Gavin A. Begg, and Gudrun Marteinsdottir. Discrimination between Icelandic cod (Gadus morhua L.) populations from adjacent spawning areas based on otolith growth and shape. Fisheries Research, 80(2–3):182–189, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001731.

Powell:2003:VTA

[PBMB03] Eric N. Powell, Allison J. Bonner, Bruce Muller, and Eleanor A. Bochenek. Vessel time allocation in the US Illex illecebrosus fishery. Fisheries Research, 61(1–3):35– 55, March 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360200245X.

Power:2005:PIP

[PBR05] A. M. Power, J. A. Balbuena, and J. A. Raga. Parasite infracommunities as predictors of harvest location of bogue (Boops boops L.): a pilot study using statistical classifiers. Fisheries Research, 72(2–3):229–239, May 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002486.

Paramo:2007:TDA

[PBVG07] Jorge Paramo, Sophie Bertrand, Héctor Villalobos, and François Gerlotto. A three-dimensional approach to school typology using vertical scanning multibeam sonar. Fisheries Research, 84(2):171–179, April 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003778.

Pitcher:2009:LSB

[PBW^{+}09] C. R. Pitcher, C. Y. Burridge, T. J. Wassenberg, B. J. Hill, and I. R. Poiner. A large scale BACI experiment to test the effects of prawn trawling on seabed biota in a closed area of the Great Barrier Reef Marine Park, Australia. Fisheries Research, 99(3):168–183, August 15, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001453.

Parrack:2003:ETL

[PC03] Michael L. Parrack and Nancie J. Cummings. Errors in transforming length samples to age frequencies without age samples. Fisheries Research, 63(2):235–243, August 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000407.

Pino:2004:OWE

[PCAS04] Claudio A. Pino, Luis A. Cubillos, Miguel Araya, and Aquiles Sepúlveda. Otolith weight as an estimator of age in the Patagonian grenadier, Macruronus magellanicus, in central-south Chile. Fisheries Research, 66(2–3):145–156, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360300225X.

Pineiro:2001:DWF

[PCB01] C. G. Piñeiro, M. Casas, and R. Bañón. The deepwater fisheries exploited by Spanish fleets in the Northeast Atlantic: a review of the current status. Fisheries Research, 51(2–3):311–320, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002545.

Perez-Castaneda:2003:RMM

[PCD03] Roberto Pérez-Castañeda and Omar Defeo. A reciprocal model for mortality at length in juvenile pink shrimps (Farfantepenaeus duorarum) in a coastal lagoon of Mexico. Fisheries Research, 63(2):283–287, August 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000584.

Poulsen:2007:AEL

[PCHM07] René Taudal Poulsen, Andrew B. Cooper, Poul Holm, and Brian R. MacKenzie. An abundance estimate of ling (Molva molva) and cod (Gadus morhua) in the Skagerrak and the northeastern North Sea, 1872. Fisheries Research, 87(2–3):196–207, November 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001609.

Poulsen:2008:BMR

[PCHM08] René Taudal Poulsen, Andrew B. Cooper, Poul Holm, and Brian R. MacKenzie. Beware the misapplication of results: Response to Cardinale and Svedäng (2008). Fisheries Research, 89(3):307–308, March 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003001. See [CS08a].

Pajaro:2007:ECN

[PCM07] Marcelo Pájaro, Jéssica Curelovich, and Gustavo J. Macchi. Egg cannibalism in the northern population of the Argentine anchovy, Engraulis anchoita (Clupeidae). Fisheries Research, 83(2–3):253–262, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003481.

Pol:2008:LHM

[PCMC08] Michael V. Pol, Steven J. Correia, Robert MacKinnon, and Jeffrey Carver. Longlining haddock with manufactured bait to reduce catch of Atlantic cod in a conservation zone. Fisheries Research, 94(2):199–205, November 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002622.

Prevost:2005:SVD

[PCS05a] Etienne Prévost, Walter W. Crozier, and Pieter-Jan Schön. Static versus dynamic model for forecasting salmon prefishery abundance of the River Bush: a Bayesian comparison. Fisheries Research, 73(1–2):111–122, June 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000044.

Purchase:2005:SMS

[PCS05b] C. F. Purchase, N. C. Collins, and B. J. Shuter. Sensitivity of maximum sustainable harvest rates to intra-specific life history variability of lake trout (Salvelinus namaycush) and walleye (Sander vitreus). Fisheries Research, 72(2–3):141– 148, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604003042.

Polet:2004:ESN

[PCV04] Hans Polet, Johan Coenjaerts, and Reinhart Verschoore. Evaluation of the sieve net as a selectivity-improving device in the Belgian brown shrimp (Crangon crangon) fishery. Fisheries Research, 69(1):35–48, August 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000840.

Pampoulie:2008:RSI

[PD08] Christophe Pampoulie and Anna Kristín Daníelsdóttir. Resolving species identification problems in the genus Sebastes using nuclear genetic markers. Fisheries Research, 93(1–2):54–63, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000647.

Perez:2002:BPD

[PdAO02] José Angel Alvarez Perez, Daniela Cordella de Aguiar, and Uirá Cavalcante Oliveira. Biology and population dynamics of the long-finned squid Loligo plei (Cephalopoda: Loliginidae) in southern Brazilian waters. Fisheries Research, 58(3):267–279, November 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003976.

Punt:2008:ETM

[PDH08] André E. Punt, Martin W. Dorn, and Melissa A. Haltuch. Evaluation of threshold management strategies for groundfish off the U.S. West Coast. Fisheries Research, 94(3):251–266, December 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003566.

Polet:2005:EFBa

[PDV05a] Hans Polet, Fernand Delanghe, and Reinhart Verschoore. On electrical fishing for brown shrimp (Crangon crangon): I. Laboratory experiments. Fisheries Research, 72(1):1– 12, April 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604002814.

Polet:2005:EFBb

[PDV05b] Hans Polet, Fernand Delanghe, and Reinhart Verschoore. On electrical fishing for brown shrimp (Crangon crangon): II. Sea trials. Fisheries Research, 72(1):13–27, April 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002826.

Perez-Espana:2006:FTW

[PEAAdLJB06] Horacio Pérez-España, Luis Gerardo Abarca-Arenas, and María de Lourdes Jiménez-Badillo. Is fishing-down trophic web a generalized phenomenon?: the case of Mexican fisheries. Fisheries Research, 79(3):349–352, July 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001317.

Perez-Espana:2007:RAF

[PEAAdLJB07] Horacio Pérez-España, Luis G. Abarca-Arenas, and María de Lourdes Jiménez-Badillo. Reply: the absence of “fishing-down food web” does not mean a healthy fishery. Fisheries Research, 85(1–2):235–236, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000045.

Palsson:2003:SEU

[PEB03] Ólafur K. Pálsson, Haraldur A. Einarsson, and Höskuldur Björnsson. Survival experiments of undersized cod in a hand-line fishery at Iceland. Fisheries Research, 61(1– 3):73–86, March 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783602002485.

Penczak:2000:FSA

[PF00] Tadeusz Penczak and Nuno Formigo. Fish stocks in the Âncora River, north Portugal: Food consumption estimated with a bioenergetics model. Fisheries Research, 49(2):117–128, December 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002022.

Parente:2008:SIF

[PFHC08] Joaquim Parente, Paulo Fonseca, Victor Henriques, and Aida Campos. Strategies for improving fuel efficiency in the Portuguese trawl fishery. Fisheries Research, 93(1–2):117–124, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000866.

Parry:2006:ESS

[PG06] Gregory D. Parry and Anne Gason. The effect of seismic surveys on catch rates of rock lobsters in western Victoria, Australia. Fisheries Research, 79(3):272–284, July 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001330.

Pon:2007:ELC

[PGF07] J. P. Seco Pon, P. A. Gandini, and M. Favero. Effect of longline configuration on seabird mortality in the Argentine semi-pelagic kingclip Genypterus blacodes fishery. Fisheries Research, 85(1–2):101–105, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000112.

Pilling:2003:UOW

[PGK03] G. M. Pilling, E. M. Grandcourt, and G. P. Kirkwood. The utility of otolith weight as a predictor of age in the emperor Lethrinus mahsena and other tropical fish species. Fisheries Research, 60(2–3):493–506, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000875.

Pinho:2001:SAB

[PGMM01] Mário R. Pinho, João M. Gonçalves, Helen R. Martins, and Gui M. Menezes. Some aspects of the biology of the deepwater crab, Chaceon affinis (Milne-Edwards and Bouvier, 1894) off the Azores. Fisheries Research, 51(2–3):283–295, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601002521.

Pajuelo:2008:BPB

[PGS^{+}08] José G. Pajuelo, José A. González, José I. Santana, José M. Lorenzo, Antonio García-Mederos, and Víctor Tuset. Biological parameters of the bathyal fish black scabbardfish (Aphanopus carbo Lowe, 1839) off the Canary Islands, Central-east Atlantic. Fisheries Research, 92(2–3):140–147, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608000039.

Pinhorn:2001:REP

[PH01] A. T. Pinhorn and R. G. Halliday. The regulation of exploitation pattern in North Atlantic groundfish stocks. Fisheries Research, 53(1):25–37, September 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360000271X.

Pascoe:2004:ECF

[PH04a] Sean Pascoe and Ines Herrero. Estimation of a composite fish stock index using data envelopment analysis. Fisheries Research, 69(1):91–105, August 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000797.

Peres:2004:AGS

[PH04b] Mônica Brick Peres and Manuel Haimovici. Age and growth of southwestern Atlantic wreckfish Polyprion americanus. Fisheries Research, 66(2–3):157–169, February 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002078.

Paul:2009:AGS

[PH09] L. J. Paul and P. L. Horn. Age and growth of sea perch (Helicolenus percoides) from two adjacent areas off the east coast of South Island, New Zealand. Fisheries Research, 95(2–3):169–180, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002580.

Punt:2006:MGR

[PHGT06] André E. Punt, David Hobday, Joy Gerhard, and Vladimir S. Troynikov. Modelling growth of rock lobsters, Jasus edwardsii, off Victoria, Australia using models that allow for individual variation in growth parameters. Fisheries Research, 82(1–3):119–130, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002918.

Poulsen:2007:LTP

[PHM07] Bo Poulsen, Poul Holm, and Brian R. MacKenzie. A long-term (1667–1860) perspective on impacts of fishing and environmental variability on fisheries for herring, eel, and whitefish in the Limfjord, Denmark. Fisheries Research, 87(2–3):181–195, November 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001853.

Perez:2005:FBA

[PIdlHGV05] Juliana Pérez, Jorge I. Izquierdo, Jerónimo de la Hoz, and Eva Garcia-Vazquez. Female biased angling harvests of Atlantic salmon in Spain. Fisheries Research, 74(1–3):

127–133, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783605000871.

Pierce:2009:BRD

[Pie09] Graham J. Pierce. Book review: La Descomposición de la Ecología (The Decomposition of Ecology), A. Guerra Sierra, S. Pascual del Hierro. Netbiblo, Spain (2008), 129 pp., hardback, ISBN: 978-84-9745-235-9. Fisheries Research, 99(2):134, August 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001106.

Prasad:2005:GMY

[PJR^{+}05] R. Raja Prasad, A. K. Jaiswar, Suneel B. Reddy, S. K. Chakraborty, R. Palaniswamy, and Pranaya Parida. Growth, mortality and yield per recruit of Polynemus heptadactylus (Cuvier) (2Teleostei: Polynemidae) from Mumbai waters, India. Fisheries Research, 76(1):155–161, October 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605001761.

Papaconstantinou:2003:BGR

[PK03a] Costas Papaconstantinou and Konstantinos Kapiris. The biology of the giant red shrimp (Aristaeomorpha foliacea) at an unexploited fishing ground in the Greek Ionian Sea. Fisheries Research, 62(1):37–51, April 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002540.

Pitt:2003:TVV

[PK03b] Kylie A. Pitt and Michael J. Kingsford. Temporal variation in the virgin biomass of the edible jellyfish, Catostylus mosaicus (Scyphozoa, Rhizostomeae). Fisheries Research, 63(3):303–313, September 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000791.

Purcell:2006:RSC

[PK06] Steven W. Purcell and David S. Kirby. Restocking the sea cucumber Holothuria scabra: Sizing no-take zones through individual-based movement modelling. Fisheries Research, 80(1):53–61, August 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001202.

Power:2006:PBA

[PKK^{+}06] G. R. Power, P. A. King, C. J. Kelly, D. McGrath, E. Mullins, and Ole Gullaksen. Precision and bias in the age determination of blue whiting, Micromesistius poutassou (Risso, 1810), within and between age-readers. Fisheries Research, 80(2–3):312–321, September 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001342.

Pickett:2004:PRS

[PKP04] G. D. Pickett, D. F. Kelley, and M. G. Pawson. The patterns of recruitment of sea bass, Dicentrarchus labrax L. from nursery areas in England and Wales and implications for fisheries management. Fisheries Research, 68(1–3):329– 342, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604000190.

Prchalova:2008:OPF

[PKŘ^{+}08] Marie Prchalová, Jan Kubečka, Milan Říha, Radek Litvín, Martin Čech, Jaroslava Frouzová, Milan Hladík, Eva Hohausová, Jiří Peterka, and Mojmír Vašek. Overestimation of percid fishes (Percidae) in gillnet sampling. Fisheries Research, 91(1):79–87, May 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003165.

Prchalova:2009:SSS

[PKŘ^{+}09] Marie Prchalová, Jan Kubečka, Milan Říha, Tomáš Mrkvička, Mojmír Vašek, Tomáš Jůza, Michal Kratochvíl, Jiří Peterka, Vladislav Draštík, and Josef Křížek. Size selectivity of standardized multimesh gillnets in sampling coarse European species. Fisheries Research, 96(1):51–57, February 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608002804.

Papaconstantinou:2002:DSC

[PKT^{+}02] C. Papaconstantinou, S. Kavadas, N. Tsimenidis, A. N. Economou, and G. Bazigos. Design of a system for the collection and compilation of basic fisheries statistics in the Mediterranean: a case study for Greece. Fisheries Research, 57(2):155–167, August 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003484.

Popper:2000:SFR

[PL00]  Arthur N. Popper and Zhongmin Lu.  Structure–

function relationships in fish otolith organs. Fisheries Research, 46(1–3):15–25, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001296.

Pajuelo:2002:GAE

[PL02] José G. Pajuelo and José M. Lorenzo. Growth and age estimation of Diplodus sargus cadenati (Sparidae) off the Canary Islands. Fisheries Research, 59(1–2):93–100, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601004210.

Pradhan:2004:MTC

[PL04] Naresh C. Pradhan and PingSun Leung. Modeling trip choice behavior of the longline fishers in Hawaii. Fisheries Research, 68(1–3):209–224, July 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360400030X.

Pradhan:2006:PNB

[PL06] Naresh C. Pradhan and PingSun Leung. A Poisson and negative binomial regression model of sea turtle interactions in Hawaii’s longline fishery. Fisheries Research, 78(2–3):309– 322, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606000166.

Psuty-Lipska:2005:FPV

[PLD05] Iwona Psuty-Lipska and Bohdan Draganik. Fishery practice versus experimental design: Preliminary results of the introduction of protective sieves in the eel fyke-net fishery of the Vistula Lagoon, Poland. Fisheries Research, 76(1):146–154, October 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001724.

Pajuelo:2003:AEG

[PLDS03] José G. Pajuelo, José M. Lorenzo, and Rosa Domínguez-Seoane. Age estimation and growth of the zebra seabream Diplodus cervinus cervinus (Lowe, 1838) on the Canary Islands shelf (Central-east Atlantic). Fisheries Research, 62(1):97–103, April 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002497.

Psuty-Lipska:2006:GNS

[PLMDB06] Iwona Psuty-Lipska, Niels Madsen, Bohdan Draganik, and Wiesław Blady. Gill net selectivity for perch (Perca fluviatilis) in the Szczecin Lagoon, Poland. Fisheries Research, 80(2–3):339–344, September 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001445.

Pereira:2009:CTI

[PLPN09] Júlio César Pereira, Roseli Aparecida Leandro, Miguel Petrere, and Tom Nishida. Comparing three indices of catch per unit effort using Bayesian geostatistics. Fisheries Research, 100(3):200–209, November 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002045.

Prellezo:2009:QQA

[PLSI09] R. Prellezo, I. Lazkano, M. Santurtún, and A. Iriondo. A qualitative and quantitative analysis of selection of fishing area by Basque trawlers. Fisheries Research, 97(1–2):24– 31, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608004001.

Priede:2009:BSC

[PM09] Imants G. Priede and Peter I. Miller. A basking shark (Cetorhinus maximus) tracked by satellite together with simultaneous remote sensing II: New analysis reveals orientation to a thermal front. Fisheries Research, 95(2–3):370–372, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003202.

Park:2007:SSD

[PMAK07] Hae-Hoon Park, Russell B. Millar, Heui-Chun An, and Hyun-Young Kim. Size selectivity of drum-net traps for whelk (Buccinum opisoplectum dall) in the Korean coastal waters of the East Sea. Fisheries Research, 86(2–3):113–119, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001154.

Potier:2007:FFD

[PMC^{+}07] Michel Potier, Francis Marsac, Yves Cherel, Vincent Lucas, Richard Sabatié, Olivier Maury, and Frédéric Ménard. Forage fauna in the diet of three large pelagic fishes (lancetfish, swordfish and yellowfin tuna) in the western equatorial Indian Ocean. Fisheries Research, 83(1):60–72, January 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003079.

Pajuelo:2006:GPA

[PMG^{+}06] José G. Pajuelo, Iciar Martínez, José A. González, José M. Lorenzo, Antonio García-Mederos, Rosa Domínguez-Seoane, and Carmen M. Hernández-Cruz. Growth pattern and age estimation of the coastal sparid fish Pagrus auriga on the Canary Islands shelf. Fisheries Research, 82(1–3):7–13, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003158.

Pajaro:2005:RPP

[PMM05] Marcelo Pájaro, Gustavo J. Macchi, and Patricia Martos. Reproductive pattern of the Patagonian stock of Argentine hake (Merluccius hubbsi). Fisheries Research, 72(1):97– 108, April 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604002607.

Phillips:2003:EER

[PMSC03] Bruce F. Phillips, Roy Melville-Smith, and Yuk W. Cheng. Estimating the effects of removing Panulirus cygnus pueruli on the fishery stock. Fisheries Research, 65(1–3):89–101, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603002376.

Pindaro:2006:STM

[PMSD06] Díaz-Jaimes Píndaro, Uribe-Alcocer Manuel, Ortega-García Sofía, and Jean-Dominique Durand. Spatial and temporal mitochondrial DNA genetic homogeneity of dolphinfish populations (Coryphaena hippurus) in the eastern central Pacific. Fisheries Research, 80(2–3):333– 338, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606001688.

Peltonen:2006:HST

[PMT06] Heikki Peltonen, Tommi Malinen, and Antti Tuomaala. Hydroacoustic in situ target strength of smelt (Osmerus eperlanus (L.)). Fisheries Research, 80(2–3):190–195, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360600172X.

Potter:2003:MEM

[PMWC03] E. C. E. Potter, J. C. MacLean, R. J. Wyatt, and R. N. B. Campbell. Managing the exploitation of migratory salmonids. Fisheries Research, 62(2):127–142, May 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602001595.

Pedersen:2006:ARK

[PNJS06] O. P. Pedersen, E. M. Nilssen, L. L. Jørgensen, and D. Slagstad. Advection of the red king crab larvae on the coast of North Norway — a Lagrangian model study. Fisheries Research, 79(3):325–336, July 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001081.

Polet:2000:CWT

[Pol00] H. Polet. Codend and whole trawl selectivity of a shrimp beam trawl used in the North Sea. Fisheries Research, 48(2):167–183, September 1, 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001259.

Polet:2002:SES

[Pol02] H. Polet. Selectivity experiments with sorting grids in the North Sea brown shrimp (Crangon crangon) fishery. Fisheries Research, 54(2):217–233, January 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002897.

Poulard:2001:DHM

[Pou01] Jean-Charles Poulard. Distribution of hake (Merluccius merluccius, Linnaeus, 1758) in the Bay of Biscay and the Celtic sea from the analysis of French commercial data. Fisheries Research, 50(1–2):173–187, February 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002502.

Pitcher:2001:RRA

[PP01] Tony J. Pitcher and David Preikshot. RAPFISH: a rapid appraisal technique to evaluate the sustainability status of fisheries. Fisheries Research, 49(3):255–270, January 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002058.

Pedersen:2003:SMM

[PP03] Torstein Pedersen and John G. Pope. Sampling and a mortality model of a Norwegian fjord cod (Gadus morhua L.) population. Fisheries Research, 63(1):1–20, July 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000614.

Perez:2005:BAM

[PPdA05] José Angel Alvarez Perez, Paulo Ricardo Pezzuto, and Humber Agrelli de Andrade. Biomass assessment of the monkfish Lophius gastrophysus stock exploited by a new deep-water fishery in southern Brazil. Fisheries Research, 72(2–3):149–162, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604003029.

Piatkowski:2001:ICF

[PPdC01a] U. Piatkowski, G. J. Pierce, and M. Morais da Cunha. Impact of cephalopods in the food chain and their interaction with the environment and fisheries: an overview. Fisheries Research, 52(1–2):5–10, June 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002260.

Piatkowski:2001:P

[PPdC01b] U. Piatkowski, G. J. Pierce, and M. Morais da Cunha. Preface. Fisheries Research, 52(1–2):1, June 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002247.

Prince:2008:NUH

[PPGD08] Jeremy D. Prince, Harry Peeters, Harry Gorfine, and Robert W. Day. The novel use of harvest policies and rapid visual assessment to manage spatially complex abalone resources (genus Haliotis). Fisheries Research, 94(3):330–338, December 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002440.

Piatkowski:2001:CPK

[PPH01] Uwe Piatkowski, Klemens Pütz, and Heidrun Heinemann. Cephalopod prey of king penguins (Aptenodytes patagonicus) breeding at Volunteer Beach, Falkland Islands, during austral winter 1996. Fisheries Research, 52(1–2):79– 90, June 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601002326.

Palmer:2004:DBG

[PPL04] M. Palmer, G. X. Pons, and M. Linde. Discriminating between geographical groups of a Mediterranean commercial clam (Chamelea gallina (L.): Veneridae) by shape analysis. Fisheries Research, 67(1):93–98, March 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002650.

Pawson:2005:RTM

[PPS05] M. G. Pawson, G. D. Pickett, and M. T. Smith. The role of technical measures in the recovery of the UK sea bass (Dicentrarchus labrax) fishery 1980–2002. Fisheries Research, 76(1):91–105, October 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001670.

Paramo:2003:AGA

[PR03] Jorge Páramo and Rubén Roa. Acoustic–geostatistical assessment and habitat–abundance relations of small pelagic fish from the Colombian Caribbean. Fisheries Research, 60(2–3):309–319, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360200142X.

Prager:2002:CLG

[Pra02] Michael H. Prager. Comparison of logistic and generalized surplus-production models applied to swordfish, Xiphias gladius, in the north Atlantic Ocean. Fisheries Research, 58(1):41–57, October 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003587.

Prager:2003:RLE

[Pra03] Michael H. Prager. Reply to the letter to the Editor by Maunder. Fisheries Research, 61(1–3):151–154, March 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602002746.

Pineiro:2007:TRE

[PRdPG07] Carmen Piñeiro, Javier Rey, Helene de Pontual, and Raquel Goñi. Tag and recapture of European hake (Merluccius merluccius L.) off the Northwest Iberian Peninsula: First results support fast growth hypothesis. Fisheries Research, 88(1–3):150–154, December 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360700207X.

Pineiro:2008:GNI

[PRdPG08] C. Piñeiro, J. Rey, H. de Pontual, and A. García. Growth of Northwest Iberian juvenile hake estimated by combining sagittal and transversal otolith microstructure analyses. Fisheries Research, 93(1–2):173–178, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001136.

Priour:2009:NOT

[Pri09] Daniel Priour. Numerical optimisation of trawls design to improve their energy efficiency. Fisheries Research, 98 (1–3):40–50, June 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783609000770.

Poole:2004:EPL

[PRM04] W. R. Poole, J. D. Reynolds, and C. Moriarty. Early postlarval growth and otolith patterns in the eel Anguilla anguilla. Fisheries Research, 66(1):107–114, January 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360300105X.

Poot-Salazar:2009:API

[PSPCVCD09] Alicia Poot-Salazar, Roberto Pérez-Castañeda, María Eugenia Vega-Cendejas, and Omar Defeo. Assessing patterns of ichthyofauna discarded by an artisanal shrimp fishery through selectivity experiments in a coastal lagoon. Fisheries Research, 97(3):155–162, May 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000411.

Pierce:2007:HTI

[PSS^{+}07] G. J. Pierce, M. B. Santos, C. Smeenk, A. Saveliev, and A. F. Zuur. Historical trends in the incidence of strandings of sperm whales (Physeter macrocephalus) on North Sea coasts: an association with positive temperature anomalies. Fisheries Research, 87(2–3):219–228, November 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001257.

Punt:2006:IDD

[PSTM06] André E. Punt, David C. Smith, Geoff N. Tuck, and Richard D. Methot, Jr. Including discard data in fisheries stock assessments: Two case studies from south-eastern Australia. Fisheries Research, 79(3):239–250, July 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001482.

Pet-Soede:2001:IIC

[PSvDPM01] C. Pet-Soede, W. L. T. van Densen, J. S. Pet, and M. A. M. Machiels. Impact of Indonesian coral reef fisheries on fish community structure and the resultant catch composition. Fisheries Research, 51(1):35–51, April 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002368.

Probst:2009:HOS

[PTE09] Wolfgang Nikolaus Probst, Gregor Thomas, and Reiner Eckmann. Hydroacoustic observations of surface shoaling behaviour of young-of-the-year perch Perca fluviatilis (Linnaeus, 1758) with a towed upward-facing transducer. Fisheries Research, 96(2–3):133–138, March 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003342.

Pedersen:2009:ARL

[PTPS09] O. P. Pedersen, K. S. Tande, T. Pedersen, and D. Slagstad. Advection and retention as life trait modulators of capelin larvae — a case study from the Norwegian coast and the Barents Sea. Fisheries Research, 97(3):234–242, May 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000484.

Pinheiro:2005:GMV

[PTR^{+}05] Ana Pinheiro, Célia M. Teixeira, Ana Luísa Rego, Joana F. Marques, and Henrique N. Cabral. Genetic and morphological variation of Solea lascaris (Risso, 1810) along the Portuguese coast. Fisheries Research, 73(1–2):67–78, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605000202.

Punt:2003:PSS

[Pun03] André E. Punt. The performance of a size-structured stock assessment method in the face of spatial heterogeneity in growth. Fisheries Research, 65(1–3):391–409, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603002571.

Puga:2005:BMR

[PVMdL05] Rafael Puga, Sergio Hernández Vázquez, Juana López Martinez, and María E. de León. Bioeconomic modelling and risk assessment of the Cuban fishery for spiny lobster Panulirus argus. Fisheries Research, 75(1–3):149–163, September 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605001359.

Punzon:2004:AHF

[PVP04] Antonio Punzón, Begoña Villamor, and Izaskun Preciado. Analysis of the handline fishery targeting mackerel (Scomber scombrus, L.) in the North of Spain (ICES Division VIIIbc). Fisheries Research, 69(2):189–204, September 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001225.

Perez:2005:BAG

[PW05] José Angel Alvarez Perez and Roberto Wahrlich. A bycatch assessment of the gillnet monkfish Lophius gastrophysus fishery off southern Brazil. Fisheries Research, 72(1):81– 95, April 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604002620.

Pope:2005:MSF

[PWB05] Kevin L. Pope, Gene R. Wilde, and Daryl L. Bauer. Maximum size of fish caught with standard gears and recreational angling. Fisheries Research, 76(1):117–122, October 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605001694.

Piner:2006:EAA

[PWHM06] Kevin R. Piner, John R. Wallace, Owen S. Hamel, and Robert Mikus. Evaluation of ageing accuracy of bocaccio (Sebastes paucispinis) rockfish using bomb radiocarbon.

Fisheries Research, 77(2):200–206, February 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002900.

Punt:2002:AMR

[PWP02] André E. Punt, Terence I. Walker, and Jeremy D. Prince. Assessing the management-related benefits of fixed-station fishery-independent surveys in Australia’s southern shark fishery. Fisheries Research, 55(1–3):281–295, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002764.

Punt:2000:SCE

[PWTP00] André E. Punt, Terence I. Walker, Bruce L. Taylor, and Fred Pribac. Standardization of catch and effort data in a spatially-structured shark fishery. Fisheries Research, 45(2):129–145, March 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369900106X.

Perry:2002:DGS

[PZH02] R. Ian Perry, Ziyang Zhang, and Rick Harbo. Development of the green sea urchin (Strongylocentrotus droebachiensis) fishery in British Columbia, Canada — back from the brink using a precautionary framework. Fisheries Research, 55(1–3):253–266, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002831.

Queirolo:2009:CBD

[QDH09] Dante Queirolo, Harold DeLouche, and Carlos Hurtado. Comparison between dynamic simulation and physical model testing of new trawl design for Chilean crustacean fisheries. Fisheries Research, 97(1–2):86–94, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000198.

Quinn:2005:MSD

[QDV05] Thomas P. Quinn, Bobette R. Dickerson, and Leif Asbjørn Vøllestad. Marine survival and distribution patterns of two Puget Sound hatchery populations of Coho (Oncorhynchus kisutch) and Chinook (Oncorhynchus tshawytscha) salmon. Fisheries Research, 76(2):209–220, November 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001797.

Quist:2004:FRG

[QGBS04] M. C. Quist, C. S. Guy, R. J. Bernot, and J. L. Stephen. Factors related to growth and survival of larval walleyes: implications for recruitment in a southern Great Plains reservoir. Fisheries Research, 67(2):215–225, April 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002947.

Quintanilla:2000:SFS

[QP00] Luis Fco Quintanilla and Nélida Pérez. Spawning frequency of Sardina pilchardus (Walb.) off the Spanish North Atlantic coast in 1997. Fisheries Research, 45(1):73–79, February 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699001034.

Quirijns:2008:SCC

[QPR08] F. J. Quirijns, J. J. Poos, and A. D. Rijnsdorp. Standardizing commercial CPUE data in monitoring stock dynamics: Accounting for targeting behaviour in mixed fisheries. Fisheries Research, 89(1):1–8, January 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002147.

Quiros:2007:TCC

[Qui07] Angela L. Quiros. Tourist compliance to a code of conduct and the resulting effects on whale shark (Rhincodon typus) behavior in Donsol, Philippines. Fisheries Research, 84(1):102–108, March 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004061.

Quinonez-Velazquez:2000:GHD

[QVNMGM00] Casimiro Quiñonez-Velázquez, Manuel Otilio Nevarez-Martínez, and Marıá Georgina Gluyas-Millán. Growth and hatching dates of juvenile Pacific sardine Sardinops caeruleus in the Gulf of California. Fisheries Research, 48(2):99–106, September 1, 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360000179X.

Rosas-Alayola:2002:DCS

[RAHHGM^{+}02] José Rosas-Alayola, Agustín Hernández-Herrera, Felipe Galvan-Magaña, L. Andres Abitia-Cárdenas, and Arturo F. Muhlia-Melo. Diet composition of sailfish (Istiophorus platypterus) from the southern Gulf of California, Mexico. Fisheries Research, 57(2):185–195, August 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003447.

Roel:2000:ASA

[RB00] B. A. Roel and D. S. Butterworth. Assessment of the South African chokka squid Loligo vulgaris reynaudii: Is disturbance of aggregations by the recent jig fishery having a negative impact on recruitment? Fisheries Research, 48(3):213–228, October 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001867.

Rosenkranz:2004:VSS

[RB04] Gregg E. Rosenkranz and Susan C. Byersdorfer. Video scallop survey in the eastern Gulf of Alaska, USA. Fisheries Research, 69(1):131–140, August 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000785.

Ryer:2006:IIT

[RB06] Clifford H. Ryer and Lewis A. K. Barnett. Influence of illumination and temperature upon flatfish reactivity and herding behavior: Potential implications for trawl capture efficiency. Fisheries Research, 81(2–3):242–250, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606002621.

Raemaekers:2009:PIA

[RB09] Serge J.-P. N. Raemaekers and Peter J. Britz. Profile of the illegal abalone fishery (Haliotis midae) in the Eastern Cape Province, South Africa: Organised pillage and management failure. Fisheries Research, 97(3):183–195, May 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609000435.

Ragonese:2002:TCE

[RBD02] S. Ragonese, M. L. Bianchini, and L. Di Stefano. Trawl cod-end selectivity for deepwater red shrimp (Aristaeomorpha foliacea, Risso 1827) in the Strait of Sicily (Mediterranean Sea). Fisheries Research, 57(2):131–144, August 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601003423.

Rabaut:2008:EBT

[RBH^{+}08] Marijn Rabaut, Ulrike Braeckman, Frederik Hendrickx, Magda Vincx, and Steven Degraer. Experimental beamtrawling in Lanice conchilega reefs: Impact on the associated fauna. Fisheries Research, 90(1–3):209–216, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607002767.

Reyes-Bonilla:2003:PPE

[RBHP03] Héctor Reyes-Bonilla and Marıá Dinorah Herrero-Pérezrul. Population parameters of an exploited population of Isostichopus fuscus (Holothuroidea) in the southern Gulf of California, México. Fisheries Research, 59(3):423–430, January 3, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602000231.

Rahman:2006:LPG

[RC06] M. J. Rahman and I. G. Cowx. Lunar periodicity in growth increment formation in otoliths of Hilsa shad (Tenualosa ilisha, Clupeidae) in Bangladesh waters. Fisheries Research, 81(2–3):342–344, November 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002396.

Revill:2007:SGN

[RCA^{+}07] Andrew Revill, John Cotter, Mike Armstrong, Jon Ashworth, Rob Forster, Gus Caslake, and Rene Holst. The selectivity of the gill-nets used to target hake (Merluccius merluccius) in the Cornish and Irish offshore fisheries. Fisheries Research, 85(1–2):142–147, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360700032X.

Rapp:2008:ESF

[RCA08] Tobias Rapp, Steven J. Cooke, and Robert Arlinghaus. Exploitation of specialised fisheries resources: the importance of hook size in recreational angling for large common carp (Cyprinus carpio L.). Fisheries Research, 94(1):79–83, October 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608001987.

Revill:2007:RWI

[RCD07] A. S. Revill, T. L. Catchpole, and G. Dunlin. Recent work to improve the efficacy of square-mesh panels used in a North Sea Nephrops norvegicus directed fishery. Fisheries Research, 85(3):321–327, July 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000884.

Romanelli:2002:GME

[RCG02] M. Romanelli, F. Colloca, and O. Giovanardi. Growth and mortality of exploited Sardina pilchardus (Walbaum) larvae along the western coast of Italy. Fisheries Research, 55(1–3):205–218, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002867.

Ramirez:2004:SVR

[RCGC04] Teodoro Ramírez, Dolores Cortés, Alberto Garcıá, and Angel Carpena. Seasonal variations of RNA/DNA ratios and growth rates of the Alboran Sea sardine larvae (Sardina pilchardus). Fisheries Research, 68(1–3):57–65, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000724.

Ribot-Carballal:2005:AGS

[RCGMQV05] M. C. Ribot-Carballal, F. Galván-Magaña, and C. Quiñónez-Velázquez. Age and growth of the shortfin mako shark, Isurus oxyrinchus, from the western coast of Baja California Sur, Mexico. Fisheries Research, 76(1):14–21, October 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605001414.

Rodriguez-Cabello:2004:LSD

[RCSFO04] C. Rodríguez-Cabello, F. Sánchez, A. Fernández, and I. Olaso. Is the lesser spotted dogfish (Scyliorhinus canicula) population from the Cantabrian Sea a unique stock? Fisheries Research, 69(1):57–71, August 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000815.

Revill:2003:FCG

[RD03] Andrew S. Revill and Gary Dunlin. The fishing capacity of gillnets lost on wrecks and on open ground in UK coastal waters. Fisheries Research, 64(2–3):107–113, November 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603002091.

Rodriguez-Dowdell:2007:PRB

[RDEACT07] Natalie Rodríguez-Dowdell, Roberto Enríquez-Andrade, and Nirari Cárdenas-Torres. Property rights-based management: Whale shark ecotourism in Bahia de los Angeles, Mexico. Fisheries Research, 84(1):119–127, March 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004097.

Reid:2001:RSN

[RdFBS01] Philip C. Reid, Maria de Fatima Borges, and Einar Svendsen. A regime shift in the North Sea circa 1988 linked to changes in the North Sea horse mackerel fishery. Fisheries Research, 50(1–2):163–171, February 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002496.

Revill:2005:SDL

[RDH05] Andrew S. Revill, Nicholas K. Dulvy, and Rene Holst. The survival of discarded lesser-spotted dogfish (Scyliorhinus canicula) in the Western English Channel beam trawl fishery. Fisheries Research, 71(1):121–124, January 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001870.

Revill:2006:SPC

[RDH06] Andrew Revill, Gary Dunlin, and René Holst. Selective properties of the cutaway trawl and several other commercial trawls used in the Farne Deeps North Sea Nephrops fishery. Fisheries Research, 81(2–3):268–275, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606002451.

Richmond:2009:RRB

[RDM^{+}09] Marshall C. Richmond, Zhiqun Deng, Craig A. McKinstry, Robert P. Mueller, Thomas J. Carlson, and Dennis D. Dauble. Response relationships between juvenile salmon and an autonomous sensor in turbulent flow.

Fisheries Research, 97(1–2):134–139, April 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360900040X.

Rowat:2007:SCS

[RE07] David Rowat and Udo Engelhardt. Seychelles: a case study of community involvement in the development of whale shark ecotourism and its socio-economic impact. Fisheries Research, 84(1):109–113, March 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004073.

Reid:2001:SSE

[Rei01] D. G. Reid. SEFOS — shelf edge fisheries and oceanography studies: an overview. Fisheries Research, 50(1–2):1–15, February 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600002381.

Ressler:2009:DCV

[RFWH09] Patrick H. Ressler, Guy W. Fleischer, Vidar G. Wespestad, and John Harms. Developing a commercial-vessel-based stock assessment survey methodology for monitoring the U.S. west coast widow rockfish (Sebastes entomelas) stock. Fisheries Research, 99(2):63–73, August 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360900099X.

Rowat:2007:RSH

[RG07] D. Rowat and M. Gore. Regional scale horizontal and local scale vertical movements of whale sharks in the Indian Ocean off Seychelles. Fisheries Research, 84(1):32–40, March 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606003985.

Robinson:2007:EHM

[RGGB07] Tamara B. Robinson, Anesh Govender, Charles L. Griffiths, and George M. Branch. Experimental harvesting of Mytilus galloprovincialis: Can an alien mussel support a small-scale fishery? Fisheries Research, 88(1–3):33–41, December 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360700166X.

Rosenkranz:2008:DHS

[RGSB08] Gregg E. Rosenkranz, Scott M. Gallager, Richard W. Shepard, and Mark Blakeslee. Development of a high-speed, megapixel benthic imaging system for coastal fisheries research in Alaska. Fisheries Research, 92(2–3):340–344, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608000982.

Revill:2004:SPS

[RH04a] Andrew Revill and Rene Holst. The selective properties of some sieve nets. Fisheries Research, 66(2–3):171–183, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360300198X.

Revill:2004:RDN

[RH04b] Andrew S. Revill and Rene Holst. Reducing discards of North Sea brown shrimp (C. crangon) by trawl modification. Fisheries Research, 68(1–3):113–122, July 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000487.

Richards:2006:ENG

[RH06] Anne Richards and Lisa Hendrickson. Effectiveness of the nordmore grate in the Gulf of Maine northern shrimp fishery. Fisheries Research, 81(1):100–106, October 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001810.

Rakowitz:2008:TMI

[RHF^{+}08] Georg Rakowitz, Wolfgang Herold, Christian Fesl, Hubert Keckeis, Jan Kubečka, and Helge Balk. Two methods to improve the accuracy of target-strength estimates for horizontal beaming. Fisheries Research, 93(3):324–331, September 28, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608001847.

Revill:2005:CBR

[RJ05] Andrew S. Revill and Simon Jennings. The capacity of benthos release panels to reduce the impacts of beam trawls on benthic communities. Fisheries Research, 75(1–3):73–85, September 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001451.

Richardson:2006:TSA

[RKEJR06] Elizabeth A. Richardson, Michel J. Kaiser, Gareth Edwards-Jones, and Kirsten Ramsay. Trends in sea anglers’ catches of trophy fish in relation to stock size. Fisheries Research, 82(1–3):253–262, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002177.

Ratz:2003:VFC

[RL03] Hans-Joachim Rätz and Josep Lloret. Variation in fish condition between Atlantic cod (Gadus morhua) stocks, the effect on their productivity and management implications. Fisheries Research, 60(2–3):369–380, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001327.

Rico:2001:AGA

[RLG^{+}01] V. Rico, J. M. Lorenzo, J. A. González, H. M. Krug, A. Mendonça, E. Gouveia, and M. Afonso Dias. Age and growth of the alfonsino Beryx splendens Lowe, 1834 from the Macaronesian archipelagos. Fisheries Research, 49(3): 233–240, January 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360000206X.

Rudorff:2009:MSL

[RLGLO09] Camila A. G. Rudorff, João A. Lorenzzetti, Douglas F. M. Gherardi, and Jorge E. Lins-Oliveira. Modeling spiny lobster larval dispersion in the tropical Atlantic. Fisheries Research, 96(2–3):206–215, March 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003548.

Rios-Lara:2007:DPS

[RLSJIA07] Veronica Rios-Lara, Silvia Salas, Bello-Pineda Javier, and Peniche Irene-Ayora. Distribution patterns of spiny lobster (Panulirus argus) at Alacranes reef, Yucatan: Spatial analysis and inference of preferential habitat. Fisheries Research, 87(1):35–45, October 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001531.

Reid:2005:CSB

[RM05] D. D. Reid and S. S. Montgomery. Creel survey based estimation of recreational harvest of penaeid prawns in four southeastern Australian estuaries and comparison with commercial catches. Fisheries Research, 74(1–3):169–185, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605001086.

Rogan:2007:MBD

[RM07] E. Rogan and M. Mackey. Megafauna bycatch in drift nets for albacore tuna (Thunnus alalunga) in the NE Atlantic. Fisheries Research, 86(1):6–14, August 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000422.

Rufino:2005:EMO

[RMA^{+}05] M. M. Rufino, F. Maynou, P. Abelló, L. Gil de Sola, and A. B. Yule. The effect of methodological options on geostatistical modelling of animal distribution: a case study with Liocarcinus depurator (Crustacea: Brachyura) trawl survey data. Fisheries Research, 76(2):252–265, November 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002006.

Radebe:2002:AGR

[RMBG02] P. V. Radebe, B. Q. Mann, L. E. Beckley, and A. Govender. Age and growth of Rhabdosargus sarba (Pisces: Sparidae), from KwaZulu-Natal, South Africa. Fisheries Research, 58(2):193–201, November 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003836.

Ribas:2006:HDN

[RMCdS06] David Ribas, Marta Muñoz, Margarida Casadevall, and Luis Gil de Sola. How does the northern Mediterranean population of Helicolenus dactylopterus dactylopterus resist fishing pressure? Fisheries Research, 79(3):285–293, July 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606001329.

Relini:2006:BFE

[RMF^{+}06] L. Orsi Relini, A. Mannini, F. Fiorentino, G. Palandri, and G. Relini. Biology and fishery of Eledone cirrhosa in the Ligurian Sea. Fisheries Research, 78(1):72–88, April 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003334.

Robertson:2003:HFD

[RMHW03] Graham Robertson, Egil Moe, Roy Haugen, and Barbara Wienecke. How fast do demersal longlines sink? Fisheries Research, 62(3):385–388, June 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360200276X.

Rosa:2006:WLR

[RMMP06] A. Rosa, G. Menezes, O. Melo, and M. R. Pinho. Weight– length relationships of 33 demersal fish species from Azores archipelago. Fisheries Research, 80(2–3):329–332, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606001718.

Reglero:2008:RBF

[RMN08] Patricia Reglero and Beatriz Morales-Nin. Relationship between first sale price, body size and total catch of trammelnet target species in Majorca (NW Mediterranean). Fisheries Research, 92(1):102–106, July 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000519.

Rifflart:2006:SRV

[RMRB06] Renaud Rifflart, Frédéric Marchand, Etienne Rivot, and Jean-Luc Bagliniere. Scale reading validation for estimating age from tagged fish recapture in a brown trout (Salmo trutta) population. Fisheries Research, 78(2–3):380–384, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606000300.

Ramirez-Macias:2007:VMC

[RMVJGMMV07] Dení Ramírez-Macías, Ricardo Vázquez-Juárez, Felipe Galván-Magaña, and Adrián Munguía-Vega. Variations of the mitochondrial control region sequence in whale sharks (Rhincodon typus) from the Gulf of California, Mexico. Fisheries Research, 84(1):87–95, March 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004048.

Ryer:2000:AAN

[RO00] C. H. Ryer and B. L. Olla. Avoidance of an approaching net by juvenile walleye pollock Theragra chalcogramma in the laboratory: the influence of light intensity. Fisheries Research, 45(2):195–199, March 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001137.

Rodhouse:2001:MFS

[Rod01] P. G. Rodhouse. Managing and forecasting squid fisheries in variable environments. Fisheries Research, 54(1):3–8, December 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601003708.

Ryer:2004:BIA

[ROS04] Clifford H. Ryer, Michele L. Ottmar, and Erick A. Sturm. Behavioral impairment after escape from trawl codends may not be limited to fragile fish species. Fisheries Research, 66(2–3):261–269, February 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001978.

Rowe:2002:PPA

[Row02] Sherrylynn Rowe. Population parameters of American lobster inside and outside no-take reserves in Bonavista Bay, Newfoundland. Fisheries Research, 56(2):167–175, May 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601003216.

Rowat:2007:OWS

[Row07] David Rowat. Occurrence of whale shark (Rhincodon typus) in the Indian Ocean: a case for regional conservation. Fisheries Research, 84(1):96–101, March 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600405X.

Rakocinski:2006:DAF

[RPC^{+}06] Chet F. Rakocinski, Mark S. Peterson, Bruce H. Comyns, Glenn A. Zapfe, and Gregory L. Fulling. Do abiotic factors drive the early growth of juvenile spot (Leiostomus xanthurus)? Fisheries Research, 82(1–3):186–193, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606002323.

Royer:2006:ECS

[RPFR06] J. Royer, G. J. Pierce, E. Foucher, and J. P. Robin. The English Channel stock of Sepia officinalis: Modelling variability in abundance and impact of the fishery. Fisheries Research, 78(1):96–106, April 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003280.

Rahikainen:2004:UMN

[RPP04] Mika Rahikainen, Heikki Peltonen, and Jukka Pönni. Unaccounted mortality in northern Baltic Sea herring fishery — magnitude and effects on estimates of stock dynamics. Fisheries Research, 67(2):111–127, April 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360300300X.

Robichaud:2003:SDC

[RR03a] David Robichaud and George A. Rose. Sex differences in cod residency on a spawning ground. Fisheries Research, 60(1):33–43, January 30, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000656.

Ruggerone:2003:MYE

[RR03b] Gregory T. Ruggerone and Donald E. Rogers. Multiyear effects of high densities of sockeye salmon spawners on juvenile salmon growth and survival: a case study from the Exxon Valdez oil spill. Fisheries Research, 63(3):379–392, September 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000997.

Ramirez-Rodriguez:2003:LHS

[RRAS03] Mauricio Ramírez-Rodríguez and Francisco Arreguín-Sánchez. Life history stage duration and natural mortality for the pink shrimp Farfantepenaeus duorarum (Burkenroad, 1939) in the southern Gulf of Mexico, using the gnomonic model for time division. Fisheries Research, 60(1):45–51, January 30, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000644.

Ramirez-Rodriguez:2003:RPP

[RRASLB03] Mauricio Ramírez-Rodríguez, Francisco Arreguín-Sánchez, and Daniel Lluch-Belda. Recruitment patterns of the pink shrimp Farfantepenaeus duorarum in the southern Gulf of Mexico. Fisheries Research, 65(1–3):81–88, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002364.

Robillard:2008:RBB

[RRJ08] Eric Robillard, Christian S. Reiss, and Cynthia M. Jones. Reproductive biology of bluefish (Pomatomus saltatrix) along the East Coast of the United States. Fisheries Research, 90(1–3):198–208, April 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002743.

Robillard:2009:AVG

[RRJ09] Eric Robillard, Christian S. Reiss, and Cynthia M. Jones. Age-validation and growth of bluefish (Pomatomus saltatrix) along the East Coast of the United States. Fisheries Research, 95(1):65–75, January 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002245.

Rose:2005:UHF

[RSM05] Craig S. Rose, Allan W. Stoner, and Keith Matteson. Use of high-frequency imaging sonar to observe fish behaviour near baited fishing gears. Fisheries Research, 76(2):291–304, November 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605002043.

Rufino:2006:APL

[RSMZ06] Marta M. Rufino, Vanessa Stelzenmüller, Francesc Maynou, and Gert-Peter Zauke. Assessing the performance of linear geostatistical tools applied to artificial fisheries data. Fisheries Research, 82(1–3):263–279, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002165.

Reid:2000:SPA

[RSP^{+}00] D. Reid, C. Scalabrin, P. Petitgas, J. Masse, R. Aukland, P. Carrera, and S. Georgakarakos. Standard protocols for the analysis of school based data from echo sounder surveys. Fisheries Research, 47(2–3):125–136, July 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001648.

Rockmann:2007:TIP

[RSST07] Christine Röckmann, Michael A. St. John, Uwe A. Schneider, and Richard S. J. Tol. Testing the implications of a permanent or seasonal marine reserve on the population dynamics of Eastern Baltic cod under varying environmental conditions. Fisheries Research, 85(1–2):1–13, June 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004280.

Ruttan:2007:SSS

[RT07] Lore M. Ruttan and Peter H. Tyedmers. Skippers, spotters and seiners: Analysis of the “skipper effect” in US menhaden (Brevoortia spp.) purse-seine fisheries. Fisheries Research, 83(1):73–80, January 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003080.

Rueda:2007:ESP

[Rue07] Mario Rueda. Evaluating the selective performance of the encircling gillnet used in tropical estuarine fisheries from Colombia. Fisheries Research, 87(1):28–34, October 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001543.

Rocha:2003:OCF

[RV03] Francisco Rocha and Marco A. Vega. Overview of cephalopod fisheries in Chilean waters. Fisheries Research, 60(1):151–159, January 30, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000802.

Rudstam:2002:CAM

[RVS02] Lars G. Rudstam, Anthony J. VanDeValk, and Mark D. Scheuerell. Comparison of acoustic and Miller high-speed sampler estimates of larval fish abundance in Oneida Lake, New York. Fisheries Research, 57(2):145–154, August 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003472.

Rotherham:2003:CMS

[RW03] Douglas Rotherham and Ronald J. West. Comparison of methods for sampling populations of ghost shrimp, Trypaea australiensis (Decapoda: Thalassinidea: Callianassidae). Fisheries Research, 60(2–3):585–591, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001376.

Replinger:2007:PIU

[RW07] S. E. Replinger and J. B. Wood. A preliminary investigation of the use of subcutaneous tagging in Caribbean reef squid Sepioteuthis sepioidea (Cephalopoda: Loliginidae). Fisheries Research, 84(3):308–313, May 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003900.

Rodhouse:2006:FBH

[RWMBHH06] Paul G. Rodhouse, Claire M. Waluda, Enrique Morales-Bojórquez, and Agustín Hernández-Herrera. Fishery biology of the Humboldt squid, Dosidicus gigas, in the Eastern Pacific Ocean. Fisheries Research, 79(1–2):13–15, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606000622.

Reid:2001:HMD

[RWT01] D. G. Reid, M. Walsh, and W. R. Turrell. Hydrography and mackerel distribution on the shelf edge west of the Norwegian deeps. Fisheries Research, 50(1–2):141–150, February 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600002472.

Ryer:2008:RFB

[Rye08] Clifford H. Ryer. A review of flatfish behavior relative to trawls. Fisheries Research, 90(1–3):138–146, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002573.

Rodhouse:2001:PA

[RYG01] P. G. Rodhouse, C. Yamashiro, and C. Goss. Preface and acknowledgements. Fisheries Research, 54(1):1, December 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601003691.

Robotham:2008:JME

[RYSN08] Hugo Robotham, Zaida I. Young, and Juan C. Saavedra-Nievas. Jackknife method for estimating the variance of the age composition using two-phase sampling with an application to commercial catches of swordfish (Xiphias gladius). Fisheries Research, 93(1–2):135–139, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608000908.

Smith:2003:MSA

[SA03] Michael T. Smith and Julian T. Addison. Methods for stock assessment of crustacean fisheries. Fisheries Research, 65(1–3):231–256, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002467.

Sherman:2003:SLM

[SAA^{+}03] Kenneth Sherman, Thomas Ajayi, Emilia Anang, Philippe Cury, M. C. M. Pierre Freon, Nicholas J. Hardman-Mountford, Chidi A. Ibe, Kwame A. Koranteng, Jacqueline McGlade, C. E. C. Cornelia Nauen, Daniel Pauly, Peter A. G. M. Scheren, Hein R. Skjoldal, Qisheng Tang, Soko Guillaume Zabi, and Antonio J. Diaz de Leon. Suitability of the large marine ecosystem concept. Fisheries Research, 64(2–3):197–204, November 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002297.

Sypitkowski:2009:CSB

[SABW09] Eben Sypitkowski, William G. Ambrose, Curtis Bohlen, and Joseph Warren. Catch statistics in the bloodworm fishery in Maine. Fisheries Research, 96(2–3):303–307, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608002890.

Sinclair:2002:RFM

[SAC^{+}02] M. Sinclair, R. Arnason, J. Csirke, Z. Karnicki, J. Sigurjonsson, H. Rune Skjoldal, and G. Valdimarsson. Responsible fisheries in the marine ecosystem. Fisheries Research, 58(3):255–265, November 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001686.

Sousa:2005:DAP

[SAG05] Pedro Sousa, Manuela Azevedo, and Manuel C. Gomes. Demersal assemblages off Portugal: Mapping, seasonal, and temporal patterns. Fisheries Research, 75(1–3):120– 137, September 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783605001335.

Suda:2005:IEF

[SAK05] Maki Suda, Tatsuro Akamine, and Tatsu Kishida. Influence of environment factors, interspecific-relationships and fishing mortality on the stock fluctuation of the Japanese sardine, Sardinops melanostictus, off the Pacific coast of Japan. Fisheries Research, 76(3):368–378, December 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002134.

Salthaug:2002:DTC

[Sal02] Are Salthaug. Do triggers in crab traps affect the probability of entry? Fisheries Research, 58(3):403–405, November 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601003915.

Shelmerdine:2007:SVC

[SALnM07] Richard L. Shelmerdine, Jennifer Adamson, Chevonne H. Laurenson, and Beth Leslie (neé Mouat). Size variation of the common whelk, Buccinum undatum, over large and small spatial scales: Potential implications for micro-management within the fishery. Fisheries Research, 86(2–3):201–206, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001270.

Santos:2000:FOU

[San00] A. Miguel P. Santos. Fisheries oceanography using satellite and airborne remote sensing methods: a review. Fisheries Research, 49(1):1–20, November 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002010.

Schonhuth:2005:MIB

[SÁR^{+}05] Susana Schönhuth, Yolanda Álvarez, Vicente Rico, José A. González, José I. Santana, Elizabete Gouveia, José M. Lorenzo, and José M. Bautista. Molecular identification and biometric analysis of Macaronesian archipelago stocks of Beryx splendens. Fisheries Research, 73(3):299–309, July 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360500038X.

Santos:2001:TRD

[SB01a] Jorge Santos and Teresa Borges. Trophic relationships in deep-water fish communities off Algarve, Portugal. Fisheries Research, 51(2–3):337–341, May 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002570.

Silvano:2001:SDF

[SB01b] R. A. M. Silvano and A. Begossi. Seasonal dynamics of fishery at the Piracicaba River (Brazil). Fisheries Research, 51(1):69–86, April 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002290.

Sanchez:2003:BLC

[SB03] A. Sánchez and E. J. Belda. Bait loss caused by seabirds on longline fisheries in the northwestern Mediterranean: is night setting an effective mitigation measure? Fisheries Research, 60(1):99–106, January 30, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000553.

Silvano:2005:LKC

[SB05] R. A. M. Silvano and A. Begossi. Local knowledge on a cosmopolitan fish: Ethnoecology of Pomatomus saltatrix (Pomatomidae) in Brazil and Australia. Fisheries Research, 71(1):43–59, January 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001882.

Scholten:2007:LSS

[SB07] George D. Scholten and Phillip W. Bettoli. Lack of size selectivity for paddlefish captured in hobbled gillnets. Fisheries Research, 83(2–3):355–359, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003912.

Sharma:2008:SDH

[SB08] C. M. Sharma and R. Borgstrøm. Shift in density, habitat use, and diet of perch and roach: an effect of changed predation pressure after manipulation of pike. Fisheries Research, 91(1):98–106, May 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003189.

Svenning:2005:IMF

[SBD^{+}05] M.-A. Svenning, R. Borgstrøm, T. O. Dehli, G. Moen, R. T. Barrett, T. Pedersen, and W. Vader. The impact of marine fish predation on Atlantic salmon smolts (Salmo salar) in the Tana estuary, North Norway, in the presence of an alternative prey, lesser sandeel (Ammodytes marinus). Fisheries Research, 76(3):466–474, December 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605002031. See corrigen-

dum [SBD^{+}06].

Svenning:2006:CIM

[SBD^{+}06] M.-A. Svenning, R. Borgstrøm, T. O. Dehli, G. Moen, R. T. Barrett, T. Pedersen, and W. Vader. Corrigendum to “The impact of marine fish predation on Atlantic salmon smolts (Salmo salar) in the Tana estuary, North Norway, in the presence of an alternative prey, lesser sandeel (Ammodytes marinus)” [Fish. Res. 76 (2005) 466– 474]. Fisheries Research, 77(1):124, January 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002778. See [SBD^{+}05].

Stransky:2008:SNC

[SBF^{+}08] Christoph Stransky, Hannes Baumann, Svein-Erik Fevolden, Alf Harbitz, Hans Høie, Kjell H. Nedreaas, Arnt-Børre Salberg, and Tuula H. Skarstein. Separation of Norwegian coastal cod and Northeast Arctic cod by outer otolith shape analysis. Fisheries Research, 90(1–3):26–35, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607002305.

Salini:2000:ABD

[SBFR00] John Salini, David Brewer, Margaret Farmer, and Nick Rawlinson. Assessment and benefits of damage reduction in prawns due to use of different bycatch reduction devices in the Gulf of Carpentaria, Australia. Fisheries Research, 45(1):1–8, February 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001046.

Sobrino:2005:EEF

[SBGG^{+}05] I. Sobrino, F. Baldó, D. García-González, J. A. Cuesta, A. Silva-García, C. Fernández-Delgado, A. M. Arias, A. Rodríguez, and P. Drake. The effect of estuarine fisheries on juvenile fish observed within the Guadalquivir Estuary (SW Spain). Fisheries Research, 76(2):229–242, November 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605001980.

Suryanarayana:2008:NNF

[SBR^{+}08] Iragavarapu Suryanarayana, Antonio Braibanti, Rupenaguntla Sambasiva Rao, Veluri Anantha Ramam, Duvvuri Sudarsan, and Gollapalli Nageswara Rao. Neural networks in fisheries research. Fisheries Research, 92(2–3):115–139, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608000520.

Sigler:2007:SAT

[SC07] Michael F. Sigler and David J. Csepp. Seasonal abundance of two important forage species in the North Pacific Ocean, Pacific herring and walleye pollock. Fisheries Research, 83(2–3):319–331, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003699.

Seijo:2008:PLI

[SC08] J. C. Seijo and J. F. Caddy. Port location for inshore fleets affects the sustainability of coastal source–sink resources: Implications for spatial management of metapopulations. Fisheries Research, 91(2–3):336–348, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003578.

Scandol:2003:UCS

[Sca03] James P. Scandol. Use of cumulative sum (CUSUM) control charts of landed catch in the management of fisheries. Fisheries Research, 64(1):19–36, October 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001048.

Schweder:2001:PWD

[Sch01] Tore Schweder. Protecting whales by distorting uncertainty: non-precautionary mismanagement? Fisheries Research, 52(3):217–225, July 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002630.

Schmidt:2009:RVC

[Sch09] Marc Bodo Schmidt. Reactions of vendace (Coregonus albula, Linnaeus 1758) towards an approaching pelagic pair-trawl observed by split-beam echosounding. Fisheries Research, 96(1):95–101, February 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002865.

Santos:2006:LGR

[SCLM06] M. N. Santos, A. Canas, P. G. Lino, and C. C. Monteiro. Length–girth relationships for 30 marine fish species. Fisheries Research, 78(2–3):368–373, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000397.

Silva:2008:GVS

[SCM^{+}08] A. Silva, P. Carrera, J. Massé, A. Uriarte, M. B. Santos, P. B. Oliveira, E. Soares, C. Porteiro, and Y. Stratoudakis. Geographic variability of sardine growth across the northeastern Atlantic and the Mediterranean Sea. Fisheries Research, 90(1–3):56–69, April 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002330.

Steingrund:2009:HFA

[SCM09] Petur Steingrund, Dagunn H. Clementsen, and Rógvi Mouritsen. Higher food abundance reduces the catchability of cod (Gadus morhua) to longlines on the Faroe Plateau.

Fisheries Research, 100(3):230–239, November 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002124.

Santos:2001:AIC

[SCP01] M. B. Santos, M. R. Clarke, and G. J. Pierce. Assessing the importance of cephalopods in the diets of marine mammals and other top predators: problems and solutions. Fisheries Research, 52(1–2):121–139, June 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002363.

Salas:2007:CAM

[SCSC07] Silvia Salas, Ratana Chuenpagdee, Juan Carlos Seijo, and Anthony Charles. Challenges in the assessment and management of small-scale fisheries in Latin America and the Caribbean. Fisheries Research, 87(1):5–16, October 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001555.

Sandoval-Castellanos:2007:PGS

[SCUADJ07] Edson Sandoval-Castellanos, Manuel Uribe-Alcocer, and Píndaro Díaz-Jaimes. Population genetic structure of jumbo squid (Dosidicus gigas) evaluated by RAPD analysis. Fisheries Research, 83(1):113–118, January 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003365.

Scheirer:2004:CSA

[SCW04] Kevin Scheirer, Yong Chen, and Carl Wilson. A comparative study of American lobster fishery sea and port sampling programs in Maine: 1998–2000. Fisheries Research, 68(1–3):343–350, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360300345X.

Storbeck:2001:FSR

[SD01] Frank Storbeck and Berent Daan. Fish species recognition using computer vision and a neural network. Fisheries Research, 51(1):11–15, April 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360000254X.

Simpfendorfer:2000:SAR

[SDH00] Colin A. Simpfendorfer, Kevin Donohue, and Norm G. Hall. Stock assessment and risk analysis for the whiskery shark (Furgaleus macki (Whitley)) in south-western Australia. Fisheries Research, 47(1):1–17, June 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001090.

Sanchez:2004:CDG

[SDM04] P. Sánchez, M. Demestre, and P. Martín. Characterisation of the discards generated by bottom trawling in the northwestern Mediterranean. Fisheries Research, 67(1):71–80, March 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603002698.

Santana:2009:DAN

[SDNL09] Francisco Marcante Santana, Paulo Duarte-Neto, and Rosangela Lessa. Demographic analysis of the night shark (Carcharhinus signatus, Poey, 1868) in the equatorial Southwestern Atlantic Ocean. Fisheries Research, 100(3): 210–214, November 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783609002057.

Sala:2009:PIS

[SdPFAL09] Antonello Sala, Joana d’Arc Prat Farran, Josefina Antonijuan, and Alessandro Lucchetti. Performance and impact on the seabed of an existing- and an experimentalotterboard: Comparison between model testing and fullscale sea trials. Fisheries Research, 100(2):156–166, October 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609001908.

Steward:2009:UOM

[SDS09] Cheree A. Steward, Karen D. DeMaria, and Jonathan M. Shenker. Using otolith morphometrics to quickly and inexpensively predict age in the gray angelfish (Pomacanthus arcuatus). Fisheries Research, 99(2):123–129, August 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609001271.

Stergiou:2002:CFG

[SE02] Konstantinos I. Stergiou and Karim Erzini. Comparative fixed gear studies in the Cyclades (Aegean Sea): size selectivity of small-hook longlines and monofilament gill nets. Fisheries Research, 58(1):25–40, October 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003630.

Sasso:2006:SST

[SE06a] Christopher R. Sasso and Sheryan P. Epperly. Seasonal sea turtle mortality risk from forced submergence in bottom trawls. Fisheries Research, 81(1):86–88, October 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002190.

Stobberup:2006:AMS

[SE06b] Kim A. Stobberup and Karim Erzini. Assessing mackerel scad, Decapterus macarellus, in Cape Verde: Using a Bayesian approach to biomass dynamic modelling in a datalimited situation. Fisheries Research, 82(1–3):194–203, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360600230X.

Seijo:2007:CMM

[Sei07] J. C. Seijo. Considerations for management of metapopulations in small-scale fisheries of the Mesoamerican barrier reef ecosystem. Fisheries Research, 87(1):86–91, October 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607001567.

Sun:2002:AYS

[SEPY02] Chi-Lu Sun, Nelson M. Ehrhardt, Clay E. Porch, and Su-Zan Yeh. Analyses of yield and spawning stock biomass per recruit for the South Atlantic albacore (Thunnus alalunga). Fisheries Research, 56(2):193–204, May 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003204.

Saly:2009:ALM

[SET^{+}09] Péter Sály, Tibor Erős, Péter Takács, András Specziár, István Kiss, and Péter Bíró. Assemblage level monitoring of stream fishes: the relative efficiency of single-pass vs. double-pass electrofishing. Fisheries Research, 99(3): 226–233, August 15, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783609001775.

Stewart:2003:MSN

[SF03] John Stewart and Douglas J. Ferrell. Mesh selectivity in the New South Wales demersal trap fishery. Fisheries Research, 59(3):379–392, January 3, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000243.

Sarvas:2005:SLG

[SF05] Tuula H. Sarvas and Svein-Erik Fevolden. The scnDNA locus Pan I reveals concurrent presence of different populations of Atlantic cod (Gadus morhua L.) within a single fjord. Fisheries Research, 76(3):307–316, December 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002183.

Strelcheck:2003:OFS

[SFCK03] Andrew J. Strelcheck, Gary R. Fitzhugh, Felicia C. Coleman, and Christopher C. Koenig. Otolith–fish size relationship in juvenile gag (Mycteroperca microlepis) of the eastern Gulf of Mexico: a comparison of growth rates between laboratory and field populations. Fisheries Research, 60(2–3):255–265, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001716.

Salthaug:2001:SCC

[SG01a] A. Salthaug and O. R. Godø. Standardisation of commercial CPUE. Fisheries Research, 49(3):271–281, January 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600002046.

Swan:2001:RAE

[SG01b] Sarah C. Swan and John D. M. Gordon. A review of age estimation in macrourid fishes with new data on age validation of juveniles. Fisheries Research, 51(2–3):177–195, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601002442.

Silvestre:2004:PPE

[SG04] Geronimo T. Silvestre and Len R. Garces. Population parameters and exploitation rate of demersal fishes in Brunei Darussalam (1989–1990). Fisheries Research, 69(1):73–90, August 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604000852.

Schmidt:2006:ISD

[SG06] Marc Bodo Schmidt and Hubert Gassner. Influence of scuba divers on the avoidance reaction of a dense vendace (Coregonus albula L.) population monitored by hydroacoustics. Fisheries Research, 82(1–3):131–139, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002888.

Sobrino:2000:TGS

[SGB00] Ignacio Sobrino, Teresa Garcıá, and Jorge Baro. Trawl gear selectivity and the effect of mesh size on the deep-water rose shrimp (Parapenaeus longirostris, Lucas, 1846) fishery off the Gulf of Cádiz (SW Spain). Fisheries Research, 44(3):235–245, January 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000909.

Sistiaga:2008:SSP

[SGL08] M. Sistiaga, E. Grimaldo, and R. B. Larsen. Size selectivity patterns in the North-east Arctic cod and haddock fishery with sorting grids of 55, 60, 70 and 80 mm. Fisheries Research, 93(1–2):195–203, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001318.

Schramm:2002:CSF

[SGP02] Harold L. Schramm, Stephen C. Grado, and Lawrence L. Pugh. The costs of sampling fishes in riverine habitats of a large river. Fisheries Research, 56(1):51–57, April 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003162.

Shephard:2009:HRN

[SGRK09] Samuel Shephard, Clifford A. Goudey, Andrew Read, and Michel J. Kaiser. Hydrodredge: Reducing the negative impacts of scallop dredging. Fisheries Research, 95(2–3):206–209, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002646.

Silva:2002:DFC

[SGS02] Luis Silva, Juan Gil, and Ignacio Sobrino. Definition of fleet components in the Spanish artisanal fishery of the Gulf of Cádiz (SW Spain ICES division IXa). Fisheries Research, 59(1–2):117–128, December 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601004209.

Silberschneider:2009:AGM

[SGS09] V. Silberschneider, C. A. Gray, and J. Stewart. Age, growth, maturity and the overfishing of the iconic sciaenid, Argyrosomus japonicus, in south-eastern, Australia. Fisheries Research, 95(2–3):220–229, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002713.

Santos:2002:WLR

[SGVM02] M. N. Santos, M. B. Gaspar, P. Vasconcelos, and C. C. Monteiro. Weight–length relationships for 50 selected fish species of the Algarve coast (southern Portugal). Fisheries Research, 59(1–2):289–295, December 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601004015.

Steele:2003:FS

[SH03] John Steele and Porter Hoagland. Are fisheries “sustainable”? Fisheries Research, 64(1):1–3, October 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002224.

Steele:2004:RZR

[SH04] John Steele and Porter Hoagland. Reply to Zeller and Russ. Fisheries Research, 67(2):247–248, April 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000165.

Stewart:2009:BFC

[SH09] John Stewart and Julian M. Hughes. Biological and fishery characteristics of rubberlip morwong Nemadactylus douglasii (Hector, 1875) in eastern Australia. Fisheries Research, 96(2–3):267–274, March 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003925.

Subbey:2007:RFM

[SHBÅ07] Sam Subbey, Daniel Howell, Bjarte Bogstad, and Morten Åsnes. Reducing fisheries model complexity using a finite Fourier series reparameterization. Fisheries Research, 84 (3):390–394, May 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783607000069.

Simpfendorfer:2002:RFI

[SHBC02] Colin A. Simpfendorfer, Robert E. Hueter, Ulrika Bergman, and Stephen M. H. Connett. Results of a fisheryindependent survey for pelagic sharks in the western North Atlantic, 1977–1994. Fisheries Research, 55(1–3):175–192, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601002880.

Slotte:2004:AMP

[SHDO04] Aril Slotte, Kaare Hansen, John Dalen, and Egil Ona. Acoustic mapping of pelagic fish distribution and abundance in relation to a seismic shooting area off the Norwegian west coast. Fisheries Research, 67(2):143–150, April 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360300290X.

Shelton:2002:BRS

[She02] Richard G. J. Shelton. Book review: Stock Enhancement and Sea Ranching: Fishing News Books; Bari R. Howell, Erlend Mokness, Terje Svåsand (Eds.); Blackwell Science, Oxford, 1999, 606 pages, hardcover, ISBN 0-85238-246-4 (£79.50). Fisheries Research, 57(2):205–206, August 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000346.

Shepherd:2003:FEC

[She03] J. G. Shepherd. Fishing effort control: could it work under the common fisheries policy? Fisheries Research, 63(2):149–153, August 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001255.

Shelton:2009:ECS

[She09] Peter A. Shelton. Eco-certification of sustainably managed fisheries — redundancy or synergy? Fisheries Research, 100(3):185–190, November 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609002276.

Scharf:2009:DAS

[SHKB09] W. R. Scharf, L. Heermann, U. König, and J. Borcherding. Development of abundance and size structure of young-of-the-year perch populations using three methods. Fisheries Research, 96(1):77–87, February 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002841.

Sistiaga:2009:IPG

[SHL09] M. Sistiaga, B. Herrmann, and R. B. Larsen. Investigation of the paired-gear method in selectivity studies. Fisheries Research, 97(3):196–205, May 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000447.

Shono:2008:ATD

[Sho08] Hiroshi Shono. Application of the Tweedie distribution to zero-catch data in CPUE analysis. Fisheries Research, 93(1–2):154–162, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000945.

Shono:2009:RLE

[Sho09] Hiroshi Shono. In reply to “Letter to the editor: Improving comparisons between models for CPUE (by Peter K. Dunn)”. Fisheries Research, 97(1–2):150, April 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000204. See [Dun09].

Siddeek:2003:DBR

[Sid03] M. S. M. Siddeek. Determination of biological reference points for Bristol Bay red king crab. Fisheries Research, 65(1–3):427–451, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002595.

Sumpton:2005:EIT

[SJ05] Wayne Sumpton and Stuart Jackson. The effects of incidental trawl capture of juvenile snapper (Pagrus auratus) on yield of a sub-tropical line fishery in Australia: an assessment examining habitat preference and early life history characteristics. Fisheries Research, 71(3):335–347, March 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001535.

Sale:2003:DES

[SK03a] Peter F. Sale and Jacob P. Kritzer. Determining the extent and spatial scale of population connectivity: decapods and coral reef fishes compared. Fisheries Research, 65(1–3):153–172, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002418.

Stergiou:2003:LGR

[SK03b] Konstantinos I. Stergiou and Vasiliki S. Karpouzi. Length– girth relationships for several marine fishes. Fisheries Research, 60(1):161–168, January 30, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000772.

Shiell:2008:DOS

[SK08a] Glenn R. Shiell and Brenton Knott. Diurnal observations of sheltering behaviour in the coral reef sea cucumber Holothuria whitmaei. Fisheries Research, 91(1):112– 117, May 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607003608.

Stoner:2008:REB

[SK08b] Allan W. Stoner and Stephen M. Kaimmer. Reducing elasmobranch bycatch: Laboratory investigation of rare earth metal and magnetic deterrents with spiny dogfish and Pacific halibut. Fisheries Research, 92(2–3):162–168, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608000209.

Sudo:2008:PSC

[SKF08] Hiroyuki Sudo, Naoto Kajihara, and Tetsuo Fujii. Predation by the swimming crab Charybdis japonica and piscivorous fishes: a major mortality factor in hatcheryreared juvenile Japanese flounder Paralichthys olivaceus released in Mano Bay, Sado Island, Japan. Fisheries Research, 89(1):49–56, January 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002135.

Stransky:2005:RAV

[SKKP05] Christoph Stransky, Günter Kanisch, Alois Krüger, and Stefan Purkl. Radiometric age validation of golden redfish (Sebastes marinus) and deep-sea redfish (S. mentella) in the Northeast Atlantic. Fisheries Research, 74(1–3): 186–197, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783605000846.

Somarakis:2002:ADE

[SKMT02] Stylianos Somarakis, Constantin Koutsikopoulos, Athanassios Machias, and Nikolaos Tsimenides. Applying the daily egg production method (DEPM) to small stocks in highly heterogeneous seas. Fisheries Research, 55(1–3):193–204, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601002843.

Solmundsson:2003:SDS

[SKP03] Jon Solmundsson, Hjalti Karlsson, and Jonbjorn Palsson. Sexual differences in spawning behaviour and catchability of plaice (Pleuronectes platessa) west of Iceland. Fisheries Research, 61(1–3):57–71, March 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002126.

Skrobe:2008:DCP

[Skr08] Laura Skrobe. Dedication to Captain Phil Ruhle, Sr. Fisheries Research, 94(2):119, November 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360800310X.

Sauer:2000:TRS

[SLA00] W. H. H. Sauer, M. R. Lipinski, and C. J. Augustyn. Tag recapture studies of the chokka squid Loligo vulgaris reynaudii d’Orbigny, 1845 on inshore spawning grounds on the south-east coast of South Africa. Fisheries Research, 45(3):283–289, April 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001186.

Sala:2004:CPP

[SLB04] Antonello Sala, Alessandro Lucchetti, and Gabriele Buglioni. The change in physical properties of some nylon (PA) netting samples before and after use. Fisheries Research, 69(2):181–188, September 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360400133X.

Sala:2007:ITT

[SLB07] Antonello Sala, Alessandro Lucchetti, and Gabriele Buglioni. The influence of twine thickness on the size selectivity of polyamide codends in a Mediterranean bottom trawl. Fisheries Research, 83(2–3):192–203, February 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003407.

Serrano:2009:MPW

[SLE09] Ignacio Serrano, Stefan Larsson, and Lars-Ove Eriksson. Migration performance of wild and hatchery sea trout (Salmo trutta L.) smolts — implications for compensatory hatchery programs. Fisheries Research, 99(3):210–215, August 15, 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609001635.

Suuronen:2005:EMT

[SLJ05] Petri Suuronen, Esa Lehtonen, and Pekka Jounela. Escape mortality of trawl caught Baltic cod (Gadus morhua)

— the effect of water temperature, fish size and codend catch. Fisheries Research, 71(2):151–163, February 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001936.

Steer:2007:AFU

[SLJ07] Mike A. Steer, Matt T. Lloyd, and W. Bruce Jackson. Assessing the feasibility of using ‘by-product’ data as a pre-recruit index in South Australia’s southern calamary (Sepioteuthis australis) fishery. Fisheries Research, 88(1–3):42–50, December 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001671.

Sala:2008:SSD

[SLPF08] Antonello Sala, Alessandro Lucchetti, Corrado Piccinetti, and Mario Ferretti. Size selection by diamond- and squaremesh codends in multi-species Mediterranean demersal trawl fisheries. Fisheries Research, 93(1–2):8–21, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608000568.

Setala:2008:SIF

[SLV^{+}08] Jari Setälä, Jukka Laitinen, Jarno Virtanen, Kaija Saarni, Max Nielsen, and Asmo Honkanen. Spatial integration of freshwater fish markets in the Northern Baltic Sea area. Fisheries Research, 92(2–3):196–206, August 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000404.

Stephens:2004:MAS

[SM04a] Andi Stephens and Alec MacCall. A multispecies approach to subsetting logbook data for purposes of estimating CPUE. Fisheries Research, 70(2–3):299–310, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001730.

Stuart:2004:DAD

[SM04b] Ivor G. Stuart and Stephen McKillup. Dubious analyses or different rivers? A comment on the growth of barramundi. Fisheries Research, 68(1–3):373–374, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000633.

Sumpton:2008:IMF

[SMB^{+}08] Wayne Sumpton, David Mayer, Ian Brown, Bill Sawynok, Mark McLennan, Adam Butcher, and John Kirkwood. Investigation of movement and factors influencing postrelease survival of line-caught coral reef fish using recreational tag-recapture data. Fisheries Research, 92(2–3): 189–195, August 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608000398.

Solomon:2003:IAS

[SMD03] D. J. Solomon, G. W. Mawle, and W. Duncan. An integrated approach to salmonid management. Fisheries Research, 62(2):229–234, May 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001583.

Sanchez:2004:MCE

[SMD04] Pilar Sánchez, Francesc Maynou, and Montserrat Demestre. Modelling catch, effort and price in a juvenile Eledone cirrhosa fishery over a 10-year period. Fisheries Research, 68(1–3):319–327, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000232.

Stergiou:2002:GNL

[SME02] Konstantinos I. Stergiou, Dimitrios K. Moutopoulos, and Karim Erzini. Gill net and longlines fisheries in Cyclades waters (Aegean Sea): species composition and gear competition. Fisheries Research, 57(1):25–37, July 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003344.

Smith:2003:SMT

[Smi03] Kimberley A. Smith. A simple multivariate technique to improve the design of a sampling strategy for age-based fishery monitoring. Fisheries Research, 64(1):79–85, October 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603001802.

Stobutzki:2001:BDV

[SMJS01] Ilona C. Stobutzki, Margaret J. Miller, Peter Jones, and John P. Salini. Bycatch diversity and variation in a tropical Australian penaeid fishery; the implications for monitoring. Fisheries Research, 53(3):283–301, November 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002733.

Steffe:2007:SRF

[SMM07] Aldo S. Steffe, William G. Macbeth, and Jeff J. Murphy. Status of the recreational fisheries in two Australian coastal estuaries following large fish-kill events. Fisheries Research, 85(3):258–269, July 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000380.

Salini:2004:AMV

[SMRH04] J. P. Salini, D. A. Milton, M. J. Rahman, and M. G. Hussain. Allozyme and morphological variation throughout the geographic range of the tropical shad, Hilsa Tenualosa ilisha. Fisheries Research, 66(1):53–69, January 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001243.

Shepherd:2002:EEV

[SMS02] Gary R. Shepherd, Christopher W. Moore, and Richard J. Seagraves. The effect of escape vents on the capture of black sea bass, Centropristis striata, in fish traps. Fisheries Research, 54(2):195–207, January 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600003052.

Stergiou:2006:TNC

[SMS^{+}06] Konstantinos I. Stergiou, Dimitrios K. Moutopoulos, Milagrosa C. Soriguer, Esteban Puente, Pedro G. Lino, Cristina Zabala, Pedro Monteiro, Luis A. Errazkin, and Karim Erzini. Trammel net catch species composition, catch rates and métiers in southern European waters: a multivariate approach. Fisheries Research, 79(1–2):170–182, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001068.

Stergiou:2003:CWD

[SMSK03] Konstantinos I. Stergiou, Athanassios Machias, Stylianos Somarakis, and Argyris Kapantagakis. Can we define target species in Mediterranean trawl fisheries? Fisheries Research, 59(3):431–435, January 3, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000164.

Stransky:2008:OSA

[SMSZ08] Christoph Stransky, Alberto G. Murta, Jürgen Schlickeisen, and Christopher Zimmermann. Otolith shape analysis as a tool for stock separation of horse mackerel (Trachurus trachurus) in the Northeast Atlantic and Mediterranean. Fisheries Research, 89(2):159–166, February 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002433.

Shukhgalter:2001:PHJ

[SN01] Olga A. Shukhgalter and Chingis M. Nigmatullin. Parasitic helminths of jumbo squid Dosidicus gigas (Cephalopoda: Ommastrephidae) in open waters of the central east Pacific. Fisheries Research, 54(1):95–110, December 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003745.

Saito:2009:RSR

[SN09] Toshihiko Saito and Kazuya Nagasawa. Regional synchrony in return rates of chum salmon (Oncorhynchus keta) in Japan in relation to coastal temperature and size at release. Fisheries Research, 95(1):14–27, January 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002014.

Sarkar:2008:BPE

[SND^{+}08] U. K. Sarkar, R. S. Negi, P. K. Deepak, W. S. Lakra, and S. K. Paul. Biological parameters of the endangered fish Chitala chitala (Osteoglossiformes: Notopteridae) from some Indian rivers. Fisheries Research, 90(1–3):170–177, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607002706.

Stoner:2006:TAA

[SOH06] Allan W. Stoner, Michele L. Ottmar, and Thomas P. Hurst. Temperature affects activity and feeding motivation in Pacific halibut: Implications for bait-dependent fishing. Fisheries Research, 81(2–3):202–209, November 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002773.

Sumpton:2008:ESD

[SOKS08] Wayne D. Sumpton, Jennifer R. Ovenden, Clive P. Keenan, and Raewyn Street. Evidence for a stock discontinuity of snapper (Pagrus auratus) on the east coast of Australia. Fisheries Research, 94(1):92–98, October 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001999.

Solow:2002:FRN

[Sol02] Andrew R. Solow. Fisheries recruitment and the North Atlantic Oscillation. Fisheries Research, 54(2):295–297, January 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600003088.

Solomon:2003:F

[Sol03] David J. Solomon. Foreword. Fisheries Research, 62(2): 109, May 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360200156X.

Solomon:2005:BRS

[Sol05] David J. Solomon. Book review: Salmon at the Edge, D. Mills (Ed.). Blackwell Science, Oxford (2003), 307 pp., hard cover, £85, ISBN: 0-632-06457-9. Fisheries Research, 74(1–3):312–313, August 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000433.

Somerton:2003:BES

[Som03] David A. Somerton. Bridle efficiency of a survey trawl for flatfish: measuring the length of the bridles in contact with the bottom. Fisheries Research, 60(2–3):273–279, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602001789.

Somerton:2002:CCT

[SOS02] David A. Somerton, Robert S. Otto, and Stephen E. Syrjala. Can changes in tow duration on bottom trawl surveys lead to changes in CPUE and mean size? Fisheries Research, 55(1–3):63–70, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002934.

Symes:2009:WBS

[SP09] David Symes and Jeremy Phillipson. Whatever became of social objectives in fisheries policy? Fisheries Research, 95(1):1–5, January 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002270.

Sancho:2003:CRM

[SPB^{+}03] Gorka Sancho, Esteban Puente, Araitz Bilbao, Eva Gomez, and Luis Arregi. Catch rates of monkfish (Lophius spp.) by lost tangle nets in the Cantabrian Sea (northern Spain). Fisheries Research, 64(2–3):129–139, November 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002121.

Sepulveda:2007:OIB

[SPS^{+}07] M. Sepúlveda, M. J. Pérez, W. Sielfeld, D. Oliva, L. R. Durán, L. Rodríguez, V. Araos, and M. Buscaglia. Operational interaction between South American sea lions Otaria flavescens and artisanal (small-scale) fishing in Chile: Results from interview surveys and on-board observations. Fisheries Research, 83(2–3):332–340, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003717.

Secor:2000:OSU

[SR00] David H. Secor and Jay R. Rooker. Is otolith strontium a useful scalar of life cycles in estuarine fishes? Fisheries Research, 46(1–3):359–371, May 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001594.

Shephard:2006:SDO

[SR06] Samuel Shephard and Emer Rogan. Seasonal distribution of orange roughy (Hoplostethus atlanticus) on the Porcupine Bank west of Ireland. Fisheries Research, 77(1):17–23, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605002912.

Solomon:2007:RBW

[SR07] Farahnaz N. Solomon and Indar W. Ramnarine. Reproductive biology of white mullet, Mugil curema (Valenciennes) in the Southern Caribbean. Fisheries Research, 88(1–3):133–138, December 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001579.

Sarda:2002:CFT

[SRA^{+}02] F. Sardà, L. Recasens, P. Abelló, G. Rotllant, and B. Molí. Commercial feasibility trial for a singlewarp deep-water “Maireta” (OTMS) trawl gear. Fisheries Research, 55(1–3):121–130, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002958.

Smith:2002:MTD

[SRH^{+}02] Peter J. Smith, Simon G. Robertson, Peter L. Horn, Brian Bull, Owen F. Anderson, Basil R. Stanton, and Catherine S. Oke. Multiple techniques for determining stock relationships between orange roughy, Hoplostethus atlanticus, fisheries in the eastern Tasman Sea. Fisheries Research, 58(2):119–140, November 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003897.

Svane:2008:FCD

[SRS08] I. Svane, S. Roberts, and T. Saunders. Fate and consumption of discarded by-catch in the Spencer Gulf prawn fishery, South Australia. Fisheries Research, 90(1–3):158–169, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360700269X.

Stetter:2006:HSR

[SRTP06] S. L. Parker Stetter, L. G. Rudstam, J. L. Stritzel Thomson, and D. L. Parrish. Hydroacoustic separation of rainbow smelt (Osmerus mordax) age groups in Lake Champlain. Fisheries Research, 82(1–3):176–185, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002335.

Steele:2000:ESB

[SS00a] J. H. Steele and M. Schumacher. Ecosystem structure before fishing. Fisheries Research, 44(3):201–205, January 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600900031.

Steele:2000:RDC

[SS00b] John Steele and Mary Schumacher. A reply to Dr. Caddy. Fisheries Research, 47(1):102, June 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001910.

Sbrana:2003:AFA

[SSB03] Mario Sbrana, Paolo Sartor, and Paola Belcari. Analysis of the factors affecting crustacean trawl fishery catch rates in the northern Tyrrhenian Sea (western Mediterranean). Fisheries Research, 65(1–3):271–284, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002480.

Swain:2003:DVT

[SSC^{+}03] D. P. Swain, A. F. Sinclair, M. Castonguay, G. A. Chouinard, K. F. Drinkwater, L. P. Fanning, and D. S. Clark. Density-versus temperature-dependent growth of Atlantic cod (Gadus morhua) in the Gulf of St. Lawrence and on the Scotian Shelf. Fisheries Research, 59(3):327–341, January 3, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000279.

Stobutzki:2006:KIC

[SSG06] Ilona C. Stobutzki, Geronimo T. Silvestre, and Len R. Garces. Key issues in coastal fisheries in South and Southeast Asia, outcomes of a regional initiative. Fisheries Research, 78(2–3):109–118, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000567.

Santos:2003:CRN

[SSGM03a] M. N. Santos, H. Saldanha, M. B. Gaspar, and C. C. Monteiro. Causes and rates of net loss off the Algarve (southern Portugal). Fisheries Research, 64(2–3):115–118, November 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603002108.

Santos:2003:HMM

[SSGM03b] M. N. Santos, H. J. Saldanha, M. B. Gaspar, and C. C. Monteiro. Hake (Merluccius merluccius L., 1758) ghost fishing by gill nets off the Algarve (southern Portugal). Fisheries Research, 64(2–3):119–128, November 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360300211X.

Skaret:2006:PSH

[SSH^{+}06] Georg Skaret, Aril Slotte, Nils Olav Handegard, Bjørn Erik Axelsen, and Roar Jørgensen. Pre-spawning herring in a protected area showed only moderate reaction to a surveying vessel. Fisheries Research, 78(2–3):359–367, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002894.

Siira:2006:SAS

[SSIE06] Antti Siira, Petri Suuronen, Erkki Ikonen, and Jaakko Erkinaro. Survival of Atlantic salmon captured in and released from a commercial trap-net: Potential for selective harvesting of stocked salmon. Fisheries Research, 80(2–3):280–294, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001378.

Suuronen:2006:RSI

[SSK^{+}06] Petri Suuronen, Antti Siira, Tiina Kauppinen, Raimo Riikonen, Esa Lehtonen, and Hannu Harjunpää. Reduction of seal-induced catch and gear damage by modification of trap-net design: Design principles for a seal-safe trapnet. Fisheries Research, 79(1–2):129–138, June 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000701.

Stottrup:2002:URR

[SSML02] J. G. Støttrup, C. R. Sparrevohn, J. Modin, and K. Lehmann. The use of releases of reared fish to enhance natural populations: a case study on turbot Psetta maxima (Linné, 1758). Fisheries Research, 59(1–2):161– 180, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783601004131.

Saldanha:2003:UBA

[SSS^{+}03] H. J. Saldanha, G. Sancho, M. N. Santos, E. Puente, M. B. Gaspar, A. Bilbao, C. C. Monteiro, E. Gomez, and L. Arregi. The use of biofouling for ageing lost nets: a case study. Fisheries Research, 64(2–3):141–150, November 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002133.

Stobutzki:2006:DDC

[SST^{+}06] I. C. Stobutzki, G. T. Silvestre, A. Abu Talib, A. Krongprom, M. Supongpan, P. Khemakorn, N. Armada, and L. R. Garces. Decline of demersal coastal fisheries resources in three developing Asian countries. Fisheries Research, 78(2–3):130–142, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000580.

Smith:2008:EIH

[SST^{+}08] Anthony D. M. Smith, David C. Smith, Geoffrey N. Tuck, Neil Klaer, André E. Punt, Ian Knuckey, Jeremy Prince, Alexander Morison, Rudy Kloser, Malcolm Haddon, Sally Wayte, Jemery Day, Gavin Fay, Fred Pribac, Mike Fuller, Bruce Taylor, and L. Richard Little. Experience in implementing harvest strategies in Australia’s south-eastern fisheries. Fisheries Research, 94(3):373–379, December 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001835.

Schirripa:2000:EMG

[ST00] Michael J. Schirripa and Joel C. Trexler. Effects of mortality and gear selectivity on the fish otolith radius–total length relation. Fisheries Research, 46(1–3):83–89, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001351.

Smith:2003:FIT

[ST03] Stephen J. Smith and M. John Tremblay. Fisheryindependent trap surveys of lobsters (Homarus americanus): design considerations. Fisheries Research, 62 (1):65–75, April 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783602002515.

Shimose:2009:ADG

[STCH09] Tamaki Shimose, Toshiyuki Tanabe, Kuo-Shu Chen, and Chien-Chung Hsu. Age determination and growth of Pacific bluefin tuna, Thunnus orientalis, off Japan and Taiwan. Fisheries Research, 100(2):134–139, October 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001829.

Steele:2003:BRH

[Ste03a] John H. Steele. Book review: Hierarchical Perspectives on Marine Complexities: Searching for Systems in the Gulf of Maine: Spencer Apollonio (Ed.); Columbia University Press, New York, 2002, Hardback, ISBN 0-231-12488-0, US $59.50; Paperback, ISBN 0-231-12489-9, US $29.50. Fisheries Research, 64(1):94–95, October 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002200.

Stevens:2003:TAL

[Ste03b] Bradley G. Stevens. Timing of aggregation and larval release by Tanner crabs, Chionoecetes bairdi, in relation to tidal current patterns. Fisheries Research, 65(1–3):201–216, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603002443.

Stewart:2003:LTR

[Ste03c] John Stewart. Long-term recaptures of tagged scyllarid lobsters (Ibacus peronii) from the east coast of Australia. Fisheries Research, 63(2):261–264, August 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360300078X.

Steele:2005:BRL

[Ste05] J. H. Steele. Book review: Large Marine Ecosystems of the World: Trends in Exploitation, Protection and Research, G. Hempel, K. Sherman (Eds.). Elsevier B.V. (2003), 440 pp., EUR 105, US $105, Hardback, ISBN 0-444-51027-3. Fisheries Research, 76(1):162, October 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001736.

Steele:2006:TEM

[Ste06] John H. Steele. Are there eco-metrics for fisheries? Fisheries Research, 77(1):1–3, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360500278X.

Stevens:2007:WSR

[Ste07] J. D. Stevens. Whale shark (Rhincodon typus) biology and ecology: a review of the primary literature. Fisheries Research, 84(1):4–9, March 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003948.

Stewart:2008:CDR

[Ste08] John Stewart. Capture depth related mortality of discarded snapper (Pagrus auratus) and implications for management. Fisheries Research, 90(1–3):289–295, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003062.

Suzuki:2003:VAF

[STH03] Katsuya Suzuki, Tsutomu Takagi, and Tomonori Hiraishi. Video analysis of fish schooling behavior in finite space using a mathematical model. Fisheries Research, 60(1):3–10, January 30, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000814.

Satoh:2008:VIM

[STI08] Keisuke Satoh, Yosuke Tanaka, and Masayuki Iwahashi. Variations in the instantaneous mortality rate between larval patches of Pacific bluefin tuna Thunnus orientalis in the northwestern Pacific Ocean. Fisheries Research, 89(3):248–256, March 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002226.

Shimizu:2005:ANF

[STK^{+}05] Takashi Shimizu, Tsutomu Takagi, Holger Korte, Tomonori Hiraishi, and Katsutaro Yamamoto. Application of NaLA, a fishing net configuration and loading analysis system, to drift gill nets. Fisheries Research, 76(1):67–80, October 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605001633.

Sutcliffe:2002:BRB

[Sut02] David Sutcliffe. Book review: Biology of Freshwater Crayfish: David M. Holdich (Ed.); Blackwell Science, Oxford, 2002, xvii + 702 pages, ISBN 0-632-05431-X, £110.00 (hbk). Fisheries Research, 59(1–2):297–299, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602001303.

Sifner:2004:PSM

[ŠV04] Svjetlana Krstulović Šifner and Nedo Vrgoč. Population structure, maturation and reproduction of the European squid, Loligo vulgaris, in the Central Adriatic Sea. Fisheries Research, 69(2):239–249, September 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001316.

Svane:2005:ODS

[Sva05] Ib Svane. Occurrence of dolphins and seabirds and their consumption of by-catch during prawn trawling in Spencer Gulf, South Australia. Fisheries Research, 76(3):317–327, December 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605002171.

Svedang:2014:CFA

[Sve14] Henrik Svedäng. Comments to Froese et al. (2008): Size matters: Ne quid nimis. Fisheries Research, 149(??):74–75, January 2014. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783613002038.  See

[FSPWG08, FWG15].

Svedang:2015:SSH

[Sve15] Henrik Svedäng. On size selectivity and L_{opt} as a harvest strategy: Reply to Froese et al. (2014). Fisheries Research, 164(??):331–332, April 2015. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783614003646. See [FSPWG08, Sve14].

Stanley:2000:VDS

[SW00] David R. Stanley and Charles A. Wilson. Variation in the density and species composition of fishes associated with three petroleum platforms using dual beam hydroacoustics. Fisheries Research, 47(2–3):161–172, July 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001673.

Somerton:2001:AST

[SW01] David A. Somerton and Kenneth L. Weinberg. The affect of speed through the water on footrope contact of a survey trawl. Fisheries Research, 53(1):17–24, September 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002721.

Smedbol:2002:MTN

[SW02] R. Kent Smedbol and J. S. Wroblewski. Metapopulation theory and northern cod population structure: interdependency of subpopulations in recovery of a groundfish population. Fisheries Research, 55(1–3):161–174, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002892.

Seung:2009:MEL

[SW09] Chang K. Seung and Edward C. Waters. Measuring the economic linkage of Alaska fisheries: a supplydriven social accounting matrix (SDSAM) approach. Fisheries Research, 97(1–2):17–23, April 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003998.

Suda:2008:TSP

[SWA08] Maki Suda, Chikako Watanabe, and Tatsuro Akamine. Two-species population dynamics model for Japanese sardine Sardinops melanostictus and chub mackerel Scomber japonicus off the Pacific coast of Japan. Fisheries Research, 94(1):18–25, October 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608001902.

Swartzman:2005:RAP

[SWC^{+}05] Gordon Swartzman, Andreas Winter, Kenneth Coyle, Richard Brodeur, Troy Buckley, Lorenzo Ciannelli, George Hunt, James Ianelli, and Allen Macklin. Relationship of age-0 pollock abundance and distribution around the Pribilof Islands, to other shelf regions of the eastern Bering Sea. Fisheries Research, 74(1–3):273–287, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001062.

Siddeek:2009:PHC

[SWM09] M. S. M. Siddeek, E. Wade, and M. Moriyasu. Proposed harvest control rule for managing the snow crab stock in the southwestern Gulf of St. Lawrence, Canada. Fisheries Research, 95(2–3):268–279, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003159.

Sun:2005:SSY

[SWPY05] Chi-Lu Sun, Sheng-Ping Wang, Clay E. Porch, and Su-Zan Yeh. Sex-specific yield per recruit and spawning stock biomass per recruit for the swordfish, Xiphias gladius, in the waters around Taiwan. Fisheries Research, 71(1):61–69, January 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360400195X.

Shertzer:2009:SST

[SWT09] Kyle W. Shertzer, Erik H. Williams, and J. Christopher Taylor. Spatial structure and temporal patterns in a large marine ecosystem: Exploited reef fishes of the southeast United States. Fisheries Research, 100(2):126–133, October 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609001817.

Symes:2005:ACF

[Sym05a] David Symes. Altering course: future directions for Europe’s fisheries policy. Fisheries Research, 71(3):259–265, March 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604003091.

Symes:2005:FEM

[Sym05b] David Symes. Fishing, the environment and the media. Fisheries Research, 73(1–2):13–19, June 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000652.

Symes:2006:FGC

[Sym06] David Symes. Fisheries governance: a coming of age for fisheries social science? Fisheries Research, 81(2–3): 113–117, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606002360.

Symes:2007:BRF

[Sym07] David Symes. Book review: Fisheries Management — Progress Towards Sustainability, T. McClanahan, J. C.

Castilla (Eds.). Blackwell Publishing Ltd. (2007), 332 pp., Hardback, Price: £99.50, ISBN-13:978-1-4051-3932-8. Fisheries Research, 88(1–3):155–156, December 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001956.

Symes:2008:BRA

[Sym08a] David Symes. Book review: Advances in Fisheries Economics: Festschrift in Honour of Professor Gordon R. Munro, T. Bjørndal, D. V. Gordon, R. Arnason, U. R. Sumaila (Eds.). Blackwell Publishing Ltd. (2007), 308 pp., Hardback, Price: £99.50, ISBN-13:978-1-4051-4161-1. Fisheries Research, 90(1–3):316–317, April 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003207.

Symes:2008:BRFb

[Sym08b] David Symes. Book review: Fisheries Buybacks, R. Curtis, D. Squires (Eds.). Blackwell Publishing (2007), 267 pp., Hardback, Price: £89.50, ISBN-13: 978-0-8138-2546-5. Fisheries Research, 93(1–2):255, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001409.

Symes:2008:BRFa

[Sym08c] David Symes. Book review: Fishers’ Knowledge in Fisheries Science and Management, N. Haggan, B. Neis, I. G. Baird (Eds.), in: Coastal Management Sourcebooks, Vol. 4. Unesco Publishing (2007), 437 pp., Softback, Price: EUR 30, ISBN: 978-92-3-104029-0. Fisheries Research, 89(3):309–310, March 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002524.

Symes:2009:REU

[Sym09] David Symes. Reform of the European Union’s Common Fisheries Policy: Making fisheries management work. Fisheries Research, 100(2):99–102, October 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001787.

Su:2008:SCE

[SYS^{+}08] Nan-Jay Su, Su-Zan Yeh, Chi-Lu Sun, André E. Punt, Yong Chen, and Sheng-Ping Wang. Standardizing catch and effort data of the Taiwanese distant-water longline fishery in the western and central Pacific Ocean for bigeye tuna, Thunnus obesus. Fisheries Research, 90(1–3):235– 246, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607002962.

Tallack:2009:RGE

[Tal09] Shelly M. L. Tallack. Regional growth estimates of Atlantic cod, Gadus morhua: Applications of the maximum likelihood GROTAG model to tagging data in the Gulf of Maine (USA/Canada) region. Fisheries Research, 99(3):137–150, August 15, 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001404.

Tian:2003:VAP

[TAS03] Yongjun Tian, Tatsuro Akamine, and Maki Suda. Variations in the abundance of Pacific saury (Cololabis saira) from the northwestern Pacific in relation to oceanic-climate changes. Fisheries Research, 60(2–3):439–454, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602001431.

Tosunoglu:2009:SDH

[TASF09] Zafer Tosunoğlu, Celalettin Aydın, Alp Salman, and Paulo Fonseca. Selectivity of diamond, hexagonal and square mesh codends for three commercial cephalopods in the Mediterranean. Fisheries Research, 97(1–2):95–102, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609000216.

Tovar-Avila:2009:USM

[TÁTWD09] J. Tovar-Ávila, V. S. Troynikov, T. I. Walker, and R. W. Day. Use of stochastic models to estimate the growth of the Port Jackson shark, Heterodontus portusjacksoni, off eastern Victoria, Australia. Fisheries Research, 95(2–3):230–235, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003068.

Taylor:2007:RFF

[Tay07] J. Geoff Taylor. Ram filter-feeding and nocturnal feeding of whale sharks (Rhincodon typus) at Ningaloo Reef, Western Australia. Fisheries Research, 84(1):65–70, March 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004012.

Tewfik:2004:BGN

[TB04] Alexander Tewfik and Christophe Béné. “The Big Grab”: non-compliance with regulations, skewed fishing effort allocation and implications for a spiny lobster fishery. Fisheries Research, 69(1):21–33, August 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000864.

Troadec:2000:UDT

[TBRL00] H. Troadec, A. Benzinou, V. Rodin, and J. Le Bihan. Use of deformable template for two-dimensional growth ring detection of otoliths by digital image processing: Application to plaice (Pleuronectes platessa) otoliths. Fisheries Research, 46(1–3):155–163, May 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001429.

Teixeira:2009:TSA

[TC09] Célia M. Teixeira and Henrique N. Cabral. Time series analysis of flatfish landings in the Portuguese coast. Fisheries Research, 96(2–3):252–258, March 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003895.

Tian:2009:EHS

[TCC^{+}09] Siquan Tian, Xinjun Chen, Yong Chen, Liuxiong Xu, and Xiaojie Dai. Evaluating habitat suitability indices derived from CPUE and fishing effort data for Ommatrephes bratramii in the northwestern Pacific Ocean. Fisheries Research, 95(2–3):181–188, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002592.

Tamini:2006:BCT

[TCPC06] L. L. Tamini, G. E. Chiaramonte, J. E. Perez, and H. L. Cappozzo. Batoids in a coastal trawl fishery of Argentina. Fisheries Research, 77(3):326–332, March 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002997.

Teh:2005:RFP

[TCS05] Louise Teh, Annadel S. Cabanban, and U. Rashid Sumaila. The reef fisheries of Pulau Banggi, Sabah: a preliminary profile and assessment of ecological and socio-economic sustainability. Fisheries Research, 76(3):359–367, December 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605002146.

Tzeng:2001:MVR

[TCY01] Tzong-Der Tzeng, Chin-Shiang Chiu, and Shean-Ya Yeh. Morphometric variation in red-spot prawn (Metapenaeopsis barbata) in different geographic waters off Taiwan. Fisheries Research, 53(3):211–217, November 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002861.

Tzanatos:2005:CTD

[TDK^{+}05] Evangelos Tzanatos, Evangelos Dimitriou, George Katselis, Michalis Georgiadis, and Constantin Koutsikopoulos. Composition, temporal dynamics and regional characteristics of small-scale fisheries in Greece. Fisheries Research, 73(1–2):147–158, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604003108.

Tuck:2003:ETR

[TdlMH^{+}03] G. N. Tuck, W. K. de la Mare, W. S. Hearn, R. Williams, A. D. M. Smith, X. He, and A. Constable. An exact time of release and recapture stock assessment model with an application to Macquarie Island Patagonian toothfish (Dissostichus eleginoides). Fisheries Research, 63(2):179–191, August 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783603000730.

Thomas:2003:CEG

[TEG03] Saly N. Thomas, Leela Edwin, and V. C. George. Catching efficiency of gill nets and trammel nets for penaeid prawns. Fisheries Research, 60(1):141–150, January 30, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000577.

Turkmen:2002:RTA

[TEYA02] Mustafa Türkmen, Orhan Erdoğan, Ayhan Yıldırım, and İhsan Akyurt. Reproduction tactics, age and growth of Capoeta capoeta umbla Heckel 1843 from the Aşkale Region of the Karasu River, Turkey. Fisheries Research, 54(3):317–328, February 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002661.

Thorpe:2009:BMA

[TF09] Teresa Thorpe and Dargan Frierson. Bycatch mitigation assessment for sharks caught in coastal anchored gillnets. Fisheries Research, 98(1–3):102–112, June 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000964.

Tully:2003:CF

[TFA03] Oliver Tully, Juan Freire, and Julian Addison. Crustacean fisheries. Fisheries Research, 65(1–3):1–2, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002315.

Tonks:2008:SCT

[TGH^{+}08] M. L. Tonks, S. P. Griffiths, D. S. Heales, D. T. Brewer, and Q. Dell. Species composition and temporal variation of prawn trawl bycatch in the Joseph Bonaparte Gulf, northwestern Australia. Fisheries Research, 89(3):276–293, March 2008. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607002263.

Tutman:2000:ISD

[TGS^{+}00] P. Tutman, B. Glamuzina, B. Skaramuca, V. Kožul, N. Glavić, and D. Lučić. Incidence of spinal deformities in natural populations of sandsmelt, Atherina boyeri (Risso, 1810) in the Neretva river estuary, middle Adriatic. Fisheries Research, 45(1):61–64, February 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699000983.

Temming:2003:GEC

[TH03] A. Temming and J.-P. Herrmann. Gastric evacuation in cod: Prey-specific evacuation rates for use in North Sea, Baltic Sea and Barents Sea multi-species models. Fisheries Research, 63(1):21–41, July 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000419.

Thomas:2006:ENS

[TH06] Saly N. Thomas and C. Hridayanathan. The effect of natural sunlight on the strength of polyamide 6 multifilament and monofilament fishing net materials. Fisheries Research, 81(2–3):326–330, November 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002359.

terHofstede:2006:ISA

[tHDC06] Remment ter Hofstede and Mark Dickey-Collas. An investigation of seasonal and annual catches and discards of the Dutch pelagic freezer-trawlers in Mauritania, Northwest Africa. Fisheries Research, 77(2):184–191, February 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360500295X.

Therkildsen:2007:SVL

[The07] Nina O. Therkildsen. Small- versus large-scale fishing operations in New England, USA. Fisheries Research, 83(2–3): 285–296, February 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606003651.

Thorstad:2004:ECR

[THN^{+}04] Eva B. Thorstad, Clinton J. Hay, Tor F. Næsje, Ben Chanda, and Finn Økland. Effects of catch-and-release angling on large cichlids in the subtropical Zambezi River. Fisheries Research, 69(1):141–144, August 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000876.

Tian:2009:IIV

[Tia09] Yongjun Tian. Interannual–interdecadal variations of spear squid Loligo bleekeri abundance in the southwestern Japan Sea during 1975–2006: Impact of the trawl fishing and recommendations for management under the different climate regimes. Fisheries Research, 100(1):78–85, September 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001593.

Thorarinsdottir:2005:FBB

[TJ05] Gudrun G. Thorarinsdóttir and Larry D. Jacobson. Fishery biology and biological reference points for management of ocean quahogs (Arctica islandica) off Iceland. Fisheries Research, 75(1–3):97–106, September 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360500130X.

Thomas:2000:NPA

[TK00] G. L. Thomas and Jay Kirsch. Nekton and plankton acoustics: an overview. Fisheries Research, 47(2–3):107–113, July 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001624.

Tuser:2009:FOA

[TKFJ09] Michal Tušer, Jan Kubečka, Jaroslava Frouzová, and Oldřich Jarolím. Fish orientation along the longitudinal profile of the Římov reservoir during daytime: Consequences for horizontal acoustic surveys. Fisheries Research, 96(1):23–29, February 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002774.

Tschernij:2003:GFL

[TL03] Vesa Tschernij and P.-O. Larsson. Ghost fishing by lost cod gill nets in the Baltic Sea. Fisheries Research, 64(2–3):151–162, November 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002145.

Tracey:2006:AEF

[TLD06] Sean R. Tracey, Jeremy M. Lyle, and Guy Duhamel. Application of elliptical Fourier analysis of otolith form as a tool for stock identification. Fisheries Research, 77(2):138–147, February 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360500305X.

Tracey:2007:RBP

[TLH07] Sean R. Tracey, Jeremy M. Lyle, and Malcolm Haddon. Reproductive biology and per-recruit analyses of striped trumpeter (Latris lineata) from Tasmania, Australia: Implications for management. Fisheries Research, 84(3):358– 367, May 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606004255.

Torres:2000:VSA

[TLMN00] Gabriel J. Torres, Antoni Lombarte, and Beatriz Morales-Nin. Variability of the sulcus acusticus in the sagittal otolith of the genus Merluccius (Merlucciidae). Fisheries Research, 46(1–3):5–13, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001284.

Torisawa:2006:TCB

[TMK^{+}06] Shinsuke Torisawa, Kazushi Miyashita, Ryo Kawabe, Yasuzumi Fujimori, Tatsuki Oshima, Satoshi Honda, and Katsufumi Sato. A technique for calculating bearing and tilt angles of walleye pollock photographed in trawls with digital still-picture loggers. Fisheries Research, 77(1):4–9, January 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783605002985.

Tsehaye:2007:RSC

[TMN07] Iyob Tsehaye, Marcel A. M. Machiels, and Leopold A. J. Nagelkerke. Rapid shifts in catch composition in the artisanal Red Sea reef fisheries of Eritrea. Fisheries Research, 86(1):58–68, August 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360700094X.

Tserpes:2006:SSF

[TMP^{+}06] George Tserpes, Dimitrios K. Moutopoulos, Panagiota Peristeraki, George Katselis, and Constantin Koutsikopoulos. Study of swordfish fishing dynamics in the eastern Mediterranean by means of machine-learning approaches. Fisheries Research, 78(2–3):196–202, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000373.

Thorstad:2003:EHR

[TNFF03] Eva B. Thorstad, Tor F. Næsje, Peder Fiske, and Bengt Finstad. Effects of hook and release on Atlantic salmon in the River Alta, northern Norway. Fisheries Research, 60(2–3):293–307, February 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001765.

Thorstad:2007:LTE

[TNL07] Eva B. Thorstad, Tor F. Næsje, and Ivar Leinan. Longterm effects of catch-and-release angling on ascending Atlantic salmon during different stages of spawning migration. Fisheries Research, 85(3):316–320, July 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607000549.

Tuya:2006:EFP

[TOBSJH06] F. Tuya, L. Ortega-Borges, P. Sanchez-Jerez, and R. J. Haroun. Effect of fishing pressure on the spatio-temporal variability of the parrotfish, Sparisoma cretense (Pisces: Scaridae), across the Canarian Archipelago (eastern Atlantic). Fisheries Research, 77(1):24–33, January 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002845.

Takasuka:2005:TIR

[TOK^{+}05] Akinori Takasuka, Yoshioki Oozeki, Hiroshi Kubota, Yoshinari Tsuruta, and Tetsuichiro Funamoto. Temperature impacts on reproductive parameters for Japanese anchovy: Comparison between inshore and offshore waters. Fisheries Research, 76(3):475–482, December 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002080.

Turan:2006:MMV

[TOÖD06] Cemal Turan, Mustafa Oral, Bayram Öztürk, and Ertuğ Düzgüneş. Morphometric and meristic variation between stocks of bluefish (Pomatomus saltatrix) in the Black, Marmara, Aegean and northeastern Mediterranean Seas. Fisheries Research, 79(1–2):139–147, June 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000713.

Tokac:2004:EPP

[TÖT04] Adnan Tokaç, Hüseyin Özbilgin, and Zafer Tosunoğlu. Effect of PA and PE material on codend selectivity in Turkish bottom trawl. Fisheries Research, 67(3):317–327, May 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603003023.

Tomas:2000:OME

[TP00] Javier Tomás and Jacques Panfili. Otolith microstructure examination and growth patterns of Vinciguerria nimbaria (Photichthyidae) in the tropical Atlantic Ocean. Fisheries Research, 46(1–3):131–145, May 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001405.

Tesfamichael:2006:MES

[TP06] D. Tesfamichael and T. J. Pitcher. Multidisciplinary evaluation of the sustainability of Red Sea fisheries using Rapfish. Fisheries Research, 78(2–3):227–235, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000312.

Tingley:2005:FAT

[TPC05] Diana Tingley, Sean Pascoe, and Louisa Coglan. Factors affecting technical efficiency in fisheries: stochastic production frontier versus data envelopment analysis approaches. Fisheries Research, 73(3):363–376, July 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000251.

Tingley:2003:ECU

[TPM03] Diana Tingley, Sean Pascoe, and Simon Mardle. Estimating capacity utilisation in multi-purpose, multi-métier fisheries. Fisheries Research, 63(1):121–134, July 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002837.

Trevorrow:2001:ESS

[Tre01] Mark V. Trevorrow. An evaluation of a steerable sidescan sonar for surveys of near-surface fish. Fisheries Research, 50(3):221–234, March 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002149.

Trisak:2005:AGT

[Tri05] Jiraporn Trisak. Applying game theory to analyze the influence of biological characteristics on fishers’ cooperation in fisheries co-management. Fisheries Research, 75(1–3):164–174, September 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001360.

Tringali:2006:BAG

[Tri06] Michael D. Tringali. A Bayesian approach for the genetic tracking of cultured and released individuals. Fisheries Research, 77(2):159–172, February 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003012.

Tuset:2006:SOS

[TRL06] Víctor M. Tuset, Paul L. Rosin, and Antonio Lombarte. Sagittal otolith shape used in the identification of fishes of the genus Serranus. Fisheries Research, 81(2–3):316–325, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606002372.

Tae:2006:AAV

[TS06] Jong-Wan Tae and Hyeon-Ok Shin. Acoustic analysis of volume variation in a bag-net within a set-net. Fisheries Research, 80(2–3):263–269, September 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600141X.

Thorson:2009:GSS

[TS09] James T. Thorson and Colin A. Simpfendorfer. Gear selectivity and sample size effects on growth curve selection in shark age and growth studies. Fisheries Research, 98(1–3):75–84, June 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000915.

Tschernij:2004:MAA

[TSJ04] Vesa Tschernij, Petri Suuronen, and Pekka Jounela. A modelling approach for assessing short-term catch losses as a consequence of a mesh size increase. Fisheries Research, 69(3):399–406, October 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001432.

Tirado:2003:RVV

[TSM03] C. Tirado, C. Salas, and I. Márquez. Reproduction of Venus verrucosa L., 1758 (Bivalvia: Veneridae) in the littoral of Málaga (southern Spain). Fisheries Research, 63(3):437–445, September 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001061.

Trigueros-Salmeron:2001:SSV

[TSOG01] Jose Angel Trigueros-Salmeron and Sofia Ortega-Garcia. Spatial and seasonal variation of relative abundance of the skipjack tuna Katsuwonus pelamis (Linnaeus, 1758) in the Eastern Pacific Ocean (EPO) during 1970–1995. Fisheries Research, 49(3):227–232, January 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002125.

Tidu:2004:MRE

[TSP^{+}04] C. Tidu, R. Sardá, M. Pinna, A. Cannas, M. F. Meloni, E. Lecca, and R. Savarino. Morphometric relationships of the European spiny lobster Palinurus elephas from northwestern Sardinia. Fisheries Research, 69(3):371–379, October 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604001249.

Takagi:2004:VLN

[TSS^{+}04] Tsutomu Takagi, Takashi Shimizu, Katsuya Suzuki, Tomonori Hiraishi, and Katsutaro Yamamoto. Validity and layout of “NaLA”: a net configuration and loading analysis system. Fisheries Research, 66(2–3):235–243, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603002042.

Tzanatos:2006:ICS

[TSTK06] Evangelos Tzanatos, Stylianos Somarakis, George Tserpes, and Constantin Koutsikopoulos. Identifying and classifying small-scale fisheries métiers in the Mediterranean: a case study in the Patraikos Gulf, Greece. Fisheries Research, 81(2–3):158–168, November 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002815.

Tzanatos:2008:CLA

[TSTK08] Evangelos Tzanatos, Stylianos Somarakis, George Tserpes, and Constantin Koutsikopoulos. Catch length analysis, relation to minimum landing sizes and management implications from a Mediterranean small-scale fishery (Patraikos Gulf, Greece). Fisheries Research, 93(1–2):125– 134, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608000891.

Tserpes:2001:AGM

[TT01] G. Tserpes and N. Tsimenides. Age, growth and mortality of Serranus cabrilla (Linnaeus, 1758) on the Cretan shelf. Fisheries Research, 51(1):27–34, April 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360000237X.

Tserpes:2009:BEE

[TTP^{+}09] George Tserpes, Evangelos Tzanatos, Panagiota Peristeraki, Vincenzo Placenti, and Laurence Kell. A bioeconomic evaluation of different management measures for the Mediterranean swordfish. Fisheries Research, 96(2– 3):160–166, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783608003378.

Toscas:2009:STM

[TVB^{+}09] Peter J. Toscas, David J. Vance, Charis Y. Burridge, Cathy M. Dichmont, Shijie Zhou, William N. Venables, Robert C. Pendrey, and Anthea Donovan. Spatio-temporal modelling of prawns in Albatross Bay, Karumba and Mornington Island. Fisheries Research, 96(2–3):173–187, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608003391.

Tafur:2001:DMS

[TVY01] Ricardo Tafur, Piero Villegas, and Miguel Rabıánd Carmen Yamashiro. Dynamics of maturation, seasonality of reproduction and spawning grounds of the jumbo squid Dosidicus gigas (Cephalopoda: Ommastrephidae) in Peruvian waters. Fisheries Research, 54(1):33–50, December 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003794.

Tzeng:2002:MAC

[TY02] Tzong-Der Tzeng and Shean-Ya Yeh. Multivariate allometric comparisons for kuruma shrimp (Penaeus japonicus) off Taiwan. Fisheries Research, 59(1–2):279–288, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601004039.

Taipe:2001:DCJ

[TYM^{+}01] Anatolio Taipe, Carmen Yamashiro, Luis Mariategui, Pablo Rojas, and Cecilia Roque. Distribution and concentrations of jumbo flying squid (Dosidicus gigas) off the Peruvian coast between 1991 and 1999. Fisheries Research, 54(1):21–32, December 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003770.

Tzeng:2004:MVB

[Tze04] Tzong-Der Tzeng. Morphological variation between populations of spotted mackerel (Scomber australasicus) off Taiwan. Fisheries Research, 68(1–3):45–55, July 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000700.

Ungfors:2007:MAE

[UHN07] Anette Ungfors, Hans Hallbäck, and Per G. Nilsson. Movement of adult edible crab (Cancer pagurus L.) at the Swedish West Coast by mark-recapture and acoustic tracking. Fisheries Research, 84(3):345–357, May 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004243.

Uki:2006:SEJ

[Uki06] Nagashisa Uki. Stock enhancement of the Japanese scallop Patinopecten yessoensis in Hokkaido. Fisheries Research, 80(1):62–66, August 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001214.

Uriarte:2001:MAM

[UL01] Andrés Uriarte and Paulino Lucio. Migration of adult mackerel along the Atlantic European shelf edge from a tagging experiment in the south of the Bay of Biscay in 1994. Fisheries Research, 50(1–2):129–139, February 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002460.

Ulrich:2002:MSM

[ULDG02] Clara Ulrich, Bertrand Le Gallic, Matthew R. Dunn, and Didier Gascuel. A multi-species multi-fleet bioeconomic simulation model for the English Channel artisanal fisheries. Fisheries Research, 58(3):379–401, November 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783601003939.

Ulrich:2002:SSB

[UM02] Clara Ulrich and Paul Marchal. Sensitivity of some biological reference points to shifts in exploitation patterns and inputs uncertainty for three North Sea demersal stocks. Fisheries Research, 58(2):153–169, November 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003812.

Ueda:2001:WBP

[UMK01] Yuji Ueda, Takashi Matsuishi, and Yasuji Kanno. Weightbased population analysis: an estimation method. Fisheries Research, 50(3):271–278, March 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002162.

Urpanen:2009:PSE

[UMV^{+}09] Olli Urpanen, Timo J. Marjomäki, Markku Viljanen, Hannu Huuskonen, and Juha Karjalainen. Population size estimation of larval coregonids in large lakes: Stratified sampling design with a simple prediction model for vertical distribution. Fisheries Research, 96(1):109–117, February 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608002889.

Ueno:2004:DOM

[USKK04] Y. Ueno, S. Suyama, Y. Kurita, and T. Kumazawa. Design and operation methods of a mid-water trawl for quantitative sampling of a surface pelagic fish, Pacific saury (Cololabis saira). Fisheries Research, 66(1):3–17, January 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001784.

Uusitalo:2005:SSA

[UVKS05] Laura Uusitalo, Kimmo Vehkalahti, Sakari Kuikka, and Pirkko Söderkultalahti. Studying species associations from commercial catch data: a Baltic Sea application. Fisheries Research, 72(2–3):301–310, May 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002541.

Villamor:2004:GVM

[VAF04] Begoña Villamor, Pablo Abaunza, and A. Celso Fariña. Growth variability of mackerel (Scomber scombrus) off north and northwest Spain and a comparative review of the growth patterns in the northeast Atlantic. Fisheries Research, 69(1):107–121, August 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000682.

VanMuiswinkel:2001:BRC

[Van01] W. B. Van Muiswinkel. Book review: Carp: Biology and Culture: R. Billard (Ed.); Springer–Praxis Series in Aquaculture and Fisheries, Praxis Publishers Ltd., Chichester, UK, 1999, 342 pages, ISBN 1-85233-118-6 (£55). Fisheries Research, 51(1):95–97, April 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360000312X.

Vasconcellos:2003:AHS

[Vas03] Marcelo Vasconcellos. An analysis of harvest strategies and information needs in the purse seine fishery for the Brazilian sardine. Fisheries Research, 59(3):363–378, January 3, 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602000267.

VanderHaegen:2004:SSC

[VAYD04] G. E. Vander Haegen, C. E. Ashbrook, K. W. Yi, and J. F. Dixon. Survival of spring Chinook salmon captured and released in a selective commercial fishery using gill nets and tangle nets. Fisheries Research, 68(1–3):123–133, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604000645.

Vitale:2008:SDC

[VBSC08] F. Vitale, P. Börjesson, H. Svedäng, and M. Casini. The spatial distribution of cod (Gadus morhua L.) spawning grounds in the Kattegat, eastern North Sea. Fisheries Research, 90(1–3):36–44, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002317.

Volk:2000:OCR

[VBSK00] Eric C. Volk, Ann Blakley, Steven L. Schroder, and Scott. M. Kuehner. Otolith chemistry reflects migratory characteristics of Pacific salmonids: Using otolith core chemistry to distinguish maternal associations with sea and freshwaters. Fisheries Research, 46(1–3):251–266, May 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600001508.

Venables:2004:GLM

[VD04a] W. N. Venables and C. M. Dichmont. A generalised linear model for catch allocation: an example from Australia’s Northern Prawn Fishery. Fisheries Research, 70(2–3):409–426, December 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001821.

Venables:2004:GGG

[VD04b] W. N. Venables and C. M. Dichmont. GLMs, GAMs and GLMMs: an overview of theory for applications in fisheries research. Fisheries Research, 70(2–3):319–337, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001754.

Vianna:2006:ECR

[VD06] Marcelo Vianna and Fernando D’Incao. Evaluation of by-catch reduction devices for use in the artisanal pink shrimp (Farfantepenaeus paulensis) fishery in Patos Lagoon, Brazil. Fisheries Research, 81(2–3):331–336, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606002347.

Volpedo:2003:EPS

[VE03] Alejandra Volpedo and Dinorah Diana Echeverría. Ecomorphological patterns of the sagitta in fish on the continental shelf off Argentine. Fisheries Research, 60(2–3):551–560, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783602001704.

Velazquez:2003:RSS

[Vel03] Armando Vega Velázquez. Reproductive strategies of the spiny lobster Panulirus interruptus related to the marine environmental variability off central Baja California, Mexico: management implications. Fisheries Research, 65(1–3):123–135, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360300239X.

Vinagre:2006:HSI

[VFCC06] Catarina Vinagre, Vanessa Fonseca, Henrique Cabral, and Maria José Costa. Habitat suitability index models for the juvenile soles, Solea solea and Solea senegalensis, in the Tagus estuary: Defining variables for species management. Fisheries Research, 82(1–3):140–149, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002839.

Vasconcellos:2001:FCC

[VG01a] Marcelo Vasconcellos and Maria A. Gasalla. Fisheries catches and the carrying capacity of marine ecosystems in southern Brazil. Fisheries Research, 50(3):279–295, March 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600002174.

Vecchione:2001:CSC

[VG01b] Michael Vecchione and John Galbraith. Cephalopod species collected by deepwater exploratory fishing off New England. Fisheries Research, 51(2–3):385–391, May 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002624.

Voges:2002:EPD

[VGBF02] Elizabeth Voges, Ana Gordoa, Chris Hans Bartholomae, and John G. Field. Estimating the probability of different levels of recruitment for Cape hakes Merluccius capensis off Namibia, using environmental indices. Fisheries Research, 58(3):333–340, November 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601004003.

Vasconcellos:2006:SWC

[VH06] Marcelo Vasconcellos and Manuel Haimovici. Status of white croaker Micropogonias furnieri exploited in southern Brazil according to alternative hypotheses of stock discreetness. Fisheries Research, 80(2–3):196–202, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360600169X.

vanHerwerden:2009:CPG

[vHAN^{+}09] Lynne van Herwerden, William J. Aspden, Stephen J. Newman, Graham G. Pegg, Leica Briskey, and William Sinclair. A comparison of the population genetics of Lethrinus miniatus and Lutjanus sebae from the east and west coasts of Australia: Evidence for panmixia and isolation. Fisheries Research, 100(2):148–155, October 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609001842.

vanHerwerden:2006:DAM

[vHMAO^{+}06] Lynne van Herwerden, Jenny McIlwain, Hamed Al-Oufi, Wahida Al-Amry, and Alejandro Reyes. Development and application of microsatellite markers for Scomberomorus commerson (Perciformes; Teleostei) to a population genetic study of Arabian Peninsula stocks. Fisheries Research, 79(3):258–266, July 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001457.

Villegas:2001:GLC

[Vil01] Piero Villegas. Growth, life cycle and fishery biology of Loligo gahi (d’Orbigny, 1835) off the Peruvian coast. Fisheries Research, 54(1):123–131, December 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003769.

Vasek:2009:DVG

[VKČ^{+}09] M. Vašek, J. Kubečka, M. Čech, V. Draštík, J. Matěna, T. Mrkvička, J. Peterka, and M. Prchalová. Diel variation in gillnet catches and vertical distribution of pelagic fishes in a stratified European reservoir. Fisheries Research, 96(1):64–69, February 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002828.

Voss:2003:CFH

[VKD03] Rüdiger Voss, Friedrich W. Köster, and Miriam Dickmann. Comparing the feeding habits of co-occurring sprat (Sprattus sprattus) and cod (Gadus morhua) larvae in the Bornholm Basin, Baltic Sea. Fisheries Research, 63(1):97–111, July 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602002825.

Valero:2008:ESD

[VL08] Juan Luis Valero and Mario Lasta. Estimating survival of discarded scallops in the Patagonian scallop fishery: Comment on “Survival of Patagonian scallop (Zygochlamys patagonica, King and Broderip, 1832) after the size selection process on commercial fishing vessels”, by Bremec et al. 2004. Fisheries Research, 90(1–3):313–315, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002585.

Vega:2009:EAS

[VL09] Rodrigo Vega and Roberto Licandeo. The effect of American and Spanish longline systems on target and nontarget species in the eastern South Pacific swordfish fishery. Fisheries Research, 98(1–3):22–32, June 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000757.

Vigliano:2000:HFS

[VLD^{+}00] P. H. Vigliano, G. Lippolt, A. Denegri, M. Alonso, P. Macchi, and C. O. Dye. The human factors of the sport and recreational fishery of San Carlos de Bariloche, Rio Negro, Argentina. Fisheries Research, 49(2):141–153, December 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600002009.

vanMarlen:2003:ISB

[vM03] B. van Marlen. Improving the selectivity of beam trawls in The Netherlands: the effect of large mesh top panels on the catch rates of sole, plaice, cod and whiting. Fisheries Research, 63(2):155–168, August 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000754.

vanMarlen:2005:NAR

[vMBGF05] B. van Marlen, M. J. N. Bergman, S. Groenewold, and M. Fonds. New approaches to the reduction of non-target mortality in beam trawling. Fisheries Research, 72(2– 3):333–345, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604003005.

Vincent:2007:CSS

[VMPP07] Amanda C. J. Vincent, Jessica J. Meeuwig, Marivic G. Pajaro, and Nelson C. Perante. Characterizing a smallscale, data-poor, artisanal fishery: Seahorses in the central Philippines. Fisheries Research, 86(2–3):207–215, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607001282.

Vogler:2008:IEV

[VMQ08] Rodolfo Vögler, Andrés C. Milessi, and Renato A. Quiñones. Influence of environmental variables on the distribution of Squatina guggenheim (Chondrichthyes, Squatinidae) in the Argentine–Uruguayan common fishing zone. Fisheries Research, 91(2–3):212–221, June 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360700344X.

Vilizzi:2008:ELT

[VMSR08] Lorenzo Vilizzi, Shaun N. Meredith, Clayton P. Sharpe, and Rohan Rehwinkel. Evaluating light trap efficiency by application of mesh to prevent inter- and intra-specific in situ predation on fish larvae and juveniles. Fisheries Research, 93(1–2):146–153, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000921.

Vallin:2000:MEE

[VN00] Lars Vallin and Anders Nissling. Maternal effects on egg size and egg buoyancy of Baltic cod, Gadus morhua: Implications for stock structure effects on recruitment. Fisheries Research, 49(1):21–37, November 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001946.

Vabo:2002:EVA

[VOH02] Rune Vabø, Kjell Olsen, and Ingvar Huse. The effect of vessel avoidance of wintering Norwegian spring spawning herring. Fisheries Research, 58(1):59–77, October 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003605.

vanOostenbrugge:2002:SBU

[vOPvDM02] J. A. E. van Oostenbrugge, J. J. Poos, W. L. T. van Densen, and M. A. M. Machiels. In search of a better unit of effort in the coastal liftnet fishery with lights for small pelagics in Indonesia. Fisheries Research, 59(1–2):43–56, December 30, 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602000097.

Velasco:2001:RCF

[VOS01] F. Velasco, I. Olaso, and F. Sánchez. The role of cephalopods as forage for the demersal fish community in the southern Bay of Biscay. Fisheries Research, 52(1–2):65– 77, June 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601002314.

Veinott:2005:UOM

[VP05] Geoff Veinott and Rex Porter. Using otolith microchemistry to distinguish Atlantic salmon (Salmo salar) parr from different natal streams. Fisheries Research, 71(3):349–355, March 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604002310.

vonSzalay:2004:EIF

[vS04] Paul G. von Szalay. The effect of improved footrope contact measurements on trawl area swept estimates. Fisheries Research, 69(3):389–398, October 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001420.

Vestergaard:2003:MCC

[VSK03] Niels Vestergaard, Dale Squires, and Jim Kirkley. Measuring capacity and capacity utilization in fisheries: the case of the Danish gill-net fleet. Fisheries Research, 60(2–3): 357–368, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783602001418.

Vivekanandan:2005:FMF

[VSK05] E. Vivekanandan, M. Srinath, and Somy Kuriakose. Fishing the marine food web along the Indian coast. Fisheries Research, 72(2–3):241–252, May 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002498.

vonSzalay:2002:ESVa

[vSM02a] Paul G. von Szalay and Robert A. McConnaughey. The effect of slope and vessel speed on the performance of a single beam acoustic seabed classification system. Fisheries Research, 54(2):181–194, January 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360000309X. See erratum [vSM02c].

vonSzalay:2002:ESVb

[vSM02b] Paul G. von Szalay and Robert A. McConnaughey. The effect of slope and vessel speed on the performance of a single beam acoustic seabed classification system. Fisheries Research, 56(1):99–112, April 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000206.

vonSzalay:2002:EES

[vSM02c] Paul G. von Szalay and Robert A. McConnaughey. Erratum to “The effect of slope and vessel speed on the performance of a single beam acoustic seabed classification system”: [Fish. Res. 54 (2002) 181–194]. Fisheries Research, 56(1): 97, April 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360200019X.  See

[vSM02a].

vonSzalay:2005:ENS

[vSS05] Paul G. von Szalay and David A. Somerton. The effect of net spread on the capture efficiency of a demersal survey trawl used in the eastern Bering Sea. Fisheries Research, 74(1–3):86–95, August 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605001104.

Vaughan:2007:GMB

[VSS07] Douglas S. Vaughan, Kyle W. Shertzer, and Joseph W. Smith. Gulf menhaden (Brevoortia patronus) in the U.S. Gulf of Mexico: Fishery characteristics and biological reference points for management. Fisheries Research, 83(2–3):263–275, February 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003638.

vonSzalay:2009:CAB

[vSS09] Paul G. von Szalay and David A. Somerton. A comparison of acoustic backscatter between a trawling and a freerunning vessel for eastern Bering Sea walleye pollock (Theragra chalcogramma). Fisheries Research, 96(2–3):223–229, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608003822.

vonSzalay:2007:CTA

[vSSK07] Paul G. von Szalay, David A. Somerton, and Stan Kotwicki. Correlating trawl and acoustic data in the eastern Bering Sea: a first step toward improving biomass estimates of walleye pollock (Theragra chalcogramma) and Pacific cod (Gadus macrocephalus)? Fisheries Research, 86(1):77–83, August 2007. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783607001117.

vanTamelen:2004:COF

[vT04] Peter G. van Tamelen. A comparison of obtaining field data using electronic and written methods. Fisheries Research, 69(1):123–130, August 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000773.

Valentinsson:2008:SSN

[VU08] D. Valentinsson and M. Ulmestrand. Species-selective Nephrops trawling: Swedish grid experiments. Fisheries Research, 90(1–3):109–117, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002536.

Vilchez:2004:OCP

[VVH04] Ma Luisa Vílchez, Francisco Velasco, and Ines Herrero. An optimal control problem with Hopf bifurcations: an application to the striped venus fishery in the Gulf of Cádiz. Fisheries Research, 67(3):295–306, May 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603003047.

Vargas-Yanez:2009:RBO

[VYMGM^{+}09] M. Vargas-Yáñez, F. Moya, M. García-Martínez, J. Rey, M. González, and P. Zunino. Relationships between Octopus vulgaris landings and environmental factors in the northern Alboran Sea (Southwestern Mediterranean). Fisheries Research, 99(3):159–167, August 15, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360900143X.

vanZwieten:2002:EIA

[vZRMvD02] P. A. M. van Zwieten, F. C. Roest, M. A. M. Machiels, and W. L. T. van Densen. Effects of inter-annual variability, seasonality and persistence on the perception of long-term trends in catch rates of the industrial pelagic purse-seine fishery of northern Lake Tanganyika (Burundi). Fisheries Research, 54(3):329–348, February 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002673.

Wahle:2003:RSR

[Wah03] Richard A. Wahle. Revealing stock–recruitment relationships in lobsters and crabs: is experimental ecology the key? Fisheries Research, 65(1–3):3–32, December 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002327.

Ward:2006:IIS

[War06] Robert D. Ward. The importance of identifying spatial population structure in restocking and stock enhancement programmes. Fisheries Research, 80(1):9–18, August 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606001159.

Wilberg:2008:PDI

[WB08] Michael J. Wilberg and James R. Bence. Performance of deviance information criterion model selection in statistical catch-at-age analysis. Fisheries Research, 93(1–2):212–221, September 1, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001343.

Wells:2008:SSS

[WBCP08] R. J. David Wells, Kevin M. Boswell, James H. Cowan, and William F. Patterson. Size selectivity of sampling gears targeting red snapper in the northern Gulf of Mexico. Fisheries Research, 89(3):294–299, March 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002718.

Woll:2001:CRH

[WBHG01] Astrid K. Woll, Jesper Boje, Rene Holst, and Agnes C. Gundersen. Catch rates and hook and bait selectivity in longline fishery for Greenland halibut (Reinhardtius hippoglossoides, Walbaum) at East Greenland. Fisheries Research, 51(2–3):237–246, May 2001. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002491.

Weyl:2005:MRC

[WBMM05] Olaf L. F. Weyl, Anthony J. Booth, Kissa R. Mwakiyongo, and Davis. S. Mandere. Management recommendations for Copadichromis chrysonotus (Pisces: Cichlidae) in Lake Malombe, Malawi, based on per-recruit analysis. Fisheries Research, 71(2):165–173, February 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001894.

White:2007:WSL

[WC07] William T. White and Rachel D. Cavanagh. Whale shark landings in Indonesian artisanal shark and ray fisheries. Fisheries Research, 84(1):128–131, March 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004103.

Wu:2009:PSA

[WCC^{+}09] Georgiana Cho-Chen Wu, Hsin-Chieh Chiang, Kuo-Shu Chen, Chien-Chung Hsu, and Hsi-Yuan Yang. Population structure of albacore (Thunnus alalunga) in the Northwestern Pacific Ocean inferred from mitochondrial DNA. Fisheries Research, 95(1):125–131, January 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002221.

Wagner:2004:SAR

[WCM04] Tyler Wagner, James L. Congleton, and Douglas M. Marsh. Smolt-to-adult return rates of juvenile Chinook salmon transported through the Snake–Columbia River hydropower system, USA, in relation to densities of cotransported juvenile steelhead. Fisheries Research, 68(1– 3):259–270, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783604000177.

Wilson:2002:ELM

[WD02] Douglas C. Wilson and Poul Degnbol. The effects of legal mandates on fisheries science deliberations: the case of Atlantic bluefish in the United States. Fisheries Research, 58(1):1–14, October 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003666.

Wassenberg:2002:IFT

[WDC02] T. J. Wassenberg, G. Dews, and S. D. Cook. The impact of fish trawls on megabenthos (sponges) on the northwest shelf of Australia. Fisheries Research, 58(2):141–151, November 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601003824.

Ward:2009:ECH

[WEK^{+}09] Peter Ward, Sheree Epe, David Kreutz, Emma Lawrence, Carolyn Robins, and Annette Sands. The effects of circle hooks on bycatch and target catches in Australia’s pelagic longline fishery. Fisheries Research, 97(3):253–262, May 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783609000678.

West:2002:CFS

[Wes02] R. J. West. Comparison of fish and shrimp trawls for sampling deep-water estuarine fish in a large coastal river in eastern Australia. Fisheries Research, 54(3):409–417, February 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601002776.

Westgaard:2007:ACG

[WF07] Jon-Ivar Westgaard and Svein-Erik Fevolden. Atlantic cod (Gadus morhua L.) in inner and outer coastal zones of northern Norway display divergent genetic signature at non-neutral loci. Fisheries Research, 85(3):306–315, July 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607000872.

Winfield:2009:AFP

[WFJB09] Ian J. Winfield, Janice M. Fletcher, J. Ben James, and Colin W. Bean. Assessment of fish populations in still waters using hydroacoustics and survey gill netting: Experiences with Arctic charr (Salvelinus alpinus) in the UK. Fisheries Research, 96(1):30–38, February 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002786.

Walsh:2003:QAF

[WG03] Stephen J. Walsh and Olav Rune Godø. Quantitative analysis of fish reaction to towed fishing gears — what responses are important? Fisheries Research, 63(2):289–292, August 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603001309.

Watkinson:2005:SDL

[WG05] Douglas A. Watkinson and Darren M. Gillis. Stock discrimination of Lake Winnipeg walleye based on Fourier and wavelet description of scale outline signals. Fisheries Research, 72(2–3):193–203, May 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604002590.

White:2006:DBF

[WGDP06] William T. White, Jenny Giles, Dharmadi, and Ian C. Potter. Data on the bycatch fishery and reproductive biology of mobulid rays (Myliobatiformes) in Indonesia. Fisheries Research, 82(1–3):65–73, December 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606002992.

Wright:2006:FAC

[WGGN06] P. J. Wright, E. Galley, I. M. Gibb, and F. C. Neat. Fidelity of adult cod to spawning grounds in Scottish waters. Fisheries Research, 77(2):148–158, February 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003024.

Waluda:2008:RSS

[WGR08] Claire M. Waluda, Huw J. Griffiths, and Paul G. Rodhouse. Remotely sensed spatial dynamics of the Illex argentinus fishery, Southwest Atlantic. Fisheries Research, 91(2–3):196–202, June 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003426.

Willis:2008:ABI

[WH08] Jay Willis and Alistair J. Hobday. Application of bioelectrical impedance analysis as a method for estimating composition and metabolic condition of southern bluefin tuna (Thunnus maccoyii) during conventional tagging. Fisheries Research, 93(1–2):64–71, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000659.

Wetherbee:2004:UMR

[WHML04] Bradley M. Wetherbee, Kim N. Holland, Carl G. Meyer, and Christopher G. Lowe. Use of a marine reserve in Kaneohe Bay, Hawaii by the giant trevally, Caranx ignobilis. Fisheries Research, 67(3):253–263, May 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603003461.

Walter:2007:EEM

[WHWM07] John Walter III, John M. Hoenig, Robert J. Wood, and Karen Marti. An estimator of episodic mortality in bivalves with an application to sea scallops (Placopecten magellanicus). Fisheries Research, 86(2–3):85–91, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001075.

Wieland:2004:LST

[Wie04] Kai Wieland. Length at sex transition in northern shrimp (Pandalus borealis) off West Greenland in relation to changes in temperature and stock size. Fisheries Research, 69(1):49–56, August 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000827.

Wilberg:2008:ESS

[WIJB08] Michael J. Wilberg, Brian J. Irwin, Michael L. Jones, and James R. Bence. Effects of source–sink dynamics on harvest policy performance for yellow perch in southern Lake Michigan. Fisheries Research, 94(3):282–289, December 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608001410.

Walsh:2005:ALA

[WIKM05] William A. Walsh, Russell Y. Ito, Kurt E. Kawamoto, and Marti McCracken. Analysis of logbook accuracy for blue marlin (Makaira nigricans) in the Hawaii-based longline fishery with a generalized additive model and commercial sales data. Fisheries Research, 75(1–3):175–192, September 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604003017.

Wilberg:2009:ERB

[Wil09] Michael J. Wilberg. Estimation of recreational bag limit noncompliance using contact creel survey data. Fisheries Research, 99(3):239–243, August 15, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360900174X.

Winfield:2000:BRM

[Win00] Ian J. Winfield. Book review: Management of Freshwater Fisheries: J. Arrignon, Balkema, Rotterdam, 1999, 582 pages, 85 EUR, ISBN 90-5410-721-9. Fisheries Research, 47(1):99–100, June 2000. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001223.

Winfield:2002:BRM

[Win02] Ian Winfield. Book review: Management and Ecology of River Fisheries: I. G. Cox (Ed.); Fishing News Books, Blackwell Scientific, Oxford, 2000, xi + 444 pages, hardback, ISBN 0-85238-250-2, £79.50. Fisheries Research, 58(3):409–410, November 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602001662.

Waknitz:2003:IAS

[WIS03] F. William Waknitz, Robert N. Iwamoto, and Mark S. Strom. Interactions of Atlantic salmon in the Pacific Northwest: IV. Impacts on the local ecosystems. Fisheries Research, 62(3):307–328, June 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000663.

Walsh:2001:GAM

[WK01] William A. Walsh and Pierre Kleiber. Generalized additive model and regression tree analyses of blue shark (Prionace glauca) catch rates by the Hawaii-based commercial longline fishery. Fisheries Research, 53(2):115–131, October 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783600003064.

Weinberg:2008:FIN

[WK08] Kenneth L. Weinberg and Stan Kotwicki. Factors influencing net width and sea floor contact of a survey bottom trawl. Fisheries Research, 93(3):265–279, September 28, 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608001422.

Watters:2006:RAC

[WKCC06] Diana L. Watters, Donna E. Kline, Kenneth H. Coale, and Gregor M. Cailliet. Radiometric age confirmation and growth of a deep-water marine fish species: the bank rockfish, Sebastes rufus. Fisheries Research, 81(2–3):251–257, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606002475.

Walsh:2002:CLR

[WKM02] William A. Walsh, Pierre Kleiber, and Marti McCracken. Comparison of logbook reports of incidental blue shark catch rates by Hawaii-based longline vessels to fishery observer data by application of a generalized additive model. Fisheries Research, 58(1):79–94, October 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003617.

Wang:2006:CAI

[WL06a] Yanjun Wang and Qun Liu. Comparison of Akaike information criterion (AIC) and Bayesian information criterion (BIC) in selection of stock–recruitment relationships. Fisheries Research, 77(2):220–225, February 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605002870.

Wang:2006:ENM

[WL06b] Yingbin Wang and Qun Liu. Estimation of natural mortality using statistical analysis of fisheries catch-at-age data. Fisheries Research, 78(2–3):342–351, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000129.

Ward:2008:LSE

[WLDH08] Peter Ward, Emma Lawrence, Rebecca Darbyshire, and Sheree Hindmarsh. Large-scale experiment shows that nylon leaders reduce shark bycatch and benefit pelagic longline fishers. Fisheries Research, 90(1–3):100–108, April 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607002512.

Wang:2008:PMD

[WLL08] Kae-Yih Wang, Cheng-Hsin Liao, and Kuo-Tien Lee. Population and maturation dynamics of the swordtip squid (Photololigo edulis) in the southern East China Sea. Fisheries Research, 90(1–3):178–186, April 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360700272X.

Walmsley:2007:BDS

[WLS07] Sarah A. Walmsley, Rob W. Leslie, and Warwick H. H. Sauer. Bycatch and discarding in the South African demersal trawl fishery. Fisheries Research, 86(1):15–30, August 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783607000720.

Wang:2006:BAA

[WLY06] Yanjun Wang, Qun Liu, and Zhenjiang Ye. A Bayesian analysis on the anchovy stock (Engraulis japonicus) in the Yellow Sea. Fisheries Research, 82(1–3):87–94, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606002979.

Weidman:2000:HRS

[WM00] Christopher R. Weidman and Richard Millner. Highresolution stable isotope records from North Atlantic cod. Fisheries Research, 46(1–3):327–342, May 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001570.

Watson:2004:SRC

[WM04] Maggie Watson and John L. Munro. Settlement and recruitment of coral reef fishes in moderately exploited and overexploited Caribbean ecosystems: implications for marine protected areas. Fisheries Research, 69(3):415–425, October 2004. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604001493.

Ward:2007:BLP

[WM07] Peter Ward and Ransom A. Myers. Bait loss and its potential effects on fishing power in pelagic longline fisheries. Fisheries Research, 86(1):69–76, August 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001087.

Waddington:2009:CBL

[WM09] Kris I. Waddington and Jessica J. Meeuwig. Contribution of bait to lobster production in an oligotrophic marine ecosystem as determined using a mass balance model. Fisheries Research, 99(1):1–6, July 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000927.

Welch:2008:STV

[WMB08] David J. Welch, Bruce D. Mapstone, and Gavin A. Begg. Spatial and temporal variation and effects of changes in management in discard rates from the commercial reef line fishery of the Great Barrier Reef, Australia.

Fisheries Research, 90(1–3):247–260, April 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002974.

Williams:2007:SPC

[WMD07] A. J. Williams, B. D. Mapstone, and C. R. Davies. Spatial patterns in cohort-specific mortality of red throat emperor, Lethrinus miniatus, on the Great Barrier Reef. Fisheries Research, 84(3):328–337, May 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004127.

Wang:2009:IMD

[WMdS09] Sheng-Ping Wang, Mark N. Maunder, and Alexandre Aires da Silva. Implications of model and data assumptions: an illustration including data for the Taiwanese longline fishery into the eastern Pacific Ocean bigeye tuna (Thunnus obesus) stock assessment. Fisheries Research, 97(1–2):118– 126, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609000344.

Wang:2009:EFI

[WMdSB09] Sheng-Ping Wang, Mark N. Maunder, Alexandre Aires da Silva, and William H. Bayliff. Evaluating fishery impacts: Application to bigeye tuna (Thunnus obesus) in the eastern Pacific Ocean. Fisheries Research, 99(2):106–111, August 1, 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783609001155.

Woods:2004:VIF

[WMS04] Chris M. C. Woods and Keith M. Martin-Smith. Visible implant fluorescent elastomer tagging of the bigbellied seahorse, Hippocampus abdominalis. Fisheries Research, 66(2–3):363–371, February 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603001838.

Wakefield:2007:CSJ

[WMTJ07] Corey B. Wakefield, Michael J. Moran, Nadia E. Tapp, and Gary Jackson. Catchability and selectivity of juvenile snapper (Pagrus auratus, Sparidae) and western butterfish (Pentapodus vitta, Nemipteridae) from prawn trawling in a large marine embayment in Western Australia. Fisheries Research, 85(1–2):37–48, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004334.

Wildman:2003:CSE

[WN03] Timothy L. Wildman and Robert M. Neumann. Comparison of snorkeling and electrofishing for estimating abundance and size structure of brook trout and brown trout in two southern New England streams. Fisheries Research, 60(1):131–139, January 30, 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602000607.

Wroblewski:2000:RIS

[WNRd00] J. S. Wroblewski, Bruce G. Nolan, George A. Rose, and Brad deYoung. Response of individual shoaling Atlantic cod to ocean currents on the northeast Newfoundland Shelf. Fisheries Research, 45(1):51–59, February 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578369900096X.

Wieland:2009:EBT

[WPOB09] Kai Wieland, Eva Maria Fenger Pedersen, Hans J. Olesen, and Jan E. Beyer. Effect of bottom type on catch rates of North Sea cod (Gadus morhua) in surveys with commercial fishing vessels. Fisheries Research, 96(2–3):244–251, March 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360800386X.

Wang:2007:RSL

[WPS^{+}07] J. Wang, G. J. Pierce, M. Sacau, J. Portela, M. B. Santos, X. Cardoso, and J. M. Bellido. Remotely sensed local oceanic thermal features and their influence on the distribution of hake (Merluccius hubbsi) at the Patagonian shelf edge in the SW Atlantic. Fisheries Research, 83(2–3):133–144, February 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003146.

Wienecke:2002:SSF

[WR02] Barbara Wienecke and Graham Robertson. Seabird and seal — fisheries interactions in the Australian Patagonian toothfish Dissostichus eleginoides trawl fishery. Fisheries Research, 54(2):253–265, January 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600003076.

Windle:2005:MRF

[WR05] Matthew J. S. Windle and George A. Rose. Migration route familiarity and homing of transplanted Atlantic cod (Gadus morhua). Fisheries Research, 75(1–3):193–199, September 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S016578360500144X.

Wade:2009:RDF

[WRGS09] O. Wade, A. S. Revill, A. Grant, and M. Sharp. Reducing the discards of finfish and benthic invertebrates of UK beam trawlers. Fisheries Research, 97(1–2):140–147, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000356.

Watson:2006:FGAa

[WRK06a] R. Watson, C. Revenga, and Y. Kura. Fishing gear associated with global marine catches: I. Database development. Fisheries Research, 79(1–2):97–102, June 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000427.

Watson:2006:FGAb

[WRK06b] R. Watson, C. Revenga, and Y. Kura. Fishing gear associated with global marine catches: II. trends in trawling and dredging. Fisheries Research, 79(1–2):103–111, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000439.

Weinberg:2002:ETD

[WRWK02] James R. Weinberg, Paul J. Rago, W. Waldo Wakefield, and Charles Keith. Estimation of tow distance and spatial heterogeneity using data from inclinometer sensors: an example using a clam survey dredge. Fisheries Research, 55(1–3):49–61, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002922.

White:2008:BPR

[WSC08] Alaina J. White, Jason F. Schreer, and Steven J. Cooke. Behavioral and physiological responses of the congeneric largemouth (Micropterus salmoides) and smallmouth bass (M. dolomieu) to various exercise and air exposure durations. Fisheries Research, 89(1):9–16, January 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002056.

Weinberg:2002:ETS

[WSM02] Kenneth L. Weinberg, David A. Somerton, and Peter T. Munro. The effect of trawl speed on the footrope capture efficiency of a survey trawl. Fisheries Research, 58(3):303–313, November 30, 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003952.

Wieland:2006:ETD

[WSP06] Kai Wieland and Marie Storr-Paulsen. Effect of tow duration on catch rate and size composition of northern shrimp (Pandalus borealis) and Greenland halibut (Reinhardtius hippoglossoides) in the West Greenland Bottom Trawl Survey. Fisheries Research, 78(2–3):276–285, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000191.

Wang:2005:ESS

[WSPY05] Sheng-Ping Wang, Chi-Lu Sun, André E. Punt, and Su-Zan Yeh. Evaluation of a sex-specific age-structured assessment method for the swordfish, Xiphias gladius, in the North Pacific Ocean. Fisheries Research, 73(1–2):79–97, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605000032.

Wang:2007:ASS

[WSPY07] Sheng-Ping Wang, Chi-Lu Sun, André E. Punt, and Su-Zan Yeh. Application of the sex-specific age-structured assessment method for swordfish, Xiphias gladius, in the North Pacific Ocean. Fisheries Research, 84(3):282–300, May 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606003882.

Ward:2002:CSP

[WSS02] Tim Ward and Jonathan Staunton-Smith. Comparison of the spawning patterns and fisheries biology of the sardine, Sardinops sagax, in temperate South Australia and sub-tropical southern Queensland. Fisheries Research, 56(1):37–49, April 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003149.

White:2006:AMS

[WSS06] D. J. White, C. Svellingen, and N. J. C. Strachan. Automated measurement of species and length of fish by computer vision. Fisheries Research, 80(2–3):203–210, September 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783606001512.

Wang:2000:TMG

[WT00] C. H. Wang and W. N. Tzeng. The timing of metamorphosis and growth rates of American and European eel leptocephali: a mechanism of larval segregative migration. Fisheries Research, 46(1–3):191–205, May 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001466.

Wilson:2008:ESI

[WVW^{+}08] David T. Wilson, Duncan Vaughan, Shaun K. Wilson, Carrie N. Simon, and Kathy Lockhart. Efficacy of a starch-iodide swab technique to detect the illegal use of bleach in a spiny lobster (Panulirus argus) fishery. Fisheries Research, 90(1–3):86–91, April 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002494.

Wallace:2006:MDF

[WW06] John R. Wallace and Charles W. West. Measurements of distance fished during the trawl retrieval period. Fisheries Research, 77(3):285–292, March 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000099.

Woodroffe:2003:VAI

[WWG03] D. A. Woodroffe, P. J. Wright, and J. D. M. Gordon. Verification of annual increment formation in the white anglerfish, Lophius piscatorius using the illicia and sagitta otoliths. Fisheries Research, 60(2–3):345–356, February 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602001741.

Watanabe:2006:LTC

[WY06] Chikako Watanabe and Akihiko Yatsu. Long-term changes in maturity at age of chub mackerel (Scomber japonicus) in relation to population declines in the waters off northeastern Japan. Fisheries Research, 78(2–3):323–332, May 2006.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000154.

Waluda:2006:IEC

[WYR06] Claire M. Waluda, Carmen Yamashiro, and Paul G. Rodhouse. Influence of the ENSO cycle on the light-fishery for Dosidicus gigas in the Peru Current: an analysis of remotely sensed data. Fisheries Research, 79(1–2):56–63, June 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606000671.

Wang:2006:SET

[WZDY06] Qingyin Wang, Zhimeng Zhuang, Jingyao Deng, and Yimin Ye. Stock enhancement and translocation of the shrimp Penaeus chinensis in China. Fisheries Research, 80(1):67–79, August 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606001226.

Xinjun:2007:FBP

[XBS^{+}07] Chen Xinjun, Liu Bilin, Tian Siquan, Qian Weiguo, and Zhao Xiaohu. Fishery biology of purpleback squid, Sthenoteuthis oualaniensis, in the northwest Indian Ocean. Fisheries Research, 83(1):98–104, January 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003171.

Xiao:2004:UGL

[Xia04a] Yongshun Xiao. Use of generalized linear models in analyzing the catch and effort data on the western king prawn Penaeus latisulcatus Kishinouye in the Gulf St. Vincent, Australia. Fisheries Research, 68(1–3):67–82, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000748.

Xiao:2004:UIT

[Xia04b] Yongshun Xiao. Use of individual types of fishing effort in analyzing catch and effort data by use of a generalized linear model. Fisheries Research, 70(2–3):311–318, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001742.

Xiao:2006:STS

[Xia06] Yongshun Xiao. Several traps in size-dependent population models published in fisheries journals. Fisheries Research, 77(3):265–270, March 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003073.

Xu:2005:VFC

[XJ05] Binduo Xu and Xianshi Jin. Variations in fish community structure during winter in the southern Yellow Sea over the period 1985–2002. Fisheries Research, 71(1):79–91, January 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783604001961.

Xiao:2004:SRP

[XK04] Yongshun Xiao and Martin Kumar. Sex ratio, and probability of sexual maturity of females at size, of the blue swimmer crab, Portunus pelagicus Linnaeus, off southern Australia. Fisheries Research, 68(1–3):271–282, July 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S016578360400027X.

Xiao:2000:UAT

[XM00] Y. Xiao and P. McShane. Use of age- and time-dependent seasonal growth models in analysis of tag/recapture data on the western king prawn Penaeus latisulcatus in the Gulf St. Vincent, Australia. Fisheries Research, 49(1):85–92, November 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600001107.

Xiao:2004:MFR

[XPMQ04] Yongshun Xiao, André E. Punt, Russell B. Millar, and Terrence J. Quinn. Models in fisheries research: GLMs, GAMS and GLMMs. Fisheries Research, 70(2–3):137–139, December 2004. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783604001626.

Xiao:2007:RPC

[XW07] Yongshun Xiao and You-Gan Wang. A revisit to Pope’s cohort analysis. Fisheries Research, 86(2–3):153– 158, September 2007. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783607001208.

Ye:2001:UGL

[YAHAB01] Yimin Ye, M. Al-Husaini, and A. Al-Baz. Use of generalized linear models to analyze catch rates having zero values: the Kuwait driftnet fishery. Fisheries Research, 53(2):151–168, October 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783600002873.

Ye:2000:BDK

[YAM00] Yimin Ye, A. H. Alsaffar, and H. M. A. Mohammed. Bycatch and discards of the Kuwait shrimp fishery. Fisheries Research, 45(1):9–19, February 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783699001058.

Yau:2001:EAP

[YCB^{+}01] C. Yau, M. A. Collins, P. M. Bagley, I. Everson, C. P. Nolan, and I. G. Priede. Estimating the abundance of Patagonian toothfish Dissostichus eleginoides using baited cameras: a preliminary study. Fisheries Research, 51(2–3):403– 412, May 2001. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601002648.

Yanase:2007:IWT

[YEA07] Kazutaka Yanase, Steve Eayrs, and Takafumi Arimoto. Influence of water temperature and fish length on the maximum swimming speed of sand flathead, Platycephalus bassensis: Implications for trawl selectivity. Fisheries Research, 84(2):180–188, April 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606003791.

Yanase:2009:QAB

[YEA09] Kazutaka Yanase, Steve Eayrs, and Takafumi Arimoto. Quantitative analysis of the behaviour of the flatheads (Platycephalidae) during the trawl capture process as determined by real-time multiple observations. Fisheries Research, 95(1):28–39, January 1, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002154.

Yemane:2008:ICS

[YFL08] Dawit Yemane, John G. Field, and Rob W. Leslie. Indicators of change in the size structure of fish communities: a case study from the south coast of South Africa. Fisheries Research, 93(1–2):163–172, September 1, 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608000957.

Yoneda:2002:RCS

[YFT^{+}02] M. Yoneda, K. Futagawa, M. Tokimura, H. Horikawa, S. Matsuura, and M. Matsuyama. Reproductive cycle, spawning frequency and batch fecundity of the female whitefin jack Kaiwarinus equula in the East China Sea. Fisheries Research, 57(3):297–309, September 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601003496.

Yilmaz:2002:RBT

[Yil02] Fevzi Yilmaz. Reproductive biology of the tench Tinca tinca (L., 1758) inhabiting Porsuk Dam Lake (Kutahya, Turkey). Fisheries Research, 55(1–3):313–317, March 2002. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002946.

Youngson:2003:MSF

[YJV^{+}03] A. F. Youngson, W. C. Jordan, E. Verspoor, P. McGinnity, T. Cross, and A. Ferguson. Management of salmonid fisheries in the British Isles: towards a practical approach based on population genetics. Fisheries Research, 62(2):193–209, May 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783602001625.

Yokota:2006:SCP

[YKM06] Kosuke Yokota, Masashi Kiyota, and Hiroshi Minami. Shark catch in a pelagic longline fishery: Comparison of circle and tuna hooks. Fisheries Research, 81(2–3):337–341, November 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606002931.

Yokota:2009:EBS

[YKO09] Kosuke Yokota, Masashi Kiyota, and Hiroshi Okamura. Effect of bait species and color on sea turtle bycatch and fish catch in a pelagic longline fishery. Fisheries Research, 97(1–2):53–58, April 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783609000022.

Yunkai-Li:2009:ESF

[YLCS^{+}09] Yunkai-Li, Yong Chen, Bing Song, Derek Olson, Na Yu, and Liqiao-Chen. Ecosystem structure and functioning of Lake Taihu (China) and the impacts of fishing. Fisheries Research, 95(2–3):309–324, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608003226.

Ye:2005:BMR

[YLD^{+}05] Yimin Ye, Neil Loneragan, David Die, Reg Watson, and Bronwyn Harch. Bioeconomic modelling and risk assessment of tiger prawn (Penaeus esculentus) stock enhancement in Exmouth Gulf, Australia. Fisheries Research, 73 (1–2):231–249, June 2005. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S016578360400308X.

YingbinWang:2009:EIR

[YLYL09] YingbinWang, Qun Liu, Cungen Yu, and Liping Lian. Estimation of instantaneous rate of natural mortality from catch-at-age data and an abundance index based on extended survivors method. Fisheries Research, 97(1–2):127– 133, April 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S016578360900037X.

Yatsu:2000:EGA

[YM00] Akihiko Yatsu and Junta Mori. Early growth of the autumn cohort of neon flying squid, Ommastrephes bartramii, in the North Pacific Ocean. Fisheries Research, 45(2):189–194, March 2000. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783699001125.

Yamamoto:2006:GET

[YMT^{+}06] Shoichiro Yamamoto, Koji Maekawa, Tsuyoshi Tamate, Itsuro Koizumi, Koh Hasegawa, and Hitoshi Kubota. Genetic evaluation of translocation in artificially isolated populations of white-spotted charr (Salvelinus leucomaenis). Fisheries Research, 78(2–3):352–358, May 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606000117.

Yousif:2003:DVS

[You03] Aly Yousif. Diel variability of size and catch rate of three fish species and three penaeid prawns in the NW Red Sea trawl fishery. Fisheries Research, 63(2):265–274, August 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783603000717.

Young:2004:ADM

[YPD^{+}04] I. A. G. Young, G. J. Pierce, H. I. Daly, M. B. Santos, L. N. Key, N. Bailey, J.-P. Robin, A. J. Bishop, G. Stowasser, M. Nyegaard, S. K. Cho, M. Rasero, and J. M. F. Pereira. Application of depletion methods to estimate stock size in the squid Loligo forbesi in Scottish waters (UK). Fisheries Research, 69(2):211–227, September 2004.

CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604001390.

Ye:2005:CAI

[YPDS05] Yimin Ye, Roland Pitcher, Darren Dennis, and Tim Skewes. Constructing abundance indices from scientific surveys of different designs for the Torres Strait ornate rock lobster (Panulirus ornatus) fishery, Australia. Fisheries Research, 73(1–2):187–200, June 2005. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604003145.

Young:2006:AGM

[YPM^{+}06] I. A. G. Young, G. J. Pierce, J. Murphy, H. I. Daly, and N. Bailey. Application of the Gómez–Muñoz model to estimate catch and effort in squid fisheries in Scotland. Fisheries Research, 78(1):26–38, April 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003310.

Young:2006:MFD

[YPS^{+}06] I. A. G. Young, G. J. Pierce, G. Stowasser, M. B. Santos, J. Wang, P. R. Boyle, P. W. Shaw, N. Bailey, I. Tuck, and M. A. Collins. The Moray Firth directed squid fishery. Fisheries Research, 78(1):39–43, April 2006. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783605003346.

Yule:2009:CPF

[YSS^{+}09] D. L. Yule, J. D. Stockwell, D. R. Schreiner, L. M. Evrard, M. Balge, and T. R. Hrabik. Can pelagic forage fish and spawning cisco (Coregonus artedi) biomass in the western arm of Lake Superior be assessed with a single summer survey? Fisheries Research, 96(1):39–50, February 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002798.

Yoneda:2002:AGL

[YST^{+}02] M. Yoneda, T. Sakai, M. Tokimura, H. Horikawa, and M. Matsuyama. Age and growth of the lizardfish Saurida sp. 1 in the East China Sea using otolith ring marks. Fisheries Research, 55(1–3):231–238, March 2002. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783601002855.

Yamaguchi:2006:RCS

[YTK06] Atsuko Yamaguchi, Taikai Todoroki, and Gen Kume. Reproductive cycle, sexual maturity and diel-reproductive periodicity of white croaker, Pennahia argentata (Sciaenidae), in Ariake Sound, Japan. Fisheries Research, 82(1–3): 95–100, December 2006. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0165783606002967.

Zhang:2000:UOM

[ZB00] Z. Zhang and R. J. Beamish. Use of otolith microstructure to study life history of juvenile Chinook salmon in the Strait of Georgia in 1995 and 1996. Fisheries Research, 46(1–3):239–250, May 2000. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600001491.

Zhang:2002:ASS

[ZC02] Z. Zhang and A. Campbell. Application of a stochastic spawning stock biomass per recruit model for the horse clam fishery in British Columbia. Fisheries Research, 57(1): 9–23, July 2002. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783601003332.

Zeeberg:2006:BRP

[ZCdG06] JaapJan Zeeberg, Ad Corten, and Erik de Graaf. Bycatch and release of pelagic megafauna in industrial trawler fisheries off Northwest Africa. Fisheries Research, 78(2–3):186– 195, May 2006. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783606000403.

Zeeberg:2008:CME

[ZCTA^{+}08] JaapJan Zeeberg, Ad Corten, Pablo Tjoe-Awie, Josep Coca, and Bambay Hamady. Climate modulates the effects of Sardinella aurita fisheries off Northwest Africa. Fisheries Research, 89(1):65–75, January 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607002160.

Zhou:2007:CDD

[ZDB^{+}07] Shijie Zhou, Cathy Dichmont, Charis Y. Burridge, William N. Venables, Peter J. Toscas, and David Vance. Is catchability density-dependent for schooling prawns? Fisheries Research, 85(1–2):23–36, June 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783606004322.

Zeller:2008:WAR

[ZDB^{+}08] D. Zeller, M. Darcy, S. Booth, M. K. Lowe, and S. Martell. What about recreational catch?: Potential impact on stock assessment for Hawaii’s bottomfish fisheries. Fisheries Research, 91(1):88–97, May 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003177.

Ziegler:2003:STV

[ZFJ03] P. E. Ziegler, S. D. Frusher, and C. R. Johnson. Space– time variation in catchability of southern rock lobster Jasus edwardsii in Tasmania explained by environmental, physiological and density-dependent processes. Fisheries Research, 61(1–3):107–123, March 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783602002400.

Zhou:2008:SAF

[ZG08] Shijie Zhou and Shane P. Griffiths. Sustainability Assessment for Fishing Effects (SAFE): a new quantitative ecological risk assessment method and its application to elasmobranch bycatch in an Australian trawl fishery. Fisheries Research, 91(1):56–68, May 2008. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003141.

Zheng:2003:UNM

[Zhe03] Jie Zheng. Uncertainties of natural mortality estimates for eastern Bering Sea snow crab, Chionoecetes opilio. Fisheries Research, 65(1–3):411–425, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002583.

Zhou:2007:DAS

[Zho07] Shijie Zhou. Discriminating alternative stock–recruitment models and evaluating uncertainty in model structure. Fisheries Research, 86(2–3):268–279, September 2007. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607001592.

Zheng:2003:SRR

[ZK03] Jie Zheng and Gordon H. Kruse. Stock–recruitment relationships for three major Alaskan crab stocks. Fisheries Research, 65(1–3):103–121, December 2003. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603002388.

Zhang:2009:EBF

[ZKG^{+}09] Chang Ik Zhang, Suam Kim, Donald Gunderson, Richard Marasco, Jae Bong Lee, Hee Won Park, and Jong Hee Lee. An ecosystem-based fisheries assessment approach for Korean fisheries. Fisheries Research, 100(1):26–41, September 2009. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-

6763 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0165783608003871.

Zhang:2009:UBH

[ZLC09] Zane Zhang, Joanne Lessard, and Alan Campbell. Use of Bayesian hierarchical models to estimate northern abalone, Haliotis kamtschatkana, growth parameters from tagrecapture data. Fisheries Research, 95(2–3):289–295, January 14, 2009. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783608003184.

Zhang:2005:USS

[ZP05] Zane Zhang and R. Ian Perry. Use of state–space modeling with a Bayesian approach to estimate target reference points for green sea urchin (Strongylocentrotus droebachiensis) stocks in the Queen Charlotte Strait region, British Columbia, Canada. Fisheries Research, 74(1–3):253–264, August 2005. CODEN FISRDJ. ISSN 0165-7836 (print),

1872-6763 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0165783605000391.

Zheng:2001:SPW

[ZPR01] X. Zheng, G. J. Pierce, and D. G. Reid. Spatial patterns of whiting abundance in Scottish waters and relationships with environmental variables. Fisheries Research, 50(3):259–270, March 2001. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783600002198.

Zeller:2004:FSC

[ZR04] Dirk Zeller and Garry R. Russ. Are fisheries ‘sustainable’? A counterpoint to Steele and Hoagland. Fisheries Research, 67(2):241–245, April 2004. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783604000153.

Zapfe:2008:CGD

[ZR08] Glenn A. Zapfe and Chet F. Rakocinski. Coherent growth and diet patterns in juvenile spot (Leiostomus xanthurus Lacépède) reflect effects of hydrology on access to shoreline habitat. Fisheries Research, 91(1):107–111, May 2008. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783607003591.

Zlokovitz:2003:PMH

[ZSP03] Erik R. Zlokovitz, David H. Secor, and Philip M. Piccoli. Patterns of migration in Hudson River striped bass as determined by otolith microchemistry. Fisheries Research, 63(2):245–259, August 2003. CODEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783603000699.

Zeeh:2009:IVI

[ZW09] Kimberly M. Zeeh and James B. Wood. Impact of visible implant elastomer tags on the growth rate of captive Caribbean reef squid Sepioteuthis sepioidea. Fisheries Research, 95(2–3):362–364, January 14, 2009. CO-DEN FISRDJ. ISSN 0165-7836 (print), 1872-6763 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0165783608002294.